A Complete Bibliography of Publications in IEEE Transactions on Visualization and Computer Graphics (2020–2029)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: https://www.math.utah.edu/~beebe/

01 May 2024 Version 1.30

Title word cross-reference

(SGD)^{2} [ADD^{+}22]. 2 [LBB^{+}20, RMB^{+}21, XYF^{+}21, ZLG^{+}21a]. 3 [CLL^{+}20, CZY^{+}20, FHR^{+}21, GZM^{+}21, HZZ^{+}20, HB20, HCL20, JDM^{+}22, JSA^{+}20, KCA^{+}21, KNM^{+}21, KRK21, KKF20, KWFK20, LWC22, LZX^{+}21, LCK^{+}21, MDJV21, MEB^{+}20, NDF^{+}21, PALW20, PL21, SYW^{+}20, SZF^{+}21, SSX^{+}20, SLR20b, STA^{+}21, SKNŽ20, SS21, TGM21, TIHS20, WZH20, WL20, WCX21b, XJZ^{+}21, XF21, XYF^{+}21, YRL^{+}20, YZJ^{+}20, YCB^{+}21, YGE^{+}21, ZHDX20, ZHL^{+}21a, ZHL^{+}20, ZH20, iIK^{+}21]. 4 [LWWY21, WZH20]. ^{∘} [ACL^{+}24, BBMM^{+}23, CZF^{+}23, GFCM23, MLC^{+}20, MP21, MSB^{+}22, CSM^{+}23]. d [ZSL21]. ℓ_{0} [TYPC20]. K [KBV24, JSF20, ZSL21]. Z [CNAA^{+}22].

-Nearest [JSF20]. -Set [KBV24].

//vgtc.org/

[Ano20j, Ano21f, Ano22o, Ano20k, Ano21g].

0-dimensional [DTS^{+}21].

1 [SESH24]. 10.1109 [FCFC22a]. 10.1109/TVCG.2020.3021534 [FCFC22a]. 100 [ZYL^{+}24]. 100-Phones [ZYL^{+}24]. 133s [HWMI23]. 19 [JDZK22, NSK^{+}21, PFCB23, ZSG^{+}23]. 1D [CQHP22].

2 [ZBNS21]. 2.0 [ESP20]. 2.5D [FPK^{+}24, MRS22]. 20 [ZWV21]. 2019 [SAB20]. 2020 [BSW20, Fek21, MB20a, MB20b, SW21]. 2021 [Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, CTW21, MB21a]. 2022 [Ano22l, Ano22n, Ano22q, Ano23a, Ano23x, Ano23b, Ano23c, Ano23d, Ano23e, CRZ22, MF22, MB22a, NAW^{+}22, OK22, SES23]. 2023 [Ano23f, Ano23j, Ano23g, Ano23n, Ano23o, Ano23p, Ano23q, Ano24w, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, BPQW23, CRH23, SK23a]. 2024 [Ano24m, IPPZ24]. 2D [AMY^{+}22, ACT^{+}24, FPK^{+}24, KLB24, KPO^{+}23, ZHTR22].

360-Degree [HRX23, NBE^{+}21, VH23, LDB^{+}21, LCW^{+}23b]. 3D [NIK24, AJVS^{+}23, AeSL^{+}23, Ano23f, BPA22, COFJ23, CSWZ24, CDS^{+}22, CLL^{+}23, CWL23, DMJ^{+}22, DHLA22, DHF^{+}22, FHTB23, FZZX22, FPK^{+}24, GWL^{+}22, GBM^{+}22, GLL^{+}22, HYH^{+}23, HNS23, HZQ22, HSC^{+}22, HCH^{+}23, HZYZ22, HZZ24, KSHW22, LFO23, LLL^{+}22, LWL^{+}23a, LMGY22, LLZ^{+}23, LSD^{+}23, LSL^{+}22, MDL^{+}23, MSK23a, MFS^{+}23, NSW24, NL24, QGY^{+}22, SA22, SDK23a, SWZ^{+}23, VH23, WCH^{+}23, WTY^{+}22, XLY^{+}22, YKF22, YXS^{+}23, YCHL24, YFM^{+}23, ZGX^{+}22, ZLY22, ZZX^{+}22, ZYL^{+}22, ZWL23, ZYC^{+}23, ZNF^{+}23, ZCSS23]. 3D-CariGAN [YXS^{+}23]. 3D-Enabled [NIK24].

4D [GXH^{+}24].

6 [CZF^{+}23]. 6-DoF [CZF^{+}23]. 6D [CLZ^{+}24]. 6DOF [MSK23b, JVRL24].

Ability [BBSC23, HKB^{+}22, RBLT^{+}22]. Ablate [CPCS20]. Abstract [PSY^{+}20, WJKN21, YCB^{+}21, ZLG^{+}21a]. Abstraction [DXX^{+}21, JSS^{+}20, KRK21]. Abstractions [BWI21, ZGX^{+}23]. AC [TSS22]. Academic [GTC^{+}23, NKWW22, WPL^{+}22]. Academy [Ano22x]. Accelerated [RGG20a]. Accelerating [MWUP22]. Acceleration [HPT^{+}23]. Access [VH23]. Accessibility [CRJ^{+}24, ENM24, SB23]. Accessibility-Centered [ENM24]. Accessible [CPR^{+}22, LS22]. According [NE24]. Accuracy [KCA^{+}21, YTHL23]. Accurate [CXW^{+}23, DHM^{+}22, FiMH21, SA22, WCX21b]. Achievement [Ano22b, Ano22e, Ano23c, Ano23e, Ano24c, Ano24f, Fek21, Kiy22b, Kli22]. Achieving [HJL^{+}23]. Acoustic [GVN^{+}20, TBL^{+}20]. Acquisition [Tet24, WPNK21, ZYR^{+}20]. Across [BSG^{+}20, HKB^{+}22, HMK^{+}20, KSHW22, KPO^{+}23]. ActFloor [WZC^{+}23]. ActFloor-GAN [WZC^{+}23]. Action [BvdPLH22, BKR^{+}24, CXZ^{+}24, KVM^{+}22, LZL^{+}23, CXZ^{+}24]. Action-Evaluator [CXZ^{+}24]. Action-Specific [KVM^{+}22]. Actionable [KBM21, PJHY20]. Activation [HPRC20]. Active [CZG^{+}22, JLCZ22, TIHS20, VVR^{+}24, ZHL^{+}21a]. Activities [XLZ24]. Activity [SYHS20, VCO^{+}23, WZC^{+}23]. Activity-Guided [WZC^{+}23]. Actors [LXH^{+}21]. Actuated [SOL^{+}22]. Acuity [WSL^{+}24]. Ad [SWS^{+}23, Wu22]. Ad-hoc [SWS^{+}23]. Adapt [QSC^{+}21]. Adaptable [ERB^{+}21, ieSM^{+}23]. Adaptation [LJCL24, SCR^{+}23, SWS^{+}23]. Adapted [LHWW22]. Adapter [JGH^{+}24]. Adapting [LHL^{+}20, PHB^{+}22]. Adaption [WYS^{+}22]. Adaptive [AYGR22, BHM^{+}22, CGD^{+}24, LTC21, LWSY20, MDL^{+}23, MTVS23, NI22, NSW24, SXW^{+}22, eSYKW23, WQP^{+}22, WITW22, WGH^{+}24, XF21, YLJ^{+}22, YWB20, ZHQH20, ZWL23, ZCZ^{+}23a, ZCR21, ZZ23]. Adaptively [WXS^{+}24]. Add [ZHKY23].

Add-on [ZHKY23]. Adding [LALG22]. Addressing [ZLVV22]. Adjacency [KKhCM23]. Adjustment [YLL^{+}22a, ZZW^{+}22a]. Administrative [RBSN22]. Advanced [GG21, HLL^{+}24]. Advances [FLS23, TG24, WCWQ22]. Adversarial [CFZ^{+}23, HGO21, LB22, MXLM20, OYK^{+}21, WCX21a, WZC^{+}23, ZHDX20]. Aesthetic [HIDI23]. Aesthetics [KJS^{+}23]. Affect [PLD^{+}23, VVR^{+}23b]. Affective [AR22, ECR22, LSZC21, LSW^{+}22, LWC24, LRA23, LLMA22]. AffectiveTDA [ECR22]. Affects [GKC^{+}24]. Affordances [VVR^{+}23b]. Afforded [FFB24]. Against [WBL^{+}22]. Age [SES20]. Agencies [ZKS^{+}20]. Agency [FOH^{+}21, LKL23]. Agenda [SBW21, YBVI22]. Agent [MM23, RLB^{+}21, RXX^{+}21, WGM^{+}24]. AgentDress [WCC^{+}21]. Agents [GGP^{+}24, WBA^{+}23, WCC^{+}21]. AgentVis [RLB^{+}21]. Aggregate [HE24]. Aggregated [BPA22, CWW^{+}23, LZM20, VBP^{+}21, ZFCG23]. Aggregation [GZW^{+}20, XF21]. Agnostic [BvdPLH22, HRS^{+}22, PBM^{+}24, Tsa21]. AI [CRM^{+}24, HMTI24, SBe^{+}21, SKK^{+}22, WTL24]. AI-Assisted [CRM^{+}24]. AI-Powered [WTL24]. AI4VIS [WWS^{+}22]. aid [RZH^{+}23]. Aided [DMBK21]. Aids [TAL24, YQN^{+}21]. Air [BS22, DWC^{+}20, DZG^{+}23, FHR^{+}21, HXL^{+}24, HDFK21, KG24, PM23, SDK23a, WBG^{+}23]. Aircraft [NFN^{+}23]. AirVis [DWC^{+}20]. Algebra [Wu22]. Algorithm [BD22, FPG^{+}23, GVT24, JSF20, LHWW22, SRBP20, SPN23, ZCZ^{+}23a, ZCH^{+}21]. Algorithmic [GM23, HM22, XMK^{+}22]. Algorithms [ADM^{+}22, CMKK21, ZMD^{+}22]. Align [RRG23]. Alignment [CMKK21, CFZZ23, MWJ22, RBLT^{+}22, WWW^{+}24, WCCS24, WBM21a, WHBC23, YJ21, BNWvW21]. Alignment-based [WBM21a]. Alignments [KTL24]. Alive [EGMP24]. Alleviate [HHC^{+}21]. Almost [NIK24]. Aloud [FWZ^{+}20, SKFZ22]. Alpha [KLCK20]. Alpha-complex [KLCK20]. Alps [BSB^{+}20]. Alternative [CNK^{+}24, JMK^{+}22]. Alternatives [LBE20]. always [LLS^{+}20]. am [GWF^{+}23]. Ambient [ER21, GG21]. Ambiguities [XPK^{+}24]. Ambiguity [JQL^{+}24, NB24, XYF^{+}21]. Ambiguous [WAA^{+}22]. Ambisonic [BBMM^{+}23]. AMM [BHM^{+}22]. Among [FOH^{+}21, KKS^{+}22]. Amount [RBRG21]. AMR [WZU^{+}21]. Analyses [LKAH21]. Analysis [ABE^{+}22, AAA^{+}21, APSB23, BSB^{+}20, BS22, BS21, BFS^{+}24, BDL^{+}21, BER^{+}23, CSJ^{+}21, CJS^{+}22, CBB23, CSL^{+}23, CYL^{+}21, CKQ^{+}23, CLL^{+}23, CSBK20, CNY22, CvW23, CFGT21, DSKE21, DPR^{+}20, DWX^{+}22, EBPF21, ECR22, EBW23, FLZ^{+}21, FW22, FNB^{+}22, FKSP^{+}24, FKM20, FWZM22, FJK^{+}20, GZRP^{+}22, GSS^{+}20, GZW^{+}20, GGPL20, GGJ^{+}22, HW20, HW22, HLW^{+}20, HWN^{+}23, HPT^{+}23, HRS^{+}22, HCM^{+}22, HCL20, JLK20, JGC^{+}21, KGQ^{+}24, KHL21, Kel23, KRZ^{+}20, KRW^{+}21, KKJ^{+}21, KKT^{+}22, KBJ^{+}20, KJI^{+}21, KSHW22, LM20, LCC^{+}20, LWY^{+}20, LPG^{+}22, LLC^{+}22, LCS^{+}24, LPL^{+}24, LWWH20, LAC^{+}24, LWM^{+}20, LBE20, LWT^{+}23, LAS^{+}20, MTW^{+}20, MM21, MBB20, MGWM22, MGU^{+}21, NQE21, NCWE22, NRA^{+}23, NDLW20, NFN^{+}23, NSG^{+}20, OWN^{+}21, PGS21, PBM^{+}24, PVT23, QRZZ21, RBK^{+}21, RPD21, RLB^{+}21, RBF^{+}23, SUS^{+}21, SBT^{+}23, SSS20, SFNRZ^{+}23, SBk^{+}24, SYW^{+}20, SHL^{+}21, SCC^{+}24, SFS^{+}22, SKFZ22, SaT^{+}23]. Analysis [SN23, SKR^{+}24, TBL^{+}20, TWC^{+}24, WHX^{+}23, WEM^{+}21, WAP^{+}21, WFY^{+}21, WXC^{+}21, WHJ^{+}22, WMH^{+}22, WBL^{+}22, WCX^{+}23, WYIS24, WKN^{+}23, WLT^{+}24, WQ20, XRN^{+}23, XZS^{+}22, XHL^{+}23, XDW21, XWY^{+}20, XGS^{+}21, XDH^{+}22, YST^{+}20, YGT^{+}23, YLTL23, YBL^{+}20, ZHN24, ZWW^{+}20, ZWW^{+}23a, ZSM21, ZSZ^{+}24, ZFF22]. Analytic [MHS^{+}22, NSS21, NCWE22, WR23, ZWW^{+}23b, ZWG^{+}23, ZWV21]. Analytical [GGT^{+}20, RQ21, WWZ^{+}22, ZeB^{+}21]. Analytics [AL20, ADL^{+}22, BCC^{+}20, BBRE24, BPW^{+}21, BJR21, CPCS20, CXD^{+}21, CAGM22, CCEA^{+}24, CMKK22, CLA^{+}20, CYP^{+}20, COZ^{+}23, CWO^{+}24, CCM20, CCW^{+}21, DSD^{+}23, DWC^{+}20, DWL^{+}22, DMBK21, DV23, EGP^{+}21, EAS^{+}23, FAS^{+}21, FCH^{+}23, FWM^{+}24, FP21, FSS^{+}21, GZL^{+}21, GLL^{+}24, HPK^{+}22, HZS^{+}22, HWW^{+}24, HZP^{+}24, HHS^{+}23, HLCY22, HZJ^{+}24, HHM^{+}24, JCZ^{+}24, JWW^{+}23, JLP^{+}23, JYLS21, KKEG20, KKZE20, KKJ^{+}21, KAS^{+}21, LLT24, LKJ^{+}20, LMF^{+}24, LWS^{+}21, LM21, LWL^{+}22b, LWD^{+}23, LYBP23, MXLM20, MFH^{+}21, MBS^{+}21, MHN^{+}24, MFS^{+}23, NKWW22, NDP^{+}21, NHC^{+}20, NB24, OCL^{+}21, PNB^{+}21, PNKC21, PMCM24, PYSJ21, PJHY20, QTW^{+}24, SDC^{+}24, SPM24, SPZS24, SEAB^{+}22, SFS^{+}22, SJK^{+}23, SCEA23, SSSEA20, SHC^{+}20, TSHI22, VKT^{+}24, WSN21, WSN22, WZD^{+}20, WMW^{+}22, WZY^{+}22, WWZ^{+}23a, WMH^{+}23, WLZ^{+}23, WQQ^{+}24, WZW^{+}23, WZD^{+}21, WWS20, WLG^{+}22, WDC^{+}23, WDX^{+}23, WXG^{+}24, XVW^{+}21, XMT^{+}21, XDW21, XMK^{+}22, XNF^{+}23, XZKM22, YZF^{+}23, ZLL^{+}21, ZWZ^{+}23]. Analytics [ZKS^{+}20, ZLL^{+}20, ZWW^{+}23c, leáB^{+}21]. Analyze [CCM20, JVRL24]. Analyzing [CLS^{+}21, DWB21, EBJ^{+}22, FYE^{+}22, FS24, NMS^{+}23, SPZS24, VBP^{+}21, YCC^{+}21, dSBdO^{+}24]. Anatomies [KSK^{+}23]. Anchors [RAFSA23a, RAFSA23b]. Angiographic [DGKOC20]. Anglewise [QCCC23]. Angular [ZGL^{+}21]. Animal [DHF^{+}22]. Animal-Like [DHF^{+}22]. Animals [NBNC20, NBE^{+}21]. Animatable [MBHE24]. Animated [CRP^{+}20, KH21, SKH^{+}23, ZPWS23]. Animation [BLIC20, CSWZ24, CFZ^{+}23, CSCM21, DCC22, GFSHO20, JJKJ20, LSW^{+}22, LZX^{+}21, LHL^{+}22, SZZW24, ZPWS23]. Animations [KVGM20, LFW^{+}22, XQXL23]. Anisomorphic [DHLA22]. Anisotropic [XF21]. Ankle [OF22]. Annotated [CCS^{+}24, DWS^{+}23]. Annulus [XWH^{+}23]. Annulus-Constrained [XWH^{+}23]. Anomalies [YST^{+}20]. Anomalous [XWY^{+}20]. Anomaly [GJC^{+}22]. Answering [WJBB22, YCHL24]. Anthropographics [MJAD22]. Anti [GLY^{+}23, ZSZ^{+}24]. Anti-blink [ZSZ^{+}24]. Anti-Bullying [GLY^{+}23]. Antisplat [NNF^{+}20]. APF [CHSC24]. APF-S2T [CHSC24]. Appear [KBPR22]. Appearance [AYA^{+}21, FALH20, MXF^{+}21, ONH21, XWZB21]. Applicability [KCWK20]. Application [JGG21, PRKM20, XWZ^{+}22]. Applications [ADM^{+}22, Ano22c, BBVS^{+}24, BWCT23, CGAG20, CFL^{+}23, FMBN23, GGPL20, HLL^{+}24, LWM^{+}20, SDMK22, XDW21, ZGL^{+}21, DJBJ23]. Applied [BDRW21, MBB20]. Applying [NDLW20, WCWQ22, YXL^{+}22]. Approach [AKS22, BWCT23, BLE^{+}23, CXZ^{+}24, CCEA^{+}24, CWO^{+}24, CCL^{+}24, CWH^{+}22, HBH^{+}20, HZX^{+}21, HZS^{+}22, HTY^{+}23, HWW^{+}24, HHK^{+}24, HHS^{+}23, HBS^{+}21, HZC^{+}20, HLCY22, JWW^{+}23, KRS^{+}22, KLB24, LWZ^{+}22, LI24, NDP^{+}21, PLL^{+}24, PWK21, QFWS22, RWJ^{+}23, SCRL20, SFS^{+}22, SHC^{+}20, TZT^{+}22, VGT21, WAA^{+}22, WJW^{+}20, WPL^{+}22, WBL^{+}22, WCX^{+}23, WLZ^{+}23, WQQ^{+}24, WLT^{+}24, WZD^{+}21, XZF^{+}22, XDW21, YIIW24, ZOS^{+}23, ZLG^{+}21a, ZWW^{+}23c, dSBdO^{+}24, DJHBJ21, KMWD21]. Approaches [CMKK22, RBF^{+}23, WWS^{+}22, BTv^{+}23]. Approximate [JSF20]. Approximated [FPG^{+}23, JKU^{+}22]. Approximation [AA21, LSL^{+}22, Pat22]. AR-Loupe [QSUK22]. Arbitrary [AMAS21, TGM21, YCH^{+}22, ZZZ24]. Arbitrary-Scale [YCH^{+}22]. ARC [WBM21a]. Archaeological [PVPM22]. ARCHIE [LELT23]. Archimedean [WCZ^{+}20]. Architectural [BHU^{+}21, XCSJ22, ZPF^{+}22]. Architecture [BvOR21, HM24, LZZ^{+}21, LWSY20, VSBY22, YITS23, YLTL23]. Architectures [CPCS20, WYIS24]. ArchiText [KDEP21]. Area [Ano22y, DTPG21, GMX23, JdJTC24, LIDM20, WNV22]. Area-Aware [LIDM20]. Areal [ZLL^{+}21]. Areas [MPNF24]. Argumentation [KRW^{+}21]. Arguments [SEAKC21]. ARGUS [CRM^{+}24, WEM^{+}21]. Arigat [WQP^{+}22]. Arjun [Ano22a]. Arm [FRL^{+}23]. Arrangement [FFB24, ZHL^{+}21a]. Arrangements [XSB^{+}22]. Array [BJCL21, RDH23]. Arrays [SZH^{+}20, RSAA20]. Arrows [RLG^{+}23]. Art [PSS^{+}22, PTS^{+}20, RCD^{+}23, SZC^{+}23, WH23, YLL^{+}22b]. Artery [PSY^{+}20]. Articles [LWY^{+}20, YQN^{+}21]. Articulated [LLK22]. Artifact [JSA^{+}20, KBM21]. Artifact-Based [JSA^{+}20]. Artificial [CHSC24, GVN^{+}20, WWS^{+}22, WHBC23]. ASD [LZZ^{+}24]. Asking [HXHT23]. Aspect [CMXF21]. Assembling [XZS^{+}23]. Assembly [GEU^{+}22, SMSK22, WAV^{+}21]. Assess [GZL^{+}21]. Assessing [WBWL24]. Assessment [CMK20, CGD^{+}24, FRiM^{+}23, JKW^{+}22, LSD^{+}23, MSK23a, NB24, QCCC23, QLC^{+}24, SMYF22, ZCM^{+}23b]. Assessments [LRHA22]. Asset [ZPF^{+}22]. Assets [GEU^{+}22]. Assigning [San20]. Assignment [LXF^{+}22]. Assistance [FYC^{+}23, LALG22, MSBV^{+}23, PMN^{+}23, WWZP22]. Assisted [DSD^{+}23, HLW^{+}20, HRX23, JGH^{+}24, KKGMH21, LGL^{+}23a, MMG^{+}21, MCSAL23, RZLX24, CRM^{+}24]. Associate [MB20c, MB21b, MB22b, SK23b, SK24]. Associated [CWSJ23]. Association [HYC^{+}23, LWL^{+}20]. Associations [RLLS20, SCL^{+}23a, TAL24, YWL^{+}20]. Assumption [ZZW21]. ASTF [ZGX^{+}23]. Astrographics [BAC^{+}20]. Asymmetric [FKSP^{+}24, HZYZ22, KRZ^{+}20, YGE^{+}21]. Asynchronous [LFCH24, XGS^{+}21]. Atlanta [PBAG23]. Atlanta-World [PBAG23]. Atmospheric [CBE^{+}21, PSH21]. Atrium [NGBA^{+}20]. Attacks [WBL^{+}22]. Attempt [LWD^{+}23]. Attention [DWB21, DCWD23, JLP^{+}23, LLSM24, LB22, LWD^{+}23, LZL^{+}23, MCS^{+}23, NBE^{+}21, QCCC23, TQW^{+}24, YGT^{+}23, YCW^{+}24]. Attention-Based [TQW^{+}24, YGT^{+}23, JLP^{+}23]. Attentional [JKL24]. AttentionViz [YCW^{+}24]. Attenuation [HHI24, KHI20]. Attribute [CMSK23, HLM^{+}20, LYZ^{+}21, LRZ^{+}23, LGWL21, MZS^{+}24, RPG23a, RGDG23, SSJ^{+}22]. Attribute-Aware [MZS^{+}24]. Attribute-based [SSJ^{+}22]. Attribute-Conditioned [LYZ^{+}21]. Attributes [BZSD21, CNY22, HJZ^{+}21, LGWL21, MM20, SLW^{+}24, PGL^{+}20]. Attribution [HPRC20]. Audio [BBVS^{+}24, CSWZ24, FPG^{+}23, GPR^{+}24, RRK^{+}22, TBL^{+}20, WWR^{+}20]. Audio-Driven [CSWZ24, WWR^{+}20]. Audio-Visual [BBVS^{+}24]. Auditing [XMT^{+}21]. Auditory [GTH20, GKN^{+}23, KL22, KG24]. Augmentation [CRWA22, CXD^{+}21, UIHS20, ZJJH21]. Augmentation-Driven [ZJJH21]. Augmented [Ano23g, APP^{+}22, BBL^{+}22, BFS^{+}23, CSW^{+}20, EBW23, FPG^{+}23, FCFC22a, FCFC22b, FKSP^{+}24, GSM^{+}22, GEU^{+}22, GLH22, GPR^{+}24, HZZ^{+}20, HFS^{+}21, HSJ^{+}20, JLH24, KNL23, KPW20, LSS24, LTLB22, LZY^{+}23, LZX^{+}22, LXZ^{+}23, LWL^{+}23b, LWW^{+}24, MA23, MTE^{+}20, MSA^{+}22, MMN^{+}22, MHIS23, PIS20, QSUK22, RS23, RMW^{+}24, RFD21, RTM^{+}20, RLG^{+}23, SWB^{+}22, SML^{+}23, SDMK22, SMPJ^{+}20, SES20, SMC^{+}21, SKS^{+}23, SMSK22, TNJ^{+}22, TCX^{+}23, TBW^{+}23, TME^{+}22, VVR^{+}23a, VBV^{+}23, VLM^{+}23, WZL22, WYS^{+}22, WQP^{+}22, WG21, WGO20, WGO22, WR23, YEP^{+}22, ZZW21, ZHH22, ZYL^{+}24, ZSCRB23, ZG20, ZWP21, ZCM^{+}23b, ZLG^{+}21b]. Augmented-Reality [ZZW21]. Augmenting [ASP^{+}22, BKS22, CYC^{+}22, CYX^{+}23, CZGF21, LSW^{+}22, LCY^{+}23]. Authentication [JJKJ20]. Authenticity [LCW^{+}23a]. Author [Ano22f]. Authoring [BGS^{+}22, CSW^{+}20, KRHH24, LKAH21, MMG^{+}21, MBS23, OBCT24, RAP21, SLR^{+}20a, SMSK22, THS^{+}21, WHS^{+}23, WTL24]. Authors [HCWK23, Hul20, PSH20]. Autism [GGP^{+}24, RNO^{+}22]. Auto [CWW^{+}20]. Auto-Extraction [CWW^{+}20]. Autoencoder [LWL^{+}23a]. Autoencoders [DMJ^{+}22, YGT^{+}23]. Autographic [Off20]. Automated [CAGM22, CWW^{+}20, KRS^{+}22, PTS^{+}20, SKH^{+}23, TGM21, ZXC^{+}23]. Automatic [CMKK22, CDZ^{+}23, CCS^{+}24, CZW^{+}20, HPM^{+}24, LYZ^{+}21, LFR^{+}21, LSW^{+}23, MZX^{+}21, PDD^{+}22, PAAG22, QSC^{+}21, QZZ22, QNWR23, SSB^{+}22, SZZW24, SXS^{+}21, WSZ^{+}20, WTY^{+}22, XLL^{+}20, YISG21, YTL^{+}22, YSD^{+}23]. Automatically [BvOR21, YGH^{+}23]. Automating [SWS20, ZTL^{+}23]. Automation [DV23]. AutoML [OCL^{+}21]. Automultiscopic [MSK23a]. Autonomous [HWW^{+}22, JZHA22, SWF^{+}24, WLZ^{+}23]. AutoSweep [CLL^{+}20]. Autotuning [HPT^{+}23]. Auvil [Ano22c]. Auxiliary [RLG^{+}23]. Avatar [AI21, BBH24, BCB22, CFDN24, DPM24, FALH20, GLH22, GFSHO20, KBF22, LKL23, LCW^{+}23a, MWD^{+}23, ONH21, PT20, SAMB^{+}23, WYS^{+}22, YeSiK^{+}23, YGE^{+}21, ZCZ22, MSAM^{+}22, VAWL24]. Avatarization [VVC^{+}24]. Avatarizing [VVR^{+}23a]. Avatars [BVV^{+}22, BVV^{+}23, CFL^{+}23, JJKJ20, KPL23, KSK^{+}23, PTC^{+}24, PLD^{+}23, SWF^{+}24, WBA^{+}23, YYK^{+}22, YGE^{+}21]. Average [HWS22, MPNF24, XCLF20, YWM^{+}20]. Averaging [AMAS21]. aversion [WKMD22]. Avoidance [MMN^{+}22]. Avoidant [HWT^{+}24]. Avoiding [ZA21]. Award [Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Pla21, Ano23a, Ano23b, Ano23c, Ano23d, Ano23e, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Fek21, GF22, Kiy22b, Kiy22a, Kli22, SW21, SAB20]. Awards [FG22]. Aware [AYGR22, BHY^{+}23, GSS^{+}20, GLL^{+}22, GZG^{+}24, HKW23, HBMK23, HSC^{+}22, JZHA22, KKE21, KLSB22, LIDM20, MCW^{+}20, MXF^{+}21, MZS^{+}24, NP24b, PFN22, QHL^{+}20, STD^{+}23, SJL^{+}23, TBL^{+}20, XXM^{+}21, YBR^{+}23, ZIX^{+}24, ZLX23, ZYC^{+}24b, ZHF^{+}20, SPW^{+}22, TWW^{+}22, WXW^{+}20, WCZ^{+}20, XLL^{+}22, ZSS^{+}21]. Awareness [BRLR24, FTWP22, HTP^{+}23, JLH24, LFG^{+}23, NCWE22, RBR20, RWJ^{+}23, SLK^{+}20, TME^{+}22, WR23, ZLL^{+}20]. away [XLL^{+}22]. Axes [TEK^{+}23]. Axis [EML^{+}23, LLL^{+}22].

b [MNPP23]. b/Surf [MNPP23]. BA [YLL^{+}22a]. Bach [Ano22d]. Back [XCG^{+}24]. Back-and-forth [XCG^{+}24]. Backed [SKK^{+}22]. Background [ZSL^{+}22]. Bacon [MGWM22]. Bad [Red23]. Badminton [CXY^{+}22, LAC^{+}24]. Bag [RAFSA23a]. Balance [CCS^{+}21, MCQ23, STD^{+}23, SSC^{+}23]. Balance-Aware [STD^{+}23]. Balanced [HJL^{+}23, XGS^{+}21, XGS^{+}23]. Balancing [LLC^{+}20, WNV22]. ball [WDX^{+}23]. Bandwidth [LCW^{+}23b]. Banners [BSB^{+}20]. Bar [FFB24, PRJ^{+}23, XCLF20, XSB^{+}22, YDM^{+}21]. BarcodeTree [LZD^{+}20]. Bare [LCM^{+}23]. Barriers [ASSB^{+}23]. Barycenters [PVDT22, VBT20, WPTG24]. Bas [YCZ^{+}22, ZLZ^{+}23b]. Bas-Relief [YCZ^{+}22, ZLZ^{+}23b]. Base [SPM24]. Based [ASA^{+}23, ACL^{+}24, AYA^{+}21, AMK^{+}21, AYG^{+}22, BMWD20, BPA22, BJR21, CLH^{+}23, CSW^{+}20, CWW^{+}21, CFZZ23, CHSC24, DWQW23, DFP^{+}20, DXX^{+}21, DHM^{+}22, DZG^{+}23, DTPG21, EBW23, FiMH21, FWZZ20, FE21, FW22, FGS^{+}22, GM23, GFCM23, GPR^{+}24, GL20, HW23, HYH^{+}23, HLA^{+}22, HAK^{+}23, HLJ^{+}22, HHK^{+}24, HLZ^{+}20, HWL^{+}22, HLCY22, HZQ22, HLH^{+}23, JZHA22, JKA^{+}24, JZCZ20, JSA^{+}20, KCGZ23, KVGM20, KW23, KAS^{+}22, KMM24, KGR^{+}24, LLD^{+}21, LQWQ21, LWZ^{+}22, LS23, LCS^{+}24, LGH^{+}24, LYZ^{+}24, LFR^{+}21, LCWL23, LPP^{+}23, LXZ^{+}23, LIFD23, LSW^{+}20, MBB20, MXT^{+}22, MA20, MHN^{+}24, MSK23b, NLPW22, OBCT24, OKM21, PFC20, PTD^{+}21, PTS^{+}20, QGL^{+}23, QTW^{+}24, QR22, RNO^{+}22, SCRL20, SYW^{+}20, SYHS20, SZF^{+}21, SS24, SDR22, SZC^{+}23, SSX^{+}20, SYX^{+}22, SMPJ^{+}20, SAK20, SDK23b, SMS^{+}22, TNJ^{+}22, TNZ^{+}24, TPH22, TQW^{+}24, TEK^{+}23, WHX^{+}23, WWSS20, WCX21b, WZY^{+}22, WHS^{+}23, WWW^{+}24, WCTW21, WH22b, WWZ^{+}22]. Based [XZWY20, XLL^{+}20, XCSJ22, XXM^{+}21, XLY^{+}22, XQXL23, XNF^{+}23, XHFZ24, XWPG^{+}24, XPK^{+}24, YGT^{+}23, YWB20, ZLY22, ZWH^{+}22, ZLZ^{+}23a, ZLW^{+}21b, ZZW^{+}22b, ZOF^{+}23, ZZZ24, ZJJH21, ZCM^{+}23b, BNWvW21, CDX^{+}20, CWW^{+}20, CXXZ21, CZF^{+}22, DMTD22, FBW21, HWC23, HJL^{+}23, HZC^{+}20, JLP^{+}23, KHD^{+}22, LHA^{+}21, LZL^{+}23, NDP^{+}21, NSK^{+}21, PCZ^{+}21, PSS23, QSC^{+}21, RPD22, SXL^{+}23, SPW^{+}22, SSJ^{+}22, VGK^{+}22, WZD^{+}20, WAP^{+}21, WMH^{+}22, WHC^{+}23, WHL^{+}20, WBM21a, XMT^{+}21, YLL^{+}22a, YHC^{+}24, YTL^{+}22, YSD^{+}23, ZWW^{+}23b, ZWZ^{+}22, WMH^{+}23]. Baseline [GTHC20]. Basketball [FS24, LCY^{+}23, TPH22, WDX^{+}23]. Baton [CFGT21]. Bayesian [Hei21, KMWD21, KKGMH21, MGO21, MSBV^{+}23, RZLX24]. Bayesian-Assisted [KKGMH21]. BCI [SMPJ^{+}20]. BCI-Based [SMPJ^{+}20]. Be [KBPR22, KPL23]. Beamforming [GVN^{+}20]. Beaming [ATHI24, HWMI23, IKi21]. beating [MSAM^{+}22]. Beautification [XYF^{+}21]. BeauVis [HIDI23]. Become [BPL23]. Behavior [BCN^{+}20, HPdlG20, JVRL24, KVB20, MMS^{+}23, NCWE22, RLB^{+}21, SPJ^{+}23, WGO20, WGO22, YBOB24, CNB^{+}22]. Behavioral [BFS^{+}24, JWE^{+}22]. Behaviors [CAA^{+}20, LGH^{+}24, SUB^{+}22, XVW^{+}21]. Behaviour [HDFK21, NHC^{+}20, SPZS24]. Behaviours [GKN^{+}23]. Behind [SBe^{+}21, ZSG^{+}23]. Being [GLH22, KSK^{+}23]. belief [KMWD21, XSKF23]. Believe [ISBP22, XSKF23]. Believing [XSKF23]. Bell [NCE23]. Benchmark [GVT24, WGM^{+}24]. Bending [PIS20]. Benefit [MGM^{+}22, NF20]. Benjamin [Ano22d]. Berkeley [NGW^{+}24]. Best [Ano22z, Ano23v, Ano24s, SW21, SAB20]. Beta [KLCK20]. Beta-complex [KLCK20]. Better [CWW^{+}22, DWX^{+}22, LWY^{+}20, WZD^{+}21, XWL^{+}21, YCB^{+}21]. Between [AKS22, CLD^{+}22, FALH20, JLH24, MA20, RAC22, SHOP23, VLM^{+}23, YDM^{+}21, CLA^{+}20, KVB20, LZM20, SMKN20]. Between-View [SHOP23]. Beyond [AMK^{+}21, CFL^{+}23, DUWW22, KHL21, PvSvdE^{+}24, RMW^{+}24, WZZ^{+}23, WCH^{+}22, ZZW21]. Bézier [MNPP23]. BGP [RCD^{+}23]. Bhattacharyya [JTT^{+}23]. Bi [SDR22]. Bi-Scale [SDR22]. Bias [BWZ^{+}20, BZKG21, CFQ21, CQ22, GKC^{+}24, GM23, HMGO23, HLCY22, PGS21, WMZ^{+}20, GM23]. Biased [MSWI22, MPNF24, XCLF20]. Biases [CMXF21, DFP^{+}20, HWS22, ISBP22, PMS^{+}22, WNC^{+}22, XSKF23, MHFF21]. Biclusters [SSAZ22]. Bidirectional [QLFG22, XCG^{+}24]. Big [THS^{+}21]. Bilateral [XZXY22]. BIM [SML^{+}23]. Bimanual [GST^{+}24]. Binary [DAK^{+}21]. Binned [ZSL21]. Binning [QBW^{+}20]. Binocular [APP^{+}22, DMTD22, ZA21]. Biologging [MBS^{+}21]. Biologists [DSKE21]. Biology [LT20]. Biomechanics [EBKL21]. Biosignals [GZG^{+}24]. Biosignals-Driven [GZG^{+}24]. Bipartite [ASA^{+}23, LWH^{+}22, ZSCC22]. BiRD [XCG^{+}24]. Bisector [YZEN22]. Bitcoin [TSHI22]. Bivariate [AJSP23]. Black [HZII24, VE21]. Black-and-White [HZII24]. Black-Hole [VE21]. Blank [Ano22g, Ano24g]. Blazars [SUB^{+}22]. Blending [GSM^{+}22, WXW^{+}20, ZGX^{+}22]. Blendshape [WLF^{+}22]. Blind [GKN^{+}23]. Blindness [MSGM23]. blink [ZSZ^{+}24]. Blinking [LH23]. Block [LCWL23, SEK^{+}24]. Blockchain [THFI21]. Blocks [YLJ^{+}22]. Blood [MNB^{+}23]. Blur [APP^{+}22, LSE20, PALW20]. Blurring [NDLW20]. Board [FMBN23]. Boba [LKAH21]. Bodies [CBE^{+}21]. Body [ASP^{+}22, BBVS^{+}24, CGAG20, CFL^{+}23, DCC22, DHLA22, FOH^{+}21, FJK^{+}20, GMTD23, JZCZ20, LHZ^{+}23, SWSK23, UVL^{+}23, ZLC^{+}22, JLM^{+}21]. Body-Centric [GMTD23]. Body-Supporting [ZLC^{+}22]. Bokeh [JBS^{+}22]. Bonded [LLCH22]. Book [XDBAR24]. Borkin [Ano22d]. borne [SaT^{+}23]. Both [IAI^{+}23]. Bounce [XZXY22]. Boundaries [BSG^{+}20, DAB^{+}23, GTH20, LWWH20]. Boundary [BKWK20, MA20, ZLZ21]. Bounded [YLGW24]. Box [ZWZ^{+}22]. BOXRR [NGW^{+}24]. BOXRR-23 [NGW^{+}24]. Brachial [CNC^{+}20]. Braiding [BZSD21]. Brain [GBM^{+}22, SHT^{+}22, VBV^{+}23]. BRDF [CBP22, GGT^{+}20, GGPL20]. BRDFs [SDR22]. Breaking [BSG^{+}20, BWWL22, GWD^{+}24, SGJC23]. breathing [MSAM^{+}22]. Bridges [CAA^{+}21]. Bridging [CLA^{+}20]. Brightly [YIIW24]. Brightness [KLSB22, SES20]. Broad [LHWW22]. Browser [CNC^{+}20]. Browsing [OM22, iIK^{+}21]. Bubble [WCW^{+}22]. Bubbles [WCW^{+}22]. Buckles [ZH20]. Budget [ZWH^{+}22]. Builder [CLX^{+}23]. Building [EHB^{+}23, IMQ^{+}20, dSBdO^{+}24]. built [BXQ^{+}22]. Bullying [GLY^{+}23]. Bundle [YLL^{+}22a]. Bundles [OM22]. Bundling [GMS^{+}21, LLC^{+}20, WAA^{+}22, WXW^{+}20, ZPG21]. Bus [WZD^{+}21]. Business [YDMP22].

C [MSM^{+}22]. C.DOT [TNJ^{+}22]. C3

[LXF^{+}22]. CAEVR [GZG^{+}24].

Calculations [DWH^{+}23]. Calibrate

[XRN^{+}23]. Calibrated

[CLH^{+}23, GWC^{+}23]. Calibration [BVV^{+}22, BVV^{+}23, HLCY22, KBPR22, SII^{+}21]. Call [BKR^{+}24, KBB^{+}23]. CallFlow [NBJ^{+}21]. Calliope [CCS^{+}24, SXS^{+}21]. Calliope-Net [CCS^{+}24]. Camera [CZF^{+}23, HGO21, HL21, LXF^{+}22, LCC^{+}23, SWY^{+}22, SMYF22, WLH24, XCG^{+}21, YITS23, TWA22]. Cameras [LCC^{+}23, ZSZ^{+}24, ZGL^{+}21]. Can [BVV^{+}23, CBB23, CBW23, HZS^{+}22, HHC^{+}21, HX23, HB24, RS23, MSAM^{+}22]. Cancer [FNB^{+}22]. Canonical [Liv21]. Can’t [YPW23, LLS^{+}20]. Capabilities [WHJ^{+}24]. Capable [LNB^{+}21, WH22a, KPW20, ZHKY23]. Capacitive [dPLM21]. Capacity [JKL24]. Caption [CWW^{+}22, KCK^{+}24]. Captioning [ZXC^{+}23]. Capture [CWL23, HLL^{+}24, KGX^{+}23, KJI^{+}21, NGW^{+}24, RZH^{+}23, SWY^{+}22, WCX21b, WPZ^{+}23, YHC^{+}24]. Cardiac [ENvBC23, MNB^{+}23]. CardioGenesis4D [SKW^{+}23]. Care [FWM^{+}24, GKC^{+}24]. Career [Pla21, WLS^{+}22]. Careers [WPL^{+}22]. Caricature [HHD^{+}20, YXS^{+}23]. CaricatureShop [HHD^{+}20]. CariGAN [YXS^{+}23]. Carlo [CS23, RGG20a, ŠK20]. Carpet [MSRJ20]. Cartograms [DTPG21, NSM^{+}22]. Cartographic [JHS^{+}21]. Cartoon [CZG^{+}22, LZLS22, XQXL23]. Cartoonization [SYX^{+}22]. Cascade [DWL^{+}22, LWY^{+}20, NSG^{+}20]. Cascaded [KHI20, LCK^{+}21]. Case [BBH24, FPG^{+}23, GVN^{+}20, HMH^{+}21, PCQ^{+}20, PVPM22]. Cast [ASCR^{+}22, JRM22]. Casting [LME^{+}23]. Casual [EGP^{+}21, CZF^{+}23]. Categorical [AR22, LFC^{+}21]. Categorization [CAA^{+}20, LST^{+}22, QTW^{+}24, ZWV21]. Category [CLZ^{+}24, LWF23, MHFF21]. Category-Level [CLZ^{+}24]. Category-Specific [LWF23]. catheter [BTv^{+}23]. Causal [DWX^{+}22, KWH22, WM23, XDW21]. Causality [CSC^{+}21, GM23, HM22, JGC^{+}21, LCS^{+}24, XSHF20]. Causality-Based [GM23, LCS^{+}24]. Cause [PYSJ21]. Cautious [HWT^{+}24]. CAVA [CXD^{+}21]. CAVE [HXL^{+}24, JK23]. CAVE-Enabled [HXL^{+}24]. cavity [BTv^{+}23]. CcNav [DAK^{+}21]. Celestial [CBE^{+}21]. Cell [CKQ^{+}23, HLW^{+}20, HMTI24, PGL^{+}20, SIL^{+}21]. Cells [HBS^{+}21]. Cellular [SVK^{+}21]. Center [Ano22c]. Centered [ALR23, ENM24, YSM^{+}20]. Centeredness [LWC22]. Centric [GMTD23, WZC^{+}23, WHC^{+}23, WSL^{+}24, YeSiK^{+}23]. Centroid [LYH^{+}23]. Cerebral [PSY^{+}20]. CerebroVis [PSY^{+}20]. Chaff [DWOB20]. Chain [ŠK20, SKNŽ20]. Chair [FG22]. Chairs [BPQW23, IMKP21, IGMW22, MZX^{+}21, NAW^{+}22, SES23, Ano21k, FGL^{+}23, HKVZ20, IPPZ24, PK21]. Challenges [BKR^{+}24, WFW^{+}20, ZLVV22]. Chamber [GSS^{+}23]. Champaign [Ano22c]. Change [LCYQ24, MSGM23]. Changes [CCPM23, WFW^{+}20]. Changing [PPE23, WLG^{+}22, ZJS^{+}22]. Channel [KBJ^{+}20, LZP^{+}20]. Channels [MYBF22]. Character [DKM24, DHF^{+}22, ZYC^{+}23]. Character-Oriented [DKM24]. Characteristics [Kel23]. Characterization [CFL^{+}23, HZII24]. Characterize [ZYC24a]. Characterizing [HMH^{+}21, LH22]. Characters [LLK22, RBK^{+}21]. Chart [FZC^{+}21, KCK^{+}24, LLY^{+}24, XHL^{+}24, YDM^{+}21, SNBC23]. Chartem [FZC^{+}21]. Charts [CLW^{+}24, CWH^{+}22, FFB24, LSW^{+}22, LZK^{+}22, LFW^{+}22, LLW^{+}22b, PRJ^{+}23, RQ21, SCL^{+}24, SSC^{+}23, WDG^{+}20, XSB^{+}22, ZXC^{+}23]. ChartSeer [ZFF22]. ChartStory [ZXC^{+}23]. ChartWalk [SNBC23]. Checking [HCM^{+}22, KGQ^{+}24]. Chemical [SUS^{+}21]. Chemicals [PWB21]. Chemistry [SNH^{+}23]. ChemVA [SUS^{+}21]. Chest [JDZK22]. Chi [TNZ^{+}24]. Chief [Mue20b, MB20c, Mue21a, MB21b, Mue22b, MB22b, Mue23b, She23a, SK23b, She24, SK24]. Children [BVY^{+}23, BZP^{+}20, DGB^{+}22, HNGC21, JCZ^{+}24, LZZ^{+}24]. Chinese [GLL^{+}24]. Chip [VSBY22, WYIS24]. Choice [OCW^{+}24, SJK^{+}23]. Choices [CMSK23, HB24]. ChordLink [ADM^{+}22]. Choropleth [LFMM24, YDM^{+}21]. Choropleths [HLM^{+}20]. Cinematic [HLL^{+}24]. Cinematic-Quality [HLL^{+}24]. Circuit [WLT^{+}24]. Circular [SWZ^{+}23, YTL^{+}22]. Circumferential [MSWI22]. Citations [HCWK23, RPNP23]. Cities [CMF^{+}22]. Citywide [YZF^{+}23]. CLAMS [JQL^{+}24]. Clarity [WSL^{+}24]. Class [HXL^{+}24, MM21, MvdEPV24, CGZ^{+}20, HSV^{+}20, LWL^{+}20]. Class-Constrained [MvdEPV24]. Classes [JKA^{+}24]. Classics [GLL^{+}24].

Classification [APBB24, CGT^{+}24, DAB^{+}23, GEU^{+}22, JLCZ22, KKV22, LWY^{+}22, NP21, PvSvdE^{+}24, RNO^{+}22, SLR20b, SLR21, SBK^{+}20, YTHL23]. Classification-for-Saliency [SLR21]. Classifier [HRS^{+}22]. Classifiers [SMS^{+}22, YYZ^{+}22]. Classroom [ZSW^{+}21]. Cliff [HHC^{+}21]. Climate [DPR^{+}20, MEHD24]. Clinical [LLP^{+}23, MSBV^{+}23, SNBC23]. ClinicalPath [LLP^{+}23]. Clipart [SC22]. ClipGen [SC22]. Clipped [LMGY22]. Close [ALC22, XDBAR24]. Closing [SH24b]. Cloth [MXF^{+}21]. Clothing [CK22, LHZ^{+}23, WCC^{+}21]. Cloud [CWS^{+}20, HPP^{+}20, LELT23, LLC^{+}22, LB22, LFO23, LLDW24, MDH^{+}23, RPHJ20, XWY^{+}20, YCH^{+}22, ZLQH21, ZZD^{+}23a, ZZD^{+}23b, ZZW^{+}22b, ZP24]. Cloud-Enabled [LELT23]. CloudDet [XWY^{+}20]. Clouds [BPA22, CZY^{+}20, HB20, HCL20, LSD^{+}23, LMD^{+}22, QHL^{+}20, WCZ^{+}23, XF21, ZHDX20]. CloVR [ZP24]. Cluster [DSKE21, JKJ^{+}22, JQL^{+}24, KWO^{+}20, QR21, RPD20, ROM^{+}23, XZS^{+}22, XHL^{+}23, ZYC^{+}24b, DSKE21]. Cluster-Aware [ZYC^{+}24b]. Clustered [WNV22]. Clustering [HTW20, JQL^{+}24, PBF^{+}21, QFWS22, QNWR23, RPD21, SLC21, XCSJ22, YWL^{+}21]. Clusters [MSuG^{+}23, JKA^{+}24]. Clutter [LIDM20]. Cluttered [MM23]. CMed [PNB^{+}21]. CNN [BvOR21, CZMR21, HLZ^{+}20, LZZ^{+}21, SYW^{+}20, SLR21, SS21, WTS^{+}21, XZWY20, XZKM22]. CNN-Based [HLZ^{+}20, XZWY20]. CNN-Rendering [SS21]. CNNPruner [LWS^{+}21]. CNNs [WL20]. Co [ATHI24, CMSK23, CNK^{+}24, FPG^{+}23, FOH^{+}21, JLH24, KMH^{+}23, LHC^{+}21, YEP^{+}22, CAA^{+}21]. Co-Bridges [CAA^{+}21]. Co-Dependent [CMSK23]. Co-Embodiment [CNK^{+}24, FOH^{+}21, KMH^{+}23]. Co-Immersion [FPG^{+}23]. Co-Located [JLH24, LHC^{+}21, YEP^{+}22]. Co-Operative [ATHI24]. Coaching [LAC^{+}24]. Coal [LWT^{+}23]. Coal-fired [LWT^{+}23]. Code [BB21, DAK^{+}21, MLBW20, SKNŽ20]. Coded [BKS22]. Codes [HYF^{+}20, NBJ^{+}21]. Cognition [KAL^{+}23, LGY^{+}22, KMWD21]. Cognitive [DFP^{+}20, DST^{+}23, JKL24, LHA^{+}21, LPL^{+}24, LLMA22, PMS^{+}22, YYD^{+}21]. Coherence [SBW21, XWPG^{+}24, ZWW^{+}20]. Coherence-Based [XWPG^{+}24]. Coherent [MSS21, NMC21, NWMC23]. Cohort [MNB^{+}23, WHL^{+}20, SIL^{+}21]. Cohort-based [WHL^{+}20]. Cohorts [ZWW^{+}23b]. CohortVA [ZWW^{+}23b]. CoLi [YLL^{+}22a]. CoLi-BA [YLL^{+}22a]. Collaboration [BK22, GTL^{+}23, iKYOW23, LKL23, SFNRZ^{+}23, SKK^{+}22, TLBB23, XDBAR24, YeSiK^{+}23, YEP^{+}22]. Collaborative [BER^{+}23, CWS^{+}20, CFDN24, DHF^{+}22, EGP^{+}21, FKSP^{+}24, GSK^{+}20, GF24, KMH^{+}23, KWFK20, LHC^{+}21, LIB24, LTLB22, MSA^{+}22, RS23, RVB^{+}22, SSZ^{+}21, SWB^{+}22, SKFZ22]. Collage [STD^{+}23]. Collected [SRKK21]. Collection [CLL^{+}21]. Collections [FRiM^{+}23, KKE21, OM22, PTD^{+}21, STD^{+}23, WQ20, vBMS22, SNBC23, ZWV21]. Collective [YST^{+}20]. Collimated [VM23]. Collisions [BGB^{+}22]. Collocation [PT20]. Collusive [ZWW^{+}23c]. Color [BKS22, DPR^{+}20, DLH^{+}22, EBW23, GSM^{+}22, HYC^{+}23, KHI20, LXF^{+}22, LZZ^{+}20, PPE23, RLLS20, RSRG23, SCL^{+}24, SWS20, YZZ^{+}22, ZWP^{+}22, ZLY22, ZYC^{+}23, bÇ22]. Color-Based [EBW23]. Color-Changing [PPE23]. Color-Coded [BKS22]. Color-Concept [HYC^{+}23, RLLS20]. Color-Encoded [RSRG23]. Color-Matching [GSM^{+}22]. Colored [LSD^{+}23]. Coloring [FFB24]. Colorization [CZG^{+}22, DLH^{+}22, FWZZ20, HYC^{+}23, LFC^{+}21, SZC^{+}23, SHS^{+}22, XWZ^{+}22, XPK^{+}24]. Colormap [RS21, YZF^{+}22, ZZW^{+}22a]. Colormaps [CZC^{+}20, NCS^{+}21, NCB^{+}21, Red23]. Colors [MSY22, NDF^{+}21]. Columnar [CXD^{+}21]. Comantics [XLF^{+}23]. Combination [LRZ^{+}23]. Combined [NVRS^{+}21]. Combining [KKG^{+}20, KBJ^{+}20, LIDM20, MvdEPV24, QSUK22, WSN21, WCX21a, ZUK21]. Comfortable [MP21]. Comic [ZXC^{+}23]. Comic-Style [ZXC^{+}23]. Comics [WRZ^{+}21, WRC^{+}22]. Command [SDK23b]. Commerce [TZT^{+}22, TWW^{+}22, WXG^{+}24, ZWZ^{+}23]. Commercial [WDSM24]. Committee [Ano20m, Ano20j, Ano20k, Ano21j, Ano21g, Ano21f, Ano22n, Ano22m, Ano22o, Ano22p, Ano22s, Ano23l, Ano23k, Ano23p, Ano23q, Ano24j, Ano24o, Ano20f, Ano20l, Ano20o, Ano20q, Ano20p, Ano21i, Ano21l, Ano21o, Ano21t, Ano21q, Ano22r, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-31, Ano23v, Ano23w, Ano23y, Ano23z, Ano23-28, Ano24k, Ano24l, Ano24n, Ano24s, Ano24t, Ano24u, Ano24v, Ano24x]. Committees [Ano20e, Ano21s]. Common [CRP^{+}20]. Commonsense [WHJ^{+}24]. CommonsenseVIS [WHJ^{+}24]. Communicate [CCPM23]. Communicating [JMK^{+}22, PLP^{+}23, PM23]. Communication [ALX^{+}22, AI21, BK22, CQHP22, SLL21, SCL^{+}23a, XVF20, YLS^{+}23, ZYL^{+}22]. Communicative [AL21, LRA23]. Community [Ano23m, LPP^{+}23, PWB21]. Compact [HHI24, RAFSA23b, YLL^{+}22a, RSAA20]. Comparable [WCJW22]. Comparative [ASA^{+}23, ATK^{+}24, BLIC20, BC21, DWOB20, FMP23, FWZM22, HKB^{+}22, HZJ^{+}24, KBB^{+}23, LJS21, LST^{+}22, MMS^{+}23, SLC21, SN23, TARB23, XZKM22, ZTWW21]. Comparing [DUWW22, DWB21, FS24, GLK^{+}23, GTL^{+}23, HMTI24, LFCH24, MTE^{+}20, NMR^{+}23, RRK^{+}22, SPNG23, WSN21, WBPC23, WMMB23]. Comparison [BBL^{+}22, BWZ^{+}20, CAA^{+}21, GSS^{+}23, GJC^{+}22, HMGO23, HKMG22, HSJ^{+}20, JOEF20, JJHS^{+}22, KNM^{+}21, KMM24, LZD^{+}20, LZM20, MFS^{+}23, PAPB20, PTS^{+}20, QBW^{+}20, SCR^{+}23, SYHS20, SHT^{+}22, WDG^{+}20, WWZ^{+}23a, WAV^{+}21, WBWL24, Wu22, SIL^{+}21]. Comparisons [SHOP23, XSB^{+}22, YBR^{+}23]. Compass [DWX^{+}22]. Compensation [HL21, IAI^{+}23, WLH24]. Competing [DWOB20, MGO21]. Compiler [DAK^{+}21]. Complementing [SEAB^{+}22, SDK23b]. Completion [LWF23, ZZD^{+}23a, ZHL^{+}20, ZZW^{+}22b]. Complex [AMY^{+}22, CNC^{+}20, HDFK21, LMD^{+}22, MMF20, WBM21a, KLCK20]. Complexes [FJK^{+}20, SPN23]. Complexities [KTL24]. Complexity [BCN^{+}20, PTM^{+}20, RRK^{+}22, eSYKW23]. COMPO*SED [CMSK23]. Component [GSS^{+}20, YGT^{+}23, ZHN24]. Components [RLG^{+}23, VVC^{+}24]. Composed [ISKM23]. Composite [CMSK23, DCM^{+}23]. Compositing [MRS22]. Composition [CZL^{+}21, PSS23, WYZ^{+}21, Wu22]. Compound [BD22, SUS^{+}21]. Comprehensive [FLS23]. Compressed [MSK23a, PD24, SKNŽ20]. Compressing [ATAS21, OM22, WDN23]. Compression [BRLP20, HRX23, KBPR22, LDC^{+}23, LLDW24, ML24, XWZ^{+}22, YLGW24]. Computation [AT23, CWS^{+}21, GVT24, LWWH20, LMGY22, LWW^{+}24, MLT^{+}24]. Computational [BFY^{+}24, CCW^{+}21, FLW^{+}21, ILZ^{+}21, KBM21, LLY^{+}24, PLD^{+}23, SOL^{+}22]. Compute [CWR21, HRX23]. Computer [Ano20a, Ano20b, Ano20c, Ano20d, Ano21b, Ano24h, CGT^{+}24, DMBK21, WRZ^{+}21]. Computer-Aided [DMBK21]. Computing [BTL23, ECR22, LMM^{+}21, MXT^{+}22, OKM21, RWJ^{+}23, SPN23, WHL^{+}20, XWY^{+}20]. Concentric [ZWP21]. Concept [EAKC^{+}20, HHS^{+}23, HYC^{+}23, HMKB23, RLLS20, WTL24, YZEN22]. Concept-Driven [WTL24]. ConceptExplainer [HMKB23]. Concepts [PDD^{+}22, ZXSR22]. Conceptual [LCC^{+}20, MSA^{+}22, WWP22, leáB^{+}21]. Conceptualize [HGB22]. Concerns [WBI20]. Concert [PCL24]. Concise [CWL23, QFWS22]. Condition [SBe^{+}21]. Conditioned [LYZ^{+}21, SYC^{+}23]. Conditioning [CIA24]. Conditions [KKS^{+}22, SA22, ZCR21]. Conducting [LELT23]. Conductivity [PSG^{+}22]. ConeSpeech [YLS^{+}23]. Conference [Ano20q, Ano21t, Ano22-27, Ano23f, Ano23w, Ano24t, CLL^{+}21, Ano20f]. Confidence [IRR^{+}22, KKLS21, PVPM22]. Confidence-Controlled [KKLS21]. Configuration [CZL^{+}21]. Conflict [KKW23]. Conflicting [MSMX23]. Confluent [ZPG21]. Conformal [NGBA^{+}20, ZWW22]. Conforming [WWG21]. Confusion [HRS^{+}22]. ConfusionFlow [HRS^{+}22]. Congestion [LKJ^{+}20, PYSJ21]. Congruence [AR22, KL22, MWD^{+}23]. congruent [WMB23]. Connecting [CLD^{+}22]. Connection [CAA^{+}21]. Connectivity [TWC^{+}24]. Consensus [FGS^{+}22]. Consent [ZC23]. Conserving [KLTB21]. Considerations [SNK^{+}22]. Considering [YPW23]. Consistency [DHM^{+}22]. Consistent [HLJ^{+}22, HCH^{+}23, SFL^{+}22, ZZD^{+}23b]. Constancy [JdJTC24]. Constrained [HYSL23, KKEW23, KMG^{+}21, LWWF21, LF23, LCSA22, MvdEPV24, WMMB23, XWH^{+}23]. Constraint [BD22, CWS^{+}20, ZNAN20]. Constraints [HLJ^{+}22, HZQ22, WWW^{+}24, ZLZ21, ZWW^{+}23d]. Construct [BZP^{+}20].

Construct-A-Vis [BZP^{+}20]. Constructed [LCYQ24]. Constructing [AAA^{+}21]. Construction [CWW^{+}21, KKV22, LHS^{+}22, MM20, RL20, SWS20, SMCL24, WCW^{+}22, WBL^{+}22, WG21, XLY^{+}22, ZSL21]. Constructive [MEHD24]. Consumer [STA^{+}21, ZX22]. Consumption [LCW^{+}23b]. Contact [ASCR^{+}22, dPLM21]. Contemplate [CPCS20]. Content [CSWZ24, KLSB22, LS22, OKM21, PTD^{+}21, XQXL23, ZLZ21, ZWG^{+}23]. Content-Aware [KLSB22]. Content-Based [OKM21, PTD^{+}21]. Content-Preserving [ZLZ21]. Contents [Ano20g, Ano20h, Ano20i, Ano21c, Ano21d, Ano22t, Ano22u, Ano22v, Ano22w, Ano23s, Ano23t, Ano23u, Ano23x, Ano24q, Ano24r]. Context [APP^{+}22, AYGR22, GZG^{+}24, HLJ^{+}22, HCL20, JZHA22, KKE21, MGWK^{+}22, MNK23, MYS^{+}22, XHL^{+}24, ZIX^{+}24, ZSS^{+}21, HTJ^{+}20, HBS^{+}21, JKW^{+}22, NWW21, NSW24]. Context- [KKE21]. Context-Aware [AYGR22, GZG^{+}24, ZIX^{+}24, ZSS^{+}21]. Context-Consistent [HLJ^{+}22]. Context-dependent [MNK23]. Contextual [LLZ^{+}23, YHC^{+}24, ZFCG23, ZCM23a]. Contextualization [LMF^{+}24]. Contextualized [PSY^{+}20]. Contextualizing [EJS^{+}23]. Contiguous [DTPG21]. Contingent [CZT^{+}21, CDS^{+}22]. Continuous [AZA^{+}23, EGMP24, LLDW24, NCS^{+}21, NCB^{+}21, SMNK21, ZS21]. Contour [CWS^{+}21, CRWA22, HZQ22, KKV22, LYL^{+}23, ZLC^{+}23]. Contraction [XHFZ24]. Contrails [NFN^{+}23]. Contrast [ZWP^{+}22]. Contrastive [FKM20, XHL^{+}23]. Control [AYA^{+}21, BWCT23, FALH20, FOH^{+}21, HTBL22, LPL^{+}24, LWT^{+}23, MHIS23, NLPW22, WCX21a, YZF^{+}23, VAWL24]. Controllable [LH23, LSE20, ZLZ^{+}23a, ZZS^{+}23].

Controlled [KKLS21, LCM^{+}23, WRZ^{+}21, WZL22]. Controller [DSD^{+}23, LJCL24, SPNG23, WBM21a]. Controllers [GLK^{+}23, LCM^{+}23, WHBC23]. Conundrum [GSS^{+}23]. Convection [PSH21]. Convention [WWP22]. Convergence [CK22, ESP20]. Conversational [GGP^{+}24]. Conversations [LLSM24]. Convertible [XFF^{+}21]. Convex [AKS22]. Convey [KHR21]. Convolution [LS23, RZW^{+}24]. Convolutional [DLH^{+}22, KIS22, LWS^{+}21, LCK^{+}21, LZL^{+}23, PSL23, TNJ^{+}22, WTS^{+}21, WCZ^{+}23, XZXY22, XZKM22, ZJJH21]. Convolutions [AFB22]. Cooks [ZYM^{+}24]. Cooperative [SCST24, SCC^{+}23]. Coordinate [HW23, HCX^{+}21, TEK^{+}23]. Coordinate-Based [HW23]. Coordinates [BKS22, CMSK23, ZS21]. Coordinating [CSC^{+}22]. CoordNet [HW23]. Copyright [Ano21e]. Core [LXL21, MZS^{+}24, RZW^{+}24, SCRL20, WLC^{+}23, ZSL21]. Cores [MWUP22]. CorGIE [LWBM22]. Coronary [LCH^{+}21]. Corr [CLZ^{+}24]. Corr-Track [CLZ^{+}24]. Corrected [BZKG21]. Correcting [SZH^{+}20]. Correlation [EBJ^{+}22, HT22, OWN^{+}21, PAPB20, XSKF23, YST^{+}20, KMWD21]. Correlations [NSW24]. Correspondence [CLZ^{+}24, LWWF21, NIK24]. Corresponding [LWBM22]. Corruption [LMM^{+}21]. Cosmological [ABE^{+}22, EBPF21]. CosmoVis [ABE^{+}22]. Cost [BLE^{+}23, GBL^{+}22, ZZS^{+}23, CXXZ21]. Couette [NMC21]. Coulomb [DS20]. Counterfactual [CMQ21, HSYZ24]. Coupled [BXQ^{+}22]. Coupling [DS20, HWN^{+}23, LSL^{+}23]. Courses [CYP^{+}20]. CourtTime [PJHY20]. CoUX [SKFZ22]. Cover [Ano22k, Ano23h, Ano24i, HSV^{+}20, WTY^{+}22, Ano22h, Ano22i, Ano22j]. COVID [JDZK22, NSK^{+}21, PFCB23, ZSG^{+}23]. COVID-19 [JDZK22, NSK^{+}21, PFCB23, ZSG^{+}23]. COVID-view [JDZK22]. Crafting [SWS20, WAV^{+}21]. Creases [ZH20]. Creating [BJR21, TLW^{+}21]. Creation [ACL^{+}24, FWW^{+}24, TG24, YKF22]. Creative [GWF^{+}23, LXF^{+}22]. Creators [ZSG^{+}23]. CreatureShop [ZYC^{+}23]. Credible [GMVRB20]. Creek [PWB21]. CRF [DHM^{+}22]. CRF-Based [DHM^{+}22]. CrimAnalyzer [GSP^{+}21]. Crime [GZRP^{+}22, GSP^{+}21]. Criminal [ZeB^{+}21]. CriPAV [GZRP^{+}22]. Crisis [ZSG^{+}23]. Criteria [BDL^{+}21, MD20]. Criterion [CAR^{+}23]. Critical [DS22, KKV22, OCW^{+}24, SLR^{+}20a]. Critics [WDC^{+}23]. Cross [GGP^{+}24, GGT^{+}20, PBBH20, RDH23, SPZS24, SSAZ22, SNK^{+}22, TLW^{+}23, WCH^{+}23]. Cross-Domain [WCH^{+}23]. Cross-Language [RDH23]. Cross-National [TLW^{+}23]. Cross-Reality [GGP^{+}24]. Cross-Renderer [GGT^{+}20]. Cross-Scale [PBBH20]. Cross-View [SSAZ22, SNK^{+}22]. Cross-Virtuality [SPZS24]. CrowbarLimbs [BTHL23]. Crowd [BGB^{+}22, HHB^{+}23, JWE^{+}22, PNB^{+}21, YHC^{+}22, YHC^{+}24]. Crowded [AeSL^{+}23]. Crowds [YHC^{+}24]. Crowdsourced [SCHE23, YXW^{+}23]. Cryo [NBE^{+}23]. Cryo-Electron [NBE^{+}23]. CT [JDZK22]. Cube [FSN20, LCC^{+}20]. Cubemap [LXF^{+}22]. Cue [YTHL23]. Cues [DWOB20, EMM^{+}22, EML^{+}23, EBW23, FTWP22, JKL24, MCQ23, MWR^{+}22, RMW^{+}24, TME^{+}22, VBV^{+}23, WPNK21, ZSCRB23]. Culling [CNAA^{+}22, IRR^{+}22]. Cultivating [BVY^{+}23]. Curated [PSS^{+}22]. Curating [ZJX^{+}24]. Curation [Ano22y]. Curiosity [BVY^{+}23]. Current [SHM23]. Curse [XVF20]. Curvature [SMNK21]. Curvatures [UDH23]. Curve [CWL23, LWC22, QHL^{+}20, SLC21, XCK20, YKF22, YWM^{+}23]. Curved [KLB24, RMB^{+}21]. Curves [HZYZ22, LZH^{+}21, LWH^{+}22, NCE23, ZJW21]. Cyberattack [RCD^{+}23]. Cybersecurity [leáB^{+}21]. Cybersickness [JLM^{+}21, KAL^{+}23, LGL^{+}23a, TIDQ24, VVR^{+}24, TB24]. Cycles [Iur22]. Cyclical [CDBM22]. cycling [MCSAL23]. Cyclone [HXL^{+}24]. Cyclones [YGP^{+}24]. Cylinder [LMD^{+}22]. Cylindrical [CDBM22]. Cytometry [SVK^{+}21].

D [BBMM^{+}23, CLL^{+}20, CZY^{+}20, FHR^{+}21, FYL^{+}23, GZM^{+}21, GM23, HZZ^{+}20, HB20, HCL20, JDM^{+}22, JSA^{+}20, KCA^{+}21, KNM^{+}21, KRK21, KKF20, KWFK20, LBB^{+}20, LWC22, LZX^{+}21, LCK^{+}21, LWWY21, MDJV21, MEB^{+}20, NDF^{+}21, PALW20, PL21, RMB^{+}21, SYW^{+}20, SZF^{+}21, SSX^{+}20, SLR20b, STA^{+}21, SKNŽ20, SS21, TGM21, TIHS20, WZH20, WL20, WCX21b, XJZ^{+}21, XLL^{+}20, XF21, XYF^{+}21, YRL^{+}20, YZJ^{+}20, YCB^{+}21, YGE^{+}21, ZHDX20, ZHL^{+}21a, ZHL^{+}20, ZLG^{+}21a, ZH20, iIK^{+}21]. D-BIAS [GM23]. D-Kernel [MDJV21]. D-SAV360 [BBMM^{+}23]. D3 [HA20, HPM^{+}24]. Daily [WDG^{+}20]. Dance [LZZ^{+}24]. Dancer [AYA^{+}23]. Daniel [Ano22e]. Dash [bÇ22]. Dashboard [BFAR^{+}23, DWQW23, MMG^{+}21, PSS23, SCST24, ZSG^{+}23]. Dashboarding [WDSM24]. Dashboards [ERB^{+}21, KNAR^{+}22, WWZ^{+}22, ZCM22]. DashBot [DWQW23]. Data [ALX^{+}22, ATAS21, AA21, AJSP23, BKR^{+}24, BLBL23, BRLP20, BCT22, BS21, BWI21, BDL^{+}21, BER^{+}23, BWZ^{+}20, BBSvL24, BK22, CBL^{+}24, COFJ23, CWS^{+}21, CRWA22, CXD^{+}21, CMKK21, CAA^{+}21, CZF^{+}22, CDBM22, CSX^{+}22, CWH^{+}23, CCS^{+}24, CLD^{+}22, CKQ^{+}23, CRJ^{+}24, CNY22, CBP22, CFGT21, CFGM22, DPC23, DPR^{+}20, DMMF21, DKM24, DWL^{+}22, DP20, DZTF22, DTS^{+}21, ENM24, ECR22, ESB^{+}24, EGMP24, EKC^{+}23, EAS^{+}23, FLZ^{+}21, FSN20, FKSP^{+}24, FZC^{+}21, FCS^{+}20, FSS^{+}21, GLY^{+}21, GMS^{+}21, GZW^{+}20, GG21, GVT24, GGJ^{+}22, GTC^{+}23, HMGO23, HW20, HZX^{+}21, HW22, HW23, HTY^{+}23, HWT^{+}24, HM24, HSB^{+}21, HSV^{+}20, HSC^{+}22, HXHT23, IRR^{+}22, ISBP22, JGG21, JLK20, JZHA22, JH20, JKW^{+}22, JCZ^{+}24, JGC^{+}21, KKG^{+}20, KHL21, KBM21, KKGMH21, KGBP20, KVGM20, KLKE21, KW23, KBJ^{+}20, LG23, LSW^{+}22, LWC24, LPJT^{+}22, LZK^{+}22, LME^{+}23, LBL^{+}21, LHC^{+}21, LIB24, LM20, LQWQ21]. Data [LMM^{+}21, LBW^{+}22, LLW^{+}23, LqZ23, LS23, LAML23, LZZ^{+}20, LWSY20, LWM^{+}20, LBE20, LZX^{+}21, LRZ^{+}23, LIFD23, LI24, LAS^{+}20, LFC^{+}21, LFW^{+}22, LGMT21, ML24, MM20, MSMX23, MCW^{+}20, MvdEPV24, MNB^{+}23, MJAD22, MEHD24, MM23, MGU^{+}21, NIK24, NSS21, NQE21, NCWE22, NP24a, NCE23, NdCS21, NF20, Off20, OWN^{+}21, PLW^{+}23, PNB^{+}21, PFN22, Par22, PPE23, PBM^{+}24, QBW^{+}20, RBK^{+}21, RPD20, RPD21, RPD22, RXX^{+}21, RGDG23, SGH^{+}23, SCRL20, SHOP23, SBk^{+}24, SRKK21, SESH24, SLX^{+}23, SSL^{+}23, SZZW24, SS24, SXS^{+}21, SHT^{+}22, SGJC23, SCC^{+}24, SFS^{+}22, SCHE23, SWT^{+}21, SSJ^{+}22, SVK^{+}21, SDXR22, SB23, SaT^{+}23, SLS21, SKK^{+}22, SSAZ22, SNK^{+}22, SCC^{+}23, SS21, TAL24, TNJ^{+}22, TCX^{+}23, THFI21, Tsa21, TLW^{+}23, VKT^{+}24, WSN21, WHX^{+}23, WZU^{+}21, WMZ22, WFW^{+}20, WSZ^{+}20, WWSS20, WRZ^{+}21, WYZ^{+}21, WRC^{+}22, WLS^{+}22, WCWQ22, WCX^{+}23, WKN^{+}23, WN21, WBI20, WIP^{+}24, WWS^{+}22, WGM^{+}24]. Data [WGH^{+}24, WGS^{+}23, WXS^{+}24, XCK20, XDBAR24, XVF20, XSHF20, XLF^{+}23, XFD^{+}23, XGS^{+}21, YBR^{+}23, YCB^{+}21, YXL^{+}22, YLLW24, YDMP22, YS20, ZSL21, ZZW^{+}22a, ZSM21, ZCL^{+}21, ZWW^{+}23b, ZCM23a, ZOS^{+}23, ZYP^{+}24, ZWZ^{+}22, ZJW21, ZZW^{+}22c, DJHBJ21, MHFF21, PWB21, SIL^{+}21, ENM24, THS^{+}21, WTL24]. Data-Agnostic [Tsa21]. Data-Aware [PFN22]. Data-Backed [SKK^{+}22]. Data-Driven [BDL^{+}21, CBP22, LZZ^{+}20, LWM^{+}20, LZX^{+}21, LFW^{+}22, MM23, OWN^{+}21, RBK^{+}21, RXX^{+}21, TAL24, ZZW^{+}22a, ZJW21]. Data-GIF [SWT^{+}21]. Database [BLE^{+}23]. Dataflow [UWF^{+}23, YS20]. Dataless [RDHH21]. Dataopsy [HE24]. Dataset [BBMM^{+}23, DWS^{+}23, PSS^{+}22, PCJ23, WGS^{+}24, ZJX^{+}24, ZYL^{+}24, LFG^{+}23]. Datasets [BHA^{+}23, EBPF21, GRi^{+}21, LWSY20, MCW^{+}20, PFC20, SPJ^{+}23, WCJW22, DJBJ23]. DataShot [WSZ^{+}20]. Dating [ZC23]. DCNN [WCX21b]. DCNN-Based [WCX21b]. DDLVis [LBW^{+}22]. Dead [EGMP24]. Deadeye [KCWK20]. Deaf [MKK20]. Deblurring [KIS22]. DECE [CMQ21]. Decentering [RBRG21]. Decision [AL20, CMQ21, DAB^{+}23, DS22, DZTF22, DIPJ22, HM22, LLP^{+}23, MSBV^{+}23, OCW^{+}24, PRJ^{+}23, SWF^{+}24, SMS^{+}22, ZLVV22]. Decision-Making [LLP^{+}23, SWF^{+}24]. Decisions [ABGG21, KKH21, PGS21, TAL24, WKMD22]. Declarative [CWH^{+}23, LM21]. Declutter [ALX^{+}22]. Decoder [ZLQH21, ZLW21a]. Decoding [GFCM23, MMF20, SDK23a]. Decomposed [SCL^{+}23b]. Decomposing [SII^{+}21, WMW^{+}22]. Decomposition [KBV24, KLKE21, KKEW23, LHL^{+}20, NWMC23, Pat22, SLJ^{+}20]. Decompositions [BKPB21]. Deconstructing [CLW^{+}24, LST^{+}22]. Decorating [LFR^{+}21]. DecorIn [LFR^{+}21]. Deep [BWM23, BMWD20, CLS^{+}21, CZY^{+}20, CWW^{+}20, CJY^{+}23, CFZ^{+}23, CCL^{+}24, DWQW23, DLH^{+}22, DCWD23, ER21, GSK^{+}20, GLA^{+}24a, HTW20, HZX^{+}21, HWG^{+}20, HPRC20, HSC^{+}22, HMKB23, JZCZ20, KM20, KKhCM23, LLZ^{+}21, LZLS22, LLC^{+}22, LWY^{+}22, LQS^{+}23, MTW^{+}20, MMG^{+}21, MXC^{+}20, MRS22, PPYW20, PNKC21, PDD^{+}22, PBAG23, PvSvdE^{+}24, SZF^{+}21, SC22, SZC^{+}23, TBL^{+}20, TNJ^{+}22, TXM23, WJW^{+}20, WZH20, WYC^{+}20, WYZ^{+}21, WZY^{+}22, WH23, WCTW21, WWZ^{+}22, WGS^{+}23, XWZ^{+}22, XZKM22, YZF^{+}22, ZCX^{+}24, ZLL^{+}21, ZLY22, ZXSR22, ZZW^{+}22c]. DeepDrawing [WJW^{+}20]. Deeper [LBL^{+}21]. Deeply [SCHE23]. DeepOrganNet [WZH20]. DeepSketchHair [SZF^{+}21]. Defence [WDC^{+}23]. Defense [WBL^{+}22]. Defocusing [UIHS20]. Deformable [BPL23, IGM24, SOL^{+}22]. Deformation [CYD^{+}23, GLY^{+}21, LSW^{+}20, MDL^{+}23, QLFG22, RHLC22, SKA21, WZH20, YGT^{+}23]. Deformation-Driven [SKA21]. Deformations [HS23, WCC^{+}21, ZH20]. deforming [PGL^{+}20]. Degree [AMAS21, HRX23, NBE^{+}21, VH23, LDB^{+}21, LCW^{+}23b, RSAA20]. DeHumor [WMW^{+}22]. Delay [KZD^{+}23]. Delimiter [SDK23b]. Delimiter-Free [SDK23b]. Delivery [FWM^{+}24, LFCH24]. Deltas [NF20]. Demand [SCRL20]. Demands [VVR^{+}24]. Dementia [PMN^{+}23]. Demers [NSM^{+}22]. Demonstration [SMSK22, ZBNS21]. Dendritic [CLL^{+}23]. Dendrograms [LZM20]. DendroMap [BHA^{+}23]. Denoising [AFB22, CWS^{+}20, IGMM22, LLZ^{+}21, LLW^{+}22a, XF21, ZWXM23]. Dense [GBNH21, LHL^{+}22, MM23, PWK21, YLL^{+}20, YFS22, iIK^{+}21]. Density [CWB^{+}20, FLZ^{+}21, HGSP20, HLCY22, KLTB21, LBW^{+}22, QR21, XPK^{+}24]. Density-Based [HLCY22]. Dental [LZY^{+}23, SEK^{+}24]. Dependent [CMSK23, ENvBC23, LWK20, MMS^{+}23, MBHE24, RG20, SXL^{+}23, WM23, MNK23]. Depending [JKL24]. Deployment [ZYL^{+}24]. DeProCams [HL21]. Depth [AKS22, DGKOC20, ENK^{+}20, FRL^{+}23, KBPR22, KHR21, MGWK^{+}22, MWR^{+}22, MF24, MSK23a, SES20, WBWL24, XCG^{+}21, YFS22, ZSLL20]. Depth-Compressed [MSK23a]. Dermatomyositis [ZCM^{+}23b]. Descent [ADD^{+}22]. Description [RAFSA23b, YXW^{+}23]. Descriptions [LS22]. Descriptor [LFO23]. Design [ALX^{+}22, AR22, BFAR^{+}23, BFY^{+}24, BDRW21, BHU^{+}21, BKH22, CWW^{+}20, CFGM22, DKM24, DCM^{+}23, DLP^{+}23, DIPJ22, EBJ^{+}22, ENXS21, FLW^{+}21, GSS^{+}23, GKC^{+}24, GKN^{+}23, HBH^{+}20, HZII24, HGB22, HB24, HLH^{+}23, KJS^{+}23, KGB22, LJS21, LG23, LSZC21, LSW^{+}22, LWC24, LRHA22, LSS24, LYZ^{+}21, LCH^{+}21, LQWQ21, LY23, LXF^{+}22, LCYQ24, MD20, MJAD22, PLL^{+}24, PSS^{+}22, Par22, QCL23, QNWR23, RS21, RR23, RPH^{+}21, SDC^{+}24, SCST24, SLC^{+}23, SMPJ^{+}20, SOL^{+}22, SNK^{+}22, TSH21, Tsa21, WSL^{+}20, WFW^{+}20, WZC^{+}23, WHC^{+}23, WYIS24, WBI20, WH22a, WWG21, WHBC23, XVW^{+}21, XDBAR24, XFF^{+}21, YXL^{+}22, YSM^{+}20, YLL^{+}22b, YLF^{+}20, YLTL23, ZSG^{+}23, ZBG^{+}24, ZZZ24, ZWG^{+}23, ZC23]. Designer [SWS20]. Designers [BLBL23]. Designing [BXF^{+}22, HNGC21, LVR^{+}24, LDT^{+}21, NB24, PSY^{+}20, WYIS24, WWS20, WWZ^{+}22, YVBI24, ZBNS21, PWB21]. Designs [BBSvL24, HPP^{+}20, KNAR^{+}22, KJS^{+}23, LCYQ24]. Detail [FZZX22, LQWQ21, LSW^{+}23, SIA^{+}23, SKNŽ20, YCB^{+}21, ZZD^{+}23a]. Detailed [BWZ^{+}20]. Details [LWL^{+}23a]. Detecting [HMGO23, JRM22, ROM^{+}23, VCO^{+}23]. Detection [BRLR24, CWW^{+}22, CDZ^{+}23, CGT^{+}24, CWR21, GZM^{+}21, GZL^{+}21, GJC^{+}22, JGH^{+}24, LCWL23, LZX^{+}22, LMD^{+}22, MLC^{+}20, NFN^{+}23, NdCS21, NNF^{+}20, ONH21, QGL^{+}23, WLZ^{+}23, XCK20, XDH^{+}22, YDMP22]. Deterministic [LCSA22]. Development [KKJ^{+}21, MSBV^{+}23, RBF^{+}23, SKW^{+}23]. Device [XJZ^{+}21, ZJS^{+}22]. Devices [NIK24, TGM21, BNWvW21]. Dexterous [HSC^{+}22]. DGaze [HLZ^{+}20]. Diabetes [SESH24]. Diachronic [XTYL20]. Diagnosing [HSYZ24, HZS^{+}22, MFH^{+}21, YYZ^{+}22]. Diagnosis [DMBK21, JDZK22, JWW^{+}23]. Diagnostic [OWW^{+}24]. Diagnostics [DWOB20]. Diagram [SCLK21]. Diagrams [CPD20, CCS^{+}24, GVT24, KGWD22, LMGY22, PCZ^{+}21, PVT23, RSRG23, RAP21, SDT24, VBT20, WDN23]. Diatoms [BKH22]. Dichoptic [EBW23]. Dichotomous [HHC^{+}21]. Dicing [GTC^{+}23]. Dictionaries [SDT24]. Dictionary [CDX^{+}20, LBW^{+}22]. Dictionary-based [CDX^{+}20]. Did [BVV^{+}22]. Diegetic [NBE^{+}21]. DIEL [WCH^{+}22]. Differences [CQHP22, DPD^{+}24, HKB^{+}22, LLW^{+}22b, XCG^{+}24]. Different [BS22, CSL^{+}21, EJS^{+}23, GLK^{+}23, GMVRB20, KKS^{+}22, UDH23]. Differentiable [FWW^{+}23, LWL^{+}23b, PLL^{+}24, SS21, WW22b, ZWL23]. Differential [JKL24, ZZW21, ZWW^{+}23d]. Differentially [ZSM21]. Differentials [ZWW22]. Difficulty [DKuH20]. Diffuse [NDF^{+}21, XZXY22]. Diffusion [LZH^{+}21, LWH^{+}22]. Digital [CLG21, OWN^{+}21, PLP^{+}23, SKFZ22]. Dilemmas [SWF^{+}24]. Dimension [EMK^{+}21, ZCX^{+}24]. Dimensional [BWZ^{+}20, EHA^{+}23, EAS^{+}23, GMS^{+}21, GZW^{+}20, KKG^{+}20, MGU^{+}21, SSJ^{+}22, VKT^{+}24, WCJW22, WN21, XZF^{+}22, ZZW^{+}22c, ZZC^{+}22, vOVR23, DTS^{+}21]. Dimensionality [ACT^{+}24, FCS^{+}20, FKM20, FSS^{+}21, FWZM22, JKA^{+}24, KKG^{+}20, MBC^{+}23, XZS^{+}22, XHL^{+}23, XHS^{+}24, ZCH^{+}21]. Diminishes [KBPR22]. DimLift [GMS^{+}21]. Direct [AMS^{+}21, EML^{+}23, GRi^{+}21, IGMM22, JKU^{+}22, KBV24, NF20, RRK^{+}22, SHPE20, SCL^{+}23a, WW22b]. Direct-View [EML^{+}23]. Direct-View/Near-Eye [EML^{+}23]. Directed [SHW^{+}20]. Direction [NMS^{+}23, SNH^{+}23]. Directional [GGPL20, YLS^{+}23]. Directionally [SII^{+}21]. Directions [KPF^{+}22, LJS21, STA^{+}21, TBW^{+}23]. Disabilities [CFQ21]. Disability [CFQ21, CQ22]. Disagreement [WWZ^{+}23a]. Disciplines [HKB^{+}22, LZJZ20]. Discomfort [CWSJ23]. Discontinuous [MDS24]. Discourse [BDRW21, SEAB^{+}22]. Discovering [CPCS20]. Discovery [NKWW22]. Discrete [CPD20, GVT24, LLCH22, MXT^{+}22, PWK21, ZNAN20]. Discriminability [MYS^{+}22, SLL21, VC20]. Discriminable [LFC^{+}21]. Discrimination [DGDC21, WXC^{+}21]. Discussion [BDRW21]. Disease [KAS^{+}21, WMH^{+}22, XNF^{+}23]. Disentangled [JZCZ20, LWL^{+}23a, LH23, WCH^{+}23]. Disentanglement [CSWZ24]. Disks [SCLK21]. Disocclusion [LBP24]. Disparity [HX23]. Dispersion [HX23]. Displacement [PL21, SKR^{+}24]. Display [BJCL21, CXXZ21, CCS^{+}21, FKSP^{+}24, GSM^{+}22, HSJ^{+}20, HWMI23, HHI24, IAI^{+}23, KHI20, KCA^{+}21, Kel23, KPW20, MSK23a, MHIS23, NMS^{+}23, NIM^{+}21, QSUK22, SA22, San20, WO22, WH22a, YLJ^{+}22, YB20, ZZ23, ZZS^{+}23, ZJS^{+}22, MSAM^{+}22]. displayed [WBA^{+}23]. Displays [ATHI24, CZT^{+}21, EML^{+}23, FGF^{+}21, FCFC22a, FCFC22b, FBB^{+}21, IGM24, IKi21, KTL24, LAN21, LSG24, LIB24, MTE^{+}20, MGWK^{+}22, MMN^{+}22, MML21, PGE^{+}22, RFD21, SPWW^{+}24, SPNG23, SBK^{+}20, TGM21, TME^{+}22, TEK^{+}23, UDH23, WSL^{+}24, WBA^{+}23, ZWP^{+}22, ZHKY23, ZSCRB23, RSAA20, TWA22]. Dissertation [Ano22a, Ano23b, Ano24b, SW21, SAB20].

Dissimilar [CFL^{+}23, YYK^{+}22].

Dissimilarities [KLCK20]. Distance [APP^{+}22, BTL23, Kel23, LVV^{+}21, LBP24, SES20, SMKN20, SN23, WBWL24, XZF^{+}22]. Distances [PVDT22, YBR^{+}23]. Distant [ALC22, LTLB22]. Distinction [SLR20b]. Distortion [CFL21, CWSJ23, DST^{+}23, SPWW^{+}24]. Distributed [FGS^{+}22, HTY^{+}23, HCJ21, PFN22, SSZ^{+}21, SYY^{+}24, WBF20, WF21, XGS^{+}21, XWPG^{+}24]. Distributed-Memory [PFN22]. Distribution [CYL^{+}21, LBW^{+}22, MCW^{+}20, RSD^{+}23, SFL^{+}22, YB20]. Distribution-Aware [MCW^{+}20]. Distributions [Kay24, KSHW22, NCE23, WWSS20, ZLG^{+}21a]. Diverse [ACL^{+}24, BHU^{+}21, LQS^{+}23, NMS^{+}23, WXW^{+}20, ZCR21]. DIVI [SH24a]. DL4GD [GLA^{+}24a]. DL4SciVis [WH23]. DMRI [CZC^{+}20]. DMT [LZZ^{+}24, ZCX^{+}24]. DMT-EV [ZCX^{+}24]. DNA [KSB^{+}22]. DNF [LLZ^{+}21]. DNF-Net [LLZ^{+}21]. Do [DHLA22, DS22, GKC^{+}24, ISBP22, MSMX23, MGM^{+}22]. DocFlow [QTW^{+}24]. Docking [MCFKF24, SBk^{+}24]. Document [HTJ^{+}20, KKE21, QTW^{+}24]. Documentaries [ieSM^{+}23]. Documenting [LLY^{+}24]. Documents [BDRW21, LZK^{+}22, MLBW20, RPNP23]. Does [LWD^{+}23, WLZ^{+}23]. DoF [CZF^{+}23]. DOI [FCFC22a]. Domain [BCC^{+}20, CFGM22, KGX^{+}23, QFWS22, TARB23, WCH^{+}23, WYIS24, XDBAR24, YSM^{+}20]. Domain-Oblivious [QFWS22]. Domain-Specific [CFGM22]. Domains [YWM^{+}23]. Dominance [CCM20, RBK^{+}21, MDV20]. Dominant [XZS^{+}23]. DOMINO [WM23]. Don’t [Hul20]. Door [BVV^{+}22, HFT^{+}24]. Door-Opening [HFT^{+}24]. Dot [RPHJ20, RSD^{+}23]. Dot-Label [RPHJ20]. Dots [CLD^{+}22]. Doubt [BBH24]. Down [DXX^{+}21]. Downside [FWM^{+}24]. Downward [NIM^{+}21]. DPVis [KAS^{+}21]. DPVisCreator [ZWW^{+}23d]. Dr. [KBV24]. Draco [SPM24, ZYM^{+}24]. Drawing [ADD^{+}22, GLA^{+}24a, WJW^{+}20, ZYC^{+}23]. DrawingInStyles [SYC^{+}23]. Drawings [CZG^{+}22, ZPG21]. Dream [LZP^{+}20]. Dream-Experiment [LZP^{+}20]. Dream360 [ACL^{+}24]. DRGraph [ZCH^{+}21]. Drift [BBM^{+}21, SLJ^{+}20, YDMP22]. Drilldown [FJK^{+}20]. Drive [WYS^{+}22]. Driven [AYA^{+}23, BDL^{+}21, CSWZ24, CFZ^{+}23, CBP22, CCW^{+}21, DWQW23, DGKOC20, FHTB23, FYE^{+}22, GLY^{+}21, GNMQ22, GZG^{+}24, HBH^{+}20, HCX^{+}21, LTC21, LBB^{+}20, LZZ^{+}20, LWM^{+}20, LZX^{+}21, LHL^{+}22, LFW^{+}22, MM23, OWN^{+}21, PNKC21, QHL^{+}20, RBK^{+}21, RXX^{+}21, SKA21, SUB^{+}22, SLX^{+}23, SHS^{+}22, TAL24, WTL24, XDH^{+}22, ZZW^{+}22a, ZOS^{+}23, ZJW21, ZJJH21, WWR^{+}20]. Driving [HWW^{+}22, JZHA22, JLM^{+}21, WLZ^{+}23]. Drug [WHC^{+}23]. Dry [ZHL^{+}21b]. DSLs [McN23]. DTexFusion [ZX22]. DTI [XNF^{+}23]. Dual [ATHI24, CML24, CCEA^{+}24, DHLA22, DLH^{+}22, LSL^{+}23, MGU^{+}21, PCQ^{+}20, QZZ22, SJL^{+}23, WQQ^{+}24, WCZ^{+}23, ZZ23, vdRBE22]. Dual-Hierarchy [vdRBE22]. Dual-Lens [DLH^{+}22]. Dual-Modal [CML24]. Dual-mode [ZZ23]. Dual-Task [PCQ^{+}20]. Duplicated [YEP^{+}22]. Dupo [KRHH24]. During [BFS^{+}24, CWSJ23, DHLA22, DKuH20, LKL23, MCS^{+}23, PGS21, SACB^{+}23, WMZ^{+}20, YB20, ZOF^{+}23, ZYR^{+}20, ZCL23, BBM^{+}21, NCWE22, XCG^{+}24]. DXplorer [CLL^{+}23]. Dynamic [AYA^{+}21, BWM24, BZKG21, CSJ^{+}21, CWB^{+}20, CCL^{+}24, CSCM21, DHM^{+}22, FHR^{+}21, FTWP22, GMX23, GZW^{+}20, Hei21, HFT^{+}24, HLZ^{+}20, JBS^{+}22, JRM22, KJL24, KKEW23, LAN21, LJCL24, LXH^{+}21, LGW^{+}20, MSWI22, MBS23, NDP^{+}21, NWMC23, NSM^{+}22, NLPW22, SAK20, SMC^{+}21, SCLK21, SCL^{+}23b, TIHS20, VBP^{+}21, VVR^{+}23b, WPZ^{+}23, WBM21b, XTYL20, XCG^{+}21, ZUK21, ZCR21, ZOF^{+}23, ZX22, ZLX23, ZRJW20, ZP23, VAWL24]. Dynamical [NDLW20]. Dynamically [IWT^{+}20, MSK23a, SPNG23, SH24a, YZP^{+}23]. Dynamics [COZ^{+}23, HZP^{+}24, KVGM20, LLK22, NDP^{+}21, PBBH20, UWF^{+}23, WPL^{+}22, XWL^{+}21, ZNF^{+}23].

E-Commerce

[TZT^{+}22, TWW^{+}22, WXG^{+}24, ZWZ^{+}23]. E-ffective [MHS^{+}22]. Ear [MKK20]. Earcut [LCSA22]. Early [CNAA^{+}22, PRJ^{+}23, SBW21]. Early- [CNAA^{+}22]. Earth [MRS22]. EarVR [MKK20]. Easy [ZCZ22]. Eating [WMB23]. EBBE [DAB^{+}23]. EBBE-Text [DAB^{+}23]. EC [KCK^{+}24]. ECG [TB24]. ECoalVis [LWT^{+}23]. Economy [TSHI22]. Ecosystem [NDP^{+}21]. Edge [FGS^{+}21, HRX23, WAA^{+}22, WXW^{+}20, ZPG21]. Edge-Compute [HRX23]. Edge-Path [WAA^{+}22]. Edinburgh [Ano22d]. Edit [SMKN20, SN23]. Editable [CLL^{+}20, XEXW24]. Editing [BMWD20, HBS^{+}21, SYC^{+}23, SCC^{+}23, WFY^{+}21, WCW^{+}22, ZGX^{+}22]. Editor [She23b, Mue20a, Mue20b, MB20c, Mue21b, Mue21a, MB21b, Mue22b, MB22b, Mue23b, She23a, SK23b, She24, SK24]. Editor-in-Chief [Mue20b, MB20c, Mue21a, MB21b, Mue22b, MB22b, Mue23b, She23a, SK23b, She24, SK24]. Editorial [CRH23, Mue22a, She23a]. Editors [BPQW23, NAW^{+}22, Ano21k, BSW20, CTW21, CRZ22, CRH23, FGL^{+}23, HKVZ20, IPPZ24, IMKP21, IGMW22]. Education [ALR23, BKR^{+}24, PLD^{+}23, RS23, SEK^{+}24]. EEG [TB24]. eFESTA [HGSP20]. Effect [APP^{+}22, DPR^{+}20, HHC^{+}21, JLK20, KKH21, Kel23, MWD^{+}23, MF24, NMS^{+}23, ONH21, RBLT^{+}22, SAMB^{+}23, SES20, SZH^{+}20, UVL^{+}23, WSN22, WSZ^{+}23, WKMD22, XQXL23, YHC^{+}24]. Effective [ALX^{+}22, BS22, CLX^{+}23, DIPJ22, MNZ^{+}20, RS21, ZCM23a]. Effectiveness [ABL^{+}22, BCT22, CCPM23, DTPG21, HHM^{+}24, LAN21, MHS^{+}22, NMR^{+}23, RQ21, VVR^{+}23a, VC20]. Effectors [VVR^{+}23a]. Effects [BRLR24, BKN^{+}22, BZW^{+}21, CZC^{+}20, CBE^{+}21, DPM24, EHB^{+}23, ENK^{+}20, GWL^{+}22, GMTD23, GWW^{+}21, GLY^{+}23, JKL24, JWE^{+}22, KOL^{+}20, KZD^{+}23, KL22, KPL23, KJL24, KMH^{+}23, KAL^{+}23, LKL23, LSE20, MNK23, PGE^{+}22, RMW^{+}24, SCL^{+}23a, eSYKW23, TARB23, VVR^{+}24, VVC^{+}24, WQP^{+}22, WZW^{+}23, WBPC23, YTHL23, YeSiK^{+}23, YGE^{+}21, ZSL^{+}22, ZCL23, BVV^{+}22, BVV^{+}23, HHC^{+}21]. Efficiency [CNAA^{+}22]. Efficient [AKS22, BLE^{+}23, CYD^{+}23, FWW^{+}23, FMBN23, GLA^{+}24a, GRi^{+}21, HSB^{+}21, HWL^{+}22, JBS^{+}22, LWL^{+}22a, MSuG^{+}23, NVRS^{+}21, SKNŽ20, WMZ22, WW22a, WWZP22, YZP^{+}23, ZWP21, ZCH^{+}21, vdRBE22]. Effort [CQHP22, VBV^{+}23, MSAM^{+}22]. EGG [TB24]. Egocentric [MMK^{+}23]. EHTask [HBLW23]. EIC [Mue23a]. Eiffel [HSS^{+}20]. Eigenvalues [HZFH20]. Eigenvectors [HZFH20]. Eindhoven [Ano22b]. Elastic [XCK20]. Election [YCM^{+}24]. Electrical [EBJ^{+}22]. Electroactive [HWN^{+}23]. Electrolytes [PSG^{+}22]. Electromagnetic [TSS22, ZLL^{+}20]. Electromagnetism [RS23]. Electromechanical [HWN^{+}23]. Electron [NBE^{+}23]. Electronic [LLP^{+}23, SNBC23]. Electrovibration [ZLS21]. Element [LLCH22, SKA21]. Elements [CNY22, MUM^{+}21]. Elevation [ZCL23]. Eleven [CBL^{+}24]. Elicitation [WGO20, WGO22, WO22, WGH^{+}24]. Eliciting [LIB24]. Ellipse [KCB^{+}21]. EllSeg [KCB^{+}21]. embComp [HKMG22]. Embedded [CYD^{+}23, GL20, LCY^{+}23, NMS^{+}23]. Embedding [EAKC^{+}20, FZC^{+}21, HHKN21, HH20, HBMK23, JZCZ20, LLWF23, LWBM22, MHIS23, MBC^{+}23, RWB^{+}23, TIHS20, WHX^{+}23, XZF^{+}22, XHL^{+}24, YVBI24, YLLW24, ZLW21a, ZZW^{+}22c]. Embedding-Based [WHX^{+}23]. Embeddings [EHA^{+}23, GNMQ22, HKMG22, LqZ23, SSJ^{+}22, XTYL20, vOVR23]. Embellished [ACPB24]. Embellishments [ACPB24]. Embodied [BBH24, KSK^{+}23, PTX^{+}22, SWF^{+}24, YCB^{+}21, MCSAL23]. Embodiment [BCB22, CNK^{+}24, DGB^{+}22, DPM24, FALH20, FOH^{+}21, GLH22, MWD^{+}23, PT20, PHB^{+}22, RL20, SAMB^{+}23, SFPM22, UVL^{+}23, YGE^{+}21, KMH^{+}23]. Embodying [GWF^{+}23, VAWL24]. Embryonic [SKW^{+}23]. Emergent [HDFK21]. Emerging [NP24a]. EmoCo [ZWW^{+}20]. Emojis [GTL^{+}23]. Emotion [LPG^{+}22, LPL^{+}24, LFG^{+}23, MHS^{+}22, WLF^{+}22, YCM^{+}24, ZWW^{+}20, ZSW^{+}21]. Emotion-Oriented [ZSW^{+}21]. Emotion-Preserving [WLF^{+}22]. Emotional [JWE^{+}22, LWC24, PZC^{+}23]. EmotionCues [ZSW^{+}21]. Empathy [GZG^{+}24, LSG24]. Emphasis [KCK^{+}24]. Empirical [ESB^{+}24, MMF20, NDP^{+}21, PSH21, RBLE22, WAV^{+}21, XZS^{+}22]. Empirically [ALX^{+}22]. Employee [ZKS^{+}20]. Empowering [GGP^{+}24]. Empowerment [WAC^{+}22]. Emptying [PAAG22]. Enabled [HXL^{+}24, LELT23, NIK24]. Enabling [CXXZ21, LBL^{+}21, LCW^{+}23b]. Encoded [RSRG23]. Encoder [ZLQH21, ZLW21a]. Encoder-Decoder [ZLQH21]. Encoding [EKC^{+}23, FBB^{+}21, HSB^{+}21, MYS^{+}22, NF20, QGY^{+}22, WMZ22]. Encodings [CMXF21, SHPE20]. EnConVis [ZLC^{+}23]. Encountered [MML21]. Encountered-Type [MML21]. Encourage [HX23]. Encrypted [MPV21]. End [HZCW22, VVR^{+}23a, YXS^{+}23]. End-to-End [HZCW22, YXS^{+}23]. End-User [CCEA^{+}24]. Energy [CNAA^{+}22]. Energy-Efficiency [CNAA^{+}22]. Enfacement [GFSHO20]. Enforcement [RHLC22]. Engagement [WQP^{+}22]. Engaging [JSA^{+}20]. Engine [NFN^{+}23]. Engineering [CMKK22, FWW^{+}24, HFS^{+}21, SFSA24, SWS^{+}23]. Engines [EBJ^{+}22]. Enhance [CSM^{+}23, LKS22]. Enhanced [iKYOW23, NP24b, PTS^{+}20, ZUK21, ZWP^{+}22]. Enhancement [DGKOC20, LQWQ21, SS21]. Enhancements [CBW23, ZCM23a]. Enhancing [CWL23, KGR^{+}24, LFW^{+}22, TNZ^{+}24]. Enjoyment [RBLE22]. Enrich [OWW^{+}24]. Enriching [LVR^{+}24]. Ensemble [CMKK21, HGSP20, HWG^{+}20, SXL^{+}23, YYZ^{+}22, YC23, ZCL^{+}21, ZLC^{+}23]. Ensembles [AMY^{+}22, CAA^{+}20, CZC^{+}20, FJK^{+}20, KBB^{+}23, KSHW22, NP21, dSBdO^{+}24]. Enthusiastic [HWT^{+}24]. ENTROPiA [MML21]. Entry [AV21, BTHL23, FMBN23, KG24, LWL^{+}22a]. Environment [AYA^{+}21, EHB^{+}23, KSB^{+}22, LHC^{+}21, LJCL24, MWD^{+}23, NCB^{+}21, NP24b, RR23, SKW^{+}23, KOL^{+}20]. Environment-Aware [NP24b]. Environmental [TKIS24, WCCS24]. Environmentally [FYE^{+}22]. Environments [AV21, AeSL^{+}23, AI21, BCB22, ENK^{+}20, FTWP22, FYC^{+}23, FGS^{+}22, GSH21, GMVRB20, GLY^{+}23, HLJ^{+}22, HGB22, HBLW21, KCA^{+}21, KVB20, KRK21, KJL24, KCWK20, LFCH24, LGY^{+}22, LF23, LXZ^{+}23, MMS^{+}23, MSGM23, MNK23, MM23, NBNC20, NMR^{+}23, SVK^{+}21, SFPCW23, TIDQ24, WSL^{+}24, WQP^{+}22, WBF20, WF21, WBG^{+}23, WBM21a, YDM^{+}21, YYK^{+}22, YPW23, YLF^{+}20, ZCZ^{+}23a, ZTL^{+}23, KOL^{+}20]. Envision [BXF^{+}22]. Epicentric [APSB23]. Epidemic [YZF^{+}23]. Epidemiological [RBF^{+}23]. EpiMob [YZF^{+}23]. Epipolar [KJI^{+}21]. Equalizer [ESP20]. Equations [ZHL^{+}21b]. Equilibrium [KSHH22]. equity [WKMD22]. Erato [SCC^{+}23]. ERGOBOSS [ZLC^{+}22]. ErgoExplorer [FRiM^{+}23]. Ergonomic [FRiM^{+}23, MZX^{+}21, ZLC^{+}22]. Errata [FCFC22a, WGO22]. Error [ABL^{+}22, JWW^{+}23, KLSB22, MNK23, PVPM22, SPW^{+}22, YLGW24]. Error-aware [SPW^{+}22]. Error-Bounded [YLGW24]. Errors [ISBP22, TSS22]. Essays [KRW^{+}21]. Estimate [DKuH20, JRM22]. Estimates [HGSP20, MPNF24, RRG23, XCLF20, XF21]. Estimating [CBP22, JQL^{+}24, RLLS20]. Estimation [AYA^{+}21, CLZ^{+}24, CXW^{+}23, JKU^{+}22, LWL^{+}23b, PPYW20, RBRG21, XSKF23, YTHL23, ZHDX20, ZCR21]. Estimative [HT22]. Ethnicity [DPM24]. Euclidean [ZZZ24]. Euler [GLHQ21, KGWD22]. EV [LCC^{+}23, ZCX^{+}24]. EV-LFV [LCC^{+}23]. Evaluate [WAV^{+}21]. Evaluating [ALX^{+}22, BCC^{+}20, BMFE20, CWO^{+}24, CCS^{+}21, DZG^{+}23, EHB^{+}23, EKC^{+}23, FSN20, GWL^{+}22, HZJ^{+}24, LVR^{+}24, LWM^{+}20, NL24, RQ21, SGH^{+}23, WMZ^{+}20, ZKS^{+}20, ZSL^{+}22, ZSCRB23, ZBG^{+}24, ZCM22]. Evaluation [ASA^{+}23, ACT^{+}24, AYG^{+}22, BTv^{+}23, BLIC20, BC21, CCP^{+}21, CXZ^{+}24, CCEA^{+}24, CGT^{+}24, FOH^{+}21, GBL^{+}22, GTC^{+}23, HPP^{+}20, HWW^{+}22, JKA^{+}24, KCA^{+}21, KKEG20, KMM24, LCH^{+}21, LTLB22, LLP^{+}23, LCM^{+}23, NDF^{+}21, PCQ^{+}20, PSS^{+}22, RRK^{+}22, RBF^{+}23, SLC21, SMPJ^{+}20, WAP^{+}21, WHBC23, YLF^{+}20, YXX^{+}21, ZMD^{+}22, ZZD^{+}23b]. Evaluation-Focused [ZMD^{+}22]. Evaluations [GNMQ22, YBVI22].

Evaluator [CXZ^{+}24]. Evasion [LWW^{+}21]. Even [GKC^{+}24]. Event [AMK^{+}21, APSB23, BZSD21, GZW^{+}20, GJC^{+}22, GGJ^{+}22, JGC^{+}21, LCC^{+}23, MSM^{+}22, RBLT^{+}22, SSZ^{+}21, SAK20, VBV^{+}23, WLGW23, YDMP22, ZSZ^{+}24]. Event-Based [AMK^{+}21, SAK20]. Events [FYE^{+}22, NNF^{+}20, ZTC^{+}23]. Ever [WLG^{+}22]. Ever-Changing [WLG^{+}22]. Everything [WFW^{+}20]. Evidence [BWWL22, PGS21]. EVis [FYE^{+}22]. EVM [KGQ^{+}24]. EvoBio [RPH^{+}21]. Evolution [BB21, DLW^{+}23, GLL^{+}22, LZJZ20, XTYL20]. Evolutionary [HSS^{+}20, HZJ^{+}24]. Evolving [BNRB21, WBI20]. ExaBricks [WZU^{+}21]. Examination [WN21]. Examine [RMW^{+}24]. Examining [BRLR24, CQHP22, GSM^{+}22, APBB24]. Example [HMTI24, QSC^{+}21, YTL^{+}22]. Example-based [QSC^{+}21, YTL^{+}22]. Examples [BLBL23, DV23, JGH^{+}24, OYK^{+}21, YXW^{+}23]. Excavations [SFSA24]. Excessive [PALW20]. Exchange [GLY^{+}23]. Execution [BS22, SYY^{+}24]. Executive [Ano22-28, Ano23y, Ano24u]. Exemplar [HYH^{+}23, MSK23b, PCZ^{+}21, ZLY22, ZCM23a]. Exemplar-Based [HYH^{+}23, MSK23b, ZLY22, PCZ^{+}21]. Exemplars [LPJT^{+}22]. Exercise [SACB^{+}23]. Exergames [GWD^{+}24]. Exertion [LCM^{+}23]. Exocentric [BWCT23]. Expansions [JKU^{+}22]. Experience [ASP^{+}22, CFDN24, GWD^{+}24, HXL^{+}24, KAL^{+}23, LVR^{+}24, MMK^{+}23, MGM^{+}22, eSYKW23, WSZ^{+}23, WIP^{+}24]. Experiences [BJR21, JVRL24, JLM^{+}21, KL22, KPL23, KJL24, LCY^{+}23, LDZ^{+}21, MCS^{+}23, NBE^{+}21, TNZ^{+}24, VVR^{+}24, WBLW23, ZHH22]. Experiment [LZP^{+}20]. Experimental [SNH^{+}23, SGB^{+}22]. Experiments [ADM^{+}22, LZP^{+}20, RPH^{+}21]. Expert [BCC^{+}20, DWS^{+}23, WYIS24, CCEA^{+}24].

Expert-Annotated [DWS^{+}23].

Expert-Led [WYIS24]. Expert/End [CCEA^{+}24]. Expert/End-User [CCEA^{+}24]. Experts [SML^{+}23, XDBAR24, YSM^{+}20]. Explain [CLD^{+}22, KLKE21, RBSN22]. Explainability [ECR22]. Explainable [JLCZ22, NP21, PvSvdE^{+}24, SSSEA20, TXM23, WTD^{+}21, ZCX^{+}24]. Explainer [WTS^{+}21, SSSEA20]. Explaining [DAB^{+}23, JKJ^{+}22, LCYQ24, MXLM20, MFH^{+}21, MWJ22, WHJ^{+}22, YXW^{+}23]. Explanation [CXY^{+}22, HMKB23, NP24a, SSJ^{+}22]. Explanations [CMQ21, HSYZ24, LCYQ24]. Explanatory [RBSN22]. Explicit [CK22]. Explorable [LCYQ24]. Exploranative [MLBW20]. Exploration [ASSB^{+}23, BHU^{+}21, BHA^{+}23, BB21, CWB^{+}20, CAA^{+}20, CXY^{+}22, CCW^{+}21, EHA^{+}23, FSN20, FAS^{+}21, GSS^{+}23, GMS^{+}21, GBNH21, GLY^{+}23, GL20, HMGO23, HTJ^{+}20, HGSP20, HWG^{+}20, HGB22, HE24, Iur22, JTT^{+}23, KKG^{+}20, KPL23, LZW^{+}23, LME^{+}23, LZC^{+}21, LIB24, LS23, LWW^{+}21, LPP^{+}23, LRZ^{+}23, MMS^{+}23, MGO21, OCL^{+}21, PFC20, PBBH20, PSG^{+}22, RBKM24, RGG20b, SSS20, SHOP23, SUB^{+}22, SESH24, SXW^{+}22, SXL^{+}23, SVK^{+}21, SLS21, UWF^{+}23, VVR^{+}24, WSN22, WWSS20, WLS^{+}22, WN21, WHL^{+}20, WQ20, WWZP22, XDW21, YS20, YBOB24, ZeB^{+}21, ZMK^{+}20, ZCL^{+}21, ZHTR22, ZWW^{+}23b, ZWZ^{+}22, vOVR23]. Exploratory [CXD^{+}21, FPK^{+}24, KGQ^{+}24, MRS22, SCC^{+}24, WYIS24, XDW21, ZFF22]. Explore [FW22, KKE21, LLT24, LY23, LLZ^{+}23, PLL^{+}24]. Explorer [CMQ21, HM22]. Exploring [ABE^{+}22, BGB^{+}22, BZP^{+}20, BBSC23, CSL^{+}21, CFDN24, CSIP22, DPC23, DLW^{+}23, DAB^{+}23, DWH^{+}23, FLS23, GSL21, GST^{+}24, GPR^{+}24, GWD^{+}24, HBS^{+}21, HLM^{+}20, JCZ^{+}24, KJS^{+}23, KG24, LKJ^{+}20, LZJZ20, LQS^{+}23, LZZ^{+}24, MHS^{+}22, MGWM22, RPNP23, SSAZ22, TCX^{+}23, UDH23, WNC^{+}22, WBLW23, XMK^{+}22, XTYL20, YLS^{+}23, YCC^{+}21, ZZW^{+}22a, ZYM^{+}24]. Expo [TEK^{+}23]. Exponents [RGG20a]. Exposition [BDRW21]. Exposure [ZCL23]. Expression [NMS^{+}23, ZCZ22]. Expressions [JYZW20]. Expressive [JSA^{+}20, PTC^{+}24, RAP21, SLC^{+}23, TLW^{+}21]. Expressiveness [LSW^{+}22]. Extended [ATHI24, CBHR^{+}23, NGW^{+}24, ZGL^{+}21]. Extending [WHC^{+}23]. Extensible [CWW^{+}20, MBBB21, VKT^{+}24]. extensions [BNWvW21]. External [GBNH21, MPWN22]. Extract [SLSW23, WWZ23b, ZYC24a]. Extracting [CWL23]. Extraction [BSB^{+}20, CMKK22, CWW^{+}20, KLB24, LWC22, NWMC23, QHL^{+}20, QRZZ21, RZW^{+}24, TSH21, YISG21, YZF^{+}22, ZXSR22, ZPF^{+}22, PGL^{+}20]. ExTreeM [LWW^{+}24]. Extremely [Liv21]. Extremum [LWW^{+}24]. Eye [AMK^{+}21, EML^{+}23, FMP23, HWMI23, HBLW21, HBLW23, JJKJ20, LH23, MDV20, RZLX24, SPWW^{+}24, SCHE23, WCX21b, ZSZ^{+}24, CNB^{+}22, DJHBJ21, DJBJ23, RSAA20, ZSZ^{+}24]. Eye-dominance-guided [MDV20]. Eye-Hand [RZLX24]. eye-images [CNB^{+}22]. eye-tracking [DJHBJ21, DJBJ23]. Eyeglasses [ILZ^{+}21, UIHS20]. Eyes [ZYR^{+}20, CNB^{+}22]. Eyes-free [ZYR^{+}20].

F2 [WCW^{+}22]. F2-Bubbles [WCW^{+}22]. Fabric [KAS^{+}22, PIS20]. Fabricate [BFY^{+}24]. Fabrics [XWZB21]. Face [CAGM22, DMJ^{+}22, LH23, PSL23, WLF^{+}22, WCH^{+}23, YXS^{+}23, ZNF^{+}23]. Facet [AFB22]. Faceted [YST^{+}20]. Facetto [KBJ^{+}20]. Facial [CSWZ24, CBW23, GFSHO20, LWL^{+}23a, LHL^{+}22, NMS^{+}23, ZCZ22]. Fact [SCC^{+}23, WSZ^{+}20]. Factored [KPW20]. Factories [SHC^{+}20]. Factors [CIA24, KHL21]. Facts [CCS^{+}24]. Fail [WLZ^{+}23]. Fairness [AL20, XMK^{+}22]. FairRankVis [XMK^{+}22]. FairSight [AL20]. Faithful [RSD^{+}23, SLC^{+}23, WCW^{+}22]. Faithfulness [LLC^{+}20]. Fallback [SPW^{+}22]. FAME [GLL^{+}22]. Farewell [Mue23a]. Farther [KBPR22]. Fast [AT23, BD22, BWM23, CXW^{+}23, FiMH21, GFCM23, GVT24, LWWH20, NGBA^{+}20, PD24, SSB^{+}22, SHM23, SDK23a, XZF^{+}22, ZZX^{+}22, ZP24]. Fast-Startup [ZP24]. Fate [CRP^{+}20]. Fatigue [BTHL23]. Fatigue-Reducing [BTHL23]. fCoSE [BD22]. Fear [LH22]. Feasibility [KKEG20, SMPJ^{+}20, TPH22]. Feature [ATK^{+}24, CMKK22, GLX^{+}21, HGSP20, HZYZ22, JH20, LS23, LDC^{+}23, LH23, LYH^{+}23, PC22, PSL23, SHD^{+}20, SIA^{+}23, ZLW^{+}21b, ZCZ^{+}21, PGL^{+}20]. Feature-Preserving [LDC^{+}23, SIA^{+}23, ZLW^{+}21b, ZCZ^{+}21]. FeatureEnVi [CMKK22]. Features [AT23, CLD^{+}22, CLL^{+}23, CWL23, DCC22, MvdEPV24]. Federated [CWW^{+}23, LWL^{+}22b, WCX^{+}23]. Feedback [FMBN23, GST^{+}24, GPR^{+}24, KCGZ23, KZD^{+}23, KVM^{+}22, MCQ23, SWSK23, TWT^{+}22, WGH^{+}24, YKJ^{+}23, ZUK21, ZZS^{+}23]. FEM [JLK20]. Female [PSH20]. Ferrofluids [SHM23]. Few [JGH^{+}24, YYZ^{+}22]. Few-Shot [YYZ^{+}22]. ffective [MHS^{+}22]. FibAR [TIHS20]. Fiber [AJSP23, XNF^{+}23]. Fibers [TIHS20, ZS21]. Fibonacci [MBBB21]. Fidelity [CDX^{+}20, KHI20, MSRJ20, MNK23, RTM^{+}20, Tet24]. Fiducial [YHD21]. Field [BWM24, BKPB21, CJY^{+}23, FHTB23, FMP23, HWN^{+}23, HSB^{+}21, JLK20, LDC^{+}23, LCC^{+}23, LWWY21, MTE^{+}20, MDS24, NIM^{+}21, QCCC23, RG20, SES20, SLR20b, SWZ^{+}23, TME^{+}22, VGT21, WWS20, WHBC23, XEXW24, YIIW24, ZSLL20, ZGL^{+}21, ZHL^{+}21b, VVR^{+}23a, VVC^{+}24]. Fields [CHSC24, DHY^{+}22, EIB23, FV24, HCWK23, HZYZ22, HZZ24, KRZ^{+}20, KSHW22, MDJV21, Pat22, PL21, QRZZ21, SCL^{+}23b, WPZ^{+}23, ZZW^{+}22a, ZLZ^{+}23a]. Figures [CLL^{+}21, ZTC^{+}23]. Filling [KAS^{+}22, ZJW21]. Filter [WFC21]. Filtered [QFWS22]. Filtering [HSC^{+}22, KHI20, LLZ^{+}21, LSL^{+}22, ZLQH21, ZCR21, ZLW^{+}21b]. Find [BLBL23, XGS^{+}21]. Finding [NBE^{+}23]. Findings [LLY^{+}24]. Fine [DWS^{+}23, KZD^{+}23, PCZ^{+}21, SJL^{+}23, WWZP22, ZG20]. Fine-Grain [WWZP22]. Fine-Grained [DWS^{+}23, SJL^{+}23, ZG20]. Fine-tuning [PCZ^{+}21]. FineStyle [SJL^{+}23]. Finger [RZLX24]. Fingers [XDH^{+}22]. Finite [RGG20a, VM23]. Finite-Time [RGG20a]. Fire [CWO^{+}24]. fired [LWT^{+}23]. Firefighters [NMR^{+}23]. First [SAMB^{+}23]. First- [SAMB^{+}23]. Fit [RBRG21, RRG23]. Fitting [NCE23, SDR22, SFSA24]. Fitts [BS22, KVM^{+}22]. FixationNet [HBLW21]. Fixations [HBLW21]. Fixed [GBNH21]. Fixer [CSX^{+}22]. Flash [Tet24]. Flat [BJCL21]. Flattened [EEL^{+}20]. Flattening [NGBA^{+}20, ZTWW21]. FleXeen [PIS20]. Flexible [HSB^{+}21, HSC^{+}22, MCW^{+}20, OBCT24, SLS21, Tsa21, VKT^{+}24, WCW^{+}22]. Flight [CBHR^{+}23]. Flipped [BBSC23]. Floor [JLM^{+}21]. Floor-vibration [JLM^{+}21]. Floorplan [WZC^{+}23]. Flow [HSS^{+}20, HXHT23, JTT^{+}23, JGG21, LLD^{+}21, MNB^{+}23, NSG^{+}20, PWK21, RMB^{+}21, SPWW^{+}24, SS24, SLC21, SJL^{+}23, SKR^{+}24, XQXL23, ZHTR22, MDH^{+}23]. Flow-Based [SS24, XQXL23]. Flower [LWL^{+}22a, vOVR23]. FlowNet [HTW20].

FlowNL [HXHT23]. Flows

[DWB21, MDH^{+}23, NWMC23, NNF^{+}20, RMB^{+}21, RG20, ZYC24a, ZHL^{+}21b]. FlowSense [YS20]. Fluid [CLFL22, FLX^{+}23, HWC23, HE24, JGG21, RHLC22, XZWY20, ZHQH20]. Fluids [KCGZ23, LHWW22]. Fly [WZH20]. FlyFusion [XCG^{+}21]. Flying [HLA^{+}22, MSRJ20, XCG^{+}21]. fNIRS [GSL21]. foams [PGL^{+}20]. Focal [APP^{+}22, ISKM23, SES20, UIHS20, YLJ^{+}22]. Focus [ALX^{+}22, EMM^{+}22, EML^{+}23, MGWK^{+}22, UIHS20, HTJ^{+}20, HBS^{+}21, JKW^{+}22, NWW21, NSW24]. Focused [HNGC21, ZMD^{+}22]. Folded [KLB24]. Folding [MOA21]. Folds [ZH20]. Foliation [WWZ23b]. Follow [HAK^{+}23]. Follow-the-Path [HAK^{+}23]. Following [LETF21]. Follows [BBH24]. Football [AAA^{+}21]. Force [CLH^{+}23, DS20, KCGZ23, iKYOW23, SHW^{+}20, XWZ^{+}23]. Force-Based [CLH^{+}23]. Force-Directed [SHW^{+}20]. Forecast [HLCY22, PFCB23, YCM^{+}24]. Forecasting [HSYZ24, HBLW21, JLP^{+}23, LKJ^{+}20]. Forest [KGBP20, NP21]. Forgetting [MNK23]. Form [BZP^{+}20]. Formation [HXL^{+}24]. Formats [LRHA22]. Formulator [WTL24]. forth [XCG^{+}24]. forward [DGD^{+}23]. Four [LS22]. Four-Level [LS22]. Fourier [MBB20, RGG20b]. Fourth [SGJC23]. FOV [BJCL21, RSAA20, DHY^{+}22]. FoV-NeRF [DHY^{+}22]. Fovea [KCA^{+}21]. Foveated [BWM23, CZT^{+}21, DHY^{+}22, FBB^{+}21, HRX23, LDB^{+}21, MDV20, MDJV21, SWWY21, SWW^{+}22, SBK^{+}20]. FoVolNet [BWM23]. Fractures [LLCH22]. Fragments [YCC^{+}23]. Frame [FHTB23, MNZ^{+}20, RG20, WSN22, WSZ^{+}23, ZSCRB23]. Framerate [MOA21]. Frames [SFL^{+}22]. Framework [BMWD20, BMA^{+}23, BJR21, CSX^{+}22, CKQ^{+}23, CLL^{+}23, DLH^{+}22, ESP20, ESB^{+}24, FSS^{+}21, GLX^{+}21, HTW20, HZCW22, HGB22, HZJ^{+}24, KBM21, KCB^{+}21, LVV^{+}21, LELT23, LFMM24, LMF^{+}24, LZC^{+}21, LCC^{+}20, LLC^{+}22, LYL^{+}23, MFH^{+}21, ML24, MHN^{+}24, NMC21, PvSvdE^{+}24, RQ21, SCEA23, SSSEA20, TWW^{+}22, TLW^{+}23, VKT^{+}24, WWZ^{+}23a, WTD^{+}21, XHS^{+}24, XMK^{+}22, YGP^{+}24, YWM^{+}23, ZMD^{+}22, ZZW21, ZLC^{+}23, ZJJH21, ZCM22]. Fraud [ZWW^{+}23c]. FraudAuditor [ZWW^{+}23c]. Free [BZP^{+}20, CZF^{+}23, FCFC22a, FCFC22b, JLX^{+}23, LTLB22, LSL^{+}23, LXZ^{+}23, MMS^{+}23, SDK23b, SMSK22, XLY^{+}22, ZLZ^{+}23a, MGC^{+}21, ZYR^{+}20, DGD^{+}23]. Free-Form [BZP^{+}20]. FREE-RDW [DGD^{+}23]. Free-Viewpoint [CZF^{+}23]. Freeform [DLP^{+}23]. Frequency [LWWH20, ZSZ^{+}24]. Fresnel [BJCL21]. Freytag [YXL^{+}22]. Friction [DS20, TWT^{+}22]. Frictional [BKWK20]. FrictShoes [TWT^{+}22]. Front [Ano22i, Ano22j, Ano22k, Ano23h, Ano23i, Ano24i, Ano22h]. Frontal [BVV^{+}22]. Frontiers [WQQ^{+}24]. Frustration [LH22]. FSLens [CWO^{+}24]. FTK [GLX^{+}21]. Full [CGAG20, NVRS^{+}21]. Full-Body [CGAG20]. Fully [LCK^{+}21, MM20, SCRL20, YZN^{+}20]. Fully-Occluded [YZN^{+}20]. Fumos [LLDW24]. Functional [NQE21, SFS^{+}22]. Functionality [GLL^{+}22]. Functionality-Aware [GLL^{+}22]. Functions [GMTD23, PLL^{+}24, SFL^{+}22, ZH20]. Further [ZPG21]. Fusion [SJL^{+}23, ZX22]. Future [LJS21, MA23, TBW^{+}23]. Fuzzy [DWH^{+}23].

Gain [CS23, XCG^{+}24]. Gaining [WBG^{+}23]. Gains [WZZ^{+}23]. Gait [RBK^{+}21]. Galex [LZJZ20]. Galleries [PLL^{+}24]. Galvanic [GTH^{+}22]. Game [ALR23, LZW^{+}23, LCY^{+}23]. Game-Centered [ALR23]. Gameplay [LNB^{+}21]. Games [CB22, HNGC21, STA^{+}21]. Gamified [MWR^{+}22]. Gaming [KAL^{+}23]. GAN [LYZ^{+}21, LB22, SYX^{+}22, WZC^{+}23]. GAN-Based [SYX^{+}22]. Gap [CLA^{+}20, MGWM22, PSH20]. Gaps [SH24b]. Gaseous [LQWQ21]. Gathering [GF24]. Gaussian [LME^{+}23, NdCS21]. Gaze [AMK^{+}21, BBMM^{+}23, CZT^{+}21, CDS^{+}22, ENK^{+}20, HLZ^{+}20, KCB^{+}21, LTC21, LWK20, NMS^{+}23, RZLX24, SCHE23, SPW^{+}22, SPNG23, WCX21b, WZL22, WIP^{+}24, CNB^{+}22]. Gaze-based [SPW^{+}22]. Gaze-Contingent [CZT^{+}21, CDS^{+}22]. Gaze-Dependent [LWK20]. Gaze-Driven [LTC21]. Gaze-Vergence-Controlled [WZL22]. GCN [LB22]. GDR [FZZX22]. GDR-Net [FZZX22]. GE [XZF^{+}22]. Gemini [KH21]. Gender [CBL^{+}24, DPM24, KAL^{+}23, SFNRZ^{+}23, TDI22, WNC^{+}22]. Genealogy [WYC^{+}20]. General [ML24, PK21, SES23]. Generalizability [KKEG20, SH24b]. Generalized [LLW^{+}22a]. Generalizing [JH20]. Generated [QLC^{+}24]. Generating [BvOR21, LH23, MP21, NSS21, PJHY20]. Generation [BLE^{+}23, CCS^{+}24, CZW^{+}20, DMJ^{+}22, DWQW23, ERB^{+}21, FHTB23, GLL^{+}22, GL20, HW23, HLJ^{+}22, LMF^{+}24, LB22, LLZ^{+}23, MZX^{+}21, QSC^{+}21, SLX^{+}23, SZZW24, SXS^{+}21, SKK^{+}22, SYC^{+}23, WSZ^{+}20, WTY^{+}22, WGH^{+}24, YCZ^{+}22, YXS^{+}23, YHC^{+}22, YTL^{+}22, YSD^{+}23, ZYP^{+}24]. Generative [BKH22, CWB^{+}20, CFZ^{+}23, DHF^{+}22, HZCW22, HGO21, KM20, KKhCM23, MSB^{+}22, SC22, XHL^{+}24]. GenerativeMap [CWB^{+}20]. Generator [AeSL^{+}23]. Generic [LZC^{+}21]. Genes [DMMF21]. Genius [GWF^{+}23]. GenNI [SKK^{+}22]. Genome [HiM^{+}22]. Genomic [RVB^{+}22]. Genomics [LG23, PLW^{+}23]. GenoREC [PLW^{+}23]. Genus [Liv21]. Geo [JZHA22, HPAB23]. Geo-Context [JZHA22]. Geo-Storylines [HPAB23]. GeoChron [DCS^{+}24]. Geodesic [MXT^{+}22, PVT23, XZF^{+}22]. GeodesicEmbedding [XZF^{+}22]. Geodesics [MXT^{+}22, PVDT22, WPTG24]. GeoDualCNN [WCZ^{+}23]. GeoExplainer [LMF^{+}24]. GeoLinter [LFMM24]. Geological [SIA^{+}23]. Geometric [CFL21, CWR21, FZZX22, HYSL23, KMG^{+}21, LS23, LYH^{+}23, SKNŽ20, TGM21, WMZ^{+}20, WFC21, YWM^{+}23]. Geometrically [KAS^{+}22]. Geometries [EEL^{+}20]. Geometry [HLJ^{+}22, HBMK23, LHL^{+}22, LSL^{+}22, WCZ^{+}23, XDH^{+}22, YBR^{+}23]. Geometry-Aware [HBMK23, YBR^{+}23]. Geometry-Driven [XDH^{+}22]. Geometry-Guided [LHL^{+}22]. Geometry-Supporting [WCZ^{+}23]. Geono [DSKE21]. Geono-Cluster [DSKE21]. Geophysical [RMB^{+}21]. GeoSynth [PCJ23]. Geovisualization [FSN20]. Germany [Ano22e]. Gestalt [LCWL23]. GestOnHMD [CXXZ21]. Gesture [CXXZ21, KNL23, SDMK22, SDK23a, SDK23b, WGO20, WGO22, WO22]. Gesture-Based [SDK23b, CXXZ21]. GestureLens [ZWW^{+}23a]. Gestures [FHR^{+}21, GTL^{+}23, ZWW^{+}23a]. GestureSurface [XZS^{+}23]. Getting [WBF20]. GEViTRec [CFGM22]. ggdist [Kay24]. GIF [SWT^{+}21]. Gigantic [BPA22]. Gigapixel [LDZ^{+}21]. Git [KKJ^{+}21]. Githru [KKJ^{+}21]. Give [VVR^{+}23a]. Glance [HiM^{+}22]. Glints [FWW^{+}23]. Global [AKS22, AYA^{+}23, BHY^{+}23, BWM24, GMX23, HZZ24, NP21, XYF^{+}21, YCW^{+}24, ZWW22]. Globally [HCH^{+}23]. GLoG [NdCS21]. Glossiness [SA22]. Glyph [CSW^{+}20, FW22, HCX^{+}21, LGL^{+}23b, YSD^{+}23, ZBG^{+}24]. Glyph-Based [CSW^{+}20, FW22, YSD^{+}23].

Glyphboard [KKG^{+}20]. GlyphCreator [YTL^{+}22]. Glyphs [BKH22, KKG^{+}20, PL21, RLB^{+}21, YTL^{+}22, vOVR23]. GNN [SXW^{+}22, WHC^{+}23]. GNN-based [WHC^{+}23]. GNN-Surrogate [SXW^{+}22]. GNNLens [JWW^{+}23]. Good [MSK23c, Red23]. GoTreeScape [LY23]. government [PWB21]. GPGPU [PTM^{+}20]. GPU [CLFL22, FK20, HCJ21, MSuG^{+}23, PC22, SCRL20, SPN23, ZSL21]. GPU-Based [SCRL20]. GPU-Optimized [HCJ21]. GPU-Parallel [MSuG^{+}23]. GPUs [ZHQH20]. Gradient [ADD^{+}22, RHLC22, TYPC20]. Gradients [SWZ^{+}23]. Grading [HPM^{+}24]. Grafting [YISG21]. Grain [WWZP22]. Grained [DWS^{+}23, SYY^{+}24, SJL^{+}23, ZG20]. Grammar [BvOR21, CSC^{+}22, Kay24, KH21, McN23, MHN^{+}24, ZPWS23]. Grammar-Based [MHN^{+}24]. Granular [GLHQ21, ZHL^{+}21b, VGK^{+}22]. Graph [ADD^{+}22, ATAS21, ADM^{+}22, AFB22, BD22, BPA22, CWS^{+}20, CWW^{+}21, CSL^{+}21, CCS^{+}24, GLA^{+}24a, HSS^{+}20, JWW^{+}23, KM20, LWZ^{+}22, LCWL23, LIDM20, LWBM22, LLZ^{+}23, LZL^{+}23, SXW^{+}22, SAK20, SDXR22, SHW^{+}20, WJW^{+}20, WWW^{+}24, XMT^{+}21, XMK^{+}22, XWH^{+}23, XWZ^{+}23, ZLW^{+}21b, ZJC^{+}21, ZSL^{+}22, ZCSS23, ZSS^{+}21, ZCH^{+}21, vBMS22]. Graph-Based [BPA22, CWW^{+}21, LWZ^{+}22, LCWL23, ZLW^{+}21b, XMT^{+}21]. Graphic [KJS^{+}23, LYZ^{+}21]. Graphical [DPD^{+}24, EKC^{+}23, RS21, SHPE20, Tsa21, WWP22, ZYM^{+}24, ZCM23a]. Graphics [Ano20j, Ano20k, Ano21g, Ano21f, Ano22o, Ano22s, Ano23m, Ano23q, Ano24j, Ano24o, CSC^{+}22, CNAA^{+}22, GSM^{+}22, HSF^{+}20, Kay24, KH21, SSB^{+}22, TG24, ZPWS23, Ano20a, Ano20b, Ano20c, Ano20d, Ano21b, Ano24h]. Graphies [RAP21]. Graphs [CSJ^{+}21, CXD^{+}21, CSL^{+}21, HBS^{+}21, JLX^{+}23, KBB^{+}23, LAN21, LGW^{+}20, LWW^{+}24, MPNF24, RWB^{+}23, SLSW23, SKR^{+}24, WWZ23b, WWP22, WNV22, XCLF20, YST^{+}20, ZAH22, ZLZ^{+}23a, ZCH^{+}21]. Grasping [BPL23, GLK^{+}23]. Gravitational [XDH^{+}22]. GRay [LME^{+}23]. Grayscale [LCL^{+}22]. Greedy [DXX^{+}21]. Grid [APSB23, MSS21, RWB^{+}23, STD^{+}23, WW22a, WXS^{+}24, ZYC^{+}24b]. Grids [IWT^{+}20, MBBB21, MSuG^{+}23, WW22a]. GridSet [CNY22]. Ground [ASCR^{+}22, HWC23, HAK^{+}23, WBG^{+}23]. Ground-Based [HAK^{+}23]. Grounded [HWT^{+}24]. GroundFlow [HWC23]. Group [LLSM24, TNZ^{+}24, WBF20, WF21]. Group-Based [TNZ^{+}24]. Grouped [HPP^{+}20]. Grouping [WN21]. Groupings [MM20]. Groups [CCM20, EJS^{+}23, LWW^{+}21]. Growing [YZP^{+}23]. Growth [HS23]. Guaranteed [GLGB24]. Guarantees [CLG21]. Guardian [JLH24]. GUCCI [MNB^{+}23]. Guest [BPQW23, FGL^{+}23, HKVZ20, IMKP21, IGMW22, NAW^{+}22, Ano21k, BSW20, CTW21, CRZ22, CRH23, IPPZ24]. Guidance [BBH24, CRM^{+}24, CAGM22, CCEA^{+}24, GPR^{+}24, HTJ^{+}20, LLSM24, LQWQ21, MTE^{+}20, NBE^{+}21, SEK^{+}24, SCEA23, WCH^{+}23, WQP^{+}22, WFC21]. Guide [FHR^{+}21]. Guided [AeSL^{+}23, BSP20, EAKC^{+}20, HM22, KGB22, LZLS22, LSL^{+}23, LHL^{+}22, LZZ^{+}24, MNB^{+}23, MBS23, NBE^{+}23, PSH21, PMCM24, QCL23, RPG23a, SCC^{+}24, SHW^{+}20, WZC^{+}23, WF21, WGH^{+}24, MDV20, ZZ23]. Guided-Tours [AeSL^{+}23]. Guidelines [ALX^{+}22, BWI21, LJS21, RR23]. Guiding [KCK^{+}24, VLM^{+}23, WSL^{+}20]. GUIRO [BSP20]. Guitar [SKS^{+}23]. guitARhero [SKS^{+}23]. Guzheng [CFZ^{+}23].

H [SYW^{+}20]. H-CNN [SYW^{+}20]. Habits [ABGG21]. Hair [LCCZ22, SZF^{+}21]. Hairpin [ZYC24a]. Half [CPD20]. Half-Space [CPD20]. Hamiltonian [CSBK20]. Hand [BER^{+}23, DHLA22, DZG^{+}23, FGF^{+}21, GWC^{+}23, GPR^{+}24, iKYOW23, ONH21, PALW20, PT20, PGS21, RZLX24, VVR^{+}23a, VVR^{+}23b, WSN21, WH22b, XZS^{+}23]. Hand-Based [DZG^{+}23, GPR^{+}24]. Hand-off [BER^{+}23]. Handed [FGF^{+}21]. Handheld [CZF^{+}23, ZP23, ZJS^{+}22]. Handling [APHD24, BKWK20, MA23]. Handoff [WFW^{+}20]. Hands [BRLR24, LCM^{+}23, MGC^{+}21, leáB^{+}21]. Hands-free [MGC^{+}21]. Hands-on [leáB^{+}21]. Handwriting [FMBN23, KTL24]. Haptic [BGB^{+}22, CZGF21, DSD^{+}23, GST^{+}24, GTH20, KVM^{+}22, KG24, MTVS23, MML21, MHIS23, OF22, SEK^{+}24, WPNK21, YB20, ZUK21, ZZS^{+}23, ZJS^{+}22]. Haptic-to-Auditory [KG24]. Haptics [HWC23, HFT^{+}24, MKK20, ZUK21, ZP23]. HaptoMapping [MHIS23]. Hard [BBRE24, MKK20]. Hard-of-Hearing [MKK20]. Hardware [MSY22, SFS^{+}22]. Harnessing [JSA^{+}20]. Hashing [SYW^{+}20, YKJ^{+}23]. Head [ATHI24, BBVS^{+}24, CXXZ21, CCS^{+}21, FGF^{+}21, GSM^{+}22, HBLW23, IAI^{+}23, KCA^{+}21, Kel23, LSG24, LGL^{+}23a, MGWK^{+}22, NMS^{+}23, NIM^{+}21, PGS21, QSUK22, SPNG23, WSL^{+}24, WBA^{+}23, XCG^{+}24, YLJ^{+}22, ZWP^{+}22, ZHKY23, TWA22]. Head-Body [BBVS^{+}24]. Head-Motion [LGL^{+}23a]. Head-Mounted [CXXZ21, CCS^{+}21, KCA^{+}21, Kel23, LSG24, MGWK^{+}22, NMS^{+}23, NIM^{+}21, QSUK22, SPNG23, WSL^{+}24, WBA^{+}23, YLJ^{+}22, ZWP^{+}22, IAI^{+}23, ZHKY23, TWA22]. Head-Up [GSM^{+}22]. HeadJoystick [HLA^{+}22]. Heads [CFDN24, DHF^{+}22]. HEADSET [LFG^{+}23]. Headsets [CML24, CWSJ23, SDK23a]. Health [JCZ^{+}24, LLP^{+}23, ZWW^{+}23c, SNBC23]. Healthcare [CLD^{+}22, ERB^{+}21, ZCM22]. HealthPrism [JCZ^{+}24]. Hearing [GVN^{+}20, MKK20]. Heart [NMR^{+}23, SKW^{+}23, MSAM^{+}22]. Heartbeat [ENvBC23]. Heatmaps [SCHE23]. Height [ZHL^{+}21b, ZCL23]. Heightfield [CZGW23]. Helmholtz [Pat22]. Helmholtz-Hodge [Pat22]. Help [HZS^{+}22, RS23, RRG23]. Here [WCH^{+}22]. Heter [RXX^{+}21]. Heter-Sim [RXX^{+}21]. HeteroFusion [YLL^{+}20]. Heterogeneity [WCX^{+}23]. Heterogeneous [CJY^{+}23, HHB^{+}23, JKU^{+}22, RSAA20, RXX^{+}21, TKIS24, WYS^{+}22]. HetVis [WCX^{+}23]. Heuristic [CCEA^{+}24]. Heuristics [SCST24]. Hex [WLC^{+}23]. Hex-Core-MK1 [WLC^{+}23]. Hexahedra [KKV22]. Hexahedral [NWW21]. Hidden [KAS^{+}21]. Hide [DS22]. Hiding [HX23]. Hierarchical [GMS^{+}21, GZW^{+}20, HJL^{+}23, HSB^{+}21, JZCZ20, JLX^{+}23, KDEP21, KBJ^{+}20, LLW^{+}23, MM20, NHC^{+}20, NBJ^{+}21, RLB^{+}21, SXW^{+}22, WWW^{+}24, WGS^{+}23, YWL^{+}21, YBOB24, iIK^{+}21, YWL^{+}20]. Hierarchies [BNRB21, CCM20, LZD^{+}20]. Hierarchy [vdRBE22]. High [BXQ^{+}22, BWZ^{+}20, CLFL22, DTS^{+}21, EAS^{+}23, FBB^{+}21, GZW^{+}20, IRR^{+}22, JLK20, KKG^{+}20, KKT^{+}22, LMM^{+}21, LAC^{+}24, LCK^{+}21, Liv21, MGU^{+}21, PGL^{+}20, RTM^{+}20, SSJ^{+}22, Tet24, UIHS20, VSBY22, VKT^{+}24, WCJW22, WBG^{+}23, WN21, XZF^{+}22, XJZ^{+}21, YST^{+}20, ZFCG23, ZSZ^{+}24, ZRJW20, ZZW^{+}22c, ZLVV22, vOVR23]. High-Dimensional [BWZ^{+}20, DTS^{+}21, EAS^{+}23, GZW^{+}20, KKG^{+}20, MGU^{+}21, SSJ^{+}22, VKT^{+}24, WCJW22, WN21, XZF^{+}22, ZZW^{+}22c, vOVR23]. High-Dynamic-Range [ZRJW20]. High-Fidelity [RTM^{+}20, Tet24]. High-Frequency [ZSZ^{+}24]. High-Order [JLK20, YST^{+}20]. High-Performance [LMM^{+}21, LAC^{+}24]. High-Quality [CLFL22, IRR^{+}22, LCK^{+}21, XJZ^{+}21]. High-Resolution [FBB^{+}21, ZFCG23].

High-Speed [UIHS20]. High-Stakes [ZLVV22]. High-throughput [PGL^{+}20]. High-Volume [KKT^{+}22]. Higher [HNS23, LBP24, MPNF24]. Higher-Order [HNS23, LBP24]. Highly [JLM^{+}21, LDT^{+}21, WYZ^{+}21]. Hinder [RS23]. Histograms [ATAS21, BKS22]. Historic [PTS^{+}20]. Historical [OWW^{+}24, WLS^{+}22, XDBAR24, ZWW^{+}23b, ZTC^{+}23, ZJX^{+}24]. History [HPK^{+}22, KKJ^{+}21]. HisVA [HPK^{+}22]. Hit [BVV^{+}22]. HiTailor [LLW^{+}23]. HMD [FLS23]. Hoc [Wu22, SWS^{+}23]. Hodge [Pat22]. Hold [BFS^{+}24]. Hole [VE21]. HoloCamera [HLL^{+}24]. Holographic [CZT^{+}21]. Holomorphic [ZWW22]. Homogeneous [CBP22, PPYW20, VM23]. Homogeneous-Material [PPYW20]. Homology [DTS^{+}21, Iur22, SHW^{+}20]. Homomorphic [MPV21]. Homomorphic-Encrypted [MPV21]. Homophily [RSRG23]. HoopInSight [FS24]. Horizon [BBSvL24, WHSZ21]. Horizontal [LWL^{+}22b, WCX^{+}23]. Horizontally [BBSC23]. Hospitals [BPW^{+}21]. HotGestures [SDK23b]. HPUI [FGF^{+}21]. http [Ano20j, Ano20k, Ano21g, Ano21f, Ano22o]. Human [APP^{+}22, BBL^{+}22, CIA24, CYD^{+}23, EBKL21, FNB^{+}22, GM23, HMTI24, HHS^{+}23, HZC^{+}20, JGH^{+}24, JZCZ20, KJS^{+}23, LNB^{+}21, LqZ23, LXH^{+}21, LZL^{+}23, LFG^{+}23, MCFKF24, MBHE24, RZH^{+}23, SYHS20, SKK^{+}22, SWY^{+}22, VCO^{+}23, WNC^{+}22, WCX21a, WRZ^{+}21, WHSZ21, WZC^{+}23, WPZ^{+}23, YCZ^{+}22, YZF^{+}23, ZHDX20, ZCZ22, ZXSR22, ZCM^{+}23b]. Human-AI [SKK^{+}22]. Human-Assisted [JGH^{+}24]. Human-Centric [WZC^{+}23]. Human-Computer [WRZ^{+}21]. Human-in-the-Loop [GM23, ZCM^{+}23b, ZXSR22]. Human-Machine [FNB^{+}22]. Humane [SBe^{+}21]. Humanistic [LVV^{+}21]. Humanities [PLP^{+}23]. Humanness [PGE^{+}22]. Humans [JWE^{+}22, LCK^{+}23, MBHE24, RPHJ20]. Humor [WMW^{+}22]. Hunches [LAML23]. Hybrid [ADM^{+}22, BBL^{+}22, BLE^{+}23, EML^{+}23, GLHQ21, HPT^{+}23, HHK^{+}24, RBLE22, XWZB21]. Hydrodynamic [ABE^{+}22]. Hygroscopicity [MDX^{+}23]. Hyper [GT20, TNZ^{+}24]. Hyper-Objective [GT20]. Hyper-Realistic [TNZ^{+}24]. Hypergraph [FAS^{+}21, LWWY21, OZZ24, VBP^{+}21]. Hypergraphs [QZZ22, VBP^{+}21, ZRPW23]. HyperJump [AZA^{+}23]. HyperLabels [iIK^{+}21]. Hyperparameter [PNKC21]. Hyperproperties [HCM^{+}22]. HyperTendril [PNKC21]. Hypnotic [YLL^{+}22b]. HypoML [WAP^{+}21]. Hypothesis [CCW^{+}21, WAP^{+}21]. Hypothesis-based [WAP^{+}21]. Hypothesis-Driven [CCW^{+}21]. Hypothetical [ZAH22]. Hz [AMK^{+}21].

Icicle [JdJTC24]. Icon [RDH23]. Idea [GL20]. Identification [ENK^{+}20, GFSHO20, KKV22, KWO^{+}20, LYH^{+}23, QNWR23, ZCM23a]. Identifying [BFS^{+}24, PAPB20, SPJ^{+}23, SLK^{+}20, TAL24, WCX^{+}23]. Idiosyncratic [OBCT24]. IDLat [SLX^{+}23]. IEEE [Ano23f, Ano23g, Ano23a, Ano20j, Ano20k, Ano21g, Ano21f, Ano22l, Ano22n, Ano22m, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano23j, Ano23l, Ano23k, Ano23m, Ano23n, Ano23o, Ano23p, Ano23q, Ano24k, Ano24j, Ano24l, Ano24m, Ano24n, Ano24o, BSW20, BPQW23, CTW21, CRZ22, CLL^{+}21, CRH23, HCWK23, IPPZ24, MF22, MB20a, MB20b, MB21a, MB22a, NAW^{+}22, OK22, SFNRZ^{+}23, SK23a, Ano20a, Ano20b, Ano20c, Ano20d, Ano21b, Ano24h, TDI22]. If [WPB^{+}20]. II [ZWV21]. II-20 [ZWV21]. Illinois [Ano22c]. Illuminated [UIHS20].

Illumination [BWM24, CXW^{+}23, GMX23, JRM22, KGX^{+}23, XZXY22, XLL^{+}20, ZWXM23]. Illusion [BZW^{+}21, CZGF21, GFSHO20, HWS22, LGY^{+}22, XSHF20]. Illustrative [HMK^{+}20, SMCL24]. Illustratively [HiM^{+}22]. ImaCytE [SVK^{+}21]. Image [ACL^{+}24, APBB24, BHA^{+}23, BLE^{+}23, CFZZ23, FWZZ20, FWW^{+}24, HWG^{+}20, HYC^{+}23, HLH^{+}23, JKW^{+}22, KBJ^{+}20, LZZ^{+}20, LHL^{+}20, MLC^{+}20, MSK23a, MHIS23, PTD^{+}21, PMCM24, PBAG23, RPD22, RLLS20, SLJ^{+}20, STD^{+}23, SYC^{+}23, TG24, WYZ^{+}21, XWZ^{+}22, XHL^{+}24, XLL^{+}20, XPK^{+}24, ZFCG23, ZZW^{+}20, ZLZ21, dAWi^{+}23, ZWV21]. Image-Based [CFZZ23, HLH^{+}23, RPD22]. Image-Space [XPK^{+}24]. Imagery [ZPF^{+}22]. Images [CXW^{+}23, FZC^{+}21, GBM^{+}22, GXY^{+}20, HJZ^{+}21, LSE20, MMG^{+}21, MP21, MSB^{+}22, SLSW23, SMYF22, ZLW21a, CNB^{+}22]. Imaging [CJY^{+}23, PNB^{+}21, SVK^{+}21, WKN^{+}23]. Imma [LGWL21]. Immersion [FPG^{+}23, HBH^{+}20, KWO^{+}20, PLD^{+}23, SBW21, WAC^{+}22]. Immersive [ACL^{+}24, AeSL^{+}23, ASP^{+}22, BCC^{+}20, BBRE24, BWCT23, BCB22, BJR21, CGAG20, CXY^{+}22, DSD^{+}23, DLW^{+}23, EGP^{+}21, FLS23, FSN20, FP21, GSL21, GTL^{+}23, GSH21, GLY^{+}23, HZZ^{+}20, HBLW23, JK23, KCA^{+}21, KVB20, KBF22, LHC^{+}21, LGY^{+}22, LYBP23, LAC^{+}24, LDZ^{+}21, MMS^{+}23, MP21, MSGM23, MCS^{+}23, PGS21, PRKM20, SDC^{+}24, SPZS24, SFPCW23, WSN21, WSN22, WHY^{+}23, WBG^{+}23, WFK24, WZW^{+}23, WWS20, WIP^{+}24, YCB^{+}21, YDM^{+}21, YTHL23, YCC^{+}21, YCC^{+}23, ZMK^{+}20, ZYL^{+}22, ZYR^{+}20, MGC^{+}21]. Impact [APBB24, BBVS^{+}24, BCB22, CIA24, CB22, FRL^{+}23, FGS^{+}21, GMVRB20, GWD^{+}24, KHL21, KTL24, KSK^{+}23, KWO^{+}20, LWC24, LRHA22, LPL^{+}24, MWD^{+}23, MGWK^{+}22, PT20, PMS^{+}22, RMW^{+}24, SFPM22, SML^{+}23, VBV^{+}23]. Impacts [CFL^{+}23, FFB24, LCW^{+}23a, WO22, TWA22, VAWL24, YCM^{+}24]. Impairments [JMK^{+}22, MCQ23]. Imperceptible [MHIS23]. Imperfect [RMW^{+}24]. Implementation [HWMI23, ZWW22]. Implications [PGE^{+}22, RBR20]. Implicit [BKWK20, BXZ^{+}21, BBH24, CFQ21, CQ22, KLTB21, LZY^{+}23, LYZ^{+}24, LGL^{+}23b, PVPM22]. Importance [BDL^{+}21, DGDC21, SLX^{+}23]. Importance-Driven [SLX^{+}23]. Impression [SAMB^{+}23, SACB^{+}23]. Improve [CBW23, CNAA^{+}22, ECR22, GZL^{+}21, HAK^{+}23, HMTI24, KG24, LqZ23, LLP^{+}23, LETF21, RSRG23, SRKK21, TLBB23, BNWvW21]. Improved [KHI20, MCQ23, MCFKF24, SKA21, TIDQ24, ZZW21]. Improvement [ERB^{+}21, OSC22]. Improves [KPO^{+}23, NIM^{+}21]. Improving [ABL^{+}22, BBSC23, BZW^{+}21, DIPJ22, HLA^{+}22, HZS^{+}22, JKA^{+}24, JLP^{+}23, MUM^{+}21, TME^{+}22, VVR^{+}23a, WWL^{+}20, XVW^{+}21, dPLM21]. Impulse [FLX^{+}23]. Imputations [SGH^{+}23]. ImTooth [LZY^{+}23]. in-Vehicle [FTWP22]. In-View [RLG^{+}23]. In-Visualization [CBB23]. Inbetweening [LZLS22]. Inclusive [CBL^{+}24]. Incomplete [QHL^{+}20]. Incompressibility [RHLC22]. Incompressible [LHWW22]. Incorporating [FTWP22, KGQ^{+}24, LAML23, ZWW^{+}23d]. Incorporation [LqZ23]. InCorr [OWN^{+}21]. Increase [GFSHO20, HB24, MSAM^{+}22]. Increased [MOA21, NIM^{+}21]. Increasing [NCWE22]. Incremental [CZGW23, FCS^{+}20, MA20, MEB^{+}20]. Index [Ano20a, Ano22f]. Indexing [JSF20]. Indicators [GTH20, LCW^{+}23a]. Indirect [GMX23, KJI^{+}21, RRK^{+}22, XZXY22, ZWXM23]. Individual [CQHP22, CNY22, DPD^{+}24, TB24]. Individuals [FOH^{+}21]. Indoor [BHY^{+}23, GZY^{+}22, HLJ^{+}22, JRM22, KRK21, LZZ^{+}20, LXZ^{+}23, LHL^{+}20, PAAG22, PCJ23, SHS^{+}22, XLL^{+}20, XFF^{+}21, YFS22, YYK^{+}22, ZHL^{+}21a, ZZX^{+}22]. Induced [NDLW20, TSS22]. Induction [LPG^{+}22]. Industrial [COZ^{+}23, FMBN23, GEU^{+}22, LAS^{+}20]. Inertial [BXQ^{+}22, RZH^{+}23]. Infer [LHZ^{+}23]. Inference [GSK^{+}20, KKGMH21, RS21, WBL^{+}22]. Inferences [KWH22]. Inferential [NE24]. Inferring [MGO21, RPNP23]. Infill [XLY^{+}22]. Infinite [MML21]. Influence [ADL^{+}22, AI21, BER^{+}23, CFDN24, FMBN23, HSS^{+}20, LNB^{+}21, QR21, RBRG21, SACB^{+}23, WWP22, XSB^{+}22, ZYM^{+}24, GWF^{+}23]. Influences [AR22, CPD^{+}24, GSS^{+}23]. Info [Ano21h]. InfoColorizer [YZZ^{+}22]. Infographic [CWW^{+}20, CWH^{+}22, LSZC21, LCWL23]. Infographics [CZW^{+}20, QSC^{+}21, YZZ^{+}22]. Informatics [MGWM22]. Information [ALR23, APBB24, CLCY20, DFP^{+}20, ENK^{+}20, HH20, IMQ^{+}20, LWY^{+}20, LCWL23, LLZ^{+}23, MSMX23, MGO21, OCW^{+}24, RFD21, SFPM22, TQW^{+}24, WBL^{+}22, WJBB22, YXW^{+}23, ZLW21a]. Informed [ABGG21, YGP^{+}24]. INFOVIS [Ano21i, Ano20l]. infrared [CNB^{+}22]. Initialization [CBB^{+}21]. Initiative [CWH^{+}22, KRHH24, PBF^{+}21]. Injecting [HHS^{+}23]. Injection [YRL^{+}20]. Injured [KSK^{+}23]. Injuries [CNC^{+}20, KSK^{+}23]. Ink [PPE23]. InkSight [LLY^{+}24]. Inner [LHZ^{+}23]. Inner-Body [LHZ^{+}23]. InnovationInsights [WQQ^{+}24]. Inpaint [MSK23c]. InpaintFusion [MEB^{+}20]. Inpainting [BHY^{+}23, GTHC20, MEB^{+}20, MSK23b, ZFCG23, dAWi^{+}23]. Input [DZG^{+}23, FMP23, KNL23, KTL24, LCM^{+}23]. Inquiry [OCW^{+}24]. Insertion [HSJ^{+}20, BTv^{+}23]. Insets [LBB^{+}20]. Insight [ABGG21, DWQW23, HMH^{+}21, KHL21, LH22, SEAB^{+}22]. Insight-Driven [DWQW23]. Insights [KRS^{+}22, LRHA22, PJHY20, RPH^{+}21]. InSituNet [HWG^{+}20]. Inspecting [COZ^{+}23, LWL^{+}22b, ZWZ^{+}23]. Inspection [NWW21]. Inspiration [BKH22, WAC^{+}22]. Inspirational [MHS^{+}22]. inspired [PPYW20]. Instance [CGT^{+}24, HCL20]. Instant [CDS^{+}22, CBB^{+}21, PJYW21, PAAG22, RAFSA23a, SFSA24]. Instantaneous [DMTD22]. Instructions [SMSK22]. instrument [BNWvW21]. Insurance [ZWW^{+}23c]. Integer [HLH^{+}23]. Integral [PBBH20, RZW^{+}24, SLC21]. Integrals [MBB20]. Integrated [ALC22, MGU^{+}21, XXM^{+}21]. Integrating [AZA^{+}23, FWZ^{+}20, HPAB23, IAI^{+}23, LM21, MPWN22, PBF^{+}21, SSC^{+}23, YLL^{+}20]. Integration [KPO^{+}23, WMB23, XLZ24]. Intelligence [FWZ^{+}20, HHS^{+}23, WWS^{+}22, ZFF22]. Intelligent [ZWV21]. Intended [CCL^{+}24]. IntenSelect [KGR^{+}24]. Intensity [HHI24, KHI20, NLPW22]. Intent [PSS23]. Intent-based [PSS23]. Intention [YCM^{+}24]. Intentional [LGL^{+}23a]. Inter [JKJ^{+}22]. Inter-Cluster [JKJ^{+}22]. Interaction [BSG^{+}20, BBSC23, CXXZ21, DP20, DGKOC20, FMP23, GSK^{+}20, HMGO23, HDFK21, KWFK20, LNB^{+}21, LQS^{+}23, LLY^{+}24, LZP^{+}20, LCM^{+}23, MTVS23, MSRJ20, MM23, SHPE20, SSS20, SYHS20, SGJC23, SPW^{+}22, SFPCW23, SLS21, VVC^{+}24, WHX^{+}23, WNC^{+}22, WRZ^{+}21, WMMB23, WH22b, YLS^{+}23, MGC^{+}21]. Interaction-Based [SYHS20]. Interactions [FGF^{+}21, GSL21, GVN^{+}20, HMH^{+}21, LIB24, TCX^{+}23, VVR^{+}23a, VVC^{+}24, WSN21, ZHL^{+}20]. Interactive [ABE^{+}22, AJVS^{+}23, BS21, BHU^{+}21, BMFE20, CMK20, CAA^{+}20, CWW^{+}21, CYL^{+}21, CGT^{+}24, CKQ^{+}23, CLL^{+}23, CBE^{+}21, DSKE21, DTPG21, EBPF21, EGMP24, FWW^{+}24, FRiM^{+}23, FWZM22, GMX23, GSL^{+}24, GMS^{+}21, GWW^{+}21, GBNH21, GNMQ22, GTC^{+}23, HWW^{+}24, HKMG22, HB20, HCJ21, HM22, HJZ^{+}21, HMKB23, KDEP21, LZK^{+}22, LME^{+}23, LAN21, LZC^{+}21, LM20, LLW^{+}23, LWW^{+}21, LDT^{+}21, MM20, MNPP23, MBHE24, MZS^{+}24, MRS22, MBS23, NWW21, NHC^{+}20, OWN^{+}21, OF22, PLW^{+}23, PPE23, RDHH21, RZW^{+}24, RFD21, RXX^{+}21, SUS^{+}21, SCRL20, SBk^{+}24, SKW^{+}23, SSB^{+}22, SNH^{+}23, SKS^{+}23, SLK^{+}20, SH24a, SDXR22, SJL^{+}23, SB23, SSSEA20, SaT^{+}23, SMS^{+}22, SWS^{+}23, TEK^{+}23, VE21, WSL^{+}20, WXW^{+}20, WEM^{+}21, WTS^{+}21, WRC^{+}22, WLS^{+}22, WPB^{+}20, WIP^{+}24, WCH^{+}22, WLGW23, XRN^{+}23, XHL^{+}23, XWZ^{+}22, XZXY22, XWY^{+}20, XYF^{+}21, YWL^{+}21, YDM^{+}21, YZF^{+}23, YGW^{+}24, YKJ^{+}23, YZZ^{+}22, ZHTR22, ZYC^{+}23, ZWW^{+}23b, ZFF22, ZWZ^{+}22]. Interactive [ZZZ24, ZBNS21, ZPWS23, CDBM22]. Interactive-Flow [SJL^{+}23]. Intercultural [LKS22]. Interdisciplinary [SFNRZ^{+}23, SNH^{+}23]. Interface [CQ22, HLA^{+}22, HM22, LZ21, LZP^{+}20, SRKK21, STA^{+}21, YS20]. Interfaces [Ano23f, FGF^{+}21, GPR^{+}24, HAK^{+}23, SSL^{+}23, SMPJ^{+}20, SKFZ22, KHD^{+}22]. Interference [PT20]. Interior [LXF^{+}22]. Interiors [PBAG23]. Interleaving [YBR^{+}23]. Interlocking [KAS^{+}22]. Internal [MPWN22, SFL^{+}22]. International [Ano20l, Ano20o, Ano20p, Ano22p, Ano23g, Ano23n, Ano24l, CB22, Ano20m, Ano21j]. Internet [SRKK21]. Interoception [ENvBC23]. Interoceptive [HTP^{+}23]. Interoperable [HM24]. Interpersonal [AI21, LCK^{+}23]. Interplay [SZZW24]. Interpolation [JGG21, LCCZ22, SCC^{+}23]. Interpretability [CvW23, HPRC20, LqZ23, NP21, bÇ22]. Interpretable [GJC^{+}22, HM22, LGWL21, ROM^{+}23, WLZ^{+}23, ZXSR22]. Interpretation [CMK20, WWP22]. Interpreting [TAL24]. Intersection [LZJZ20]. Intersectional [BMWD20]. Intervention [LCH^{+}21]. Interventions [HPdlG20, LTC21]. Interview [CPR^{+}22, NE24]. Interweaving [SLS21]. Intrinsic [LHL^{+}20, QGL^{+}23, SLJ^{+}20]. Introducing [LVV^{+}21, MCSAL23, MB20a, MB20b, MB21a, MB22a, QCCC23, SK23a, VH23]. Introduction [BSW20, CTW21, CRZ22, CRH23, Mue23a]. Introductory [YXW^{+}23]. Intuitive [FSN20]. Intuitiveness [bÇ22]. InVADo [SBk^{+}24]. Invariant [LLC^{+}22]. Inverse [EAS^{+}23, KRK21, SDR22]. Inverse-Projections [EAS^{+}23]. Invertible [AMAS21, LCL^{+}22, YLLW24]. Investigate [SWF^{+}24]. Investigating [CPD^{+}24, HTP^{+}23, JLH24, KNL23, KKS^{+}22, RLG^{+}23, SHPE20, SPJ^{+}23, SML^{+}23, TIDQ24, VVC^{+}24, WMB23, XLZ24, ZSM21, ZPG21, DHLA22, SSS20]. Investigation [ASCR^{+}22, FWZ^{+}20, JMK^{+}22, KCWK20, MNB^{+}23, MF24, TB24]. Investigations [SGB^{+}22, ZeB^{+}21]. investment [WKMD22]. Invisible [GTH20]. Inviwo [JSS^{+}20]. InvVis [YLLW24]. Inward [HTP^{+}23]. Ionic [PSG^{+}22]. IoT [WMH^{+}23]. IoT-based [WMH^{+}23]. Iris [JJKJ20, WCX21b]. Irradiance [NI22]. IRVINE [EBJ^{+}22]. island [MHFF21]. ISMAR [Ano22l, Ano22n, Ano22m, Ano23j, Ano23l, Ano23k, FGL^{+}23, HKVZ20, IMKP21, IGMW22]. Isn’t [FW22]. Iso [JH20]. Iso-Surfaces [JH20]. Isolation [GWD^{+}24]. Isometric [CYD^{+}23, KBV24]. Isomorphic [GWL^{+}22]. Isosurface [WCTW21]. Isosurfacing [KKLS21]. IsoTrotter [PSH21]. Issue [MB20a, MB20b, MB21a, MB22a, SK23a]. Issues [YGW^{+}24]. item [CAA^{+}21]. Items [San20]. Iterative [KLB24, KMG^{+}21, LZZ^{+}20, XDBAR24].

Jam [BK22]. Jarke [Ano22b]. Joint [LKL23, LWF23, MNZ^{+}20, WCJW22, ZWL23]. Joints [XCSJ22]. Journal [Ano20m, Ano20n, Ano21j, Ano21m, Ano22l, Ano22m, Ano22p, Ano22q, Ano23j, Ano23k, FGL^{+}23, IMKP21, IGMW22, Mue20c, NAW^{+}22]. Journalism [KBM21, MEHD24]. Journey [GLL^{+}24]. Journeys [RBSN22]. Joystick [BC21]. JPEG [LCL^{+}22]. JSON [McN23]. JSON-style [McN23]. Judge [RPHJ20, RBLT^{+}22]. judgement [KMWD21]. Judgment [WBWL24]. Judgments [KKH21, WBPC23]. Jumping [NP24a]. Just [LLW^{+}22b]. Justice [RBSN22]. Juvenile [ZCM^{+}23b].

Kaleidoscopic [ZZZ24]. KD [ZWZ^{+}22]. KD-Box [ZWZ^{+}22]. KD-tree [ZWZ^{+}22]. Kernel [BKPB21, LZL^{+}23, MDJV21]. Key [SDMK22]. Keyboard [AV21, DZG^{+}23, HDFK21]. Keyboards [KNL23]. Keyframes [MSK23c]. Keyphrases [TQW^{+}24]. Keyword [CML24, WBG^{+}22]. KG4Vis [LWZ^{+}22]. KID [KBV24]. Kinetic [CLFL22, LLD^{+}21]. Kinetic-Based [LLD^{+}21]. Kineticharts [LSW^{+}22]. KiriPhys [DPC23]. Knit [WTY^{+}22]. Knitting [WTY^{+}22]. Knots [LZ21]. Knowledge [BSG^{+}20, CXD^{+}21, EJS^{+}23, KJL24, LWZ^{+}22, LqZ23, LAML23, LALG22, LHL^{+}20, PBF^{+}21, SPM24, XVF20, YQN^{+}21]. Known [CBP22, MA20]. Kori [LZK^{+}22]. Kyrix [THS^{+}21]. Kyrix-S [THS^{+}21].

L [CCL^{+}24]. lab [KHD^{+}22]. lab-based [KHD^{+}22]. Label [CCL^{+}24, JKA^{+}24, RPHJ20, YGW^{+}24, ZWP21]. Label-Based [JKA^{+}24]. Labeled [MM20, YWM^{+}20].

Labeling [GBNH21, HCL20, KKZE20, LYBP23, MPWN22, PC22, VCO^{+}23]. Labels [MPWN22, PC22]. Lack [WMMB23]. LADV [MMG^{+}21]. Lagrange [GLHQ21]. Lambertian [ZZW21]. Lambertian-World [ZZW21]. Landmark [WWW^{+}24, ZSCRB23]. Landscapes [HTBL22]. Language [CYX^{+}23, CZW^{+}20, DWB21, GSS^{+}23, HTY^{+}23, JKV^{+}22, LM21, LS22, LTL^{+}22, NSS21, RDH23, SCR^{+}23, SSL^{+}23, SCL^{+}24, SWS^{+}23, WHS^{+}23, WHJ^{+}24, YS20, HZC^{+}20]. Language-Based [WHS^{+}23]. Languages [HXHT23]. Laplacian [NdCS21, QGY^{+}22]. Large [ACT^{+}24, BHA^{+}23, BBSvL24, CB22, CSIP22, DWL^{+}22, DCS^{+}24, FBB^{+}21, FJK^{+}20, HS23, IRR^{+}22, JLX^{+}23, KKE21, KLSB22, LT20, LPJT^{+}22, LAN21, LIB24, LF23, LPP^{+}23, LXZ^{+}23, LZL^{+}23, LMD^{+}22, MCW^{+}20, NRA^{+}23, NWMC23, PFC20, PJYW21, RPD20, RPD21, RPD22, RFD21, RZH^{+}23, RAFSA23a, SCRL20, SIA^{+}23, SEAKC21, SWS^{+}23, UDH23, WMZ22, WWSS20, WYIS24, WLT^{+}24, WXS^{+}24, YWL^{+}20, YLLW24, ZYL^{+}24, ZWZ^{+}22, ZBG^{+}24, ZSS^{+}21, ZCH^{+}21, ZPF^{+}22, SNBC23]. Large-kernel [LZL^{+}23]. Large-Magnitude-Range [ZBG^{+}24]. Large-Scale [ACT^{+}24, BHA^{+}23, DWL^{+}22, DCS^{+}24, KLSB22, LPJT^{+}22, LXZ^{+}23, LMD^{+}22, MCW^{+}20, NWMC23, PJYW21, RAFSA23a, SIA^{+}23, WYIS24, WLT^{+}24, WXS^{+}24, YLLW24, ZPF^{+}22, RZH^{+}23, ZWZ^{+}22, ZCH^{+}21]. LargeNetVis [LPP^{+}23]. larger [VAWL24]. Lasso [CZY^{+}20]. Lasso-Selection [CZY^{+}20]. LassoNet [CZY^{+}20]. Late [FPG^{+}23]. Latency [BXQ^{+}22, BCN^{+}20, FE21, HWMI23, IAI^{+}23, MF24, ZP24]. Latent [CFZZ23, LS23, SLX^{+}23, SXL^{+}23]. Latent-based [SXL^{+}23]. Lateral [BVV^{+}23]. Latitude [HXL^{+}24]. Lattice [WWG21]. Law [KVM^{+}22]. Layer [BWWL22, ISKM23, KLB24, SWY^{+}22, XLY^{+}22]. Layered [DKuH20, EML^{+}23, GRi^{+}21, dBRGD22]. Layers [XXM^{+}21, LVV^{+}21]. laying [dBRGD22]. Layout [BD22, CWO^{+}24, CLW^{+}24, KM20, LYZ^{+}21, MMN^{+}22, PCZ^{+}21, QCL23, QZZ22, WZW^{+}23, XFF^{+}21, XWZ^{+}23, ZXC^{+}23, ZWP21, ZYC^{+}24b, ZCH^{+}21, dBRGD22]. Layouts [GLA^{+}24b, HSB^{+}21, LJS21, San20, SHW^{+}20, XYF^{+}21]. LDA [CAA^{+}20]. LDR [CXW^{+}23]. Leaning [BC21, HLA^{+}22, HAK^{+}23]. Leaning-Based [HLA^{+}22, HAK^{+}23]. Learned [NBE^{+}23, YXW^{+}23]. Learning [AL21, BvdPLH22, BHA^{+}23, BMWD20, CPD^{+}24, CMKK21, CWB^{+}20, CWW^{+}20, CYP^{+}20, CWW^{+}21, CCL^{+}24, CMQ21, CKQ^{+}23, CB22, CvW23, DWQW23, DCWD23, EHB^{+}23, FKM20, GKC^{+}24, GSK^{+}20, GLA^{+}24a, HTW20, HZX^{+}21, HTY^{+}23, HXL^{+}24, HDFK21, HPRC20, HMTI24, HCX^{+}21, HHB^{+}23, HJZ^{+}21, HSC^{+}22, JGG21, JLCZ22, KMH^{+}23, KBJ^{+}20, LRHA22, LRA23, LQWQ21, LM21, LBW^{+}22, LWL^{+}22b, LQS^{+}23, LHZ^{+}23, LFO23, LXH^{+}21, LGL^{+}23b, MXLM20, MMG^{+}21, MFH^{+}21, MBHE24, OWW^{+}24, PLD^{+}23, PHB^{+}22, PvSvdE^{+}24, QGY^{+}22, QGL^{+}23, QFWS22, RS23, RBLE22, RMW^{+}24, RR23, SKW^{+}23, SDR22, SCC^{+}24, SSX^{+}20, SLK^{+}20, SDXR22, SWZ^{+}23, SLSW23, SSSEA20, SRBP20, TLW^{+}21, TFE22, WJW^{+}20, WTS^{+}21, WAP^{+}21, WHSZ21, WXC^{+}21, WCWQ22, WZY^{+}22, WH23, WCX^{+}23, WWZ^{+}23a, WQP^{+}22, WBG^{+}22, WCTW21, WITW22, WPB^{+}20, WTD^{+}21, WWZ^{+}22, XEXW24, XGS^{+}23, YBOB24, ZHDX20, ZZX^{+}22, ZWH^{+}22, ZOS^{+}23, ZXSR22, ZSS^{+}21, ZJJH21, ZLVV22]. Learning-Based [BMWD20, QGL^{+}23, SDR22, WCTW21, ZWH^{+}22, CWW^{+}20]. Learning-From-Disagreement [WWZ^{+}23a]. Leaves [WWZ23b]. Led [WYIS24]. Left [NGBA^{+}20, WNC^{+}22]. Leg [KCGZ23, VAWL24]. Leg-Based [KCGZ23]. Legal [RPNP23]. LegalVis [RPNP23]. Length [FRL^{+}23]. Lens [ACPB24, BJCL21, CWSJ23, DLH^{+}22, LSZC21, RZW^{+}24, TWA22]. LensIet [BJCL21]. Lenslet [BJCL21]. Leonardo [GWF^{+}23]. Lesions [DMBK21]. Less [WAA^{+}22]. Lessons [NB24, YXW^{+}23]. Let [XHL^{+}24]. Level [CLZ^{+}24, GZRP^{+}22, JH20, JKL24, LSW^{+}23, LETF21, LS22, MDX^{+}23, NRA^{+}23, SKNŽ20, VLM^{+}23, WNV22, MLC^{+}20]. Level-of-Detail [SKNŽ20]. Level-Sets [JH20]. Leveling [FMP23]. Levels [CSM^{+}23, DV23, JSS^{+}20, SIA^{+}23]. Leveraging [EJS^{+}23, EAS^{+}23, LWC24, LLY^{+}24, OWW^{+}24, WTD^{+}21]. LF2MV [XEXW24]. LFACon [QCCC23]. LFV [LCC^{+}23]. LiberRoad [GLL^{+}24]. Library [LZC^{+}21]. LiDAR [RZH^{+}23]. LiDAR-aid [RZH^{+}23]. Lifetime [Ano22b, Ano23c, Ano24c, Kli22]. Light [ASCR^{+}22, AYA^{+}21, CJY^{+}23, FV24, GZL^{+}21, HHI24, KHI20, KPW20, KJI^{+}21, LCC^{+}23, LWK20, LWWY21, MDJV21, NI22, PPYW20, QCCC23, ŠK20, SWZ^{+}23, SII^{+}21, VM23, WGH20, XEXW24, YIIW24, ZCR21, ZGL^{+}21]. Light-Attenuation [HHI24, KHI20]. Lightcuts [SWW^{+}22, Yuk21]. LightGuider [WSL^{+}20]. Lighting [BHY^{+}23, CBP22, KKS^{+}22, LWL^{+}23b, PPYW20, TKIS24, WSL^{+}20, ZCR21, ZLX23]. Lights [GMX23, TKIS24, Yuk21]. Lightweight [XZXY22]. Like [AT23, DHF^{+}22]. Likely [RDH23]. Limb [CNC^{+}20, CNK^{+}24]. Limitations [WZZ^{+}23]. Limited [NVRS^{+}21]. Line [CZG^{+}22, GG21, KNM^{+}21, MPNF24, PTS^{+}20, RZW^{+}24, RQ21, SZC^{+}23, SLW^{+}24, WWP22, XCLF20, XPK^{+}24, YLL^{+}22b, ZWZ^{+}22]. Line-Based [PTS^{+}20, XPK^{+}24]. Line-segment-based [ZWZ^{+}22]. Lineage [HMTI24]. Linear [LCSA22, MM21, MLT^{+}24, PTM^{+}20, SSJ^{+}22, SS21, WDG^{+}20, WDN23]. Linearization [YLL^{+}22a]. Lines [HNS23, LXL21]. LineSmooth [RQ21]. Lineup [CLX^{+}23]. Link [ASA^{+}23, CCS^{+}24, RSRG23, RAP21, PCZ^{+}21, dBRGD22]. Links [RSRG23, ZSCC22]. LinSets.zip [WDN23]. Linter [CSX^{+}22]. Linting [LFMM24]. Liquid [HWC23]. Liquid-based [HWC23]. Liquids [IWT^{+}20, KLTB21]. List [Ano24a, San20]. List* [Ano21a]. Lit [YIIW24]. Lite [ZPWS23]. Literacy [BVY^{+}23, CGD^{+}24, HNGC21]. Literature [CJS^{+}22, CBHR^{+}23, KPF^{+}22, NKWW22, PCL24, PMW23]. Live [FBW21, HZZ^{+}20, LZX^{+}22]. LiveObj [FBW21]. LiveRetro [WXG^{+}24]. Livestream [WXG^{+}24]. Load [JKL24, XGS^{+}21, XGS^{+}23, YYD^{+}21]. Load-Balanced [XGS^{+}21, XGS^{+}23]. Loads [LHA^{+}21]. Lobby [LVR^{+}24]. Local [BHY^{+}23, BXZ^{+}21, BPA22, HYSL23, KKLS21, KLB24, LS23, LFO23, LYH^{+}23, NP21, SN23, TFE21]. Local-to-Global [BHY^{+}23]. Locale [BHY^{+}23]. Locale-Aware [BHY^{+}23]. Locality [YKJ^{+}23]. Locality-Sensitive [YKJ^{+}23]. Localization [DMTD22, LXZ^{+}23, RAFSA23a, ZHL^{+}20, ZWP21]. Localize [RLG^{+}23]. Localized [KGBP20, LIFD23, LI24, LGMT21, SWSK23]. Locally [MM21, YLJ^{+}22]. Located [JLH24, LHC^{+}21, YEP^{+}22]. Location [ISBP22, MWUP22, NIM^{+}21]. Locations [San20]. Locomotion [AZA^{+}23, AMF20, BC21, CCP^{+}21, CQ22, GMTD23, HAK^{+}23, KJL24, LGL^{+}23a, NVRS^{+}21, PMW23, SACB^{+}23, ZOF^{+}23, KHD^{+}22]. Log [LDB^{+}21]. Log-Rectilinear [LDB^{+}21]. Lollipops [RRG23]. LoM [LVV^{+}21]. Long [DHM^{+}22, SMC^{+}21, WHSZ21]. Long-Horizon [WHSZ21]. Long-Range [SMC^{+}21]. Long-Term [DHM^{+}22].

Longitudinal [FNB^{+}22, LKS22, WLS^{+}22]. Look [JBS^{+}22]. Look-Up [JBS^{+}22]. Looking [LSZC21]. Loon [LPJT^{+}22]. Loop [GM23, LDT^{+}21, SHM23, SH24b, ZCM^{+}23b, ZXSR22]. Loretta [Ano22c]. loss [WKMD22]. Lossy [YLGW24]. Lot [MGWM22]. Lotse [SCEA23]. Loupe [QSUK22, QSUK22]. Low [CWS^{+}20, CXXZ21, EHA^{+}23, GBL^{+}22, HWMI23, LSL^{+}22, LCW^{+}23b, MF24, ZZS^{+}23, ZP24]. Low-Cost [GBL^{+}22, ZZS^{+}23, CXXZ21]. Low-Dimensional [EHA^{+}23]. Low-Latency [HWMI23, MF24, ZP24]. Low-Rank [CWS^{+}20]. LSTM [FiMH21, QLFG22]. Ludicrous [DMTD22]. Lumos [NCWE22]. Lung [WZH20]. Lyapunov [RGG20a]. Lyra [ZBNS21].

M2Lens [WHJ^{+}22]. Mach [VSBY22]. Mach-RT [VSBY22]. Machine [BHA^{+}23, CMQ21, CvW23, FWZ^{+}20, FNB^{+}22, GKC^{+}24, HMTI24, JGG21, LM21, LTL^{+}22, MXLM20, RMW^{+}24, SSSEA20, WAP^{+}21, WXC^{+}21, WCWQ22, WPB^{+}20, WTY^{+}22, WGH^{+}24, YBOB24, ZOS^{+}23, ZFF22, ZLVV22]. Machine-Guided [WGH^{+}24]. Made [ZCZ22]. Magazine [LTC21]. Magazine-Style [LTC21]. Magic [BBRE24, MSRJ20]. MagLoc [LXZ^{+}23]. MagLoc-AR [LXZ^{+}23]. Magnetic [HSC^{+}22, LXZ^{+}23]. Magnetic-Based [LXZ^{+}23]. Magnified [QSUK22]. Magnitude [ZBG^{+}24]. Mail [NIK24]. Maintenance [GEU^{+}22]. Major [TBW^{+}23]. Makers [DZTF22]. Makes [SWT^{+}21]. Making [AL20, DS22, DIPJ22, HM22, LLP^{+}23, NCS^{+}21, OCW^{+}24, PRJ^{+}23, RVB^{+}22, SWF^{+}24, XLL^{+}22, ZLVV22]. Man [YHC^{+}22]. Management [BS21, LWSY20, ZLL^{+}20]. Managing [dSBdO^{+}24]. ManhattanFusion [YFS22]. Manifold [HYSL23, WHSZ21]. Manifold-Constrained [HYSL23].

Manifolds [KLB24]. Manipulating [PIS20]. Manipulation [CFZZ23, DHLA22, GLK^{+}23, GWL^{+}22, GSH21, KKF20, SHPE20, SMNK21, WSN21, WCH^{+}23, WGO20, WGO22]. ManiVault [VKT^{+}24]. ManuKnowVis [EJS^{+}23]. Many [LZM20, NI22, VSBY22, Yuk21, ZYM^{+}24]. Many-Light [NI22]. Map [CLCY20, CBP22, FLZ^{+}21, GBNH21, HSS^{+}20, JGG21, TSH21, YDM^{+}21, CLCY20]. Mapping [CWH^{+}23, EIB23, HIK^{+}23, KIPS21, KLSB22, LF23, MSWI22, NLPW22, PJYW21, RCD^{+}23, SWWY21, TKIS24, TIHS20, YIIW24]. Mappings [WPTG24]. Maps [AR22, AMY^{+}22, BXQ^{+}22, CCPM23, CZMR21, CWB^{+}20, FK20, GBNH21, HPAB23, IRR^{+}22, JWKN21, KKE21, KW23, LFMM24, MSS21, PM23, RPG23b]. Marching [ZYP^{+}24]. Marjorie [SESH24]. mark [MHFF21]. Marker [DMTD22, YHD21]. Marker-based [DMTD22]. Markers [MSWI22, TIHS20]. Markov [KAS^{+}21, ŠK20, SLR20b]. Marks [FFB24]. MARVisT [CSW^{+}20]. Masked360 [LCW^{+}23b]. Masking [SWSK23]. Mass [CPD^{+}24, QHL^{+}20, SVK^{+}21, YB20, ZYL^{+}24]. Mass-Driven [QHL^{+}20]. Massive [CYP^{+}20, SPJ^{+}23, XWPG^{+}24]. MAT [LLL^{+}22]. Match [LAC^{+}24]. Matched [CZT^{+}21, DPM24, WH22a]. Matches [PJHY20, WWC^{+}21]. Matching [GSM^{+}22, GJC^{+}22, PJYW21, SES20]. Material [Off20, PPYW20, SDR22, SOL^{+}22]. Materials [DS20, GLHQ21, ZYP^{+}24, VGK^{+}22]. matExplorer [PSG^{+}22]. Mathematical [LZ21]. Matrices [KKhCM23]. Matrix [ASA^{+}23, BSP20, CLZ^{+}24, HBS^{+}21, LSL^{+}22, NP21, vBMS22]. Matrix-Based [ASA^{+}23]. Matrix-Visualization [NP21]. Matter [Ano23i]. Matters [MYS^{+}22].

Maximise [RR23]. Maximization [ADL^{+}22]. May [KBPR22, LLSM24]. Maze [LCK^{+}23]. MD [JK23]. MD-Cave [JK23]. Me [CAGM22, KPL23, VVR^{+}23a]. Mean [HWS22]. Meaning [LVV^{+}21]. Measure [CDX^{+}20, JQL^{+}24, LWC22, RAC22]. Measured [KVGM20]. Measurement [LSG24, MWR^{+}22]. Measures [NF20, PBBH20]. Measuring [ABL^{+}22, CCPM23, CZC^{+}20, FE21, HIDI23, JKJ^{+}22, LCK^{+}23, TARB23, PGL^{+}20, ENvBC23]. Mechanics [MXF^{+}21]. Mechanics-Aware [MXF^{+}21]. Mechanisms [DWB21, NBE^{+}21]. MedChemLens [SNH^{+}23]. Media [CLCY20, GWW^{+}21, ISBP22, JSA^{+}20, JKU^{+}22, KKT^{+}22, LWWH20, RBLE22, SKR^{+}24, VM23, WGH20]. Medial [LLL^{+}22]. Medical [HSJ^{+}20, OWW^{+}24, PNB^{+}21, SBe^{+}21]. Medicinal [SNH^{+}23]. Medley [PSS23]. Meeting [ABGG21]. Members [Ano22n, Ano22r, Ano23p, Ano24n, ROM^{+}23, Ano22m, Ano23l, Ano23k]. Memorability [bÇ22]. Memory [CQHP22, DST^{+}23, MMS^{+}23, PCQ^{+}20, PFN22, SOL^{+}22, WMZ22, YQN^{+}21]. Mental [JCZ^{+}24, VBV^{+}23]. Merge [BTL23, KW23, LWW^{+}24, PVDT22, PVT23, SMKN20, SN23, WG21, WPTG24, YWM^{+}20, YBR^{+}23]. Merging [TYW^{+}22]. Merging-and-Splitting [TYW^{+}22]. Mesh [AFB22, DLP^{+}23, DHF^{+}22, GLY^{+}21, LLZ^{+}21, LSW^{+}23, LLW^{+}22a, Liv21, MSY22, MWUP22, NWW21, PSL23, QLFG22, SXW^{+}22, SLR21, TYPC20, YGT^{+}23, ZYP^{+}24, ZLW^{+}21b, ZCZ^{+}21, ZZC^{+}22]. Meshes [BHM^{+}22, CDZ^{+}23, KKV22, MNPP23, NDF^{+}21, QGY^{+}22, ZH20]. Meshing [GLX^{+}21]. Meshless [Pat22]. Mesoscale [NSK^{+}21]. Message [Ano21k, BPQW23, FGL^{+}23, FG22, HKVZ20, IPPZ24, IMKP21, IGMW22, Mue20b, MB20c, Mue21a, MB21b, Mue22b, MB22b, Mue23b, NAW^{+}22, PK21, She23a, SK23b, She24, SK24, SES23]. Messages [FFB24]. Meta [AYGR22, HH20, Kel23, XEXW24, YCH^{+}22]. Meta-Analysis [Kel23]. Meta-PU [YCH^{+}22]. Meta-Strategy [AYGR22]. Meta-View [XEXW24]. MeTACAST [ZIX^{+}24]. Metadata [KKJ^{+}21]. Metagenomic [DMMF21]. MetaGlyph [YSD^{+}23]. Metameric [dAWi^{+}23]. Metaphor [BTHL23, BNRB21, CLCY20, WWP22]. Metaphoric [YSD^{+}23]. Metaverse [ZC23]. Meteorology [HXL^{+}24]. Method [CAGM22, DGD^{+}23, FCS^{+}20, GSM^{+}22, GLA^{+}24b, HTP^{+}23, LF23, LFR^{+}21, LLCH22, SHPE20, SLX^{+}23, TEK^{+}23]. Methods [ACT^{+}24, CWR21, DZG^{+}23, ENXS21, GBNH21, KPF^{+}22, KKEG20, MCQ23, MTE^{+}20, MBB20, QLC^{+}24, RRK^{+}22, ŠK20, WAV^{+}21, YXX^{+}21, ZLG^{+}21b]. Metric [FHTB23, LIDM20, WSL^{+}24]. Metric-Driven [FHTB23]. Metrics [CMKK21, KMM24, TEK^{+}23]. Metrics-Based [KMM24, TEK^{+}23]. MetricsVis [ZKS^{+}20]. Metro [JWKN21]. Metropolitan [SHL^{+}21]. Metropolitan-Scale [SHL^{+}21]. MetroSets [JWKN21]. MFVs [PFCB23]. Michelle [Ano22d]. Micro [SVK^{+}21, XWZB21]. Micro-Appearance [XWZB21]. Micro-Environments [SVK^{+}21]. Micrography [HYF^{+}20]. microlens [RSAA20]. Microsaccades [DKuH20]. Microscopy [GBM^{+}22, LPJT^{+}22]. Microservice [RHHH20]. Microtubule [KVGM20]. Mid [BS22, DZG^{+}23, FHR^{+}21, HDFK21, KG24, SDK23a, WBG^{+}23]. Mid-Air [BS22, DZG^{+}23, FHR^{+}21, HDFK21, KG24, SDK23a, WBG^{+}23]. Midair [AV21]. Middle [HXL^{+}24, PLD^{+}23]. Middle-Latitude [HXL^{+}24]. Midterms [YCM^{+}24]. Million [NGW^{+}24]. Mind [PSH20]. Miner [WWC^{+}21]. MineTime [ABGG21]. Minimization [TYPC20, vdRBE22]. Minimizing [BZW^{+}21]. Mining [FWM^{+}24, LWW^{+}21, TSHI22, WWC^{+}21, WLGW23, XMK^{+}22]. MiningVis [TSHI22]. Minkowski [CWR21]. Minority [ZJC^{+}21]. Minutes [LLT24]. Mirror [CNK^{+}24, YZEN22]. MissBiN [ZSCC22]. Missing [FW22, SGH^{+}23, ZSCC22]. Mitigate [WNC^{+}22]. Mitigating [JLM^{+}21, YGW^{+}24]. Mitigation [KLSB22, NDLW20, ZOF^{+}23]. Mitral [EEL^{+}20]. Mixed [Ano23g, CZMR21, CIA24, CLH^{+}23, CFDN24, CWH^{+}22, EMM^{+}22, EML^{+}23, FYC^{+}23, GWL^{+}22, HPdlG20, HLH^{+}23, KKW23, iKYOW23, KRHH24, LNB^{+}21, LFCH24, LZZ^{+}24, MPWN22, PPYW20, PBF^{+}21, TLBB23, VVC^{+}24, WBLW23, YIIW24, YeSiK^{+}23, ZYR^{+}20]. Mixed-Initiative [CWH^{+}22, KRHH24, PBF^{+}21]. Mixed-Integer [HLH^{+}23]. Mixture [ATAS21, LME^{+}23]. MK1 [WLC^{+}23]. ML4VIS [WCWQ22]. Mobile [BSG^{+}20, BGS^{+}22, BLIC20, CSW^{+}20, CBB^{+}21, FBW21, WWS20, WTD^{+}21, XJZ^{+}21, YZJ^{+}20, YKF22, ZYL^{+}24, ZG20]. Mobile3DRecon [YZJ^{+}20]. Mobile3DScanner [XJZ^{+}21]. MobileVisFixer [WTD^{+}21]. Mobility [WLS^{+}22, YZF^{+}23]. Modal [CML24, LLSM24]. Modalities [SPW^{+}22]. Modality [ENvBC23, FMP23]. Modality-Dependent [ENvBC23]. Mode [KKEW23, NWMC23, QRZZ21, WSN22, ZZ23]. Model [CLH^{+}23, CWB^{+}20, CFZ^{+}23, CGT^{+}24, DPD^{+}24, DV23, EAKC^{+}20, FWZZ20, FAS^{+}21, GKC^{+}24, GSL^{+}24, GWW^{+}21, GLL^{+}22, HRS^{+}22, HCM^{+}22, HJZ^{+}21, KGQ^{+}24, KM20, KKhCM23, LTLB22, LSL^{+}23, LZZ^{+}20, LS22, MTW^{+}20, MSA^{+}22, MSB^{+}22, MXC^{+}20, MGO21, PRKM20, RBRG21, SCR^{+}23, SHM23, SC22, SXL^{+}23, SMSK22, UWF^{+}23, VGK^{+}22, WHC^{+}23, WWZ^{+}23a, XRN^{+}23, XHL^{+}24, XWH^{+}23, YBOB24, ZMK^{+}20, ZLY22, ZOS^{+}23, ZLX23, ZCZ^{+}21, leáB^{+}21, dBRGD22]. Model-Agnostic [HRS^{+}22]. Model-Free [LTLB22, SMSK22]. Modeling [BBVS^{+}24, BBM^{+}21, COFJ23, GWC^{+}23, GXY^{+}20, HMGO23, Hei21, IMQ^{+}20, JYZW20, KWH22, KDEP21, KMG^{+}21, KSB^{+}22, LMF^{+}24, LWL^{+}23a, LZZ^{+}20, LZX^{+}21, LLW^{+}22b, MXF^{+}21, NDP^{+}21, NSK^{+}21, OWW^{+}24, QR21, RBK^{+}21, RPHJ20, RS21, RBF^{+}23, SZF^{+}21, SJK^{+}23, SMS^{+}22, TSS22, WHSZ21, XWZB21, ZZW^{+}20, ZLQH21, ZYC^{+}23, ZLZ^{+}23b]. Modelling [PSH21]. Models [AMS^{+}21, AJSP23, ACT^{+}24, CMKK21, CMQ21, CLD^{+}22, DWOB20, DWB21, GZM^{+}21, GSK^{+}20, JZHA22, JLP^{+}23, KAS^{+}21, LME^{+}23, LWY^{+}22, LWM^{+}20, MGO21, NSK^{+}21, ROM^{+}23, RRG23, RLG^{+}23, SS24, SIA^{+}23, SKNŽ20, SWS^{+}23, TFE21, WZH20, WAP^{+}21, WHJ^{+}22, WHJ^{+}24, WPB^{+}20, XMK^{+}22, ZGX^{+}22, ZXSR22, iIK^{+}21]. Moderation [TWW^{+}22]. Modifiable [ZLL^{+}21]. Modified [MMK^{+}23]. Modular [YITS23, dBRGD22]. Modulated [ILZ^{+}21]. Modulator [KPW20, ZOF^{+}23]. Module [SS21]. Molecular [HZP^{+}24, MCFKF24, PBBH20, SBk^{+}24, SFL^{+}22, UWF^{+}23, iIK^{+}21, ieSM^{+}23]. Molecumentary [ieSM^{+}23]. MolSieve [HZP^{+}24]. Moments [RPD22]. Monitoring [LKJ^{+}20, SFS^{+}22, TZT^{+}22, ZLL^{+}20, MNK23]. Monochrome [DLH^{+}22]. Monochrome-Color [DLH^{+}22]. Monocular [APP^{+}22, LHZ^{+}23, MWR^{+}22, SWY^{+}22, YITS23, YZJ^{+}20]. Monotonic [LGWL21]. Monte [CS23, ŠK20, RGG20a]. Monte-Carlo [CS23]. Mood [SHS^{+}22, YC23]. Mood-Driven [SHS^{+}22]. Moral [SWF^{+}24]. MoReVis [VFP24].

Morphing [LLC^{+}20]. Morphological [CLL^{+}23, SKW^{+}23]. Morse [AMY^{+}22, GVT24, MLT^{+}24, SPN23]. Mosaic [HM24]. MosaicSets [RWB^{+}23]. Motifs [TWC^{+}24]. Motion [AYA^{+}23, CGAG20, CNC^{+}20, EKC^{+}23, HWMI23, HSC^{+}22, LGL^{+}23a, LSE20, NGW^{+}24, NDLW20, PALW20, RZH^{+}23, SJL^{+}23, WCX21a, WYS^{+}22, WBS21, WGS^{+}24, XFF^{+}21, YBVI22, YVBI24, YHC^{+}24, ZZ23, ZLZ^{+}23a, ZOF^{+}23, MCSAL23]. Motion-adaptive [ZZ23]. Motion-Based [ZOF^{+}23]. Motion-Blur [LSE20]. Motion-to-Photon [HWMI23]. Motions [PGS21, WHSZ21, YHC^{+}22]. Motor [KZD^{+}23, KMH^{+}23, KAL^{+}23, KKF20, VAWL24]. Mounted [CXXZ21, CCS^{+}21, FGF^{+}21, KCA^{+}21, Kel23, LSG24, MGWK^{+}22, NMS^{+}23, NIM^{+}21, QSUK22, SPNG23, WSL^{+}24, WBA^{+}23, YLJ^{+}22, ZWP^{+}22, IAI^{+}23, TWA22, ZHKY23]. Movable [HZS^{+}22]. Movement [BvdPLH22, CNC^{+}20, DGB^{+}22, HTBL22, LZZ^{+}24, SCHE23, TSS22, WHY^{+}23, YCC^{+}21, ZSZ^{+}24]. Movement-Induced [TSS22]. Movements [HBLW23, LH23, ONH21, WDX^{+}23]. Moving [SCLK21, SMYF22, VFP24, WBI20]. MR [LZP^{+}20]. MulayCap [SWY^{+}22]. Multi [BDL^{+}21, BWWL22, CWS^{+}20, CGZ^{+}20, CAA^{+}21, CMSK23, DGD^{+}23, DIPJ22, FJK^{+}20, GLHQ21, GWC^{+}23, HHK^{+}24, HHKN21, HSV^{+}20, HZJ^{+}24, IGM24, ISKM23, JH20, KBM21, KRZ^{+}20, KVGM20, KGB22, KBJ^{+}20, KSHW22, KWFK20, KLSB22, LG23, LQS^{+}23, LETF21, LRZ^{+}23, LWL^{+}20, LMD^{+}22, LZP^{+}20, LGWL21, MSAM^{+}22, NMC21, NLPW22, RAC22, RXX^{+}21, SA22, SYY^{+}24, SHT^{+}22, SYX^{+}22, SWY^{+}22, TGM21, VLM^{+}23, WFK24, WNV22, WXS^{+}24, XVW^{+}21, XWZB21, XQXL23, XLW^{+}24, YLS^{+}23, YLL^{+}20, YWB20, ZSL21, ZHH22, LLSM24].

Multi-Agent [RXX^{+}21]. Multi-Attribute [CMSK23, LRZ^{+}23, LGWL21]. Multi-Body [FJK^{+}20]. Multi-Channel [KBJ^{+}20, LZP^{+}20]. Multi-class [CGZ^{+}20, HSV^{+}20, LWL^{+}20]. Multi-Core [ZSL21]. Multi-Criteria [BDL^{+}21]. Multi-Cylinder [LMD^{+}22]. Multi-Grained [SYY^{+}24]. Multi-Grid [WXS^{+}24]. Multi-item [CAA^{+}21]. Multi-Layer [BWWL22, ISKM23, SWY^{+}22]. Multi-Level [LETF21, VLM^{+}23, WNV22]. Multi-Modal [LLSM24]. Multi-Objective [DIPJ22, HZJ^{+}24]. Multi-Parameter [KSHW22]. Multi-Patch [CWS^{+}20]. Multi-Person [YLS^{+}23]. Multi-Perspective [HHKN21]. Multi-Projection [KLSB22, NLPW22]. Multi-Projector [IGM24, TGM21]. Multi-Rate [YWB20]. Multi-Resolution [RAC22]. Multi-Scale [GLHQ21, KRZ^{+}20, KVGM20, WFK24, XWZB21, XQXL23, NMC21]. Multi-Sensor [GWC^{+}23]. Multi-Sensors [YLL^{+}20]. Multi-sensory [MSAM^{+}22]. Multi-step [XVW^{+}21]. Multi-Style [SYX^{+}22]. Multi-Table [KBM21]. Multi-User [XLW^{+}24, DGD^{+}23, ZHH22]. Multi-Variate [JH20, SHT^{+}22]. Multi-View [KGB22, LG23, SA22, SHT^{+}22, HSV^{+}20, LQS^{+}23]. Multi-Volume [HHK^{+}24]. Multi-Window [KWFK20]. Multiclass [LYL^{+}23]. Multicriteria [ADD^{+}22, NSM^{+}22]. Multidimensional [BRLP20, BXZ^{+}21, ESB^{+}24, FCS^{+}20, HKW23, JTT^{+}23, JKJ^{+}22]. Multidisciplinary [GF24]. Multifocal [KIPS21]. Multilayer [CSIP22, FPK^{+}24, ZSLL20]. Multilayer-Neighborhood [ZSLL20]. Multilevel [MSM^{+}22, TWT^{+}22]. Multilinear [BHM^{+}22]. Multimodal [HHM^{+}24, JCZ^{+}24, KKW23, LIB24, LPL^{+}24, LFG^{+}23, OWW^{+}24, SSS20, SLS21, TWW^{+}22, WHJ^{+}22, WGO20, WGO22, WQ20]. Multiphase [LLD^{+}21]. Multiple [CZL^{+}21, GTHC20, GXY^{+}20, HJZ^{+}21, LZD^{+}20, LCC^{+}23, MBS^{+}21, PFCB23, QBW^{+}20, RHLC22, SHOP23, WGH20, WWC^{+}21, WCJW22, WZW^{+}23, XF21, ZYP^{+}24]. Multiple-Fluid [RHLC22]. Multiple-View [CZL^{+}21, WZW^{+}23]. Multiples [BLIC20, LZC^{+}21]. Multiplexed [WKN^{+}23]. Multipole [HNS23]. Multiresolution [AMAS21]. Multiscale [BKPB21, CSJ^{+}21, CJS^{+}22, FJK^{+}20, HiM^{+}22, HCL20, KSB^{+}22, LBB^{+}20, YGT^{+}23]. MultiSegVA [MBS^{+}21]. Multisensory [CSM^{+}23, JKL24, MGM^{+}22, TME^{+}22, WMB23]. Multithread [SSB^{+}22]. Multivariate [AJVS^{+}23, FSS^{+}21, HZX^{+}21, HBS^{+}21, JKW^{+}22, KLKE21, MM20, NP24a, RPD21, RGDG23, WLGW23, XCK20, YKJ^{+}23]. Multiverse [LKAH21]. MultiVision [WWZ^{+}22]. Muscle [iKYOW23, ZCM^{+}23b]. MUSE [CSL^{+}23]. Museum [MMF20, RBLE22]. Museums [DMMF21]. Music [AYA^{+}23, CFZ^{+}23, KAL^{+}23, PCL24, YC23]. Music-Driven [AYA^{+}23, CFZ^{+}23]. Musical [CSL^{+}23]. MusiKeys [KG24]. MVNet [SHT^{+}22]. My [DGB^{+}22, ENvBC23, GKC^{+}24, LCYQ24, CFL^{+}23]. myopic [WKMD22]. Mystique [CLW^{+}24].

Nano [NBE^{+}23]. Nano-Ötzi [NBE^{+}23]. Nanophotonic [SZH^{+}20]. Nanostructures [KSB^{+}22]. Nanotilus [AeSL^{+}23]. Narrated [ieSM^{+}23]. Narration [SZZW24]. Narration-Animation [SZZW24]. Narrative [HNGC21, LTC21, eSYKW23]. Narrative-Focused [HNGC21]. Narratives [CSC^{+}21, MBS23, ZXC^{+}23]. Narrow [MTE^{+}20, TME^{+}22]. NAS [TXM23]. NAS-Navigator [TXM23]. NASA [CQHP22]. NASA-TLX [CQHP22]. National [Ano22c, TLW^{+}23]. Natural [CYX^{+}23, CBP22, CZW^{+}20, HTY^{+}23, HZC^{+}20, HXHT23, JSA^{+}20, LS22, LZP^{+}20, LTL^{+}22, NSS21, SSL^{+}23, SCL^{+}24, WHS^{+}23, WHJ^{+}24, XHFZ24, YS20]. Natural-language-based [HZC^{+}20]. Nature [APBB24, BWWL22]. NaviBoard [NVRS^{+}21]. NaviChair [NVRS^{+}21]. Navigate [LY23]. Navigating [CMF^{+}22, SBT^{+}23, SNBC23]. Navigation [BGB^{+}22, BBM^{+}21, ENM24, LAN21, LBB^{+}20, PMN^{+}23, WBF20, WF21, YCB^{+}21]. Navigator [ENM24, TXM23]. Near [AMK^{+}21, HWMI23, NSG^{+}20, SPWW^{+}24, SES20, VVR^{+}23a, VVC^{+}24, ZSZ^{+}24, RSAA20]. Near-Eye [AMK^{+}21, EML^{+}23, HWMI23, SPWW^{+}24, ZSZ^{+}24, RSAA20]. Near-Field [SES20, VVR^{+}23a, VVC^{+}24]. Near-Wall [NSG^{+}20]. Nearest [JSF20]. Nebula [CSC^{+}22]. Neck [AT23, LZX^{+}21]. Neck-Like [AT23]. Need [DHLA22, ZC23]. Needle [HSJ^{+}20]. Needs [DZTF22]. Negative [HMK^{+}20, RMW^{+}24]. Neighbor [LYH^{+}23]. Neighborhood [WKN^{+}23, ZSLL20]. Neighbors [JSF20]. NeRF [DHY^{+}22, QLC^{+}24, YZP^{+}23]. NeRF-NQA [QLC^{+}24]. NerfCap [WPZ^{+}23]. NeRFPlayer [SCL^{+}23b]. Nerve [SEK^{+}24]. Nested [LGW^{+}20, WHC^{+}23]. Net [GZM^{+}21, CCS^{+}24, FZZX22, GZM^{+}21, LLZ^{+}21, LHL^{+}20, MDL^{+}23, WYZ^{+}21]. Net2Vis [BvOR21]. Netgrams [XWH^{+}23]. NetHOPs [ZAH22]. Network [ASA^{+}23, ASSB^{+}23, BFY^{+}24, DLH^{+}22, DHF^{+}22, FZZX22, FPK^{+}24, GWW^{+}21, HW23, HLW^{+}20, JZCZ20, KIS22, LLT24, LLZ^{+}21, LHL^{+}22, LZL^{+}23, LSW^{+}20, MDH^{+}23, MDL^{+}23, PSY^{+}20, PBF^{+}21, SSS20, SXW^{+}22, SHT^{+}22, SLR20b, SLSW23, SEAKC21, TXM23, WZH20, WWW^{+}24, WYIS24, WCZ^{+}23, YCH^{+}22, YYD^{+}21, YFM^{+}23, ZCX^{+}24, ZLW21a, ZAH22, ZPG21, dBRGD22].

Network-Based [LSW^{+}20]. Network-on-Chip [WYIS24]. Networking [FGS^{+}22]. Networks [BWM23, BWM24, CLS^{+}21, CSCM21, CSIP22, DAB^{+}23, FiMH21, Hei21, HGO21, HMKB23, JHS^{+}21, JWW^{+}23, JJHS^{+}22, LLT24, LWS^{+}21, LPP^{+}23, LCK^{+}21, LWBM22, MSBV^{+}23, PNKC21, PDD^{+}22, PSL23, WYC^{+}20, WTS^{+}21, WCX21a, WYZ^{+}21, WZC^{+}23, WZD^{+}21, WXS^{+}24, XZXY22, XTYL20, XZKM22, ZSCC22, ZCSS23, ZSS^{+}21, ZJJH21]. Neural [BWM23, CLS^{+}21, CPCS20, DAB^{+}23, DHY^{+}22, EIB23, FV24, GMX23, GWW^{+}21, HW23, HLW^{+}20, HMKB23, JGG21, JHS^{+}21, JZCZ20, JWW^{+}23, JYF^{+}20, KIS22, KGX^{+}23, KLKE21, LLT24, LWS^{+}21, LZY^{+}23, LLDW24, LXH^{+}21, LWBM22, LSW^{+}20, LTL^{+}22, LCCZ22, PNKC21, PDD^{+}22, QLC^{+}24, RPG23a, SXW^{+}22, SCL^{+}23b, TXM23, WYC^{+}20, WTS^{+}21, WCX21a, WPZ^{+}23, WCZ^{+}23, XZXY22, XZKM22, YLTL23, ZWXM23, ZLZ^{+}23a, ZLZ^{+}23b, ZCSS23, ZJJH21]. NeuRegenerate [BMA^{+}23]. Neurites [GBM^{+}22]. NeuroCartography [PDD^{+}22]. NeuroConstruct [GBM^{+}22]. Neurodegeneration [BMA^{+}23]. Neurodegenerative [XNF^{+}23]. Neuron [MUM^{+}21]. Neuronal [TWC^{+}24]. Neurophysiological [LPG^{+}22]. Newtonian [LGH^{+}24]. NIID [LHL^{+}20]. NIID-Net [LHL^{+}20]. NL2Color [SCL^{+}24]. NL4DV [NSS21]. NLP [LWY^{+}22]. NNVA [HLW^{+}20]. NNWarp [LSW^{+}20]. No [BCC^{+}20, MHFF21, McN23, QCCC23, QLC^{+}24]. No-Reference [QCCC23, QLC^{+}24]. Node [ASA^{+}23, CCS^{+}24, JdJTC24, PCZ^{+}21, RSRG23, RAP21, XTYL20, dBRGD22]. Node-Link [ASA^{+}23, CCS^{+}24, RSRG23, RAP21, PCZ^{+}21, dBRGD22]. Noise [AJSP23, CLS^{+}21, CBW23, HGO21, RBRG21, RWJ^{+}23, SaT^{+}23]. Noisy [WYZ^{+}21, WCZ^{+}23]. Non [APHD24, DGD^{+}23, GWL^{+}22, GTL^{+}23, KNL23, KJI^{+}21, LGH^{+}24, MTE^{+}20, San20, SML^{+}23, SSJ^{+}22, SS21, XZS^{+}23, YWL^{+}20, ZZZ24]. Non-Dominant [XZS^{+}23]. Non-Epipolar [KJI^{+}21]. Non-Euclidean [ZZZ24]. Non-Experts [SML^{+}23]. Non-forward [DGD^{+}23]. Non-hierarchical [YWL^{+}20]. Non-Isomorphic [GWL^{+}22]. Non-Linear [SSJ^{+}22, SS21]. Non-List [San20]. Non-Newtonian [LGH^{+}24]. Non-Stationary [KNL23]. Non-Verbal [GTL^{+}23]. Non-Visible [APHD24]. Non-Visual [MTE^{+}20]. Nonlinear [LSW^{+}20, RRG23, SIA^{+}23]. Nonparametric [AMS^{+}21, AJSP23]. Nonuniform [TWT^{+}22]. Normal [LLZ^{+}21, LHL^{+}20, WFC21, YXS^{+}23, ZLW^{+}21b]. Normalization [WBS21]. Normalizing [MDH^{+}23]. Northeastern [Ano22d]. Notations [KMM24]. Note [Mue20a, Mue21b, QCL23, She23b]. Notebooks [LLY^{+}24]. notes [SNBC23]. notice [Ano21e]. Noticeable [LLW^{+}22b]. Novel [BBSvL24, GXH^{+}24, HLA^{+}22, LZZ^{+}24, SPJ^{+}23, WHBC23, XZWY20, YWL^{+}20]. NQA [QLC^{+}24]. Numbers [KHL21]. Numerous [KNAR^{+}22].

OA [HLH^{+}23]. OA-Style [HLH^{+}23]. Object [BVV^{+}22, BVV^{+}23, CWW^{+}22, CGT^{+}24, FBW21, GLK^{+}23, GSH21, HZQ22, JGH^{+}24, KKF20, LTLB22, LZZ^{+}20, MLC^{+}20, PPYW20, PJYW21, TNJ^{+}22, WLZ^{+}23, WGO20, WGO22, XJZ^{+}21, YB20, YGH^{+}23, ZHL^{+}20]. Object-level [MLC^{+}20]. Objective [DPR^{+}20, DIPJ22, GT20, HZJ^{+}24, NDF^{+}21, PCQ^{+}20, RMB^{+}21, ZHTR22]. Objectives [LRHA22, LRA23]. Objects [CLL^{+}20, CZGF21, FZZX22, GKN^{+}23, HZS^{+}22, HB20, KKS^{+}22, KBPR22, LYBP23, SLR20b, TIHS20, WTY^{+}22, YKF22, YPW23, ZHL^{+}21a, ZZX^{+}22, ZJS^{+}22, dPLM21]. Oblivious [QFWS22]. Observable [SUB^{+}22, ZHTR22]. Observation [SACB^{+}23]. Observations [WDSM24]. Observed [RZW^{+}24]. Observer [RMB^{+}21]. Observer-Relative [RMB^{+}21]. Obstacle [ZCZ^{+}23a]. Obstacle-Ridden [ZCZ^{+}23a]. Obstacles [ZA21]. Obstetrical [CNC^{+}20]. OBTracker [WDX^{+}23]. Occluded [YLF^{+}20, YZN^{+}20]. Occluding [XQXL23]. Occlusion [ER21, FRL^{+}23, GG21, IRR^{+}22, KPW20, MA23, PTX^{+}22, SMC^{+}21, WH22a, WWZP22, ZHKY23, ZLX23]. Occlusion-Capable [WH22a, KPW20, ZHKY23]. Occlusions [LFG^{+}23]. Ocean [SXW^{+}22]. OctoPocus [FHR^{+}21]. Octree [HHK^{+}24]. Ocular [CWSJ23, MF24]. OD [LLC^{+}20]. Odometry [BXQ^{+}22, CBB^{+}21]. Off [CBB23, EML^{+}23, KKV22, WDX^{+}23, YLF^{+}20, ZHKY23, BER^{+}23, JJKJ20, ZGL^{+}21]. Off-Axis [EML^{+}23]. Off-ball [WDX^{+}23]. Off-Screen [YLF^{+}20]. Off-the-Shelf [ZHKY23]. Off-Vertex [KKV22]. offs [PFCB23]. Offsets [CPD20]. OldVisOnline [ZJX^{+}24]. Omega [CNAA^{+}22]. Omega-Test [CNAA^{+}22]. omics [SIL^{+}21]. omics-data [SIL^{+}21]. Omnidirectional [GTH^{+}22, MLC^{+}20, MP21, PBAG23, SMYF22, WLC^{+}23, WSL^{+}24]. Omnioculars [LCY^{+}23]. On-Demand [SCRL20]. On-the-Fly [WZH20]. On-Tube [RGDG23]. Once [CSC^{+}21]. Oncological [FWM^{+}24]. One [BKN^{+}22, BBH24, FGF^{+}21, TXM23, YHC^{+}22, ZCZ23b]. One-Handed [FGF^{+}21]. One-Man-Crowd [YHC^{+}22]. One-Shot [TXM23]. One-Step [ZCZ23b]. Online [CYP^{+}20, CSCM21, HHM^{+}24, KIS22, XJZ^{+}21, XLW^{+}24, YFS22, ZWZ^{+}23]. Only [MWR^{+}22]. onto [YGH^{+}23]. Ontologies [YWL^{+}20]. OntoPlot [YWL^{+}20].

OoDAnalyzer [CYL^{+}21]. Opacity [RGG20b]. Open [CYP^{+}20, JGH^{+}24, NGW^{+}24, PWB21, PGL^{+}20]. open-cell [PGL^{+}20]. Open-World [JGH^{+}24]. Opening [HFT^{+}24]. OpenSpace [BAC^{+}20]. Opera [LVR^{+}24]. Operations [CWR21]. Operative [ATHI24]. Operator [CSBK20]. Operators [PSL23]. Opinion [HB24]. Opportunities [BKR^{+}24, DPC23, WDSM24, WFW^{+}20]. Optic [SPWW^{+}24]. Optical [APP^{+}22, EBW23, FCFC22a, FCFC22b, GBM^{+}22, KPW20, MGWK^{+}22, PGE^{+}22, QSUK22, TIHS20, WH22a, WH22b, ZWP^{+}22, ZHKY23, BNWvW21]. Optics [HIK^{+}23, ZZZ24]. Optics-Based [ZZZ24]. Optimal [FK20]. Optimisation [LDT^{+}21]. Optimization [AKS22, BvdPLH22, CRWA22, CLFL22, CDS^{+}22, FYL^{+}23, HTBL22, HSV^{+}20, HWL^{+}22, HYSL23, HZJ^{+}24, LWH^{+}22, NSM^{+}22, PNKC21, PTM^{+}20, QCL23, QNWR23, RXX^{+}21, RGG20b, SHD^{+}20, WWG21, ZSLL20, ZWL23, ZCZ^{+}23a, ZLC^{+}22, dBRGD22]. Optimizations [DAK^{+}21]. Optimized [HCJ21, WHBC23]. Optimizer [KMG^{+}21]. Optimizing [CWO^{+}24]. Orbifolds [ZZZ24]. Order [HWN^{+}23, HNS23, JLK20, KKE21, LBP24, PFC20, QCL23, TZT^{+}22, YST^{+}20]. Order-Aware [KKE21]. Ordered [VBP^{+}21]. Ordering [FFB24, HCX^{+}21]. Orderings [vBMS22]. Organic [KBV24]. Organization [LCWL23, WN21]. Organizations [BK22, DZTF22]. Organized [VCO^{+}23]. Organizing [Ano21l]. Orientation [ATHI24, FMBN23]. Oriented [DKM24, HBLW21, KRS^{+}22, ZSW^{+}21, ZCL^{+}21]. OrthoAligner [CFZZ23]. Othello [ALC22]. Ötzi [NBE^{+}23]. Out-of-Core [SCRL20]. Out-of-Distribution [CYL^{+}21]. Out-of-Place [ZCZ23b]. Out-of-View [LYBP23]. Outcome [HM22, ZAH22]. Outcome-Explorer [HM22]. Outcomes [CPD^{+}24, HXL^{+}24, HX23]. Outcrop [OWN^{+}21]. Outdoor [ACL^{+}24, LTLB22, LZX^{+}22]. Outlier [XCK20]. Outpainting [ACL^{+}24]. Output [KBF22, XRN^{+}23]. Output-Sensitive [KBF22]. Overestimation [XCLF20]. Overlap [CWR21, LSL^{+}23]. Overlap-Free [LSL^{+}23]. Overview [MSM^{+}22, YCB^{+}21]. Overviews [BZSD21]. OW [JGH^{+}24]. OW-Adapter [JGH^{+}24]. Oz [RMW^{+}24].

P5 [LM20]. P6 [LM21]. PACE [MM23]. PacificVis [BSW20, CTW21, CRZ22, CRH23]. Packing [SKA21]. Page [Ano22g, Ano24g, Ano21p]. Painting [LZH^{+}21]. Pair [CAA^{+}21, NF20, YWB20]. Pair-Based [YWB20]. Pair-wise [CAA^{+}21]. Pairing [HJL^{+}23]. Palace [YQN^{+}21]. Palettailor [LFC^{+}21]. Palettes [SCL^{+}24, YZZ^{+}22]. Pancreatic [DMBK21]. Pancreatography [JDM^{+}22]. Panels [OWN^{+}21]. PANENE [JSF20]. Panorama [BHY^{+}23, CZF^{+}23]. Panoramas [LDZ^{+}21]. Panoramic [CXW^{+}23, MP21, PJYW21, PAAG22, ZPF^{+}22]. Paper [Ano20n, Ano21m, Ano21n, Ano22l, Ano22q, Ano23j, Ano23o, Ano24m, HLH^{+}23, NAW^{+}22]. Papers [Ano20n, Ano21m, Ano22l, Ano22m, Ano22z, Ano23j, Ano23k, Ano23v, Ano24s, Ano20m, Ano21j, Ano22p]. Paradigm [AZA^{+}23, LPG^{+}22, LCK^{+}23, LZZ^{+}24, PCQ^{+}20]. Parallax [FCFC22b, IAI^{+}23, MF24, FCFC22a]. Parallel [BKS22, CWS^{+}21, CRWA22, CMSK23, ESP20, LM20, LKS22, LMGY22, LI24, MLT^{+}24, NBJ^{+}21, NLPW22, SBT^{+}23, SPN23, TEK^{+}23, VBP^{+}21, WG21, WBI20, XHS^{+}24, XGS^{+}21, XGS^{+}23, ZS21, MSuG^{+}23]. Parameter [FiMH21, GGT^{+}20, HWG^{+}20, KSHW22, SXW^{+}22, SXL^{+}23]. Parameter-Sharing [FiMH21]. Parameterization [ZWW22]. Parameterizations [HYSL23].

Parameters [CBP22, HPT^{+}23, LQWQ21]. Parametric [AJSP23, HHB^{+}23, RAFSA23b]. Part [WL20]. Partial [LWF23, LFG^{+}23]. Partially [ZWP21]. Partially-Sorted [ZWP21]. Participants [PSH20]. Participating [GWW^{+}21, JKU^{+}22, LWWH20, WGH20]. Particle [IRR^{+}22, LS23, LGH^{+}24, MZS^{+}24, NL24, TYW^{+}22, XGS^{+}23]. Particle-Based [LGH^{+}24]. Particles [ZHQH20]. Partitioning [FFB24, MSuG^{+}23, ZXC^{+}23]. Parts [GMTD23, LB22]. Passability [BVV^{+}22, BVV^{+}23]. Passage [SAPW23]. Passenger [MMN^{+}22]. Passes [XWL^{+}21]. Passing [CFGT21]. Passive [ZUK21, ZZS^{+}23]. Passthrough [GWD^{+}24]. PassVizor [XWL^{+}21]. Past [MA23]. Patch [CWS^{+}20, HCL20, MSY22]. Path [CAR^{+}23, DST^{+}23, FK20, HAK^{+}23, IGMM22, LETF21, MSuG^{+}23, WAA^{+}22, WPTG24, KOL^{+}20]. Path-Following [LETF21]. Path-Traced [CAR^{+}23]. Pathogen [BPW^{+}21]. Pathology [CCW^{+}21, JKW^{+}22]. Paths [RBSN22]. Pathways [BPW^{+}21, KAS^{+}21]. Patient [BPW^{+}21]. Pattern [LBB^{+}20, LqZ23, NdCS21, SESH24, SDXR22, WLGW23, YKJ^{+}23, ZWW^{+}23d]. Pattern-Driven [LBB^{+}20]. Patterns [BFAR^{+}23, BFS^{+}24, CZL^{+}21, DCM^{+}23, EBJ^{+}22, GZRP^{+}22, GSP^{+}21, LG23, LSS24, MGO21, NP24a, SPZS24, VCO^{+}23, WDG^{+}20, ZGX^{+}23, ZLG^{+}21b]. Paulo [GSP^{+}21]. PBM [YZEN22]. PC [TEK^{+}23]. PC-Expo [TEK^{+}23]. Peak [CWS^{+}21]. PeckVis [CCM20]. Penalty [DS20]. Penetration [AKS22]. People [GKC^{+}24, GKN^{+}23, JMK^{+}22, MCQ23, MKK20, MJAD22, PMN^{+}23, RDH23, RBLT^{+}22, ZNAN20]. Per-Ray [WWSS20]. Perceived [ASCR^{+}22, BVV^{+}22, BVV^{+}23, CZGF21, GKC^{+}24, HWS22, LCW^{+}23a, PGE^{+}22, PIS20, SA22, TWA22].

Perception [BBL^{+}22, BBVS^{+}24, CBW23, DPD^{+}24, DGKOC20, EKC^{+}23, FRL^{+}23, HZZ^{+}20, Kel23, KVM^{+}22, LWK20, MGWK^{+}22, MF24, MCFKF24, NF20, QR21, QR22, RSRG23, SLW^{+}24, UVL^{+}23, WBWL24, WBPC23, WAC^{+}22, YGE^{+}21, ZYM^{+}24, ZSL^{+}22, BNWvW21]. Perception-Based [QR22, BNWvW21]. Perceptions [ACPB24, HXL^{+}24, VVR^{+}23b]. Perceptual [APBB24, ATK^{+}24, BPL23, CDS^{+}22, GSM^{+}22, JOEF20, JQL^{+}24, LSD^{+}23, MSK23a, MYS^{+}22, OYK^{+}21, SPWW^{+}24, SMYF22, WZZ^{+}23, WMZ^{+}20, XCLF20, XXM^{+}21, ZLS21]. Perceptual-Aware [XXM^{+}21]. Perceptually [CZT^{+}21, GGT^{+}20, SMCL24, ZZ23]. Perceptually-guided [ZZ23]. Perceptually-Matched [CZT^{+}21]. Percutaneous [LCH^{+}21]. Perform [PvSvdE^{+}24]. Performance [APP^{+}22, BBL^{+}22, BCC^{+}20, BMFE20, CMKK21, CFDN24, DPR^{+}20, DPD^{+}24, DZG^{+}23, ENK^{+}20, FMBN23, FS24, GWF^{+}23, GTC^{+}23, GWD^{+}24, HKB^{+}22, JKL24, KVB20, KSK^{+}23, KVM^{+}22, KKF20, KPO^{+}23, LMM^{+}21, LAC^{+}24, LETF21, LCM^{+}23, MBHE24, NBJ^{+}21, OSC22, PFCB23, PRKM20, RMW^{+}24, SBT^{+}23, SFPM22, SWY^{+}22, TARB23, UDH23, VSBY22, WSZ^{+}23, WPZ^{+}23, WQP^{+}22, WBWL24, ZKS^{+}20]. Performances [XWY^{+}20]. Periodic [ZH20]. Peripersonal [FCFC22a, FCFC22b]. Periphery [KCA^{+}21]. Persistence [GVT24, Iur22, PVT23, SDT24, VBT20]. Persistent [Iur22, SHW^{+}20]. Person [BVV^{+}22, BVV^{+}23, SAMB^{+}23, YLS^{+}23]. Person-Plus-Virtual-Object [BVV^{+}22, BVV^{+}23]. Personal [BCB22, ISBP22, JLH24, LAML23, MGWM22, RFD21, SRKK21]. Personalization [TIDQ24]. Personalized [CSWZ24, HHD^{+}20, HHM^{+}24, LCH^{+}21, LHS^{+}22, SWF^{+}24, SEAB^{+}22, ZNF^{+}23]. Persons [CFQ21, GGP^{+}24]. Perspective [HHKN21, HMKB23, LHL^{+}22, YYD^{+}21]. Perspectives [BKT^{+}22, CFL^{+}23]. Pet [XLZ24]. PetPresence [XLZ24]. Phase [DGDC21, ILZ^{+}21]. Phase-Modulated [ILZ^{+}21]. Phased [SZH^{+}20]. Phenomena [WM23]. Phenotype [KBJ^{+}20]. Phoenixmap [ZLG^{+}21a]. Phone [YZJ^{+}20]. Phones [BLIC20, WTD^{+}21, ZYL^{+}24, ZYL^{+}24]. Photo [NMS^{+}23, SYX^{+}22]. Photo-Reflective [NMS^{+}23]. Photograph [CLL^{+}20, YCZ^{+}22, ZLZ^{+}23b]. Photographic [ZRJW20]. Photographs [Tet24]. Photography [ZTL^{+}23]. Photometric [SUB^{+}22]. Photometry [RPG23a]. Photometry-Guided [RPG23a]. Photon [BWM24, HWMI23, SWWY21]. Photorealistic [HHD^{+}20, PCJ23, WWR^{+}20]. Photos [MSK23b, YXS^{+}23]. Photosensitive [SB23]. PhraseMap [TQW^{+}24]. Phylogenetic [LZM20]. Physarum [EBPF21]. Physical [DUWW22, JCZ^{+}24, KKS^{+}22, LNB^{+}21, LQWQ21, VVC^{+}24, WCCS24, WBPC23, XLW^{+}24, YKF22, ZCL23, VGK^{+}22]. Physicalization [DPC23, KBV24, PWB21]. Physicalizations [BFY^{+}24]. Physically [PPYW20, ZHTR22]. Physically-inspired [PPYW20]. Physically-Observable [ZHTR22]. Physics [FGS^{+}21, YGP^{+}24]. Physics-Informed [YGP^{+}24]. Physiological [LPL^{+}24, LH22, ZCL23, MSAM^{+}22]. PICO [KMG^{+}21]. Pictographs [BXF^{+}22]. Pictorial [SLC^{+}23]. Piecewise [MLT^{+}24, ZLZ21]. Piling [LZC^{+}21]. PinNPivot [GSH21]. Pins [GSH21]. Pipeline [BFY^{+}24, CNAA^{+}22, MSS21]. PipelineProfiler [OCL^{+}21]. Pipelines [LM20, OCL^{+}21]. Pixel [HZQ22, LWWY21]. Pixel-Wise [HZQ22]. Place [BC21, ZCZ23b, MCSAL23]. Placed [WXS^{+}24]. Placement [CCL^{+}24, SFPM22, YYK^{+}22]. Places [SGB^{+}22]. Planar [DLP^{+}23, GZM^{+}21, GZY^{+}22, HBMK23, KRK21]. Plane [LFR^{+}21, vOVR23]. Plane-Based [LFR^{+}21]. PlaneFusion [GZY^{+}22]. Planning [DST^{+}23, RBKM24, SHC^{+}20, XFF^{+}21]. PlanningVis [SHC^{+}20]. Plant [GXY^{+}20]. Plants [LWT^{+}23]. Platform [APSB23, CSIP22, OF22]. Platforms [ZWG^{+}23]. Plausibility [BWWL22, DCC22, ISBP22, MWD^{+}23, WBLW23]. Plausible [DMJ^{+}22, WCC^{+}21]. Play [BVY^{+}23, CIA24]. Player [CXZ^{+}24, GWD^{+}24, SZZW24]. Players [CB22]. Playing [CFZ^{+}23, FMP23, GLY^{+}23, HNGC21, LGY^{+}22]. Pleasure [HIDI23]. Plexus [CNC^{+}20]. Plots [BKS22, KNAR^{+}22, RSD^{+}23, XPK^{+}24, ZAH22]. PlotThread [TLW^{+}21]. PLUME [JVRL24]. Plus [BVV^{+}22, BVV^{+}23]. PMU [APSB23]. POI [XLL^{+}22]. Poincaré [TSH21]. Point [BPA22, CZY^{+}20, CWS^{+}20, CDZ^{+}23, FALH20, GBNH21, HCL20, LLC^{+}22, LB22, LLWF23, LWF23, LFO23, LLDW24, LSD^{+}23, LMD^{+}22, MDH^{+}23, MXT^{+}22, MWUP22, PC22, QHL^{+}20, RPHJ20, WW22a, WCZ^{+}23, XF21, YCH^{+}22, ZHDX20, ZLQH21, ZZD^{+}23a, ZZD^{+}23b, ZZW^{+}22b]. Point-Feature [PC22]. Point-in-Polygon [WW22a]. Pointer [SPW^{+}22, WSN21]. Pointfilter [ZLQH21]. Pointing [KCA^{+}21, LTLB22, UDH23]. PointRetrieval [ZZD^{+}23b]. Points [CFL21, KKV22]. Poisson [BKPB21, HSF^{+}20, XZWY20]. Polarimetric [SUB^{+}22]. Polarization [HLW^{+}20, HB24]. Polarizing [HB24]. Pole [DXX^{+}21]. Political [HB24]. Polls [HB24]. Pollution [DWC^{+}20]. polycephalum} [EBPF21]. Polygon [QZZ22, WW22a].

Polygonal [Liv21]. Polygons [WBM21b]. Polyhedra [AKS22]. Polymers [HWN^{+}23]. Polyphony [CKQ^{+}23]. Polyphorm [EBPF21]. Pooling [QGY^{+}22]. Poorly [LCYQ24]. Pop [HLH^{+}23]. Pop-Up [HLH^{+}23]. Porous [HWL^{+}22, SKR^{+}24]. Portable [LM20]. Portfolios [YBL^{+}20]. Portrait [HYH^{+}23, SYC^{+}23, ZZW^{+}20, ZLZ^{+}23b]. Portraits [WWR^{+}20]. Pose [CLZ^{+}24, MZX^{+}21, MBHE24, ZHDX20, ZNF^{+}23]. Pose-Dependent [MBHE24]. Poser [RZH^{+}23]. Position [CMXF21, HWS22, XCLF20]. Positional [BBM^{+}21]. Post [CNK^{+}24]. Post-Stroke [CNK^{+}24]. Posts [KKT^{+}22]. Posture [MSWI22, SACB^{+}23]. Potential [BSB^{+}20, CHSC24, WHBC23]. PoVRPoint [BGS^{+}22]. Power [APSB23, CPD20, LWT^{+}23, PPE23, WEM^{+}21, ZHQH20, ZWH^{+}22, ZPG21]. Power-Budget [ZWH^{+}22]. Power-Confluent [ZPG21]. Powered [WTL24]. PowerNet [ZWH^{+}22]. Practical [SCEA23]. Practice [KPO^{+}23, Par22]. Practices [LBE20, ZSG^{+}23]. pragmatic [ZWV21]. Pre [BXQ^{+}22]. Pre-built [BXQ^{+}22]. Preattentiveness [KCWK20]. Precedent [RPNP23]. Precise [ZSZ^{+}24]. Precision [HSB^{+}21, MHFF21]. Precomputed [WGH20]. Precueing [LETF21]. Precues [VLM^{+}23]. Predict [WYS^{+}22]. Predict-and-Drive [WYS^{+}22]. Predicting [BCN^{+}20, CWSJ23, HMGO23, MBC^{+}23, PSG^{+}22, SCHE23]. Prediction [BHY^{+}23, CFZZ23, DCWD23, FBW21, GMX23, HLZ^{+}20, JWW^{+}23, LWY^{+}20, NB24, TIDQ24, TFE21, WHL^{+}20, ZLL^{+}21]. Predictions [HMTI24, YC23]. Predictive [CNAA^{+}22, PFN22, VBV^{+}23, XNF^{+}23]. Preface [Ano23r, Ano24p, EFL^{+}21, LMY^{+}22, MGJ^{+}20]. Preference [FALH20, KJS^{+}23, WBWL24]. Preferences [MBC^{+}23, SSC^{+}23]. premium [WKMD22].

Prepare [CB22, DMTD22]. Presence [BCC^{+}20, BWWL22, CIA24, CBW23, CFDN24, CSM^{+}23, GMVRB20, HPdlG20, KVB20, MWD^{+}23, MNK23, NIM^{+}21, PTX^{+}22, PALW20, RR23, WBLW23, WBI20, YeSiK^{+}23]. Present [MA23]. Presentation [WQ20, ZWW^{+}20, ZWW^{+}23a]. Presentations [BGS^{+}22]. Preserving [CWW^{+}23, DTS^{+}21, LDC^{+}23, MDL^{+}23, SIA^{+}23, WLF^{+}22, WIP^{+}24, YLGW24, ZLZ21, ZLW^{+}21b, ZJC^{+}21, ZCZ^{+}21, DJHBJ21, DJBJ23, KGWD22, ZWW^{+}23d]. Prevalence [CFGM22]. Previews [LBP24, OM22]. PreVR [LBP24]. Primal [QZZ22]. Primal-Dual [QZZ22]. Primitives [RAFSA23b]. Principal [GSS^{+}20, PVT23, ZHN24]. Principles [KJS^{+}23, LCWL23, YLTL23]. Printed [TIHS20, TCX^{+}23]. Printing [XLY^{+}22, YRL^{+}20]. Prior [GZY^{+}22, PBF^{+}21]. priori [WEM^{+}21]. Priors [HCH^{+}23, ZZX^{+}22]. Prism [YDM^{+}21]. Privacy [CWW^{+}23, DJBJ23, WIP^{+}24, ZWW^{+}23d, DJHBJ21]. Privacy-Preserving [CWW^{+}23, WIP^{+}24, DJBJ23, ZWW^{+}23d, DJHBJ21]. Private [ZSM21]. Probabilistic [AJVS^{+}23, BDL^{+}21, DZG^{+}23, IRR^{+}22, PWK21, RPD21, SS24, XRN^{+}23, ZAH22]. Probabilities [MvdEPV24, RDH23]. Probability [HLCY22]. Probablement [RDH23]. Probing [GLL^{+}24, WPB^{+}20]. Problem [AL21, HBH^{+}20, LDT^{+}21, XVW^{+}21, ZLL^{+}21]. Problem-Driven [HBH^{+}20]. Problem-Solving [LDT^{+}21, XVW^{+}21]. Problems [ZNAN20]. Procedural [GXY^{+}20, JYZW20, KVGM20, KMG^{+}21]. Procedure [SHOP23]. Process [CLCY20, GSL21, LWL^{+}22b, LAS^{+}20, ZeB^{+}21, ZMK^{+}20]. Processes [BZP^{+}20, Hei21, HZJ^{+}24, MFH^{+}21, RZLX24, SKR^{+}24, YDMP22]. Processing [BPA22, CFL21, HTY^{+}23, HCL20, LM20, SCRL20, TZT^{+}22]. Production [SHC^{+}20]. Productivity [MMN^{+}22]. Professional [HKB^{+}22]. Profiles [JCZ^{+}24, NHC^{+}20, NBJ^{+}21]. Profiling [EGMP24]. Program [Ano20l, Ano20m, Ano20o, Ano20p, Ano21i, Ano21j, Ano21o, Ano21q, Ano22m, Ano22p, Ano22-29, Ano23k, Ano23n, Ano23z, Ano24l, Ano24v, BPQW23, FGL^{+}23, FG22, HKVZ20, IMKP21, IGMW22, Ano21k, IPPZ24]. Programmable [KJI^{+}21]. Programming [HWW^{+}24, HHS^{+}23, HLH^{+}23, PLD^{+}23]. Programs [WBI20]. Progression [KAS^{+}21, WMH^{+}22]. Progressive [CZF^{+}22, HZP^{+}24, JSF20, JYLS21, LM20, LLDW24, ML24, PRJ^{+}23, PMS^{+}22, VBT20, VGT21]. Project [ALR23]. Projected [MHIS23]. Projecting [YGH^{+}23]. Projection [AYA^{+}21, ATHI24, BXZ^{+}21, DTS^{+}21, EIB23, GBL^{+}22, HIK^{+}23, KIPS21, KLTB21, KLSB22, LRZ^{+}23, MSWI22, TKIS24, TIHS20, YIIW24, NLPW22]. Projection-Based [AYA^{+}21]. Projections [CMK20, EAKC^{+}20, EAS^{+}23, JKJ^{+}22, WZH20, WCJW22]. Projective [LLK22, YZEN22]. Projector [HL21, IGM24, KIS22, SII^{+}21, TGM21, UIHS20, WLH24, YITS23]. Projector-Camera [HL21, WLH24, YITS23]. Projectors [ATHI24, TKIS24]. Prolonged [BFS^{+}24]. Prominent [CWL23, WFY^{+}21]. Promote [ABGG21]. Promoting [NKWW22]. Promotion [ZWZ^{+}23]. PromotionLens [ZWZ^{+}23]. Prompt [FWW^{+}24, SWS^{+}23]. PromptMagician [FWW^{+}24]. Propagating [KNAR^{+}22, ZHN24]. Propagation [APSB23, CAR^{+}23, DWC^{+}20, LMM^{+}21, LWH^{+}22, MXT^{+}22]. Propeller [KCGZ23]. Propeller-Based [KCGZ23]. PropelWalker [KCGZ23]. Properties [CSL^{+}21]. Proportion [CZW^{+}20, QSC^{+}21]. Proportion-Related [CZW^{+}20, QSC^{+}21].

Proportional [LT20]. Props [MML21]. ProReveal [JYLS21]. Protein [FJK^{+}20]. Proteus [MWD^{+}23]. ProtoColVR [GF24]. ProtoSteer [MXC^{+}20]. Prototypes [MXC^{+}20]. Prototyping [GVN^{+}20, GF24, LLT24, SDMK22]. Provectories [WHX^{+}23]. Provenance [BER^{+}23, CBB23, WSL^{+}20, WHX^{+}23]. Provide [CAGM22]. Providing [TWT^{+}22]. Proxies [JOEF20, OYK^{+}21]. Proximate [FGF^{+}21]. Proximity [PT20, SZH^{+}20, TME^{+}22]. Proxy [CZGF21, Off20, OWW^{+}24, WWSS20]. PRS [GZM^{+}21]. PRS-Net [GZM^{+}21]. Pruning [CWS^{+}21, LWS^{+}21]. Pseudo [BPL23, YB20, YKJ^{+}23]. Pseudo-Haptic [YB20]. Pseudo-Stiffness [BPL23]. PSRFlow [SS24]. Psychoacoustic [CAR^{+}23]. PU [MDH^{+}23, YCH^{+}22]. PU-Flow [MDH^{+}23]. Public [CB22, HWT^{+}24, HB24, HHM^{+}24, LQS^{+}23, YCM^{+}24, ZKS^{+}20]. Publication [BvOR21]. Publication-Tailored [BvOR21]. Publications [CLL^{+}21, SFNRZ^{+}23]. Pull [XCLF20]. Pupil [WCX21b, WH22a, ZSZ^{+}24]. Pupil-Matched [WH22a]. Pupillometry [PCQ^{+}20]. Puppetry [PZC^{+}23]. Purely [TNJ^{+}22]. Purposes [WAV^{+}21]. Pursuing [BWI21]. Putting [KKW23]. puzzle [WKMD22]. PuzzleFixer [YCC^{+}23]. PVG [HSF^{+}20]. PW [dSBdO^{+}24]. Pyramid [CZF^{+}22, YXL^{+}22]. Pyramid-based [CZF^{+}22]. Pyramids [SAPW23]. PyramidTags [KKE21].

QEVIS [SYY^{+}24]. QLens [XVW^{+}21]. QR [HYF^{+}20]. Quad [FLW^{+}21]. Quad-Robots [FLW^{+}21]. Quadratic [ZWW22]. Quadrature [MBB20]. Quadrilateral [DLP^{+}23]. QuadStack [GRi^{+}21]. QualDash [ERB^{+}21]. Qualitative [HTP^{+}23, KHL21, MGU^{+}21]. Quality [CAR^{+}23, CLFL22, ERB^{+}21, FGS^{+}21, HTJ^{+}20, HMH^{+}21, HLL^{+}24, IRR^{+}22, LCK^{+}21, LSD^{+}23, MSK23a, MBC^{+}23, MLBW20, NDF^{+}21, PM23, QCCC23, QLC^{+}24, RRK^{+}22, SWS20, SMYF22, WSL^{+}20, YGW^{+}24, XJZ^{+}21]. Quantifiable [WWZP22]. Quantifying [BKN^{+}22, WJBB22]. Quantitative [EKC^{+}23, EMK^{+}21, MGU^{+}21]. Quantivine [WLT^{+}24]. Quantum [LXL21, RWJ^{+}23, WLT^{+}24, ZBG^{+}24]. Quasi [NGBA^{+}20]. Quasi-Conformal [NGBA^{+}20]. Queries [AJVS^{+}23, MZS^{+}24, NSS21, XZF^{+}22]. Query [CWW^{+}23, HE24, HZC^{+}20, LLS^{+}20, LBW^{+}22, MCW^{+}20, SUB^{+}22, SYY^{+}24]. Query-Driven [SUB^{+}22]. Querying [ATAS21, CSIP22, FiMH21, JSF20, KGBP20, TLW^{+}23]. Quest [LCY^{+}23]. Question [QTW^{+}24, WJBB22, XVW^{+}21, YCHL24]. Question-Based [QTW^{+}24]. Questionnaire [LCS^{+}24, RL20]. Quick [MSuG^{+}23]. Quiver [RLG^{+}23]. Qwerty [DZG^{+}23].

R [CLCY20]. R-Map [CLCY20]. R2CNN [LZZ^{+}21]. Racial [PGS21]. Racing [HAK^{+}23]. Racket [WMH^{+}23, WLG^{+}22, WLGW23]. Radial [JdJTC24, SPWW^{+}24, WDG^{+}20, XWH^{+}23]. Radiance [DHY^{+}22, SCL^{+}23b, WPZ^{+}23, ZLZ^{+}23a]. Radiation [WHL^{+}20]. Radio [ZLL^{+}20, ZGX^{+}23]. Radiologists [JK23]. RadViz [ABL^{+}22]. RagRug [FCH^{+}23]. RainBio [LT20]. Rainbow [Red23, bÇ22, bÇ22]. Rainbows [RS21]. Ramps [SWS20]. Random [NP21, SGH^{+}23, SLR20b]. Randomised [LKS22]. Range [AJVS^{+}23, ATHI24, LHWW22, MZS^{+}24, SMC^{+}21, ZBG^{+}24, ZRJW20]. Ranges [BBSvL24]. Rank [CWS^{+}20, LSL^{+}22, ZZD^{+}23b].

Rank-PointRetrieval [ZZD^{+}23b]. RankAxis [LRZ^{+}23]. Ranking [DPD^{+}24, LRZ^{+}23, MLC^{+}20, XMT^{+}21]. Ranks [MYBF22]. Rapid [GF24, LLT24, PC22, SDMK22, WGH20]. RASIPAM [WLGW23]. Rasterization [LZZ^{+}21, TG24]. Rate [NMR^{+}23, WSZ^{+}23, YWB20]. Rated [San20]. Ratings [MZX^{+}21]. Ratio [CMXF21]. Rational [WGM^{+}24]. Rationale [YZEN22]. Rationalizing [XCSJ22]. Ray [HPT^{+}23, HJZ^{+}21, LME^{+}23, LTLB22, PFN22, VSBY22, WSN21, WZU^{+}21, WMZ22, WWSS20, XWPG^{+}24]. Ray-Based [WWSS20]. raying [JKV^{+}22]. RBFs [MZS^{+}24]. RCMVis [SJK^{+}23]. RDW [DGD^{+}23]. Reach [HAK^{+}23]. Reach-the-Target [HAK^{+}23]. Reaction [CIA24, GTL^{+}23, MWR^{+}22]. Readability [BZW^{+}21, LIDM20, WJKN21, YPW23]. Readable [GLA^{+}24b]. Reader [SSC^{+}23]. Reading [ALC22, AR22, NL24]. Real [BWM24, COFJ23, CFL^{+}23, GLH22, GTH20, GZY^{+}22, HCH^{+}23, IGM24, IGMM22, JBS^{+}22, KVB20, KKT^{+}22, LBW^{+}22, LWSY20, LXL21, LZX^{+}22, LWL^{+}23b, LCCZ22, MP21, NMR^{+}23, PFC20, SIA^{+}23, SGB^{+}22, SLK^{+}20, WLF^{+}22, WGS^{+}24, WBPC23, WTY^{+}22, XLZ24, YZJ^{+}20, YGH^{+}23, ZWH^{+}22, ZLX23, ZHL^{+}21b, CNB^{+}22]. Real-Time [BWM24, GZY^{+}22, HCH^{+}23, IGM24, IGMM22, JBS^{+}22, KKT^{+}22, LWSY20, LXL21, LZX^{+}22, LWL^{+}23b, LCCZ22, MP21, PFC20, SIA^{+}23, WLF^{+}22, ZWH^{+}22, ZLX23, ZHL^{+}21b, LBW^{+}22, SLK^{+}20, YZJ^{+}20]. Real-World [COFJ23, WTY^{+}22, XLZ24, WGS^{+}24]. Realism [WBPC23]. Realistic [CXW^{+}23, GXY^{+}20, JLM^{+}21, MSB^{+}22, TNZ^{+}24]. Reality [Ano22x, APP^{+}22, BTHL23, BBL^{+}22, BJCL21, BS22, BRLR24, BBVS^{+}24, BWCT23, BSG^{+}20, BGS^{+}22, BBSC23, BBM^{+}21, BFS^{+}23, BC21, CCP^{+}21, CGAG20, CZMR21, CIA24, CSW^{+}20, CDS^{+}22, CMF^{+}22, CFL^{+}23, CZGF21, CFDN24, CBHR^{+}23, CNK^{+}24, DUWW22, DCC22, DHY^{+}22, DPM24, EMM^{+}22, EML^{+}23, EHB^{+}23, ENvBC23, EBW23, FE21, FBW21, FHR^{+}21, FTWP22, FCFC22a, FCFC22b, FPK^{+}24, FYC^{+}23, FKSP^{+}24, FG22, GSM^{+}22, GWL^{+}22, GMTD23, GST^{+}24, GGP^{+}24, GEU^{+}22, GLH22, GBL^{+}22, GF22, GTH^{+}22, GPR^{+}24, GKN^{+}23, GZG^{+}24, HZZ^{+}20, HWC23, HFS^{+}21, HPdlG20, HSJ^{+}20, HBLW23, HRX23, JLH24, KKW23, Kel23, KNL23, KL22, KPL23, iKYOW23, Kiy22b, Kiy22a, KBPR22, KVM^{+}22, KPW20, KGR^{+}24, KWFK20, LBL^{+}21, LNB^{+}21, LSS24, LSG24, LWL^{+}22a, LFCH24, LTLB22, LKS22, LZY^{+}23, LCW^{+}23a, LETF21, LZX^{+}22, LXZ^{+}23, LWL^{+}23b, LZZ^{+}24, LWK20, LH22, LCM^{+}23, MA23, MCQ23, MWD^{+}23]. Reality [MTE^{+}20, MSA^{+}22, MMK^{+}23, MUM^{+}21, MMN^{+}22, MWR^{+}22, MGM^{+}22, MML21, MF24, MCS^{+}23, MKK20, MCFKF24, MHIS23, MSK23b, MB20a, MB20b, MB21a, NGW^{+}24, NDF^{+}21, NP24b, OSC22, OF22, PTX^{+}22, PPYW20, PCL24, PSH20, PGS21, PRKM20, PMW23, PIS20, QSUK22, RS23, RBR20, RMW^{+}24, RFD21, RNO^{+}22, RRK^{+}22, RLG^{+}23, SWB^{+}22, SML^{+}23, SDMK22, SMPJ^{+}20, SES20, SMC^{+}21, SKS^{+}23, SDK23b, SW21, SMSK22, SAB20, TIDQ24, TNJ^{+}22, TLBB23, TWA22, TCX^{+}23, TBW^{+}23, TME^{+}22, TPH22, UDH23, VVR^{+}23a, VVR^{+}23b, VVR^{+}24, VVC^{+}24, VH23, VBV^{+}23, VLM^{+}23, WZL22, WYS^{+}22, WSZ^{+}23, WSL^{+}24, WQP^{+}22, WFK24, WGS^{+}24, WBLW23, WGO20, WGO22, WMMB23, WIP^{+}24, WR23, WH22b, WPNK21, XLZ24, XHFZ24, YLS^{+}23, YQN^{+}21, YHC^{+}22, YeSiK^{+}23, YLF^{+}20, YZN^{+}20, YB20, YEP^{+}22, ZZ23, ZMK^{+}20, ZUK21, ZZW21, ZHH22, ZYL^{+}24, ZSCRB23, ZYR^{+}20, ZG20, ZWP21]. Reality [ZCM^{+}23b, ZCL23, ZLG^{+}21b, KHD^{+}22, MGC^{+}21, MCSAL23, Ano23f, Ano23g]. Realtime [WCX21b, WCC^{+}21, XCG^{+}21, XWZB21]. Rearrangement [NP24b]. Reasoning [JKV^{+}22, KKH21, KRS^{+}22, WM23, WHJ^{+}24, ZTC^{+}23]. Reassembly [YCC^{+}23]. Recall [CMXF21]. Recallability [WJBB22]. Receptivity [HWT^{+}24]. Recital [BK22]. Reclaiming [BBSvL24]. Recognition [LZZ^{+}21, LZL^{+}23, NMS^{+}23, ZG20]. Recognizing [HBLW23]. Recommendation [CFGM22, LST^{+}22, LWZ^{+}22, LQS^{+}23, PLW^{+}23, TQW^{+}24, WWZ^{+}22, YZZ^{+}22, ZMD^{+}22]. Recommendations [OKM21, PSS23, ZYM^{+}24, ZWG^{+}23]. Recommender [KH21]. Reconnaissance [CFGM22]. Reconstructing [CZMR21]. Reconstruction [CGAG20, FYL^{+}23, GBM^{+}22, GZY^{+}22, HL21, HSC^{+}22, HCH^{+}23, LCK^{+}21, MDL^{+}23, MA20, SWZ^{+}23, WZH20, WITW22, XJZ^{+}21, XCG^{+}21, YLL^{+}20, YZJ^{+}20, YFS22, YFM^{+}23, ZWL23, ZLZ^{+}23a, ZLX23]. Record [JVRL24]. Recordings [NGW^{+}24]. Records [LLP^{+}23, OWW^{+}24, SNBC23]. Recovering [CLL^{+}20, FZZX22]. Recovery [CWS^{+}20, CCS^{+}21, XLL^{+}20]. Rectangular [ZLZ21]. Rectilinear [LDB^{+}21]. Recurrent [WCX21a]. Recursive [CGZ^{+}20, ZZW^{+}22c, YZP^{+}23]. Recursive-NeRF [YZP^{+}23]. Recycling [TLW^{+}23]. Redesign [SLSW23]. Redirect [XCG^{+}24]. Redirected [AYGR22, AYG^{+}22, CS23, DGD^{+}23, FLS23, SMNK21, SRBP20, WCCS24, WBM21a, WBM21b, XLL^{+}22, XLW^{+}24, ZCZ23b]. RedirectedDoors [HFT^{+}24]. Redirection [AYGR22, CHSC24, HFT^{+}24, LJCL24, LF23, WBM21a, WHBC23, ZP23]. Redress [RBSN22]. Reduce [LVV^{+}21]. Reduced [MWR^{+}22]. Reduces [CFQ21, CQ22, MCS^{+}23]. Reducing [BTHL23, LGL^{+}23a, XPK^{+}24]. Reduction [ACT^{+}24, EMK^{+}21, FCS^{+}20, FKM20, FSS^{+}21, FWZM22, JKA^{+}24, KKG^{+}20, MBC^{+}23, XZS^{+}22, XHL^{+}23, XHS^{+}24, ZCX^{+}24, ZCH^{+}21]. Reeb [RAC22, WWZ23b]. Reevaluation [FMP23]. Reference [QCCC23, QLC^{+}24, RG20, SZC^{+}23, SFL^{+}22, WSN22, ZSCRB23]. Reference-Based [SZC^{+}23]. Referencing [LTLB22]. Referent [WO22]. Referents [APHD24]. Referral [SWSK23]. Refinement [EAKC^{+}20, LLDW24, Liv21]. Refining [KJS^{+}23, SCL^{+}24]. Reflectance [AYA^{+}21, CBP22, EIB23, JRM22, KGX^{+}23]. Reflection [CZMR21, DMMF21, DS22, HHM^{+}24]. Reflectional [QGL^{+}23]. Reflections [JRM22, SLR^{+}20a]. Reflective [GZM^{+}21, NMS^{+}23]. Refocusable [LDZ^{+}21]. Reforming [PBBH20]. Refractive [LWWH20]. Region [HZQ22]. Region-Based [HZQ22]. Regional [COZ^{+}23, NGBA^{+}20]. Regions [VFP24]. Registration [CYD^{+}23, FCFC22a, FCFC22b, HZFH20, LWF23, LFO23, RAFSA23b, TGM21, ZZD^{+}23b]. Regularization [CYD^{+}23, CLG21, FWW^{+}23]. Rehabilitation [CNK^{+}24, WHY^{+}23]. Reinforcement [CCL^{+}24, DWQW23, HDFK21, HCX^{+}21, HHB^{+}23, PHB^{+}22, SCC^{+}24, SRBP20, TLW^{+}21, WZY^{+}22, WTD^{+}21, XGS^{+}23]. Related [CZW^{+}20, VBV^{+}23, QSC^{+}21]. Relation [NF20, ZZX^{+}22]. Relational [SCL^{+}23a]. Relations [RPHJ20]. Relationship [ZZW^{+}22b]. Relationship-Based [ZZW^{+}22b]. Relationships [EHA^{+}23, SSAZ22, SNK^{+}22]. Relative [FALH20, PRKM20, RMB^{+}21]. Relaxed [RSD^{+}23]. Relevance [MGO21, YKJ^{+}23]. Relevant [SLK^{+}20]. Reliability [JKJ^{+}22, RMW^{+}24]. Reliable [SWZ^{+}23]. Relief [JHS^{+}21, YCZ^{+}22, ZZW^{+}20, ZLZ^{+}23b]. Relighting [HL21]. Remapped [BRLR24, ONH21]. Remapping [GGT^{+}20]. Remeshing [KPF^{+}22, KBV24]. Remote [FYC^{+}23, KHD^{+}22, KVB20, iKYOW23, LVR^{+}24, TLBB23, YLS^{+}23, YeSiK^{+}23]. Removal [WFC21, WWZP22]. Renderer [GGT^{+}20]. Rendering [ATAS21, AMS^{+}21, BWM23, BGB^{+}22, CZMR21, CXW^{+}23, ESP20, FWW^{+}23, GRi^{+}21, GG21, HHK^{+}24, IRR^{+}22, IGMM22, JBS^{+}22, JSA^{+}20, KNM^{+}21, KRK21, LXH^{+}21, LWL^{+}23b, MP21, MPV21, MDV20, MDJV21, MF24, NI22, NWW21, PGE^{+}22, RGG20a, SSB^{+}22, SMC^{+}21, SS21, TBL^{+}20, WCTW21, WW22b, XWZB21, ZSLL20, ZZW21, ZWP^{+}22, ZWH^{+}22, ZWL23, dPLM21]. Reordering [BSP20, KKhCM23, TEK^{+}23]. Repeated [AMAS21, KBPR22]. Repetitive [BBM^{+}21, MOA21]. Replacing [TKIS24, VGK^{+}22]. Replay [JVRL24]. Replica [CPD^{+}24]. Replicas [SGB^{+}22, TLBB23]. Replication [KHD^{+}22]. Replies [WDC^{+}23]. Report [ASSB^{+}23, LMF^{+}24]. Reporting [WRZ^{+}21]. Reports [SEAB^{+}22]. Repositories [EJS^{+}23, OKM21]. Reposting [CLCY20]. Represent [ZNAN20]. Representation [AI21, CBL^{+}24, GRi^{+}21, GKN^{+}23, HWL^{+}22, ISKM23, LS23, LYZ^{+}24, LGL^{+}23b, LWWY21, MCW^{+}20, SDXR22, SCL^{+}23b, SKNŽ20, TG24, WLT^{+}24, WXS^{+}24, XEXW24, XWZ^{+}23, ZSS^{+}21]. Representations [ASA^{+}23, BER^{+}23, DHLA22, HSYZ24, HIDI23, HTY^{+}23, HHC^{+}21, KBF22, MBHE24, QFWS22, SFPM22, VVR^{+}23b, WZW^{+}23]. Representative [OM22]. Representing [LCW^{+}23a]. Reproduction [PJYW21, SA22, WHY^{+}23]. reproductions [MHFF21]. repulsion [MHFF21]. RepulsionPak [SKA21]. Repurposing [WHC^{+}23]. Requirements [GF24]. Reranking [ZZD^{+}23b]. Research [DS22, ENXS21, GL20, KPF^{+}22, LVV^{+}21, PLP^{+}23, PCL24, PSH20, RVB^{+}22, SNH^{+}23, SBW21, STA^{+}21, SH24b, TBW^{+}23, WGS^{+}24, YBVI22, KHD^{+}22]. Researcher [Ano22d, Ano23d, Ano24e, GF22]. Reset [MTVS23]. Resets [XLL^{+}22]. Resetting [ZCZ^{+}23a, ZCZ23b]. Residency [HHK^{+}24]. Residential [ZTL^{+}23]. Resolution [CK22, CJY^{+}23, CSM^{+}23, FBB^{+}21, HW20, HZCW22, HW22, HSB^{+}21, RAC22, SS24, WCTW21, WGS^{+}23, XYF^{+}21, ZFCG23, ZGL^{+}21]. Resolutions [WW22a]. Resorts [RBKM24]. Response [CWSJ23]. Responses [JWE^{+}22, LCS^{+}24, LH22, VBV^{+}23]. Responsive [HBS^{+}21, KRS^{+}22, KRHH24, LG23, YHC^{+}24]. Restoration [YCC^{+}23]. Restraint [WMMB23]. Restricted [NRA^{+}23]. Restrictions [YZF^{+}23]. Results [FKM20, HCM^{+}22, KKW23, SBW21]. Retarders [HHI24]. Retargeted [MTVS23]. Retargeting [YYK^{+}22, ZUK21, ZCZ22]. Rethinking [MYBF22]. Retina [NDLW20]. Retinal [CZT^{+}21]. retirement [WKMD22]. Retrieval [LZZ^{+}20, ML24, QTW^{+}24, YQN^{+}21, ZZD^{+}23b, ZJJH21]. Retrieve [QSC^{+}21]. Retrieve-Then-Adapt [QSC^{+}21]. Retrofitting [SML^{+}23]. Retrospect [WXG^{+}24]. Retrospective [CFGT21]. Retrotransmissive [HIK^{+}23]. Reusable [Tsa21]. Reuse [CLW^{+}24]. Reusing [CWH^{+}22]. Reveal [VBV^{+}23]. Revealed [SPNG23]. Revealing [BWI21, LMM^{+}21, OYK^{+}21, XLF^{+}23]. Reverberation [FPG^{+}23]. Review [BS21, BFS^{+}23, CWR21, CBHR^{+}23, FLS23, GEU^{+}22, IMQ^{+}20, JYF^{+}20, KPF^{+}22, LSG24, MGM^{+}22, PCL24, PMN^{+}23, PMW23, SDC^{+}24, SFPCW23, THFI21, WBA^{+}23, WR23, ZZC^{+}22, MGC^{+}21, SNBC23]. Reviewers [Ano20n, Ano21h, Ano21m, Ano21a, Ano21n, Ano21r, Ano22l, Ano22q, Ano23j, Ano23o, Ano24m, Ano24a, Ano24w, Ano22-30, Ano23-27]. Reviewing [FSS^{+}21]. Revisited [ACPB24, HFS^{+}21, KCWK20, LJS21, RS21, WAV^{+}21]. Revisiting [DPD^{+}24, DCM^{+}23, XZS^{+}22, ZLL^{+}21, ZGL^{+}21]. Reviving [FZC^{+}21]. Reward [LKL23]. Reweighted [LYH^{+}23]. Reweighting [BZKG21, YGW^{+}24]. RGB [FYL^{+}23, LCC^{+}23, MEB^{+}20, XLL^{+}20]. RGB-D [FYL^{+}23, MEB^{+}20, XLL^{+}20]. RGBD [ZX22]. Rhythm [AYA^{+}23]. RIAS [AMAS21]. Rich [BDRW21]. Ridden [ZCZ^{+}23a]. Rigel [CWH^{+}23]. Right [DPM24, WNC^{+}22]. Rigor [MD20, RPH^{+}21]. RISeer [COZ^{+}23]. Risk [FRiM^{+}23, HPdlG20, NB24, PM23, TWW^{+}22]. Risk-aware [TWW^{+}22]. Risk-Taking [HPdlG20]. Risks [DPD^{+}24]. RIT [JdJTC24, CNB^{+}22]. RIT-Eyes [CNB^{+}22]. RL [CCL^{+}24]. RL-L [CCL^{+}24]. RNN [LZZ^{+}21, WZY^{+}22]. RNN-Based [WZY^{+}22]. RNN-Rasterization-CNN [LZZ^{+}21]. Road [LKJ^{+}20]. Roadmaps [LF23]. RoboHapalytics [DSD^{+}23]. Robot [BWCT23, DSD^{+}23, EHB^{+}23]. Robots [FLW^{+}21]. Roboviz [ALR23]. Robust [AYA^{+}21, BXQ^{+}22, CML24, GLX^{+}21, KCB^{+}21, LCL^{+}22, LCW^{+}23b, MSWI22, SDK23a, WH22b, YGP^{+}24, YHD21, ZSZ^{+}24]. Robustly [CWL23]. Robustness [CLS^{+}21, WWL^{+}20, WIP^{+}24]. Rocks [LALG22]. Role [BCN^{+}20, DGB^{+}22, GLY^{+}23, HNGC21, KVM^{+}22, LGY^{+}22, LBE20, NE24, RVB^{+}22, ZA21]. Role-Exchange [GLY^{+}23]. Role-Playing [GLY^{+}23, HNGC21, LGY^{+}22]. Rolling [DMTD22]. Room [HFT^{+}24, TKIS24, WYS^{+}22, WLC^{+}23].

Room-Scale

[HFT^{+}24, WYS^{+}22, WLC^{+}23]. Roses [FWM^{+}24]. Roslingifier [SKH^{+}23]. Rotate [CDBM22]. Rotating [MML21]. Rotation [GWL^{+}22, LLC^{+}22, NVRS^{+}21, WZZ^{+}23, XCG^{+}24, YB20]. Rotation-Invariant [LLC^{+}22]. Rotations [BBVS^{+}24]. Route [RBKM24, SJK^{+}23]. RSATree [MCW^{+}20]. RT [MWUP22, MZS^{+}24, VSBY22]. Rule [FWM^{+}24, McN23, NSK^{+}21, SMS^{+}22, YBOB24]. Rule-Based [SMS^{+}22, NSK^{+}21]. Rules [MBB20]. Running [LWL^{+}22b]. Runtime [BMFE20, dSBdO^{+}24].

S [THS^{+}21]. S2T [CHSC24]. S4 [TFE22]. Sacrificing [BXF^{+}22]. Safeguards [JYLS21]. Safety [CB22, HPdlG20, LHS^{+}22, NQE21, NP24b, ZKS^{+}20]. SafetyLens [NQE21]. Saliency [EBW23, MLC^{+}20, SLR21, ZHF^{+}20]. Saliency-Aware [ZHF^{+}20]. Salient [MLC^{+}20]. Same [CSL^{+}21]. Sample [CB22]. Samples [CYL^{+}21, RSD^{+}23, DJBJ23]. Sampling [BDL^{+}21, CGZ^{+}20, CZF^{+}22, HSV^{+}20, JBS^{+}22, JLX^{+}23, NI22, NSW24, RPD20, WITW22, YXX^{+}21, Yuk21, ZJC^{+}21, ZSS^{+}21]. Sand [ZHL^{+}21b]. Sandwich [GVT24]. SAniHead [DHF^{+}22]. Sans [PPE23]. São [GSP^{+}21]. SAV360 [BBMM^{+}23]. says [TB24]. Scaffold [XZS^{+}23]. Scagnostics [WWL^{+}20]. Scalability [VC20, YYD^{+}21]. Scalable [AJVS^{+}23, ADD^{+}22, CWS^{+}21, GLA^{+}24b, GLX^{+}21, HM24, HHK^{+}24, HE24, KBB^{+}23, KBV24, LBB^{+}20, LZD^{+}20, LWM^{+}20, Liv21, LWW^{+}24, OZZ24, PDD^{+}22, RGG20b, SSB^{+}22, THS^{+}21, YHD21, YFM^{+}23, ZYP^{+}24]. Scalar [CZC^{+}20, GVT24, HSB^{+}21, JLK20, KW23, LGMT21, VGT21, ZZW^{+}22a, ZRJW20]. Scale [ACT^{+}24, BHA^{+}23, DWL^{+}22, DCS^{+}24, GLHQ21, GBNH21, HIDI23, HHS^{+}23, HFT^{+}24, JLX^{+}23, KRZ^{+}20, KVGM20, KLSB22, LPJT^{+}22, LXZ^{+}23, LMD^{+}22, MCW^{+}20, NWMC23, PBBH20, PJYW21, RAFSA23a, SHL^{+}21, SDR22, SIA^{+}23, WYS^{+}22, WLC^{+}23, WYIS24, WFK24, WLT^{+}24, WXS^{+}24, XWZB21, XQXL23, YCH^{+}22, YLLW24, ZPF^{+}22, ZH20, KOL^{+}20, NRA^{+}23, NMC21, RZH^{+}23, ZWZ^{+}22, ZCH^{+}21, HMK^{+}20]. Scale-Free [JLX^{+}23]. Scale-Space [PBBH20]. Scales [DPR^{+}20, HMK^{+}20, MBS^{+}21]. Scaling [HKW23, HPRC20, WMZ^{+}20, ZHH22]. ScanGAN360 [MSB^{+}22]. Scanned [FZZX22]. Scanner [SCHE23, XJZ^{+}21]. Scanpaths [BBMM^{+}23, MSB^{+}22]. Scattered [RPD20, RPD21]. Scattering [GWW^{+}21, HJZ^{+}21, LWWH20, NI22, VM23, WGH20]. ScatterNet [MTW^{+}20]. Scatterplot [LSL^{+}23, QNWR23, THS^{+}21]. Scatterplots [CGZ^{+}20, CZF^{+}22, GLGB24, HT22, HWS22, HSV^{+}20, LWL^{+}20, MTW^{+}20, QR21, RBRG21, RRG23, SGH^{+}23, SKH^{+}23, WMZ^{+}20, YCB^{+}21, YTHL23, YXX^{+}21, ZS21]. Scenario [BER^{+}23]. Scenarios [CCL^{+}24, LHS^{+}22, XLW^{+}24]. Scene [ACL^{+}24, GZY^{+}22, ISKM23, LQWQ21, LZZ^{+}20, LLZ^{+}23, MSK23a, NP24b, PJYW21, PBAG23, PCJ23, RBR20, RAFSA23b, SCL^{+}23b, TBL^{+}20, WBS21, WXS^{+}24, XFF^{+}21, XCG^{+}21, YLL^{+}20, ZZX^{+}22]. Scene-Aware [TBL^{+}20]. Scenes [CWSJ23, HLZ^{+}20, JRM22, LHL^{+}20, MEB^{+}20, NL24, PAAG22, QLC^{+}24, RRK^{+}22, SHS^{+}22, SLR20b, WBM21b, XLL^{+}20, XWPG^{+}24, YFS22, ZHL^{+}21a, ZSG^{+}23]. SceneViewer [ZTL^{+}23]. Schedule [LKL23]. Scheduling [ABGG21, PFN22]. Schelling [CDZ^{+}23]. Schemas [Liv21]. Scheme [BWCT23, bÇ22]. Scholarly [GTC^{+}23, YQN^{+}21]. School [PLD^{+}23, RR23]. Science [Ano22m, Ano23k, CFGT21, ENXS21, EGMP24, FGL^{+}23, HKVZ20, IMKP21, IGMW22, MMF20, WQQ^{+}24]. Scientific [HTY^{+}23, LWM^{+}20, MBS23, PD24, RHHH20, SLX^{+}23, SS24, WH23, WGS^{+}23, XGS^{+}21, YSM^{+}20]. Scientists [DPR^{+}20]. Scientometric [SFNRZ^{+}23]. SciVis [Ano20o, Ano21o, Ano21n]. Scope2Screen [JKW^{+}22]. Score [KGR^{+}24]. Score-Based [KGR^{+}24]. Scowl [LSZC21]. Scraps [KBM21]. Screen [BSG^{+}20, GBL^{+}22, LWL^{+}23b, YLF^{+}20]. Screen-Space [LWL^{+}23b]. Screening [RNO^{+}22, SUS^{+}21]. Scribble [SSX^{+}20]. Scribble-Based [SSX^{+}20]. Scripts [XLF^{+}23, XFD^{+}23]. Scrollytelling [MBS23]. ScrollyVis [MBS23]. Sculpting [HE24]. SD2 [GTC^{+}23]. SDRQuerier [TLW^{+}23]. Sea [DMMF21]. Seamless [AZA^{+}23, FYL^{+}23, IGM24, QRZZ21]. SeamlessGAN [RPG23b]. SEAR [ZHH22]. Search [BCN^{+}20, BPW^{+}21, DKuH20, GST^{+}24, KKS^{+}22, MMS^{+}23, RMW^{+}24, SDXR22, YKJ^{+}23]. Searching [HA20, LYBP23]. Seated [MMK^{+}23, WZZ^{+}23]. Seating [WMMB23]. Secondary [VVR^{+}24]. Section [BSW20, CTW21, CRZ22, CRH23]. Security [JJKJ20, LAS^{+}20]. Security-Utility [JJKJ20]. See [APP^{+}22, EMM^{+}22, EBW23, FCFC22a, FCFC22b, IAI^{+}23, KPW20, MGWK^{+}22, PGE^{+}22, QSUK22, WZL22, WBWL24, WH22a, XSKF23, YPW23, ZWP^{+}22, ZHKY23, BNWvW21]. See-Through [APP^{+}22, EMM^{+}22, EBW23, FCFC22a, FCFC22b, MGWK^{+}22, PGE^{+}22, QSUK22, WZL22, WBWL24, WH22a, ZWP^{+}22, IAI^{+}23, KPW20, ZHKY23, BNWvW21]. Seeing [WMB23, XSKF23]. Seek [WPL^{+}22]. Seeking [TQW^{+}24]. SEG [LLL^{+}22]. SEG-MAT [LLL^{+}22]. Segment [MBS^{+}21, ZWZ^{+}22]. Segmentation [ATAS21, CGT^{+}24, HZS^{+}22, HCL20, KCB^{+}21, LLL^{+}22, LZZ^{+}20, LWWY21, NBE^{+}23, PD24, SSX^{+}20, SFSA24, TYPC20, WL20, WCX21b, WYZ^{+}21, ZGX^{+}22, ZCSS23]. Segmentations [MLT^{+}24]. Segmented [LF23]. Segments [MM21]. Selection [BWZ^{+}20, BZKG21, CLX^{+}23, CMKK22, CZY^{+}20, CS23, DXX^{+}21, HTW20, HZX^{+}21, JYZW20, KGR^{+}24, MGO21, PMCM24, SNH^{+}23, SPNG23, SDK23b, YZN^{+}20, ZIX^{+}24]. Selection-Bias-Corrected [BZKG21]. Selective [LJCL24, WFC21]. Self [BVV^{+}22, BVV^{+}23, BBH24, CNK^{+}24, DGB^{+}22, GFSHO20, GWC^{+}23, HYC^{+}23, KPL23, LB22, LGY^{+}22, LLWF23, LZZ^{+}24, NIM^{+}21, PLD^{+}23, PT20, RPG23b, SWF^{+}24, SOL^{+}22, TFE22, ZLX23, MSAM^{+}22, VAWL24]. Self-Actuated [SOL^{+}22]. Self-Attention [LB22]. Self-Avatar [BBH24, PT20, MSAM^{+}22, VAWL24]. Self-Avatars [BVV^{+}22, BVV^{+}23, PLD^{+}23]. Self-Calibrated [GWC^{+}23]. Self-Embedding [LLWF23]. Self-Guided [LZZ^{+}24]. Self-Identification [GFSHO20]. Self-Illusion [LGY^{+}22]. Self-Location [NIM^{+}21]. Self-Occlusion [ZLX23]. Self-Similar [KPL23]. Self-Supervised [HYC^{+}23, RPG23b, TFE22]. Semantic [CSL^{+}23, EAKC^{+}20, GSK^{+}20, HZZ^{+}20, HZS^{+}22, HCL20, HCH^{+}23, KGB22, LVV^{+}21, LS22, MYS^{+}22, PJYW21, SLL21, SJL^{+}23, WL20, XHL^{+}24, ZCSS23]. Semantic-Aware [SJL^{+}23]. Semantical [MLC^{+}20]. Semantically [HPP^{+}20, LWL^{+}23a]. Semantics [FBW21, KGWD22, LCC^{+}20, XLF^{+}23]. Semantics-based [FBW21]. Semantics-preserving [KGWD22]. Semi [CMKK22, CWW^{+}21, MOA21, MWJ22, SWF^{+}24, SKH^{+}23]. Semi-Automated [SKH^{+}23]. Semi-Automatic [CMKK22]. Semi-Autonomous [SWF^{+}24]. Semi-Repetitive [MOA21]. Semi-Supervised [CWW^{+}21, MWJ22]. sensation [MSAM^{+}22]. Sense [DGB^{+}22, DPM24, FALH20, FOH^{+}21, NIM^{+}21, RVB^{+}22]. Sense-Making [RVB^{+}22]. Sensemaking [DLW^{+}23, LLS^{+}20, PPE23, GSL21]. Sensing [BFY^{+}24, iKYOW23, dPLM21]. Sensitive [KBF22, WBL^{+}22, YKJ^{+}23]. Sensitivity [ATK^{+}24, HLM^{+}20, LKS22, PBM^{+}24, RBF^{+}23, XMT^{+}21]. Sensor [GWC^{+}23, HSC^{+}22, ZX22]. Sensors [NMS^{+}23, RZH^{+}23, YLL^{+}20]. Sensory [CPR^{+}22, CRJ^{+}24, FMBN23, GMVRB20, KG24, MSAM^{+}22]. Sentence [KTL24]. Sentiment [WHJ^{+}22]. Separate [XLW^{+}24]. Separating [DWOB20]. Separation [JdJTC24, SLJ^{+}20, XQXL23, TWA22]. Separations [MM21]. Sequen [MSM^{+}22]. Sequen-C [MSM^{+}22]. Sequence [BZSD21, CSL^{+}23, CYP^{+}20, GZW^{+}20, GJC^{+}22, GGJ^{+}22, JGC^{+}21, MXC^{+}20, RBLT^{+}22, YDMP22]. Sequences [BZSD21, GJC^{+}22, MSM^{+}22, MOA21, QBW^{+}20, QLFG22, WLGW23, YFS22, CNB^{+}22]. Sequential [FWM^{+}24]. Serendipitous [NKWW22]. Series [BBSvL24, DWX^{+}22, DCS^{+}24, FiMH21, FSS^{+}21, HSYZ24, MBS^{+}21, MZS^{+}24, SCC^{+}24, YKJ^{+}23, ZWZ^{+}22]. Serious [CB22]. Service [Ano22c, Ano23a, Ano24d, FGS^{+}21, Kiy22a]. Session [BK22]. Sessions [FWZ^{+}20]. Set [CNY22, HSV^{+}20, JGG21, KBV24, LLWF23, LWF23, NRA^{+}23, RWB^{+}23, WJKN21, WDN23, WCW^{+}22]. Set-Typed [CNY22]. Sets [GBNH21, HCX^{+}21, JWKN21, JH20, KNM^{+}21, KLKE21, LT20, YBOB24]. Setup [KKW23]. SG [LB22]. SG-GAN [LB22]. Shader [LSW^{+}23]. Shading [JHS^{+}21, MBB20, SLJ^{+}20, YWB20]. Shadow [LZX^{+}22]. Shadowless [HIK^{+}23]. ShadowMover [YGH^{+}23]. Shadows [ASCR^{+}22, JRM22, YGH^{+}23]. Shallow [ZHL^{+}21b]. Shape [CSBK20, DXX^{+}21, GLL^{+}22, HB24, HCX^{+}21, HL21, LWWF21, LWF23, LLL^{+}22, LCK^{+}21, LGL^{+}23b, SYW^{+}20, SSX^{+}20, SOL^{+}22, WCZ^{+}20, ZJJH21, ZJS^{+}22]. Shape-aware [WCZ^{+}20]. Shape-changing [ZJS^{+}22]. Shape-Driven [HCX^{+}21]. Shaped [WTY^{+}22]. Shapes [KBV24, Liv21, TGM21, WL20]. ShapeWordle [WCZ^{+}20]. Share [JVRL24]. Shared [ENK^{+}20, GWD^{+}24, LHC^{+}21, MMN^{+}22]. Sharing [FiMH21, FOH^{+}21, RDHH21]. Shear [NWMC23]. Shedding [ASCR^{+}22]. Sheets [WSZ^{+}20, WTY^{+}22]. Shelf [ZHKY23]. Shells [HS23]. Shielding [MMN^{+}22]. Shift [MEHD24]. Shock [MEHD24]. Shoes [DPM24, TWT^{+}22]. Shooting [FS24, PGS21]. Shortcuts [SDK23b]. Shortening [YWM^{+}23]. Shot [JLCZ22, TXM23, YYZ^{+}22]. Show [CAGM22]. Showing [MJAD22, ZBG^{+}24]. Shutter [DMTD22]. ShuttleSpace [YCC^{+}21]. Sibyl [ZLVV22]. sick [TB24]. Sickness [DCWD23, MCS^{+}23, NDLW20, RBR20, WSZ^{+}23, WGS^{+}24, ZOF^{+}23]. Side [BBL^{+}22]. Side-by-Side [BBL^{+}22]. SightBi [SSAZ22]. Signals [GTH20, MHIS23, ZGX^{+}23]. Significance [LqZ23]. Significant [Ano22d, Ano23d, Ano24e, GF22]. Significantly [CBW23]. Silent [LMM^{+}21]. Sim [RXX^{+}21]. Similar [KPL23]. Similarities [KLCK20]. Similarity [DGDC21, MTW^{+}20, OKM21, RAC22, RBLT^{+}22, SUS^{+}21, TFE22]. Simple [FE21, LWC22, MSS21, LHWW22]. Simplicial [GLX^{+}21]. Simplicity [LLC^{+}20]. Simplification [LGMT21, SLC21, XXM^{+}21]. Simplifications [ZRPW23]. Simplified [LCSA22]. SimuExplorer [LZW^{+}23]. Simulated [CS23]. Simulating [BBM^{+}21, GLHQ21, HWC23, HDFK21, IWT^{+}20, LLCH22]. Simulation [AYG^{+}22, BMWD20, CK22, CLH^{+}23, CLFL22, CFQ21, CQ22, FLX^{+}23, GWL^{+}22, GWW^{+}21, HLW^{+}20, HHB^{+}23, LZW^{+}23, LCH^{+}21, LLD^{+}21, LGH^{+}24, LWK20, LCCZ22, MDX^{+}23, RXX^{+}21, SHM23, ŠK20, WGH20, XZWY20, ZHQH20, ZHL^{+}21b, dSBdO^{+}24]. Simulation-Based [AYG^{+}22]. Simulations [ABE^{+}22, FGS^{+}21, HWG^{+}20, HZP^{+}24, NFN^{+}23, SXW^{+}22, SXL^{+}23]. Simulative [WZD^{+}20]. Simulator [DCWD23, WSZ^{+}23]. Simulators [CBHR^{+}23, GF24]. Simultaneous [HHKN21, HL21, LMD^{+}22, vBMS22]. Simur [WZD^{+}20]. Sine [BZW^{+}21]. SineStream [BZW^{+}21]. Single [CLL^{+}20, CKQ^{+}23, DLP^{+}23, KPW20, LWWH20, LZZ^{+}20, MDL^{+}23, PBAG23, WZH20, XZXY22, XLL^{+}20, YCZ^{+}22, YHC^{+}22, ZZW^{+}20, ZYC^{+}23, ZLZ^{+}23b, SIL^{+}21]. Single-Bounce [XZXY22]. Single-Cell [CKQ^{+}23, SIL^{+}21]. Single-View [MDL^{+}23, WZH20]. Sitting [MZX^{+}21]. Situ [BLE^{+}23]. Situated [APHD24, BKT^{+}22, FCH^{+}23, LSS24, LYBP23, WZW^{+}23, PWB21]. Situatedness [BKT^{+}22]. Situation [BKT^{+}22, FTWP22, ZLL^{+}20]. Situational [SLK^{+}20, WR23]. Situations [BXQ^{+}22]. Size [HJL^{+}23, KKH21, WFK24, WBPC23]. Size-based [HJL^{+}23]. SizePairs [HJL^{+}23]. Sizing [BCB22]. Skeleton [COFJ23, LWC22, LZL^{+}23, QHL^{+}20, YISG21, ZZD^{+}23a]. Skeleton-based [LZL^{+}23]. Skeleton-Detail [ZZD^{+}23a]. Skeletons [CWL23]. Sketch [DLP^{+}23, FiMH21, LZZ^{+}21, LZLS22, LLY^{+}24, OBCT24, SZF^{+}21, XXM^{+}21, ZJJH21, LLS^{+}20, LZZ^{+}21]. Sketch-Based [FiMH21, OBCT24, SZF^{+}21]. Sketch-Guided [LZLS22]. Sketch-R2CNN [LZZ^{+}21]. Sketch2PQ [DLP^{+}23]. Sketches [MMG^{+}21]. Sketching [DHF^{+}22, HHD^{+}20, XZS^{+}23]. Sketchplan [SPJ^{+}23]. Ski [RBKM24]. Skill [KMH^{+}23, OSC22, WPNK21]. Skills [KAL^{+}23]. Skin [PGE^{+}22]. Skinned [FLW^{+}21]. SkiVis [RBKM24]. SLAM [DHM^{+}22, ZYL^{+}24]. slant [TWA22]. Slice [ZOS^{+}23]. Slice-Driven [ZOS^{+}23]. SliceTeller [ZOS^{+}23]. Slicing [GTC^{+}23, LMD^{+}22]. Slicing-Tracking-Detection [LMD^{+}22]. Smale [MLT^{+}24, SPN23]. Small [BLIC20, CCM20, JWE^{+}22, LZC^{+}21, STD^{+}23, YPW23, ZHL^{+}21a, ZH20]. Small-Scale [ZH20]. Smart [SHC^{+}20]. SmartCube [LWSY20]. SmartSpring [ZZS^{+}23]. Smelling [WMB23]. Smile [LSZC21]. sMolBoxes [UWF^{+}23]. Smooth [TG24]. Smoothing [MNZ^{+}20, RQ21, ZHF^{+}20]. Snapping [KGB22]. Snapshots [CSJ^{+}21]. SNE [CMK20, MvdEPV24, PTM^{+}20, WCJW22, vdRBE22]. Soccer [CLX^{+}23, CXZ^{+}24, SPJ^{+}23, XWL^{+}21]. Social [AI21, CLCY20, CFDN24, HGB22, HX23, ISBP22, JJKJ20, KKZE20, KPL23, KKT^{+}22, LNB^{+}21, LLSM24, LCW^{+}23a, PBF^{+}21, YeSiK^{+}23, ZC23]. Socially [LVR^{+}24]. Socio [WDSM24]. Socio-Technical [WDSM24]. Socrates [WGH^{+}24]. Soft [CLZ^{+}24, LLK22, ZJS^{+}22]. Soft-Correspondence [CLZ^{+}24]. Softer [MMK^{+}23]. Softness [CZGF21]. Software [KKJ^{+}21, Ano22a]. Solid [HS23, PSG^{+}22]. Solid-Shells [HS23]. Solid-state [PSG^{+}22]. Solids [SOL^{+}22]. Solution [BHU^{+}21, YXS^{+}23]. Solutions [JLK20]. Solver [GLHQ21, LGH^{+}24, XZWY20, XWZ^{+}23, YLL^{+}22a]. Solving [KLSB22, LDT^{+}21, XVW^{+}21]. Somnus [XFD^{+}23]. Sort [LGWL21]. Sorted [ZWP21]. Sorting [LGWL21]. Sound [CAR^{+}23, CRJ^{+}24]. Source [BB21, MNK23]. Source-monitoring [MNK23]. Space [BCC^{+}20, BCB22, CPD20, CSL^{+}21, CFGM22, DLW^{+}23, ENXS21, FCFC22a, FCFC22b, FSN20, GKN^{+}23, HWG^{+}20, KHR21, KAS^{+}22, LJS21, LCC^{+}20, LSL^{+}23, LY23, LWL^{+}23b, MA20, MJAD22, MBHE24, PBBH20, PLL^{+}24, PAPB20, PMCM24, QCCC23, RBLE22, SDC^{+}24, SXW^{+}22, SXL^{+}23, eSYKW23, XPK^{+}24, YXL^{+}22, ZJW21, ZWG^{+}23]. Space-Adaptive [eSYKW23]. Space-Filling [KAS^{+}22, ZJW21]. Space-Time [FSN20]. Spacecraft [TSH21]. Spaces [AAA^{+}21, BvdPLH22, EAKC^{+}20, FLS23, GWD^{+}24, LHC^{+}21, QFWS22, RAC22, RMB^{+}21, WYS^{+}22, XLW^{+}24, YLTL23]. Spacetime [GLX^{+}21, PBBH20]. Spacing [FFB24]. Spambot [KKZE20]. Spanning [AA21]. Spark [XHL^{+}24, HCJ21]. Sparse [GLY^{+}21, RZH^{+}23, SHL^{+}21, WYZ^{+}21, ZSL21]. Spatial [BBVS^{+}24, BBSC23, CJY^{+}23, CWO^{+}24, DPR^{+}20, DCS^{+}24, DST^{+}23, GST^{+}24, HKB^{+}22, HW22, JZHA22, KHI20, KVB20, KJL24, KPW20, LMF^{+}24, MCFKF24, NIK24, NSW24, NRA^{+}23, NdCS21, PIS20, RBR20, RR23, SYW^{+}20, eSYKW23, TARB23, UIHS20, VBV^{+}23, VLM^{+}23, WLZ^{+}23, WKN^{+}23, WN21, XHFZ24, ZZW^{+}22a, ZCL^{+}21, ZZX^{+}22, ZLG^{+}21a, ZIX^{+}24, KOL^{+}20, SIL^{+}21]. Spatial-Temporal [WLZ^{+}23]. Spatialization [ACT^{+}24]. Spatializing [CRJ^{+}24]. Spatially [PSY^{+}20, SYC^{+}23, XLL^{+}20]. Spatio [NdCS21, SKR^{+}24, VCO^{+}23, WHSZ21, ZTC^{+}23, ZGL^{+}21]. Spatio-Angular [ZGL^{+}21]. Spatio-Temporal [NdCS21, SKR^{+}24, VCO^{+}23, WHSZ21, ZTC^{+}23]. Spatiotemporal [DWL^{+}22, HZCW22, LBW^{+}22, LWSY20, MFS^{+}23, PFC20, TFE21, TFE22, VFP24]. Speak [LLSM24]. Speaking [HHM^{+}24]. Special [BSW20, CTW21, CRZ22, CRH23, MB20a, MB20b, MB21a, MB22a, SK23a]. Specific [CFGM22, KVM^{+}22, LWF23]. Specification [LRHA22]. Specifications [NSS21]. Speckle [CZT^{+}21]. Spectral [CSBK20, MBB20, TYPC20]. Specular [FWW^{+}23, JRM22, Tet24]. Speech [AV21, CML24, CBW23, LHL^{+}22, SSS20, WGO20, WGO22, WO22, YLS^{+}23, SSS20]. Speech-Driven [LHL^{+}22]. Speeches [MHS^{+}22]. SpeechMirror [HHM^{+}24]. Speed [DMTD22, NL24, UIHS20]. SPEULER [KGWD22]. SPH [BKWK20, LHWW22, RHLC22]. Sphere [MSWI22]. Spherical [CZC^{+}20, MBBB21]. Spinal [EBKL21]. Spine [CLL^{+}23, EBKL21]. SPinPong [WPNK21]. Spins [ZBG^{+}24]. Spirals [WCZ^{+}20]. Splat [NSG^{+}20, NNF^{+}20]. Splatting [PBBH20]. Splines [MNPP23]. SplitStreams [BNRB21]. Splitting [TYW^{+}22]. Spoon [BCC^{+}20]. sPortfolio [YBL^{+}20]. Sporthesia [CYX^{+}23]. Sports [CYC^{+}22, CYX^{+}23, WMH^{+}23, WLG^{+}22, WLGW23]. SpotSDC [LMM^{+}21]. Spotter [SDMK22]. Spotting [CML24, SDMK22]. Sprawl [LIDM20]. Sprawlter [LIDM20]. Spreading [ZCL^{+}21]. Spreadsheet [DWH^{+}23, SXS^{+}21]. Spurious [TAL24]. Square [CMXF21]. Squeeze [BVV^{+}23]. Srinivasan [Ano22a]. SSIM [WHL^{+}20]. SSR [HW22]. SSR-TVD [HW22]. SSRNet [YFM^{+}23]. stability [VAWL24]. Stabilization [MNZ^{+}20]. Stable [BTL23, HJL^{+}23, LWC22]. Stacked [BKS22, CZMR21]. Stacked-CNN [CZMR21]. StackGenVis [CMKK21]. Stacking [CMKK21]. Stacks [ISKM23]. Stage [MLC^{+}20]. Stage-wise [MLC^{+}20]. Staged [CSCM21]. Staggered [ZHQH20]. StainedSweeper [HHI24]. StainedView [KHI20]. Stakes [ZLVV22]. Standardized [SBK^{+}20]. Standing [MCQ23]. Star [HCX^{+}21, ROM^{+}23, vOVR23]. Startup [ZP24]. State [Mue20c, PSS^{+}22, RCD^{+}23, WMH^{+}22, WH23, MSAM^{+}22, PSG^{+}22]. State-based [WMH^{+}22]. State-of-the-Art [PSS^{+}22, WH23]. Statement [Ano22s, Ano23q, Ano24o]. Statements [CZW^{+}20]. Static [CCPM23, FPG^{+}23, HRX23, KW23, LFW^{+}22, WBM21b]. Stationary [KNL23]. Stations [CWO^{+}24]. Statistical [AMY^{+}22, KH21, NRA^{+}23, NSK^{+}21, RRG23, SHOP23]. Statisticians [NE24]. Statistics [GGPL20, NE24, PFC20, RLLS20]. Stats [CSL^{+}21]. Status [COZ^{+}23]. STBins [QBW^{+}20]. STD [MDL^{+}23]. STD-Net [MDL^{+}23]. Steadier [MCQ23]. Steady [RG20]. Steering [Ano21s, Ano22n, Ano22r, Ano22-31, Ano23l, Ano23p, Ano23-28, Ano24n, Ano24x, ATHI24, BBM^{+}21, CHSC24, LHA^{+}21, MXC^{+}20, SRBP20, TXM23, YWL^{+}21, ZFF22]. Steering-based [LHA^{+}21]. Steganalysis [ZCZ^{+}21, ZZC^{+}22]. Steganography [ZZC^{+}22]. Steiner [MXT^{+}22]. Step [CBL^{+}24, SHD^{+}20, SLJ^{+}20, ZCZ23b, XVW^{+}21]. Stepping [DPM24]. Steps [DGD^{+}23]. Stepwise [CMKK22]. Stereo [MP21, YWB20, TWA22]. Stereopsis [MWR^{+}22]. Stereoscopic [DCWD23, KIPS21, SPWW^{+}24]. Stereotyping [HX23]. Stick [ZP23]. Stiffness [BPL23, PIS20]. Still [LSE20]. Stimulation [GTH^{+}22]. Stimuli [GMVRB20, JLM^{+}21, KL22, MGM^{+}22, WMB23]. Stitching [ZLZ21]. STNet [HZCW22]. Stochastic [AKS22, ADD^{+}22, SWW^{+}22, Yuk21]. Stock [YBL^{+}20]. Stories [LSW^{+}22, SGJC23, SBe^{+}21, YXL^{+}22, ZSL^{+}22]. Story [CLA^{+}20, ISBP22, SXS^{+}21, SCC^{+}23, WGH^{+}24]. Storyline [HPAB23, TLW^{+}21]. Storylines [HPAB23]. Storytelling [CLA^{+}20, DKM24, eSYKW23, SKH^{+}23]. Strategic [WXG^{+}24]. Strategies [BS22, BER^{+}23, CSCM21, DLW^{+}23, KKH21, LWT^{+}23, ZWZ^{+}23]. Strategy [AYGR22, CWW^{+}23, CvW23, LJCL24]. StrategyAtlas [CvW23]. Stratified [YBL^{+}20]. STRATISFIMAL [dBRGD22]. Stream [HTW20, ZTWW21]. Streamable [SCL^{+}23b]. Streamgraphs [BZW^{+}21].

Streaming [CDS^{+}22, CZF^{+}22, FBW21, FCS^{+}20, HRX23, LDB^{+}21, LLDW24, LCW^{+}23b, SFS^{+}22, WIP^{+}24, XHS^{+}24, DJHBJ21]. Streamlines [HTW20, PL21]. Streams [CAA^{+}21, LCC^{+}23]. Street [GZRP^{+}22]. Street-Level [GZRP^{+}22]. Strength [ZCM^{+}23b]. Stress [PL21, XWH^{+}23, ZCL23]. Striking [SSC^{+}23]. Stroke [CNK^{+}24]. Strolling [WLC^{+}23]. Strong [YRL^{+}20]. Strongly [KLB24]. Stroop [PT20]. StructGraphics [Tsa21]. Structural [EBPF21, YWM^{+}20]. Structure [ATAS21, AYA^{+}23, BPA22, COZ^{+}23, EHA^{+}23, HPT^{+}23, HA20, HSC^{+}22, LIFD23, MDL^{+}23, NWMC23, SaT^{+}23, WXW^{+}20, WWW^{+}24, XLY^{+}22, YXL^{+}22]. Structure-Aware [HSC^{+}22, WXW^{+}20]. Structure-borne [SaT^{+}23]. Structure-Preserving [MDL^{+}23]. Structured [BS21, CJS^{+}22, SII^{+}21, WZU^{+}21]. Structures [AA21, HWL^{+}22, HBMK23, KGBP20, KLB24, LI24, NMC21, Tsa21, WFY^{+}21, WWG21, ZJC^{+}21]. Student [HXL^{+}24]. Studies [QR22, WRZ^{+}21, XZKM22, ZYM^{+}24]. Study [ATK^{+}24, CB22, CPR^{+}22, EBJ^{+}22, ESB^{+}24, FPG^{+}23, FPK^{+}24, FKSP^{+}24, GVN^{+}20, HKB^{+}22, HCWK23, HPdlG20, HMH^{+}21, HJZ^{+}21, JZHA22, KBM21, LST^{+}22, LGY^{+}22, LKS22, MMF20, MMS^{+}23, MSGM23, MD20, NE24, PCQ^{+}20, PVPM22, RS23, RBLE22, RDH23, RPH^{+}21, TARB23, TPH22, WHC^{+}23, WYIS24, WBI20, XZS^{+}22, XDBAR24, YSM^{+}20, KHD^{+}22]. Studying [BPL23, FALH20, HPK^{+}22, HT22, KL22, PRJ^{+}23, XNF^{+}23]. Stuttgart [Ano22e]. Style [CSWZ24, CFZZ23, HA20, HLH^{+}23, JYF^{+}20, LTC21, SLC^{+}23, SYX^{+}22, SJL^{+}23, ZXC^{+}23, McN23]. Style-Content [CSWZ24]. StyleGAN [SYC^{+}23]. Stylization [HYH^{+}23]. Subdivision [CGZ^{+}20]. Subjective [CWSJ23, DPR^{+}20, LPG^{+}22, MTW^{+}20, NDF^{+}21]. Subsequent [EHB^{+}23]. Subset [BWZ^{+}20]. Subspaces [BXZ^{+}21, FV24]. Substitution [CPR^{+}22, CRJ^{+}24, KG24]. Subsurface [HFS^{+}21, NI22, SFSA24]. Success [WPL^{+}22]. Suggestions [WSL^{+}20]. Suggestive [LZ21]. Summaries [CSJ^{+}21, RPD21]. Summarization [PTD^{+}21, PDD^{+}22, WBS21, ZSW^{+}21]. Summarizations [HPRC20]. Summarizing [YLTL23]. Summary [AMY^{+}22, VFP24]. Summative [KKEG20]. Summit [HPRC20]. Super [Ano24l, CJY^{+}23, HW20, HZCW22, HW22, LWWY21, SS24, WCTW21, WGS^{+}23, ZGL^{+}21]. Super-Pixel [LWWY21]. Super-Resolution [CJY^{+}23, HW20, HZCW22, HW22, WCTW21, ZGL^{+}21]. Supercommittee [Ano23n]. Supercomputing [Ano22c]. Superhuman [GVN^{+}20]. Superpixel [FWZZ20]. Superpixel-Based [FWZZ20]. Superpowers [WAC^{+}22]. Superquadric [PL21]. Supervised [CWW^{+}21, HYC^{+}23, KBJ^{+}20, LFO23, MWJ22, RPG23b, SSX^{+}20, SLR21, TFE22, ZHDX20]. Supervision [CWW^{+}22]. Support [BD22, CBB23, EJS^{+}23, FWZ^{+}20, HSJ^{+}20, KWH22, LYBP23, MSY22, MSBV^{+}23, PSS23, RBF^{+}23, SESH24, SNH^{+}23, SLK^{+}20, XLY^{+}22, ZeB^{+}21, ZG20, vOVR23]. Support-Free [XLY^{+}22]. Supported [LAS^{+}20, WBG^{+}22]. Supporting [CLA^{+}20, DGD^{+}23, FKM20, LDT^{+}21, SCST24, SLC^{+}23, SCC^{+}24, WCZ^{+}23, WAV^{+}21, ZLC^{+}22]. Suppression [CZT^{+}21]. Surf [LYZ^{+}24, MNPP23]. Surface [AMAS21, CPD20, CLG21, DHF^{+}22, HZFH20, HGSP20, KTL24, KPF^{+}22, LYZ^{+}24, LHL^{+}20, MNPP23, MML21, XF21, XZS^{+}23, YFM^{+}23, ZGX^{+}22, ZH20]. Surface-Adaptive [XF21]. Surfaces [CYD^{+}23, DKuH20, HTW20, HZYZ22, IGM24, JH20, KRZ^{+}20, KHR21, LHC^{+}21, QRZZ21, WWZ23b, XCSJ22, ZTWW21, ZLC^{+}22]. Surfacing [ZWG^{+}23]. Surfel [MA20]. Surfel-Based [MA20]. SurfRiver [ZTWW21]. Surgery [EBKL21, LCH^{+}21]. Surgical [KZD^{+}23]. Surrogate [YBOB24, SXW^{+}22, SXL^{+}23]. Surround [GBL^{+}22, MP21]. Surround-Screen [GBL^{+}22]. Surround-View [MP21]. Surrounding [YIIW24]. Survey [AMF20, BKT^{+}22, COFJ23, CGAG20, DV23, EMK^{+}21, FYC^{+}23, FP21, GLH22, GGJ^{+}22, LLMA22, McN23, PSS^{+}22, QR22, SWB^{+}22, SSL^{+}23, ŠK20, TG24, TLW^{+}23, WCWQ22, WH23, WHY^{+}23, WWS^{+}22, YJ21]. susceptibility [TB24]. Suspicious [LWW^{+}21]. SVBRDF [Tet24]. SVG [CLW^{+}24]. SVGs [SSB^{+}22]. Swaying [YCM^{+}24]. Sweep [UIHS20]. Sweeping [HHI24]. Swift [ZSZ^{+}24]. Swift-Eye [ZSZ^{+}24]. Swimming [YVBI24]. Switching [APP^{+}22, LJCL24]. Symmetric [HZZ24, QRZZ21]. Symmetry [GZM^{+}21, QGL^{+}23]. Symposium [Ano23g]. Symptoms [FNB^{+}22]. Synchronization [PTS^{+}20, SSZ^{+}21]. Synchronous [BK22, LFCH24]. Synthesis [AYA^{+}23, CLA^{+}20, GXH^{+}24, HWG^{+}20, LZK^{+}22, MBHE24, PBAG23, QLC^{+}24, RPG23b, SC22, TXM23, WCX21a, WCC^{+}21, ZZX^{+}22]. Synthesize [MSMX23]. Synthesizing [HZCW22, HGO21, LHS^{+}22, LCC^{+}23, QLFG22]. Synthetic [PCJ23, TNJ^{+}22, CNB^{+}22]. System [BBRE24, BVV^{+}22, BVV^{+}23, BAC^{+}20, CXD^{+}21, DMBK21, EGP^{+}21, GM23, GBL^{+}22, GZL^{+}21, HPK^{+}22, HZP^{+}24, HHM^{+}24, JCZ^{+}24, JLP^{+}23, JSS^{+}20, KCGZ23, KH21, LKJ^{+}20, LELT23, LCH^{+}21, MHS^{+}22, PLW^{+}23, PRKM20, QTW^{+}24, RNO^{+}22, SPZS24, SFS^{+}22, SJK^{+}23, TNZ^{+}24, WDSM24, WMH^{+}23, XNF^{+}23, XZKM22, YITS23, YCC^{+}23, YS20, YHD21, ZZ23, ZYL^{+}22, ZWW^{+}23b, ZKS^{+}20, ZCM^{+}23b]. Systematic [BFS^{+}23, GEU^{+}22, IMQ^{+}20, KPF^{+}22, LSG24, LCC^{+}20, LRZ^{+}23, MGM^{+}22, PCL24, PMN^{+}23, PMW23, SDC^{+}24, SNK^{+}22, THFI21, WBA^{+}23, MGC^{+}21]. Systems [AYG^{+}22, BFS^{+}23, DLH^{+}22, HL21, LLS^{+}20, LMM^{+}21, LDT^{+}21, LALG22, MYS^{+}22, PMN^{+}23, RXX^{+}21, RWB^{+}23, SLR^{+}20a, WLH24, WDC^{+}23, XWY^{+}20, ZSL21].

T [WHL^{+}20, CMK20, MvdEPV24, PTM^{+}20, WCJW22, vdRBE22]. t-SNE [CMK20, MvdEPV24, PTM^{+}20, WCJW22, vdRBE22]. T-SSIM [WHL^{+}20]. t-viSNE [CMK20]. Table [Ano22t, Ano22u, Ano22v, Ano22w, Ano23s, Ano23t, Ano23u, Ano24q, Ano24r, JBS^{+}22, KBM21, LZW^{+}23, OSC22, WZD^{+}20, WWC^{+}21, WPNK21]. Tableau [Ano22a]. Tables [BCT22, CLL^{+}21]. Tabletop [EGP^{+}21, LNB^{+}21]. Tabular [CWH^{+}23, DWH^{+}23, LLW^{+}23, MCW^{+}20, SDR22, WSZ^{+}20]. Tac [WZD^{+}20, WWC^{+}21, WMH^{+}23]. Tac-Miner [WWC^{+}21]. Tac-Simur [WZD^{+}20]. Tac-Trainer [WMH^{+}23]. Tackling [GM23]. Tactic [TPH22, WZD^{+}20, WWC^{+}21]. Tactic-based [WZD^{+}20]. Tactical [AAA^{+}21, SPJ^{+}23]. TacticFlow [WLG^{+}22]. Tactics [CXY^{+}22, WLG^{+}22]. Tactile [JLM^{+}21, MHIS23, SWSK23, ZLS21]. TactualPlot [CRJ^{+}24]. Tag [KKE21]. Tai [TNZ^{+}24]. Tailored [BvOR21]. Tailoring [WCZ^{+}20, WTD^{+}21]. Takeaways [SSC^{+}23, XSB^{+}22]. Taking [HPdlG20]. Talk [JWE^{+}22]. Talking [LH23, PTC^{+}24, ZNF^{+}23]. Talks [WQ20]. Tangible [EGP^{+}21, dPLM21]. Tangibles [BPL23]. Tanglement [CK22]. Tap [KNL23]. TapeTouch [ZJS^{+}22]. Target [CHSC24, ENK^{+}20, HAK^{+}23, LHA^{+}21, TME^{+}22, WBI20, XWH^{+}23, YZN^{+}20, ZIX^{+}24, ZYR^{+}20]. Target- [ZIX^{+}24]. Target-based [LHA^{+}21]. Targeting [YSM^{+}20]. Targets [SPNG23, WBG^{+}23, XLL^{+}22, YLF^{+}20]. Task [BvdPLH22, BS22, BCN^{+}20, BPL23, CRM^{+}24, DFP^{+}20, DKuH20, DTPG21, ENK^{+}20, HKB^{+}22, HBLW21, JKL24, KRS^{+}22, KVM^{+}22, KPO^{+}23, LFCH24, LI24, MMS^{+}23, NP24b, PCQ^{+}20, PSS^{+}22, QR22, SBT^{+}23, SMS^{+}22, SWS^{+}23, TARB23, VVR^{+}24, WG21, WBI20, YTHL23, YeSiK^{+}23]. Task-Agnostic [BvdPLH22]. Task-Based [DFP^{+}20, DTPG21, SMS^{+}22]. Task-Centric [YeSiK^{+}23]. Task-Dependent [MMS^{+}23]. Task-Oriented [HBLW21, KRS^{+}22]. Task-Parallel [LI24, WG21]. Tasks [BS22, BBM^{+}21, CFDN24, DPR^{+}20, DS22, GWL^{+}22, HAK^{+}23, HBLW23, KZD^{+}23, KWO^{+}20, LETF21, PSS^{+}22, VLM^{+}23, WN21]. Taurus [XWZ^{+}23]. Tax [LWW^{+}21]. Taxonomies [LPP^{+}23, PMW23]. Taxonomizer [MM20]. Taxonomy [DFP^{+}20, DV23, LST^{+}22, MSA^{+}22, ZLG^{+}21b]. TaxThemis [LWW^{+}21]. Taylor [JKU^{+}22, NMC21]. Team [CLX^{+}23]. Team-Builder [CLX^{+}23]. Teaming [HMTI24]. Teams [GF24]. Technical [Ano20j, Ano20k, Ano21g, Ano21f, Ano22o, Ano22s, Ano22e, Ano23m, Ano23q, Ano23e, Ano24j, Ano24o, Ano24f, Fek21, Kiy22b, WDSM24, ZG20]. Technique [BRLR24, CGZ^{+}20, FE21, PTS^{+}20]. Techniques [CCS^{+}21, EMK^{+}21, HMGO23, HTJ^{+}20, HSJ^{+}20, JKW^{+}22, KNM^{+}21, KJL24, LHA^{+}21, LAN21, LTLB22, LCM^{+}23, MGWK^{+}22, MTVS23, RCD^{+}23, RQ21, SFPCW23, TIDQ24, VVC^{+}24, WMMB23, WQ20, XZS^{+}22, YLF^{+}20, YGE^{+}21, ZCZ^{+}23a, ZLG^{+}21b]. Technology [Ano22m, Ano22b, Ano23k, BS21, FGL^{+}23, HKVZ20, IMKP21, IGMW22, PCL24, WQQ^{+}24]. TED [WQ20]. Teeth [CFZZ23, ZCSS23]. TeethGNN [ZCSS23]. Telecollaboration [RTM^{+}20]. Teleconsultation [YGE^{+}21]. Teleoperation [EHB^{+}23, FE21]. Teleport [BC21]. Teleportation [RBR20, RTM^{+}20, WBG^{+}23, WFK24]. Teleporting [AZA^{+}23, KOL^{+}20, KHD^{+}22]. Telepresence [ASP^{+}22, HRX23, KBF22, WYS^{+}22, YGE^{+}21]. Telling [SBe^{+}21]. Temporal [BZSD21, CSJ^{+}21, EEL^{+}20, FAS^{+}21, HW20, HJL^{+}23, HRS^{+}22, KW23, LZH^{+}21, LPP^{+}23, MSM^{+}22, NdCS21, QBW^{+}20, SKR^{+}24, SAPW23, VCO^{+}23, WHSZ21, WLZ^{+}23, WPNK21, XWPG^{+}24, ZWXM23, ZTC^{+}23, CNB^{+}22]. Tennis [LZW^{+}23, OSC22, PJHY20, WZD^{+}20, WWC^{+}21, WPNK21]. Tensor [BRLP20, HWN^{+}23, HZYZ22, HZZ24, KRZ^{+}20, PL21, QRZZ21, ZCZ^{+}21]. Tensors [HNS23]. Term [DHM^{+}22]. Terms [CSL^{+}21]. Tessellation [CZGW23, NRA^{+}23]. Test [SKFZ22, CNAA^{+}22]. Testbed [CCP^{+}21]. Testing [LELT23, NCB^{+}21, SHOP23, VGK^{+}22]. Tests [VC20, WW22a]. Text [AV21, ACT^{+}24, BTHL23, DAB^{+}23, DZG^{+}23, FWW^{+}24, FMBN23, JMK^{+}22, KNL23, KTL24, KG24, LZK^{+}22, LWL^{+}22a, LWY^{+}22, MWJ22, NL24, OKM21, SSC^{+}23, SKK^{+}22, XHL^{+}24, YJ21, SNBC23, CZW^{+}20, DAB^{+}23]. Text-Based [OKM21]. Text-Entry [KG24]. Text-to-Image [FWW^{+}24, XHL^{+}24]. Text-to-Viz [CZW^{+}20]. Textiles [MDX^{+}23]. Texts [LCC^{+}20]. Textual [CSC^{+}21]. Texture [DGDC21, FYL^{+}23, PJYW21, RPG23b, WFC21, ZX22, ZHF^{+}20]. Textured [LCK^{+}21]. Textures [HZII24, LXH^{+}21, MSY22, SMCL24]. Texturing [ZYC^{+}23]. THALIS [FNB^{+}22]. Their [GLHQ21, GKN^{+}23, ISBP22, LCW^{+}23a, RSD^{+}23]. Them [McN23].

Theory [MYS^{+}22]. Therapy [CNK^{+}24, FNB^{+}22, LZZ^{+}24, WHL^{+}20]. There [BCC^{+}20, FW22, MGWM22, WBF20]. Thermal [SWSK23]. Thin [BJCL21, HS23, KLB24, XWZB21]. Thin-Layer [KLB24]. Think [DLW^{+}23, FWZ^{+}20, SKFZ22]. Think-Aloud [FWZ^{+}20, SKFZ22]. Thinking [HHC^{+}21, PLD^{+}23]. ThinVR [RSAA20]. Third [HWN^{+}23, SAMB^{+}23]. Third-Order [HWN^{+}23]. Third-Person [SAMB^{+}23]. Thirty [HCWK23]. Thirty-Two [HCWK23]. Thorns [FWM^{+}24]. ThreadStates [WMH^{+}22]. Three [TBW^{+}23, ZZC^{+}22]. Three-Dimensional [ZZC^{+}22]. Thresholds [BPL23, ONH21, SPWW^{+}24, ZLS21]. Throughput [BS22, PGL^{+}20]. Thumbnails [VH23]. Tight [BFS^{+}24]. Tightly [BXQ^{+}22]. Tightly-Coupled [BXQ^{+}22]. Tile [YWB20]. Tileable [RPG23b]. Tiling [KAS^{+}22]. Tilt [YDM^{+}21]. Time [BWM24, BBSvL24, CCPM23, CIA24, CSC^{+}21, DWX^{+}22, DCS^{+}24, FiMH21, FSN20, FSS^{+}21, GZY^{+}22, HW20, HZX^{+}21, HW22, HW23, HSYZ24, HCH^{+}23, IGM24, IGMM22, JBS^{+}22, KKE21, KKT^{+}22, KBPR22, LAN21, LCC^{+}20, LWSY20, LXL21, LZX^{+}22, LWL^{+}23b, LCSA22, LCCZ22, MP21, MWR^{+}22, MBS^{+}21, MZS^{+}24, NSK^{+}21, PFC20, PAPB20, PMCM24, RGG20a, RG20, SCC^{+}24, SIA^{+}23, SAPW23, UVL^{+}23, WWSS20, WLF^{+}22, WM23, YBR^{+}23, YKJ^{+}23, ZWH^{+}22, ZWZ^{+}22, ZGX^{+}23, ZLX23, ZHL^{+}21b, LBW^{+}22, SLK^{+}20, YZJ^{+}20]. Time- [KKE21]. Time-Dependent [RG20, WM23]. Time-Series [BBSvL24, FSS^{+}21, HSYZ24, ZWZ^{+}22]. Time-Varying [HW20, HZX^{+}21, HW22, HW23, WWSS20, YBR^{+}23, ZGX^{+}23].

Timeline [CWW^{+}20]. Timelines [OBCT24]. Timely [CAGM22]. Times [AAA^{+}21]. TimeSplines [OBCT24]. TimeTubesX [SUB^{+}22]. TimeTuner [HSYZ24]. Tissue [HS23, WKN^{+}23]. Title [Ano21p]. TIVEE [CXY^{+}22]. TLX [CQHP22]. Together [WBF20]. Tolerance [GLGB24]. Tomography [NBE^{+}23]. Tones [PGE^{+}22]. Too [ZYM^{+}24]. Tool [ABE^{+}22, CPD^{+}24, CCM20, KRHH24, KCK^{+}24, LLP^{+}23, OCL^{+}21, SDMK22, SNH^{+}23, WPB^{+}20]. Toolkit [ENM24, FCH^{+}23, GNMQ22, KKZE20, LLT24, MHN^{+}24, NSS21]. Tools [OCW^{+}24, RCD^{+}23]. Tooth [LZY^{+}23, WWW^{+}24]. Top [DXX^{+}21]. Top-Down [DXX^{+}21]. Topic [ACT^{+}24, EAKC^{+}20, GL20, KDEP21]. Topic-Based [GL20]. Topics [BXF^{+}22]. TopoCluster [LIFD23]. Topological [ATK^{+}24, ECR22, JLK20, KRZ^{+}20, KGBP20, LB22, LWM^{+}20, LI24, LGMT21, MUM^{+}21, QFWS22, QRZZ21, RAC22, YHD21, ZRPW23]. Topologically [YGP^{+}24]. Topologies [CDBM22]. Topology [FLZ^{+}21, HZZ24, KW23, LWWF21, LIFD23, MDL^{+}23, MDS24, QHL^{+}20, QR21, RG20, TSH21, VGK^{+}22, VGT21, YLGW24]. Topology-Adaptive [MDL^{+}23]. Topology-Aware [QHL^{+}20]. Topology-Based [KW23, LIFD23, VGK^{+}22]. TopoMap [DTS^{+}21]. TopoSZ [YLGW24]. TopoTag [YHD21]. Toroidal [CDBM22]. Total [LLW^{+}22a]. Touch [DUWW22, KKW23, SSS20, SSS20]. Touchscreen [BSG^{+}20, CRJ^{+}24]. Tours [AeSL^{+}23, WF21]. TPMS [HWL^{+}22, YRL^{+}20]. TPMS-Based [HWL^{+}22]. Traced [CAR^{+}23]. Traces [HGB22, Off20, SBT^{+}23, WNC^{+}22]. Tracing [HPT^{+}23, IGMM22, MUM^{+}21, MSuG^{+}23, PFN22, VSBY22, WZU^{+}21, WMZ22, XGS^{+}23, XWPG^{+}24]. Track [MXT^{+}22, CLZ^{+}24]. Track-Based [MXT^{+}22]. Tracker [TNJ^{+}22, APSB23]. Trackers [TSS22]. Tracking [AMK^{+}21, BWZ^{+}20, CLZ^{+}24, FMP23, GWC^{+}23, GLX^{+}21, HZQ22, KCB^{+}21, LMD^{+}22, LGW^{+}20, MSWI22, PALW20, QBW^{+}20, SHD^{+}20, SFL^{+}22, WLF^{+}22, XDH^{+}22, YGP^{+}24, ZeB^{+}21, ZSZ^{+}24, DJHBJ21, DJBJ23]. Tracks [MXT^{+}22]. Tracts [XNF^{+}23]. Trade [CBB23, JJKJ20, PFCB23, XDBAR24, ZGL^{+}21]. Trade-Off [CBB23, JJKJ20, ZGL^{+}21]. Trade-offs [PFCB23]. Traditional [JSA^{+}20]. Traffic [BMWD20, CDX^{+}20, CLH^{+}23, GZL^{+}21, JLP^{+}23, LKJ^{+}20, PYSJ21, ZLL^{+}21]. Trafficking [VCO^{+}23]. TrafficVis [VCO^{+}23]. Trainer [WMH^{+}23]. Training [BFS^{+}23, CPD^{+}24, FYC^{+}23, GEU^{+}22, GF24, KMH^{+}23, LHS^{+}22, NMR^{+}23, OSC22, SML^{+}23, TNZ^{+}24, TPH22, WCX21a, WMH^{+}23, leáB^{+}21]. Traits [RBK^{+}21]. Trajectories [HZC^{+}20, RPD20, SHL^{+}21]. Trajectory [FSN20, MNZ^{+}20, RGDG23, SDK23a, TSH21, YCC^{+}21]. Transactions [Ano20a, Ano20b, Ano20c, Ano20d, Ano21b, Ano24h]. Transdisciplinary [HBH^{+}20]. Transfer [CKQ^{+}23, GMTD23, JYF^{+}20, MFH^{+}21, OSC22, PLL^{+}24, RPG23a, SLC^{+}23, SJL^{+}23, ZLY22]. Transferable [WCCS24]. TransforLearn [GSL^{+}24]. Transform [LLL^{+}22, PvSvdE^{+}24]. Transform-and-Perform [PvSvdE^{+}24]. Transformation [LDB^{+}21, LLW^{+}23]. Transformations [JLK20, ZFCG23]. Transformer [ACL^{+}24, GSL^{+}24, YCW^{+}24, ZZD^{+}23a]. Transformer-Based [ACL^{+}24]. Transformers [JKV^{+}22, LWD^{+}23, NKWW22]. Transforming [CWH^{+}23]. Transient [APP^{+}22, CZGF21]. Transit [MMN^{+}22]. Transition [TME^{+}22]. Transitioning [WDSM24]. Transitions [CRP^{+}20, GLHQ21, KH21, RBR20, SKW^{+}23, YDM^{+}21]. Translation [HZX^{+}21, LTL^{+}22, NVRS^{+}21, SSB^{+}22]. Translations [ALC22]. Transmission [BPW^{+}21]. Transmittance [JKU^{+}22]. Transparency [GG21, KNM^{+}21, PGE^{+}22, SRKK21, YeSiK^{+}23]. Transparent [DWOB20]. Transport [KJI^{+}21, ŠK20, WMMB23]. TransVis [ALC22]. Travel [LHA^{+}21]. Traveler [SBT^{+}23]. Travels [HMK^{+}20]. Treadmill [WLC^{+}23]. Tree [COFJ23, CWS^{+}21, GLA^{+}24b, JdJTC24, KGBP20, KKV22, KW23, LY23, LWW^{+}24, PSS^{+}22, SN23, SKNŽ20, WG21, WPTG24, YBR^{+}23, ZSL21, ZWZ^{+}22]. Treelike [HBMK23]. Treemaps [BHA^{+}23, HJL^{+}23]. Trees [AA21, BTL23, CRWA22, LZM20, PVDT22, PVT23, SMKN20, SN23, SMS^{+}22, YWM^{+}20]. Trenches [ASSB^{+}23]. Trend [BLIC20]. Trends [PCL24, SFNRZ^{+}23, TBW^{+}23]. Triage [LAS^{+}20]. Triangulation [LCSA22]. triggers [TB24]. TROPHY [YGP^{+}24]. Tropical [YGP^{+}24]. Trotter [HMK^{+}20]. Trouble [WW22a]. Truncation [MBB20]. Truncation-Based [MBB20]. Trust [ESB^{+}24, GKC^{+}24, LCK^{+}23, PFCB23, YCM^{+}24]. Trustworthiness [LCW^{+}23a]. Truth [CMXF21]. Try [WFK24]. TSR [HW20]. TSR-TVD [HW20]. TTHRESH [BRLP20]. TU [Ano22b]. TU/e [Ano22b]. Tube [RGDG23]. Tubes [CZC^{+}20]. Tumor [JKW^{+}22]. Tunable [HHI24]. Tuning [WW22a, PCZ^{+}21]. Turbine [NSG^{+}20]. Turbulence [LXL21, NMC21]. Turbulent [NNF^{+}20, ZYC24a]. Turning [ZHKY23]. Turns [XCG^{+}24]. Tutorial [GSL^{+}24]. Tutorials [SKS^{+}23]. TVCG [IMKP21, IGMW22, FGL^{+}23, HKVZ20]. TVCG.2020.3021534 [FCFC22a]. TVD [HW20, HW22]. Tweets [SLK^{+}20]. Two [AKS22, FOH^{+}21, HCWK23, LGWL21, RBF^{+}23, SHD^{+}20, YBVI22, vOVR23].

Two-Dimensional [vOVR23]. Two-Step [SHD^{+}20]. Type [MML21, PBM^{+}24, SESH24]. Typed [CNY22]. Types [DV23, FKSP^{+}24, RCD^{+}23]. Typing [HDFK21, RZLX24, SDK23a].

U.S. [YCM^{+}24]. Ubiquitous [BBRE24]. Ultra [LCW^{+}23b]. Umbra [WBL^{+}22]. Unambiguous [ZPG21]. Uncalibrated [TGM21]. Uncertain [HZC^{+}20, SHT^{+}22]. Uncertainties [DWH^{+}23]. Uncertainty [AMS^{+}21, AMY^{+}22, AJSP23, CQHP22, GSS^{+}20, HKW23, HLM^{+}20, Hul20, Kay24, KPO^{+}23, LWL^{+}20, PVPM22, PLP^{+}23, PM23, SGH^{+}23, YWM^{+}20, ZHN24, ZCL^{+}21, ZAH22, ZTC^{+}23, ZS21, KMWD21, WKMD22]. Uncertainty-Aware [GSS^{+}20, HKW23]. Uncertainty-Oriented [ZCL^{+}21]. Uncover [ROM^{+}23]. Underestimation [XCLF20]. Underrepresentation [PSH20]. Undersampled [SWZ^{+}23]. Understand [GZL^{+}21, RBSN22]. Understandable [SWT^{+}21]. Understanding [ASSB^{+}23, BLBL23, BBM^{+}21, BXF^{+}22, CBB23, CNC^{+}20, CIA24, CSC^{+}21, CPR^{+}22, DAK^{+}21, DWH^{+}23, ENvBC23, FWM^{+}24, GSP^{+}21, HWT^{+}24, HCM^{+}22, KZD^{+}23, KKJ^{+}21, LLT24, LLS^{+}20, LBL^{+}21, LWY^{+}22, LBE20, PTX^{+}22, Par22, PCJ23, SPM24, WPL^{+}22, WHJ^{+}24, WQQ^{+}24, WGO20, WGO22, XWL^{+}21, ZSCC22, ZLVV22]. Unfair [GKC^{+}24]. Unfamiliar [EHB^{+}23]. Unfolding [HiM^{+}22]. Unified [CGT^{+}24, CLL^{+}23, HMGO23, HSB^{+}21, LWY^{+}22, LGH^{+}24, XWZ^{+}23, ZLC^{+}23]. Uniform [MSWI22, SMCL24]. Uniform/Biased [MSWI22]. Unifying [SCL^{+}23a, ZPWS23]. Unintended [BBM^{+}21]. Union [XGS^{+}21]. Unit [SLS21, ZLL^{+}21]. Universal [XWZ^{+}23]. University [Ano22b, Ano22c, Ano22d, Ano22e].

Unknown [MA20]. Unmet [DZTF22]. Unmodified [FMP23]. UNOC [PTX^{+}22]. Unordered [WG21]. UnProjection [EAS^{+}23]. Unraveling [SDC^{+}24]. Unreasonable [BCT22]. Unsteady [RMB^{+}21, ZHTR22]. Unstructured [KKV22, MWUP22, MSuG^{+}23, SXW^{+}22, WMZ22]. Unstructured-Mesh [SXW^{+}22]. Unsupervised [KBJ^{+}20, LWF23]. Untangling [LZ21, SEAKC21]. Untidy [BCT22]. Update [WLF^{+}22]. updating [KMWD21, KOL^{+}20]. Uplift [EGP^{+}21]. Upon [CSC^{+}21]. Upper [CNC^{+}20, CNK^{+}24, SWSK23]. Upsampling [MDH^{+}23, YCH^{+}22]. Urban [DWX^{+}22, FLZ^{+}21, MHN^{+}24, MFS^{+}23, PJYW21]. Urbana [Ano22c]. UrbanMotion [SHL^{+}21]. UrbanRama [CMF^{+}22]. Usability [LHA^{+}21, MUM^{+}21, SKFZ22, ZLVV22]. Usage [JSS^{+}20]. Use [BLBL23, BCC^{+}20, CSC^{+}21, LKS22, RR23, RLG^{+}23, SGH^{+}23, SDK23b]. Used [GEU^{+}22]. User [Ano23f, BSP20, CFDN24, DPM24, EJS^{+}23, FGF^{+}21, GNMQ22, HMGO23, HBLW23, JVRL24, JWE^{+}22, KJL24, LZP^{+}20, MHIS23, MBC^{+}23, NHC^{+}20, NDLW20, PNKC21, SACB^{+}23, SHPE20, SFPM22, SPZS24, STA^{+}21, VVR^{+}23a, WRZ^{+}21, WSZ^{+}23, WHC^{+}23, WSL^{+}24, WGO20, WGO22, WIP^{+}24, WGH^{+}24, XLW^{+}24, YSM^{+}20, YHC^{+}22, YGE^{+}21, DGD^{+}23, ZHH22, CCEA^{+}24]. User-Centered [YSM^{+}20]. User-Centric [WHC^{+}23, WSL^{+}24]. User-Driven [GNMQ22, PNKC21]. User-Guided [BSP20]. User-Imperceptible [MHIS23]. User-Matched [DPM24]. Users [BCC^{+}20, CCS^{+}21, FTWP22, JLH24, LQS^{+}23, NGW^{+}24, UDH23, XCG^{+}24]. Using [AMY^{+}22, BBL^{+}22, BJCL21, BKPB21, COFJ23, CXD^{+}21, CQHP22, CZMR21, CMKK21, CMKK22, CZF^{+}23, CJY^{+}23, CYX^{+}23, CXW^{+}23, CLL^{+}23, CRJ^{+}24, CFGM22, CBHR^{+}23, DMJ^{+}22, DLH^{+}22, DHM^{+}22, DHF^{+}22, DKuH20, ECR22, EIB23, FiMH21, FHR^{+}21, GWW^{+}21, GSH21, GFSHO20, HLA^{+}22, HGO21, HCJ21, HCL20, HHB^{+}23, HZQ22, HS23, JTT^{+}23, JKU^{+}22, JLM^{+}21, KIS22, KMH^{+}23, LPJT^{+}22, LBL^{+}21, LWWH20, LLL^{+}22, LETF21, LCSA22, MML21, MBS^{+}21, MSWI22, MKK20, MCFKF24, MZS^{+}24, NMR^{+}23, NCE23, NBJ^{+}21, PCQ^{+}20, PL21, PPE23, PGS21, PGE^{+}22, PTS^{+}20, PJHY20, QBW^{+}20, QR21, RBK^{+}21, RPD22, RHLC22, RAFSA23b, SWF^{+}24, SCHE23, SN23, SRBP20, SKNŽ20, SWY^{+}22, TGM21, TLBB23, WZU^{+}21, WWSS20, WWZ23b, WGO20, WBM21b, WGO22, WR23, WHBC23, XCG^{+}21, XCG^{+}24, YITS23, YST^{+}20, YIIW24, YHC^{+}22, ZSLL20, ZX22, ieSM^{+}23, AV21, BWM23, BNWvW21, EAKC^{+}20, GLK^{+}23, HJL^{+}23]. using [HDFK21, HIK^{+}23, HE24, IRR^{+}22, JDZK22, LQS^{+}23, RPD21, RZW^{+}24, SMNK21, SDK23a, SKR^{+}24, TLW^{+}21, TB24, UIHS20, WSL^{+}20, WCZ^{+}20, WBA^{+}23, WCC^{+}21, WPNK21, ZCL^{+}21, ZLW21a, ZCR21]. Utilitarian [SWF^{+}24]. Utility [HFS^{+}21, JJKJ20, SFSA24]. UV [MBHE24].

V [NIK24]. V-Mail [NIK24]. V2V [HZX^{+}21]. V4D [GXH^{+}24]. VAC [XZKM22]. VAC-CNN [XZKM22]. VACSEN [RWJ^{+}23]. Validated [GGT^{+}20, HIDI23]. Validating [AYG^{+}22]. Validation [MWR^{+}22, MSBV^{+}23, XDW21, ZOS^{+}23]. Validity [KKEG20]. Value [BBSvL24]. Values [FW22]. Valve [EEL^{+}20]. Variability [HX23, JQL^{+}24, MPNF24, NMR^{+}23]. Variable [HZX^{+}21, HHI24, KHI20, LBP24, ZCL^{+}21]. Variable-Distance [LBP24]. Variable-Intensity [HHI24, KHI20].

Variable-to-Variable [HZX^{+}21]. Variant [YQN^{+}21]. Variate [CPCS20, JH20, SHT^{+}22]. Variation [LLW^{+}22a, XHFZ24]. Variational [DMJ^{+}22, FWZZ20, LWL^{+}23a, YWM^{+}23]. Variations [ATK^{+}24, ZZW^{+}22a]. Various [YWM^{+}23]. Varying [CSM^{+}23, HW20, HZX^{+}21, HW22, HW23, RRK^{+}22, WWSS20, XLL^{+}20, YBR^{+}23, ZGX^{+}23]. VASABI [NHC^{+}20]. VASSL [KKZE20]. VAST [Ano20p, Ano21q, Ano21r]. VATLD [GZL^{+}21]. VBridge [CLD^{+}22]. VC [WYZ^{+}21]. VC-Net [WYZ^{+}21]. VDL [SXL^{+}23]. VDL-Surrogate [SXL^{+}23]. VDOM [SSB^{+}22]. VeCHArt [PTS^{+}20]. Vector [BKPB21, HKMG22, HSF^{+}20, LZZ^{+}21, LDC^{+}23, MDS24, Pat22, RG20, SSB^{+}22, TG24]. Vectorization [LWH^{+}22, SC22, TG24]. Vectorized [LZH^{+}21]. Vectorizing [LXL21]. Vega [ZPWS23]. Vega-Lite [ZPWS23]. Vehicle [FTWP22, JLM^{+}21, SWF^{+}24]. Vehicles [NQE21]. Velocity [XHFZ24]. ventricle [BTv^{+}23]. Venues [TBW^{+}23]. VEQ [RL20]. Verbal [GTL^{+}23, WBPC23]. Verbalizations [HT22]. Verbalize [RDH23]. Vergence [WZL22]. versus [KLCK20, WBPC23, vOVR23]. Vertex [KKV22, PSL23]. Vertical [GPR^{+}24, LIB24]. VERTIGo [CSIP22]. Very [CB22]. Vessel [KHR21]. Vestibular [GTH^{+}22]. Vestibulo [CWSJ23]. Vestibulo-Ocular [CWSJ23]. VG [HPM^{+}24]. VGG [XXM^{+}21]. VGTC [Ano20j, Ano20k, Ano21g, Ano21f, Ano22o, Ano22s, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano23m, Ano23a, Ano23q, Ano23b, Ano23c, Ano23d, Ano23e, Ano24j, Ano24o, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, SW21, SAB20, Ano22x, FG22, GF22, Kiy22b, Kiy22a, Kli22]. vgtc.org [Ano20j, Ano20k, Ano21g, Ano21f, Ano22o]. VI [ZYL^{+}24]. VI-SLAM [ZYL^{+}24]. Via [ZZD^{+}23a, ADD^{+}22, ACL^{+}24, BDL^{+}21, BZKG21, CSWZ24, CWB^{+}20, CWS^{+}20, CYD^{+}23, CFZZ23, COZ^{+}23, CCS^{+}24, DLP^{+}23, EBPF21, GZW^{+}20, GLL^{+}22, GPR^{+}24, GJC^{+}22, HW23, HSV^{+}20, HCX^{+}21, HYSL23, HYC^{+}23, HLH^{+}23, LBW^{+}22, LWL^{+}22b, LWH^{+}22, LWF23, LWL^{+}23b, LZL^{+}23, LWW^{+}24, LS22, MNZ^{+}20, MLC^{+}20, MGO21, OWW^{+}24, RZLX24, SSX^{+}20, SLR20b, SLR21, SDXR22, SOL^{+}22, SCC^{+}23, TBL^{+}20, TYPC20, UIHS20, WHSZ21, WJBB22, WBS21, WGH^{+}24, XTYL20, ZLQH21, ZWZ^{+}23, ZZD^{+}23b, ZWW22, ZSS^{+}21, ZWW^{+}23d]. Vibration [CZGF21, JLM^{+}21]. Vibrations [SAMB^{+}23]. ViComp [WLH24]. Video [BFS^{+}24, EMM^{+}22, FE21, FRiM^{+}23, GTHC20, HWW^{+}24, HRX23, IAI^{+}23, JZHA22, LDB^{+}21, LZLS22, LWH^{+}22, LHZ^{+}23, LLDW24, LAC^{+}24, LCW^{+}23b, MNZ^{+}20, MOA21, SZC^{+}23, SWY^{+}22, SAPW23, TWW^{+}22, WFY^{+}21, WLH24, WWR^{+}20, WBWL24, WQ20, ZYL^{+}22, ZLZ^{+}23a, ZCM^{+}23b]. Video-Based [FE21, ZCM^{+}23b]. VideoModerator [TWW^{+}22]. VideoPro [HWW^{+}24]. Videos [BBMM^{+}23, CYC^{+}22, CYX^{+}23, CSM^{+}23, GFCM23, LZX^{+}22, SZZW24, SKFZ22, SMYF22, VH23, YVBI24, ZWW^{+}20, ZSW^{+}21, ZWW^{+}23a]. View [CZL^{+}21, DHF^{+}22, FALH20, GXH^{+}24, GXY^{+}20, KGX^{+}23, KGB22, LG23, LYBP23, MDL^{+}23, MTE^{+}20, MP21, NIM^{+}21, QLC^{+}24, RLG^{+}23, SA22, SHOP23, SHT^{+}22, SXL^{+}23, SSAZ22, SNK^{+}22, TME^{+}22, WZH20, WZW^{+}23, Wu22, XEXW24, YCW^{+}24, HSV^{+}20, LQS^{+}23, JDZK22]. View-Dependent [SXL^{+}23]. View-Illumination [KGX^{+}23]. View-Surface [DHF^{+}22]. View/Near [EML^{+}23]. Viewer [ACPB24, XSB^{+}22]. Viewers [MSMX23]. Viewing [APP^{+}22, HB20, LCY^{+}23, MP21, WBPC23]. Viewpoint [CZF^{+}23, GTHC20, MMK^{+}23, ZLZ^{+}23a].

Viewport [FBW21]. Views

[EEL^{+}20, HM24]. VIMO [TWC^{+}24]. Vinci [GWF^{+}23]. VIPurPCA [ZHN24]. VIRD [LAC^{+}24]. Virtual [AV21, ACL^{+}24, AMF20, Ano22x, Ano23f, ASP^{+}22, AI21, BTHL23, BBL^{+}22, BJCL21, BS22, BRLR24, BNWvW21, BBVS^{+}24, BWCT23, BVV^{+}22, BVV^{+}23, BSG^{+}20, BGS^{+}22, BBSC23, BBM^{+}21, BCB22, BC21, CCP^{+}21, CGAG20, CPD^{+}24, CDX^{+}20, CDS^{+}22, CMF^{+}22, CFL^{+}23, CZGF21, CFDN24, CNK^{+}24, DUWW22, DCC22, DHY^{+}22, DGB^{+}22, DPM24, DZG^{+}23, EHB^{+}23, ENvBC23, FLS23, FE21, FBW21, FHR^{+}21, FTWP22, FPK^{+}24, FOH^{+}21, FGS^{+}22, FG22, GMTD23, GST^{+}24, GLH22, GVN^{+}20, GSH21, GMVRB20, GBL^{+}22, GF22, GTH^{+}22, GLY^{+}23, GKN^{+}23, GZG^{+}24, HWC23, HLJ^{+}22, HDFK21, HGB22, HB20, HBLW21, HBLW23, HRX23, JDM^{+}22, JWE^{+}22, JJKJ20, Kel23, KNL23, KVB20, KKS^{+}22, KL22, KPL23, Kiy22b, Kiy22a, KSK^{+}23, KMH^{+}23, KBPR22, KVM^{+}22, KCWK20, KGR^{+}24, KKF20, KWFK20, KSB^{+}22, LBL^{+}21, LNB^{+}21, LSG24, LWL^{+}22a, LGY^{+}22, LKS22, LF23, LCK^{+}23, LCW^{+}23a, LETF21, LWK20, LZP^{+}20]. Virtual [LH22, LCM^{+}23, MCQ23, MWD^{+}23, MMK^{+}23, MUM^{+}21, MMN^{+}22, MWR^{+}22, MGM^{+}22, MML21, MF24, MCS^{+}23, MKK20, MCFKF24, MNK23, MBHE24, MSK23b, MB20a, MB20b, MB21a, MM23, NBNC20, NMR^{+}23, NDF^{+}21, NVRS^{+}21, NP24b, NBE^{+}21, OSC22, OF22, PTX^{+}22, PCL24, PSH20, PGS21, PRKM20, PHB^{+}22, PMW23, RBR20, RBK^{+}21, RTM^{+}20, RR23, RNO^{+}22, RRK^{+}22, SUS^{+}21, SACB^{+}23, SWF^{+}24, SKW^{+}23, SDMK22, SGB^{+}22, SHS^{+}22, SDK23b, SW21, SAB20, TIDQ24, TLBB23, TWA22, TPH22, UDH23, UVL^{+}23, VVR^{+}23b, VVR^{+}24, VH23, WSN21, WSZ^{+}23, WHY^{+}23, WSL^{+}24, WCCS24, WBF20, WF21, WBG^{+}23, WFK24, WGS^{+}24, WMMB23, WIP^{+}24, WH22b, WCC^{+}21, WPNK21, XLZ24, XHFZ24, YLS^{+}23, YQN^{+}21, YHC^{+}22, YYK^{+}22, YPW23, YLF^{+}20, YZN^{+}20, YB20, YGH^{+}23, ZZ23, ZMK^{+}20, ZUK21, ZTL^{+}23, ZA21, ZSCRB23, ZCL23, KOL^{+}20, KHD^{+}22, MGC^{+}21, MSAM^{+}22, MCSAL23, RL20]. Virtual-Physical [WCCS24]. VirtualCube [ZYL^{+}22]. Virtuality [FPG^{+}23, SPZS24]. Virtually [GWF^{+}23]. VIS [Ano20q, Ano21t, Ano21s, Ano22y, Ano22z, Ano22-27, Ano22-28, Ano22-29, Ano22-30, Ano22-31, Ano23v, Ano23x, Ano23w, Ano23y, Ano23z, Ano23-27, Ano23-28, Ano24s, Ano24t, Ano24u, Ano24v, Ano24w, Ano24x, HCWK23, PK21, SFNRZ^{+}23, SES23, Ano21h, BZP^{+}20, BB21]. Vis-a-Vis [BB21]. VIS30K [CLL^{+}21]. VIS4ML [SH24b]. Visceralization [LBL^{+}21]. VisCode [ZLW21a]. VisCommentator [CYC^{+}22]. VisConnect [SSZ^{+}21]. Viscosity [LHWW22]. Viscous [XDH^{+}22]. ViSeq [CYP^{+}20]. VisExPreS [GNMQ22]. Visibility [GLGB24, LKL23, WBM21b]. Visible [APHD24]. Visilant [ZeB^{+}21]. VisImages [DWS^{+}23]. Visinity [WKN^{+}23]. VisInReport [SEAB^{+}22]. Vision [CGT^{+}24, ENXS21, JZHA22, JKV^{+}22, KKW23, LWD^{+}23, UIHS20, WZL22, ZA21, ZOF^{+}23]. Vision-Based [JZHA22]. viSNE [CMK20]. Vispur [TAL24]. VisQA [JKV^{+}22]. VisRecall [WJBB22]. VisTA [FWZ^{+}20]. Vistrust [ESB^{+}24]. Visual [ABE^{+}22, AL20, ABL^{+}22, ADL^{+}22, BSB^{+}20, BRLP20, BXQ^{+}22, BZSD21, BCN^{+}20, BvOR21, BPW^{+}21, BBVS^{+}24, BHA^{+}23, BER^{+}23, BNRB21, BB21, CSJ^{+}21, CPCS20, CXD^{+}21, CAGM22, CCEA^{+}24, CSL^{+}23, CMKK22, CLA^{+}20, CYP^{+}20, CAA^{+}21, CWW^{+}23, COZ^{+}23, CWO^{+}24, CBW23, CXY^{+}22, CCM20, CBB^{+}21, CCW^{+}21, CSIP22, DSKE21, DWOB20, DKM24, DWC^{+}20, DWL^{+}22, DMBK21, DV23, DKuH20, EHA^{+}23, EGP^{+}21, ENK^{+}20, EKC^{+}23, EAS^{+}23, FAS^{+}21, FWM^{+}24, FKSP^{+}24, FSS^{+}21, FJK^{+}20, GSS^{+}23, GSL^{+}24, GEU^{+}22, GSK^{+}20, GNMQ22, GZW^{+}20, GZL^{+}21, GJC^{+}22, GGJ^{+}22, GLL^{+}24, HMK^{+}20, HTJ^{+}20, HPK^{+}22, HLW^{+}20, HZS^{+}22, HIDI23, HWW^{+}24, HKMG22, HHC^{+}21, HWN^{+}23, HZP^{+}24, HA20, HM22, HHS^{+}23, HE24, HCM^{+}22, HWW^{+}22, HZC^{+}20, HJZ^{+}21, HLCY22, HZJ^{+}24, HHM^{+}24, Iur22, JOEF20, JQL^{+}24, JLCZ22, JCZ^{+}24, JGC^{+}21, JWW^{+}23, JLP^{+}23, JYLS21, JSA^{+}20, JJHS^{+}22, JMK^{+}22, KKH21, KGQ^{+}24, KKG^{+}20, KHL21, KNAR^{+}22]. Visual [KZD^{+}23, KKEG20, KKZE20, KRW^{+}21, KKJ^{+}21, KL22, KLKE21, KKT^{+}22, KSHW22, KAS^{+}21, LLT24, LZW^{+}23, LKJ^{+}20, LLS^{+}20, LMF^{+}24, LWY^{+}20, LWS^{+}21, LM21, LBW^{+}22, LWL^{+}22b, LWD^{+}23, LCS^{+}24, LYBP23, LCW^{+}23a, LPP^{+}23, LZM20, LWT^{+}23, LXZ^{+}23, MXLM20, MTW^{+}20, MM21, MFH^{+}21, MHS^{+}22, MCQ23, MMS^{+}23, MTE^{+}20, MYBF22, MCW^{+}20, MBS^{+}21, MNK23, MHN^{+}24, MBS23, MSBV^{+}23, NQE21, NCWE22, NKWW22, NDP^{+}21, NDF^{+}21, NHC^{+}20, NFN^{+}23, NB24, OCL^{+}21, PTD^{+}21, PNKC21, PDD^{+}22, PMCM24, PRKM20, PYSJ21, PBM^{+}24, PJHY20, PSG^{+}22, QBW^{+}20, QCL23, QTW^{+}24, QR21, RMW^{+}24, RPD21, RBKM24, RLB^{+}21, RBRG21, RRG23, RPG23a, RBF^{+}23, SUS^{+}21, SSS20, SHOP23, SUB^{+}22, SBk^{+}24, SLL21, SPM24, SCL^{+}23a, SEAB^{+}22, SCR^{+}23, SXS^{+}21, SHL^{+}21, SNH^{+}23, SLC^{+}23, SCC^{+}24, SFS^{+}22, SJK^{+}23, SBe^{+}21, SVK^{+}21, SIL^{+}21, SDXR22, SKFZ22, SCEA23, SSSEA20, SaT^{+}23, SKR^{+}24, SMS^{+}22, SWS^{+}23, SHC^{+}20, TAL24]. Visual [TSHI22, TWC^{+}24, TLW^{+}23, TXM23, VFP24, VKT^{+}24, WZD^{+}20, WYC^{+}20, WAP^{+}21, WWC^{+}21, WXC^{+}21, WLS^{+}22, WMW^{+}22, WMH^{+}22, WBL^{+}22, WZY^{+}22, WM23, WCX^{+}23, WWZ^{+}23a, WMH^{+}23, WLZ^{+}23, WQQ^{+}24, WSL^{+}24, WYIS24, WKN^{+}23, WZD^{+}21, WHL^{+}20, WWS20, WBPC23, WQ20, WPNK21, WLG^{+}22, WDC^{+}23, WDX^{+}23, WXG^{+}24, XVW^{+}21, XZS^{+}22, XHL^{+}23, XMT^{+}21, XDW21, XMK^{+}22, XVF20, XSB^{+}22, XWY^{+}20, XDH^{+}22, XNF^{+}23, XZKM22, YST^{+}20, YZF^{+}23, YTHL23, YCC^{+}23, YDMP22, YS20, YLTL23, YBOB24, YBL^{+}20, ZeB^{+}21, ZWW^{+}20, ZSW^{+}21, ZLL^{+}21, ZWW^{+}23a, ZSM21, ZWW^{+}23b, ZWZ^{+}23, ZTC^{+}23, ZKS^{+}20, ZLL^{+}20, ZFF22, ZGX^{+}23, ZG20, ZWG^{+}23, ZWW^{+}23c, dSBdO^{+}24, leáB^{+}21, GKC^{+}24]. Visual-Free [LXZ^{+}23]. Visual-Inertial [BXQ^{+}22]. Visualisation [ERB^{+}21, LHC^{+}21, RVB^{+}22, SAK20, WJBB22, YWL^{+}20, YYD^{+}21]. Visualisations [ACPB24, CDBM22, FPK^{+}24, SGH^{+}23, WBG^{+}22]. Visualising [DMMF21, RBSN22]. Visualization [ALR23, AJVS^{+}23, ASSB^{+}23, AR22, Ano20a, Ano20b, Ano20c, Ano20d, Ano20j, Ano20k, Ano21b, Ano21g, Ano21f, Ano22o, Ano22s, Ano22a, Ano22b, Ano22c, Ano22d, Ano22e, Ano23m, Ano23a, Ano23q, Ano23b, Ano23c, Ano23d, Ano23e, Ano24h, Ano24j, Ano24o, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, APSB23, APBB24, AMY^{+}22, AJSP23, BKR^{+}24, BVY^{+}23, BLBL23, BBRE24, BS21, BDRW21, BZP^{+}20, BB21, BZKG21, BBSvL24, BLIC20, BKT^{+}22, CBL^{+}24, CJS^{+}22, CXZ^{+}24, CRM^{+}24, CBB23, CRP^{+}20, CWB^{+}20, CSW^{+}20, CLL^{+}21, CZF^{+}22, CSX^{+}22, CWW^{+}23, CLL^{+}23, CSC^{+}21, CPR^{+}22, CBE^{+}21, CFGM22, CGD^{+}24, DPD^{+}24, DWS^{+}23, DFP^{+}20, DP20, DS22, DZTF22, DIPJ22, EBPF21, ENXS21, EBKL21, FWZ^{+}20, Fek21, FLZ^{+}21, FW22, FS24, GSL21, GZRP^{+}22, GBM^{+}22, HKB^{+}22, HW20, HW22, HW23, HZII24, Hei21, HHC^{+}21, HRS^{+}22, HB24].

Visualization

[HCJ21, HSS^{+}20, HNGC21, ISBP22,

IMQ^{+}20, JK23, JLX^{+}23, JSS^{+}20, JMK^{+}22, KSHH22, KBB^{+}23, KRS^{+}22, KRHH24, KW23, KCWK20, KGB22, KMM24, KSB^{+}22, LG23, LVV^{+}21, LT20, LWC24, LME^{+}23, LAN21, LST^{+}22, LSS24, LM20, LWZ^{+}22, LQS^{+}23, LYL^{+}23, LY23, LLW^{+}23, LCY^{+}23, LLP^{+}23, LWSY20, LWM^{+}20, LXL21, LIFD23, LCYQ24, LALG22, LS22, LTL^{+}22, MMF20, MGWK^{+}22, McN23, MWJ22, MD20, MDS24, MEHD24, MZS^{+}24, MRS22, NSS21, NP21, NP24a, NSW24, NCE23, NE24, NMC21, NBE^{+}23, NSG^{+}20, NNF^{+}20, OZZ24, OKM21, OM22, OCW^{+}24, OWW^{+}24, PVPM22, PLP^{+}23, PSY^{+}20, PLW^{+}23, Par22, PL21, PPE23, PD24, Pla21, PWK21, PMS^{+}22, QZZ22, QR22, RHHH20, RDHH21, RPD22, RMB^{+}21, RCD^{+}23, RFD21, RSD^{+}23, RWJ^{+}23, RGDG23, SCRL20, SLR^{+}20a, SRKK21, SSZ^{+}21, SSL^{+}23, SYY^{+}24, SLC21, SLC^{+}23, SIA^{+}23, SH24a, SLSW23, SLW^{+}24, TARB23, TZT^{+}22]. Visualization [TFE21, THFI21, TSH21, Tsa21, VBP^{+}21, VC20, VE21, WSL^{+}20, WFW^{+}20, WZH20, WTS^{+}21, WYZ^{+}21, WCWQ22, WPL^{+}22, WH23, WHS^{+}23, WTL24, WBA^{+}23, WITW22, WZW^{+}23, WLT^{+}24, WAC^{+}22, WCH^{+}22, WWS^{+}22, WGM^{+}24, WXS^{+}24, XCK20, XGS^{+}21, XWH^{+}23, YWM^{+}20, YCB^{+}21, YBVI22, YVBI24, YSM^{+}20, YCC^{+}21, YLLW24, YSD^{+}23, YJ21, YLF^{+}20, ZMD^{+}22, ZYM^{+}24, ZTWW21, ZCL^{+}21, ZLW21a, ZLC^{+}23, ZSG^{+}23, ZPG21, ZZZ24, ZRJW20, ZLG^{+}21b, bÇ22, ieSM^{+}23, BTv^{+}23, TDI22]. Visualizations [AL21, ADM^{+}22, APHD24, ATK^{+}24, BvOR21, BMFE20, CZL^{+}21, CXY^{+}22, DUWW22, DCM^{+}23, ESB^{+}24, GSS^{+}23, GLK^{+}23, GTL^{+}23, GL20, HBH^{+}20, HWT^{+}24, HA20, HH20, HPM^{+}24, HPAB23, JSA^{+}20, JMK^{+}22, KWH22, Kay24, KHR21, KPO^{+}23, LTC21, LRA23, LBB^{+}20, LAML23, MMG^{+}21, MSMX23, MFS^{+}23, NL24, Off20, PCQ^{+}20, PFCB23, PAPB20, RDH23, RPHJ20, SCHE23, SB23, SLS21, TLW^{+}21, THS^{+}21, TCX^{+}23, WJKN21, WXW^{+}20, WAV^{+}21, WTD^{+}21, YXW^{+}23, YCM^{+}24, YVBI24, YZF^{+}22, ZCM23a, ZJX^{+}24, ZWW^{+}23d, ZBNS21, KMWD21, WKMD22, dBRGD22]. Visualize [Hul20, LPJT^{+}22, SEAKC21, ZLG^{+}21a]. Visualized [KKGMH21, XSHF20]. Visualizing [ABGG21, BMA^{+}23, CNC^{+}20, CLCY20, CNY22, DCS^{+}24, FCS^{+}20, HiM^{+}22, HTBL22, HNS23, HPRC20, HX23, HMTI24, JWKN21, iKYOW23, LKAH21, LWBM22, LWL^{+}20, NBJ^{+}21, SESH24, SNK^{+}22, SAPW23, VCO^{+}23, WDG^{+}20, WHJ^{+}22, WHJ^{+}24, WBI20, XDBAR24, XFD^{+}23, YC23, ZHN24, ZAH22]. Visually [FYE^{+}22, LAS^{+}20, NDLW20, PSH21, PTS^{+}20, PIS20, ZZD^{+}23b, ZNAN20]. Visuals [JMK^{+}22]. Visuo [CZGF21, MHIS23, SEK^{+}24]. Visuo-Haptic [CZGF21, MHIS23, SEK^{+}24]. VITALITY [NKWW22]. Vivern [KSB^{+}22]. VividGraph [SLSW23]. Viz [CZW^{+}20]. VizCommender [OKM21]. VizLinter [CSX^{+}22]. VizSnippets [OM22]. Vocabulary [WBG^{+}22, WBG^{+}22]. Voice [PZC^{+}23]. Void [KHR21, RPD20]. Void-and-Cluster [RPD20]. Vol [Ano20a]. Volume [ATHI24, AMS^{+}21, BWM23, HHK^{+}24, IGMM22, JTT^{+}23, KKT^{+}22, KLTB21, MPV21, SCRL20, TFE21, VM23, WWSS20, WYZ^{+}21, WW22b, WITW22, ZSL21]. Volume-Composition [WYZ^{+}21]. Volumes [DGKOC20, HZCW22, HW23, MZS^{+}24, PD24]. Volumetric [BWM24, ER21, HLL^{+}24, IGMM22, MSuG^{+}23, MRS22, RPD22, WL20, WCTW21, YLJ^{+}22, ZYP^{+}24]. Voronoi [LMGY22, NRA^{+}23, SCLK21]. Vortex [LXL21, RZW^{+}24]. Vortex-Core [LXL21]. Vortices [GT20, ZYC24a, ZHTR22].

Vorticity [BSB^{+}20]. Voting

[CFL21, ZCZ^{+}21]. Vox [LYZ^{+}24]. Vox-Surf [LYZ^{+}24]. Voxel [GXH^{+}24, LYZ^{+}24]. Voxel-Based [LYZ^{+}24]. voxel2vec [HTY^{+}23]. VoxSegNet [WL20]. VR [Ano22p, Ano22q, Ano22r, Ano22s, Ano23n, Ano23o, Ano23p, Ano23q, Ano24k, Ano24l, Ano24m, Ano24n, Ano24o, BPQW23, IPPZ24, MF22, MB22a, NAW^{+}22, OK22, SK23a, ASCR^{+}22, AZA^{+}23, BJCL21, BGB^{+}22, BVV^{+}22, BVV^{+}23, BKN^{+}22, BFS^{+}24, BPL23, CWSJ23, CXXZ21, CFQ21, CQ22, CCS^{+}21, FHR^{+}21, FMP23, FMBN23, FGS^{+}21, GLK^{+}23, GTH20, GF24, GWD^{+}24, HTP^{+}23, HLA^{+}22, HAK^{+}23, HLL^{+}24, HFT^{+}24, JKL24, JJHS^{+}22, JLM^{+}21, KCGZ23, LKL23, LVR^{+}24, LLSM24, LPG^{+}22, LHS^{+}22, LPL^{+}24, LBP24, LLMA22, LDZ^{+}21, MSRJ20, MSAM^{+}22, NL24, PLD^{+}23, RSAA20, SACB^{+}23, SEK^{+}24, SFPM22, STA^{+}21, TWT^{+}22, WZZ^{+}23, WLC^{+}23, WMB23, WBA^{+}23, WBWL24, WBPC23, WWZP22, XZS^{+}23, XLW^{+}24, YHC^{+}24, ZCZ23b, ZZS^{+}23, ZLS21, ZOF^{+}23, ZIX^{+}24, ZP23, ZP24, ZC23, dPLM21]. VR-based [YHC^{+}24]. VR.net [WGS^{+}24]. VRIA [BJR21]. vs [GTL^{+}23, HX23, NMR^{+}23]. Vulnerabilities [MXLM20].

Wahrscheinlich [RDH23]. Walking [AYGR22, AYG^{+}22, CHSC24, CS23, DGD^{+}23, FLS23, KCGZ23, OF22, SAMB^{+}23, SMNK21, SRBP20, WCCS24, WBM21a, WBM21b, XLL^{+}22, XCG^{+}24, XLW^{+}24, XHFZ24, ZCZ23b, ZA21, ZYR^{+}20, MCSAL23]. walking-by-cycling [MCSAL23]. walking-in-place [MCSAL23]. Walks [CS23]. Wall [NSG^{+}20, SGJC23]. want [LLS^{+}20]. Warehouse [TZT^{+}22]. Warping [IWT^{+}20, MNZ^{+}20, dAWi^{+}23]. Wasserstein [PVDT22, SDT24, VBT20]. Watching [LVR^{+}24]. Watertight [CZGW23]. Wavelet [GFCM23].

Wavelet-Based [GFCM23]. Way [SPJ^{+}23]. Weakly [LFO23, SSX^{+}20, SLR21, ZHDX20]. Weakly-Supervised [SSX^{+}20]. Wearable [GWC^{+}23, HWMI23, KCGZ23, iKYOW23, TBW^{+}23, WH22a, WTY^{+}22, ZZS^{+}23]. Wearing [CCS^{+}21]. Weather [dSBdO^{+}24]. Weaves [KAS^{+}22]. Web [WTD^{+}21, BJR21, HHK^{+}24]. Web-Based [BJR21, HHK^{+}24]. WeChat [LWY^{+}20]. Week [BKN^{+}22]. Weighted [HWS22, HZQ22, SPW^{+}22]. Weiskopf [Ano22e]. Welcome [DGW24, SN22, SES23]. Well [ISBP22, KSK^{+}23]. Well-Being [KSK^{+}23]. WeSeer [LWY^{+}20]. What-If [WPB^{+}20]. Wheat [DWOB20]. Wheelchair [CQ22]. Where [DS22, HZS^{+}22, WLZ^{+}23]. Whether [CPD^{+}24]. While [FOH^{+}21, ZA21, WMB23]. White [HZII24]. Who [TB24]. Whole [HiM^{+}22, JLM^{+}21, VM23]. Whole-body [JLM^{+}21]. Wide [BJCL21, GTHC20]. Wide-Baseline [GTHC20]. Wide-FOV [BJCL21]. Wijk [Ano22b]. Wild [LELT23]. Window [KWFK20]. Windows [ZYP^{+}24]. Winglets [LWL^{+}20]. Wise [HZQ22, CAA^{+}21, MLC^{+}20]. Within [DZTF22, KPO^{+}23, SHOP23, VLM^{+}23, WZZ^{+}23, YS20]. Within-View [SHOP23]. Without [BWWL22, BXF^{+}22, WW22a, YHC^{+}24, JMK^{+}22]. Wizard [RMW^{+}24]. Wizualization [BBRE24]. Word [EAKC^{+}20, HPP^{+}20, KNL23, KKE21]. Word-Embedding [EAKC^{+}20]. Word-Gesture [KNL23]. Wordles [WCZ^{+}20]. Words [HT22]. Work [BFS^{+}24, CFGT21, LWD^{+}23, SWB^{+}22]. Workbench [JK23]. Workers [BSG^{+}20, CFGT21]. Working [BKN^{+}22, CQHP22, PCQ^{+}20]. Workspace [JLH24]. Workspaces [KNL23, VLM^{+}23]. World [COFJ23, GTH20, JGH^{+}24, PBAG23, RAFSA23a, RAFSA23b, WBPC23, WTY^{+}22, XLZ24, ZZW21, WGS^{+}24].

Woven [MDX^{+}23]. Wrangling [KBM21, XLF^{+}23, XFD^{+}23]. Wrap [CDBM22]. Wrinkle [DMJ^{+}22]. Wrinkles [ZH20]. WSDesc [LFO23]. WYSIWYG [YLL^{+}22b].

X [HJZ^{+}21, JKV^{+}22]. X-Ray [HJZ^{+}21]. X-raying [JKV^{+}22]. XAI [WHC^{+}23]. XR [CML24, DJBJ23, JVRL24, KNL23, KTL24, NGW^{+}24].

Yarn [MDX^{+}23]. Yarn-Level [MDX^{+}23]. Years [CBL^{+}24, HCWK23, SFNRZ^{+}23, TDI22]. Yeast [HLW^{+}20]. Young [HNGC21].

Zero [BPW^{+}21, JLCZ22]. Zero-Shot [JLCZ22]. Zooming [YCB^{+}21]. Zoomless [GBNH21].

References

Alcaide:2021:STA

[AA21] Daniel Alcaide and Jan Aerts. Spanning trees as approximation of data structures. IEEE Transactions on Visualization and Computer Graphics, 27 (10):3994–4008, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Andrienko:2021:CST

[AAA^{+}21] G. Andrienko, N. Andrienko, G. Anzer, P. Bauer, G. Budziak, G. Fuchs, D. Hecker, H. Weber, and S. Wrobel. Constructing spaces and times for tactical analysis in football. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2280–2297, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Abramov:2022:CIV

[ABE^{+}22] David Abramov, Joseph N.

Burchett, Oskar Elek, Cameron Hummels, J. Xavier Prochaska, and Angus G. Forbes. Cosmo-Vis: an interactive visual analysis tool for exploring hydrodynamic cosmological simulations. IEEE Transactions on Visualization and Computer Graphics, 28(8):2909–2925, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Ancona:2021:MIV

[ABGG21] M. Ancona, M. Beyeler, M. Gross, and T. Günther. MineTime insight: Visualizing meeting habits to promote informed scheduling decisions. IEEE Transactions on Visualization and Computer Graphics, 27(3):1986–1999, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Angelini:2022:EEM

[ABL^{+}22] Marco Angelini, Graziano Blasilli, Simone Lenti, Alessia Palleschi, and Giuseppe Santucci. Effectiveness error: Measuring and improving Rad-Viz visual effectiveness. IEEE Transactions on Visualization and Computer Graphics, 28(12):4770–4786, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ai:2024:DDI

[ACL^{+}24] Hao Ai, Zidong Cao, Haonan Lu, Chen Chen, Jian Ma, Pengyuan Zhou, Tae-Kyun Kim, Pan Hui, and Lin Wang. Dream360: Diverse and immersive outdoor virtual scene creation via transformer-based 360^{∘} image outpainting. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2734–2744, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Alebri:2024:ERP

[ACPB24] Muna Alebri, Enrico Costanza, Georgia Panagiotidou, and Duncan P. Brumby. Embellishments revisited: Perceptions of embellished visualisations through the viewer’s lens. IEEE Transactions on Visualization and Computer Graphics, 30(1):1424–1434, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Atzberger:2024:LSE

[ACT^{+}24] Daniel Atzberger, Tim Cech, Matthias Trapp, Rico Richter, Willy Scheibel, Jürgen Döllner, and Tobias Schreck. Largescale evaluation of topic models and dimensionality reduction methods for 2D text spatialization. IEEE Transactions on Visualization and Computer Graphics, 30(1):902–912, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Ahmed:2022:MSG

[ADD^{+}22] Reyan Ahmed, Felice De Luca, Sabin Devkota, Stephen Kobourov, and Mingwei Li. Multicriteria scalable graph drawing via stochastic gradient descent, (SGD)^{2}. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2388–2399, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Arleo:2022:IMV

[ADL^{+}22] Alessio Arleo, Walter Didimo, Giuseppe Liotta, Silvia Miksch, and Fabrizio Montecchiani. Influence maximization with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 28(10): 3428–3440, October 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Angori:2022:HGV

[ADM^{+}22] Lorenzo Angori, Walter Didimo, Fabrizio Montecchiani, Daniele Pagliuca, and Alessandra Tappini. Hybrid graph visualizations with ChordLink: Algorithms, experiments, and applications. IEEE Transactions on Visualization and Computer Graphics, 28(2):1288– 1300, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Alharbi:2023:NGI

[AeSL^{+}23] Ruwayda Alharbi, Ond ej Strnad, Laura R. Luidolt, Manuela Waldner, David Kou il, Ciril Bohak, Tobias Klein, Eduard Gröller, and Ivan Viola. Nanotilus: Generator of immersive guided-tours in crowded 3D environments. IEEE Transactions on Visualization and Computer Graphics, 29(3):1860– 1875, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Armando:2022:MDF

[AFB22] Matthieu Armando, Jean-Sébastien Franco, and Edmond Boyer. Mesh denoising with facet graph convolutions. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2999–3012, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Aseeri:2021:IAR

[AI21] Sahar Aseeri and Victoria Interrante. The influence of avatar representation on interpersonal communication in virtual social environments. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2608–2617, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Athawale:2023:FUV

[AJSP23] Tushar M. Athawale, Chris R. Johnson, Sudhanshu Sane, and David Pugmire. Fiber uncertainty visualization for bivariate data with parametric and nonparametric noise models. IEEE Transactions on Visualization and Computer Graphics, 29(1): 613–623, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Ageeli:2023:MPR

[AJVS^{+}23] Amani Ageeli, Alberto Jaspe-Villanueva, Ronell Sicat, Florian Mannuss, Peter Rautek, and Markus Hadwiger. Multivariate probabilistic range queries for scalable interactive 3D visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):646– 656, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Abramson:2022:PDB

[AKS22] Mark A. Abramson, Griffin D.

Kent, and Gavin W. Smith. Penetration depth between two convex polyhedra: an efficient stochastic global optimization approach. IEEE Transactions on Visualization and Computer Graphics, 28(12):4267– 4273, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ahn:2020:FVA

[AL20] Y. Ahn and Y. Lin. Fair-Sight: Visual analytics for fairness in decision making. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1086–1095, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Adar:2021:CVL

[AL21] E. Adar and E. Lee. Communicative visualizations as a learning problem. IEEE Transactions on Visualization and Computer Graphics, 27(2):946– 956, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Alharbi:2022:TID

[ALC22] Mohammad Alharbi, Robert S.

Laramee, and Tom Cheesman. TransVis: Integrated distant and close reading of Othello translations. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1397–1414, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Adar:2023:RGC

[ALR23] Eytan Adar and Elsie Lee-Robbins. Roboviz: a gamecentered project for information visualization education. IEEE Transactions on Visualization and Computer Graphics, 29(1): 268–277, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Ajani:2022:DFE

[ALX^{+}22] Kiran Ajani, Elsie Lee, Cindy Xiong, Cole Nussbaumer Knaflic, William Kemper, and Steven Franconeri. Declutter and focus: Empirically evaluating design guidelines for effective data communication. IEEE Transactions on Visualization and Computer Graphics, 28(10): 3351–3364, October 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Alderson:2021:RRI

[AMAS21] Troy Alderson, Ali Mahdavi-Amiri, and Faramarz Samavati. RIAS: Repeated invertible averaging for surface multiresolution of arbitrary degree. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3546–3557, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

AlZayer:2020:VLS

[AMF20] M. Al Zayer, P. MacNeilage, and E. Folmer. Virtual locomotion: a survey. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2315–2334, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Angelopoulos:2021:EBN

[AMK^{+}21] Anastasios N. Angelopoulos, Julien N. P. Martel, Amit P. Kohli, Jörg Conradt, and Gordon Wetzstein. Event-based near-eye gaze tracking beyond 10,000 Hz. IEEE Transactions on Visualization and Computer Graphics, 27(5):2577– 2586, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Athawale:2021:DVR

[AMS^{+}21] T. M. Athawale, B. Ma, E. Sakhaee, C. R. Johnson, and A. Entezari. Direct volume rendering with nonparametric models of uncertainty. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1797–1807, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Athawale:2022:UVM

[AMY^{+}22] Tushar M. Athawale, Dan Maljovec, Lin Yan, Chris R. Johnson, Valerio Pascucci, and Bei Wang. Uncertainty visualization of 2D Morse complex ensembles using statistical summary maps. IEEE Transactions on Visualization and Computer Graphics, 28(4):

1955–1966, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:IIT

[Ano20a] Anonymous. 2019 index IEEE Transactions on Visualization and Computer Graphics vol. 25. IEEE Transactions on Visualization and Computer Graphics, 26(2):1–40, February 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:ITVa

[Ano20b]  Anonymous.  IEEE Trans-

actions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 26(1): i, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:ITVb

[Ano20c]  Anonymous.  IEEE Trans-

actions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 26(5): i, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:ITVc

[Ano20d]  Anonymous.  IEEE Trans-

actions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 26(5): i, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:Cb

[Ano20e]  Anonymous.  Committees.

IEEE Transactions on Visualization and Computer Graphics, 26(1):xxiii, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:CCa

[Ano20f] Anonymous. Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 26(5):viii, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:Ca

[Ano20g] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 26(1): ii–ix, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:Cc

[Ano20h] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 26(5): ii–iii, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:Cd

[Ano20i] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 26(5): ii–iii, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:IVGa

[Ano20j] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC): http://vgtc.org/. IEEE Transactions on Visualization and Computer Graphics, 26(1): xvii, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:IVGb

[Ano20k] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC):

http://vgtc.org/.  IEEE

Transactions on Visualization and Computer Graphics, 26 (5):vii, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:IIP

[Ano20l] Anonymous. InfoVis International Program Committee. IEEE Transactions on Visualization and Computer Graphics, 26(1):xxi, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:IPC

[Ano20m] Anonymous. International Program Committee for Journal Papers. IEEE Transactions on Visualization and Computer Graphics, 26(5):ix, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:PRJ

[Ano20n] Anonymous. Paper reviewers for journal papers. IEEE Transactions on Visualization and Computer Graphics, 26(5):x–xi, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2020:SIP

[Ano20o]  Anonymous.  SciVis Inter-

national Program Committee. IEEE Transactions on Visualization and Computer Graphics, 26(1):xxii, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:VIP

[Ano20p]  Anonymous.  VAST Inter-

national Program Committee. IEEE Transactions on Visualization and Computer Graphics, 26(1):xx, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2020:VCC

[Ano20q] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 26(1):xviii–xix, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Anonymous:2021:RL

[Ano21a]  Anonymous.  2021 review-

ers list*. IEEE Transactions on Visualization and Computer Graphics, 27(12):4533– 4539, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:ITV

[Ano21b]  Anonymous.  IEEE Trans-

actions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 27(5): i, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:Ca

[Ano21c] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 27(2): iii–xv, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:Cb

[Ano21d] Anonymous. Contents. IEEE Transactions on Visualization and Computer Graphics, 27(5): ii–iii, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:CN

[Ano21e] Anonymous. Copyright notice. IEEE Transactions on Visualization and Computer Graphics, 27(2):C2, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:IVGb

[Ano21f] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC): http://vgtc.org/. IEEE Transactions on Visualization and Computer Graphics, 27(5):vi, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:IVGa

[Ano21g] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC):

http://vgtc.org/.  IEEE

Transactions on Visualization and Computer Graphics, 27(2): xxvi, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:IVR

[Ano21h] Anonymous. Info Vis reviewers.

IEEE Transactions on Visualization and Computer Graphics, 27(2):xl–xli, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:IPCa

[Ano21i] Anonymous. INFOVIS 2020 Program Committee. IEEE Transactions on Visualization and Computer Graphics, 27 (2):xxxii–xxxiii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:IPCb

[Ano21j] Anonymous. International Program Committee for Journal Papers. IEEE Transactions on Visualization and Computer Graphics, 27(5):viii, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:MPC

[Ano21k] Anonymous. Message from the Program Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 27(5):v, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:OC

[Ano21l] Anonymous. Organizing Committee. IEEE Transactions on Visualization and Computer Graphics, 27(5):vii, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:PRJ

[Ano21m] Anonymous. Paper reviewers for journal papers. IEEE Transactions on Visualization and Computer Graphics, 27(5):ix, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:SPR

[Ano21n] Anonymous. SciVis paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 27(2):xlii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:SPC

[Ano21o] Anonymous. SciVis Program Committee. IEEE Transactions on Visualization and Computer Graphics, 27(2):xxxiv– xxxv, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:TP

[Ano21p] Anonymous. [Title page]. IEEE Transactions on Visualization and Computer Graphics, 27(2): C1, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:VPC

[Ano21q] Anonymous. VAST Program Committee. IEEE Transactions on Visualization and Computer Graphics, 27(2):xxxi, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:VR

[Ano21r] Anonymous. VAST reviewers. IEEE Transactions on Visualization and Computer Graphics, 27(2):xxxviii–xxxix, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:VSC

[Ano21s] Anonymous. VIS 2020 Steering Committees. IEEE Transactions on Visualization and Computer Graphics, 27(2): xxxvi–xxxvii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2021:VCC

[Ano21t] Anonymous. VIS Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 27(2):xxix– xxx, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVD

[Ano22a] Anonymous. 2021 VGTC Visualization Dissertation Award Arjun Srinivasan, Tableau Software. IEEE Transactions on Visualization and Computer Graphics, 28(1):xxx, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVL

[Ano22b] Anonymous. 2021 VGTC Visualization Lifetime Achievement Award Jarke van Wijk, Eindhoven University of Technology (TU/e). IEEE Transactions on Visualization and Computer Graphics, 28(1):xxv, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVSa

[Ano22c] Anonymous. 2021 VGTC Visualization Service Award Loretta Auvil, National Center for Supercomputing Applications at the University of Illinois at Urbana Champaign. IEEE Transactions on Visualization and Computer Graphics, 28(1): xxvi, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVSb

[Ano22d]  Anonymous.  2021 VGTC

Visualization Significant New Researcher Award Michelle Borkin, Northeastern University and Benjamin Bach, University of Edinburgh. IEEE Transactions on Visualization and Computer Graphics, 28(1): xxvii–xxviii, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVT

[Ano22e] Anonymous. 2021 VGTC Visualization Technical Achievement Award Daniel Weiskopf, University of Stuttgart, Germany. IEEE Transactions on Visualization and Computer Graphics, 28(1):xxix, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:AI

[Ano22f] Anonymous. Author index. IEEE Transactions on Visualization and Computer Graphics, 28(1):1191–1197, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:BP

[Ano22g] Anonymous. Blank page. IEEE Transactions on Visualization and Computer Graphics, 28 (5):Bii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:FCa

[Ano22h]  Anonymous.  [Front cover].

IEEE Transactions on Visualization and Computer Graphics, 28(1):i, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:FCb

[Ano22i] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 28 (5):C1, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:FCc

[Ano22j] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 28 (6):i, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:FCd

[Ano22k] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 28 (11):i, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IIP

[Ano22l] Anonymous. IEEE ISMAR 2022 paper reviewers for journal papers. IEEE Transactions on Visualization and Computer Graphics, 28(11): viii–ix, November 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IISb

[Ano22m] Anonymous. IEEE ISMAR 2022 Science & Technology Program Committee Members for journal papers. IEEE Transactions on Visualization and Computer Graphics, 28 (11):viii, November 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IISa

[Ano22n] Anonymous. IEEE ISMAR 2022 steering committee members. IEEE Transactions on Visualization and Computer Graphics, 28(11):vii, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IVG

[Ano22o] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC):

http://vgtc.org/.  IEEE

Transactions on Visualization and Computer Graphics, 28(1): xxiv, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IVI

[Ano22p] Anonymous. IEEE VR 2022 International Program Committee for Journal Papers. IEEE Transactions on Visualization and Computer Graphics, 28 (5):x, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IVJ

[Ano22q] Anonymous. IEEE VR 2022 journal paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 28 (5):xi, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IVS

[Ano22r] Anonymous. IEEE VR 2022 Steering Committee members. IEEE Transactions on Visualization and Computer Graphics, 28(5):ix, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:IVV

[Ano22s] Anonymous. IEEE VR 2022 Visualization and Graphics Technical Committee (VGTC) statement. IEEE Transactions on Visualization and Computer Graphics, 28(5):viii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:TCa

[Ano22t] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 28(1):ii–xi, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:TCb

[Ano22u] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 28(5):iii–v, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:TCc

[Ano22v] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 28(6):ii–iii, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:TCd

[Ano22w] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 28(11):ii–iii, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VVR

[Ano22x] Anonymous. VGTC Virtual Reality Academy. IEEE Transactions on Visualization and Computer Graphics, 28(5):xix– xx, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VAC

[Ano22y] Anonymous. VIS 2021 Area Curation Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1): xxxvi, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VBP

[Ano22z] Anonymous. VIS 2021 Best Papers Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1):xli– xliii, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VCC

[Ano22-27] Anonymous. VIS 2021 Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1): xxxi–xxxiii, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VEC

[Ano22-28] Anonymous. VIS 2021 Executive Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1): xxxv, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VPC

[Ano22-29] Anonymous. VIS 2021 Program Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1):xxxvii–xl, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VR

[Ano22-30] Anonymous. VIS 2021 Reviewers. IEEE Transactions on Visualization and Computer Graphics, 28(1):xliv–xlviii, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2022:VSC

[Ano22-31] Anonymous. VIS 2021 Steering Committee. IEEE Transactions on Visualization and Computer Graphics, 28(1):xxxiv, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVVa

[Ano23a] Anonymous. 2022 IEEE VGTC Visualization Service Award. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxvii, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VVD

[Ano23b] Anonymous. 2022 VGTC Visualization Dissertation Award. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxxii, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VVL

[Ano23c] Anonymous. 2022 VGTC Visualization Lifetime Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxvi, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VVS

[Ano23d]  Anonymous.  2022 VGTC

Visualization Significant New Researcher Award. IEEE Transactions on Visualization and Computer Graphics, 29(1): xxviii–xxix, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VVT

[Ano23e] Anonymous. 2022 VGTC Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxx– xxxi, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:ICV

[Ano23f] Anonymous. 2023 IEEE Conference on Virtual Reality and 3D user interfaces. IEEE Transactions on Visualization and Computer Graphics, 29(5):ii, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IISa

[Ano23g] Anonymous. 2023 IEEE International Symposium on Mixed and Augmented Reality. IEEE Transactions on Visualization and Computer Graphics, 29 (11):i, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:FC

[Ano23h] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 29(1): C1, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:FM

[Ano23i] Anonymous. Front matter.

IEEE Transactions on Visualization and Computer Graphics, 29(6):i, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IIP

[Ano23j] Anonymous. IEEE ISMAR 2023 - paper reviewers for journal papers. IEEE Transactions on Visualization and Computer Graphics, 29(11):ix–x, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IISc

[Ano23k] Anonymous. IEEE ISMAR 2023 Science & Technology Program Committee Members for journal papers. IEEE Transactions on Visualization and Computer Graphics, 29 (11):viii, November 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IISb

[Ano23l] Anonymous. IEEE ISMAR 2023 Steering Committee Members. IEEE Transactions on Visualization and Computer Graphics, 29(11):vii, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVG

[Ano23m] Anonymous. IEEE Visualization and Graphics Technical Community (VGTC). IEEE Transactions on Visualization and Computer Graphics, 29(1): xxv, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVI

[Ano23n] Anonymous. IEEE VR 2023 international program supercommittee. IEEE Transactions on Visualization and Computer Graphics, 29(5):xii–xiv, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVP

[Ano23o] Anonymous. IEEE VR 2023 paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 29(5):xv– xvii, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVS

[Ano23p] Anonymous. IEEE VR 2023 steering committee members. IEEE Transactions on Visualization and Computer Graphics, 29(5):xi, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:IVVb

[Ano23q] Anonymous. IEEE VR 2023 visualization and graphics technical committee (VGTC) statement. IEEE Transactions on Visualization and Computer Graphics, 29(5):x, May 2023. CODEN ITVGEA. ISSN 1077-

Anonymous:2023:P

[Ano23r] Anonymous. Preface. IEEE Transactions on Visualization and Computer Graphics, 29(1): xvi–xxiv, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:TCa

[Ano23s] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 29(5):iii–vi, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:TCb

[Ano23t] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 29(6):ii, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:TCc

[Ano23u] Anonymous. Table of contents.

IEEE Transactions on Visualization and Computer Graphics, 29(11):ii–iv, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VBP

[Ano23v] Anonymous. VIS 2022 Best Papers Committee. IEEE Transactions on Visualization and Computer Graphics, 29(1): xlii, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VCC

[Ano23w] Anonymous. VIS 2022 Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 29(1): xxxiii–xxxv, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VC

[Ano23x] Anonymous. VIS 2022 contents. IEEE Transactions on Visualization and Computer Graphics, 29(1):iii–xiii, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VEC

[Ano23y] Anonymous. VIS 2022 Executive Committee. IEEE Transactions on Visualization and Computer Graphics, 29(1): xxxvii, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VPC

[Ano23z] Anonymous. VIS 2022 Program Committee. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxxviii– xli, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VR

[Ano23-27] Anonymous. VIS 2022 Reviewers. IEEE Transactions on Visualization and Computer Graphics, 29(1):xliii–xliv, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2023:VSC

[Ano23-28] Anonymous. VIS 2022 Steering Committee. IEEE Transactions on Visualization and Computer Graphics, 29(1):xxxvi, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:RL

[Ano24a]  Anonymous.  2023 review-

ers list. IEEE Transactions on Visualization and Computer Graphics, 30(4):2023– 2032, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VVD

[Ano24b] Anonymous. 2023 VGTC Visualization Dissertation Award. IEEE Transactions on Visualization and Computer Graphics, 30(1):xxxiii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VVL

[Ano24c] Anonymous. 2023 VGTC Visualization Lifetime Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 30(1):xxviii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VVSa

[Ano24d] Anonymous. 2023 VGTC Visualization Service Award. IEEE Transactions on Visualization and Computer Graphics, 30(1): xxix, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VVSb

[Ano24e] Anonymous. 2023 VGTC Visualization Significant New Researcher Award. IEEE Transactions on Visualization and Computer Graphics, 30(1):xxx– xxxi, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VVT

[Ano24f] Anonymous. 2023 VGTC Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 30(1): xxxii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:BP

[Ano24g] Anonymous. Blank page. IEEE Transactions on Visualization and Computer Graphics, 30 (5):Bii, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:ITV

[Ano24h]  Anonymous.  IEEE Trans-

actions on Visualization and Computer Graphics. IEEE Transactions on Visualization and Computer Graphics, 30(5): i, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:FC

[Ano24i] Anonymous. Front cover. IEEE Transactions on Visualization and Computer Graphics, 30(1): C1, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVG

[Ano24j] Anonymous. IEEE Visualization and Graphics Technical Committee (VGTC). IEEE Transactions on Visualization and Computer Graphics, 30(1): xxvii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVC

[Ano24k] Anonymous. IEEE VR 2024 Committee. IEEE Transactions on Visualization and Computer Graphics, 30(5):xiv, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVI

[Ano24l] Anonymous. IEEE VR 2024 International Program Super Committee. IEEE Transactions on Visualization and Computer Graphics, 30(5):xiii, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVP

[Ano24m] Anonymous. IEEE VR 2024 paper reviewers. IEEE Transactions on Visualization and Computer Graphics, 30(5):xvi– xviii, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVS

[Ano24n] Anonymous. IEEE VR 2024 Steering Committee members. IEEE Transactions on Visualization and Computer Graphics, 30(5):xii, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:IVV

[Ano24o] Anonymous. IEEE VR 2024 Visualization and Graphics Technical Committee (VGTC) statement. IEEE Transactions on Visualization and Computer Graphics, 30(5):xi, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:P

[Ano24p] Anonymous. Preface. IEEE Transactions on Visualization and Computer Graphics, 30(1): xvii–xxvi, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:TC

[Ano24q] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 30(1):i–xiii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:TCb

[Ano24r] Anonymous. Table of contents. IEEE Transactions on Visualization and Computer Graphics, 30(5):iii–vii, May 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VBP

[Ano24s] Anonymous. VIS 2023 Best Papers Committee. IEEE Transactions on Visualization and Computer Graphics, 30(1): xliii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VCC

[Ano24t] Anonymous. VIS 2023 Conference Committee. IEEE Transactions on Visualization and Computer Graphics, 30 (1):xxxiv–xxxvi, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VEC

[Ano24u] Anonymous. VIS 2023 Executive Committee. IEEE Transactions on Visualization and Computer Graphics, 30(1): xxxviii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VPC

[Ano24v] Anonymous. VIS 2023 Program Committee. IEEE Transactions on Visualization and Computer Graphics, 30(1):xxxix–xlii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VR

[Ano24w] Anonymous. VIS 2023 reviewers. IEEE Transactions on Visualization and Computer Graphics, 30(1):xliv–xlvii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Anonymous:2024:VSC

[Ano24x] Anonymous. VIS 2023 Steering Committee. IEEE Transactions on Visualization and Computer Graphics, 30(1):xxxvii, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Arunkumar:2024:IIE

[APBB24] Anjana Arunkumar, Lace Padilla, Gi-Yeul Bae, and Chris Bryan. Image or information? Examining the nature and impact of visualization perceptual classification. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1030–1040, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Assor:2024:HNV

[APHD24] Ambre Assor, Arnaud Prouzeau, Martin Hachet, and Pierre Dragicevic. Handling nonvisible referents in situated visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1336–1346, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Arefin:2022:ECS

[APP^{+}22] Mohammed Safayet Arefin, Nate Phillips, Alexander Plopski, Joseph L. Gabbard, and J. Edward Swan. The effect of context switching, focal switching distance, binocular and monocular viewing, and transient focal blur on human performance in optical see-through augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(5):2014–2025, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Arunkumar:2023:PTV

[APSB23] Anjana Arunkumar, Andrea Pinceti, Lalitha Sankar, and Chris Bryan. PMU Tracker: a visualization platform for epicentric event propagation analysis in the power grid. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1081–1090, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Anderson:2022:ACV

[AR22] Cary L. Anderson and Anthony C. Robinson. Affective congruence in visualization design: Influences on reading categorical maps. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2867–2878, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Abdelaal:2023:CEB

[ASA^{+}23] Moataz Abdelaal, Nathan D.

Schiele, Katrin Angerbauer, Kuno Kurzhals, Michael Sedlmair, and Daniel Weiskopf. Comparative evaluation of bipartite, node-link, and matrixbased network representations. IEEE Transactions on Visualization and Computer Graphics, 29(1):896–906, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Adams:2022:SLC

[ASCR^{+}22] Haley Adams, Jeanine Stefanucci, Sarah Creem-Regehr, Grant Pointon, William Thompson, and Bobby Bodenheimer. Shedding light on cast shadows: an investigation of perceived ground contact in AR and VR. IEEE Transactions on Visualization and Computer Graphics, 28(12):4624–4639, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Arora:2022:AIT

[ASP^{+}22] Nikunj Arora, Markku Suomalainen, Matti Pouke, Evan G. Center, Katherine J. Mimnaugh, Alexis P. Chambers, Sakaria Pouke, and Steven M. LaValle. Augmenting immersive telepresence experience with a virtual body. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2135–2145, May 2022. CODEN ITVGEA. ISSN 1077-2626.

AlKadi:2023:UBN

[ASSB^{+}23] Mashael AlKadi, Vanessa Serrano, James Scott-Brown, Catherine Plaisant, Jean-Daniel Fekete, Uta Hinrichs, and Benjamin Bach. Understanding barriers to network exploration with visualization: a report from the trenches. IEEE Transactions on Visualization and Computer Graphics, 29(1):907– 917, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Abdelrahman:2023:FCN

[AT23] Hayam Abdelrahman and Yiying Tong. Fast computation of neck-like features. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5384–5393, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Al-Thelaya:2021:MGD

[ATAS21] K. Al-Thelaya, M. Agus, and J. Schneider. The mixture graph — a data structure for compressing, rendering, and querying segmentation histograms. IEEE Transactions on Visualization and Computer Graphics, 27(2):645– 655, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Aoki:2024:TCO

[ATHI24] Hiroto Aoki, Takumi Tochimoto, Yuichi Hiroi, and Yuta Itoh. Towards co-operative beaming displays: Dual steering projectors for extended projection volume and head orientation range. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2309–2318, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Athawale:2024:CSP

[ATK^{+}24] Tushar M. Athawale, Bryan Triana, Tanmay Kotha, Dave Pugmire, and Paul Rosen. A comparative study of the perceptual sensitivity of topological visualizations to feature variations. IEEE Transactions on Visualization and Computer Graphics, 30(1):1074– 1084, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Adhikary:2021:TEV

[AV21] Jiban Adhikary and Keith Vertanen. Text entry in virtual environments using speech and a midair keyboard. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2648–2658, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Akiyama:2021:RRE

[AYA^{+}21] R. Akiyama, G. Yamamoto, T. Amano, T. Taketomi, A. Plopski, C. Sandor, and H. Kato. Robust reflectance estimation for projection-based appearance control in a dynamic light environment. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2041–2055, March 2021. CO-DEN ITVGEA. ISSN 1077-

Aristidou:2023:RDM

[AYA^{+}23] Andreas Aristidou, Anastasios Yiannakidis, Kfir Aberman, Daniel Cohen-Or, Ariel Shamir, and Yiorgos Chrysanthou. Rhythm is a dancer: Music-driven motion synthesis with global structure. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3519–3534, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Azmandian:2022:VSB

[AYG^{+}22] Mahdi Azmandian, Rhys Yahata, Timofey Grechkin, Jerald Thomas, and Evan Suma Rosenberg. Validating simulationbased evaluation of redirected walking systems. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2288–2298, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Azmandian:2022:ARC

[AYGR22] Mahdi Azmandian, Rhys Yahata, Timofey Grechkin, and Evan Suma Rosenberg. Adaptive redirection: a contextaware redirected walking metastrategy. IEEE Transactions on Visualization and Computer Graphics, 28(5):2277– 2287, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Adhikari:2023:ICT

[AZA^{+}23] Ashu Adhikari, Daniel Zielasko, Ivan Aguilar, Alexander Bretin, Ernst Kruijff, Markus von der Heyde, and Bernhard E. Riecke.

Integrating continuous and teleporting VR locomotion into a seamless HyperJump paradigm. IEEE Transactions on Visualization and Computer Graphics, 29(12):5265–5281, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Bock:2020:OSA

[BAC^{+}20] A. Bock, E. Axelsson, J. Costa, G. Payne, M. Acinapura, V. Trakinski, C. Emmart, C. Silva, C. Hansen, and A. Ynnerman. OpenSpace: a system for astrographics. IEEE Transactions on Visualization and Computer Graphics, 26(1): 633–642, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bolte:2021:VVV

[BB21] Fabian Bolte and Stefan Bruckner. Vis-a-Vis: Visual exploration of visualization source code evolution. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3153–3167, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Boban:2024:CDO

[BBH24] Loën Boban, Ronan Boulic, and Bruno Herbelin. In case of doubt, one follows one’s self: The implicit guidance of the embodied self-avatar. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2109–2118, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Banerjee:2022:SSC

[BBL^{+}22] Neil T. Banerjee, Alex J. Baughman, Shu-Yu Lin, Zoë A. Witte, David M. Klaus, and Allison P. Anderson. Side-by-side comparison of human perception and performance using augmented, hybrid, and virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):4787– 4796, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Brument:2021:UMS

[BBM^{+}21] Hugo Brument, Gerd Bruder, Maud Marchal, Anne Hélène Olivier, and Ferran Argelaguet. Understanding, modeling and simulating unintended positional drift during repetitive steering navigation tasks in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(11):4300– 4310, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Bernal-Berdun:2023:DSD

[BBMM^{+}23] Edurne Bernal-Berdun, Daniel Martin, Sandra Malpica, Pedro J. Perez, Diego Gutierrez, Belen Masia, and Ana Serrano. D-SAV360: a dataset of gaze scanpaths on 360^{∘} ambisonic videos. IEEE Transactions on Visualization and Computer Graphics, 29(11):4350– 4360, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Batch:2024:WHM

[BBRE24] Andrea Batch, Peter W. S.

Butcher, Panagiotis D. Ritsos, and Niklas Elmqvist. Wizualization: a hard magic visualization system for immersive and ubiquitous analytics. IEEE Transactions on Visualization and Computer Graphics, 30(1): 507–517, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Bozgeyikli:2023:EHF

[BBSC23] Lal Lila Bozgeyikli, Evren Bozgeyikli, Christopher Schnell, and Jaclynn Clark. Exploring horizontally flipped interaction in virtual reality for improving spatial ability. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4514–4524, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Braun:2024:RHN

[BBSvL24] Daniel Braun, Rita Borgo, Max Sondag, and Tatiana von Landesberger. Reclaiming the horizon: Novel visualization designs for time-series data with large value ranges. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1161–1171, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Bernal-Berdun:2024:MIH

[BBVS^{+}24] Edurne Bernal-Berdun, Mateo Vallejo, Qi Sun, Ana Serrano, and Diego Gutierrez. Modeling the impact of headbody rotations on audio-visual spatial perception for virtual reality applications. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2624–2632, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Buttussi:2021:LPV

[BC21] F. Buttussi and L. Chittaro. Locomotion in place in virtual reality: a comparative evaluation of joystick, teleport, and leaning. IEEE Transactions on Visualization and Computer Graphics, 27(1):125–136, January 2021. CODEN ITVGEA. ISSN 1077-2626.

biowska:2022:RDI

[bÇ22] Izabela M. Go biowska and Arzu Çöltekin. Rainbow Dash: Intuitiveness, interpretability and memorability of the rainbow color scheme in visualization. IEEE Transactions on Visualization and Computer Graphics, 28(7):2722– 2733, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Buck:2022:IEA

[BCB22] Lauren E. Buck, Soumyajit Chakraborty, and Bobby Bodenheimer. The impact of embodiment and avatar sizing on personal space in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2102–2113, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Batch:2020:TNS

[BCC^{+}20] A. Batch, A. Cunningham, M. Cordeil, N. Elmqvist, T. Dwyer, B. H. Thomas, and K. Marriott. There is no spoon: Evaluating performance, space use, and presence with expert domain users in immersive analytics. IEEE Transactions on Visualization and Computer Graphics, 26(1):536–546, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Battle:2020:RLT

[BCN^{+}20] L. Battle, R. J. Crouser, A. Nakeshimana, A. Montoly, R. Chang, and M. Stonebraker. The role of latency and task complexity in predicting visual search behavior. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1246–1255, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bartram:2022:UDU

[BCT22] Lyn Bartram, Michael Correll, and Melanie Tory. Untidy data: The unreasonable effectiveness of tables. IEEE Transactions on Visualization and Computer Graphics, 28(1):686–696, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Balci:2022:PFF

[BD22] Hasan Balci and Ugur Dogrusoz. fCoSE: a fast compound graph layout algorithm with constraint support. IEEE Transactions on Visualization and Computer Graphics, 28(12):4582–4593, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Biswas:2021:PDD

[BDL^{+}21] Ayan Biswas, Soumya Dutta, Earl Lawrence, John Patchett, Jon C. Calhoun, and James Ahrens. Probabilistic data-driven sampling via multicriteria importance analysis. IEEE Transactions on Visualization and Computer Graphics, 27(12):4439–4454, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Beecham:2021:DED

[BDRW21] Roger Beecham, Jason Dykes, Chris Rooney, and William Wong. Design exposition discussion documents for rich design discourse in applied visualization. IEEE Transactions on Visualization and Computer Graphics, 27(8):3451–3462, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Block:2023:IVP

[BER^{+}23] Jeremy E. Block, Shaghayegh Esmaeili, Eric D. Ragan, John R. Goodall, and G. David Richardson. The influence of visual provenance representations on strategies in a collaborative hand-off data analysis scenario. IEEE Transactions on Visualization and Computer Graphics, 29(1):1113–1123, January 2023. CODEN ITVGEA. ISSN 1077-

Bach:2023:DDP

[BFAR^{+}23] Benjamin Bach, Euan Freeman, Alfie Abdul-Rahman, Cagatay Turkay, Saiful Khan, Yulei Fan, and Min Chen. Dashboard design patterns. IEEE Transactions on Visualization and Computer Graphics, 29(1):342– 352, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Butaslac:2023:SRA

[BFS^{+}23] Isidro III Mendoza Butaslac, Yuichiro Fujimoto, Taishi Sawabe, Masayuki Kanbara, and Hirokazu Kato. Systematic review of augmented reality training systems. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5062–5082, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Biener:2024:HTI

[BFS^{+}24] Verena Biener, Forouzan Farzinnejad, Rinaldo Schuster, Seyedmasih Tabaei, Leon Lindlein, Jinghui Hu, Negar Nouri, John J. Dudley, Per Ola Krlstensson, Jörg Müller, and Jens Grubert. Hold tight: Identifying behavioral patterns during prolonged work in VR through video analysis. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2796–2806, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Bae:2024:CDP

[BFY^{+}24] S. Sandra Bae, Takanori Fujiwara, Anders Ynnerman, Ellen Yi-Luen Do, Michael L. Rivera, and Danielle Albers Szafir. A computational design pipeline to fabricate sensing network physicalizations. IEEE Transactions on Visualization and Computer Graphics, 30(1):913– 923, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Berton:2022:CNV

[BGB^{+}22] Florian Berton, Fabien Grzeskowiak, Alexandre Bonneau, Alberto Jovane, Marco Aggravi, Ludovic Hoyet, Anne-Hélène Olivier, Claudio Pacchierotti, and Julien Pettré. Crowd navigation in VR: Exploring haptic rendering of collisions. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2589–2601, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Biener:2022:PAP

[BGS^{+}22] Verena Biener, Travis Gesslein, Daniel Schneider, Felix Kawala, Alexander Otte, Per Ola Kristensson, Michel Pahud, Eyal Ofek, Cuauhtli Campos, Matjaž Kljun, Klen Čopič Pucihar, and Jens Grubert. PoVR-Point: Authoring presentations in mobile virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2069–2079, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Bertucci:2023:DVE

[BHA^{+}23] Donald Bertucci, Md Montaser Hamid, Yashwanthi Anand, Anita Ruangrotsakun, Delyar Tabatabai, Melissa Perez, and Minsuk Kahng. DendroMap: Visual exploration of largescale image datasets for machine learning with treemaps. IEEE Transactions on Visualization and Computer Graphics, 29(1):320–330, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Bhatia:2022:AAM

[BHM^{+}22] Harsh Bhatia, Duong Hoang, Nate Morrical, Valerio Pascucci, Peer-Timo Bremer, and Peter Lindstrom. AMM: Adaptive multilinear meshes. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2350–2363, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Berseth:2021:IAD

[BHU^{+}21] G. Berseth, B. Haworth, M. Usman, D. Schaumann, M. Khayatkhoei, M. Kapadia, and P. Faloutsos. Interactive architectural design with diverse solution exploration. IEEE Transactions on Visualization and Computer Graphics, 27(1): 111–124, January 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Bai:2023:LGP

[BHY^{+}23] Jiayang Bai, Zhen He, Shan Yang, Jie Guo, Zhenyu Chen, Yan Zhang, and Yanwen Guo. Local-to-global panorama inpainting for locale-aware indoor lighting prediction. IEEE Transactions on Visualization and Computer Graphics, 29(11):4405–4416, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Bang:2021:LVT

[BJCL21] Kiseung Bang, Youngjin Jo, Minseok Chae, and Byoungho Lee. LensIet VR: Thin, flat and wide-FOV virtual reality display using Fresnel lens and lenslet array. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2545–2554, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Butcher:2021:VWB

[BJR21] Peter W. S. Butcher, Nigel W. John, and Panagiotis D. Ritsos. VRIA: a Web-based framework for creating immersive analytics experiences. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3213–3225, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Brehmer:2022:JSR

[BK22] Matthew Brehmer and Robert Kosara. From jam session to recital: Synchronous communication and collaboration around data in organizations. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1139–1149, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Brehmer:2022:GDI

[BKH22] Matthew Brehmer, Robert Kosara, and Carmen Hull. Generative design inspiration for glyphs with diatoms. IEEE Transactions on Visualization and Computer Graphics, 28(1): 389–399, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Biener:2022:QEW

[BKN^{+}22] Verena Biener, Snehanjali Kalamkar, Negar Nouri, Eyal Ofek, Michel Pahud, John J. Dudley, Jinghui Hu, Per Ola Kristensson, Maheshya Weerasinghe, Klen Čopič Pucihar, Matja Kljun, Stephan Streuber, and Jens Grubert. Quantifying the effects of working in VR for one week. IEEE Transactions on Visualization and Computer Graphics, 28(11):3810– 3820, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Bhatia:2021:VFD

[BKPB21] Harsh Bhatia, Robert M. Kirby, Valerio Pascucci, and Peer-Timo Bremer. Vector field decompositions using multiscale Poisson kernel. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3781–3793, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Bach:2024:COD

[BKR^{+}24] Benjamin Bach, Mandy Keck, Fateme Rajabiyazdi, Tatiana Losev, Isabel Meirelles, Jason Dykes, Robert S. Laramee, Mashael AlKadi, Christina Stoiber, Samuel Huron, Charles Perin, Luiz Morais, Wolfgang Aigner, Doris Kosminsky, Magdalena Boucher, Søren Knudsen, Areti Manataki, Jan Aerts, Uta Hinrichs, Jonathan C. Roberts, and Sheelagh Carpendale. Challenges and opportunities in data visualization education: a call to action. IEEE Transactions on Visualization and Computer Graphics, 30(1): 649–660, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Bok:2022:APC

[BKS22] Jinwook Bok, Bohyoung Kim, and Jinwook Seo. Augmenting parallel coordinates plots with color-coded stacked histograms. IEEE Transactions on Visualization and Computer Graphics, 28(7):2563–2576, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Bressa:2022:WSS

[BKT^{+}22] Nathalie Bressa, Henrik Korsgaard, Aurélien Tabard, Steven Houben, and Jo Vermeulen. What’s the situation with situated visualization? A survey and perspectives on situatedness. IEEE Transactions on Visualization and Computer Graphics, 28(1):107–117, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Bender:2020:IFB

[BKWK20] J. Bender, T. Kugelstadt, M. Weiler, and D. Koschier. Implicit frictional boundary handling for SPH. IEEE Transactions on Visualization and Computer Graphics, 26(10): 2982–2993, October 2020. CO-DEN ITVGEA. ISSN 1077-2626.

Bako:2023:UHD

[BLBL23] Hannah K. Bako, Xinyi Liu, Leilani Battle, and Zhicheng Liu. Understanding how designers find and use data visualization examples. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1048–1058, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Bruder:2023:HSA

[BLE^{+}23] Valentin Bruder, Matthew Larsen, Thomas Ertl, Hank Childs, and Steffen Frey. A hybrid in situ approach for cost efficient image database generation. IEEE Transactions on Visualization and Computer Graphics, 29(9):3788– 3798, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Brehmer:2020:CEA

[BLIC20] M. Brehmer, B. Lee, P. Isenberg, and E. K. Choe. A comparative evaluation of animation and small multiples for trend visualization on mobile phones. IEEE Transactions on Visualization and Computer Graphics, 26(1):364–374, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Boorboor:2023:NFV

[BMA^{+}23] Saeed Boorboor, Shawn Mathew, Mala Ananth, David Talmage, Lorna W. Role, and Arie E.

Kaufman. NeuRegenerate: a framework for visualizing neurodegeneration. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1625–1637, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Bruder:2020:ERP

[BMFE20] V. Bruder, C. Müller, S. Frey, and T. Ertl. On evaluating runtime performance of interactive visualizations. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2848–2862, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Bi:2020:DLB

[BMWD20] H. Bi, T. Mao, Z. Wang, and Z. Deng. A deep learning-based framework for intersectional traffic simulation and editing. IEEE Transactions on Visualization and Computer Graphics, 26(7):2335–2348, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Bolte:2021:SVM

[BNRB21] Fabian Bolte, Mahsan Nourani, Eric D. Ragan, and Stefan Bruckner. SplitStreams: a visual metaphor for evolving hierarchies. IEEE Transactions on Visualization and Computer Graphics, 27(8):3571–3584, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Benmahdjoub:2021:VEI

[BNWvW21] Mohamed Benmahdjoub, Wiro J.

Niessen, Eppo B. Wolvius, and Theo van Walsum. Virtual extensions improve perceptionbased instrument alignment using optical see-through devices. IEEE Transactions on Visualization and Computer Graphics, 27(11):4332–4341, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Bletterer:2022:LGB

[BPA22] Arnaud Bletterer, Frédéric Payan, and Marc Antonini. A local graph-based structure for processing gigantic aggregated 3D point clouds. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2822–2833, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Bouzbib:2023:WTB

[BPL23] Elodie Bouzbib, Claudio Pacchierotti, and Anatole Lécuyer. When tangibles become deformable: Studying pseudostiffness perceptual thresholds in a VR grasping task. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2743–2752, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Bodenheimer:2023:IVM

[BPQW23] Bobby Bodenheimer, Voicu Popescu, John Quarles, and Lili Wang. IEEE VR 2023 message from the program chairs and guest editors. IEEE Transactions on Visualization and Computer Graphics, 29(5): viii–ix, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Baumgartl:2021:SPZ

[BPW^{+}21] T. Baumgartl, M. Petzold, M. Wunderlich, M. Hohn, D. Archambault, M. Lieser, A. Dalpke, S. Scheithauer, M. Marschollek, V. M. Eichel, N. T. Mutters, H. Consortium, and T. V. Landesberger. In search of patient zero: Visual analytics of pathogen transmission pathways in hospitals. IEEE Transactions on Visualization and Computer Graphics, 27(2):711–721, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Ballester-Ripoll:2020:TTC

[BRLP20] R. Ballester-Ripoll, P. Lindstrom, and R. Pajarola. TTHRESH: Tensor compression for multidimensional visual data. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2891–2903, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Benda:2024:EET

[BRLR24] Brett Benda, Benjamin Rheault, Yanna Lin, and Eric D. Ragan. Examining effects of technique awareness on the detection of remapped hands in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2651– 2661, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Battle:2021:SRD

[BS21] L. Battle and C. Scheidegger. A structured review of data management technology for interactive visualization and analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2):1128–1138, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Batmaz:2022:ETA

[BS22] Anil Ufuk Batmaz and Wolfgang Stuerzlinger. Effective throughput analysis of different task execution strategies for mid-air Fitts’ tasks in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3939– 3947, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Bader:2020:EVA

[BSB^{+}20] R. Bader, M. Sprenger, N. Ban, S. Rüdisühli, C. Schär, and T. Günther. Extraction and visual analysis of potential vorticity banners around the Alps. IEEE Transactions on Visualization and Computer Graphics, 26(1):259–269, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Biener:2020:BSI

[BSG^{+}20] V. Biener, D. Schneider, T. Gesslein, A. Otte, B. Kuth, P. O. Kristensson, E. Ofek, M. Pahud, and J. Grubert. Breaking the screen: Interaction across touchscreen boundaries in virtual reality for mobile knowledge workers. IEEE Transactions on Visualization and Computer Graphics, 26(12):3490–3502, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Behrisch:2020:GUG

[BSP20] M. Behrisch, T. Schreck, and H. Pfister. GUIRO: Userguided matrix reordering. IEEE Transactions on Visualization and Computer Graphics, 26(1): 184–194, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Beck:2020:GEI

[BSW20] F. Beck, J. Seo, and C. Wang. Guest Editors introduction: Special section on IEEE PacificVis 2020. IEEE Transactions on Visualization and Computer Graphics, 26(6):2142– 2143, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Bakar:2023:CFR

[BTHL23] Muhammad Abu Bakar, Yu-Ting Tsai, Hao-Han Hsueh, and Elena Carolina Li. CrowbarLimbs: a fatigue-reducing virtual reality text entry metaphor. IEEE Transactions on Visualization and Computer Graphics, 29(5):2806– 2815, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Bollen:2023:CSD

[BTL23] Brian Bollen, Pasindu Tennakoon, and Joshua A. Levine.

Computing a stable distance on merge trees. IEEE Transactions on Visualization and Computer Graphics, 29(1):1168– 1177, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Benmahdjoub:2023:EAV

[BTv^{+}23] Mohamed Benmahdjoub, Abdullah Thabit, Marie-Lise C. van Veelen, Wiro J. Niessen, Eppo B. Wolvius, and Theo van Walsum. Evaluation of AR visualization approaches for catheter insertion into the ventricle cavity. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2434–2445, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Babadi:2022:LTA

[BvdPLH22] Amin Babadi, Michiel van de Panne, C. Karen Liu, and Perttu Hämäläinen. Learning task-agnostic action spaces for movement optimization. IEEE Transactions on Visualization and Computer Graphics, 28(12):4700–4712, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Bauerle:2021:NVG

[BvOR21] Alex Bäuerle, Christian van Onzenoodt, and Timo Ropinski. Net2Vis: a visual grammar for automatically generating publication-tailored CNN architecture visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2980–2991, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Bhargava:2022:DHD

[BVV^{+}22] Ayush Bhargava, Roshan Venkatakrishnan, Rohith Venkatakrishnan, Hannah Solini, Kathryn Lucaites, Andrew C. Robb, Christopher C. Pagano, and Sabarish V. Babu. Did I hit the door? Effects of self-avatars and calibration in a person-plus-virtual-object system on perceived frontal passability in VR. IEEE Transactions on Visualization and Computer Graphics, 28(12):4198–4210, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Bhargava:2023:CST

[BVV^{+}23] Ayush Bhargava, Rohith Venkatakrishnan, Roshan Venkatakrishnan, Kathryn Lucaites, Hannah Solini, Andrew C. Robb, Christopher C. Pagano, and Sabarish V. Babu. Can I squeeze through? Effects of self-avatars and calibration in a person-plus-virtual-object system on perceived lateral passability in VR. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2348–2357, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Bae:2023:CVL

[BVY^{+}23] S. Sandra Bae, Rishi Vanukuru, Ruhan Yang, Peter Gyory, Ran Zhou, Ellen Yi-Luen Do, and Danielle Albers Szafir. Cultivating visualization literacy for children through curiosity and play. IEEE Transactions on Visualization and Computer Graphics, 29(1):257–267, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Betancourt:2023:ECS

[BWCT23] Julio Betancourt, Baptiste Wojtkowski, Pedro Castillo, and Indira Thouvenin. Exocentric control scheme for robot applications: an immersive virtual reality approach. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3392–3404, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Bigelow:2021:GPR

[BWI21] A. Bigelow, K. Williams, and K. E. Isaacs. Guidelines for pursuing and revealing data abstractions. IEEE Transactions on Visualization and Computer Graphics, 27(2):1503– 1513, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Bauer:2023:FFV

[BWM23] David Bauer, Qi Wu, and Kwan-Liu Ma. FoVolNet: Fast volume rendering using foveated deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 29(1):515–525, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Bauer:2024:PFN

[BWM24] David Bauer, Qi Wu, and Kwan-Liu Ma. Photon field networks for dynamic real-time volumetric global illumination. IEEE Transactions on Visualization and Computer Graphics, 30(1):975–985, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Brubach:2022:BPB

[BWWL22] Larissa Brübach, Franziska Westermeier, Carolin Wienrich, and Marc Erich Latoschik. Breaking plausibility without breaking presence — evidence for the multi-layer nature of plausibility. IEEE Transactions on Visualization and Computer Graphics, 28(5):2267– 2276, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Borland:2020:SBT

[BWZ^{+}20] D. Borland, W. Wang, J. Zhang, J. Shrestha, and D. Gotz. Selection bias tracking and detailed subset comparison for high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 26(1): 429–439, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Burns:2022:DPE

[BXF^{+}22] Alyxander Burns, Cindy Xiong, Steven Franconeri, Alberto Cairo, and Narges Mahyar. Designing with pictographs: Envision topics without sacrificing understanding. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4515–4530, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Bao:2022:RTC

[BXQ^{+}22] Hujun Bao, Weijian Xie, Quanhao Qian, Danpeng Chen, Shangjin Zhai, Nan Wang, and Guofeng Zhang. Robust tightlycoupled visual-inertial odometry with pre-built maps in high latency situations. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2212–2222, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Bian:2021:IMP

[BXZ^{+}21] R. Bian, Y. Xue, L. Zhou, J. Zhang, B. Chen, D. Weiskopf, and Y. Wang. Implicit multidimensional projection of local subspaces. IEEE Transactions on Visualization and Computer Graphics, 27(2):1558– 1568, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Borland:2021:SBC

[BZKG21] D. Borland, J. Zhang, S. Kaul, and D. Gotz. Selection-bias-corrected visualization via dynamic reweighting. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1481–1491, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Bishop:2020:CVE

[BZP^{+}20] F. Bishop, J. Zagermann, U. Pfeil, G. Sanderson, H. Reiterer, and U. Hinrichs. Construct-A-Vis: Exploring the freeform visualization processes of children. IEEE Transactions on Visualization and Computer Graphics, 26(1):451–460, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Bartolomeo:2021:SBV

[BZSD21] S. D. Bartolomeo, Y. Zhang, F. Sheng, and C. Dunne. Sequence braiding: Visual overviews of temporal event sequences and attributes. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1353–1363, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Bu:2021:SIR

[BZW^{+}21] C. Bu, Q. Zhang, Q. Wang, J. Zhang, M. Sedlmair, O. Deussen, and Y. Wang. SineStream: Improving the readability of streamgraphs by minimizing sine illusion effects. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1634–1643, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2020:LEI

[CAA^{+}20] S. Chen, N. Andrienko, G. Andrienko, L. Adilova, J. Barlet, J. Kindermann, P. H. Nguyen, O. Thonnard, and C. Turkay. LDA ensembles for interactive exploration and categorization of behaviors. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2775–2792, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Chen:2021:CBP

[CAA^{+}21] S. Chen, N. Andrienko, G. Andrienko, J. Li, and X. Yuan. Co-Bridges: Pair-wise visual connection and comparison for multi-item data streams. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1612–1622, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Ceneda:2022:SMY

[CAGM22] Davide Ceneda, Alessio Arleo, Theresia Gschwandtner, and Silvia Miksch. Show me your face: Towards an automated method to provide timely guidance in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 28(12):4570– 4581, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Cao:2023:PQC

[CAR^{+}23] Chunxiao Cao, Zili An, Zhong Ren, Dinesh Manocha, and Kun Zhou. A psychoacoustic quality criterion for path-traced sound propagation. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5422–5433, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Chittaro:2022:LST

[CB22] Luca Chittaro and Fabio Buttussi. Learning safety through public serious games: a study of prepare for impact on a very large, international sample of players. IEEE Transactions on Visualization and Computer Graphics, 28(3):1573– 1584, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Concha:2021:IVO

[CBB^{+}21] Alejo Concha, Michael Burri, Jesús Briales, Christian Forster, and Luc Oth. Instant visual odometry initialization for mobile AR. IEEE Transactions on Visualization and Computer Graphics, 27(11):4226– 4235, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Chakhchoukh:2023:UHV

[CBB23] Mehdi Chakhchoukh, Nadia Boukhelifa, and Anastasia Bezerianos. Understanding how in-visualization provenance can support trade-off analysis. IEEE Transactions on Visualization and Computer Graphics, 29(9):3758–3774, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Costa:2021:IVA

[CBE^{+}21] J. Costa, A. Bock, C. Emmart, C. Hansen, A. Ynnerman, and C. Silva. Interactive visualization of atmospheric effects for celestial bodies. IEEE Transactions on Visualization and Computer Graphics, 27(2):785– 795, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Cross:2023:UER

[CBHR^{+}23] Jamie Cross, Christine Boag-Hodgson, Tim Ryley, Timothy J Mavin, and Leigh Ellen Potter. Using extended reality in flight simulators: a literature review. IEEE Transactions on Visualization and Computer Graphics, 29(9):3961– 3975, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Cabric:2024:EYG

[CBL^{+}24] Florent Cabric, Margrét Vilborg Bjarnadóttir, Meng Ling, Gudbjörg Linda Rafnsdóttir, and Petra Isenberg. Eleven years of gender data visualization: a step towards more inclusive gender representation. IEEE Transactions on Visualization and Computer Graphics, 30(1):316–326, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Cooper:2022:EHD

[CBP22] Victoria L. Cooper, James C.

Bieron, and Pieter Peers. Estimating homogeneous datadriven BRDF parameters from a reflectance map under known natural lighting. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4289–4303, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Choudhary:2023:VFE

[CBW23] Zubin Datta Choudhary, Gerd Bruder, and Gregory F. Welch. Visual facial enhancements can significantly improve speech perception in the presence of noise. IEEE Transactions on Visualization and Computer Graphics, 29(11):4751– 4760, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Ceneda:2024:HAD

[CCEA^{+}24] Davide Ceneda, Christopher Collins, Mennatallah El-Assady, Silvia Miksch, Christian Tominski, and Alessio Arleo. A heuristic approach for dual Expert/End-User evaluation of guidance in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 30(1):997–1007, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Chen:2024:RDR

[CCL^{+}24] Zhutian Chen, Daniele Chiappalupi, Tica Lin, Yalong Yang, Johanna Beyer, and Hanspeter Pfister. RL-L: a deep reinforcement learning approach intended for AR label placement in dynamic scenarios. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1347–1357, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Coelho:2020:PVA

[CCM20] D. Coelho, I. Chase, and K. Mueller. PeckVis: a visual analytics tool to analyze dominance hierarchies in small groups. IEEE Transactions on Visualization and Computer Graphics, 26(4):1650– 1660, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cannavo:2021:ETL

[CCP^{+}21] A. Cannavò, D. Calandra, F. G.

Pratticò, V. Gatteschi, and F. Lamberti. An evaluation testbed for locomotion in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(3):1871– 1889, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Calvo:2023:MES

[CCPM23] Luz Calvo, Fernando Cucchietti, and Mario Pérez-Montoro. Measuring the effectiveness of static maps to communicate changes over time. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4243–4255, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Cortes:2021:EBR

[CCS^{+}21] C. A. T. Cortes, H.-T. Chen, D. L. Sturnieks, J. Garcia, S. R. Lord, and C.-T. Lin. Evaluating balance recovery techniques for users wearing headmounted display in VR. IEEE Transactions on Visualization and Computer Graphics, 27(1): 204–215, January 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2024:CNA

[CCS^{+}24] Qing Chen, Nan Chen, Wei Shuai, Guande Wu, Zhe Xu, Hanghang Tong, and Nan Cao. Calliope-Net: Automatic generation of graph data facts via annotated node-link diagrams.

IEEE Transactions on Visualization and Computer Graphics, 30(1):562–572, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Corvo:2021:VAH

[CCW^{+}21] A. Corvò, H. S. Garcia Caballero, M. A. Westenberg, M. A. van Driel, and J. J. van Wijk. Visual analytics for hypothesis-driven exploration in computational pathology. IEEE Transactions on Visualization and Computer Graphics, 27(10):3851–3866, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:RWI

[CDBM22] Kun-Ting Chen, Tim Dwyer, Benjamin Bach, and Kim Marriott. Rotate or wrap? Interactive visualisations of cyclical data on cylindrical or toroidal topologies. IEEE Transactions on Visualization and Computer Graphics, 28(1):727–736, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:IRG

[CDS^{+}22] Shaoyu Chen, Budmonde Duinkharjav, Xin Sun, Li-Yi Wei, Stefano Petrangeli, Jose Echevarria, Claudio Silva, and Qi Sun. Instant reality: Gaze-contingent perceptual optimization for 3D virtual reality streaming. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2157–2167, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Chao:2020:DBF

[CDX^{+}20] Q. Chao, Z. Deng, Y. Xiao, D. He, Q. Miao, and X. Jin. Dictionary-based fidelity measure for virtual traffic. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1490–1501, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2023:ASP

[CDZ^{+}23] Geng Chen, Hang Dai, Tao Zhou, Jianbing Shen, and Ling Shao. Automatic Schelling point detection from meshes. IEEE Transactions on Visualization and Computer Graphics, 29(6):2926–2939, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Combe:2024:EIV

[CFDN24] Theo Combe, Rebecca Fribourg, Lucas Detto, and Jean-Marie Norm. Exploring the influence of virtual avatar heads in mixed reality on social presence, performance and user experience in collaborative tasks. IEEE Transactions on Visualization and Computer Graphics, 30(5):2206–2216, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Crisan:2022:PGD

[CFGM22] Anamaria Crisan, Shannah E.

Fisher, Jennifer L. Gardy, and Tamara Munzner. GEViTRec: Data reconnaissance through recommendation using a domainspecific visualization prevalence design space. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4855–4872, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Crisan:2021:PDB

[CFGT21] A. Crisan, B. Fiore-Gartland, and M. Tory. Passing the data baton: a retrospective analysis on data science work and workers. IEEE Transactions on Visualization and Computer Graphics, 27(2):1860– 1870, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Chai:2021:VDP

[CFL21] S. Chai, X.-M. Fu, and L. Liu. Voting for distortion points in geometric processing. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2469–2480, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Cheymol:2023:BMR

[CFL^{+}23] Antonin Cheymol, Rebecca Fribourg, Anatole Lécuyer, Jean-Marie Normand, and Ferran Argelaguet. Beyond my real body: Characterization, impacts, applications and perspectives of dissimilar avatars in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4426– 4437, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Chowdhury:2021:VDS

[CFQ21] Tanvir Irfan Chowdhury, Sharif Mohammad Shahnewaz Ferdous, and John Quarles. VR disability simulation reduces implicit bias towards persons with disabilities. IEEE Transactions on Visualization and Computer Graphics, 27(6): 3079–3090, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:MDD

[CFZ^{+}23] Jiali Chen, Changjie Fan, Zhimeng Zhang, Gongzheng Li, Zeng Zhao, Zhigang Deng, and Yu Ding. A music-driven deep generative adversarial model for guzheng playing animation. IEEE Transactions on Visualization and Computer Graphics, 29(2):1400–1414, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:OIB

[CFZZ23] Beijia Chen, Hongbo Fu, Kun Zhou, and Youyi Zheng. OrthoAligner: Image-based teeth alignment prediction via latent style manipulation. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3617–3629, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Caserman:2020:SFB

[CGAG20] P. Caserman, A. Garcia-Agundez, and S. Göbel. A survey of full-body motion reconstruction in immersive virtual reality applications. IEEE Transactions on Visualization and Computer Graphics, 26 (10):3089–3108, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Cui:2024:AAV

[CGD^{+}24] Yuan Cui, Lily W. Ge, Yiren Ding, Fumeng Yang, Lane Harrison, and Matthew Kay. Adaptive assessment of visualization literacy. IEEE Transactions on Visualization and Computer Graphics, 30(1):628–637, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Chen:2024:UIM

[CGT^{+}24] Changjian Chen, Yukai Guo, Fengyuan Tian, Shilong Liu, Weikai Yang, Zhaowei Wang, Jing Wu, Hang Su, Hanspeter Pfister, and Shixia Liu. A unified interactive model evaluation for classification, object detection, and instance segmentation in computer vision. IEEE Transactions on Visualization and Computer Graphics, 30(1): 76–86, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:RST

[CGZ^{+}20] X. Chen, T. Ge, J. Zhang, B. Chen, C. Fu, O. Deussen, and Y. Wang. A recursive subdivision technique for sampling multi-class scatterplots. IEEE Transactions on Visualization and Computer Graphics, 26(1): 729–738, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2024:ASS

[CHSC24] Jun-Jie Chen, Huan-Chang Hung, Yu-Ru Sun, and Jung-Hong Chuang. APF-S2T: Steering to target redirection walking based on artificial potential fields. IEEE Transactions on Visualization and Computer Graphics, 30(5):2464– 2473, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Chandio:2024:HFP

[CIA24] Yasra Chandio, Victoria Interrante, and Fatima M. Anwar. Human factors at play: Understanding the impact of conditioning on presence and reaction time in mixed reality. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2400–2410, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Cakmak:2022:MVS

[CJS^{+}22] Eren Cakmak, Dominik Jäckle, Tobias Schreck, Daniel A. Keim, and Johannes Fuchs. Multiscale visualization: a structured literature analysis. IEEE Transactions on Visualization and Computer Graphics, 28(12):4918–4929, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:DLF

[CJY^{+}23] Yeyao Chen, Gangyi Jiang, Mei Yu, Haiyong Xu, and Yo-Sung Ho. Deep light field spatial super-resolution using heterogeneous imaging. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4183–4197, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Cha:2022:TRC

[CK22] Ick-Hoon Cha and Hyeong-Seok Ko. Tanglement resolution in clothing simulation with explicit convergence. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2764–2775, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Cheng:2023:PIT

[CKQ^{+}23] Furui Cheng, Mark S Keller, Huamin Qu, Nils Gehlenborg, and Qianwen Wang. Polyphony: an interactive transfer learning framework for single-cell data analysis. IEEE Transactions on Visualization and Computer Graphics, 29(1): 591–601, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2020:SSS

[CLA^{+}20] S. Chen, J. Li, G. Andrienko, N. Andrienko, Y. Wang, P. H. Nguyen, and C. Turkay. Supporting story synthesis: Bridging the gap between visual analytics and storytelling. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2499–2516, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:RMM

[CLCY20] S. Chen, S. Li, S. Chen, and X. Yuan. R-Map: a map metaphor for visualizing information reposting process in social media. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1204–1214, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cheng:2022:VCD

[CLD^{+}22] Furui Cheng, Dongyu Liu, Fan Du, Yanna Lin, Alexandra Zytek, Haomin Li, Huamin Qu, and Kalyan Veeramachaneni. VBridge: Connecting the dots between features and data to explain healthcare models. IEEE Transactions on Visualization and Computer Graphics, 28(1):378–388, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:GOH

[CLFL22] Yixin Chen, Wei Li, Rui Fan, and Xiaopei Liu. GPU optimization for high-quality kinetic fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3235–3251, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Coeurjolly:2021:DSR

[CLG21] David Coeurjolly, Jacques-Olivier Lachaud, and Pierre Gueth. Digital surface regularization with guarantees. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2896–2907, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Chao:2023:CFB

[CLH^{+}23] Qianwen Chao, Pengfei Liu, Yi Han, Yingying Lin, Chaoneng Li, Qiguang Miao, and Xiaogang Jin. A calibrated forcebased model for mixed traffic simulation. IEEE Transactions on Visualization and Computer Graphics, 29(3):1664– 1677, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:ARE

[CLL^{+}20] X. Chen, Y. Li, X. Luo, T. Shao, J. Yu, K. Zhou, and Y. Zheng. AutoSweep: Recovering 3D editable objects from a single photograph. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1466–1475, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2021:VCF

[CLL^{+}21] Jian Chen, Meng Ling, Rui Li, Petra Isenberg, Tobias Isenberg, Michael Sedlmair, Torsten Möller, Robert S. Laramee, Han-Wei Shen, Katharina Wünsche, and Qiru Wang. VIS30K: a collection of figures and tables from IEEE visualization conference publications. IEEE Transactions on Visualization and Computer Graphics, 27(9):3826–3833, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Choi:2023:DUV

[CLL^{+}23] JunYoung Choi, Sang-Eun Lee, YeIn Lee, Eunji Cho, Sunghoe Chang, and Won-Ki Jeong. DXplorer: a unified visualization framework for interactive dendritic spine analysis using 3D morphological features. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1424–1437, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Cao:2021:ANR

[CLS^{+}21] Kelei Cao, Mengchen Liu, Hang Su, Jing Wu, Jun Zhu, and Shixia Liu. Analyzing the noise robustness of deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 27(7):3289– 3304, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2024:MDS

[CLW^{+}24] Chen Chen, Bongshin Lee, Yunhai Wang, Yunjeong Chang, and Zhicheng Liu. Mystique: Deconstructing SVG charts for layout reuse. IEEE Transactions on Visualization and Computer Graphics, 30(1):447– 457, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Cao:2023:TBT

[CLX^{+}23] Anqi Cao, Ji Lan, Xiao Xie, Hongyu Chen, Xiaolong Zhang, Hui Zhang, and Yingcai Wu. Team-Builder: Toward more effective lineup selection in soccer. IEEE Transactions on Visualization and Computer Graphics, 29(12):5178– 5193, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Cao:2024:CTC

[CLZ^{+}24] Xin Cao, Jia Li, Panpan Zhao, Jiachen Li, and Xueying Qin. Corr-Track: Categorylevel 6D pose tracking with soft-correspondence matrix estimation. IEEE Transactions on Visualization and Computer Graphics, 30(5):2173– 2183, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:PUN

[CMF^{+}22] Shaoyu Chen, Fabio Miranda, Nivan Ferreira, Marcos Lage, Harish Doraiswamy, Corinne Brenner, Connor Defanti, Michael Koutsoubis, Luc Wilson, Ken Perlin, and Claudio Silva. UrbanRama: Navigating cities in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):4685–4699, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Chatzimparmpas:2020:VIA

[CMK20] A. Chatzimparmpas, R. M. Martins, and A. Kerren. tviSNE: Interactive assessment and interpretation of t-SNE projections. IEEE Transactions on Visualization and Computer Graphics, 26(8):2696–2714, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Chatzimparmpas:2021:SAD

[CMKK21] A. Chatzimparmpas, R. M. Martins, K. Kucher, and A. Kerren. StackGenVis: Alignment of data, algorithms, and models for stacking ensemble learning using performance metrics. IEEE Transactions on Visualization and Computer Graphics, 27(2):1547– 1557, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Chatzimparmpas:2022:PFV

[CMKK22] Angelos Chatzimparmpas, Rafael M.

Martins, Kostiantyn Kucher, and Andreas Kerren. FeatureEnVi: Visual analytics for feature engineering using stepwise selection and semi-automatic extraction approaches. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1773–1791, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Cai:2024:RDM

[CML24] Zhuojiang Cai, Yuhan Ma, and Feng Lu. Robust dualmodal speech keyword spotting for XR headsets. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2507–2516, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Cheng:2021:DDE

[CMQ21] F. Cheng, Y. Ming, and H. Qu.

DECE: Decision explorer with counterfactual explanations for machine learning models. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1438–1447, February 2021. CO-DEN ITVGEA. ISSN 1077-

Cibulski:2023:CCP

[CMSK23] Lena Cibulski, Thorsten May, Johanna Schmidt, and Jörn Kohlhammer. COMPO*SED: Composite parallel coordinates for co-dependent multiattribute choices. IEEE Transactions on Visualization and Computer Graphics, 29(10): 4047–4061, October 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Ceja:2021:TSA

[CMXF21] C. R. Ceja, C. M. McColeman, C. Xiong, and S. L. Franconeri. Truth or square: Aspect ratio biases recall of position encodings. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1054–1062, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Corbalan-Navarro:2022:POT

[CNAA^{+}22] David Corbalán-Navarro, Juan L. Aragón, Martí Anglada, Enrique de Lucas, Joan-Manuel Parcerisa, and Antonio González. Omega-Test: a predictive early-Z culling to improve the graphics pipeline energy-efficiency. IEEE Transactions on Visualization and Computer Graphics, 28(12):4375–4388, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Chaudhary:2022:PTR

[CNB^{+}22] Aayush K. Chaudhary, Nitinraj Nair, Reynold J. Bailey, Jeff B. Pelz, Sachin S.

Talathi, and Gabriel J. Diaz.

Temporal RIT-Eyes: From real

infrared eye-images to synthetic sequences of gaze behavior. IEEE Transactions on Visualization and Computer Graphics, 28(11):3948–3958, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Chan:2020:MBV

[CNC^{+}20] G. Y. Chan, L. G. Nonato, A. Chu, P. Raghavan, V. Aluru, and C. T. Silva. Motion browser: Visualizing and understanding complex upper limb movement under obstetrical brachial plexus injuries. IEEE Transactions on Visualization and Computer Graphics, 26(1): 981–990, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Curiel:2024:VRS

[CNK^{+}24] Rodrigo Cerecero Curiel, Takuto Nakamura, Hideaki Kuzuoka, Takafumi Kanaya, Cosima Prahm, and Keigo Matsumoto. Virtual reality self coembodiment: an alternative to mirror therapy for post-stroke upper limb rehabilitation. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2390–2399, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Chung:2022:GVI

[CNY22] Haeyong Chung, Santhosh Nandhakumar, and Seungwon Yang. GridSet: Visualizing individual elements and attributes for analysis of settyped data. IEEE Transactions on Visualization and Computer Graphics, 28(8):2983–2998, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Cardenas:2023:MTS

[COFJ23] José L. Cárdenas, Carlos J. Ogayar, Francisco R. Feito, and Juan M. Jurado. Modeling of the 3D tree skeleton using real-world data: a survey. IEEE Transactions on Visualization and Computer Graphics, 29(12):4920–4935, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:RIS

[COZ^{+}23] Longfei Chen, Yang Ouyang, Haipeng Zhang, Suting Hong, and Quan Li. RISeer: Inspecting the status and dynamics of regional industrial structure via visual analytics. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1070–1080, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Cashman:2020:AVC

[CPCS20] D. Cashman, A. Perer, R. Chang, and H. Strobelt. Ablate, variate, and contemplate: Visual analytics for discovering neural architectures. IEEE Transactions on Visualization and Computer Graphics, 26(1):863– 873, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2020:HSP

[CPD20] Z. Chen, D. Panozzo, and J. Dumas. Half-space power diagrams and discrete surface offsets. IEEE Transactions on Visualization and Computer Graphics, 26(10):2970– 2981, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Cauquis:2024:IWM

[CPD^{+}24] Julien Cauquis, Etienne Peillard, Lionel Dominjon, Thierry Duval, and Guillaume Moreau. Investigating whether the mass of a tool replica influences virtual training learning outcomes. IEEE Transactions on Visualization and Computer Graphics, 30(5):2411–2421, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Chundury:2022:TUS

[CPR^{+}22] Pramod Chundury, Biswaksen Patnaik, Yasmin Reyazuddin, Christine Tang, Jonathan Lazar, and Niklas Elmqvist. Towards understanding sensory substitution for accessible visualization: an interview study. IEEE Transactions on Visualization and Computer Graphics, 28(1):1084–1094, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Chowdhury:2022:WLI

[CQ22] Tanvir Irfan Chowdhury and John Quarles. A wheelchair locomotion interface in a VR disability simulation reduces implicit bias. IEEE Transactions on Visualization and Computer Graphics, 28(12):4658– 4670, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Castro:2022:EEU

[CQHP22] Spencer C. Castro, P. Samuel Quinan, Helia Hosseinpour, and Lace Padilla. Examining effort in 1D uncertainty communication using individual differences in working memory and NASA-TLX. IEEE Transactions on Visualization and Computer Graphics, 28(1):411– 421, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Choo:2023:EGE

[CRH23] Jaegul Choo, Timo Ropinski, and Yifan Hu. Editorial: Guest Editors introduction: Special section on IEEE PacificVis 2023. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2847–2848, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Chundury:2024:TSD

[CRJ^{+}24] Pramod Chundury, Yasmin Reyazuddin, J. Bern Jordan, Jonathan Lazar, and Niklas Elmqvist. TactualPlot: Spatializing data as sound using sensory substitution for touchscreen accessibility. IEEE Transactions on Visualization and Computer Graphics, 30(1): 836–846, January 2024. CO-DEN ITVGEA. ISSN 1077-

Castelo:2024:AVA

[CRM^{+}24] Sonia Castelo, Joao Rulff, Erin McGowan, Bea Steers, Guande Wu, Shaoyu Chen, Iran Roman, Roque Lopez, Ethan Brewer, Chen Zhao, Jing Qian, Kyunghyun Cho, He He, Qi Sun, Huy Vo, Juan Bello, Michael Krone, and Claudio Silva. ARGUS: Visualization of AI-Assisted task guidance in AR. IEEE Transactions on Visualization and Computer Graphics, 30(1):1313– 1323, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Chalbi:2020:CFA

[CRP^{+}20] A. Chalbi, J. Ritchie, D. Park, J. Choi, N. Roussel, N. Elmqvist, and F. Chevalier. Common fate for animated transitions in visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):386–396, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Carr:2022:OAD

[CRWA22] Hamish A. Carr, Oliver Rübel, Gunther H. Weber, and James P. Ahrens. Optimization and augmentation for data parallel contour trees. IEEE Transactions on Visualization and Computer Graphics, 28(10):3471– 3485, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Cao:2022:GEI

[CRZ22] Nan Cao, Timo Ropinski, and Jian Zhao. Guest Editors’ introduction: Special section on IEEE PacificVis 2022. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2299–2300, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Congdon:2023:MCR

[CS23] Ben J. Congdon and Anthony Steed. Monte-carlo redirected walking: Gain selection through simulated walks. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2637–2646, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Choukroun:2020:HOS

[CSBK20] Y. Choukroun, A. Shtern, A. Bronstein, and R. Kimmel. Hamiltonian operator for spectral shape analysis. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1320–1331, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Choudhry:2021:OUT

[CSC^{+}21] A. Choudhry, M. Sharma, P. Chundury, T. Kapler, D. W. S. Gray, N. Ramakrishnan, and N. Elmqvist. Once upon a time in visualization: Understanding the use of textual narratives for causality. IEEE Transactions on Visualization and Computer Graphics, 27(2):1332–1342, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:PNC

[CSC^{+}22] Ran Chen, Xinhuan Shu, Jiahui Chen, Di Weng, Junxiu Tang, Siwei Fu, and Yingcai Wu. Nebula: a coordinating grammar of graphics. IEEE Transactions on Visualization and Computer Graphics, 28(12):4127–4140, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Crnovrsanin:2021:SAS

[CSCM21] T. Crnovrsanin, Shilpika, S. Chandrasegaran, and K.-L. Ma. Staged animation strategies for online dynamic networks. IEEE Transactions on Visualization and Computer Graphics, 27(2): 539–549, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Cuenca:2022:VVP

[CSIP22] Erick Cuenca, Arnaud Sallaberry, Dino Ienco, and Pascal Poncelet. VERTIGo: a visual platform for querying and exploring large multilayer networks. IEEE Transactions on Visualization and Computer Graphics, 28(3):1634– 1647, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Cakmak:2021:MSV

[CSJ^{+}21] E. Cakmak, U. Schlegel, D. Jäckle, D. Keim, and T. Schreck. Multiscale snapshots: Visual analysis of temporal summaries in dynamic graphs. IEEE Transactions on Visualization and Computer Graphics, 27(2):517–527, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2021:SSD

[CSL^{+}21] H. Chen, U. Soni, Y. Lu, V. Huroyan, R. Maciejewski, and S. Kobourov. Same stats, different graphs: Exploring the space of graphs in terms of graph properties. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2056–2072, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Chang:2023:MVA

[CSL^{+}23] Baofeng Chang, Guodao Sun, Tong Li, Houchao Huang, and Ronghua Liang. MUSE: Visual analysis of musical semantic sequence. IEEE Transactions on Visualization and Computer Graphics, 29(9):4015– 4030, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Covaci:2023:MVU

[CSM^{+}23] Alexandra Covaci, Estêvão B. Saleme, Gebremariam Mesfin, Ioan-Sorin Comsa, Ramona Trestian, Celso A. S. Santos, and George Ghinea. Multisensory 360^{∘} videos under varying resolution levels enhance presence. IEEE Transactions on Visualization and Computer Graphics, 29(4):2093– 2101, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:MAG

[CSW^{+}20] Z. Chen, Y. Su, Y. Wang, Q. Wang, H. Qu, and Y. Wu. MARVisT: Authoring glyphbased visualization in mobile augmented reality. IEEE Transactions on Visualization and Computer Graphics, 26(8): 2645–2658, August 2020. CO-DEN ITVGEA. ISSN 1077-2626.

Chai:2024:PAD

[CSWZ24] Yujin Chai, Tianjia Shao, Yanlin Weng, and Kun Zhou. Personalized audio-driven 3D facial animation via style-content disentanglement. IEEE Transactions on Visualization and Computer Graphics, 30(3): 1803–1820, March 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2022:VLF

[CSX^{+}22] Qing Chen, Fuling Sun, Xinyue Xu, Zui Chen, Jiazhe Wang, and Nan Cao. VizLinter: a linter and fixer framework for data visualization. IEEE Transactions on Visualization and Computer Graphics, 28(1):206– 216, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Cao:2021:GEI

[CTW21] Nan Cao, Holger Theisel, and Chaoli Wang. Guest Editors’ introduction: Special section on IEEE PacificVis 2021. IEEE Transactions on Visualization and Computer Graphics, 27(6):

2768–2769, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Collaris:2023:SSA

[CvW23] Dennis Collaris and Jarke J. van Wijk. StrategyAtlas: Strategy analysis for machine learning interpretability. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2996–3008, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:GVE

[CWB^{+}20] C. Chen, C. Wang, X. Bai, P. Zhang, and C. Li. GenerativeMap: Visualization and exploration of dynamic density maps via generative learning model. IEEE Transactions on Visualization and Computer Graphics, 26(1):216–226, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Cui:2022:MIA

[CWH^{+}22] Weiwei Cui, Jinpeng Wang, He Huang, Yun Wang, Chin-Yew Lin, Haidong Zhang, and Dongmei Zhang. A mixedinitiative approach to reusing infographic charts. IEEE Transactions on Visualization and Computer Graphics, 28(1): 173–183, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2023:RTT

[CWH^{+}23] Ran Chen, Di Weng, Yanwei Huang, Xinhuan Shu, Jiayi Zhou, Guodao Sun, and Yingcai Wu. Rigel: Transforming tabular data by declarative mapping. IEEE Transactions on Visualization and Computer Graphics, 29(1):128–138, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Chu:2023:REC

[CWL23] Yiyao Chu, Wencheng Wang, and Lei Li. Robustly extracting concise 3D curve skeletons by enhancing the capture of prominent features. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3472–3488, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2024:FVA

[CWO^{+}24] Longfei Chen, He Wang, Yang Ouyang, Yang Zhou, Naiyu Wang, and Quan Li. FSLens: a visual analytics approach to evaluating and optimizing the spatial layout of fire stations. IEEE Transactions on Visualization and Computer Graphics, 30(1):847–857, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Cox:2021:RMC

[CWR21] Wesley Cox, Lyndon While, and Mark Reynolds. A review of methods to compute Minkowski operations for geometric overlap detection. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3377–3396, August 2021. CO-DEN ITVGEA. ISSN 1077-

Chen:2020:MPC

[CWS^{+}20] H. Chen, M. Wei, Y. Sun, X. Xie, and J. Wang. Multipatch collaborative point cloud denoising via low-rank recovery with graph constraint. IEEE Transactions on Visualization and Computer Graphics, 26(11):3255–3270, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Carr:2021:SCT

[CWS^{+}21] H. A. Carr, G. H. Weber, C. M. Sewell, O. Rübel, P. Fasel, and J. P. Ahrens. Scalable contour tree computation by data parallel peak pruning. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2437–2454, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Chan:2023:PSD

[CWSJ23] Tsz Tai Chan, Yixuan Wang, Richard Hau Yue So, and Jerry Jia. Predicting subjective discomfort associated with lens distortion in VR headsets during vestibulo-ocular response to VR scenes. IEEE Transactions on Visualization and Computer Graphics, 29(8):3656–3669, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:TAI

[CWW^{+}20] Z. Chen, Y. Wang, Q. Wang, Y. Wang, and H. Qu. Towards automated infographic design: Deep learning-based auto-extraction of extensible timeline. IEEE Transactions on Visualization and Computer Graphics, 26(1):917–926, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2021:IGC

[CWW^{+}21] Changjian Chen, Zhaowei Wang, Jing Wu, Xiting Wang, Lan-Zhe Guo, Yu-Feng Li, and Shixia Liu. Interactive graph construction for graph-based semi-supervised learning. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3701–3716, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:TBC

[CWW^{+}22] Changjian Chen, Jing Wu, Xiaohan Wang, Shouxing Xiang, Song-Hai Zhang, Qifeng Tang, and Shixia Liu. Towards better caption supervision for object detection. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1941–1954, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:FVP

[CWW^{+}23] Wei Chen, Yating Wei, Zhiyong Wang, Shuyue Zhou, Bingru Lin, and Zhiguang Zhou. Federated visualization: a privacypreserving strategy for aggregated visual query. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2901–2913, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Cashman:2021:CVA

[CXD^{+}21] D. Cashman, S. Xu, S. Das, F. Heimerl, C. Liu, S. R. Humayoun, M. Gleicher, A. Endert, and R. Chang. CAVA: a visual analytics system for exploratory columnar data augmentation using knowledge graphs. IEEE Transactions on Visualization and Computer Graphics, 27(2):1731– 1741, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Cheng:2023:FAI

[CXW^{+}23] Haojie Cheng, Chunxiao Xu, Jiajun Wang, Zhenxin Chen, and Lingxiao Zhao. Fast and accurate illumination estimation using LDR panoramic images for realistic rendering. IEEE Transactions on Visualization and Computer Graphics, 29(12):5235–5249, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2021:GEG

[CXXZ21] Taizhou Chen, Lantian Xu, Xianshan Xu, and Kening Zhu. GestOnHMD: Enabling gesture-based interaction on low-cost VR head-mounted display. IEEE Transactions on Visualization and Computer Graphics, 27(5):2597– 2607, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Chu:2022:TVE

[CXY^{+}22] Xiangtong Chu, Xiao Xie, Shuainan Ye, Haolin Lu, Hongguang Xiao, Zeqing Yuan, Zhutian Chen, Hui Zhang, and Yingcai Wu. TIVEE: Visual exploration and explanation of badminton tactics in immersive visualizations. IEEE Transactions on Visualization and Computer Graphics, 28(1):118– 128, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Cao:2024:AEV

[CXZ^{+}24] Anqi Cao, Xiao Xie, Mingxu Zhou, Hui Zhang, Mingliang Xu, and Yingcai Wu. Action-Evaluator: a visualization approach for player action evaluation in soccer. IEEE Transactions on Visualization and Computer Graphics, 30(1):880– 890, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:ASV

[CYC^{+}22] Zhutian Chen, Shuainan Ye, Xiangtong Chu, Haijun Xia, Hui Zhang, Huamin Qu, and Yingcai Wu. Augmenting sports videos with VisCommentator. IEEE Transactions on Visualization and Computer Graphics, 28(1):824–834, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:ERH

[CYD^{+}23] Kunyao Chen, Fei Yin, Bang Du, Baichuan Wu, and Truong Q. Nguyen. Efficient registration for human surfaces via isometric regularization on embedded deformation. IEEE Transactions on Visualization and Computer Graphics, 29(12):5020– 5032, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2021:OIA

[CYL^{+}21] Changjian Chen, Jun Yuan, Yafeng Lu, Yang Liu, Hang Su, Songtao Yuan, and Shixia Liu. OoDAnalyzer: Interactive analysis of out-of-distribution samples. IEEE Transactions on Visualization and Computer Graphics, 27(7):3335– 3349, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Chen:2020:VVA

[CYP^{+}20] Q. Chen, X. Yue, X. Plantaz, Y. Chen, C. Shi, T. Pong, and H. Qu. ViSeq: Visual analytics of learning sequence in massive open online courses. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1622–1636, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2023:SAS

[CYX^{+}23] Zhutian Chen, Qisen Yang, Xiao Xie, Johanna Beyer, Haijun Xia, Yingcai Wu, and Hanspeter Pfister. Sporthesia: Augmenting sports videos using natural language. IEEE Transactions on Visualization and Computer Graphics, 29(1): 918–928, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Chen:2020:MES

[CZC^{+}20] J. Chen, G. Zhang, W. Chiou, D. H. Laidlaw, and A. P.

Auchus. Measuring the effects of scalar and spherical colormaps on ensembles of DMRI tubes. IEEE Transactions on Visualization and Computer Graphics, 26(9):2818– 2833, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:PBS

[CZF^{+}22] Xin Chen, Jian Zhang, Chi-Wing Fu, Jean-Daniel Fekete, and Yunhai Wang. Pyramidbased scatterplots sampling for progressive and streaming data visualization. IEEE Transactions on Visualization and Computer Graphics, 28(1):593– 603, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Chen:2023:CDF

[CZF^{+}23] Rongsen Chen, Fang-Lue Zhang, Simon Finnie, Andrew Chalmers, and Taehyun Rhee. Casual 6-DoF: Free-viewpoint panorama using a handheld 360^{∘} camera. IEEE Transactions on Visualization and Computer Graphics, 29(9):3976–3988, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2022:ACC

[CZG^{+}22] Shu-Yu Chen, Jia-Qi Zhang, Lin Gao, Yue He, Shihong Xia, Min Shi, and Fang-Lue Zhang. Active colorization for cartoon line drawings. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1198–1208, February 2022. CO-DEN ITVGEA. ISSN 1077-

Choi:2021:APS

[CZGF21] Inrak Choi, Yiwei Zhao, Eric J. Gonzalez, and Sean Follmer. Augmenting perceived softness of haptic proxy objects through transient vibration and visuohaptic illusion in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(12):4387–4400, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Cornel:2023:WIH

[CZGW23] Daniel Cornel, Silvana Zechmeister, Eduard Gröller, and Jürgen Waser. Watertight incremental heightfield tessellation. IEEE Transactions on Visualization and Computer Graphics, 29(9):3888– 3899, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Chen:2021:CCP

[CZL^{+}21] X. Chen, W. Zeng, Y. Lin, H. M. Al-maneea, J. Roberts, and R. Chang. Composition and configuration patterns in multiple-view visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(2):1514–1524, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Chalmers:2021:RRM

[CZMR21] Andrew Chalmers, Junhong Zhao, Daniel Medeiros, and Taehyun Rhee. Reconstructing reflection maps using a stacked-CNN for mixed reality rendering. IEEE Transactions on Visualization and Computer Graphics, 27(10): 4073–4084, October 2021. CO-DEN ITVGEA. ISSN 1077-2626. Chakravarthula:2021:GCR

[CZT^{+}21] Praneeth Chakravarthula, Zhan Zhang, Okan Tursun, Piotr Didyk, Qi Sun, and Henry Fuchs. Gaze-contingent retinal speckle suppression for perceptually-matched foveated holographic displays. IEEE Transactions on Visualization and Computer Graphics, 27(11):4194–4203, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Cui:2020:TVA

[CZW^{+}20] W. Cui, X. Zhang, Y. Wang, H. Huang, B. Chen, L. Fang, H. Zhang, J. Lou, and D. Zhang. Text-to-Viz: Automatic generation of infographics from proportion-related natural language statements. IEEE Transactions on Visualization and Computer Graphics, 26(1): 906–916, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Chen:2020:LDL

[CZY^{+}20] Z. Chen, W. Zeng, Z. Yang, L. Yu, C. Fu, and H. Qu. LassoNet: Deep lasso-selection of 3D point clouds. IEEE Transactions on Visualization and Computer Graphics, 26(1): 195–204, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Delaforge:2023:ETE

[DAB^{+}23] Alexis Delaforge, Jérôme Azé, Sandra Bringay, Caroline Mollevi, Arnaud Sallaberry, and Maximilien Servajean. EBBE-Text: Explaining neural networks by exploring text classification decision boundaries. IEEE Transactions on Visualization and Computer Graphics, 29(10): 4154–4171, October 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Devkota:2021:CUC

[DAK^{+}21] S. Devkota, P. Aschwanden, A. Kunen, M. Legendre, and K. E. Isaacs. CcNav: Understanding compiler optimizations in binary code. IEEE Transactions on Visualization and Computer Graphics, 27(2): 667–677, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

dosAnjos:2023:MII

[dAWi^{+}23] Rafael Kuffner dos Anjos, David Walton, Kaan Ak it, Sebastian Friston, David Swapp, Anthony Steed, and Tobias Ritschel. Metameric inpainting for image warping. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5511–5522, December 2023. CODEN ITVGEA. ISSN 1077-2626.

diBartolomeo:2022:SLM

[dBRGD22] Sara di Bartolomeo, Mirek Riedewald, Wolfgang Gatterbauer, and Cody Dunne. STRATISFIMAL LAYOUT: a modular optimization model for laying out layered node-link network visualizations. IEEE Transactions on Visualization and Computer Graphics, 28(1): 324–334, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Debarba:2022:PVB

[DCC22] Henrique Galvan Debarba, Sylvain Chagué, and Caecilia Charbonnier. On the plausibility of virtual body animation features in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1880–1893, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Deng:2023:RDP

[DCM^{+}23] Dazhen Deng, Weiwei Cui, Xiyu Meng, Mengye Xu, Yu Liao, Haidong Zhang, and Yingcai Wu. Revisiting the design patterns of composite visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(12):5406– 5421, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Deng:2024:VLS

[DCS^{+}24] Zikun Deng, Shifu Chen, Tobias Schreck, Dazhen Deng, Tan Tang, Mingliang Xu, Di Weng, and Yingcai Wu. Visualizing large-scale spatial time series with GeoChron. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1194–1204, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Du:2023:LDS

[DCWD23] Minghan Du, Hui Cui, Yuan Wang, and Henry Been-Lirn Duh. Learning from deep stereoscopic attention for simulator sickness prediction. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1415–1423, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Dimara:2020:TBT

[DFP^{+}20] E. Dimara, S. Franconeri, C. Plaisant, A. Bezerianos, and P. Dragicevic. A task-based taxonomy of cognitive biases for information visualization. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1413–1432, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dewe:2022:MVS

[DGB^{+}22] Hayley Dewe, Janna M. Gottwald, Laura-Ashleigh Bird, Harry Brenton, Marco Gillies, and Dorothy Cowie. My virtual self: The role of movement in children’s sense of embodiment. IEEE Transactions on Visualization and Computer Graphics, 28(12):4061–4072, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Dong:2023:FRM

[DGD^{+}23] Tianyang Dong, Tieqi Gao, Yinyan Dong, Liming Wang, Kefan Hu, and Jing Fan. FREE-RDW: a multi-user redirected walking method for supporting non-forward steps. IEEE Transactions on Visualization and Computer Graphics, 29(5):2315–2325, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Dong:2021:IPT

[DGDC21] Xinghui Dong, Ying Gao, Junyu Dong, and Mike J. Chantler. The importance of phase to texture discrimination and similarity. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3755–3768, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Drouin:2020:IDE

[DGKOC20] S. Drouin, D. A. D. Giovanni, M. Kersten-Oertel, and D. L. Collins. Interaction driven enhancement of depth perception in angiographic volumes. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2247–2257, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Dwyer:2024:W

[DGW24] Tim Dwyer, Sarah Goodwin, and Michael Wybrow. Welcome. IEEE Transactions on Visualization and Computer Graphics, 30(1):xv–xvi, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Du:2022:SSA

[DHF^{+}22] Dong Du, Xiaoguang Han, Hongbo Fu, Feiyang Wu, Yizhou Yu, Shuguang Cui, and Ligang Liu. SAniHead: Sketching animal-like 3D character heads using a view-surface collaborative mesh generative network. IEEE Transactions on Visualization and Computer Graphics, 28(6):2415– 2429, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Dewez:2022:DYN

[DHLA22] Diane Dewez, Ludovic Hoyet, Anatole Lécuyer, and Ferran Argelaguet. Do you need another hand? Investigating dual body representations during anisomorphic 3D manipulation. IEEE Transactions on Visualization and Computer Graphics, 28(5):2047–2057, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Du:2022:ADS

[DHM^{+}22] Zheng-Jun Du, Shi-Sheng Huang, Tai-Jiang Mu, Qunhe Zhao, Ralph R. Martin, and Kun Xu. Accurate dynamic SLAM using CRF-based longterm consistency. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1745–1757, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Deng:2022:PFN

[DHY^{+}22] Nianchen Deng, Zhenyi He, Jiannan Ye, Budmonde Duinkharjav, Praneeth Chakravarthula, Xubo Yang, and Qi Sun. FoV-NeRF: Foveated neural radiance fields for virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3854–3864, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Dy:2022:IVD

[DIPJ22] Bianchi Dy, Nazim Ibrahim, Ate Poorthuis, and Sam Joyce. Improving visualization design for effective multi-objective decision making. IEEE Transactions on Visualization and Computer Graphics, 28(10): 3405–3416, October 2022. CO-DEN ITVGEA. ISSN 1077-2626.

David-John:2023:PPD

[DJBJ23] Brendan David-John, Kevin Butler, and Eakta Jain. Privacypreserving datasets of eyetracking samples with applications in XR. IEEE Transactions on Visualization and Computer Graphics, 29(5):2774– 2784, May 2023. CODEN ITVGEA. ISSN 1077-2626.

David-John:2021:PPA

[DJHBJ21] Brendan David-John, Diane Hosfelt, Kevin Butler, and Eakta Jain. A privacypreserving approach to streaming eye-tracking data. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2555–2565, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Dasu:2024:COD

[DKM24] Keshav Dasu, Yun-Hsin Kuo, and Kwan-Liu Ma. Characteroriented design for visual data storytelling. IEEE Transactions on Visualization and Computer Graphics, 30(1):98–108, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Duchowski:2020:UME

[DKuH20] A. T. Duchowski, K. Krejtz, J. urawska, and D. H. House. Using microsaccades to estimate task difficulty during visual search of layered surfaces. IEEE Transactions on Visualization and Computer Graphics, 26(9):2904–2918, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Dong:2022:CFM

[DLH^{+}22] Xuan Dong, Weixin Li, Xiaoyan Hu, Xiaojie Wang, and Yunhong Wang. A colorization framework for monochromecolor dual-lens systems using a deep convolutional network. IEEE Transactions on Visualization and Computer Graphics, 28(3):1469–1485, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Deng:2023:SFP

[DLP^{+}23] Zhi Deng, Yang Liu, Hao Pan, Wassim Jabi, Juyong Zhang, and Bailin Deng. Sketch2PQ: Freeform planar quadrilateral mesh design via a single sketch. IEEE Transactions on Visualization and Computer Graphics, 29(9):3826–3839, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Davidson:2023:EES

[DLW^{+}23] Kylie Davidson, Lee Lisle, Kirsten Whitley, Doug A. Bowman, and Chris North. Exploring the evolution of sensemaking strategies in immersive space to think. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5294–5307, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Dmitriev:2021:VAC

[DMBK21] K. Dmitriev, J. Marino, K. Baker, and A. E. Kaufman. Visual analytics of a computer-aided diagnosis system for pancreatic lesions. IEEE Transactions on Visualization and Computer Graphics, 27(3):2174– 2185, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Deng:2022:PFW

[DMJ^{+}22] Qixin Deng, Luming Ma, Aobo Jin, Huikun Bi, Binh Huy Le, and Zhigang Deng. Plausible 3D face wrinkle generation using variational autoencoders. IEEE Transactions on Visualization and Computer Graphics, 28(9):3113–3125, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Dasu:2021:SGR

[DMMF21] K. Dasu, K.-L. Ma, J. Ma, and J. Frazier. Sea of genes: a reflection on visualising metagenomic data for museums. IEEE Transactions on Visualization and Computer Graphics, 27(2): 935–945, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Dibene:2022:PLS

[DMTD22] Juan Carlos Dibene, Yazmin Maldonado, Leonardo Trujillo, and Enrique Dunn. Prepare for ludicrous speed: Marker-based instantaneous binocular rolling shutter localization. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2201–2211, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Dimara:2020:WID

[DP20] E. Dimara and C. Perin. What is interaction for data visualization? IEEE Transactions on Visualization and Computer Graphics, 26(1):119–129, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Daneshzand:2023:KEN

[DPC23] Foroozan Daneshzand, Charles Perin, and Sheelagh Carpendale. KiriPhys: Exploring new data physicalization opportunities. IEEE Transactions on Visualization and Computer Graphics, 29(1):225–235, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Davis:2024:RRR

[DPD^{+}24] Russell Davis, Xiaoying Pu, Yiren Ding, Brian D. Hall, Karen Bonilla, Mi Feng, Matthew Kay, and Lane Harrison. The risks of ranking: Revisiting graphical perception to model individual differences in visualization performance. IEEE Transactions on Visualization and Computer Graphics, 30(3):1756–1771, March 2024. CODEN ITVGEA. ISSN 1077-2626.

deTinguy:2021:CSI

[dPLM21] X. de Tinguy, C. Pacchierotti, A. Lécuyer, and M. Marchal. Capacitive sensing for improving contact rendering with tangible objects in VR. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2481–2487, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Do:2024:SRS

[DPM24] Tiffany D. Do, Camille Isabella Protko, and Ryan P. McMahan. Stepping into the right shoes: The effects of usermatched avatar ethnicity and gender on sense of embodiment in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2434–2443, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Dasgupta:2020:ECS

[DPR^{+}20] A. Dasgupta, J. Poco, B. Rogowitz, K. Han, E. Bertini, and C. T. Silva. The effect of color scales on climate scientists objective and subjective performance in spatial data analysis tasks. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1577–1591, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ding:2020:PFC

[DS20] O. Ding and C. Schroeder. Penalty force for coupling materials with Coulomb friction. IEEE Transactions on Visualization and Computer Graphics, 26(7):2443–2455, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Dimara:2022:CRV

[DS22] Evanthia Dimara and John Stasko. A critical reflection on visualization research: Where do decision making tasks hide? IEEE Transactions on Visualization and Computer Graphics, 28(1):1128–1138, January 2022. CODEN ITVGEA. ISSN 1077-2626.

deSouza:2024:PVA

[dSBdO^{+}24] Carolina Veiga Ferreira de Souza, Suzanna Maria Bonnet, Daniel de Oliveira, Marcio Cataldi, Fabio Miranda, and Marcos Lage. PW: a visual approach for building, managing, and analyzing weather simulation ensembles at runtime. IEEE Transactions on Visualization and Computer Graphics, 30(1):

738–747, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Dai:2023:RRA

[DSD^{+}23] Shaozhang Dai, Jim Smiley, Tim Dwyer, Barrett Ens, and Lonni Besancon. RoboHapalytics: a robot assisted haptic controller for immersive analytics. IEEE Transactions on Visualization and Computer Graphics, 29(1):451–461, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Das:2021:GCI

[DSKE21] Subhajit Das, Bahador Saket, Bum Chul Kwon, and Alex Endert. Geono-Cluster: Interactive visual cluster analysis for biologists. IEEE Transactions on Visualization and Computer Graphics, 27(12):4401– 4412, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Dubey:2023:CPP

[DST^{+}23] Rohit K. Dubey, Samuel S.

Sohn, Tyler Thrash, Christoph Hölscher, André Borrmann, and Mubbasir Kapadia. Cognitive path planning with spatial memory distortion. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3535–3549, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Duncan:2021:TBE

[DTPG21] I. K. Duncan, S. Tingsheng, S. T. Perrault, and M. T.

Gastner. Task-based effectiveness of interactive contiguous area cartograms. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2136–2152, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Doraiswamy:2021:TDH

[DTS^{+}21] H. Doraiswamy, J. Tierny, P. J. S. Silva, L. G. Nonato, and C. Silva. TopoMap: a 0-dimensional homology preserving projection of highdimensional data. IEEE Transactions on Visualization and Computer Graphics, 27(2):561– 571, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Danyluk:2022:TBC

[DUWW22] Kurtis Danyluk, Teoman Ulusoy, Wei Wei, and Wesley Willett. Touch and beyond: Comparing physical and virtual reality visualizations. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1930–1940, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Domova:2023:MTL

[DV23] Veronika Domova and Katerina Vrotsou. A model for types and levels of automation in visual analytics: a survey, a taxonomy, and examples. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3550–3568, August 2023. CO-DEN ITVGEA. ISSN 1077-

DeRose:2021:AFA

[DWB21] J. F. DeRose, J. Wang, and M. Berger. Attention flows: Analyzing and comparing attention mechanisms in language models. IEEE Transactions on Visualization and Computer Graphics, 27(2):1160– 1170, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Deng:2020:AVA

[DWC^{+}20] Z. Deng, D. Weng, J. Chen, R. Liu, Z. Wang, J. Bao, Y. Zheng, and Y. Wu. AirVis: Visual analytics of air pollution propagation. IEEE Transactions on Visualization and Computer Graphics, 26(1):800– 810, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Dhanoa:2023:FSU

[DWH^{+}23] Vaishali Dhanoa, Conny Walchshofer, Andreas Hinterreiter, Eduard Gröller, and Marc Streit. Fuzzy spreadsheet: Understanding and exploring uncertainties in tabular calculations. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1463–1477, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Deng:2022:VCA

[DWL^{+}22] Zikun Deng, Di Weng, Yuxuan Liang, Jie Bao, Yu Zheng, Tobias Schreck, Mingliang Xu, and Yingcai Wu. Visual cascade analytics of large-scale spatiotemporal data. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2486–2499, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Dasgupta:2020:SWC

[DWOB20] A. Dasgupta, H. Wang, N. O’Brien, and S. Burrows. Separating the wheat from the chaff: Comparative visual cues for transparent diagnostics of competing models. IEEE Transactions on Visualization and Computer Graphics, 26(1):1043– 1053, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Deng:2023:DID

[DWQW23] Dazhen Deng, Aoyu Wu, Huamin Qu, and Yingcai Wu. DashBot: Insight-driven dashboard generation based on deep reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 29(1): 690–700, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Deng:2023:VFG

[DWS^{+}23] Dazhen Deng, Yihong Wu, Xinhuan Shu, Jiang Wu, Siwei Fu, Weiwei Cui, and Yingcai Wu. VisImages: a finegrained expert-annotated visualization dataset. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3298–3311, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Deng:2022:CTB

[DWX^{+}22] Zikun Deng, Di Weng, Xiao Xie, Jie Bao, Yu Zheng, Mingliang Xu, Wei Chen, and Yingcai Wu. Compass: Towards better causal analysis of urban time series. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1051–1061, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Dou:2021:TSA

[DXX^{+}21] Zhiyang Dou, Shiqing Xin, Rui Xu, Jian Xu, Yuanfeng Zhou, Shuangmin Chen, Wenping Wang, Xiuyang Zhao, and Changhe Tu. Top-down shape abstraction based on greedy pole selection. IEEE Transactions on Visualization and Computer Graphics, 27 (10):3982–3993, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Dudley:2023:EPH

[DZG^{+}23] John J. Dudley, Jingyao Zheng, Aakar Gupta, Hrvoje Benko, Matt Longest, Robert Wang, and Per Ola Kristensson. Evaluating the performance of hand-based probabilistic text input methods on a midair virtual qwerty keyboard. IEEE Transactions on Visualization and Computer Graphics, 29(11):4567–4577, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Dimara:2022:UDV

[DZTF22] Evanthia Dimara, Harry Zhang, Melanie Tory, and Steven Franconeri. The unmet data visualization needs of decision makers within organizations. IEEE Transactions on Visualization and Computer Graphics, 28(12):4101–4112, December 2022. CODEN ITVGEA. ISSN 1077-2626.

El-Assady:2020:SCS

[EAKC^{+}20] M. El-Assady, R. Kehlbeck, C. Collins, D. Keim, and O. Deussen. Semantic concept spaces: Guided topic model refinement using wordembedding projections. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1001–1011, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Espadoto:2023:ULI

[EAS^{+}23] Mateus Espadoto, Gabriel Appleby, Ashley Suh, Dylan Cashman, Mingwei Li, Carlos Scheidegger, Erik W. Anderson, Remco Chang, and Alexandru C. Telea. UnProjection: Leveraging inverseprojections for visual analytics of high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1559–1572, February 2023. CO-DEN ITVGEA. ISSN 1077-

Eirich:2022:IDS

[EBJ^{+}22] Joscha Eirich, Jakob Bonart, Dominik Jäckle, Michael Sedlmair, Ute Schmid, Kai Fischbach, Tobias Schreck, and Jürgen Bernard. IRVINE: a design study on analyzing correlation patterns of electrical engines. IEEE Transactions on Visualization and Computer Graphics, 28(1):11–21, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Eulzer:2021:VHS

[EBKL21] P. Eulzer, S. Bauer, F. Kilian, and K. Lawonn. Visualization of human spine biomechanics for spinal surgery. IEEE Transactions on Visualization and Computer Graphics, 27(2): 700–710, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Elek:2021:PSA

[EBPF21] O. Elek, J. N. Burchett, J. X.

Prochaska, and A. G. Forbes. Polyphorm: Structural analysis of cosmological datasets via interactive Physarum polycephalum visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2): 806–816, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Erickson:2023:ASC

[EBW23] Austin Erickson, Gerd Bruder, and Gregory F. Welch. Analysis of the saliency of colorbased dichoptic cues in optical see-through augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(12):4936–4950, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Elhamdadi:2022:AUT

[ECR22] Hamza Elhamdadi, Shaun Canavan, and Paul Rosen. AffectiveTDA: Using topological data analysis to improve analysis and explainability in affective computing. IEEE Transactions on Visualization and Computer Graphics, 28(1):769– 779, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Eulzer:2020:TVF

[EEL^{+}20] P. Eulzer, S. Engelhardt, N. Lichtenberg, R. de Simone, and K. Lawonn. Temporal views of flattened mitral valve geometries. IEEE Transactions on Visualization and Computer Graphics, 26(1):971–980, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Elmqvist:2021:P

[EFL^{+}21] N. Elmqvist, B. Fisher, P. Lindstrom, R. Maciejewski, M. Meyer, S. Miksch, L. G. Nonato, N. Riche, H.-W. Shen, R. Westermann, J. Wood, and J. Yang. Preface. IEEE Transactions on Visualization and Computer Graphics, 27(2):xviii– xxv, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Epperson:2024:DAC

[EGMP24] Will Epperson, Vaishnavi Gorantla, Dominik Moritz, and Adam Perer. Dead or Alive: Continuous data profiling for interactive data science. IEEE Transactions on Visualization and Computer Graphics, 30(1): 197–207, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Ens:2021:UTI

[EGP^{+}21] B. Ens, S. Goodwin, A. Prouzeau, F. Anderson, F. Y. Wang, S. Gratzl, Z. Lucarelli, B. Moyle, J. Smiley, and T. Dwyer. Uplift: a tangible and immersive tabletop system for casual collaborative visual analytics. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1193–1203, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Eckelt:2023:VER

[EHA^{+}23] Klaus Eckelt, Andreas Hinterreiter, Patrick Adelberger, Conny Walchshofer, Vaishali Dhanoa, Christina Humer, Moritz Heckmann, Christian Steinparz, and Marc Streit. Visual exploration of relationships and structure in low-dimensional embeddings. IEEE Transactions on Visualization and Computer Graphics, 29(7):3312–3326, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Eisentrager:2023:EEV

[EHB^{+}23] Karl Eisenträger, Judith Haubner, Jennifer Brade, Wolfgang Einhäuser, Alexandra Bendixen, Sven Winkler, Philipp Klimant, and Georg Jahn. Evaluating the effects of virtual reality environment learning on subsequent robot teleoperation in an unfamiliar building. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2220–2229, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Erel:2023:NPM

[EIB23] Yotam Erel, Daisuke Iwai, and Amit H. Bermano. Neural projection mapping using reflectance fields. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4339–4349, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Eirich:2023:MHS

[EJS^{+}23] Joscha Eirich, Dominik Jäckle, Michael Sedlmair, Christoph Wehner, Ute Schmid, Jürgen Bernard, and Tobias Schreck. ManuKnowVis: How to support different user groups in contextualizing and leveraging knowledge repositories. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3441–3457, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Esmaeili:2023:EGP

[EKC^{+}23] Shaghayegh Esmaeili, Samia Kabir, Anthony M. Colas, Rhema P. Linder, and Eric D. Ragan. Evaluating graphical perception of visual motion for quantitative data encoding. IEEE Transactions on Visualization and Computer Graphics, 29(12):4845–4857, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Espadoto:2021:TQS

[EMK^{+}21] M. Espadoto, R. M. Martins, A. Kerren, N. S. T. Hirata, and A. C. Telea. Toward a quantitative survey of dimension reduction techniques. IEEE Transactions on Visualization and Computer Graphics, 27(3):2153–2173, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Ebner:2023:ALD

[EML^{+}23] Christoph Ebner, Peter Mohr, Tobias Langlotz, Yifan Peng, Dieter Schmalstieg, Gordon Wetzstein, and Denis Kalkofen. Off-axis layered displays: Hybrid direct-view/near-eye mixed reality with focus cues. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2816–2825, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Ebner:2022:VST

[EMM^{+}22] Christoph Ebner, Shohei Mori, Peter Mohr, Yifan Peng, Dieter Schmalstieg, Gordon Wetzstein, and Denis Kalkofen. Video see-through mixed reality with focus cues. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2256–2266, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Erickson:2020:EDI

[ENK^{+}20] A. Erickson, N. Norouzi, K. Kim, J. J. LaViola, G. Bruder, and G. F. Welch. Effects of depth information on visual target identification task performance in shared gaze environments. IEEE Transactions on Visualization and Computer Graphics, 26(5):1934– 1944, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Elavsky:2024:DNA

[ENM24] Frank Elavsky, Lucas Nadolskis, and Dominik Moritz. Data Navigator: an accessibilitycentered data navigation toolkit. IEEE Transactions on Visualization and Computer Graphics, 30(1):803–813, January 2024. CODEN ITVGEA. ISSN 1077-2626.

ElAli:2023:MHM

[ENvBC23] Abdallah El Ali, Rayna Ney, Zeph M. C. van Berlo, and Pablo Cesar. Is that my heartbeat? Measuring and understanding modality-dependent cardiac interoception in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4805– 4815, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Elliott:2021:DSV

[ENXS21] M. A. Elliott, C. Nothelfer, C. Xiong, and D. A. Szafir. A design space of vision science methods for visualization research. IEEE Transactions on Visualization and Computer Graphics, 27(2):1117– 1127, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Engel:2021:DVA

[ER21] D. Engel and T. Ropinski. Deep volumetric ambient occlusion. IEEE Transactions on Visualization and Computer Graphics, 27(2):1268–1278, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Elshehaly:2021:QAG

[ERB^{+}21] M. Elshehaly, R. Randell, M. Brehmer, L. McVey, N. Alvarado, C. P. Gale, and R. A. Ruddle. QualDash: Adaptable generation of visualisation dashboards for healthcare quality improvement. IEEE Transactions on Visualization and Computer Graphics, 27(2):689– 699, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Elhamdadi:2024:VMF

[ESB^{+}24] Hamza Elhamdadi, Adam Stefkovics, Johanna Beyer, Eric Moerth, Hanspeter Pfister, Cindy Xiong Bearfield, and Carolina Nobre. Vistrust: a multidimensional framework and empirical study of trust in data visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(1): 348–358, January 2024. CO-DEN ITVGEA. ISSN 1077-

Eilemann:2020:ECP

[ESP20] S. Eilemann, D. Steiner, and R. Pajarola. Equalizer 2.0 convergence of a parallel rendering framework. IEEE Transactions on Visualization and Computer Graphics, 26(2):1292– 1307, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shin:2023:ESC

[eSYKW23] Jae eun Shin, Boram Yoon, Dooyoung Kim, and Woontack Woo. The effects of spatial complexity on narrative experience in space-adaptive AR storytelling. IEEE Transactions on Visualization and Computer Graphics, 29(12):5137– 5148, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Fribourg:2020:ASE

[FALH20] R. Fribourg, F. Argelaguet, A. Lécuyer, and L. Hoyet. Avatar and sense of embodiment: Studying the relative preference between appearance, control and point of view. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2062–2072, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Fischer:2021:VAT

[FAS^{+}21] M. T. Fischer, D. Arya, D. Streeb, D. Seebacher, D. A. Keim, and M. Worring. Visual analytics for temporal hypergraph model exploration. IEEE Transactions on Visualization and Computer Graphics, 27(2): 550–560, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Friess:2021:FEL

[FBB^{+}21] F. Frieß, M. Braun, V. Bruder, S. Frey, G. Reina, and T. Ertl. Foveated encoding for large high-resolution displays. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1850–1859, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Feng:2021:LOS

[FBW21] Xianglong Feng, Zeyang Bao, and Sheng Wei. LiveObj: Object semantics-based viewport prediction for live mobile virtual reality streaming. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2736–2745, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Ferrari:2022:EPF

[FCFC22a] Vincenzo Ferrari, Nadia Cattari, Umberto Fontana, and Fabrizio Cutolo. Errata to Parallax free registration for augmented reality optical see-through displays in the peripersonal space [1] (DOI: 10.1109/TVCG.2020.3021534). IEEE Transactions on Visualization and Computer Graphics, 28(8):3069, August 2022. CODEN ITVGEA. ISSN 1077-2626. See [FCFC22b].

Ferrari:2022:PFR

[FCFC22b] Vincenzo Ferrari, Nadia Cattari, Umberto Fontana, and Fabrizio Cutolo. Parallax free registration for augmented reality optical see-through displays in the peripersonal space. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1608–1618, March 2022. CO-DEN ITVGEA. ISSN 1077-2626. See errata [FCFC22a].

Fleck:2023:RTS

[FCH^{+}23] Philipp Fleck, Aimée Sousa Calepso, Sebastian Hubenschmid, Michael Sedlmair, and Dieter Schmalstieg. RagRug: a toolkit for situated analytics. IEEE Transactions on Visualization and Computer Graphics, 29(7):3281–3297, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Fujiwara:2020:IDR

[FCS^{+}20] T. Fujiwara, J. Chou, S. Shilpika, P. Xu, L. Ren, and K. Ma. An incremental dimensionality reduction method for visualizing streaming multidimensional data. IEEE Transactions on Visualization and Computer Graphics, 26(1):418–428, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Feldstein:2021:SVB

[FE21] Ilja T. Feldstein and Stephen R. Ellis. A simple video-based technique for measuring latency in virtual reality or teleoperation. IEEE Transactions on Visualization and Computer Graphics, 27(9):3611– 3625, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Fekete:2021:VTA

[Fek21] J.-D. Fekete. The 2020 Visualization Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 27(2):xxvii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Fygenson:2024:AMI

[FFB24] Racquel Fygenson, Steven Franconeri, and Enrico Bertini. The arrangement of marks impacts afforded messages: Ordering, partitioning, spacing, and coloring in bar charts. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1008–1018, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Fuchs:2022:VVR

[FG22] Henry Fuchs and Federico Gil.

VGTC Virtual Reality Awards Program Chair message. IEEE Transactions on Visualization and Computer Graphics, 28 (5):xii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Faleel:2021:HHP

[FGF^{+}21] Shariff A. M. Faleel, Michael Gammon, Kevin Fan, Da-Yuan Huang, Wei Li, and Pourang Irani. HPUI: Hand proximate user interfaces for one-handed interactions on head mounted displays. IEEE Transactions on Visualization and Computer Graphics, 27(11):4215– 4225, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Fiorentino:2023:MIS

[FGL^{+}23] Michele Fiorentino, Joseph L. Gabbard, Gun Lee, Maud Marchal, and Guillaume Moreau. Message from the ISMAR 2023 Science and Technology Journal Program Chairs and TVCG Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 29(11): vi, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Friston:2021:QSI

[FGS^{+}21] Sebastian Friston, Elias Griffith, David Swapp, Caleb Lrondi, Fred Jjunju, Ryan Ward, Alan Marshall, and Anthony Steed. Quality of service impact on edge physics simulations for VR. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2691–2701, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Friston:2022:CBN

[FGS^{+}22] Sebastian Friston, Elias Griffith, David Swapp, Simon Julier, Caleb Irondi, Fred Jjunju, Ryan Ward, Alan Marshall, and Anthony Steed. Consensus based networking of distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 28(9):3138–3153, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Fennedy:2021:OVU

[FHR^{+}21] Katherine Fennedy, Jeremy Hartmann, Quentin Roy, Simon Tangi Perrault, and Daniel Vogel. OctoPocus in VR: Using a dynamic guide for 3D midair gestures in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(12):4425–4438, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Fang:2023:MDF

[FHTB23] Xianzhong Fang, Jin Huang, Yiying Tong, and Hujun Bao. Metric-driven 3D frame field generation. IEEE Transactions on Visualization and Computer Graphics, 29(4):1964– 1976, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Fan:2021:SBF

[FiMH21] Chaoran Fan, Kre imir Matkovi , and Helwig Hauser. Sketchbased fast and accurate querying of time series using parameter-sharing LSTM networks. IEEE Transactions on Visualization and Computer Graphics, 27(12):4495– 4506, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Furmanova:2020:MVD

[FJK^{+}20] K. Furmanová, A. Jurčík, B. Kozlíková, H. Hauser, and J. Byška. Multiscale visual drilldown for the analysis of large ensembles of multi-body protein complexes. IEEE Transactions on Visualization and Computer Graphics, 26(1):843–852, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Farias:2020:OPM

[FK20] R. Farias and M. Kallmann. Optimal path maps on the GPU. IEEE Transactions on Visualization and Computer Graphics, 26(9):2863– 2874, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Fujiwara:2020:SAD

[FKM20] T. Fujiwara, O. Kwon, and K. Ma. Supporting analysis of dimensionality reduction results with contrastive learning. IEEE Transactions on Visualization and Computer Graphics, 26(1):45–55, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Friedl-Knirsch:2024:SCV

[FKSP^{+}24] Judith Friedl-Knirsch, Christian Stach, Fabian Pointecker, Christoph Anthes, and Daniel Roth. A study on collaborative visual data analysis in augmented reality with asymmetric display types. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2633–2643, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Fan:2023:RWE

[FLS23] Linwei Fan, Huiyu Li, and Miaowen Shi. Redirected walking for exploring immersive virtual spaces with HMD: a comprehensive review and recent advances. IEEE Transactions on Visualization and Computer Graphics, 29(10):4104– 4123, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Feng:2021:CDS

[FLW^{+}21] Xudong Feng, Jiafeng Liu, Huamin Wang, Yin Yang, Hujun Bao, Bernd Bickel, and Weiwei Xu. Computational design of skinned quad-robots. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2881–2895, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Feng:2023:IFS

[FLX^{+}23] Fan Feng, Jinyuan Liu, Shiying Xiong, Shuqi Yang, Yaorui Zhang, and Bo Zhu. Impulse fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 29(6): 3081–3092, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Feng:2021:TDM

[FLZ^{+}21] Z. Feng, H. Li, W. Zeng, S.- H. Yang, and H. Qu. Topology density map for urban data visualization and analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2): 828–838, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Fourrier:2023:HET

[FMBN23] Nicolas Fourrier, Guillaume Moreau, Mustapha Benaouicha, and Jean-Marie Norm. Handwriting for efficient text entry in industrial VR applications: Influence of board orientation and sensory feedback on performance. IEEE Transactions on Visualization and Computer Graphics, 29(11):4438– 4448, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Fernandes:2023:LPF

[FMP23] Ajoy S. Fernandes, T. Scott Murdison, and Michael J. Proulx. Leveling the playing field: a comparative reevaluation of unmodified eye tracking as an input and interaction modality for VR. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2269–2279, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Floricel:2022:THM

[FNB^{+}22] Carla Floricel, Nafiul Nipu, Mikayla Biggs, Andrew Wentzel, Guadalupe Canahuate, Lisanne Van Dijk, Abdallah Mohamed, C. David Fuller, and G. Elisabeta Marai. THALIS: Humanmachine analysis of longitudinal symptoms in cancer therapy. IEEE Transactions on Visualization and Computer Graphics, 28(1):151–161, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Fribourg:2021:VCE

[FOH^{+}21] Rebecca Fribourg, Nami Ogawa, Ludovic Hoyet, Ferran Argelaguet, Takuji Narumi, Michitaka Hirose, and Anatole Lécuyer. Virtual coembodiment: Evaluation of the sense of agency while sharing the control of a virtual body among two individuals. IEEE Transactions on Visualization and Computer Graphics, 27 (10):4023–4038, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Fonnet:2021:SIA

[FP21] A. Fonnet and Y. Prié. Survey of immersive analytics. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2101–2122, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Fantini:2023:CIA

[FPG^{+}23] Davide Fantini, Giorgio Presti, Michele Geronazzo, Riccardo Bona, Alessandro Giuseppe Privitera, and Federico Avanzini. Co-immersion in audio augmented virtuality: The case study of a static and approximated late reverberation algorithm. IEEE Transactions on Visualization and Computer Graphics, 29(11):4472– 4482, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Feyer:2024:ESM

[FPK^{+}24] Stefan P. Feyer, Bruno Pinaud, Stephen Kobourov, Nicolas Brich, Michael Krone, Andreas Kerren, Michael Behrisch, Falk Schreiber, and Karsten Klein. 2D, 2.5D, or 3D? An exploratory study on multilayer network visualisations in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(1):469–479, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Fernandez:2023:EIE

[FRiM^{+}23] Manlio Massiris Fernández, Sanjin Rado , Kre imir Matkovi , M. Eduard Gröller, and Claudio Delrieux. ErgoExplorer: Interactive ergonomic risk assessment from video collections. IEEE Transactions on Visualization and Computer Graphics, 29(1):43–52, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Fischer:2023:IOD

[FRL^{+}23] Marc Fischer, Jarrett Rosenberg, Christoph Leuze, Brian Hargreaves, and Bruce Daniel. The impact of occlusion on depth perception at arm’s length. IEEE Transactions on Visualization and Computer Graphics, 29(11):4494– 4502, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Fu:2024:HAC

[FS24] Yu Fu and John Stasko. Hoop-InSight: Analyzing and comparing basketball shooting performance through visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):858–868, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Filho:2020:EIS

[FSN20] J. A. W. Filho, W. Stuerzlinger, and L. Nedel. Evaluating an immersive space-time cube geovisualization for intuitive trajectory data exploration. IEEE Transactions on Visualization and Computer Graphics, 26(1): 514–524, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fujiwara:2021:VAF

[FSS^{+}21] T. Fujiwara, Shilpika, N. Sakamoto, J. Nonaka, K. Yamamoto, and K.-L. Ma. A visual analytics framework for reviewing multivariate time-series data with dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1601–1611, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Fereydooni:2022:ISA

[FTWP22] Nadia Fereydooni, Einat Tenenboim, Bruce N. Walker, and Srinivas Peeta. Incorporating situation awareness cues in virtual reality for users in dynamic in-vehicle environments. IEEE Transactions on Visualization and Computer Graphics, 28(11):3865–3873, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Feng:2024:NSL

[FV24] Brandon Yushan Feng and Amitabh Varshney. Neural subspaces for light fields. IEEE Transactions on Visualization and Computer Graphics, 30(3): 1685–1695, March 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Fernstad:2022:EWI

[FW22] Sara Johansson Fernstad and Jimmy Johansson Westberg. To explore what isn’t there: Glyph-based visualization for analysis of missing values. IEEE Transactions on Visualization and Computer Graphics, 28 (10):3513–3529, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Floricel:2024:RTU

[FWM^{+}24] Carla Floricel, Andrew Wentzel, Abdallah Mohamed, C. David Fuller, Guadalupe Canahuate, and G. Elisabeta Marai. Roses have thorns: Understanding the downside of oncological care delivery through visual analytics and sequential rule mining. IEEE Transactions on Visualization and Computer Graphics, 30(1):1227–1237, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Fan:2023:ESG

[FWW^{+}23] Jiahui Fan, Beibei Wang, Wenshi Wu, Milo Ha an, Jian Yang, and Ling-Qi Yan. Efficient specular glints rendering with differentiable regularization. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2940–2949, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Feng:2024:PIP

[FWW^{+}24] Yingchaojie Feng, Xingbo Wang, Kam Kwai Wong, Sijia Wang, Yuhong Lu, Minfeng Zhu, Baicheng Wang, and Wei Chen. PromptMagician: Interactive prompt engineering for text-to-image creation. IEEE Transactions on Visualization and Computer Graphics, 30(1): 295–305, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Fan:2020:VIM

[FWZ^{+}20] M. Fan, K. Wu, J. Zhao, Y. Li, W. Wei, and K. N. Truong. VisTA: Integrating machine intelligence with visualization to support the investigation of think-aloud sessions. IEEE Transactions on Visualization and Computer Graphics, 26(1):343–352, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Fujiwara:2022:IDR

[FWZM22] Takanori Fujiwara, Xinhai Wei, Jian Zhao, and Kwan-Liu Ma. Interactive dimensionality reduction for comparative analysis. IEEE Transactions on Visualization and Computer Graphics, 28(1):758–768, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Fang:2020:SBV

[FWZZ20] F. Fang, T. Wang, T. Zeng, and G. Zhang. A superpixelbased variational model for image colorization. IEEE Transactions on Visualization and Computer Graphics, 26(10): 2931–2943, October 2020. CO-DEN ITVGEA. ISSN 1077-2626.

Fidalgo:2023:SRA

[FYC^{+}23] Catarina G. Fidalgo, Yukang Yan, Hyunsung Cho, Maurício Sousa, David Lindlbauer, and Joaquim Jorge. A survey on remote assistance and training in mixed reality environments. IEEE Transactions on Visualization and Computer Graphics, 29(5):2291–2303, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Feng:2022:EVA

[FYE^{+}22] Tinghao Feng, Jing Yang, Martha-Cary Eppes, Zhaocong Yang, and Faye Moser. EVis: Visually analyzing environmentally driven events. IEEE Transactions on Visualization and Computer Graphics, 28(1): 912–921, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Fu:2023:STO

[FYL^{+}23] Yanping Fu, Qingan Yan, Jie Liao, Huajian Zhou, Jin Tang, and Chunxia Xiao. Seamless texture optimization for RGB-D reconstruction. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1845–1859, March 2023. CO-DEN ITVGEA. ISSN 1077-

Fu:2021:CRC

[FZC^{+}21] J. Fu, B. Zhu, W. Cui, S. Ge, Y. Wang, H. Zhang, H. Huang, Y. Tang, D. Zhang, and X. Ma. Chartem: Reviving chart images with data embedding. IEEE Transactions on Visualization and Computer Graphics, 27(2):337–346, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Feng:2022:PGN

[FZZX22] Wanquan Feng, Juyong Zhang, Yuanfeng Zhou, and Shiqing Xin. GDR-Net: a geometric detail recovering network for 3D scanned objects. IEEE Transactions on Visualization and Computer Graphics, 28(12): 3959–3973, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Goncalves:2022:ELC

[GBL^{+}22] Afonso Gonçalves, Adrián Borrego, Jorge Latorre, Roberto Llorens, and Sergi Bermúdez i Badia. Evaluation of a low-cost virtual reality surround-screen projection system. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4452–4461, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ghahremani:2022:PNR

[GBM^{+}22] Parmida Ghahremani, Saeed Boorboor, Pooya Mirhosseini, Chetan Gudisagar, Mala Ananth, David Talmage, Lorna W. Role, and Arie E. Kaufman.

NeuroConstruct: 3D recon-

struction and visualization of neurites in optical microscopy brain images. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4951–4965, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Gedicke:2021:ZME

[GBNH21] S. Gedicke, A. Bonerath, B. Niedermann, and J.-H. Haunert. Zoomless maps: External labeling methods for the interactive exploration of dense point sets at a fixed map scale. IEEE Transactions on Visualization and Computer Graphics, 27(2):1247–1256, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Gattullo:2022:WHW

[GEU^{+}22] Michele Gattullo, Alessandro Evangelista, Antonio E. Uva, Michele Fiorentino, and Joseph L. Gabbard. What, how, and why are visual assets used in industrial augmented reality? A systematic review and classification in maintenance, assembly, and training (from 1997 to 2019). IEEE Transactions on Visualization and Computer Graphics, 28(2): 1443–1456, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Gonzalez-Franco:2022:VVR

[GF22] Mar Gonzalez-Franco. VGTC Virtual Reality Significant New Researcher Award. IEEE Transactions on Visualization and Computer Graphics, 28 (5):xv, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Gomez:2024:PRG

[GF24] Vivian Gómez and Pablo Figueroa. ProtoColVR: Requirements gathering and collaborative rapid prototyping of VR training simulators for multidisciplinary teams. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2549–2558, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Groth:2023:WBF

[GFCM23] Colin Groth, Sascha Fricke, Susana Castillo, and Marcus Magnor. Wavelet-based fast decoding of 360^{∘} videos. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2508–2516, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Gonzalez-Franco:2020:UFA

[GFSHO20] M. Gonzalez-Franco, A. Steed, S. Hoogendyk, and E. Ofek. Using facial animation to increase the enfacement illusion and avatar self-identification. IEEE Transactions on Visualization and Computer Graphics, 26(5):2023–2029, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Gross:2021:ARL

[GG21] D. Groß and S. Gumhold. Advanced rendering of line data with ambient occlusion and transparency. IEEE Transactions on Visualization and Computer Graphics, 27(2):614– 624, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Guo:2022:SVA

[GGJ^{+}22] Yi Guo, Shunan Guo, Zhuochen Jin, Smiti Kaul, David Gotz, and Nan Cao. Survey on visual analysis of event sequence data. IEEE Transactions on Visualization and Computer Graphics, 28(12):5091–5112, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Garzotto:2024:EPA

[GGP^{+}24] Franca Garzotto, Mattia Gianotti, Alberto Patti, Francesca Pentimalli, and Francesco Vona. Empowering persons with autism through cross-reality and conversational agents. IEEE Transactions on Visualization and Computer Graphics, 30(5):2591– 2601, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Guo:2020:BAD

[GGPL20] J. Guo, Y. Guo, J. Pan, and W. Lu. BRDF analysis with directional statistics and its applications. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1476–1489, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guarnera:2020:PVC

[GGT^{+}20] D. Guarnera, G. C. Guarnera, M. Toscani, M. Glencross, B. Li, J. Y. Hardeberg, and K. R. Gegenfurtner. Perceptually validated cross-renderer analytical BRDF parameter remapping. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2258–2272, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Guo:2022:IAD

[GJC^{+}22] Shunan Guo, Zhuochen Jin, Qing Chen, David Gotz, Hongyuan Zha, and Nan Cao. Interpretable anomaly detection in event sequences via sequence matching and visual comparison. IEEE Transactions on Visualization and Computer Graphics, 28(12):4531– 4545, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Gaba:2024:MMU

[GKC^{+}24] Aimen Gaba, Zhanna Kaufman, Jason Cheung, Marie Shvakel, Kyle Wm. Hall, Yuriy Brun, and Cindy Xiong Bearfield. My model is unfair, do people even care? Visual design affects trust and perceived bias in machine learning. IEEE Transactions on Visualization and Computer Graphics, 30(1):327–337, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Guerreiro:2023:DSA

[GKN^{+}23] João Guerreiro, Yujin Kim, Rodrigo Nogueira, SeungA Chung, André Rodrigues, and Uran Oh. The design space of the auditory representation of objects and their behaviours in virtual reality for blind people. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2763–2773, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Guo:2020:TBE

[GL20] H. Guo and D. H. Laidlaw. Topic-based exploration and embedded visualizations for research idea generation. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1592–1607, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Giovannangeli:2024:TED

[GLA^{+}24a] Loann Giovannangeli, Frederic Lalanne, David Auber, Romain Giot, and Romain Bourqui. Toward efficient deep learning for graph drawing (DL4GD). IEEE Transactions on Visualization and Computer Graphics, 30(2): 1516–1532, February 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Gray:2024:SMR

[GLA^{+}24b] Kathryn Gray, Mingwei Li, Reyan Ahmed, Md. Khaledur Rahman, Ariful Azad, Stephen Kobourov, and Katy Börner. A scalable method for readable tree layouts. IEEE Transactions on Visualization and Computer Graphics, 30(2):

1564–1578, February 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Giovannangeli:2024:GVS

[GLGB24] Loann Giovannangeli, Frederic Lalanne, Romain Giot, and Romain Bourqui. Guaranteed visibility in scatterplots with tolerance. IEEE Transactions on Visualization and Computer Graphics, 30(1):792–802, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Genay:2022:BAR

[GLH22] Adélaïde Genay, Anatole Lécuyer, and Martin Hachet. Being an avatar for real: a survey on virtual embodiment in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):5071– 5090, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Gao:2021:SMS

[GLHQ21] Yang Gao, Shuai Li, Aimin Hao, and Hong Qin. Simulating multi-scale, granular materials and their transitions with a hybrid Euler–Lagrange solver. IEEE Transactions on Visualization and Computer Graphics, 27(12):4483–4494, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Ganias:2023:CDG

[GLK^{+}23] Giorgos Ganias, Christos Lougiakis, Akrivi Katifori, Maria Roussou, Yannis Ioannidis, and loannis Panagiotis Ioannidis. Comparing different grasping visualizations for object manipulation in VR using controllers. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2369–2378, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Guan:2022:PFS

[GLL^{+}22] Yanran Guan, Han Liu, Kun Liu, Kangxue Yin, Ruizhen Hu, Oliver van Kaick, Yan Zhang, Ersin Yumer, Nathan Carr, Radomir Mech, and Hao Zhang. FAME: 3D shape generation via functionalityaware model evolution. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1758–1772, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Guo:2024:LPJ

[GLL^{+}24] Yuhan Guo, Yuchu Luo, Keer Lu, Linfang Li, Haizheng Yang, and Xiaoru Yuan. LiberRoad: Probing into the journey of Chinese classics through visual analytics. IEEE Transactions on Visualization and Computer Graphics, 30(1):529–539, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Guo:2021:FSS

[GLX^{+}21] Hanqi Guo, David Lenz, Jiayi Xu, Xin Liang, Wenbin He, Iulian R. Grindeanu, Han-Wei Shen, Tom Peterka, Todd Munson, and Ian Foster. FTK: a simplicial spacetime meshing framework for robust and scalable feature tracking. IEEE Transactions on Visualization and Computer Graphics, 27(8):

3463–3480, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Gao:2021:SDD

[GLY^{+}21] L. Gao, Y.-K. Lai, J. Yang, L.-X. Zhang, S. Xia, and L. Kobbelt. Sparse data driven mesh deformation. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2085–2100, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Gu:2023:REP

[GLY^{+}23] Xiang Gu, Sheng Li, Kangrui Yi, Xiaojuan Yang, Huiling Liu, and Guoping Wang. Role-exchange playing: an exploration of role-playing effects for anti-bullying in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4215–4228, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Ghai:2023:DBC

[GM23] Bhavya Ghai and Klaus Mueller.

D-BIAS: a causality-based human-in-the-loop system for tackling algorithmic bias. IEEE Transactions on Visualization and Computer Graphics, 29(1): 473–482, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Garrison:2021:DIH

[GMS^{+}21] Laura Garrison, Juliane Müller, Stefanie Schreiber, Steffen Oeltze-Jafra, Helwig Hauser, and Stefan Bruckner. Dim-Lift: Interactive hierarchical data exploration through dimensional bundling. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2908–2922, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Gao:2023:ETF

[GMTD23] BoYu Gao, Zijun Mai, Huawei Tu, and Henry Been-Lirn Duh. Effects of transfer functions and body parts on body-centric locomotion in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(8):3670–3684, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Goncalves:2020:IDS

[GMVRB20] G. Gonçalves, M. Melo, J. Vasconcelos-Raposo, and M. Bessa. Impact of different sensory stimuli on presence in credible virtual environments. IEEE Transactions on Visualization and Computer Graphics, 26(11):3231– 3240, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Gao:2023:NGI

[GMX23] Duan Gao, Haoyuan Mu, and Kun Xu. Neural global illumination: Interactive indirect illumination prediction under dynamic area lights. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5325–5341, December 2023. CODEN ITVGEA. ISSN 1077-

Ghosh:2022:VVI

[GNMQ22] Aindrila Ghosh, Mona Nashaat, James Miller, and Shaikh Quader. VisExPreS: a visual interactive toolkit for userdriven evaluations of embeddings. IEEE Transactions on Visualization and Computer Graphics, 28(7):2791– 2807, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Guarese:2024:EAI

[GPR^{+}24] Renan Guarese, Emma Pretty, Aidan Renata, Deb Polson, and Fabio Zambetta. Exploring audio interfaces for vertical guidance in augmented reality via hand-based feedback. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2818–2828, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Graciano:2021:QER

[GRi^{+}21] Alejandro Graciano, Antonio J.

Rueda, Adam Pospí il, Ji í Bittner, and Bedrich Benes. Quad-Stack: an efficient representation and direct rendering of layered datasets. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3733–3744, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Gloumeau:2021:POM

[GSH21] P. C. Gloumeau, W. Stuerzlinger, and J. Han. Pin-NPivot: Object manipulation using pins in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2488–2494, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Gehrmann:2020:VID

[GSK^{+}20] S. Gehrmann, H. Strobelt, R. Krüger, H. Pfister, and A. M. Rush. Visual interaction with deep learning models through collaborative semantic inference. IEEE Transactions on Visualization and Computer Graphics, 26(1):884–894, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Galati:2021:ESP

[GSL21] Alexia Galati, Riley Schoppa, and Aidong Lu. Exploring the SenseMaking process through interactions and fNIRS in immersive visualization. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2714–2724, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Gao:2024:TIV

[GSL^{+}24] Lin Gao, Zekai Shao, Ziqin Luo, Haibo Hu, Cagatay Turkay, and Siming Chen. TransforLearn: Interactive visual tutorial for the transformer model. IEEE Transactions on Visualization and Computer Graphics, 30(1): 891–901, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Gabbard:2022:PCM

[GSM^{+}22] Joseph L. Gabbard, Missie Smith, Coleman Merenda, Gary Burnett, and David R. Large. A perceptual color-matching method for examining color blending in augmented reality head-up display graphics. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2834–2851, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Garcia:2021:CUC

[GSP^{+}21] G. García, J. Silveira, J. Poco, A. Paiva, M. B. Nery, C. T. Silva, S. Adorno, and L. G. Nonato. CrimAnalyzer: Understanding crime patterns in São Paulo. IEEE Transactions on Visualization and Computer Graphics, 27(4):2313– 2328, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Gortler:2020:UAP

[GSS^{+}20] J. Görtler, T. Spinner, D. Streeb, D. Weiskopf, and O. Deussen. Uncertainty-aware principal component analysis. IEEE Transactions on Visualization and Computer Graphics, 26(1): 822–831, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gaba:2023:CCC

[GSS^{+}23] Aimen Gaba, Vidya Setlur, Arjun Srinivasan, Jane Hoffswell, and Cindy Xiong. Comparison conundrum and the chamber of visualizations: an exploration of how language influences visual design. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1211–1221, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Gao:2024:EBH

[GST^{+}24] BoYu Gao, Tong Shao, Huawei Tu, Qizi Ma, Zitao Liu, and Teng Han. Exploring bimanual haptic feedback for spatial search in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2422–2433, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Gunther:2020:HOV

[GT20] T. Günther and H. Theisel. Hyper-objective vortices. IEEE Transactions on Visualization and Computer Graphics, 26(3): 1532–1547, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Guo:2023:SSD

[GTC^{+}23] Zhichun Guo, Jun Tao, Siming Chen, Nitesh V. Chawla, and Chaoli Wang. SD2: Slicing and dicing scholarly data for interactive evaluation of academic performance. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3569–3585, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

George:2020:IBV

[GTH20] C. George, P. Tamunjoh, and H. Hussmann. Invisible boundaries for VR: Auditory and haptic signals as indicators for real world boundaries. IEEE Transactions on Visualization and Computer Graphics, 26(12): 3414–3422, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Groth:2022:OGV

[GTH^{+}22] Colin Groth, Jan-Philipp Tauscher, Nikkel Heesen, Max Hattenbach, Susana Castillo, and Marcus Magnor. Omnidirectional galvanic vestibular stimulation in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2234–2244, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Gilbert:2020:IWB

[GTHC20] A. Gilbert, M. Trumble, A. Hilton, and J. Collomosse. Inpainting of wide-baseline multiple viewpoint video. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2417–2428, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Giovannelli:2023:GVE

[GTL^{+}23] Alexander Giovannelli, Jerald Thomas, Logan Lane, Francielly Rodrigues, and Doug A. Bowman. Gestures vs. emojis: Comparing non-verbal reaction visualizations for immersive collaboration. IEEE Transactions on Visualization and Computer Graphics, 29(11):4772– 4781, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Geronazzo:2020:SHV

[GVN^{+}20] M. Geronazzo, L. S. Vieira, N. C. Nilsson, J. Udesen, and S. Serafin. Superhuman hearing — virtual prototyping of artificial hearing: a case study on interactions and acoustic beamforming. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1912–1922, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Guillou:2024:DMS

[GVT24] Pierre Guillou, Jules Vidal, and Julien Tierny. Discrete Morse sandwich: Fast computation of persistence diagrams for scalar data: An algorithm and a benchmark. IEEE Transactions on Visualization and Computer Graphics, 30(4): 1897–1915, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Gosala:2023:SCM

[GWC^{+}23] Nikhil Gosala, Fangjinhua Wang, Zhaopeng Cui, Hanxue Liang, Oliver Glauser, Shihao Wu, and Olga Sorkine-Hornung. Self-calibrated multisensor wearable for hand tracking and modeling. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1769–1784, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Guo:2024:BIE

[GWD^{+}24] Zixuan Guo, Hongyu Wang, Hanxiao Deng, Wenge Xu, Nilufar Baghaei, Cheng-Hung Lo, and Hai-Ning Liang. Breaking the isolation: Exploring the impact of passthrough in shared spaces on player performance and experience in VR exergames. IEEE Transactions on Visualization and Computer Graphics, 30(5):2580– 2590, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Gorisse:2023:GIV

[GWF^{+}23] Geoffrey Gorisse, Simon Wellenreiter, Sylvain Fleury, Anatole Lécuyer, Simon Richir, and Olivier Christmann. I am a genius! Influence of virtually embodying Leonardo da Vinci on creative performance. IEEE Transactions on Visualization and Computer Graphics, 29(11):4328–4338, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Gao:2022:EEN

[GWL^{+}22] Zihan Gao, Huiqiang Wang, Hongwu Lv, Moshu Wang, and Yifan Qi. Evaluating the effects of non-isomorphic rotation on 3D manipulation tasks in mixed reality simulation. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1261–1273, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Ge:2021:ISS

[GWW^{+}21] Liangsheng Ge, Beibei Wang, Lu Wang, Xiangxu Meng, and Nicolas Holzschuch. Interactive simulation of scattering effects in participating media using a neural network model. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3123–3134, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Gan:2024:VVN

[GXH^{+}24] Wanshui Gan, Hongbin Xu, Yi Huang, Shifeng Chen, and Naoto Yokoya. V4D: Voxel for 4D novel view synthesis. IEEE Transactions on Visualization and Computer Graphics, 30(2): 1579–1591, February 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Guo:2020:RPP

[GXY^{+}20] J. Guo, S. Xu, D. Yan, Z. Cheng, M. Jaeger, and X. Zhang. Realistic procedural plant modeling from multiple view images. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1372–1384, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gupta:2024:CBD

[GZG^{+}24] Kunal Gupta, Yuewei Zhang, Tamil Selvan Gunasekaran, Nanditha Krishna, Yun Suen Pai, and Mark Billinghurst. CAEVR: Biosignals-driven contextaware empathy in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2671– 2681, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Gou:2021:VVA

[GZL^{+}21] L. Gou, L. Zou, N. Li, M. Hofmann, A. K. Shekar, A. Wendt, and L. Ren. VATLD: a visual analytics system to assess, understand and improve traffic light detection. IEEE Transactions on Visualization and Computer Graphics, 27(2):261– 271, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Gao:2021:PNP

[GZM^{+}21] Lin Gao, Ling-Xiao Zhang, Hsien-Yu Meng, Yi-Hui Ren, Yu-Kun Lai, and Leif Kobbelt. PRS-Net: Planar reflective symmetry detection net for 3D models. IEEE Transactions on Visualization and Computer Graphics, 27(6):3007– 3018, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Garcia-Zanabria:2022:CSL

[GZRP^{+}22] Germain García-Zanabria, Marcos M. Raimundo, Jorge Poco, Marcelo Batista Nery, Cláudio T. Silva, Sergio Adorno, and Luis Gustavo Nonato. CriPAV: Street-level crime patterns analysis and visualization. IEEE Transactions on Visualization and Computer Graphics, 28(12):4000–4015, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Gotz:2020:VAH

[GZW^{+}20] D. Gotz, J. Zhang, W. Wang, J. Shrestha, and D. Borland. Visual analysis of highdimensional event sequence data via dynamic hierarchical aggregation. IEEE Transactions on Visualization and Computer Graphics, 26(1):440– 450, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Gong:2022:PPR

[GZY^{+}22] Bingjian Gong, Zunjie Zhu, Chenggang Yan, Zhiguo Shi,

and Feng Xu. PlaneFusion:

Real-time indoor scene reconstruction with planar prior. IEEE Transactions on Visualization and Computer Graphics, 28(12):4671–4684, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Hoque:2020:SVS

[HA20] E. Hoque and M. Agrawala. Searching the visual style and structure of D3 visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(1):1236–1245, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hashemian:2023:LBI

[HAK^{+}23] Abraham M. Hashemian, Ashu Adhikari, Ernst Kruijff, Markus von der Heyde, and Bernhard E. Riecke. Leaning-based interfaces improve groundbased VR locomotion in reachthe-target, follow-the-path, and racing tasks. IEEE Transactions on Visualization and Computer Graphics, 29(3):

1748–1768, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Hong:2020:OCV

[HB20] X. Hong and S. Brooks. 3D objects clouds: Viewing virtual objects in interactive clouds. IEEE Transactions on Visualization and Computer Graphics, 26(3):1442–1453, March 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Holder:2024:PPP

[HB24] Eli Holder and Cindy Xiong Bearfield. Polarizing political polls: How visualization design choices can shape public opinion and increase political polarization. IEEE Transactions on Visualization and Computer Graphics, 30(1):1446– 1456, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Hall:2020:DIT

[HBH^{+}20] K. W. Hall, A. J. Bradley, U. Hinrichs, S. Huron, J. Wood, C. Collins, and S. Carpendale. Design by immersion: a transdisciplinary approach to problem-driven visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(1):109–118, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2021:FFE

[HBLW21] Zhiming Hu, Andreas Bulling, Sheng Li, and Guoping Wang.

FixationNet: Forecasting eye fixations in task-oriented virtual environments. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2681–2690, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Hu:2023:ERU

[HBLW23] Zhiming Hu, Andreas Bulling, Sheng Li, and Guoping Wang. EHTask: Recognizing user tasks from eye and head movements in immersive virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(4):1992– 2004, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Hu:2023:GAP

[HBMK23] Ping Hu, Saeed Boorboor, Joseph Marino, and Arie E. Kaufman. Geometry-aware planar embedding of treelike structures. IEEE Transactions on Visualization and Computer Graphics, 29(7):3182– 3194, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Horak:2021:RMC

[HBS^{+}21] T. Horak, P. Berger, H. Schumann, R. Dachselt, and C. Tominski. Responsive matrix cells: a Focus+Context approach for exploring and editing multivariate graphs. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1644–1654, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Huang:2023:RTG

[HCH^{+}23] Shi-Sheng Huang, Haoxiang Chen, Jiahui Huang, Hongbo Fu, and Shi-Min Hu. Real-time globally consistent 3D reconstruction with semantic priors. IEEE Transactions on Visualization and Computer Graphics, 29(4):1977–1991, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Hong:2021:DIV

[HCJ21] Sumin Hong, Junyoung Choi, and Won-Ki Jeong. Distributed interactive visualization using GPU-optimized Spark. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3670–3684, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Hu:2020:SLI

[HCL20] S. Hu, J. Cai, and Y. Lai. Semantic labeling and instance segmentation of 3D point clouds using patch context analysis and multiscale processing. IEEE Transactions on Visualization and Computer Graphics, 26(7):2485–2498, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Horak:2022:VAH

[HCM^{+}22] Tom Horak, Norine Coenen, Niklas Metzger, Christopher Hahn, Tamara Flemisch, Julián Méndez, Dennis Dimov, Bernd Finkbeiner, and Raimund Dachselt. Visual analysis of hyperproperties for understanding model checking results. IEEE Transactions on Visualization and Computer Graphics, 28(1): 357–367, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Hao:2023:TTY

[HCWK23] Hongtao Hao, Yumian Cui, Zhengxiang Wang, and Yea-Seul Kim. Thirty-two years of IEEE VIS: Authors, fields of study and citations. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1016–1025, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Hu:2021:SDC

[HCX^{+}21] Ruizhen Hu, Bin Chen, Juzhan Xu, Oliver van Kaick, Oliver Deussen, and Hui Huang. Shape-driven coordinate ordering for star glyph sets via reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 27(6): 3034–3047, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Hetzel:2021:CIE

[HDFK21] Lorenz Hetzel, John Dudley, Anna Maria Feit, and Per Ola Kristensson. Complex interaction as emergent behaviour: Simulating mid-air virtual keyboard typing using reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 27(11): 4140–4149, November 2021. CODEN ITVGEA. ISSN 1077-

Hoque:2024:DSF

[HE24] Md Naimul Hoque and Niklas Elmqvist. Dataopsy: Scalable and fluid visual exploration using aggregate query sculpting. IEEE Transactions on Visualization and Computer Graphics, 30(1):186–196, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Heine:2021:TMV

[Hei21] C. Heine. Towards modeling visualization processes as dynamic Bayesian networks. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1000–1010, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Hansen:2021:ARS

[HFS^{+}21] Lasse H. Hansen, Philipp Fleck, Marco Stranner, Dieter Schmalstieg, and Clemens Arth. Augmented reality for subsurface utility engineering, revisited. IEEE Transactions on Visualization and Computer Graphics, 27(11):4119–4128, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Hoshikawa:2024:RDO

[HFT^{+}24] Yukai Hoshikawa, Kazuyuki Fujita, Kazuki Takashima, Morten Fjeld, and Yoshifumi Kitamura. RedirectedDoors+: Door-opening redirection with dynamic haptics in roomscale VR. IEEE Transactions on Visualization and Computer Graphics, 30(5):2276– 2286, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Hirsch:2022:TVE

[HGB22] Linda Hirsch, Ceenu George, and Andreas Butz. Traces in virtual environments: a framework and exploration to conceptualize the design of social virtual environments. IEEE Transactions on Visualization and Computer Graphics, 28(11):3874–3884, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Henz:2021:SCN

[HGO21] B. Henz, E. S. L. Gastal, and M. M. Oliveira. Synthesizing camera noise using generative adversarial networks. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2123–2135, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

He:2020:EEF

[HGSP20] W. He, H. Guo, H. Shen, and T. Peterka. eFESTA: Ensemble feature exploration with surface density estimates. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1716–1731, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hota:2020:EMI

[HH20] A. Hota and J. Huang. Embedding meta information into visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(11): 3189–3203, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Hu:2023:HCS

[HHB^{+}23] Kaidong Hu, Brandon Haworth, Glen Berseth, Vladimir Pavlovic, Petros Faloutsos, and Mubbasir Kapadia. Heterogeneous crowd simulation using parametric reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 29(4):2036– 2052, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Helske:2021:CVA

[HHC^{+}21] Jouni Helske, Satu Helske, Matthew Cooper, Anders Ynnerman, and Lonni Besançon. Can visualization alleviate dichotomous thinking? Effects of visual representations on the cliff effect. IEEE Transactions on Visualization and Computer Graphics, 27(8):3397–3409, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Han:2020:CPP

[HHD^{+}20] X. Han, K. Hou, D. Du, Y. Qiu, S. Cui, K. Zhou, and Y. Yu. CaricatureShop: Personalized and photorealistic caricature sketching. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2349–2361, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Hiroi:2024:SCV

[HHI24] Yuichi Hiroi, Takefumi Hiraki, and Yuta Itoh. StainedSweeper: Compact, variable-intensity light-attenuation display with sweeping tunable retarders. IEEE Transactions on Visualization and Computer Graphics, 30(5):2682–2692, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Herzberger:2024:ROH

[HHK^{+}24] Lukas Herzberger, Markus Hadwiger, Robert Krüger, Peter Sorger, Hanspeter Pfister, Eduard Gröller, and Johanna Beyer. Residency octree: a hybrid approach for scalable Webbased multi-volume rendering. IEEE Transactions on Visualization and Computer Graphics, 30(1):1380–1390, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Hossain:2021:MPS

[HHKN21] M. I. Hossain, V. Huroyan, S. Kobourov, and R. Navarrete. Multi-perspective, simultaneous embedding. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1569–1579, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Huang:2024:SMV

[HHM^{+}24] Zeyuan Huang, Qiang He, Kevin Maher, Xiaoming Deng, Yu-Kun Lai, Cuixia Ma, Sheng-Feng Qin, Yong-Jin Liu, and Hongan Wang. SpeechMirror: a multimodal visual analytics system for personalized reflection of online public speaking effectiveness. IEEE Transactions on Visualization and Computer Graphics, 30(1):606–616, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Hoque:2023:VCP

[HHS^{+}23] Md Naimul Hoque, Wenbin He, Arvind Kumar Shekar, Liang Gou, and Liu Ren. Visual concept programming: a visual analytics approach to injecting human intelligence at scale. IEEE Transactions on Visualization and Computer Graphics, 29(1): 74–83, January 2023. CODEN ITVGEA. ISSN 1077-2626.

He:2023:BVS

[HIDI23] Tingying He, Petra Isenberg, Raimund Dachselt, and Tobias Isenberg. BeauVis: a validated scale for measuring the aesthetic pleasure of visual representations. IEEE Transactions on Visualization and Computer Graphics, 29(1):363–373, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Hiratani:2023:SPM

[HIK^{+}23] Kosuke Hiratani, Daisuke Iwai, Yuta Kageyama, Parinya Punpongsanon, Takefumi Hiraki, and Kosuke Sato. Shadowless projection mapping using retrotransmissive optics. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2280–2290, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Halladjian:2022:MUI

[HiM^{+}22] Sarkis Halladjian, David Kou il, Haichao Miao, M. Eduard Gröller, Ivan Viola, and Tobias Isenberg. Multiscale unfolding: Illustratively visualizing the whole genome at a glance. IEEE Transactions on Visualization and Computer Graphics, 28(10):3456–3470, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Han:2023:SAS

[HJL^{+}23] Chang Han, Jaemin Jo, Anyi Li, Bongshin Lee, Oliver Deussen, and Yunhai Wang. SizePairs: Achieving stable and balanced temporal treemaps using hierarchical size-based pairing. IEEE Transactions on Visualization and Computer Graphics, 29(1):193–202, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Huang:2021:IVS

[HJZ^{+}21] X. Huang, S. Jamonnak, Y. Zhao, B. Wang, M. Hoai, K. Yager, and W. Xu. Interactive visual study of multiple attributes learning model of Xray scattering images. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1312–1321, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Hall:2022:PDC

[HKB^{+}22] Kyle Wm. Hall, Anthony Kouroupis, Anastasia Bezerianos, Danielle Albers Szafir, and Christopher Collins. Professional differences: a comparative study of visualization task performance and spatial ability across disciplines. IEEE Transactions on Visualization and Computer Graphics, 28(1):654– 664, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Heimerl:2022:EVI

[HKMG22] Florian Heimerl, Christoph Kralj, Torsten Möller, and Michael Gleicher. embComp: Visual interactive comparison of vector embeddings. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2953–2969, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Hu:2020:MIS

[HKVZ20] S.-M. Hu, D. Kalkofen, J. Ventura, and S. Zollmann. Message from the ISMAR 2020 Science and Technology Program Chairs and TVCG Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 26(12):3387– 3388, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Hagele:2023:UAM

[HKW23] David Hägele, Tim Krake, and Daniel Weiskopf. Uncertaintyaware multidimensional scaling. IEEE Transactions on Visualization and Computer Graphics, 29(1):23–32, January 2023. CODEN ITVGEA. ISSN 1077-

Huang:2021:DSR

[HL21] Bingyao Huang and Haibin Ling. DeProCams: Simultaneous relighting, compensation and shape reconstruction for projector-camera systems. IEEE Transactions on Visualization and Computer Graphics, 27(5):2725–2735, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Hashemian:2022:PHI

[HLA^{+}22] Abraham M. Hashemian, Matin Lotfaliei, Ashu Adhikari, Ernst Kruijff, and Bernhard E.

                                        Riecke. HeadJoystick: Im-

proving flying in VR using a novel leaning-based interface. IEEE Transactions on Visualization and Computer Graphics, 28(4):1792–1809, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Huang:2022:PDB

[HLCY22] Renpei Huang, Quan Li, Li Chen, and Xiaoru Yuan. A probability density-based visual analytics approach to forecast bias calibration. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1732–1744, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Huang:2023:IBO

[HLH^{+}23] Fei Huang, Chen Liu, Kai-Wen Hsiao, Ying-Miao Kuo, Hung-Kuo Chu, and Yong-Liang Yang. Image-based OA-style paper pop-up design via mixedinteger programming. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4269–4283, October 2023. CODEN ITVGEA. ISSN 1077-2626.

He:2022:CCG

[HLJ^{+}22] Yu He, Ying-Tian Liu, Yi-Han Jin, Song-Hai Zhang, Yu-Kun Lai, and Shi-Min Hu. Context-consistent generation of indoor virtual environments based on geometry constraints. IEEE Transactions on Visualization and Computer Graphics, 28(12):3986–3999, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Heagerty:2024:HAV

[HLL^{+}24] Jonathan Heagerty, Sida Li, Eric Lee, Shuvra Bhattacharyya, Sujal Bista, Barbara Brawn, Brandon Y. Feng, Susmija Jabbireddy, Joseph JaJa, Hernisa Kacorri, David Li, Derek Yarnell, Matthias Zwicker, and Amitabh Varshney. HoloCamera: Advanced volumetric capture for cinematic-quality VR applications. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2767–2775, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Huang:2020:ESC

[HLM^{+}20] Z. Huang, Y. Lu, E. A. Mack, W. Chen, and R. Maciejewski. Exploring the sensitivity of choropleths under attribute uncertainty. IEEE Transactions on Visualization and Computer Graphics, 26(8):2576–2590, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Hazarika:2020:NNN

[HLW^{+}20] S. Hazarika, H. Li, K. Wang, H. Shen, and C. Chou. NNVA: Neural network assisted visual analysis of yeast cell polarization simulation. IEEE Transactions on Visualization and Computer Graphics, 26(1):34– 44, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2020:DCB

[HLZ^{+}20] Z. Hu, S. Li, C. Zhang, K. Yi, G. Wang, and D. Manocha. DGaze: CNN-based gaze prediction in dynamic scenes. IEEE Transactions on Visualization and Computer Graphics, 26(5):1902–1911, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Hoque:2022:OEC

[HM22] Md Naimul Hoque and Klaus Mueller. Outcome-explorer: a causality guided interactive visual interface for interpretable algorithmic decision making. IEEE Transactions on Visualization and Computer Graphics, 28(12):4728–4740, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Heer:2024:MAS

[HM24] Jeffrey Heer and Dominik Moritz. Mosaic: an architecture for scalable & interoperable data views. IEEE Transactions on Visualization and Computer Graphics, 30(1):436– 446, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Ha:2023:UCU

[HMGO23] Sunwoo Ha, Shayan Monadjemi, Roman Garnett, and Alvitta Ottley. A unified comparison of user modeling techniques for predicting data interaction and detecting exploration bias. IEEE Transactions on Visualization and Computer Graphics, 29(1):483–492, January 2023. CODEN ITVGEA. ISSN 1077-2626.

He:2021:CQI

[HMH^{+}21] Chen He, Luana Micallef, Liye He, Gopal Peddinti, Tero Aittokallio, and Giulio Jacucci. Characterizing the quality of insight by interactions: a case study. IEEE Transactions on Visualization and Computer Graphics, 27(8):3410–3424, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Halladjian:2020:STI

[HMK^{+}20] S. Halladjian, H. Miao, D. Kouřil, M. E. Gröller, I. Viola, and T. Isenberg. Scale Trotter: Illustrative visual travels across negative scales. IEEE Transactions on Visualization and Computer Graphics, 26(1):654– 664, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Huang:2023:CIE

[HMKB23] Jinbin Huang, Aditi Mishra, Bum Chul Kwon, and Chris Bryan. ConceptExplainer: Interactive explanation for deep neural networks from a concept perspective. IEEE Transactions on Visualization and Computer Graphics, 29(1):831–841, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Hong:2024:VCM

[HMTI24] Jiayi Hong, Ross Maciejewski, Alain Trubuil, and Tobias Isenberg. Visualizing and comparing machine learning predictions to improve human–AI teaming on the example of cell lineage. IEEE Transactions on Visualization and Computer Graphics, 30(4):1956– 1969, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Huynh:2021:DNF

[HNGC21] E. Huynh, A. Nyhout, P. Ganea, and F. Chevalier. Designing narrative-focused role-playing games for visualization literacy in young children. IEEE Transactions on Visualization and Computer Graphics, 27(2):924– 934, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Hergl:2023:VHO

[HNS23] Chiara Hergl, Thomas Nagel, and Gerik Scheuermann. Visualizing higher-order 3D tensors by multipole lines. IEEE Transactions on Visualization and Computer Graphics, 29(7):

3405–3418, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Hulstein:2023:GSI

[HPAB23] Golina Hulstein, Vanessa Peña-Araya, and Anastasia Bezerianos. Geo-Storylines: Integrating maps into storyline visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):994–1004, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Hasanzadeh:2020:PMR

[HPdlG20] S. Hasanzadeh, N. F. Polys, and J. M. de la Garza. Presence, mixed reality, and risktaking behavior: a study in safety interventions. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2115–2125, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Han:2022:PHV

[HPK^{+}22] Dongyun Han, Gorakh Parsad, Hwiyeon Kim, Jaekyom Shim, Oh-Sang Kwon, Kyung A. Son, Jooyoung Lee, Isaac Cho, and Sungahn Ko. HisVA: a visual analytics system for studying history. IEEE Transactions on Visualization and Computer Graphics, 28(12):4344– 4359, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Hull:2024:VAG

[HPM^{+}24] Matthew Hull, Vivian Pednekar, Hannah Murray, Nimisha Roy, Emmanuel Tung, Susanta Routray, Connor Guerin, Justin Chen, Zijie J. Wang, Seongmin Lee, Mahdi Roozbahani, and Duen Horng Chau. VG: Automatic grading of D3 visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(1):617–627, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Hearst:2020:ESG

[HPP^{+}20] M. A. Hearst, E. Pedersen, L. Patil, E. Lee, P. Laskowski, and S. Franconeri. An evaluation of semantically grouped word cloud designs. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2748–2761, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Hohman:2020:SSD

[HPRC20] Fred Hohman, Haekyu Park, Caleb Robinson, and Duen Horng Polo Chau. Summit: Scaling deep learning interpretability by visualizing activation and attribution summarizations. IEEE Transactions on Visualization and Computer Graphics, 26(1):1096– 1106, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Herveau:2023:AAS

[HPT^{+}23] Killian Herveau, Philip Pfaffe, Martin Tillmann, Walter F. Tichy, and Carsten Dachsbacher. Analysis of acceleration structure parameters and hybrid autotuning for ray tracing.

IEEE Transactions on Visualization and Computer Graphics, 29(2):1345–1356, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Hinterreiter:2022:CMA

[HRS^{+}22] Andreas Hinterreiter, Peter Ruch, Holger Stitz, Martin Ennemoser, Jürgen Bernard, Hendrik Strobelt, and Marc Streit. ConfusionFlow: a model-agnostic visualization for temporal analysis of classifier confusion. IEEE Transactions on Visualization and Computer Graphics, 28(2):1222– 1236, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Huang:2023:VRT

[HRX23] Xincheng Huang, James Riddell, and Robert Xiao. Virtual reality telepresence: 360-degree video streaming with edgecompute assisted static foveated compression. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4525–4534, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Huang:2023:LGD

[HS23] Danny Huang and Ian Stavness.

Large growth deformations of thin tissue using solid-shells. IEEE Transactions on Visualization and Computer Graphics, 29(3):1893–1909, March 2023. CODEN ITVGEA. ISSN 1077-

Hoang:2021:EFH

[HSB^{+}21] D. Hoang, B. Summa, H. Bhatia, P. Lindstrom, P. Klacansky, W. Usher, P.-T. Bremer, and V. Pascucci. Efficient and flexible hierarchical data layouts for a unified encoding of scalar field precision and resolution. IEEE Transactions on Visualization and Computer Graphics, 27(2): 603–613, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Huang:2022:RDM

[HSC^{+}22] Jiawei Huang, Ryo Sugawara, Kinfung Chu, Taku Komura, and Yoshifumi Kitamura. Reconstruction of dexterous 3D motion data from a flexible magnetic sensor with deep learning and structure-aware filtering. IEEE Transactions on Visualization and Computer Graphics, 28(6):2400– 2414, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Hou:2020:PVG

[HSF^{+}20] F. Hou, Q. Sun, Z. Fang, Y. Liu, S. Hu, H. Qin, A. Hao, and Y. He. Poisson Vector Graphics (PVG). IEEE Transactions on Visualization and Computer Graphics, 26(2): 1361–1371, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Heinrich:2020:CAR

[HSJ^{+}20] F. Heinrich, L. Schwenderling, F. Joeres, K. Lawonn, and C. Hansen. Comparison of augmented reality display techniques to support medical needle insertion. IEEE Transactions on Visualization and Computer Graphics, 26(12): 3568–3575, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Huang:2020:EEF

[HSS^{+}20] Y. Huang, L. Shi, Y. Su, Y. Hu, H. Tong, C. Wang, T. Yang, D. Wang, and S. Liang. Eiffel: Evolutionary flow map for influence graph visualization. IEEE Transactions on Visualization and Computer Graphics, 26(10):2944–2960, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Hu:2020:DSM

[HSV^{+}20] R. Hu, T. Sha, O. Van Kaick, O. Deussen, and H. Huang. Data sampling in multi-view and multi-class scatterplots via set cover optimization. IEEE Transactions on Visualization and Computer Graphics, 26(1): 739–748, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hao:2024:TDT

[HSYZ24] Jianing Hao, Qing Shi, Yilin Ye, and Wei Zeng. Time-Tuner: Diagnosing time representations for time-series forecasting with counterfactual explanations. IEEE Transactions on Visualization and Computer Graphics, 30(1):1183– 1193, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Henkin:2022:WEC

[HT22] Rafael Henkin and Cagatay Turkay. Words of estimative correlation: Studying verbalizations of scatterplots. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1967–1981, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Hamalainen:2022:VMC

[HTBL22] Perttu Hämäläinen, Juuso Toikka, Amin Babadi, and C. Karen Liu. Visualizing movement control optimization landscapes. IEEE Transactions on Visualization and Computer Graphics, 28(3):1648– 1660, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Han:2020:VQG

[HTJ^{+}20] Q. Han, D. Thom, M. John, S. Koch, F. Heimerl, and T. Ertl. Visual quality guidance for document exploration with Focus+Context techniques. IEEE Transactions on Visualization and Computer Graphics, 26(8):2715–2731, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Haley:2023:IVT

[HTP^{+}23] Alexander C. Haley, Don Thorpe, Alex Pelletier, Svetlana Yarosh, and Daniel F. Keefe. Inward VR: Toward a qualitative method for investigating interoceptive awareness in VR. IEEE Transactions on Visualization and Computer Graphics, 29(5):2557– 2566, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Han:2020:FDL

[HTW20] J. Han, J. Tao, and C. Wang.

FlowNet: a deep learning framework for clustering and selection of streamlines and stream surfaces. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1732–1744, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

He:2023:PVN

[HTY^{+}23] Xiangyang He, Yubo Tao, Shuoliu Yang, Haoran Dai, and Hai Lin. voxel2vec: a natural language processing approach to learning distributed representations for scientific data. IEEE Transactions on Visualization and Computer Graphics, 29(10):4296–4311, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Hullman:2020:WAD

[Hul20] J. Hullman. Why authors don’t visualize uncertainty. IEEE Transactions on Visualization and Computer Graphics, 26(1): 130–139, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Han:2020:TTT

[HW20] J. Han and C. Wang. TSR-TVD: Temporal super-resolution for time-varying data analysis and visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1): 205–215, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Han:2022:STS

[HW22] Jun Han and Chaoli Wang. SSR-TVD: Spatial super-resolution for time-varying data analysis and visualization. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2445–2456, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Han:2023:CDG

[HW23] Jun Han and Chaoli Wang. CoordNet: Data generation and visualization generation for time-varying volumes via a coordinate-based neural network. IEEE Transactions on Visualization and Computer Graphics, 29(12):4951– 4963, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Han:2023:GLB

[HWC23] Ping-Hsuan Han, Tzu-Hua Wang, and Chien-Hsing Chou. GroundFlow: Liquid-based haptics for simulating fluid on the ground in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5):2670–2679, May 2023. CODEN ITVGEA. ISSN 1077-2626.

He:2020:IDI

[HWG^{+}20] W. He, J. Wang, H. Guo, K. Wang, H. Shen, M. Raj, Y. S. G. Nashed, and T. Peterka. InSituNet: Deep image synthesis for parameter space exploration of ensemble simulations. IEEE Transactions on Visualization and Computer Graphics, 26(1):23–33, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Hu:2022:ERO

[HWL^{+}22] Jiangbei Hu, Shengfa Wang, Baojun Li, Fengqi Li, Zhongxuan Luo, and Ligang Liu. Efficient representation and optimization for TPMS-based porous structures. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2615–2627, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Hiroi:2023:LLB

[HWMI23] Yuichi Hiroi, Akira Watanabe, Yuri Mikawa, and Yuta Itoh. Low-latency beaming display: Implementation of wearable, 133s motion-to-photon latency near-eye display. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4761–4771, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Hergl:2023:ECE

[HWN^{+}23] Chiara Hergl, Carina Witt, Baldwin Nsonga, Andreas Menzel, and Gerik Scheuermann.

Electromechanical coupling in electroactive polymers a visual analysis of a third-order tensor field. IEEE Transactions on Visualization and Computer Graphics, 29(12):5357– 5371, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Hong:2022:WAI

[HWS22] Matt-Heun Hong, Jessica K. Witt, and Danielle Albers Szafir. The weighted average illusion: Biases in perceived mean position in scatterplots. IEEE Transactions on Visualization and Computer Graphics, 28(1):987–997, January 2022. CODEN ITVGEA. ISSN 1077-2626.

He:2024:EGA

[HWT^{+}24] Helen Ai He, Jagoda Walny, Sonja Thoma, Sheelagh Carpendale, and Wesley Willett. Enthusiastic and grounded, avoidant and cautious: Understanding public receptivity to data and visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1435–1445, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Hou:2022:VEA

[HWW^{+}22] Yijie Hou, Chengshun Wang, Junhong Wang, Xiangyang Xue, Xiaolong Luke Zhang, Jun Zhu, Dongliang Wang, and Siming Chen. Visual evaluation for autonomous driving. IEEE Transactions on Visualization and Computer Graphics, 28(1):

1030–1039, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

He:2024:VVA

[HWW^{+}24] Jianben He, Xingbo Wang, Kam Kwai Wong, Xijie Huang, Changjian Chen, Zixin Chen, Fengjie Wang, Min Zhu, and Huamin Qu. VideoPro: a visual analytics approach for interactive video programming. IEEE Transactions on Visualization and Computer Graphics, 30(1): 87–97, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Holder:2023:DVD

[HX23] Eli Holder and Cindy Xiong.

Dispersion vs disparity: Hiding variability can encourage stereotyping when visualizing social outcomes. IEEE Transactions on Visualization and Computer Graphics, 29(1):624– 634, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Huang:2023:FAF

[HXHT23] Jieying Huang, Yang Xi, Junnan Hu, and Jun Tao. FlowNL: Asking the flow data in natural languages. IEEE Transactions on Visualization and Computer Graphics, 29(1):1200– 1210, January 2023. CODEN ITVGEA. ISSN 1077-2626.

He:2024:LML

[HXL^{+}24] Hao He, Xinhao Xu, Shangman Li, Fang Wang, Isaac Schroeder, Eric M. Aldrich, Scottie D. Murrell, Lanxin Xue, and Yuanyuan Gu. Learning middle-latitude cyclone formation up in the air: Student learning experience, outcomes, and perceptions in a CAVEenabled meteorology class. IEEE Transactions on Visualization and Computer Graphics, 30(5):2807–2817, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Hu:2023:SSC

[HYC^{+}23] Ruizhen Hu, Ziqi Ye, Bin Chen, Oliver van Kaick, and Hui Huang. Self-supervised colorconcept association via image colorization. IEEE Transactions on Visualization and Computer Graphics, 29(1):247– 256, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Hung:2020:MQC

[HYF^{+}20] S. Hung, C. Yao, Y. Fang, P. Tan, R. Lee, A. Sheffer, and H. Chu. Micrography QR codes. IEEE Transactions on Visualization and Computer Graphics, 26(9):2834–2847, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Han:2023:EBP

[HYH^{+}23] Fangzhou Han, Shuquan Ye, Mingming He, Menglei Chai, and Jing Liao. Exemplar-based 3D portrait stylization. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1371–1383, February 2023. CO-DEN ITVGEA. ISSN 1077-

Hu:2023:MCG

[HYSL23] Bo-Yi Hu, Chunyang Ye, Jian-Ping Su, and Ligang Liu. Manifold-constrained geometric optimization via local parameterizations. IEEE Transactions on Visualization and Computer Graphics, 29(2):1318– 1329, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Huang:2020:NLB

[HZC^{+}20] Z. Huang, Y. Zhao, W. Chen, S. Gao, K. Yu, W. Xu, M. Tang, M. Zhu, and M. Xu. A natural-language-based visual query approach of uncertain human trajectories. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1256–1266, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Han:2022:SEE

[HZCW22] Jun Han, Hao Zheng, Danny Z.

Chen, and Chaoli Wang. STNet: an end-to-end generative framework for synthesizing spatiotemporal super-resolution volumes. IEEE Transactions on Visualization and Computer Graphics, 28(1):270–280, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Hamidian:2020:SRE

[HZFH20] H. Hamidian, Z. Zhong, F. Fotouhi, and J. Hua. Surface registration with eigenvalues and eigenvectors. IEEE Transactions on Visualization and Computer Graphics, 26(11): 3327–3339, November 2020. CODEN ITVGEA. ISSN 1077-2626.

He:2024:DCB

[HZII24] Tingying He, Yuanyang Zhong, Petra Isenberg, and Tobias Isenberg. Design characterization for black-and-white textures in visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1019–1029, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Huang:2024:CVA

[HZJ^{+}24] Yansong Huang, Zherui Zhang, Ao Jiao, Yuxin Ma, and Ran Cheng. A comparative visual analytics framework for evaluating evolutionary processes in multi-objective optimization. IEEE Transactions on Visualization and Computer Graphics, 30(1):661–671, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Hnatyshyn:2024:MPV

[HZP^{+}24] Rostyslav Hnatyshyn, Jieqiong Zhao, Danny Perez, James Ahrens, and Ross Maciejewski. MolSieve: a progressive visual analytics system for molecular dynamics simulations. IEEE Transactions on Visualization and Computer Graphics, 30(1): 727–737, January 2024. CO-DEN ITVGEA. ISSN 1077-

Huang:2022:PWW

[HZQ22] Hong Huang, Fan Zhong, and Xueying Qin. Pixel-wise weighted region-based 3D object tracking using contour constraints. IEEE Transactions on Visualization and Computer Graphics, 28(12):4319– 4331, December 2022. CODEN ITVGEA. ISSN 1077-2626.

He:2022:WCW

[HZS^{+}22] Wenbin He, Lincan Zou, Arvind Kumar Shekar, Liang Gou, and Liu Ren. Where can we help? A visual analytics approach to diagnosing and improving semantic segmentation of movable objects. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1040–1050, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Han:2021:VDL

[HZX^{+}21] J. Han, H. Zheng, Y. Xing, D. Z. Chen, and C. Wang. V2V: a deep learning approach to variable-to-variable selection and translation for multivariate time-varying data. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1290–1300, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Hung:2022:FCS

[HZYZ22] Shih-Hsuan Hung, Yue Zhang, Harry Yeh, and Eugene Zhang. Feature curves and surfaces of 3D asymmetric tensor fields.

IEEE Transactions on Visualization and Computer Graphics, 28(1):33–42, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Han:2020:LSP

[HZZ^{+}20] L. Han, T. Zheng, Y. Zhu, L. Xu, and L. Fang. Live semantic 3D perception for immersive augmented reality. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2012–2022, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Hung:2024:GTS

[HZZ24] Shih-Hsuan Hung, Yue Zhang, and Eugene Zhang. Global topology of 3D symmetric tensor fields. IEEE Transactions on Visualization and Computer Graphics, 30(1):1282– 1291, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Ishihara:2023:IBP

[IAI^{+}23] Atsushi Ishihara, Hiroyuki Aga, Yasuko Ishihara, Hirotake Ichikawa, Hidetaka Kaji, Koichi Kawasaki, Daita Kobayashi, Toshimi Kobayashi, Ken Nishida, Takumi Hamasaki, Hideto Mori, and Yuki Morikubo. Integrating both parallax and latency compensation into video see-through head-mounted display. IEEE Transactions on Visualization and Computer Graphics, 29(5):2826– 2836, May 2023. CODEN ITVGEA. ISSN 1077-2626.

il:2023:MAN

[ieSM^{+}23] David Kou il, Ond ej Strnad, Peter Mindek, Sarkis Halladjian, Tobias Isenberg, M. Eduard Gröller, and Ivan Viola. Molecumentary: Adaptable narrated documentaries using molecular visualization. IEEE Transactions on Visualization and Computer Graphics, 29(3):1733–1747, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Ibrahim:2024:RTS

[IGM24] Muhammad Twaha Ibrahim, M. Gopi, and Aditi Majumder. Real-time seamless multi-projector displays on deformable surfaces. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2527–2537, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Iglesias-Guitian:2022:RTD

[IGMM22] Jose A. Iglesias-Guitian, Prajita Mane, and Bochang Moon. Real-time denoising of volumetric path tracing for direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2734–2747, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Iwai:2022:MIS

[IGMW22] Daisuke Iwai, Joseph L. Gabbard, Guillaume Moreau, and Lili Wang. Message from the ISMAR 2022 Science and Technology Journal Program Chairs and TVCG Guest Editors.

IEEE Transactions on Visualization and Computer Graphics, 28(11):v, November 2022. CODEN ITVGEA. ISSN 1077-2626.

il:2021:HBD

[iIK^{+}21] David Kou il, Tobias Isenberg, Barbora Kozlíková, Miriah Meyer, M. Eduard Gröller, and Ivan Viola. HyperLabels: Browsing of dense and hierarchical molecular 3D models. IEEE Transactions on Visualization and Computer Graphics, 27(8):3493–3504, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Itoh:2021:BD

[IKi21] Yuta Itoh, Takumi Kaminokado, and Kaan Ak it. Beaming displays. IEEE Transactions on Visualization and Computer Graphics, 27(5):2659– 2668, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Kim:2023:VHF

[iKYOW23] Hyung il Kim, Boram Yoon, Seo Young Oh, and Woontack Woo. Visualizing hand force with wearable muscle sensing for enhanced mixed reality remote collaboration. IEEE Transactions on Visualization and Computer Graphics, 29(11):4611–4621, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Itoh:2021:CPM

[ILZ^{+}21] Y. Itoh, T. Langlotz, S. Zollmann, D. Iwai, K. Kiyoshi, and T. Amano. Computational phase-modulated eyeglasses. IEEE Transactions on Visualization and Computer Graphics, 27(3):1916– 1928, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Iwai:2021:MIS

[IMKP21] Daisuke Iwai, Guillaume Moreau, Denis Kalkofen, and Tabitha Peck. Message from the ISMAR 2021 Science and Technology Journal Program Chairs and TVCG Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 27 (11):4086, November 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Ivson:2020:SRV

[IMQ^{+}20] P. Ivson, A. Moreira, F. Queiroz, W. Santos, and W. Celes. A systematic review of visualization in building information modeling. IEEE Transactions on Visualization and Computer Graphics, 26(10):3109– 3127, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Itoh:2024:IVM

[IPPZ24] Yuta Itoh, Voicu Popescu, Tabitha Peck, and Stefanie Zollmann. IEEE VR 2024 message from the Program Chairs and Guest Editors. IEEE Transactions on Visualization and Computer Graphics, 30(5):ix–x, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Ibrahim:2022:POC

[IRR^{+}22] Mohamed Ibrahim, Peter Rautek, Guido Reina, Marco Agus, and Markus Hadwiger. Probabilistic occlusion culling using confidence maps for high-quality rendering of large particle data. IEEE Transactions on Visualization and Computer Graphics, 28(1):573–582, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Isenberg:2022:DYB

[ISBP22] Tobias Isenberg, Zujany Salazar, Rafael Blanco, and Catherine Plaisant. Do you believe your (social media) data? A personal story on location data biases, errors, and plausibility as well as their visualization. IEEE Transactions on Visualization and Computer Graphics, 28(9):3277–3291, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Ishikawa:2023:MLS

[ISKM23] Reina Ishikawa, Hideo Saito, Denis Kalkofen, and Shohei Mori. Multi-layer scene representation from composed focal stacks. IEEE Transactions on Visualization and Computer Graphics, 29(11):4719– 4729, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Iuricich:2022:PCV

[Iur22]  Federico Iuricich.  Persis-

tence cycles for visual exploration of persistent homology. IEEE Transactions on Visualization and Computer Graphics, 28(12):4966–4979, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ibayashi:2020:SLD

[IWT^{+}20] H. Ibayashi, C. Wojtan, N. Thuerey, T. Igarashi, and R. Ando. Simulating liquids on dynamically warping grids. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2288–2302, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Jeong:2022:RTD

[JBS^{+}22] Yuna Jeong, Seung Youp Baek, Yechan Seok, Gi Beom Lee, and Sungkil Lee. Real-time dynamic bokeh rendering with efficient look-up table sampling. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1373–1384, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Jiang:2024:HVA

[JCZ^{+}24] Zhihan Jiang, Handi Chen, Rui Zhou, Jing Deng, Xinchen Zhang, Running Zhao, Cong Xie, Yifang Wang, and Edith C. H. Ngai. HealthPrism: a visual analytics system for exploring children’s physical and mental health profiles with multimodal data. IEEE Transactions on Visualization and Computer Graphics, 30(1):1205– 1215, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Jin:2024:RIT

[JdJTC24] Yuanzhe Jin, Tim J. A. de Jong, Martijn Tennekes, and Min Chen. Radial icicle tree (RIT): Node separation and area constancy. IEEE Transactions on Visualization and Computer Graphics, 30(1):251–261, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Jadhav:2022:VP

[JDM^{+}22] Shreeraj Jadhav, Konstantin Dmitriev, Joseph Marino, Matthew Barish, and Arie E. Kaufman. 3D virtual pancreatography. IEEE Transactions on Visualization and Computer Graphics, 28(3):1457– 1468, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Jadhav:2022:CVD

[JDZK22] Shreeraj Jadhav, Gaofeng Deng, Marlene Zawin, and Arie E. Kaufman. COVIDview: Diagnosis of COVID-19 using chest CT. IEEE Transactions on Visualization and Computer Graphics, 28(1):227– 237, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Jin:2021:VCA

[JGC^{+}21] Z. Jin, S. Guo, N. Chen, D. Weiskopf, D. Gotz, and N. Cao. Visual causality analysis of event sequence data. IEEE Transactions on Visualization and Computer Graphics, 27(2):1343–1352, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Jakob:2021:FFD

[JGG21] J. Jakob, M. Gross, and T. Günther. A fluid flow data set for machine learning and its application to neural flow map interpolation. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1279–1289, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Jamonnak:2024:OAH

[JGH^{+}24] Suphanut Jamonnak, Jiajing Guo, Wenbin He, Liang Gou, and Liu Ren. OW-Adapter: Human-assisted open-world object detection with a few examples. IEEE Transactions on Visualization and Computer Graphics, 30(1):694–704, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Jankowai:2020:FLS

[JH20] J. Jankowai and I. Hotz. Feature level-sets: Generalizing isosurfaces to multi-variate data. IEEE Transactions on Visualization and Computer Graphics, 26(2):1308–1319, February 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jenny:2021:CRS

[JHS^{+}21] B. Jenny, M. Heitzler, D. Singh, M. Farmakis-Serebryakova, J. C. Liu, and L. Hurni. Cartographic relief shading with neural networks. IEEE Transactions on Visualization and Computer Graphics, 27(2):1225– 1235, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Joos:2022:VCN

[JJHS^{+}22] Lucas Joos, Sabrina Jaeger-Honz, Falk Schreiber, Daniel A. Keim, and Karsten Klein. Visual comparison of networks in VR. IEEE Transactions on Visualization and Computer Graphics, 28(11):3651– 3661, November 2022. CODEN ITVGEA. ISSN 1077-2626.

John:2020:SUT

[JJKJ20] B. John, S. Jörg, S. Koppal, and E. Jain. The securityutility trade-off for iris authentication and eye animation for social virtual avatars. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1880–1890, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Jadhav:2023:MCI

[JK23] Shreeraj Jadhav and Arie E. Kaufman. MD-Cave: an immersive visualization workbench for radiologists. IEEE Transactions on Visualization and Computer Graphics, 29(12):4832–4844, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Jeon:2024:CCI

[JKA^{+}24] Hyeon Jeon, Yun-Hsin Kuo, Michaël Aupetit, Kwan-Liu Ma, and Jinwook Seo. Classes are not Clusters: Improving label-based evaluation of dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 30(1): 781–791, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Jeon:2022:MEI

[JKJ^{+}22] Hyeon Jeon, Hyung-Kwon Ko, Jaemin Jo, Youngtaek Kim, and Jinwook Seo. Measuring and explaining the inter-cluster reliability of multidimensional projections. IEEE Transactions on Visualization and Computer Graphics, 28(1):551–561, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Jeong:2024:DEM

[JKL24] Sihyun Jeong, Jinwook Kim, and Jeongmi Lee. The differential effects of multisensory attentional cues on task performance in VR depending on the level of cognitive load and cognitive capacity. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2703–2712, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Jonsson:2022:DTE

[JKU^{+}22] Daniel Jönsson, Joel Kronander, Jonas Unger, Thomas B. Schön, and Magnus Wrenninge. Direct transmittance estimation in heterogeneous participating media using approximated Taylor expansions. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2602–2614, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Jaunet:2022:VXR

[JKV^{+}22] Théo Jaunet, Corentin Kervadec, Romain Vuillemot, Grigory Antipov, Moez Baccouche, and Christian Wolf. VisQA: Xraying vision and language reasoning in transformers. IEEE Transactions on Visualization and Computer Graphics, 28(1): 976–986, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Jessup:2022:SFT

[JKW^{+}22] Jared Jessup, Robert Krueger, Simon Warchol, John Hoffer, Jeremy Muhlich, Cecily C. Ritch, Giorgio Gaglia, Shannon Coy, Yu-An Chen, Jia-Ren Lin, Sandro Santagata, Peter K. Sorger, and Hanspeter Pfister. Scope2Screen: Focus+Context techniques for pathology tumor assessment in multivariate image data. IEEE Transactions on Visualization and Computer Graphics, 28(1):259–269, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Jia:2022:TVE

[JLCZ22] Shichao Jia, Zeyu Li, Nuo Chen, and Jiawan Zhang. Towards visual explainable active learning for zero-shot classification. IEEE Transactions on Visualization and Computer Graphics, 28(1):791–801, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Jackson:2024:WGI

[JLH24] Bret Jackson, Linda Lor, and Brianna C. Heggeseth. Workspace guardian: Investigating awareness of personal workspace between co-located augmented reality users. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2724–2733, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Jallepalli:2020:EDT

[JLK20] A. Jallepalli, J. A. Levine, and R. M. Kirby. The effect of data transformations on scalar field topological analysis of highorder FEM solutions. IEEE Transactions on Visualization and Computer Graphics, 26(1): 162–172, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jung:2021:FVV

[JLM^{+}21] Sungchul Jung, Richard Li, Ryan McKee, Mary C. Whitton, and Robert W. Lindeman. Floor-vibration VR: Mitigating cybersickness using whole-body tactile stimuli in highly realistic vehicle driving experiences. IEEE Transactions on Visualization and Computer Graphics, 27(5):2669–2680, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Jin:2023:VAS

[JLP^{+}23] Seungmin Jin, Hyunwook Lee, Cheonbok Park, Hyeshin Chu, Yunwon Tae, Jaegul Choo, and Sungahn Ko. A visual analytics system for improving attentionbased traffic forecasting models.

IEEE Transactions on Visualization and Computer Graphics, 29(1):1102–1112, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Jiao:2023:HSV

[JLX^{+}23] Bo Jiao, Xin Lu, Jingbo Xia, Brij Bhooshan Gupta, Lei Bao, and Qingshan Zhou. Hierarchical sampling for the visualization of large scale-free graphs. IEEE Transactions on Visualization and Computer Graphics, 29(12):5111–5123, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Jung:2022:CVV

[JMK^{+}22] Crescentia Jung, Shubham Mehta, Atharva Kulkarni, Yuhang Zhao, and Yea-Seul Kim. Communicating visualizations without visuals: Investigation of visualization alternative text for people with visual impairments. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1095–1105, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Jardine:2020:PPV

[JOEF20] N. Jardine, B. D. Ondov, N. Elmqvist, and S. Franconeri. The perceptual proxies of visual comparison. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1012–1021, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jeon:2024:CCA

[JQL^{+}24] Hyeon Jeon, Ghulam Jilani Quadri, Hyunwook Lee, Paul Rosen, Danielle Albers Szafir, and Jinwook Seo. CLAMS: a cluster ambiguity measure for estimating perceptual variability in visual clustering. IEEE Transactions on Visualization and Computer Graphics, 30(1): 770–780, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Jiddi:2022:DSR

[JRM22] Salma Jiddi, Philippe Robert, and Eric Marchand. Detecting specular reflections and cast shadows to estimate reflectance and illumination of dynamic indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1249–1260, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Johnson:2020:ABR

[JSA^{+}20] S. Johnson, F. Samsel, G. Abram, D. Olson, A. J. Solis, B. Herman, P. J. Wolfram, C. Lenglet, and D. F. Keefe. Artifact-based rendering: Harnessing natural and traditional visual media for more expressive and engaging 3D visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(1): 492–502, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jo:2020:PPA

[JSF20] J. Jo, J. Seo, and J. Fekete. PANENE: a progressive algorithm for indexing and querying approximate k-nearest neighbors. IEEE Transactions on Visualization and Computer Graphics, 26(2):1347– 1360, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jonsson:2020:IVS

[JSS^{+}20] D. Jönsson, P. Steneteg, E. Sundén, R. Englund, S. Kottravel, M. Falk, A. Ynnerman, I. Hotz, and T. Ropinski. Inviwo: a visualization system with usage abstraction levels. IEEE Transactions on Visualization and Computer Graphics, 26(11):3241–3254, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Jadhav:2023:VEU

[JTT^{+}23] Shreeraj Jadhav, Mahsa Torkaman, Allen Tannenbaum, Saad Nadeem, and Arie E. Kaufman. Volume exploration using multidimensional Bhattacharyya flow. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1651–1663, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Javerliat:2024:PRR

[JVRL24] Charles Javerliat, Sophie Villenave, Pierre Raimbaud, and Guillaume Lavoué. PLUME:

Record, replay, analyze and share user behavior in 6DoF XR experiences. IEEE Transactions on Visualization and Computer Graphics, 30(5):2087– 2097, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Jhan:2022:EST

[JWE^{+}22] Xing-Da Jhan, Sai-Keung Wong, Elham Ebrahimi, Yuwen Lai, Wei-Chia Huang, and Sabarish V. Babu. Effects of small talk with a crowd of virtual humans on users’ emotional and behavioral responses. IEEE Transactions on Visualization and Computer Graphics, 28(11):3767–3777, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Jacobsen:2021:MVS

[JWKN21] B. Jacobsen, M. Wallinger, S. Kobourov, and M. Nöllenburg. MetroSets: Visualizing sets as metro maps. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1257–1267, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Jin:2023:GVA

[JWW^{+}23] Zhihua Jin, Yong Wang, Qianwen Wang, Yao Ming, Tengfei Ma, and Huamin Qu. GNNLens: a visual analytics approach for prediction error diagnosis of graph neural networks. IEEE Transactions on Visualization and Computer Graphics, 29(6):3024– 3038, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Jing:2020:NST

[JYF^{+}20] Y. Jing, Y. Yang, Z. Feng, J. Ye, Y. Yu, and M. Song. Neural style transfer: a review. IEEE Transactions on Visualization and Computer Graphics, 26(11):3365–3385, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Jo:2021:PPV

[JYLS21] Jaemin Jo, Sehi L. Yi, Bongshin Lee, and Jinwook Seo. ProReveal: Progressive visual analytics with safeguards. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3109–3122, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Jiang:2020:SEP

[JYZW20] H. Jiang, D. Yan, X. Zhang, and P. Wonka. Selection expressions for procedural modeling. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1775–1788, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Jiang:2020:DHB

[JZCZ20] B. Jiang, J. Zhang, J. Cai, and J. Zheng. Disentangled human body embedding based on deep hierarchical neural network. IEEE Transactions on Visualization and Computer Graphics, 26(8):2560–2575, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Jamonnak:2022:GCA

[JZHA22] Suphanut Jamonnak, Ye Zhao, Xinyi Huang, and Md Amiruzzaman. Geo-context aware study of vision-based autonomous driving models and spatial video data. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1019–1029, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Kourtesis:2023:CCM

[KAL^{+}23] Panagiotis Kourtesis, Rayaan Amir, Josie Linnell, Ferran Argelaguet, and Sarah E. MacPherson. Cybersickness, cognition, & motor skills: The effects of music, gender, and gaming experience. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2326–2336, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Kwon:2021:DVA

[KAS^{+}21] Bum Chul Kwon, Vibha Anand, Kristen A. Severson, Soumya Ghosh, Zhaonan Sun, Brigitte I. Frohnert, Markus Lundgren, and Kenney Ng. DPVis: Visual analytics with hidden Markov models for disease progression pathways. IEEE Transactions on Visualization and Computer Graphics, 27(9):3685– 3700, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Krishnamurthy:2022:GIS

[KAS^{+}22] Vinayak R. Krishnamurthy, Ergun Akleman, Sai Ganesh Subramanian, Matthew Ebert, Jiaqi Cui, Chia an Fu, and Courtney Starrett. Geometrically interlocking space-filling tiling based on fabric weaves. IEEE Transactions on Visualization and Computer Graphics, 28(10):3391–3404, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Kay:2024:GVD

[Kay24] Matthew Kay. ggdist: Visualizations of distributions and uncertainty in the grammar of graphics. IEEE Transactions on Visualization and Computer Graphics, 30(1):414–424, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Kesavan:2023:SCV

[KBB^{+}23] Suraj P. Kesavan, Harsh Bhatia, Abhinav Bhatele, Stephanie Brink, Olga Pearce, Todd Gamblin, Peer-Timo Bremer, and Kwan-Liu Ma. Scalable comparative visualization of ensembles of call graphs. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1691–1704, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Kreskowski:2022:OSA

[KBF22] Adrian Kreskowski, Stephan Beck, and Bernd Froehlich. Output-sensitive avatar representations for immersive telepresence. IEEE Transactions on Visualization and Computer Graphics, 28(7):2697– 2709, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Krueger:2020:FCU

[KBJ^{+}20] R. Krueger, J. Beyer, W. Jang, N. W. Kim, A. Sokolov, P. K. Sorger, and H. Pfister. Facetto: Combining unsupervised and supervised learning for hierarchical phenotype analysis in multi-channel image data. IEEE Transactions on Visualization and Computer Graphics, 26(1):227–237, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kasica:2021:TSA

[KBM21] S. Kasica, C. Berret, and T. Munzner. Table scraps: an actionable framework for multi-table data wrangling from an artifact study of computational journalism. IEEE Transactions on Visualization and Computer Graphics, 27(2):957– 966, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Kohm:2022:OMF

[KBPR22] Kristopher Kohm, Sabarish V.

Babu, Christopher Pagano, and Andrew Robb. Objects may be farther than they appear: Depth compression diminishes over time with repeated calibration in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3907– 3916, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Khan:2024:DKD

[KBV24] Dawar Khan, Ciril Bohak, and Ivan Viola. Dr. KID: Direct remeshing and K-set isometric decomposition for scalable physicalization of organic shapes. IEEE Transactions on Visualization and Computer Graphics, 30(1):705–715, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Katzakis:2021:EPA

[KCA^{+}21] N. Katzakis, L. Chen, O. Ariza, R. J. Teather, and F. Steinicke. Evaluation of 3D pointing accuracy in the fovea and periphery in immersive head-mounted display environments. IEEE Transactions on Visualization and Computer Graphics, 27(3): 1929–1936, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Kothari:2021:EES

[KCB^{+}21] Rakshit S. Kothari, Aayush K. Chaudhary, Reynold J. Bailey, Jeff B. Pelz, and Gabriel J. Diaz. EllSeg: an ellipse segmentation framework for robust gaze tracking. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2757–2767, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Ke:2023:PLB

[KCGZ23] Pingchuan Ke, Shaoyu Cai, Haichen Gao, and Kening Zhu. PropelWalker: a legbased wearable system with propeller-based force feedback for walking in fluids in VR. IEEE Transactions on Visualization and Computer Graphics, 29(12):5149–5164, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Kim:2024:ETG

[KCK^{+}24] Dae Hyun Kim, Seulgi Choi, Juho Kim, Vidya Setlur, and Maneesh Agrawala. EC: a tool for guiding chart and caption emphasis. IEEE Transactions on Visualization and Computer Graphics, 30(1):120–130, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Krekhov:2020:DVR

[KCWK20] A. Krekhov, S. Cmentowski, A. Waschk, and J. Krüger. Deadeye visualization revisited: Investigation of preattentiveness and applicability in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 26(1):547– 557, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2021:AIH

[KDEP21] Hannah Kim, Barry Drake, Alex Endert, and Haesun Park. ArchiText: Interactive hierarchical topic modeling. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3644–3655, September 2021. CODEN ITVGEA. ISSN 1077-

Kelly:2023:DPV

[Kel23] Jonathan W. Kelly. Distance perception in virtual reality: a meta-analysis of the effect of head-mounted display characteristics. IEEE Transactions on Visualization and Computer Graphics, 29(12):4978– 4989, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Krasner:2024:MEH

[KG24] Alexander Krasner and Joseph Gabbard. MusiKeys: Exploring haptic-to-auditory sensory substitution to improve mid-air text-entry. IEEE Transactions on Visualization and Computer Graphics, 30(5):2247– 2256, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Kristiansen:2022:SSG

[KGB22] Yngve S. Kristiansen, Laura Garrison, and Stefan Bruckner. Semantic snapping for guided multi-view visualization design. IEEE Transactions on Visualization and Computer Graphics, 28(1):43–53, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Klacansky:2020:TLT

[KGBP20] P. Klacansky, A. Gyulassy, P. Bremer, and V. Pascucci. Toward localized topological data structures: Querying the forest for the tree. IEEE Transactions on Visualization and Computer Graphics, 26(1):173–183, January 2020. CODEN ITVGEA.

ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kale:2024:EIM

[KGQ^{+}24] Alex Kale, Ziyang Guo, Xiao Li Qiao, Jeffrey Heer, and Jessica Hullman. EVM: Incorporating model checking into exploratory visual analysis. IEEE Transactions on Visualization and Computer Graphics, 30(1): 208–218, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Kruger:2024:IES

[KGR^{+}24] Marcel Krüger, Tim Gerrits, Timon Römer, Torsten Kuhlen, and Tim Weissker. IntenSelect+: Enhancing score-based selection in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2829–2838, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Kehlbeck:2022:SSP

[KGWD22] Rebecca Kehlbeck, Jochen Görtler, Yunhai Wang, and Oliver Deussen. SPEULER: Semantics-preserving Euler diagrams. IEEE Transactions on Visualization and Computer Graphics, 28(1):433–442, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Kang:2023:NRC

[KGX^{+}23] Kaizhang Kang, Minyi Gu, Cihui Xie, Xuanda Yang, Hongzhi Wu, and Kun Zhou. Neural reflectance capture in the viewillumination domain. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1450–1462, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Kim:2021:GGR

[KH21] Y. Kim and J. Heer. Gemini: a grammar and recommender system for animated transitions in statistical graphics. IEEE Transactions on Visualization and Computer Graphics, 27(2): 485–494, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Kelly:2022:RRL

[KHD^{+}22] Jonathan W. Kelly, Melynda Hoover, Taylor A. Doty, Alex Renner, Moriah Zimmerman, Kimberly Knuth, Lucia A. Cherep, and Stephen B. Gilbert. Remote research on locomotion interfaces for virtual reality: Replication of a lab-based study on teleporting interfaces. IEEE Transactions on Visualization and Computer Graphics, 28(5):2037– 2046, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Kaminokado:2020:SVI

[KHI20] T. Kaminokado, Y. Hiroi, and Y. Itoh. StainedView: Variableintensity light-attenuation display with cascaded spatial color filtering for improved color fidelity. IEEE Transactions on Visualization and Computer Graphics, 26(12):3576– 3586, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Karer:2021:IBN

[KHL21] B. Karer, H. Hagen, and D. J.

Lehmann. Insight beyond numbers: The impact of qualitative factors on visual data analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2):1011–1021, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Kreiser:2021:VSS

[KHR21] Julian Kreiser, Pedro Hermosilla, and Timo Ropinski. Void space surfaces to convey depth in vessel visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(10):3913–3925, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Kimura:2021:MSP

[KIPS21] Sorashi Kimura, Daisuke Iwai, Parinya Punpongsanon, and Kosuke Sato. Multifocal stereoscopic projection mapping. IEEE Transactions on Visualization and Computer Graphics, 27(11):4256–4266, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Kageyama:2022:OPD

[KIS22] Yuta Kageyama, Daisuke Iwai, and Kosuke Sato. Online projector deblurring using a convolutional neural network. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2223–2233, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Kiyokawa:2022:VVRb

[Kiy22a] Kiyoshi Kiyokawa. VGTC Virtual Reality Service Award. IEEE Transactions on Visualization and Computer Graphics, 28(5):xviii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Kiyokawa:2022:VVRa

[Kiy22b] Kiyoshi Kiyokawa. VGTC Virtual Reality Technical Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 28(5):xvi, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Kubo:2021:PNE

[KJI^{+}21] H. Kubo, S. Jayasuriya, T. Iwaguchi, T. Funatomi, Y. Mukaigawa, and S. G. Narasimhan. Programmable non-epipolar indirect light transport: Capture and analysis. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2421–2436, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Kim:2024:LTD

[KJL24] Hyunjeong Kim, Sang-Bin Jeon, and In-Kwon Lee. Locomotion techniques for dynamic environments: Effects on spatial knowledge and user experiences. IEEE Transactions on Visualization and Computer Graphics, 30(5):2184– 2194, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Kong:2023:ARG

[KJS^{+}23] Wenyuan Kong, Zhaoyun Jiang, Shizhao Sun, Zhuoning Guo, Weiwei Cui, Ting Liu, Jianguang Lou, and Dongmei Zhang. Aesthetics++: Refining graphic designs by exploring design principles and human preference. IEEE Transactions on Visualization and Computer Graphics, 29(6):3093– 3104, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Knittel:2021:PCT

[KKE21] Johannes Knittel, Steffen Koch, and Thomas Ertl. Pyramid-Tags: Context-, time- and word order-aware tag maps to explore large document collections. IEEE Transactions on Visualization and Computer Graphics, 27(12):4455– 4468, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Khayat:2020:VGF

[KKEG20] M. Khayat, M. Karimzadeh, D. S. Ebert, and A. Ghafoor. The validity, generalizability and feasibility of summative evaluation methods in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 26(1):353–363, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krake:2023:CDM

[KKEW23] Tim Krake, Daniel Klötzl, Bernhard Eberhardt, and Daniel Weiskopf. Constrained dynamic mode decomposition. IEEE Transactions on Visualization and Computer Graphics, 29(1):182–192, January 2023.

CODEN ITVGEA. ISSN 1077-2626.

Kulik:2020:MPV

[KKF20] A. Kulik, A. Kunert, and B. Froehlich. On motor performance in virtual 3D object manipulation. IEEE Transactions on Visualization and Computer Graphics, 26(5):2041– 2050, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Kammer:2020:GVE

[KKG^{+}20] D. Kammer, M. Keck, T. Gründer, A. Maasch, T. Thom, M. Kleinsteuber, and R. Groh. Glyphboard: Visual exploration of high-dimensional data combining glyphs with dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 26(4):1661– 1671, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2021:BAI

[KKGMH21] Y.-S. Kim, P. Kayongo, M. Grunde-McLaughlin, and J. Hullman. Bayesian-assisted inference from visualized data. IEEE Transactions on Visualization and Computer Graphics, 27(2): 989–999, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Kale:2021:VRS

[KKH21] A. Kale, M. Kay, and J. Hullman. Visual reasoning strategies for effect size judgments and decisions. IEEE Transactions on Visualization and Computer Graphics, 27(2):272– 282, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Kwon:2023:DGM

[KKhCM23] Oh-Hyun Kwon, Chiun-How Kao, Chun houh Chen, and Kwan-Liu Ma. A deep generative model for reordering adjacency matrices. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3195–3208, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Kim:2021:GVA

[KKJ^{+}21] Y. Kim, J. Kim, H. Jeon, Y.- H. Kim, H. Song, B. Kim, and J. Seo. Githru: Visual analytics for understanding software development history through Git metadata analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2):656–666, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Kim:2021:CCL

[KKLS21] D. Kim, H. Kye, J. Lee, and Y.- G. Shin. Confidence-controlled local isosurfacing. IEEE Transactions on Visualization and Computer Graphics, 27(1):29– 42, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Kim:2022:ISA

[KKS^{+}22] You-Jin Kim, Radha Kumaran, Ehsan Sayyad, Anne Milner, Tom Bullock, Barry Giesbrecht, and Tobias Höllerer. Investigating search among physical and virtual objects under different lighting conditions. IEEE Transactions on Visualization and Computer Graphics, 28(11):3788–3798, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Knittel:2022:RTV

[KKT^{+}22] Johannes Knittel, Steffen Koch, Tan Tang, Wei Chen, Yingcai Wu, Shixia Liu, and Thomas Ertl. Real-time visual analysis of high-volume social media posts. IEEE Transactions on Visualization and Computer Graphics, 28(1):879–889, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Koch:2022:ICV

[KKV22] Marius K. Koch, Paul H. J. Kelly, and Peter E. Vincent. Identification and classification of off-vertex critical points for contour tree construction on unstructured meshes of hexahedra. IEEE Transactions on Visualization and Computer Graphics, 28(12):5178– 5180, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Kang:2023:PVT

[KKW23] Hyeokmook Kang, Taeho Kang, and Christian Wallraven. Putting vision and touch into conflict: Results from a multimodal mixed reality setup. IEEE Transactions on Visualization and Computer Graphics, 29(12):5224–5234, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Khayat:2020:VVA

[KKZE20] M. Khayat, M. Karimzadeh, J. Zhao, and D. S. Ebert. VASSL: a visual analytics toolkit for social spambot labeling. IEEE Transactions on Visualization and Computer Graphics, 26(1):874–883, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kim:2022:SEC

[KL22] Hayeon Kim and In-Kwon Lee.

Studying the effects of congruence of auditory and visual stimuli on virtual reality experiences. IEEE Transactions on Visualization and Computer Graphics, 28(5):2080– 2090, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Klenert:2024:LIA

[KLB24] Nicolas Klenert, Verena Lepper, and Daniel Baum. A local iterative approach for the extraction of 2D manifolds from strongly curved and folded thin-layer structures. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1260–1270, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Kim:2020:BCV

[KLCK20] D. Kim, M. Lee, Y. Cho, and D. Kim. Beta-complex versus alpha-complex: Similarities and dissimilarities. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1686–1701, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Klinker:2022:VLA

[Kli22]  Gudrun Klinker.  VGTC

Lifetime Achievement Award. IEEE Transactions on Visualization and Computer Graphics, 28(5):xvii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Knittel:2021:VND

[KLKE21] J. Knittel, A. Lalama, S. Koch, and T. Ertl. Visual neural decomposition to explain multivariate data sets. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1374–1384, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Kurth:2022:CAB

[KLSB22] Philipp Kurth, Markus Leuschner, Marc Stamminger, and Frank Bauer. Content-aware brightness solving and error mitigation in large-scale multiprojection mapping. IEEE Transactions on Visualization and Computer Graphics, 28(11):3607–3617, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Kugelstadt:2021:IDP

[KLTB21] T. Kugelstadt, A. Longva, N. Thuerey, and J. Bender. Implicit density projection for volume conserving liquids. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2385–2395, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Kwon:2020:DGM

[KM20] O. Kwon and K. Ma. A deep generative model for graph layout. IEEE Transactions on Visualization and Computer Graphics, 26(1):665–675, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Krs:2021:PPI

[KMG^{+}21] Vojtěch Krs, Radomír Měch, Mathieu Gaillard, Nathan Carr, and Bedrich Benes. PICO: Procedural iterative constrained optimizer for geometric modeling. IEEE Transactions on Visualization and Computer Graphics, 27(10):3968– 3981, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Kodama:2023:ECT

[KMH^{+}23] Daiki Kodama, Takato Mizuho, Yuji Hatada, Takuji Narumi, and Michitaka Hirose. Effects of collaborative training using virtual co-embodiment on motor skill learning. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2304–2314, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Kruchten:2024:MBE

[KMM24] Nicolas Kruchten, Andrew M.

McNutt, and Michael J. McGuffin. Metrics-based evaluation and comparison of visualization notations. IEEE Transactions on Visualization and Computer Graphics, 30(1):425– 435, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Karduni:2021:BCA

[KMWD21] A. Karduni, D. Markant, R. Wesslen, and W. Dou. A Bayesian cognition approach for belief updating of correlation judgement through uncertainty visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(2):978– 988, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Khan:2022:PVD

[KNAR^{+}22] Saiful Khan, Phong H. Nguyen, Alfie Abdul-Rahman, Benjamin Bach, Min Chen, Euan Freeman, and Cagatay Turkay. Propagating visual designs to numerous plots and dashboards. IEEE Transactions on Visualization and Computer Graphics, 28(1):86–95, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Kern:2023:TIN

[KNL23] Florian Kern, Florian Niebling, and Marc Erich Latoschik. Text input for non-stationary XR workspaces: Investigating tap and word-gesture keyboards in virtual and augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(5):2658–2669, May 2023. CODEN ITVGEA. ISSN 1077-

Kern:2021:CRT

[KNM^{+}21] Michael Kern, Christoph Neuhauser, Torben Maack, Mengjiao Han, Will Usher, and Rüdiger Westermann. A comparison of rendering techniques for 3D line sets with transparency. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3361–3376, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Kelly:2020:TTV

[KOL^{+}20] J. W. Kelly, A. G. Ostrander, A. F. Lim, L. A. Cherep, and S. B. Gilbert. Teleporting through virtual environments: Effects of path scale and environment scale on spatial updating. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1841–1850, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Khan:2022:SRS

[KPF^{+}22] Dawar Khan, Alexander Plopski, Yuichiro Fujimoto, Masayuki Kanbara, Gul Jabeen, Yongjie Jessica Zhang, Xiaopeng Zhang, and Hirokazu Kato. Surface remeshing: a systematic literature review of methods and research directions. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1680–1713, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Kim:2023:MES

[KPL23] Hayeon Kim, Jinhyung Park, and In-Kwon Lee. To be or not to be me?: Exploration of selfsimilar effects of avatars on social virtual reality experiences. IEEE Transactions on Visualization and Computer Graphics, 29(11):4794–4804, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Kusumastuti:2023:PIP

[KPO^{+}23] Sarah A. Kusumastuti, Kimberly A. Pollard, Ashley H. Oiknine, Bianca Dalangin, Tiffany R. Raber, and Benjamin T. Files. Practice improves performance of a 2D uncertainty integration task within and across visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(9):3949– 3960, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Krajancich:2020:FOS

[KPW20] B. Krajancich, N. Padmanaban, and G. Wetzstein. Factored occlusion: Single spatial light modulator occlusion-capable optical see-through augmented reality display. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1871–1879, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Kim:2024:DMI

[KRHH24] Hyeok Kim, Ryan Rossi, Jessica Hullman, and Jane Hoffswell. Dupo: a mixed-initiative authoring tool for responsive visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):934–943, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Kim:2021:PAI

[KRK21] Young Min Kim, Sangwoo Ryu, and Ig-Jae Kim. Planar abstraction and inverse rendering of 3D indoor environments. IEEE Transactions on Visualization and Computer Graphics, 27(6):2992–3006, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Kim:2022:AAR

[KRS^{+}22] Hyeok Kim, Ryan Rossi, Abhraneel Sarma, Dominik Moritz, and Jessica Hullman. An automated approach to reasoning about task-oriented insights in responsive visualization. IEEE Transactions on Visualization and Computer Graphics, 28(1):129–139, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Kiesel:2021:VAA

[KRW^{+}21] D. Kiesel, P. Riehmann, H. Wachsmuth, B. Stein, and B. Froehlich. Visual analysis of argumentation in essays. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1139–1148, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Khan:2020:MST

[KRZ^{+}20] F. Khan, L. Roy, E. Zhang, B. Qu, S. Hung, H. Yeh, R. S. Laramee, and Y. Zhang. Multi-scale topological analysis of asymmetric tensor fields on surfaces. IEEE Transactions on Visualization and Computer Graphics, 26(1):270–279, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kutak:2022:PVV

[KSB^{+}22] David Kuťák, Matias Nicolás Selzer, Jan Byška, María Luján Ganuza, Ivan Barišić, Barbora Kozlíková, and Haichao Miao. Vivern: a virtual environment for multiscale visualization and modeling of DNA nanostructures. IEEE Transactions on Visualization and Computer Graphics, 28(12):4825– 4838, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Kayongo:2022:VE

[KSHH22] Paula Kayongo, Glenn Sun, Jason Hartline, and Jessica Hullman. Visualization equilibrium. IEEE Transactions on Visualization and Computer Graphics, 28(1):465–474, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Kumpf:2022:VAM

[KSHW22] Alexander Kumpf, Josef Stumpfegger, Patrick Fabian Härtl, and Rüdiger Westermann. Visual analysis of multi-parameter distributions across ensembles of 3D fields. IEEE Transactions on Visualization and Computer Graphics, 28(10):3530– 3545, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Kleinbeck:2023:IAI

[KSK^{+}23] Constantin Kleinbeck, Hannah Schieber, Julian Kreimeier, Alejandro Martin-Gomez, Mathias Unberath, and Daniel Roth. Injured avatars: The impact of embodied anatomies and virtual injuries on well-being and performance. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4503–4513, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Kern:2024:HTI

[KTL24] Florian Kern, Jonathan Tschanter, and Marc Erich Latoschik. Handwriting for text input and the impact of XR displays, surface alignments, and sentence complexities. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2357–2367, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Khenak:2020:SPP

[KVB20] N. Khenak, J. Vézien, and P. Bourdot. Spatial presence, performance, and behavior between real, remote, and virtual immersive environments. IEEE Transactions on Visualization and Computer Graphics, 26(12):3467–3478, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Klein:2020:MSP

[KVGM20] T. Klein, I. Viola, E. Gröller, and P. Mindek. Multi-scale procedural animations of microtubule dynamics based on measured data. IEEE Transactions on Visualization and Computer Graphics, 26(1):622–632, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Kourtesis:2022:ASP

[KVM^{+}22] Panagiotis Kourtesis, Sebastian Vizcay, Maud Marchal, Claudio Pacchierotti, and Ferran Argelaguet. Action-specific perception & performance on a Fitts’s Law task in virtual reality: The role of haptic feedback. IEEE Transactions on Visualization and Computer Graphics, 28(11):3715–3726, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Kopp:2023:TMT

[KW23] Wiebke Köpp and Tino Weinkauf. Temporal merge tree maps: a topology-based static visualization for temporal scalar data. IEEE Transactions on Visualization and Computer Graphics, 29(1):1157–1167, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Kunert:2020:MWI

[KWFK20] A. Kunert, T. Weissker, B. Froehlich, and A. Kulik. Multi-window 3D interaction for collaborative virtual reality. IEEE Transactions on Visualization and Computer Graphics, 26(11):3271–3284, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Kale:2022:CSM

[KWH22] Alex Kale, Yifan Wu, and Jessica Hullman. Causal support: Modeling causal inferences with visualizations. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1150–1160, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Kraus:2020:IIC

[KWO^{+}20] M. Kraus, N. Weiler, D. Oelke, J. Kehrer, D. A. Keim, and J. Fuchs. The impact of immersion on cluster identification tasks. IEEE Transactions on Visualization and Computer Graphics, 26(1):525–535, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Khan:2023:UEV

[KZD^{+}23] Talha Khan, Toby S. Zhu, Thomas Downes, Lucille Cheng, Nicolás M. Kass, Edward G. Andrews, and Jacob T. Biehl. Understanding effects of visual feedback delay in AR on fine motor surgical tasks. IEEE Transactions on Visualization and Computer Graphics, 29(11):4697–4707, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Lin:2024:VIM

[LAC^{+}24] Tica Lin, Alexandre Aouididi, Zhutian Chen, Johanna Beyer, Hanspeter Pfister, and Jui-Hsien Wang. VIRD: Immersive match video analysis for high-performance badminton coaching. IEEE Transactions on Visualization and Computer Graphics, 30(1):458– 468, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Lohfink:2022:KRA

[LALG22] Anna-Pia Lohfink, Simon D. Duque Anton, Heike Leitte, and Christoph Garth. Knowledge rocks: Adding knowledge assistance to visualization systems. IEEE Transactions on Visualization and Computer Graphics, 28(1):1117– 1127, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:DHI

[LAML23] Haihan Lin, Derya Akbaba, Miriah Meyer, and Alexander Lex. Data hunches: Incorporating personal knowledge into visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):504– 514, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lee:2021:EIV

[LAN21] A. Lee, D. Archambault, and M. A. Nacenta. The effectiveness of interactive visualization techniques for time navigation of dynamic graphs on large displays. IEEE Transactions on Visualization and Computer Graphics, 27(2):528– 538, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lohfink:2020:SPV

[LAS^{+}20] A. Lohfink, S. D. D. Anton, H. D. Schotten, H. Leitte, and C. Garth. Security in process: Visually supported triage analysis in industrial process data. IEEE Transactions on Visualization and Computer Graphics, 26(4):1638–1649, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2022:SGA

[LB22] Yushi Li and George Baciu. SG-GAN: Adversarial self-attention GCN for point cloud topological parts generation. IEEE Transactions on Visualization and Computer Graphics, 28 (10):3499–3512, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Lekschas:2020:PDN

[LBB^{+}20] F. Lekschas, M. Behrisch, B. Bach, P. Kerpedjiev, N. Gehlenborg, and H. Pfister. Patterndriven navigation in 2D multiscale visualizations with scalable insets. IEEE Transactions on Visualization and Computer Graphics, 26(1):611–621, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2020:URA

[LBE20] J. Liu, N. Boukhelifa, and J. R.

Eagan. Understanding the role of alternatives in data analysis practices. IEEE Transactions on Visualization and Computer Graphics, 26(1):66–76, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2021:DVE

[LBL^{+}21] B. Lee, D. Brown, B. Lee, C. Hurter, S. Drucker, and T. Dwyer. Data visceralization: Enabling deeper understanding of data using virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(2):1095–1105, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Liao:2024:PVD

[LBP24] Shuqi Liao, Vetria Byrd, and Voicu Popescu. PreVR: Variable-distance previews for higher-order disocclusion in VR. IEEE Transactions on Visualization and Computer Graphics, 30(5):2454– 2463, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:DRT

[LBW^{+}22] Chenhui Li, George Baciu, Yunzhe Wang, Junjie Chen, and Changbo Wang. DDLVis: Realtime visual query of spatiotemporal data distribution via density dictionary learning. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1062–1072, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2020:SST

[LCC^{+}20] J. Li, S. Chen, W. Chen, G. Andrienko, and N. Andrienko. Semantics–space–time cube: a conceptual framework for systematic analysis of texts in space and time. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1789–1806, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lu:2023:ELS

[LCC^{+}23] Zhicheng Lu, Xiaoming Chen, Vera Yuk Ying Chung, Weidong Cai, and Yiran Shen. EV-LFV: Synthesizing light field event streams from an event camera and multiple RGB cameras. IEEE Transactions on Visualization and Computer Graphics, 29(11):4546–4555, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Lyu:2022:RTH

[LCCZ22] Qing Lyu, Menglei Chai, Xiang Chen, and Kun Zhou. Real-time hair simulation with neural interpolation. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1894–1905, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2021:DEP

[LCH^{+}21] Shuai Li, Jiahao Cui, Aimin Hao, Shuyang Zhang, and Qinping Zhao. Design and evaluation of personalized percutaneous coronary intervention surgery simulation system. IEEE Transactions on Visualization and Computer Graphics, 27(11):4150–4160, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:HQT

[LCK^{+}21] Z.-N. Liu, Y.-P. Cao, Z.-F. Kuang, L. Kobbelt, and S.-M. Hu. High-quality textured 3D shape reconstruction with cascaded fully convolutional networks. IEEE Transactions on Visualization and Computer Graphics, 27(1):83–97, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:MIT

[LCK^{+}23] Jinghuai Lin, Johrine Cronjé, Ivo Käthner, Paul Pauli, and Marc Erich Latoschik. Measuring interpersonal trust towards virtual humans with a virtual maze paradigm. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2401–2411, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:JRI

[LCL^{+}22] Kunlin Liu, Dongdong Chen, Jing Liao, Weiming Zhang, Hang Zhou, Jie Zhang, Wenbo Zhou, and Nenghai Yu. JPEG robust invertible grayscale. IEEE Transactions on Visualization and Computer Graphics, 28(12):4403–4417, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Luong:2023:CBH

[LCM^{+}23] Tiffany Luong, Yi Fei Cheng, Max Möbus, Andreas Fender, and Christian Holz. Controllers or bare hands? A controlled evaluation of input techniques on interaction performance and exertion in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4633– 4643, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2024:CBV

[LCS^{+}24] Renzhong Li, Weiwei Cui, Tianqi Song, Xiao Xie, Rui Ding, Yun Wang, Haidong Zhang, Hong Zhou, and Yingcai Wu. Causality-based visual analysis of questionnaire responses. IEEE Transactions on Visualization and Computer Graphics, 30(1):638–648, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Livesu:2022:DLT

[LCSA22] Marco Livesu, Gianmarco Cherchi, Riccardo Scateni, and Marco Attene. Deterministic linear time constrained triangulation using simplified earcut. IEEE Transactions on Visualization and Computer Graphics, 28(12):5172–5177, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:VIR

[LCW^{+}23a] Jinghuai Lin, Johrine Cronjé, Carolin Wienrich, Paul Pauli, and Marc Erich Latoschik. Visual indicators representing avatars’ authenticity in social virtual reality and their impacts on perceived trustworthiness. IEEE Transactions on Visualization and Computer Graphics, 29(11):4589–4599, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Luo:2023:MER

[LCW^{+}23b] Zhenxiao Luo, Baili Chai, Zelong Wang, Miao Hu, and Di Wu. Masked360: Enabling robust 360-degree video streaming with ultra low bandwidth consumption. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2690–2699, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:GBI

[LCWL23] Jie Lin, Yi Cai, Xin Wu, and Jianwei Lu. Graph-based information block detection in infographic with gestalt organization principles. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1705–1718, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Lin:2023:QOE

[LCY^{+}23] Tica Lin, Zhutian Chen, Yalong Yang, Daniele Chiappalupi, Johanna Beyer, and Hanspeter Pfister. The quest for Omnioculars: Embedded visualization for augmenting basketball game viewing experiences. IEEE Transactions on Visualization and Computer Graphics, 29(1):962–971, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lo:2024:WCM

[LCYQ24] Leo Yu-Ho Lo, Yifan Cao, Leni Yang, and Huamin Qu. Why change my design: Explaining poorly constructed visualization designs with explorable explanations. IEEE Transactions on Visualization and Computer Graphics, 30(1):955– 964, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Li:2021:LRT

[LDB^{+}21] David Li, Ruofei Du, Adharsh Babu, Camelia D. Brumar, and Amitabh Varshney. A log-rectilinear transformation for foveated 360-degree video streaming. IEEE Transactions on Visualization and Computer Graphics, 27(5):2638– 2647, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Liang:2023:TFP

[LDC^{+}23] Xin Liang, Sheng Di, Franck Cappello, Mukund Raj, Chunhui Liu, Kenji Ono, Zizhong Chen, Tom Peterka, and Hanqi Guo. Toward feature-preserving vector field compression. IEEE Transactions on Visualization and Computer Graphics, 29(12):5434–5450, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:SPS

[LDT^{+}21] J. Liu, T. Dwyer, G. Tack, S. Gratzl, and K. Marriott. Supporting the problem-solving loop: Designing highly interactive optimisation systems.

IEEE Transactions on Visualization and Computer Graphics, 27(2):1764–1774, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lyu:2021:RGP

[LDZ^{+}21] W. Lyu, P. Ding, Y. Zhang, A. Chen, M. Wu, S. Yin, and J. Yu. Refocusable gigapixel panoramas for immersive VR experiences. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2028–2040, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

lejek:2021:CMV

[leáB^{+}21] Radek O. lej ek, Vít Rus ák, Karolína Burská, Valdemar vábenský, Jan Vykopal, and Jakub egan. Conceptual model of visual analytics for hands-on cybersecurity training. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3425–3437, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Lehman:2023:ACE

[LELT23] Sarah M. Lehman, Semir Elezovikj, Haibin Ling, and Chiu C. Tan. ARCHIE++: a cloud-enabled framework for conducting AR system testing in the wild. IEEE Transactions on Visualization and Computer Graphics, 29(4):2102– 2116, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:UML

[LETF21] Jen-Shuo Liu, Carmine Elvezio, Barbara Tversky, and Steven Feiner. Using multi-level precueing to improve performance in path-following tasks in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(11):4311– 4320, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:SRM

[LF23] Huiyu Li and Linwei Fan.

A segmented redirection mapping method for roadmaps of large constrained virtual environments. IEEE Transactions on Visualization and Computer Graphics, 29(12):5308– 5324, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Lu:2021:PDC

[LFC^{+}21] K. Lu, M. Feng, X. Chen, M. Sedlmair, O. Deussen, D. Lischinski, Z. Cheng, and Y. Wang. Palettailor: Discriminable colorization for categorical data. IEEE Transactions on Visualization and Computer Graphics, 27(2):475– 484, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lenz:2024:CSA

[LFCH24] Lara Sofie Lenz, Andreas Rene Fender, Julia Chatain, and Christian Holz. Comparing synchronous and asynchronous task delivery in mixed reality environments. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2776–2784, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Lohesara:2023:HHE

[LFG^{+}23] Fatemeh Ghorbani Lohesara, Davi Rabbouni Freitas, Christine Guillemot, Karen Eguiazarian, and Sebastian Knorr. HEADSET: Human emotion awareness under partial occlusions multimodal DataSET. IEEE Transactions on Visualization and Computer Graphics, 29(11):4686–4696, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Lei:2024:GLF

[LFMM24] Fan Lei, Arlen Fan, Alan M. MacEachren, and Ross Maciejewski. GeoLinter: a linting framework for choropleth maps. IEEE Transactions on Visualization and Computer Graphics, 30(2):1592–1607, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:WWS

[LFO23] Lei Li, Hongbo Fu, and Maks Ovsjanikov. WSDesc: Weakly supervised 3D local descriptor learning for point cloud registration. IEEE Transactions on Visualization and Computer Graphics, 29(7):3368– 3379, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Liang:2021:DAM

[LFR^{+}21] Yuan Liang, Lubin Fan, Peiran Ren, Xuansong Xie, and Xian-Sheng Hua. DecorIn: an automatic method for plane-based decorating. IEEE Transactions on Visualization and Computer Graphics, 27(8):3438–3450, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Lu:2022:ESC

[LFW^{+}22] Min Lu, Noa Fish, Shuaiqi Wang, Joel Lanir, Daniel Cohen-Or, and Hui Huang. Enhancing static charts with data-driven animations. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2628–2640, July 2022. CODEN ITVGEA. ISSN 1077-2626.

LYi:2023:MVD

[LG23] Sehi L’Yi and Nils Gehlenborg.

Multi-view design patterns and responsive visualization for genomics data. IEEE Transactions on Visualization and Computer Graphics, 29(1):559– 569, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2024:UPB

[LGH^{+}24] Chunlei Li, Yang Gao, Jiayi He, Tianwei Cheng, Shuai Li, Aimin Hao, and Hong Qin. A unified particle-based solver for non-Newtonian behaviors simulation. IEEE Transactions on Visualization and Computer Graphics, 30(4):1998– 2010, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:IHM

[LGL^{+}23a] Zehui Lin, Xiang Gu, Sheng Li, Zhiming Hu, and Guoping Wang. Intentional headmotion assisted locomotion for reducing cybersickness. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3458–3471, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2023:LIG

[LGL^{+}23b] Ying-Tian Liu, Yuan-Chen Guo, Yi-Xiao Li, Chen Wang, and Song-Hai Zhang. Learning implicit glyph shape representation. IEEE Transactions on Visualization and Computer Graphics, 29(10):4172– 4182, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Lukasczyk:2021:LTS

[LGMT21] J. Lukasczyk, C. Garth, R. Maciejewski, and J. Tierny. Localized topological simplification of scalar data. IEEE Transactions on Visualization and Computer Graphics, 27(2):572– 582, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lukasczyk:2020:DNT

[LGW^{+}20] J. Lukasczyk, C. Garth, G. H. Weber, T. Biedert, R. Maciejewski, and H. Leitte. Dynamic nested tracking graphs. IEEE Transactions on Visualization and Computer Graphics, 26(1):249–258, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lyu:2021:IST

[LGWL21] Y. Lyu, F. Gao, I.-S. Wu, and B. Y. Lim. Imma Sort by two or more attributes with interpretable monotonic multi-attribute sorting. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2369–2384, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:SIS

[LGY^{+}22] Sheng Li, Xiang Gu, Kangrui Yi, Yanlin Yang, Guoping Wang, and Dinesh Manocha. Self-illusion: a study on cognition of role-playing in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 28(8):3035–3049, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Luong:2022:CPR

[LH22] Tiffany Luong and Christian Holz. Characterizing physiological responses to fear, frustration, and insight in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3917– 3927, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:GTF

[LH23] Shiguang Liu and Jiaqi Hao.

Generating talking face with controllable eye movements by disentangled blinking feature. IEEE Transactions on Visualization and Computer Graphics, 29(12):5050–5061, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Lai:2021:CLU

[LHA^{+}21] Chengyuan Lai, Xinyu Hu, Afham Ahmed Aiyaz, Ann Segismundo, Ananya Phadke, and Ryan P. McMahan. The cognitive loads and usability of target-based and steering-based travel techniques. IEEE Transactions on Visualization and Computer Graphics, 27(11): 4289–4299, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Lee:2021:SSS

[LHC^{+}21] B. Lee, X. Hu, M. Cordeil, A. Prouzeau, B. Jenny, and T. Dwyer. Shared surfaces and spaces: Collaborative data visualisation in a co-located immersive environment. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1171–1181, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Luo:2020:NNA

[LHL^{+}20] J. Luo, Z. Huang, Y. Li, X. Zhou, G. Zhang, and H. Bao. NIID-Net: Adapting surface normal knowledge for intrinsic image decomposition in indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 26(12):3434– 3445, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:GGD

[LHL^{+}22] Jingying Liu, Binyuan Hui, Kun Li, Yunke Liu, Yu-Kun Lai, Yuxiang Zhang, Yebin Liu, and Jingyu Yang. Geometryguided dense perspective network for speech-driven facial animation. IEEE Transactions on Visualization and Computer Graphics, 28(12):4873– 4886, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:SPC

[LHS^{+}22] Wanwan Li, Haikun Huang, Tomay Solomon, Behzad Esmaeili, and Lap-Fai Yu. Synthesizing personalized construction safety training scenarios for VR training. IEEE Transactions on Visualization and Computer Graphics, 28(5): 1993–2002, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:ASA

[LHWW22] Shusen Liu, Xiaowei He, Wencheng Wang, and Enhua Wu. Adapted SIMPLE algorithm for incompressible SPH fluids with a broad range viscosity. IEEE Transactions on Visualization and Computer Graphics, 28(9):3168– 3179, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:LII

[LHZ^{+}23] Xiongzheng Li, Jing Huang, Jinsong Zhang, Xiaokun Sun, Haibiao Xuan, Yu-Kun Lai, Yingdi Xie, Jingyu Yang, and Kun Li. Learning to infer inner-body under clothing from monocular video. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5083–5096, December 2023.

CODEN ITVGEA. ISSN 1077-2626.

Liu:2024:TPA

[LI24] Guoxi Liu and Federico Iuricich. A task-parallel approach for localized topological data structures. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1271–1281, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Leon:2024:EMC

[LIB24] Gabriela Molina León, Petra Isenberg, and Andreas Breiter. Eliciting multimodal and collaborative interactions for data exploration on large vertical displays. IEEE Transactions on Visualization and Computer Graphics, 30(2):1624– 1637, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Liu:2020:SGR

[LIDM20] Z. Liu, T. Itoh, J. Q. Dawson, and T. Munzner. The Sprawlter graph readability metric: Combining sprawl and area-aware clutter. IEEE Transactions on Visualization and Computer Graphics, 26(6):2180– 2191, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:TLD

[LIFD23] Guoxi Liu, Federico Iuricich, Riccardo Fellegara, and Leila De Floriani. TopoCluster: a localized data structure for topology-based visualization.

IEEE Transactions on Visualization and Computer Graphics, 29(2):1506–1517, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Livesu:2021:SMR

[Liv21] M. Livesu. Scalable mesh refinement for canonical polygonal schemas of extremely high genus shapes. IEEE Transactions on Visualization and Computer Graphics, 27(1):254– 260, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Lee:2024:RSS

[LJCL24] Ho Jung Lee, Sang-Bin Jeon, Yong-Hun Cho, and In-Kwon Lee. Redirection strategy switching: Selective redirection controller for dynamic environment adaptation. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2474–2484, May 2024. CODEN ITVGEA. ISSN 1077-2626.

LYi:2021:CLR

[LJS21] Sehi L’Yi, Jaemin Jo, and Jinwook Seo. Comparative layouts revisited: Design space, guidelines, and future directions. IEEE Transactions on Visualization and Computer Graphics, 27(2):1525–1535, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:BAV

[LKAH21] Y. Liu, A. Kale, T. Althoff, and J. Heer. Boba: Authoring and visualizing multiverse analyses.

IEEE Transactions on Visualization and Computer Graphics, 27(2):1753–1763, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lee:2020:VAS

[LKJ^{+}20] C. Lee, Y. Kim, S. Jin, D. Kim, R. Maciejewski, D. Ebert, and S. Ko. A visual analytics system for exploring, monitoring, and forecasting road traffic congestion. IEEE Transactions on Visualization and Computer Graphics, 26(11):3133– 3146, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Lee:2023:ERS

[LKL23] Seung Un Lee, Jinwook Kim, and Jeongmi Lee. Effects of reward schedule and avatar visibility on joint agency during VR collaboration. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4372–4382, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:UVR

[LKS22] Chen Li, Angel Lo Lo Kon, and Horace Ho Shing Ip. Use virtual reality to enhance intercultural sensitivity: a randomised parallel longitudinal study. IEEE Transactions on Visualization and Computer Graphics, 28(11):3673– 3683, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Lyu:2020:OMB

[LLC^{+}20] Y. Lyu, X. Liu, H. Chen, A. Mangal, K. Liu, C. Chen, and B. Lim. OD morphing: Balancing simplicity with faithfulness for OD bundling. IEEE Transactions on Visualization and Computer Graphics, 26(1): 811–821, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2022:RIF

[LLC^{+}22] Xianzhi Li, Ruihui Li, Guangyong Chen, Chi-Wing Fu, Daniel Cohen-Or, and Pheng-Ann Heng. A rotation-invariant framework for deep point cloud analysis. IEEE Transactions on Visualization and Computer Graphics, 28(12):4503– 4514, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Lu:2022:SFB

[LLCH22] Jia-Ming Lu, Chen-Feng Li, Geng-Chen Cao, and Shi-Min Hu. Simulating fractures with bonded discrete element method. IEEE Transactions on Visualization and Computer Graphics, 28(12):4810– 4824, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2021:KBM

[LLD^{+}21] Wei Li, Daoming Liu, Mathieu Desbrun, Jin Huang, and Xiaopei Liu. Kinetic-based multiphase flow simulation. IEEE Transactions on Visualization and Computer Graphics, 27(7):

3318–3334, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Liang:2024:FNC

[LLDW24] Zhicheng Liang, Junhua Liu, Mallesham Dasari, and Fangxin Wang. Fumos: Neural compression and progressive refinement for continuous point cloud video streaming. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2849–2859, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:SAC

[LLK22] Jing Li, Tiantian Liu, and Ladislav Kavan. Soft articulated characters in projective dynamics. IEEE Transactions on Visualization and Computer Graphics, 28(2):1385– 1396, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Lin:2022:SMS

[LLL^{+}22] Cheng Lin, Lingjie Liu, Changjian Li, Leif Kobbelt, Bin Wang, Shiqing Xin, and Wenping Wang. SEG-MAT: 3D shape segmentation using medial axis transform. IEEE Transactions on Visualization and Computer Graphics, 28(6):2430– 2444, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Luong:2022:SAC

[LLMA22] Tiffany Luong, Anatole Lecuyer, Nicolas Martin, and Ferran Argelaguet. A survey on affective and cognitive VR. IEEE Transactions on Visualization and Computer Graphics, 28(12):5154–5171, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Linhares:2023:CVT

[LLP^{+}23] Claudio D. G. Linhares, Daniel M.

Lima, Jean R. Ponciano, Mauro M. Olivatto, Marco A. Gutierrez, Jorge Poco, Caetano Traina, and Agma J. M. Traina. ClinicalPath: a visualization tool to improve the evaluation of electronic health records in clinical decision-making. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4031–4046, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Lee:2020:YCA

[LLS^{+}20] D. J. Lee, J. Lee, T. Siddiqui, J. Kim, K. Karahalios, and A. Parameswaran. You can’t always sketch what you want: Understanding sensemaking in visual query systems. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1267–1277, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lee:2024:MSM

[LLSM24] Geonsun Lee, Dae Yeol Lee, Guan-Ming Su, and Dinesh Manocha. “May I speak?”: Multi-modal attention guidance in social VR group conversations. IEEE Transactions on Visualization and Computer Graphics, 30(5):2287– 2297, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Lai:2024:EYN

[LLT24] Shaoxuan Lai, Wanna Luan, and Jun Tao. Explore your network in minutes: a rapid prototyping toolkit for understanding neural networks with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 30(1):683–693, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:MTG

[LLW^{+}22a] Zheng Liu, Yanlei Li, Weina Wang, Ligang Liu, and Renjie Chen. Mesh total generalized variation for denoising. IEEE Transactions on Visualization and Computer Graphics, 28(12):4418–4433, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Lu:2022:MJN

[LLW^{+}22b] Min Lu, Joel Lanir, Chufeng Wang, Yucong Yao, Wen Zhang, Oliver Deussen, and Hui Huang. Modeling just noticeable differences in charts. IEEE Transactions on Visualization and Computer Graphics, 28(1): 718–726, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2023:HIT

[LLW^{+}23] Guozheng Li, Runfei Li, Zicheng Wang, Chi Harold Liu, Min Lu, and Guoren Wang. Hi-Tailor: Interactive transformation and visualization for hierarchical tabular data. IEEE Transactions on Visualization and Computer Graphics, 29(1): 139–148, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2023:PSS

[LLWF23] Ruihui Li, Xianzhi Li, Tien-Tsin Wong, and Chi-Wing Fu. Point set self-embedding. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3226–3237, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Lin:2024:ILS

[LLY^{+}24] Yanna Lin, Haotian Li, Leni Yang, Aoyu Wu, and Huamin Qu. InkSight: Leveraging sketch interaction for documenting chart findings in computational notebooks. IEEE Transactions on Visualization and Computer Graphics, 30(1): 944–954, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2021:DND

[LLZ^{+}21] Xianzhi Li, Ruihui Li, Lei Zhu, Chi-Wing Fu, and Pheng-Ann Heng. DNF-Net: a deep normal filtering network for mesh denoising. IEEE Transactions on Visualization and Computer Graphics, 27 (10):4060–4072, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:ECI

[LLZ^{+}23] Yuanyuan Liu, Chengjiang Long, Zhaoxuan Zhang, Bokai Liu, Qiang Zhang, Baocai Yin, and Xin Yang. Explore contextual information for 3D scene graph generation. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5556–5568, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2020:PPP

[LM20] J. K. Li and K. Ma. P5: Portable progressive parallel processing pipelines for interactive data analysis and visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):1151– 1160, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2021:PDL

[LM21] J. K. Li and K.-L. Ma. P6: a declarative language for integrating machine learning in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 27(2):380– 389, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lu:2022:STD

[LMD^{+}22] Zhuheng Lu, Weiwei Mao, Yuewei Dai, Weiqing Li, and Zhiyong Su. Slicing-tracking-detection: Simultaneous multicylinder detection from largescale and complex point clouds. IEEE Transactions on Visualization and Computer Graphics, 28(12):4172–4185, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Lawonn:2023:GRC

[LME^{+}23] Kai Lawonn, Monique Meuschke, Pepe Eulzer, Matthias Mitterreiter, Joachim Giesen, and Tobias Günther. GRay: Ray casting for visualization and interactive data exploration of Gaussian mixture models. IEEE Transactions on Visualization and Computer Graphics, 29(1): 526–536, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Lei:2024:GVA

[LMF^{+}24] Fan Lei, Yuxin Ma, A. Stewart Fotheringham, Elizabeth A. Mack, Ziqi Li, Mehak Sachdeva, Sarah Bardin, and Ross Maciejewski. GeoExplainer: a visual analytics framework for spatial modeling contextualization and report generation. IEEE Transactions on Visualization and Computer Graphics, 30(1):1391–1401, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:PCC

[LMGY22] Xiaohan Liu, Lei Ma, Jianwei Guo, and Dong-Ming Yan. Parallel computation of 3D clipped Voronoi diagrams. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1363–1372, February 2022. CO-DEN ITVGEA. ISSN 1077-

Li:2021:SRS

[LMM^{+}21] Zhimin Li, Harshitha Menon, Dan Maljovec, Yarden Livnat, Shusen Liu, Kathryn Mohror, Peer-Timo Bremer, and Valerio Pascucci. SpotSDC: Revealing the silent data corruption propagation in high-performance computing systems. IEEE Transactions on Visualization and Computer Graphics, 27 (10):3938–3952, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Lee:2022:P

[LMY^{+}22] Bongshin Lee, Silvia Miksch, Anders Ynnerman, Anastasia Bezerianos, Jian Chen, Wei Chen, Christopher Collins, Michael Gleicher, Eduard Gröller, Alexander Lex, Bernhard Preim, Jinwook Seo, Ruediger Westermann, Jing Yang, Xiaoru Yuan, Han-Wei Shen, Jean-Daniel Fekete, and Shixia Liu. Preface. IEEE Transactions on Visualization and Computer Graphics, 28(1): xiv–xxiii, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Lee:2021:MRT

[LNB^{+}21] Myungho Lee, Nahal Norouzi, Gerd Bruder, Pamela J. Wisniewski, and Gregory F. Welch. Mixed reality tabletop gameplay: Social interaction with a virtual human capable of physical influence. IEEE Transactions on Visualization and Computer Graphics, 27(8):

3534–3545, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2022:NSA

[LPG^{+}22] Ming Li, Junjun Pan, Yang Gao, Yang Shen, Fang Luo, Ju Dai, Aimin Hao, and Hong Qin. Neurophysiological and subjective analysis of VR emotion induction paradigm. IEEE Transactions on Visualization and Computer Graphics, 28(11):3832–3842, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Lange:2022:LUE

[LPJT^{+}22] Devin Lange, Eddie Polanco, Robert Judson-Torres, Thomas Zangle, and Alexander Lex. Loon: Using exemplars to visualize large-scale microscopy data. IEEE Transactions on Visualization and Computer Graphics, 28(1):248–258, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2024:MPA

[LPL^{+}24] Ming Li, Junjun Pan, Yu Li, Yang Gao, Hong Qin, and Yang Shen. Multimodal physiological analysis of impact of emotion on cognitive control in VR. IEEE Transactions on Visualization and Computer Graphics, 30(5):2044–2054, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Linhares:2023:LVE

[LPP^{+}23] Claudio D. G. Linhares, Jean R.

Ponciano, Diogenes S. Pedro, Luis E. C. Rocha, Agma J. M. Traina, and Jorge Poco. LargeNetVis: Visual exploration of large temporal networks based on community taxonomies. IEEE Transactions on Visualization and Computer Graphics, 29(1):203–213, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:DIR

[LQS^{+}23] Yixuan Li, Yusheng Qi, Yang Shi, Qing Chen, Nan Cao, and Siming Chen. Diverse interaction recommendation for public users exploring multiview visualization using deep learning. IEEE Transactions on Visualization and Computer Graphics, 29(1):95–105, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2021:LPP

[LQWQ21] Chen Li, Sheng Qiu, Changbo Wang, and Hong Qin. Learning physical parameters and detail enhancement for gaseous scene design based on data guidance. IEEE Transactions on Visualization and Computer Graphics, 27(10):3867–3880, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:IHK

[LqZ23] Jie Li and Chun qi Zhou. Incorporation of human knowledge into data embeddings to improve pattern significance and interpretability. IEEE Transactions on Visualization and Computer Graphics, 29(1):723– 733, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lee-Robbins:2023:ALO

[LRA23] Elsie Lee-Robbins and Eytan Adar. Affective learning objectives for communicative visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):1–11, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lee-Robbins:2022:LOI

[LRHA22] Elsie Lee-Robbins, Shiqing He, and Eytan Adar. Learning objectives, insights, and assessments: How specification formats impact design. IEEE Transactions on Visualization and Computer Graphics, 28(1): 676–685, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2023:RTS

[LRZ^{+}23] Qiangqiang Liu, Yukun Ren, Zhihua Zhu, Dai Li, Xiaojuan Ma, and Quan Li. RankAxis: Towards a systematic combination of projection and ranking in multi-attribute data exploration. IEEE Transactions on Visualization and Computer Graphics, 29(1):701–711, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lundgard:2022:AVN

[LS22] Alan Lundgard and Arvind Satyanarayan. Accessible visualization via natural language descriptions: a four-level model of semantic content. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1073–1083, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2023:LLR

[LS23] Haoyu Li and Han-Wei Shen. Local latent representation based on geometric convolution for particle data feature exploration. IEEE Transactions on Visualization and Computer Graphics, 29(7):3354– 3367, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:PQA

[LSD^{+}23] Qi Liu, Honglei Su, Zhengfang Duanmu, Wentao Liu, and Zhou Wang. Perceptual quality assessment of colored 3D point clouds. IEEE Transactions on Visualization and Computer Graphics, 29(8):3642–3655, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Luo:2020:CMB

[LSE20] X. Luo, N. Z. Salamon, and E. Eisemann. Controllable motion-blur effects in still images. IEEE Transactions on Visualization and Computer Graphics, 26(7):2362– 2372, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Lee:2024:MEV

[LSG24] Yongho Lee, Heesook Shin, and Youn-Hee Gil. Measurement of empathy in virtual reality with head-mounted displays: a systematic review. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2485–2495, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Lu:2022:LRM

[LSL^{+}22] Xuequan Lu, Scott Schaefer, Jun Luo, Lizhuang Ma, and Ying He. Low rank matrix approximation for 3D geometry filtering. IEEE Transactions on Visualization and Computer Graphics, 28(4):1835– 1847, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:DSC

[LSL^{+}23] Zeyu Li, Ruizhi Shi, Yan Liu, Shizhuo Long, Ziheng Guo, Shichao Jia, and Jiawan Zhang. Dual space coupling model guided overlap-free scatterplot. IEEE Transactions on Visualization and Computer Graphics, 29(1):657–667, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Lee:2024:DPS

[LSS24] Benjamin Lee, Michael Sedlmair, and Dieter Schmalstieg. Design patterns for situated visualization in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 30(1):1324– 1335, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Lee:2022:DCV

[LST^{+}22] Doris Jung-Lin Lee, Vidya Setlur, Melanie Tory, Karrie Karahalios, and Aditya Parameswaran. Deconstructing categorization in visualization recommendation: a taxonomy and comparative study. IEEE Transactions on Visualization and Computer Graphics, 28(12):4225–4239, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Luo:2020:NNN

[LSW^{+}20] R. Luo, T. Shao, H. Wang, W. Xu, X. Chen, K. Zhou, and Y. Yang. NNWarp: Neural network-based nonlinear deformation. IEEE Transactions on Visualization and Computer Graphics, 26(4):1745– 1759, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Lan:2022:KAA

[LSW^{+}22] Xingyu Lan, Yang Shi, Yanqiu Wu, Xiaohan Jiao, and Nan Cao. Kineticharts: Augmenting affective expressiveness of charts in data stories with animation design. IEEE Transactions on Visualization and Computer Graphics, 28(1):933– 943, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Liang:2023:AMS

[LSW^{+}23] Yuzhi Liang, Qi Song, Rui Wang, Yuchi Huo, and Hujun Bao. Automatic mesh and shader level of detail. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4284–4295, October 2023.

CODEN ITVGEA. ISSN 1077-2626.

Lan:2021:SSL

[LSZC21] Xingyu Lan, Yang Shi, Yueyao Zhang, and Nan Cao. Smile or scowl? Looking at infographic design through the affective lens. IEEE Transactions on Visualization and Computer Graphics, 27(6):2796– 2807, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Lamy:2020:RPV

[LT20] J.-B. Lamy and R. Tsopra.

RainBio: Proportional visualization of large sets in biology. IEEE Transactions on Visualization and Computer Graphics, 26(11):3285–3298, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Lalle:2021:GDA

[LTC21] Sébastien Lallé, Dereck Toker, and Cristina Conati. Gazedriven adaptive interventions for magazine-style narrative visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2941–2952, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Luo:2022:NLV

[LTL^{+}22] Yuyu Luo, Nan Tang, Guoliang Li, Jiawei Tang, Chengliang Chai, and Xuedi Qin. Natural language to visualization by neural machine translation. IEEE Transactions on Visualization and Computer Graphics, 28(1):217–226, January 2022.

CODEN ITVGEA. ISSN 1077-2626.

Li:2022:EPR

[LTLB22] Yuan Li, Ibrahim A. Tahmid, Feiyu Lu, and Doug A. Bowman. Evaluation of pointing ray techniques for distant object referencing in model-free outdoor collaborative augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3896– 3906, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Lee:2024:DEV

[LVR^{+}24] Sueyoon Lee, Irene Viola, Silvia Rossi, Zhirui Guo, Ignacio Reimat, Kinga Ławicka, Alina Striner, and Pablo Cesar. Designing and evaluating a VR lobby for a socially enriching remote opera watching experience. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2055–2065, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Lamqaddam:2021:ILM

[LVV^{+}21] H. Lamqaddam, A. Vande Moere, V. Vanden Abeele, K. Brosens, and K. Verbert. Introducing Layers of Meaning (LoM): a framework to reduce semantic distance of visualization in humanistic research. IEEE Transactions on Visualization and Computer Graphics, 27(2):1084–1094, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:VGN

[LWBM22] Zipeng Liu, Yang Wang, Jürgen Bernard, and Tamara Munzner. Visualizing graph neural networks with CorGIE: Corresponding a graph to its embedding. IEEE Transactions on Visualization and Computer Graphics, 28(6):2500– 2516, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:SSC

[LWC22] Lei Li, Wencheng Wang, and Yiyao Chu. A simple and stable centeredness measure for 3D curve skeleton extraction. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1486–1499, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Lan:2024:AVD

[LWC24] Xingyu Lan, Yanqiu Wu, and Nan Cao. Affective visualization design: Leveraging the emotional impact of data. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1–11, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:HDA

[LWD^{+}23] Yiran Li, Junpeng Wang, Xin Dai, Liang Wang, Chin-Chia Michael Yeh, Yan Zheng, Wei Zhang, and Kwan-Liu Ma. How does attention work in vision transformers? A visual analytics attempt. IEEE Transactions on Visualization and Computer Graphics, 29(6):

2888–2900, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:UCS

[LWF23] Xiang Li, Lingjing Wang, and Yi Fang. Unsupervised category-specific partial point set registration via joint shape completion and registration. IEEE Transactions on Visualization and Computer Graphics, 29(7):3251–3265, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:VVB

[LWH^{+}22] Yuanqi Li, Chuan Wang, Jing Hong, Jie Zhu, Jie Guo, Jue Wang, Yanwen Guo, and Wenping Wang. Video vectorization via bipartite diffusion curves propagation and optimization. IEEE Transactions on Visualization and Computer Graphics, 28(9):3265–3276, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Luidolt:2020:GDS

[LWK20] L. R. Luidolt, M. Wimmer, and K. Krösl. Gaze-dependent simulation of light perception in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 26(12): 3557–3567, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Lu:2020:WVA

[LWL^{+}20] M. Lu, S. Wang, J. Lanir, N. Fish, Y. Yue, D. Cohen-Or, and H. Huang. Winglets: Visualizing association with uncertainty in multi-class scatterplots. IEEE Transactions on Visualization and Computer Graphics, 26(1):770–779, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Leng:2022:EFT

[LWL^{+}22a] Jiaye Leng, Lili Wang, Xiaolong Liu, Xuehuai Shi, and Miao Wang. Efficient flower text entry in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3662–3672, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:IRP

[LWL^{+}22b] Quan Li, Xiguang Wei, Huanbin Lin, Yang Liu, Tianjian Chen, and Xiaojuan Ma. Inspecting the running process of horizontal federated learning via visual analytics. IEEE Transactions on Visualization and Computer Graphics, 28(12):4085–4100, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ling:2023:SDV

[LWL^{+}23a] Jingwang Ling, Zhibo Wang, Ming Lu, Quan Wang, Chen Qian, and Feng Xu. Semantically disentangled variational autoencoder for modeling 3D facial details. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3630–3641, August 2023. CO-DEN ITVGEA. ISSN 1077-

Liu:2023:RTL

[LWL^{+}23b] Celong Liu, Lingyu Wang, Zhong Li, Shuxue Quan, and Yi Xu. Real-time lighting estimation for augmented reality via differentiable screen-space rendering. IEEE Transactions on Visualization and Computer Graphics, 29(4):2132– 2145, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2020:STD

[LWM^{+}20] S. Liu, D. Wang, D. Maljovec, R. Anirudh, J. J. Thiagarajan, S. A. Jacobs, B. C. Van Essen, D. Hysom, J. Yeom, J. Gaffney, L. Peterson, P. B. Robinson, H. Bhatia, V. Pascucci, B. K. Spears, and P. Bremer. Scalable topological data analysis and visualization for evaluating data-driven models in scientific applications. IEEE Transactions on Visualization and Computer Graphics, 26(1): 291–300, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2021:CPC

[LWS^{+}21] G. Li, J. Wang, H.-W. Shen, K. Chen, G. Shan, and Z. Lu. CNNPruner: Pruning convolutional neural networks with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1364–1373, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2020:SAD

[LWSY20] C. Liu, C. Wu, H. Shao, and X. Yuan. SmartCube: an adaptive data management architecture for the real-time visualization of spatiotemporal datasets. IEEE Transactions on Visualization and Computer Graphics, 26(1):790–799, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Liu:2023:EVA

[LWT^{+}23] Shuhan Liu, Di Weng, Yuan Tian, Zikun Deng, Haoran Xu, Xiangyu Zhu, Honglei Yin, Xianyuan Zhan, and Yingcai Wu. ECoalVis: Visual analysis of control strategies in coal-fired power plants. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1091–1101, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Lin:2021:TIM

[LWW^{+}21] Y. Lin, K. Wong, Y. Wang, R. Zhang, B. Dong, H. Qu, and Q. Zheng. TaxThemis: Interactive mining and exploration of suspicious tax evasion groups. IEEE Transactions on Visualization and Computer Graphics, 27(2):849–859, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Lukasczyk:2024:ESA

[LWW^{+}24] Jonas Lukasczyk, Michael Will, Florian Wetzels, Gunther H. Weber, and Christoph Garth.

ExTreeM: Scalable augmented merge tree computation via extremum graphs. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1085–1094, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2021:TCS

[LWWF21] Xiang Li, Congcong Wen, Lingjing Wang, and Yi Fang. Topology constrained shape correspondence. IEEE Transactions on Visualization and Computer Graphics, 27(10): 3926–3937, October 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Liang:2020:FCS

[LWWH20] Y. Liang, B. Wang, L. Wang, and N. Holzschuch. Fast computation of single scattering in participating media with refractive boundaries using frequency analysis. IEEE Transactions on Visualization and Computer Graphics, 26 (10):2961–2969, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Lv:2021:LFS

[LWWY21] Xianqiang Lv, Xue Wang, Qing Wang, and Jingyi Yu. 4D light field segmentation from light field super-pixel hypergraph representation. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3597–3610, September 2021. CODEN ITVGEA. ISSN 1077-

Li:2020:WVA

[LWY^{+}20] Q. Li, Z. Wu, L. Yi, K. Seann, H. Qu, and X. Ma. We-Seer: Visual analysis for better information cascade prediction of WeChat articles. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1399–1412, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2022:UUD

[LWY^{+}22] Zhen Li, Xiting Wang, Weikai Yang, Jing Wu, Zhengyan Zhang, Zhiyuan Liu, Maosong Sun, Hui Zhang, and Shixia Liu. A unified understanding of deep NLP models for text classification. IEEE Transactions on Visualization and Computer Graphics, 28(12):4980– 4994, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:KKG

[LWZ^{+}22] Haotian Li, Yong Wang, Songheng Zhang, Yangqiu Song, and Huamin Qu. KG4Vis: a knowledge graph-based approach for visualization recommendation. IEEE Transactions on Visualization and Computer Graphics, 28(1):195–205, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Lin:2022:CAC

[LXF^{+}22] Juncong Lin, Pintong Xiao, Yinan Fu, Yubin Shi, Hongran Wang, Shihui Guo, Ying He, and Tong-Yee Lee. C3 assignment: Camera cubemap color assignment for creative interior design. IEEE Transactions on Visualization and Computer Graphics, 28(8):2895–2908, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:LDT

[LXH^{+}21] Lingjie Liu, Weipeng Xu, Marc Habermann, Michael Zollhöfer, Florian Bernard, Hyeongwoo Kim, Wenping Wang, and Christian Theobalt. Learning dynamic textures for neural rendering of human actors. IEEE Transactions on Visualization and Computer Graphics, 27(10):4009–4022, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2021:VQT

[LXL21] Daoming Liu, Chi Xiong, and Xiaopei Liu. Vectorizing quantum turbulence vortex-core lines for real-time visualization. IEEE Transactions on Visualization and Computer Graphics, 27(9):3794–3807, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:MAM

[LXZ^{+}23] Haomin Liu, Hua Xue, Linsheng Zhao, Danpeng Chen, Zhen Peng, and Guofeng Zhang. MagLoc-AR: Magneticbased localization for visualfree augmented reality in largescale indoor environments. IEEE Transactions on Visualization and Computer Graphics, 29(11):4383–4393, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:GNE

[LY23] Guozheng Li and Xiaoru Yuan.

GoTreeScape: Navigate and explore the tree visualization design space. IEEE Transactions on Visualization and Computer Graphics, 29(12):5451– 5467, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Lin:2023:LVO

[LYBP23] Tica Lin, Yalong Yang, Johanna Beyer, and Hanspeter Pfister. Labeling out-of-view objects in immersive analytics to support situated visual searching. IEEE Transactions on Visualization and Computer Graphics, 29(3):1831– 1844, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2023:NRL

[LYH^{+}23] Tong Liu, Zhenhua Yang, Shaojun Hu, Zhiyi Zhang, Chunxia Xiao, Xiaohu Guo, and Long Yang. Neighbor reweighted local centroid for geometric feature identification. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1545–1558, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Li:2023:FMC

[LYL^{+}23] Sihang Li, Jiacheng Yu, Mingxuan Li, Le Liu, Xiaolong Luke Zhang, and Xiaoru Yuan. A framework for multiclass contour visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):353– 362, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2021:ACL

[LYZ^{+}21] Jianan Li, Jimei Yang, Jianming Zhang, Chang Liu, Christina Wang, and Tingfa Xu. Attribute-conditioned layout GAN for automatic graphic design. IEEE Transactions on Visualization and Computer Graphics, 27(10):4039– 4048, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2024:VSV

[LYZ^{+}24] Hai Li, Xingrui Yang, Hongjia Zhai, Yuqian Liu, Hujun Bao, and Guofeng Zhang. Voxsurf: Voxel-based implicit surface representation. IEEE Transactions on Visualization and Computer Graphics, 30(3): 1743–1755, March 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2021:SIU

[LZ21] H. Liu and H. Zhang. A suggestive interface for untangling mathematical knots. IEEE Transactions on Visualization and Computer Graphics, 27(2): 593–602, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Lekschas:2021:GFL

[LZC^{+}21] F. Lekschas, X. Zhou, W. Chen, N. Gehlenborg, B. Bach, and H. Pfister. A generic framework and library for exploration of small multiples through interactive piling. IEEE Transactions on Visualization and Computer Graphics, 27(2):358– 368, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2020:BSC

[LZD^{+}20] G. Li, Y. Zhang, Y. Dong, J. Liang, J. Zhang, J. Wang, M. J. Mcguffin, and X. Yuan. BarcodeTree: Scalable comparison of multiple hierarchies. IEEE Transactions on Visualization and Computer Graphics, 26(1):1022–1032, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2021:VPT

[LZH^{+}21] Y. Li, X. Zhai, F. Hou, Y. Liu, A. Hao, and H. Qin. Vectorized painting with temporal diffusion curves. IEEE Transactions on Visualization and Computer Graphics, 27(1):228–240, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Li:2020:GEE

[LZJZ20] Z. Li, C. Zhang, S. Jia, and J. Zhang. Galex: Exploring the evolution and intersection of disciplines. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1182–1192, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Latif:2022:KIS

[LZK^{+}22] Shahid Latif, Zheng Zhou, Yoon Kim, Fabian Beck, and Nam Wook Kim. Kori: Interactive synthesis of text and charts in data documents. IEEE Transactions on Visualization and Computer Graphics, 28(1): 184–194, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2023:SBH

[LZL^{+}23] Yanan Liu, Hao Zhang, Yanqiu Li, Kangjian He, and Dan Xu. Skeleton-based human action recognition via largekernel attention graph convolutional network. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2575–2585, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Li:2022:DSG

[LZLS22] Xiaoyu Li, Bo Zhang, Jing Liao, and Pedro V. Sander. Deep sketch-guided cartoon video inbetweening. IEEE Transactions on Visualization and Computer Graphics, 28(8):2938–2952, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Liu:2020:ADV

[LZM20] Z. Liu, S. H. Zhan, and T. Munzner. Aggregated dendrograms for visual comparison between many phylogenetic trees. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2732–2747, September 2020.

CODEN ITVGEA. ISSN 1077-2626.

Luo:2020:DEM

[LZP^{+}20] T. Luo, M. Zhang, Z. Pan, Z. Li, N. Cai, J. Miao, Y. Chen, and M. Xu. Dream-experiment: a MR user interface with natural multi-channel interaction for virtual experiments. IEEE Transactions on Visualization and Computer Graphics, 26(12):3524–3534, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Lan:2023:SVE

[LZW^{+}23] Ji Lan, Zheng Zhou, Jiachen Wang, Hui Zhang, Xiao Xie, and Yingcai Wu. SimuExplorer: Visual exploration of game simulation in table tennis. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1719–1732, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Liu:2021:DDN

[LZX^{+}21] Yilong Liu, Chengwei Zheng, Feng Xu, Xin Tong, and Baining Guo. Data-driven 3D neck modeling and animation. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3226–3237, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Liu:2022:RTS

[LZX^{+}22] Yanli Liu, Xingming Zou, Songhua Xu, Guanyu Xing, Housheng Wei, and Yanci Zhang. Real-time shadow detection from live outdoor videos for augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2748–2763, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Li:2023:INI

[LZY^{+}23] Hai Li, Hongjia Zhai, Xingrui Yang, Zhirong Wu, Yihao Zheng, Haofan Wang, Jianchao Wu, Hujun Bao, and Guofeng Zhang. ImTooth: Neural implicit tooth for dental augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2837–2846, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Liu:2020:DDI

[LZZ^{+}20] M. Liu, K. Zhang, J. Zhu, J. Wang, J. Guo, and Y. Guo. Data-driven indoor scene modeling from a single color image with iterative object segmentation and model retrieval. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1702–1715, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Li:2021:SRR

[LZZ^{+}21] Lei Li, Changqing Zou, Youyi Zheng, Qingkun Su, Hongbo Fu, and Chiew-Lan Tai. Sketch-R2CNN: an RNN-Rasterization-CNN architecture for vector sketch recognition. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3745–3754, September 2021.

CODEN ITVGEA. ISSN 1077-2626.

Liu:2024:SGD

[LZZ^{+}24] Weiying Liu, Yanyan Zhang, Baiqiao Zhang, Qianqian Xiong, Hong Zhao, Sheng Li, Juan Liu, and Yulong Bian. Selfguided DMT: Exploring a novel paradigm of dance movement therapy in mixed reality for children with ASD. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2119–2128, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Monica:2020:SBI

[MA20] R. Monica and J. Aleotti.

Surfel-based incremental reconstruction of the boundary between known and unknown space. IEEE Transactions on Visualization and Computer Graphics, 26(8):2683–2695, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Macedo:2023:OHA

[MA23] Márcio C. F. Macedo and Antônio L. Apolinário. Occlusion handling in augmented reality: Past, present and future. IEEE Transactions on Visualization and Computer Graphics, 29(2):1590–1609, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2020:IIVa

[MB20a] K. Mueller and D. Bowman. Introducing the IEEE Virtual Reality 2020 special issue. IEEE Transactions on Visualization and Computer Graphics, 26(5): iv–v, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2020:IIVb

[MB20b] K. Mueller and D. Bowman. Introducing the IEEE Virtual Reality 2020 special issue. IEEE Transactions on Visualization and Computer Graphics, 26(5): iv–v, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2020:MECb

[MB20c] K. Mueller and D. Bowman. Message from the Editor-in-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 26(12): 3386, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2021:IIV

[MB21a] Klaus Mueller and Doug Bowman. Introducing the IEEE Virtual Reality 2021 special issue. IEEE Transactions on Visualization and Computer Graphics, 27(5):iv, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2021:MECb

[MB21b] Klaus Mueller and Doug Bowman. Message from the Editorin-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 27 (11):4085, November 2021. CO-DEN ITVGEA. ISSN 1077-

Mueller:2022:IVI

[MB22a] Klaus Mueller and Doug Bowman. IEEE VR 2022 introducing the special issue. IEEE Transactions on Visualization and Computer Graphics, 28 (5):vi, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2022:MECb

[MB22b] Klaus Mueller and Doug Bowman. Message from the Editorin-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 28 (11):iv, November 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Marques:2020:SAQ

[MBB20] R. Marques, C. Bouville, and K. Bouatouch. Spectral analysis of quadrature rules and Fourier truncation-based methods applied to shading integrals. IEEE Transactions on Visualization and Computer Graphics, 26(10):3022–3036, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Marques:2021:ESF

[MBBB21] R. Marques, C. Bouville, K. Bouatouch, and J. Blat. Extensible spherical Fibonacci grids. IEEE Transactions on Visualization and Computer Graphics, 27(4):2341– 2354, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Morariu:2023:PUP

[MBC^{+}23] Cristina Morariu, Adrien Bibal, Rene Cutura, Benoît Frénay, and Michael Sedlmair. Predicting user preferences of dimensionality reduction embedding quality. IEEE Transactions on Visualization and Computer Graphics, 29(1):745–755, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Morgenstern:2024:AVH

[MBHE24] Wieland Morgenstern, Milena T. Bagdasarian, Anna Hilsmann, and Peter Eisert. Animatable virtual humans: Learning posedependent human representations in UV space for interactive performance synthesis. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2644–2650, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Meschenmoser:2021:MUV

[MBS^{+}21] P. Meschenmoser, J. F. Buchmüller, D. Seebacher, M. Wikelski, and D. A. Keim. Multi-SegVA: Using visual analytics to segment biologging time series on multiple scales. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1623–1633, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Morth:2023:SIV

[MBS23] Eric Mörth, Stefan Bruckner, and Noeska N. Smit. ScrollyVis: Interactive visual authoring of guided dynamic narratives for scientific scrollytelling. IEEE Transactions on Visualization and Computer Graphics, 29(12):5165– 5177, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Mishra:2024:MDI

[MCFKF24] Shivam Mishra, Missael Corro-Flores, David Krum, and Negin Forouzesh. Molecular docking improved with human spatial perception using virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2269–2275, May 2024. CODEN ITVGEA. ISSN 1077-2626.

McNutt:2023:NGR

[McN23] Andrew M. McNutt. No grammar to rule them all: a survey of JSON-style DSLs for visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):160–170, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Mahmud:2023:VCS

[MCQ23] M. Rasel Mahmud, Alberto Cordova, and John Quarles. Visual cues for a steadier you: Visual feedback methods improved standing balance in virtual reality for people with balance impairments. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4666–4675, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Mimnaugh:2023:VRS

[MCS^{+}23] Katherine J. Mimnaugh, Evan G. Center, Markku Suomalainen, Israel Becerra, Eliezer Lozano, Rafael Murrieta-Cid, Timo Ojala, Steven M. LaValle, and Kara D. Federmeier. Virtual reality sickness reduces attention during immersive experiences. IEEE Transactions on Visualization and Computer Graphics, 29(11):4394–4404, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Moullec:2023:AWP

[MCSAL23] Yann Moullec, Mélanie Cogné, Justine Saint-Aubert, and Anatole Lécuyer. Assisted walkingin-place: Introducing assisted motion to walking-by-cycling in embodied virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2796–2805, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Mei:2020:RDA

[MCW^{+}20] H. Mei, W. Chen, Y. Wei, Y. Hu, S. Zhou, B. Lin, Y. Zhao, and J. Xia. RSATree: Distribution-aware data representation of large-scale tabular datasets for flexible visual query. IEEE Transactions on Visualization and Computer Graphics, 26(1):1161– 1171, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meyer:2020:CRV

[MD20] M. Meyer and J. Dykes. Criteria for rigor in visualization design study. IEEE Transactions on Visualization and Computer Graphics, 26(1):87–97, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mao:2023:PFP

[MDH^{+}23] Aihua Mao, Zihui Du, Junhui Hou, Yaqi Duan, Yong-Jin Liu, and Ying He. PU-Flow: a point cloud upsampling network with normalizing flows. IEEE Transactions on Visualization and Computer Graphics, 29(12):4964–4977, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Meng:2021:KFR

[MDJV21] Xiaoxu Meng, Ruofei Du, Joseph F. JaJa, and Amitabh Varshney. 3D-Kernel foveated rendering for light fields. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3350–3360, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Mao:2023:SNS

[MDL^{+}23] Aihua Mao, Canglan Dai, Qing Liu, Jie Yang, Lin Gao, Ying He, and Yong-Jin Liu. STD-Net: Structure-preserving and topology-adaptive deformation network for single-view 3D reconstruction. IEEE Transactions on Visualization and Computer Graphics, 29(3):

1785–1798, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Miftari:2024:VDV

[MDS24] Egzon Miftari, Daniel Durstewitz, and Filip Sadlo. Visualization of discontinuous vector field topology. IEEE Transactions on Visualization and Computer Graphics, 30(1):45– 54, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Meng:2020:EDG

[MDV20] X. Meng, R. Du, and A. Varshney. Eye-dominance-guided foveated rendering. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1972–1980, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Mao:2023:YLS

[MDX^{+}23] Aihua Mao, Wenbo Dong, Chaoqiang Xie, Huamin Wang, Yong-Jin Liu, Guiqing Li, and Ying He. Yarn-level simulation of hygroscopicity of woven textiles. IEEE Transactions on Visualization and Computer Graphics, 29(12):5250– 5264, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Mori:2020:IIR

[MEB^{+}20] S. Mori, O. Erat, W. Broll, H. Saito, D. Schmalstieg, and D. Kalkofen. InpaintFusion: Incremental RGB-D inpainting for 3D scenes. IEEE Transactions on Visualization and Computer Graphics, 26 (10):2994–3007, October 2020.

CODEN ITVGEA. ISSN 1077-2626.

Morini:2024:SSD

[MEHD24] Francesca Morini, Anna Eschenbacher, Johanna Hartmann, and Marian Dörk. From shock to shift: Data visualization for constructive climate journalism. IEEE Transactions on Visualization and Computer Graphics, 30(1):1413– 1423, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Machuca:2022:IV

[MF22] Mayra Donaji Barrera Machuca and Javad Fotouhi. IEEE VR 2022. IEEE Transactions on Visualization and Computer Graphics, 28(5):xiii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Mikawa:2024:LLO

[MF24] Yuri Mikawa and Taiki Fukiage. Low-latency ocular parallax rendering and investigation of its effect on depth perception in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2228–2238, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Ma:2021:VAF

[MFH^{+}21] Y. Ma, A. Fan, J. He, A. R.

Nelakurthi, and R. Maciejewski. A visual analytics framework for explaining and diagnosing transfer learning processes. IEEE Transactions on Visualization and Computer Graphics, 27(2):1385– 1395, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Mota:2023:CSV

[MFS^{+}23] Roberta Mota, Nivan Ferreira, Julio Daniel Silva, Marius Horga, Marcos Lage, Luis Ceferino, Usman Alim, Ehud Sharlin, and Fabio Miranda. A comparison of spatiotemporal visualizations for 3D urban analytics. IEEE Transactions on Visualization and Computer Graphics, 29(1):1277– 1287, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Monteiro:2021:HFI

[MGC^{+}21] Pedro Monteiro, Guilherme Gonçalves, Hugo Coelho, Miguel Melo, and Maximino Bessa. Hands-free interaction in immersive virtual reality: a systematic review. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2702–2713, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Marchal:2020:P

[MGJ^{+}20] M. Marchal, J. L. Gabbard, J. Jorge, T. W. Kuhlen, and A. Steed. Preface. IEEE Transactions on Visualization and Computer Graphics, 26(5):vi, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Melo:2022:DMS

[MGM^{+}22] Miguel Melo, Guilherme Gonçalves, Pedro Monteiro, Hugo Coelho, José Vasconcelos-Raposo, and Maximino Bessa. Do multisensory stimuli benefit the virtual reality experience? A systematic review. IEEE Transactions on Visualization and Computer Graphics, 28(2):1428– 1442, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Monadjemi:2021:CMI

[MGO21] S. Monadjemi, R. Garnett, and A. Ottley. Competing models: Inferring exploration patterns and information relevance via Bayesian model selection. IEEE Transactions on Visualization and Computer Graphics, 27(2):412–421, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Muller:2021:IDA

[MGU^{+}21] Juliane Muller, Laura Garrison, Philipp Ulbrich, Stefanie Schreiber, Stefan Bruckner, Helwig Hauser, and Steffen Oeltze-Jafra. Integrated dual analysis of quantitative and qualitative high-dimensional data. IEEE Transactions on Visualization and Computer Graphics, 27(6):2953– 2966, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Martin-Gomez:2022:IFC

[MGWK^{+}22] Alejandro Martin-Gomez, Jakob Weiss, Andreas Keller, Ulrich Eck, Daniel Roth, and Nassir Navab. The impact of focus and context visualization techniques on depth perception in optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 28(12):4156– 4171, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Moore:2022:EPI

[MGWM22] Jimmy Moore, Pascal Goffin, Jason Wiese, and Miriah Meyer. Exploring the personal informatics analysis gap: There’s a lot of bacon. IEEE Transactions on Visualization and Computer Graphics, 28(1):96– 106, January 2022. CODEN ITVGEA. ISSN 1077-2626.

McColeman:2021:NMI

[MHFF21] C. M. McColeman, L. Harrison, M. Feng, and S. Franconeri. No mark is an island: Precision and category repulsion biases in data reproductions. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1063–1072, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Miyatake:2023:HVH

[MHIS23] Yamato Miyatake, Takefumi Hiraki, Daisuke Iwai, and Kosuke Sato. HaptoMapping: Visuo-haptic augmented reality by embedding userimperceptible tactile display control signals in a projected image. IEEE Transactions on Visualization and Computer Graphics, 29(4):2005– 2019, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Moreira:2024:UTG

[MHN^{+}24] Gustavo Moreira, Maryam Hosseini, Md Nafiul Alam Nipu, Marcos Lage, Nivan Ferreira, and Fabio Miranda. The urban toolkit: a grammar-based framework for urban visual analytics. IEEE Transactions on Visualization and Computer Graphics, 30(1):1402– 1412, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Maher:2022:FVA

[MHS^{+}22] Kevin Maher, Zeyuan Huang, Jiancheng Song, Xiaoming Deng, Yu-Kun Lai, Cuixia Ma, Hao Wang, Yong-Jin Liu, and Hongan Wang. E-ffective: a visual analytic system for exploring the emotion and effectiveness of inspirational speeches. IEEE Transactions on Visualization and Computer Graphics, 28(1):508–517, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Morais:2022:SDA

[MJAD22] Luiz Morais, Yvonne Jansen, Nazareno Andrade, and Pierre Dragicevic. Showing data about people: a design space of anthropographics. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1661–1679, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Mirzaei:2020:EUE

[MKK20] M. Mirzaei, P. Kán, and H. Kaufmann. EarVR: Using ear haptics in virtual reality for deaf and hard-of-hearing people. IEEE Transactions on Visualization and Computer Graphics, 26(5):2084– 2093, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Magri:2024:GFP

[ML24] Victor A. P. Magri and Peter Lindstrom. A general framework for progressive data compression and retrieval. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1358–1368, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Mumtaz:2020:ECQ

[MLBW20] H. Mumtaz, S. Latif, F. Beck, and D. Weiskopf. Exploranative code quality documents. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1129–1139, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2020:SWS

[MLC^{+}20] G. Ma, S. Li, C. Chen, A. Hao, and H. Qin. Stage-wise salient object detection in 360^{∘} omnidirectional image via object-level semantical saliency ranking. IEEE Transactions on Visualization and Computer Graphics, 26(12):3535–3545, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Maack:2024:PCP

[MLT^{+}24] Robin G. C. Maack, Jonas Lukasczyk, Julien Tierny, Hans Hagen, Ross Maciejewski, and Christoph Garth. Parallel computation of piecewise linear Morse–Smale segmentations.

IEEE Transactions on Visualization and Computer Graphics, 30(4):1942–1955, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Mahmood:2020:TIC

[MM20] S. Mahmood and K. Mueller.

Taxonomizer: Interactive construction of fully labeled hierarchical groupings from attributes of multivariate data. IEEE Transactions on Visualization and Computer Graphics, 26(9):2875–2890, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Ma:2021:VAC

[MM21] Y. Ma and R. Maciejewski. Visual analysis of class separations with locally linear segments. IEEE Transactions on Visualization and Computer Graphics, 27(1):241–253, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Mullen:2023:PDD

[MM23] James F. Mullen and Dinesh Manocha. PACE: Data-driven virtual agent interaction in dense and cluttered environments. IEEE Transactions on Visualization and Computer Graphics, 29(5):2536– 2546, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Ma:2020:DCV

[MMF20] J. Ma, K. Ma, and J. Frazier. Decoding a complex visualization in a science museum: An empirical study. IEEE Transactions on Visualization and Computer Graphics, 26(1): 472–481, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ma:2021:LDL

[MMG^{+}21] Ruixian Ma, Honghui Mei, Huihua Guan, Wei Huang, Fan Zhang, Chengye Xin, Wenzhuo Dai, Xiao Wen, and Wei Chen. LADV: Deep learning assisted authoring of dashboard visualizations from images and sketches. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3717–3732, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Matsumuro:2023:MEV

[MMK^{+}23] Miki Matsumuro, Shohei Mori, Yuta Kataoka, Fumiaki Igarashi, Fumihisa Shibata, and Asako Kimura. Modified egocentric viewpoint for softer seated experience in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2230–2238, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Mercado:2021:ETI

[MML21] V. Mercado, M. Marchal, and A. Lécuyer. ENTROPiA: Towards infinite surface haptic displays in virtual reality using encountered-type rotating props. IEEE Transactions on Visualization and Computer Graphics, 27(3):2237– 2243, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Medeiros:2022:SAP

[MMN^{+}22] Daniel Medeiros, Mark McGill, Alexander Ng, Robert Mc-Dermid, Nadia Pantidi, Julie Williamson, and Stephen Brewster. From shielding to avoidance: Passenger augmented reality and the layout of virtual displays for productivity in shared transit. IEEE Transactions on Visualization and Computer Graphics, 28(11): 3640–3650, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Malpica:2023:TDV

[MMS^{+}23] Sandra Malpica, Daniel Martin, Ana Serrano, Diego Gutierrez, and Belen Masia. Taskdependent visual behavior in immersive environments: a comparative study of free exploration, memory and visual search. IEEE Transactions on Visualization and Computer Graphics, 29(11):4417– 4425, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Meuschke:2023:GGC

[MNB^{+}23] Monique Meuschke, Uli Niemann, Benjamin Behrendt, Matthias Gutberlet, Bernhard Preim, and Kai Lawonn. GUCCI — guided cardiac cohort investigation of blood flow data. IEEE Transactions on Visualization and Computer Graphics, 29(3):1876– 1892, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Mizuho:2023:EVF

[MNK23] Takato Mizuho, Takuji Narumi, and Hideaki Kuzuoka. Effects of the visual fidelity of virtual environments on presence, context-dependent forgetting, and source-monitoring error. IEEE Transactions on Visualization and Computer Graphics, 29(5):2607–2614, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Mancinelli:2023:BSI

[MNPP23] Claudio Mancinelli, Giacomo Nazzaro, Fabio Pellacini, and Enrico Puppo. b/Surf: Interactive Bézier splines on surface meshes. IEEE Transactions on Visualization and Computer Graphics, 29(7):3419– 3435, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Ma:2020:EVS

[MNZ^{+}20] T. Ma, Y. Nie, Q. Zhang, Z. Zhang, H. Sun, and G. Li. Effective video stabilization via joint trajectory smoothing and frame warping. IEEE Transactions on Visualization and Computer Graphics, 26(11): 3163–3176, November 2020. CODEN ITVGEA. ISSN 1077-2626.

May:2021:VFI

[MOA21] Chris May, Manuel M. Oliveira, and Daniel Aliaga. Video folding: Increased framerate for semi-repetitive sequences.

IEEE Transactions on Visualization and Computer Graphics, 27(10):3900–3912, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Marrinan:2021:RTO

[MP21] Thomas Marrinan and Michael E.

Papka. Real-time omnidirectional stereo rendering: Generating 360^{∘} surround-view panoramic images for comfortable immersive viewing. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2587–2596, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Moritz:2024:AEL

[MPNF24] Dominik Moritz, Lace M.

Padilla, Francis Nguyen, and Steven L. Franconeri. Average estimates in line graphs are biased toward areas of higher variability. IEEE Transactions on Visualization and Computer Graphics, 30(1):306–315, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Mazza:2021:HEV

[MPV21] S. Mazza, D. Patel, and I. Viola. Homomorphic-encrypted volume rendering. IEEE Transactions on Visualization and Computer Graphics, 27(2):635– 644, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Molik:2022:MLI

[MPWN22] Ladislav Molík, Václav Pavlovec, Hsiang-Yun Wu, and Martin Nöllenburg. Mixed labeling:

Integrating internal and external labels. IEEE Transactions on Visualization and Computer Graphics, 28(4):1848– 1861, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Morse:2022:EVD

[MRS22] Peter E. Morse, Anya M. Reading, and Tobias Stål. Exploratory volumetric deep earth visualization by 2.5D interactive compositing. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2641–2653, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Marques:2022:CMT

[MSA^{+}22] Bernardo Marques, Samuel Silva, João Alves, Tiago Araújo, Paulo Dias, and Beatriz Sousa Santos. A conceptual model and taxonomy for collaborative augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(12): 5113–5133, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Moullec:2022:MSD

[MSAM^{+}22] Yann Moullec, Justine Saint-Aubert, Julien Manson, Mélanie Cogné, and Anatole Lécuyer. Multi-sensory display of selfavatar’s physiological state: virtual breathing and heart beating can increase sensation of effort in VR. IEEE Transactions on Visualization and Computer Graphics, 28(11):3596– 3606, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Martin:2022:SGM

[MSB^{+}22] Daniel Martin, Ana Serrano, Alexander W. Bergman, Gordon Wetzstein, and Belen Masia. ScanGAN360: a generative model of realistic scanpaths for 360^{∘} images. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2003–2013, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Muller-Sielaff:2023:VAD

[MSBV^{+}23] Juliane Müller-Sielaff, Seyed Behnam Beladi, Stephanie W. Vrede, Monique Meuschke, Peter J. F. Lucas, Johanna M. A. Pijnenborg, and Steffen Oeltze-Jafra. Visual assistance in development and validation of Bayesian networks for clinical decision support. IEEE Transactions on Visualization and Computer Graphics, 29(8):3602–3616, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Martin:2023:SCB

[MSGM23] Daniel Martin, Xin Sun, Diego Gutierrez, and Belen Masia. A study of change blindness in immersive environments. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2446–2455, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Miyashita:2023:PAI

[MSK23a] Yamato Miyashita, Yasuhito Sawahata, and Kazuteru Komine. Perceptual assessment of image and depth quality of dynamically depth-compressed scene for automultiscopic 3D display. IEEE Transactions on Visualization and Computer Graphics, 29(6):3067–3080, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Mori:2023:EBI

[MSK23b] Shohei Mori, Dieter Schmalstieg, and Denis Kalkofen. Exemplar-based inpainting for 6DOF virtual reality photos. IEEE Transactions on Visualization and Computer Graphics, 29(11):4644–4654, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Mori:2023:GKI

[MSK23c] Shohei Mori, Dieter Schmalstieg, and Denis Kalkofen. Good keyframes to inpaint. IEEE Transactions on Visualization and Computer Graphics, 29(9):3989–4000, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Magallanes:2022:SCM

[MSM^{+}22] Jessica Magallanes, Tony Stone, Paul D. Morris, Suzanne Mason, Steven Wood, and Maria-Cruz Villa-Uriol. Sequen-C: a multilevel overview of temporal event sequences. IEEE Transactions on Visualization and Computer Graphics, 28(1): 901–911, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Mantri:2023:HDV

[MSMX23] Prateek Mantri, Hariharan Subramonyam, Audrey L.

Michal, and Cindy Xiong. How do viewers synthesize conflicting information from data visualizations? IEEE Transactions on Visualization and Computer Graphics, 29(1):1005– 1015, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Medeiros:2020:MCI

[MSRJ20] D. Medeiros, M. Sousa, A. Raposo, and J. Jorge. Magic Carpet: Interaction fidelity for flying in VR. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2793–2804, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Meulemans:2021:SPC

[MSS21] W. Meulemans, M. Sondag, and B. Speckmann. A simple pipeline for coherent grid maps. IEEE Transactions on Visualization and Computer Graphics, 27(2):1236–1246, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Morrical:2023:QCG

[MSuG^{+}23] Nate Morrical, Alper Sahistan, U ur Güdükbay, Ingo Wald, and Valerio Pascucci. Quick clusters: a GPU-Parallel partitioning for efficient path tracing of unstructured volumetric grids. IEEE Transactions on Visualization and Computer Graphics, 29(1):537–547, January 2023. CODEN ITVGEA. ISSN 1077-

Mikawa:2022:DPM

[MSWI22] Yuri Mikawa, Tomohiro Sueishi, Yoshihiro Watanabe, and Masatoshi Ishikawa. Dynamic projection mapping for robust sphere posture tracking using uniform/ biased circumferential markers. IEEE Transactions on Visualization and Computer Graphics, 28(12):4016–4031, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Mallett:2022:PTH

[MSY22] Ian Mallett, Larry Seiler, and Cem Yuksel. Patch textures: Hardware support for mesh colors. IEEE Transactions on Visualization and Computer Graphics, 28(7):2710– 2721, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Marquardt:2020:CNV

[MTE^{+}20] A. Marquardt, C. Trepkowski, T. D. Eibich, J. Maiero, E. Kruijff, and J. Schöning. Comparing non-visual and visual guidance methods for narrow field of view augmented reality displays. IEEE Transactions on Visualization and Computer Graphics, 26(12): 3389–3401, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Matthews:2023:ART

[MTVS23] Brandon J. Matthews, Bruce H. Thomas, G. Stewart Von Itzstein, and Ross T. Smith. Adaptive reset techniques for haptic retargeted interaction.

IEEE Transactions on Visualization and Computer Graphics, 29(2):1478–1490, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Ma:2020:SDS

[MTW^{+}20] Y. Ma, A. K. H. Tung, W. Wang, X. Gao, Z. Pan, and W. Chen. ScatterNet: a deep subjective similarity model for visual analysis of scatterplots. IEEE Transactions on Visualization and Computer Graphics, 26(3):1562–1576, March 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:2020:EN

[Mue20a] K. Mueller. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2135–2141, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2020:MECa

[Mue20b] K. Mueller. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1637, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:2020:SJ

[Mue20c] K. Mueller. State of the journal.

IEEE Transactions on Visualization and Computer Graphics, 26(3):1440–1441, March 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Mueller:2021:MECa

[Mue21a] K. Mueller. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 27(2): xvi, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2021:EN

[Mue21b] Klaus Mueller. Editor’s note. IEEE Transactions on Visualization and Computer Graphics, 27(12):4342–4346, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2022:E

[Mue22a] Klaus Mueller. Editorial. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1714, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2022:MECa

[Mue22b] Klaus Mueller. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 28(1): xii, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Mueller:2023:FNE

[Mue23a] Klaus Mueller. Farewell and new EIC introduction. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1299–1300, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Mueller:2023:MEC

[Mue23b] Klaus Mueller. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 29(1): xiv, January 2023. CODEN ITVGEA. ISSN 1077-2626.

McDonald:2021:IUV

[MUM^{+}21] T. McDonald, W. Usher, N. Morrical, A. Gyulassy, S. Petruzza, F. Federer, A. Angelucci, and V. Pascucci. Improving the usability of virtual reality neuron tracing with topological elements. IEEE Transactions on Visualization and Computer Graphics, 27(2): 744–754, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Meng:2024:CCS

[MvdEPV24] Linhao Meng, Stef van den Elzen, Nicola Pezzotti, and Anna Vilanova. Class-constrained t-SNE: Combining data features and class probabilities. IEEE Transactions on Visualization and Computer Graphics, 30(1):164–174, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Mal:2023:IAE

[MWD^{+}23] David Mal, Erik Wolf, Nina Döllinger, Carolin Wienrich, and Marc Erich Latoschik. The impact of avatar and environment congruence on plausibility, embodiment, presence, and the Proteus effect in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5):2358– 2368, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Meinecke:2022:ESS

[MWJ22] Christofer Meinecke, David Joseph Wrisley, and Stefan Jänicke. Explaining semi-supervised text alignment through visualization. IEEE Transactions on Visualization and Computer Graphics, 28(12):4797– 4809, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Mehringer:2022:SOV

[MWR^{+}22] Wolfgang Mehringer, Markus Wirth, Daniel Roth, Georg Michelson, and Bjoern M. Eskofier. Stereopsis only: Validation of a monocular depth cues reduced gamified virtual reality with reaction time measurement. IEEE Transactions on Visualization and Computer Graphics, 28(5):2114– 2124, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Morrical:2022:AUM

[MWUP22] Nate Morrical, Ingo Wald, Will Usher, and Valerio Pascucci. Accelerating unstructured mesh point location with RT cores. IEEE Transactions on Visualization and Computer Graphics, 28(8):2852–2866, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Ming:2020:PSD

[MXC^{+}20] Y. Ming, P. Xu, F. Cheng, H. Qu, and L. Ren. Proto-Steer: Steering deep sequence model with prototypes. IEEE Transactions on Visualization and Computer Graphics, 26(1):

238–248, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Montazeri:2021:MAM

[MXF^{+}21] Z. Montazeri, C. Xiao, Y. Fei, C. Zheng, and S. Zhao. Mechanics-aware modeling of cloth appearance. IEEE Transactions on Visualization and Computer Graphics, 27(1):137– 150, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Ma:2020:EVA

[MXLM20] Y. Ma, T. Xie, J. Li, and R. Maciejewski. Explaining vulnerabilities to adversarial machine learning through visual analytics. IEEE Transactions on Visualization and Computer Graphics, 26(1):1075– 1085, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Meng:2022:GTC

[MXT^{+}22] Wenlong Meng, Shiqing Xin, Changhe Tu, Shuangmin Chen, Ying He, and Wenping Wang. Geodesic tracks: Computing discrete geodesics with trackbased Steiner point propagation. IEEE Transactions on Visualization and Computer Graphics, 28(12):4887– 4901, December 2022. CODEN ITVGEA. ISSN 1077-2626.

McColeman:2022:RRV

[MYBF22] Caitlyn M. McColeman, Fumeng Yang, Timothy F. Brady, and Steven Franconeri. Rethinking the ranks of visual channels. IEEE Transactions on Visualization and Computer Graphics, 28(1):707–717, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Mukherjee:2022:CMT

[MYS^{+}22] Kushin Mukherjee, Brian Yin, Brianne E. Sherman, Laurent Lessard, and Karen B. Schloss. Context matters: a theory of semantic discriminability for perceptual encoding systems. IEEE Transactions on Visualization and Computer Graphics, 28(1):697–706, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Morrical:2024:AAR

[MZS^{+}24] Nate Morrical, Stefan Zellmann, Alper Sahistan, Patrick Shriwise, and Valerio Pascucci. Attribute-aware RBFs: Interactive visualization of time series particle volumes using RT core range queries. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1150–1160, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Mao:2021:ASP

[MZX^{+}21] A. Mao, H. Zhang, Z. Xie, M. Yu, Y.-J. Liu, and Y. He. Automatic sitting pose generation for ergonomic ratings of chairs. IEEE Transactions on Visualization and Computer Graphics, 27(3):1890– 1903, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Nedel:2022:IVM

[NAW^{+}22] Luciana Nedel, Ferran Argelaguet, Lili Wang, Jeanine Stefannuci, and Daisuke Iwai. IEEE VR 2022 message from the journal paper chairs and guest editors. IEEE Transactions on Visualization and Computer Graphics, 28(5):vii, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Nowak:2024:DAV

[NB24] Stan Nowak and Lyn Bartram.

Designing for ambiguity in visual analytics: Lessons from risk assessment and prediction. IEEE Transactions on Visualization and Computer Graphics, 30(1):924–933, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Norouzi:2021:VAD

[NBE^{+}21] Nahal Norouzi, Gerd Bruder, Austin Erickson, Kangsoo Kim, Jeremy Bailenson, Pamela Wisniewski, Charlie Hughes, and Greg Welch. Virtual animals as diegetic attention guidance mechanisms in 360-degree experiences. IEEE Transactions on Visualization and Computer Graphics, 27(11):4321– 4331, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Nguyen:2023:FNO

[NBE^{+}23] Ngan Nguyen, Ciril Bohak, Dominik Engel, Peter Mindek, Ond ej Strnad, Peter Wonka, Sai Li, Timo Ropinski, and Ivan Viola. Finding Nano-Ötzi: Cryo-electron tomography visualization guided by learned segmentation. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4198–4214, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Nguyen:2021:VHP

[NBJ^{+}21] H. T. Nguyen, A. Bhatele, N. Jain, S. P. Kesavan, H. Bhatia, T. Gamblin, K.-L. Ma, and P.-T. Bremer. Visualizing hierarchical performance profiles of parallel codes using CallFlow. IEEE Transactions on Visualization and Computer Graphics, 27(4):2455–2468, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Naik:2020:AVE

[NBNC20] H. Naik, R. Bastien, N. Navab, and I. D. Couzin. Animals in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2073–2083, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Nardini:2021:TEC

[NCB^{+}21] P. Nardini, M. Chen, R. Bujack, M. Bottinger, and G. Scheuermann. A testing environment for continuous colormaps. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1043–1053, February 2021. CO-DEN ITVGEA. ISSN 1077-

Newburger:2023:FBC

[NCE23] Eric Newburger, Michael Correll, and Niklas Elmqvist. Fitting bell curves to data distributions using visualization. IEEE Transactions on Visualization and Computer Graphics, 29(12):5372–5383, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Nardini:2021:MCC

[NCS^{+}21] Pascal Nardini, Min Chen, Francesca Samsel, Roxana Bujack, Michael Böttinger, and Gerik Scheuermann. The making of continuous colormaps. IEEE Transactions on Visualization and Computer Graphics, 27(6):3048–3063, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Narechania:2022:LIA

[NCWE22] Arpit Narechania, Adam Coscia, Emily Wall, and Alex Endert. Lumos: Increasing awareness of analytic behavior during visual data analysis. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1009–1018, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Nonato:2021:GLG

[NdCS21] Luis Gustavo Nonato, Fabiano Petronetto do Carmo, and Claudio T. Silva. GLoG: Laplacian of Gaussian for spatial pattern detection in spatiotemporal data. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3481–3492, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Nehme:2021:VQM

[NDF^{+}21] Y. Nehmé, F. Dupont, J.-P. Farrugia, P. Le Callet, and G. Lavoué. Visual quality of 3D meshes with diffuse colors in virtual reality: Subjective and objective evaluation. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2202–2219, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Nie:2020:AMV

[NDLW20] G. Nie, H. B. Duh, Y. Liu, and Y. Wang. Analysis on mitigation of visually induced motion sickness by applying dynamical blurring on a user’s retina. IEEE Transactions on Visualization and Computer Graphics, 26(8):2535–2545, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Natsukawa:2021:VAA

[NDP^{+}21] H. Natsukawa, E. R. Deyle, G. M. Pao, K. Koyamada, and G. Sugihara. A visual analytics approach for ecosystem dynamics based on empirical dynamic modeling. IEEE Transactions on Visualization and Computer Graphics, 27(2):506– 516, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Newburger:2024:VAS

[NE24] Eric Newburger and Niklas Elmqvist. Visualization according to statisticians: an interview study on the role of visualization for inferential statistics. IEEE Transactions on Visualization and Computer Graphics, 30(1):230–239, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Nothelfer:2020:MBD

[NF20] C. Nothelfer and S. Franconeri.

Measures of the benefit of direct encoding of data deltas for data pair relation perception. IEEE Transactions on Visualization and Computer Graphics, 26(1): 311–320, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nipu:2023:VAD

[NFN^{+}23] Nafiul Nipu, Carla Floricel, Negar Naghashzadeh, Roberto Paoli, and G. Elisabeta Marai. Visual analysis and detection of contrails in aircraft engine simulations. IEEE Transactions on Visualization and Computer Graphics, 29(1):798–808, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Nunez-Garcia:2020:FQC

[NGBA^{+}20] M. Nuñez-Garcia, G. Bernardino, F. Alarcón, G. Caixal, L. Mont, O. Camara, and C. Butakoff. Fast quasi-conformal regional flattening of the left atrium.

IEEE Transactions on Visualization and Computer Graphics, 26(8):2591–2602, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Nair:2024:BOE

[NGW^{+}24] Vivek Nair, Wenbo Guo, Rui Wang, James F. O’Brien, Louis Rosenberg, and Dawn Song. Berkeley Open Extended Reality Recordings 2023 (BOXRR-23): 4.7 million motion capture recordings from 105,000 XR users. IEEE Transactions on Visualization and Computer Graphics, 30(5):2239– 2246, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Nguyen:2020:VHU

[NHC^{+}20] P. H. Nguyen, R. Henkin, S. Chen, N. Andrienko, G. Andrienko, O. Thonnard, and C. Turkay. VASABI: Hierarchical user profiles for interactive visual user behaviour analytics. IEEE Transactions on Visualization and Computer Graphics, 26(1):77–86, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nabata:2022:AIS

[NI22] Kosuke Nabata and Kei Iwasaki. Adaptive irradiance sampling for many-light rendering of subsurface scattering. IEEE Transactions on Visualization and Computer Graphics, 28(10): 3324–3335, October 2022. CO-DEN ITVGEA. ISSN 1077-

Nam:2024:VME

[NIK24] Jung Who Nam, Tobias Isenberg, and Daniel F. Keefe. V-Mail: 3d-enabled correspondence about spatial data on (almost) all your devices. IEEE Transactions on Visualization and Computer Graphics, 30(4): 1853–1867, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Nakano:2021:HMD

[NIM^{+}21] Kizashi Nakano, Naoya Isoyama, Diego Monteiro, Nobuchika Sakata, Kiyoshi Kiyokawa, and Takuji Narumi. Head-mounted display with increased downward field of view improves presence and sense of selflocation. IEEE Transactions on Visualization and Computer Graphics, 27(11):4204– 4214, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Narechania:2022:VPS

[NKWW22] Arpit Narechania, Alireza Karduni, Ryan Wesslen, and Emily Wall. VITALITY: Promoting serendipitous discovery of academic literature with transformers & visual analytics. IEEE Transactions on Visualization and Computer Graphics, 28(1):486–496, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Novotny:2024:ETR

[NL24] Johannes Novotny and David H.

Laidlaw. Evaluating text reading speed in VR scenes and 3D particle visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2602–2612, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Nomoto:2022:DMP

[NLPW22] Takashi Nomoto, Wanlong Li, Hao-Lun Peng, and Yoshihiro Watanabe. Dynamic multiprojection mapping based on parallel intensity control. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2125–2134, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Nguyen:2021:VFM

[NMC21] D. B. Nguyen, R. O. Monico, and G. Chen. A visualization framework for multi-scale coherent structures in Taylor– Couette turbulence. IEEE Transactions on Visualization and Computer Graphics, 27(2): 902–912, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Narciso:2023:UHR

[NMR^{+}23] David Narciso, Miguel Melo, Susana Rodrigues, João Paulo Cunha, José Vasconcelos-Raposo, and Maximino Bessa. Using heart rate variability for comparing the effectiveness of virtual vs real training environments for firefighters. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3238–3250, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Nakamura:2023:AED

[NMS^{+}23] Fumihiko Nakamura, Masaaki Murakami, Katsuhiro Suzuki, Masaaki Fukuoka, Katsutoshi Masai, and Maki Sugimoto. Analyzing the effect of diverse gaze and head direction on facial expression recognition with photo-reflective sensors embedded in a head-mounted display. IEEE Transactions on Visualization and Computer Graphics, 29(10):4124–4139, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Nsonga:2020:DVS

[NNF^{+}20] B. Nsonga, M. Niemann, J. Fröhlich, J. Staib, S. Gumhold, and G. Scheuermann. Detection and visualization of splat and antisplat events in turbulent flows. IEEE Transactions on Visualization and Computer Graphics, 26(11):3147– 3162, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Neto:2021:EMV

[NP21] M. P. Neto and F. V.

Paulovich. Explainable matrixvisualization for global and local interpretability of random forest classification ensembles. IEEE Transactions on Visualization and Computer Graphics, 27(2):1427–1437, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Neto:2024:MDE

[NP24a] Mário Popolin Neto and Fernando V. Paulovich. Multivariate data explanation by jumping emerging patterns visualization. IEEE Transactions on Visualization and Computer Graphics, 30(2):1549– 1563, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Ning:2024:TEA

[NP24b] Bing Ning and Mingtao Pei. Task and environment-aware virtual scene rearrangement for enhanced safety in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2517– 2526, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Narechania:2021:SVD

[NQE21] A. Narechania, A. Qamar, and A. Endert. SafetyLens: Visual data analysis of functional safety of vehicles. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1688–1697, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Neuroth:2023:LSR

[NRA^{+}23] Tyson Neuroth, Martin Rieth, Konduri Aditya, Myoungkyu Lee, Jacqueline H Chen, and Kwan-Liu Ma. Level set restricted Voronoi tessellation for large scale spatial statistical analysis. IEEE Transactions on Visualization and Computer Graphics, 29(1):548–558, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Nsonga:2020:ANW

[NSG^{+}20] B. Nsonga, G. Scheuermann, S. Gumhold, J. Ventosa-Molina, D. Koschichow, and J. Fröhlich. Analysis of the near-wall flow in a turbine cascade by splat visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):719–728, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Nguyen:2021:MTC

[NSK^{+}21] N. Nguyen, O. Strnad, T. Klein, D. Luo, R. Alharbi, P. Wonka, M. Maritan, P. Mindek, L. Autin, D. S. Goodsell, and I. Viola. Modeling in the time of COVID-19: Statistical and rule-based mesoscale models. IEEE Transactions on Visualization and Computer Graphics, 27(2):722–732, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Nickel:2022:MOD

[NSM^{+}22] Soeren Nickel, Max Sondag, Wouter Meulemans, Stephen Kobourov, Jaakko Peltonen, and Martin Nöllenburg. Multicriteria optimization for dynamic demers cartograms. IEEE Transactions on Visualization and Computer Graphics, 28(6):2376–2387, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Narechania:2021:NTG

[NSS21] A. Narechania, A. Srinivasan, and J. Stasko. NL4DV: a toolkit for generating analytic specifications for data visualization from natural language queries. IEEE Transactions on Visualization and Computer Graphics, 27(2):369–379, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Neuhauser:2024:ASS

[NSW24] Christoph Neuhauser, Josef Stumpfegger, and Rüdiger Westermann. Adaptive sampling of 3D spatial correlations for Focus+Context visualization. IEEE Transactions on Visualization and Computer Graphics, 30(2):1608– 1623, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Nguyen-Vo:2021:NNL

[NVRS^{+}21] T. Nguyen-Vo, B. E. Riecke, W. Stuerzlinger, D.-M. Pham, and E. Kruijff. NaviBoard and NaviChair: Limited translation combined with full rotation for efficient virtual locomotion. IEEE Transactions on Visualization and Computer Graphics, 27(1):165–177, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Nguyen:2023:DMD

[NWMC23] Duong B. Nguyen, Panruo Wu, Rodolfo Ostilla Monico, and Guoning Chen. Dynamic mode decomposition for largescale coherent structure extraction in shear flows. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1531–1544, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Neuhauser:2021:IFR

[NWW21] Christoph Neuhauser, Junpeng Wang, and Rüdiger Westermann. Interactive Focus+Context rendering for hexahedral mesh inspection. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3505–3518, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Offenwanger:2024:TSB

[OBCT24] Anna Offenwanger, Matthew Brehmer, Fanny Chevalier, and Theophanis Tsandilas. Time-Splines: Sketch-based authoring of flexible and idiosyncratic timelines. IEEE Transactions on Visualization and Computer Graphics, 30(1):34–44, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Ono:2021:PVA

[OCL^{+}21] J. P. Ono, S. Castelo, R. Lopez, E. Bertini, J. Freire, and C. Silva. PipelineProfiler: a visual analytics tool for the exploration of AutoML pipelines. IEEE Transactions on Visualization and Computer Graphics, 27(2):390–400, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Oral:2024:ICC

[OCW^{+}24] Emre Oral, Ria Chawla, Michel Wijkstra, Narges Mahyar, and Evanthia Dimara. From information to choice: a critical inquiry into visualization tools for decision making. IEEE Transactions on Visualization and Computer Graphics, 30(1):359– 369, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Otaran:2022:HAP

[OF22] Ata Otaran and Ildar Farkhatdinov. Haptic ankle platform for interactive walking in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):3974– 3985, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Offenhuber:2020:DPM

[Off20] D. Offenhuber. Data by proxy material traces as autographic visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(1):98– 108, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ogawa:2022:IV

[OK22] Nami Ogawa and Katharina Krösl. IEEE VR 2022. IEEE Transactions on Visualization and Computer Graphics, 28 (5):xiv, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Oppermann:2021:VCT

[OKM21] M. Oppermann, R. Kincaid, and T. Munzner. VizCommender: Computing text-based similarity in visualization repositories for content-based recommendations. IEEE Transactions on Visualization and Computer Graphics, 27(2):495– 505, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Oppermann:2022:VCV

[OM22] Michael Oppermann and Tamara Munzner. VizSnippets: Compressing visualization bundles into representative previews for browsing visualization collections. IEEE Transactions on Visualization and Computer Graphics, 28(1):747–757, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Ogawa:2021:EAA

[ONH21] Nami Ogawa, Takuji Narumi, and Michitaka Hirose. Effect of avatar appearance on detection thresholds for remapped hand movements. IEEE Transactions on Visualization and Computer Graphics, 27(7):3182– 3197, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Oagaz:2022:PIS

[OSC22] Hawkar Oagaz, Breawn Schoun, and Min-Hyung Choi. Performance improvement and skill transfer in table tennis through training in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):4332–4343, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Ortner:2021:IID

[OWN^{+}21] T. Ortner, A. Walch, R. Nowak, R. Barnes, T. Höllt, and M. E. Gröller. InCorr: Interactive data-driven correlation panels for digital outcrop analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2):755–764, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Ouyang:2024:LHM

[OWW^{+}24] Yang Ouyang, Yuchen Wu, He Wang, Chenyang Zhang, Furui Cheng, Chang Jiang, Lixia Jin, Yuanwu Cao, and Quan Li. Leveraging historical medical records as a proxy via multimodal modeling and visualization to enrich medical diagnostic learning. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1238–1248, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Ondov:2021:RPP

[OYK^{+}21] B. D. Ondov, F. Yang, M. Kay, N. Elmqvist, and S. Franconeri. Revealing perceptual proxies with adversarial examples. IEEE Transactions on Visualization and Computer Graphics, 27(2):1073– 1083, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Oliver:2024:SHV

[OZZ24] Peter Oliver, Eugene Zhang, and Yue Zhang. Scalable hypergraph visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1): 595–605, January 2024. CO-DEN ITVGEA. ISSN 1077-

Pintore:2022:IAE

[PAAG22] Giovanni Pintore, Marco Agus, Eva Almansa, and Enrico Gobbetti. Instant automatic emptying of panoramic indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 28(11):3629–3639, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Park:2020:HTP

[PALW20] G. Park, A. Argyros, J. Lee, and W. Woo. 3D hand tracking in the presence of excessive motion blur. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1891–1901, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Pena-Araya:2020:CVI

[PAPB20] V. Peña-Araya, E. Pietriga, and A. Bezerianos. A comparison of visualizations for identifying correlation over space and time. IEEE Transactions on Visualization and Computer Graphics, 26(1):375–385, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Parsons:2022:UDV

[Par22] Paul Parsons. Understanding data visualization design practice. IEEE Transactions on Visualization and Computer Graphics, 28(1):665–675, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Patane:2022:MAH

[Pat22] Giuseppe Patané. Meshless approximation and helmholtzhodge decomposition of vector fields. IEEE Transactions on Visualization and Computer Graphics, 28(2):1328– 1341, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Pintore:2023:DSS

[PBAG23] Giovanni Pintore, Fabio Bettio, Marco Agus, and Enrico Gobbetti. Deep scene synthesis of atlanta-world interiors from a single omnidirectional image. IEEE Transactions on Visualization and Computer Graphics, 29(11):4708–4718, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Palenik:2020:SSS

[PBBH20] J. Pálenik, J. Byška, S. Bruckner, and H. Hauser. Scale-space splatting: Reforming spacetime for cross-scale exploration of integral measures in molecular dynamics. IEEE Transactions on Visualization and Computer Graphics, 26(1):643–653, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pister:2021:IPK

[PBF^{+}21] A. Pister, P. Buono, J.-D. Fekete, C. Plaisant, and P. Valdivia. Integrating prior knowledge in mixed-initiative social network clustering. IEEE Transactions on Visualization and Computer Graphics, 27(2):

1775–1785, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Piccolotto:2024:DTA

[PBM^{+}24] Nikolaus Piccolotto, Markus Bögl, Christoph Muehlmann, Klaus Nordhausen, Peter Filzmoser, Johanna Schmidt, and Silvia Miksch. Data type agnostic visual sensitivity analysis. IEEE Transactions on Visualization and Computer Graphics, 30(1):1106–1116, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Pavlovec:2022:RLP

[PC22] Vaclav Pavlovec and Ladislav Cmolik. Rapid labels: Pointfeature labeling on GPU. IEEE Transactions on Visualization and Computer Graphics, 28(1): 604–613, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Pugh:2023:GPS

[PCJ23] Brian Pugh, Davin Chernak, and Salma Jiddi. GeoSynth: a photorealistic synthetic indoor dataset for scene understanding. IEEE Transactions on Visualization and Computer Graphics, 29(5):2586– 2595, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Park:2024:RTV

[PCL24] Jieun Park, Youjin Choi, and Kyung Myun Lee. Research trends in virtual reality music concert technology: a systematic literature review. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2195–2205, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Padilla:2020:TOE

[PCQ^{+}20] L. M. K. Padilla, S. C. Castro, P. S. Quinan, I. T. Ruginski, and S. H. Creem-Regehr. Toward objective evaluation of working memory in visualizations: a case study using pupillometry and a dual-task paradigm. IEEE Transactions on Visualization and Computer Graphics, 26(1):332–342, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pan:2021:EBL

[PCZ^{+}21] J. Pan, W. Chen, X. Zhao, S. Zhou, W. Zeng, M. Zhu, J. Chen, S. Fu, and Y. Wu. Exemplar-based layout finetuning for node-link diagrams. IEEE Transactions on Visualization and Computer Graphics, 27(2):1655–1665, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Piochowiak:2024:FCS

[PD24] Max Piochowiak and Carsten Dachsbacher. Fast compressed segmentation volumes for scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):12– 22, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Park:2022:NSA

[PDD^{+}22] Haekyu Park, Nilaksh Das, Rahul Duggal, Austin P. Wright, Omar Shaikh, Fred Hohman, and Duen Horng Polo Chau. NeuroCartography: Scalable automatic visual summarization of concepts in deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 28(1):813– 823, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Pahins:2020:RTE

[PFC20] C. A. L. Pahins, N. Ferreira, and J. L. Comba. Real-time exploration of large spatiotemporal datasets based on order statistics. IEEE Transactions on Visualization and Computer Graphics, 26(11):3314– 3326, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Padilla:2023:MFV

[PFCB23] Lace Padilla, Racquel Fygenson, Spencer C. Castro, and Enrico Bertini. Multiple forecast visualizations (MFVs): Tradeoffs in trust and performance in multiple COVID-19 forecast visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):12– 22, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Park:2022:DAP

[PFN22] Hyungman Park, Donald Fussell, and Paul Navrátil. Dataaware predictive scheduling for distributed-memory ray tracing. IEEE Transactions on Visualization and Computer Graphics, 28(1):1172– 1181, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Peck:2022:ETP

[PGE^{+}22] Tabitha C. Peck, Jessica J. Good, Austin Erickson, Isaac Bynum, and Gerd Bruder. Effects of transparency on perceived humanness: Implications for rendering skin tones using optical see-through displays. IEEE Transactions on Visualization and Computer Graphics, 28(5):2179–2189, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Petruzza:2020:HTF

[PGL^{+}20] S. Petruzza, A. Gyulassy, S. Leventhal, J. J. Baglino, M. Czabaj, A. D. Spear, and V. Pascucci. High-throughput feature extraction for measuring attributes of deforming open-cell foams. IEEE Transactions on Visualization and Computer Graphics, 26(1):140– 150, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peck:2021:ERB

[PGS21] Tabitha C. Peck, Jessica J. Good, and Katharina Seitz. Evidence of racial bias using immersive virtual reality: Analysis of head and hand motions during shooting decisions. IEEE Transactions on Visualization and Computer Graphics, 27(5):2502–2512, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Porssut:2022:AVE

[PHB^{+}22] Thibault Porssut, Yawen Hou, Olaf Blanke, Bruno Herbelin, and Ronan Boulic. Adapting virtual embodiment through reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3193–3205, September 2022. CODEN ITVGEA. ISSN 1077-2626. Punpongsanon:2020:FVM

[PIS20] P. Punpongsanon, D. Iwai, and K. Sato. FleXeen: Visually manipulating perceived fabric bending stiffness in spatial augmented reality. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1433–1439, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Polk:2020:CGA

[PJHY20] T. Polk, D. Jäckle, J. Häußler, and J. Yang. CourtTime: Generating actionable insights into tennis matches using visual analytics. IEEE Transactions on Visualization and Computer Graphics, 26(1):397–406, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Park:2021:IPT

[PJYW21] Jinwoo Park, Ik-Beom Jeon, Sung-Eui Yoon, and Woontack Woo. Instant panoramic texture mapping with semantic object matching for large-scale urban scene reproduction. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2746–2756, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Pascucci:2021:MVG

[PK21] V. Pascucci and M. Kirby. Message from VIS 2020 General Chairs. IEEE Transactions on Visualization and Computer Graphics, 27(2):xvii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Patel:2021:VST

[PL21] Mohak Patel and David H. Laidlaw. Visualization of 3D stress tensor fields using superquadric glyphs on displacement streamlines. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3264–3276, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Plaisant:2021:VCA

[Pla21] C. Plaisant. The 2020 visualization career award. IEEE Transactions on Visualization and Computer Graphics, 27(2): xxviii, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Parmar:2023:HIS

[PLD^{+}23] Dhaval Parmar, Lorraine Lin, Nikeetha DSouza, Sophie Jörg, Alison E. Leonard, Shaundra B. Daily, and Sabarish V. Babu. How immersion and self-avatars in VR affect learning programming and computational thinking in middle school education. IEEE Transactions on Visualization and Computer Graphics, 29(8):3698–3713, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Pan:2024:DDG

[PLL^{+}24] Bo Pan, Jiaying Lu, Haoxuan Li, Weifeng Chen, Yiyao Wang, Minfeng Zhu, Chenhao Yu, and Wei Chen. Differentiable design galleries: a differentiable approach to explore the design space of transfer functions. IEEE Transactions on Visualization and Computer Graphics, 30(1):1369–1379, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Panagiotidou:2023:CUD

[PLP^{+}23] Georgia Panagiotidou, Houda Lamqaddam, Jeroen Poblome, Koenraad Brosens, Katrien Verbert, and Andrew Vande Moere. Communicating uncertainty in digital humanities visualization research. IEEE Transactions on Visualization and Computer Graphics, 29(1): 635–645, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Pandey:2023:GRS

[PLW^{+}23] Aditeya Pandey, Sehi L’Yi, Qianwen Wang, Michelle A. Borkin, and Nils Gehlenborg. GenoREC: a recommendation system for interactive genomics data visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):570– 580, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Preston:2023:CUR

[PM23] Annie Preston and Kwan-Liu Ma. Communicating uncertainty and risk in air quality maps. IEEE Transactions on Visualization and Computer Graphics, 29(9):3746– 3757, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Perez-Messina:2024:GVA

[PMCM24] Ignacio Pérez-Messina, Davide Ceneda, and Silvia Miksch. Guided visual analytics for image selection in time and space. IEEE Transactions on Visualization and Computer Graphics, 30(1):66–75, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Pillette:2023:SRN

[PMN^{+}23] Léa Pillette, Guillaume Moreau, Jean-Marie Normand, Manon Perrier, Anatole Lécuyer, and Mélanie Cogné. A systematic review of navigation assistance systems for people with dementia. IEEE Transactions on Visualization and Computer Graphics, 29(4):2146– 2165, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Procopio:2022:ICB

[PMS^{+}22] Marianne Procopio, Ab Mosca, Carlos Scheidegger, Eugene Wu, and Remco Chang. Impact of cognitive biases on progressive visualization. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3093–3112, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Prinz:2023:SLR

[PMW23] Lisa Marie Prinz, Tintu Mathew, and Benjamin Weyers. A systematic literature review of virtual reality locomotion taxonomies. IEEE Transactions on Visualization and Computer Graphics, 29(12):5208– 5223, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Park:2021:CCA

[PNB^{+}21] Ji Hwan Park, Saad Nadeem, Saeed Boorboor, Joseph Marino, and Arie Kaufman. CMed: Crowd analytics for medical imaging data. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2869–2880, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Park:2021:HVA

[PNKC21] H. Park, Y. Nam, J.-H. Kim, and J. Choo. HyperTendril: Visual analytics for user-driven hyperparameter optimization of deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1407–1416, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Patnaik:2023:SSP

[PPE23] Biswaksen Patnaik, Huaishu Peng, and Niklas Elmqvist.

Sensemaking sans power: Interactive data visualization using color-changing ink. IEEE Transactions on Visualization and Computer Graphics, 29(12):5282–5293, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Park:2020:PID

[PPYW20] J. Park, H. Park, S. Yoon, and W. Woo. Physically-inspired deep light estimation from a homogeneous-material object for mixed reality lighting. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2002–2011, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Patil:2023:SED

[PRJ^{+}23] Ameya Patil, Gaëlle Richer, Christopher Jermaine, Dominik Moritz, and Jean-Daniel Fekete. Studying early decision making with progressive bar charts. IEEE Transactions on Visualization and Computer Graphics, 29(1):407–417, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Perroud:2020:ARV

[PRKM20] B. Perroud, S. Régnier, A. Kemeny, and F. Mérienne. Application of a relative visual performance model in a virtual reality immersive system. IEEE Transactions on Visualization and Computer Graphics, 26(10):3128–3132, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Pu:2022:MVE

[PSG^{+}22] Jiansu Pu, Hui Shao, Boyang Gao, Zhengguo Zhu, Yanlin Zhu, Yunbo Rao, and Yong Xiang. matExplorer: Visual exploration on predicting ionic conductivity for solid-state electrolytes. IEEE Transactions on Visualization and Computer Graphics, 28(1):65–75, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Peck:2020:MGU

[PSH20] T. C. Peck, L. E. Sockol, and S. M. Hancock. Mind the gap: The underrepresentation of female participants and authors in virtual reality research. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1945–1954, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Palenik:2021:IVG

[PSH21] J. Palenik, T. Spengler, and H. Hauser. IsoTrotter: Visually guided empirical modelling of atmospheric convection. IEEE Transactions on Visualization and Computer Graphics, 27(2): 775–784, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Perez:2023:MCN

[PSL23] Daniel Perez, Yuzhong Shen, and Jiang Li. Mesh convolutional networks with face and vertex feature operators. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1678–1690, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Pandey:2022:SAS

[PSS^{+}22] Aditeya Pandey, Uzma Haque Syeda, Chaitya Shah, John A. Guerra-Gomez, and Michelle A. Borkin. A state-of-the-art survey of tasks for tree design and evaluation with a curated task dataset. IEEE Transactions on Visualization and Computer Graphics, 28 (10):3563–3584, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Pandey:2023:MIB

[PSS23] Aditeya Pandey, Arjun Srinivasan, and Vidya Setlur. Medley: Intent-based recommendations to support dashboard composition. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1135–1145, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Pandey:2020:CDA

[PSY^{+}20] A. Pandey, H. Shukla, G. S. Young, L. Qin, A. A. Zamani, L. Hsu, R. Huang, C. Dunne, and M. A. Borkin. CerebroVis: Designing an abstract yet spatially contextualized cerebral artery network visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):938–948, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Peck:2020:ISA

[PT20] T. C. Peck and A. Tutar.

The impact of a self-avatar, hand collocation, and hand proximity on embodiment and Stroop interference. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1964–1971, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Pan:2024:ETA

[PTC^{+}24] Ye Pan, Shuai Tan, Shengran Cheng, Qunfen Lin, Zijiao Zeng, and Kenny Mitchell. Expressive talking avatars. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2538–2548, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Pan:2021:CBV

[PTD^{+}21] X. Pan, F. Tang, W. Dong, C. Ma, Y. Meng, F. Huang, T.- Y. Lee, and C. Xu. Contentbased visual summarization for image collections. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2298–2312, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Pezzotti:2020:GLC

[PTM^{+}20] N. Pezzotti, J. Thijssen, A. Mordvintsev, T. Höllt, B. Van Lew, B. P. F. Lelieveldt, E. Eisemann, and A. Vilanova. GPGPU linear complexity t-SNE optimization. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1172–1181, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Pfluger:2020:VVE

[PTS^{+}20] H. Pflüger, D. Thom, A. Schütz, D. Bohde, and T. Ertl. VeCHArt: Visually enhanced comparison of historic art using an automated line-based synchronization technique. IEEE Transactions on Visualization and Computer Graphics, 26 (10):3063–3076, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Parger:2022:PUU

[PTX^{+}22] Mathias Parger, Chengcheng Tang, Yuanlu Xu, Christopher D. Twigg, Lingling Tao, Yijing Li, Robert Wang, and Markus Steinberger. UNOC: Understanding occlusion for embodied presence in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12):4240– 4251, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Pont:2022:WDG

[PVDT22] Mathieu Pont, Jules Vidal, Julie Delon, and Julien Tierny. Wasserstein distances, geodesics and barycenters of merge trees. IEEE Transactions on Visualization and Computer Graphics, 28(1):291–301, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Panagiotidou:2022:IEU

[PVPM22] Georgia Panagiotidou, Ralf Vandam, Jeroen Poblome, and Andrew Vande Moere. Implicit error, uncertainty and confidence in visualization: an archaeological case study. IEEE Transactions on Visualization and Computer Graphics, 28(12):4389–4402, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Prasad:2024:TPF

[PvSvdE^{+}24] Vidya Prasad, Ruud J. G. van Sloun, Stef van den Elzen, Anna Vilanova, and Nicola Pezzotti. The transform-and-perform framework: Explainable deep learning beyond classification. IEEE Transactions on Visualization and Computer Graphics, 30(2):1502– 1515, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Pont:2023:PGA

[PVT23] Mathieu Pont, Jules Vidal, and Julien Tierny. Principal geodesic analysis of merge trees (and persistence diagrams). IEEE Transactions on Visualization and Computer Graphics, 29(2):1573–1589, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Perovich:2021:CCD

[PWB21] L. J. Perovich, S. A. Wylie, and R. Bongiovanni. Chemicals in the creek: designing a situated data physicalization of open government data with the community. IEEE Transactions on Visualization and Computer Graphics, 27(2):913– 923, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Preuss:2021:DPA

[PWK21] Daniel Preuß, Tino Weinkauf, and Jens Krüger. A discrete probabilistic approach to dense flow visualization. IEEE Transactions on Visualization and Computer Graphics, 27(12): 4347–4358, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Pi:2021:VCA

[PYSJ21] M. Pi, H. Yeon, H. Son, and Y. Jang. Visual cause analytics for traffic congestion. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2186–2201, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Pan:2023:EVP

[PZC^{+}23] Ye Pan, Ruisi Zhang, Shengran Cheng, Shuai Tan, Yu Ding, Kenny Mitchell, and Xubo Yang. Emotional voice puppetry. IEEE Transactions on Visualization and Computer Graphics, 29(5):2527– 2535, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Qi:2020:SVT

[QBW^{+}20] J. Qi, V. Bloemen, S. Wang, J. van Wijk, and H. van de Wetering. STBins: Visual tracking and comparison of multiple data sequences using temporal binning. IEEE Transactions on Visualization and Computer Graphics, 26(1):

1054–1063, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Qu:2023:LIA

[QCCC23] Qiang Qu, Xiaoming Chen, Yuk Ying Chung, and Weidong Cai. LFACon: Introducing anglewise attention to no-reference quality assessment in light field space. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2239–2248, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Qiao:2023:DOG

[QCL23] Xiaotian Qiao, Ying Cao, and Rynson W. H. Lau. Design order guided visual note layout optimization. IEEE Transactions on Visualization and Computer Graphics, 29 (9):3922–3936, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Qin:2022:DOA

[QFWS22] Yu Qin, Brittany Terese Fasy, Carola Wenk, and Brian Summa. A domain-oblivious approach for learning concise representations of filtered topological spaces for clustering. IEEE Transactions on Visualization and Computer Graphics, 28(1):302–312, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Qiao:2023:LBI

[QGL^{+}23] Yi-Ling Qiao, Lin Gao, Shu-Zhi Liu, Ligang Liu, Yu-Kun Lai, and Xilin Chen. Learning-based intrinsic reflectional symmetry detection. IEEE Transactions on Visualization and Computer Graphics, 29(9):3799– 3808, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Qiao:2022:LML

[QGY^{+}22] Yi-Ling Qiao, Lin Gao, Jie Yang, Paul L. Rosin, Yu-Kun Lai, and Xilin Chen. Learning on 3D meshes with Laplacian encoding and pooling. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1317–1327, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Qin:2020:MDT

[QHL^{+}20] H. Qin, J. Han, N. Li, H. Huang, and B. Chen. Massdriven topology-aware curve skeleton extraction from incomplete point clouds. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2805–2817, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Qu:2024:NNN

[QLC^{+}24] Qiang Qu, Hanxue Liang, Xiaoming Chen, Yuk Ying Chung, and Yiran Shen. NeRF-NQA: No-reference quality assessment for scenes generated by NeRF and neural view synthesis methods. IEEE Transactions on Visualization and Computer Graphics, 30(5):2129– 2139, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Qiao:2022:SMD

[QLFG22] Yi-Ling Qiao, Yu-Kun Lai, Hongbo Fu, and Lin Gao. Synthesizing mesh deformation sequences with bidirectional LSTM. IEEE Transactions on Visualization and Computer Graphics, 28(4):1906– 1916, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Quadri:2023:ASD

[QNWR23] Ghulam Jilani Quadri, Jennifer Adorno Nieves, Brenton M. Wiernik, and Paul Rosen. Automatic scatterplot design optimization for clustering identification. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4312–4327, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Quadri:2021:MIV

[QR21] G. J. Quadri and P. Rosen.

Modeling the influence of visual density on cluster perception in scatterplots using topology. IEEE Transactions on Visualization and Computer Graphics, 27(2):1829–1839, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Quadri:2022:SPB

[QR22] Ghulam Jilani Quadri and Paul Rosen. A survey of perceptionbased visualization studies by task. IEEE Transactions on Visualization and Computer Graphics, 28(12):5026– 5048, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Qu:2021:MSS

[QRZZ21] B. Qu, L. Roy, Y. Zhang, and E. Zhang. Mode surfaces of symmetric tensor fields: Topological analysis and seamless extraction. IEEE Transactions on Visualization and Computer Graphics, 27(2):583– 592, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Qian:2021:RTA

[QSC^{+}21] C. Qian, S. Sun, W. Cui, J.-G. Lou, H. Zhang, and D. Zhang. Retrieve-Then-Adapt: Example-based automatic generation for proportionrelated infographics. IEEE Transactions on Visualization and Computer Graphics, 27(2): 443–452, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Qian:2022:ALM

[QSUK22] Long Qian, Tianyu Song, Mathias Unberath, and Peter Kazanzides. AR-Loupe: Magnified augmented reality by combining an optical seethrough head-mounted display and a loupe. IEEE Transactions on Visualization and Computer Graphics, 28(7):2550– 2562, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Qiu:2024:DVA

[QTW^{+}24] Rui Qiu, Yamei Tu, Yu-Shuen Wang, Po-Yin Yen, and Han-Wei Shen. DocFlow: a visual analytics system for questionbased document retrieval and categorization. IEEE Transactions on Visualization and Computer Graphics, 30(2): 1533–1548, February 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Qu:2022:APL

[QZZ22] Botong Qu, Eugene Zhang, and Yue Zhang. Automatic polygon layout for primal-dual visualization of hypergraphs. IEEE Transactions on Visualization and Computer Graphics, 28(1): 633–642, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Ramamurthi:2022:TSM

[RAC22] Yashwanth Ramamurthi, Tripti Agarwal, and Amit Chattopadhyay. A topological similarity measure between multiresolution Reeb spaces. IEEE Transactions on Visualization and Computer Graphics, 28(12):4360–4374, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Reyes-Aviles:2023:BWA

[RAFSA23a] Fernando Reyes-Aviles, Philipp Fleck, Dieter Schmalstieg, and Clemens Arth. Bag of world anchors for instant large-scale localization. IEEE Transactions on Visualization and Computer Graphics, 29(11):4730– 4739, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Reyes-Aviles:2023:CWA

[RAFSA23b] Fernando Reyes-Aviles, Philipp Fleck, Dieter Schmalstieg, and Clemens Arth. Compact world anchors: Registration using parametric primitives as scene description. IEEE Transactions on Visualization and Computer Graphics, 29(10):4140– 4153, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Romat:2021:EAN

[RAP21] H. Romat, C. Appert, and E. Pietriga. Expressive authoring of node-link diagrams with graphies. IEEE Transactions on Visualization and Computer Graphics, 27(4):2329– 2340, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Rydow:2023:DET

[RBF^{+}23] Erik Rydow, Rita Borgo, Hui Fang, Thomas Torsney-Weir, Ben Swallow, Thibaud Porphyre, Cagatay Turkay, and Min Chen. Development and evaluation of two approaches of visual sensitivity analysis to support epidemiological modeling. IEEE Transactions on Visualization and Computer Graphics, 29(1):1255– 1265, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Randhavane:2021:MDD

[RBK^{+}21] Tanmay Randhavane, Aniket Bera, Emily Kubin, Kurt Gray, and Dinesh Manocha. Modeling data-driven dominance traits for virtual characters using gait analysis. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2967–2979, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Rauscher:2024:SVE

[RBKM24] Julius Rauscher, Raphael Buchmüller, Daniel A. Keim, and Matthias Miller. SkiVis: Visual exploration and route planning in ski resorts. IEEE Transactions on Visualization and Computer Graphics, 30(1):869–879, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Rahimi:2022:NMS

[RBLE22] Farzan Baradaran Rahimi, Jeffrey E. Boyd, Richard M. Levy, and Jennifer Eiserman. New media and space: an empirical study of learning and enjoyment through museum hybrid space. IEEE Transactions on Visualization and Computer Graphics, 28(8):3013–3021, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Ruddle:2022:EAP

[RBLT^{+}22] Roy A. Ruddle, Jürgen Bernard, Hendrik Lücke-Tieke, Thorsten May, and Jörn Kohlhammer. The effect of alignment on people’s ability to judge event sequence similarity. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3070–3081, September 2022. CODEN ITVGEA. ISSN 1077-

Rahimi:2020:STT

[RBR20] K. Rahimi, C. Banigan, and E. D. Ragan. Scene transitions and teleportation in virtual reality and the implications for spatial awareness and sickness. IEEE Transactions on Visualization and Computer Graphics, 26(6):2273–2287, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Reimann:2021:VMF

[RBRG21] Daniel Reimann, Christine Blech, Nilam Ram, and Robert Gaschler. Visual model fit estimation in scatterplots: Influence of amount and decentering of noise. IEEE Transactions on Visualization and Computer Graphics, 27(9):3834– 3838, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Roberts:2022:EJV

[RBSN22] Jonathan C. Roberts, Peter Butcher, Ann Sherlock, and Sarah Nason. Explanatory journeys: Visualising to understand and explain administrative justice paths of redress. IEEE Transactions on Visualization and Computer Graphics, 28(1): 518–528, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Raynor:2023:SAB

[RCD^{+}23] Justin Raynor, Tarik Crnovrsanin, Sara Di Bartolomeo, Laura South, David Saffo, and Cody Dunne. The state of the art in BGP visualization tools: a mapping of visualization techniques to cyberattack types. IEEE Transactions on Visualization and Computer Graphics, 29(1):1059–1069, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Rakotondravony:2023:PWL

[RDH23] Noëlle Rakotondravony, Yiren Ding, and Lane Harrison. Probablement, wahrscheinlich, likely? A cross-language study of how people verbalize probabilities in icon array visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):1189– 1199, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Raji:2021:DSI

[RDHH21] Mohammad Raji, Jeremiah Duncan, Tanner Hobson, and Jian Huang. Dataless sharing of interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 27(9):3656–3669, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Reda:2023:RCW

[Red23] Khairi Reda. Rainbow colormaps: What are they good and bad for? IEEE Transactions on Visualization and Computer Graphics, 29(12): 5496–5510, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Reipschlager:2021:PAR

[RFD21] P. Reipschlager, T. Flemisch, and R. Dachselt. Personal augmented reality for information visualization on large interactive displays. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1182–1192, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Rojo:2020:VFT

[RG20] I. B. Rojo and T. Günther. Vector field topology of timedependent flows in a steady reference frame. IEEE Transactions on Visualization and Computer Graphics, 26(1):280– 290, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Russig:2023:TAV

[RGDG23] Benjamin Russig, David Graß, Raimund Dachselt, and Stefan Gumhold. On-tube attribute visualization for multivariate trajectory data. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1288–1298, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Rojo:2020:AMC

[RGG20a] I. B. Rojo, M. Gross, and T. Günther. Accelerated Monte Carlo rendering of finite-time Lyapunov exponents. IEEE Transactions on Visualization and Computer Graphics, 26(1):

708–718, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rojo:2020:FOO

[RGG20b] I. B. Rojo, M. Gross, and T. Günther. Fourier opacity optimization for scalable exploration. IEEE Transactions on Visualization and Computer Graphics, 26(11):3204– 3216, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Raji:2020:SVM

[RHHH20] M. Raji, A. Hota, T. Hobson, and J. Huang. Scientific visualization as a microservice. IEEE Transactions on Visualization and Computer Graphics, 26(4): 1760–1774, April 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ren:2022:IEM

[RHLC22] Bo Ren, Wei He, Chen-Feng Li, and Xu Chen. Incompressibility enforcement for multiplefluid SPH using deformation gradient. IEEE Transactions on Visualization and Computer Graphics, 28(10):3417– 3427, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Roth:2020:CVE

[RL20] D. Roth and M. E. Latoschik.

Construction of the Virtual Embodiment Questionnaire (VEQ). IEEE Transactions on Visualization and Computer Graphics, 26(12):3546– 3556, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Rees:2021:AVA

[RLB^{+}21] Dylan Rees, Robert S. Laramee, Paul Brookes, Tony D’Cruze, Gary A. Smith, and Aslam Miah. AgentVis: Visual analysis of agent behavior with hierarchical glyphs. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3626–3643, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Romano:2023:MAQ

[RLG^{+}23] Sara Romano, Enricoandrea Laviola, Michele Gattullo, Michele Fiorentino, and Antonio Emmanuele Uva. More arrows in the quiver: Investigating the use of auxiliary models to localize in-view components with augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4483–4493, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Rathore:2020:ECC

[RLLS20] R. Rathore, Z. Leggon, L. Lessard, and K. B. Schloss. Estimating color-concept associations from image statistics. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1226–1235, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rautek:2021:OOR

[RMB^{+}21] P. Rautek, M. Mlejnek, J. Beyer, J. Troidl, H. Pfister, T. Theußl, and M. Hadwiger. Objective observer-relative flow visualization in curved spaces for unsteady 2D geophysical flows. IEEE Transactions on Visualization and Computer Graphics, 27(2):283–293, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Raikwar:2024:BWO

[RMW^{+}24] Aditya Raikwar, Domenick Mifsud, Christopher D. Wickens, Anil Ufuk Batmaz, Amelia C. Warden, Brendan Kelley, Benjamin A. Clegg, and Francisco R. Ortega. Beyond the Wizard of Oz: Negative effects of imperfect machine learning to examine the impact of reliability of augmented reality cues on visual search performance. IEEE Transactions on Visualization and Computer Graphics, 30(5):2662–2670, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Robles:2022:VRB

[RNO^{+}22] Marta Robles, Negar Namdarian, Julia Otto, Evelyn Wassiljew, Nassir Navab, Christine Falter-Wagner, and Daniel Roth. A virtual reality based system for the screening and classification of autism. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2168–2178, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Ratzenbock:2023:UTI

[ROM^{+}23] Sebastian Ratzenböck, Verena Obermüller, Torsten Möller, João Alves, and Immanuel M. Bomze. Uncover: Toward interpretable models for detecting new star cluster members. IEEE Transactions on Visualization and Computer Graphics, 29(9):3855–3872, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Rapp:2020:VCS

[RPD20] T. Rapp, C. Peters, and C. Dachsbacher. Void-and-cluster sampling of large scattered data and trajectories. IEEE Transactions on Visualization and Computer Graphics, 26(1):780–789, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Rapp:2021:VAL

[RPD21] T. Rapp, C. Peters, and C. Dachsbacher. Visual analysis of large multivariate scattered data using clustering and probabilistic summaries. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1580–1590, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Rapp:2022:IBV

[RPD22] Tobias Rapp, Christoph Peters, and Carsten Dachsbacher. Image-based visualization of large volumetric data using moments. IEEE Transactions on Visualization and Computer Graphics, 28(6):2314– 2325, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Rodriguez-Pardo:2023:NPG

[RPG23a] Carlos Rodriguez-Pardo and Elena Garces. Neural photometryguided visual attribute transfer. IEEE Transactions on Visualization and Computer Graphics, 29(3):1818–1830, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Rodriguez-Pardo:2023:SSS

[RPG23b] Carlos Rodriguez-Pardo and Elena Garces. SeamlessGAN: Self-supervised synthesis of tileable texture maps. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2914–2925, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Rogers:2021:IER

[RPH^{+}21] J. Rogers, A. H. Patton, L. Harmon, A. Lex, and M. Meyer. Insights from experiments with rigor in an EvoBio design study. IEEE Transactions on Visualization and Computer Graphics, 27(2):1106–1116, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Reckziegel:2020:MHH

[RPHJ20] M. Reckziegel, L. Pfeiffer, C. Heine, and S. Jänicke. Modeling how humans judge dot-label relations in point cloud visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(6):

2144–2155, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Resck:2023:LEI

[RPNP23] Lucas E. Resck, Jean R. Ponciano, Luis Gustavo Nonato, and Jorge Poco. LegalVis: Exploring and inferring precedent citations in legal documents. IEEE Transactions on Visualization and Computer Graphics, 29(6):3105–3120, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Rosen:2021:LAF

[RQ21] P. Rosen and G. J. Quadri. LineSmooth: an analytical framework for evaluating the effectiveness of smoothing techniques on line charts. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1536–1546, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Rieger:2023:HMS

[RR23] Marc Bastian Rieger and Björn Risch. How to maximise spatial presence: Design guidelines for a virtual learning environment for school use. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2517–2526, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Reimann:2023:LHA

[RRG23] Daniel Reimann, Nilam Ram, and Robert Gaschler. Lollipops help align visual and statistical fit estimates in scatterplots with nonlinear models. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3436–3440, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Robotham:2022:CDI

[RRK^{+}22] Thomas Robotham, Olli S.

Rummukainen, Miriam Kurz, Marie Eckert, and Emanuël A. P. Habets. Comparing direct and indirect methods of audio quality evaluation in virtual reality scenes of varying complexity. IEEE Transactions on Visualization and Computer Graphics, 28(5):2091– 2101, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Reda:2021:RRM

[RS21] K. Reda and D. A. Szafir.

Rainbows revisited: Modeling effective colormap design for graphical inference. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1032–1042, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Radu:2023:HAR

[RS23] Iulian Radu and Bertrand Schneider. How augmented reality (AR) can help and hinder collaborative learning: a study of AR in electromagnetism education. IEEE Transactions on Visualization and Computer Graphics, 29(9):3734– 3745, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Ratcliff:2020:THM

[RSAA20] J. Ratcliff, A. Supikov, S. Alfaro, and R. Azuma. ThinVR: Heterogeneous microlens arrays for compact, 180 degree FOV VR near-eye displays. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1981–1990, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Rodrigues:2023:RDP

[RSD^{+}23] Nils Rodrigues, Christoph Schulz, Sören Döring, Daniel Baumgartner, Tim Krake, and Daniel Weiskopf. Relaxed dot plots: Faithful visualization of samples and their distribution. IEEE Transactions on Visualization and Computer Graphics, 29(1):278–287, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Reimann:2023:CEL

[RSRG23] Daniel Reimann, André Schulz, Nilam Ram, and Robert Gaschler. Color-encoded links improve homophily perception in node-link diagrams. IEEE Transactions on Visualization and Computer Graphics, 29(12):5593–5598, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Rhee:2020:AVT

[RTM^{+}20] T. Rhee, S. Thompson, D. Medeiros, R. dos Anjos, and A. Chalmers. Augmented virtual teleportation for high-fidelity telecollaboration. IEEE Transactions on Visualization and Computer Graphics, 26(5):1923– 1933, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Rittenbruch:2022:CSM

[RVB^{+}22] Markus Rittenbruch, Kellie Vella, Margot Brereton, James M. Hogan, Daniel Johnson, Julian Heinrich, and Sean O’Donoghue. Collaborative sense-making in genomic research: The role of visualisation. IEEE Transactions on Visualization and Computer Graphics, 28(12):4477– 4489, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Rottmann:2023:MES

[RWB^{+}23] Peter Rottmann, Markus Wallinger, Annika Bonerath, Sven Gedicke, Martin Nöllenburg, and Jan-Henrik Haunert. MosaicSets: Embedding set systems into grid graphs. IEEE Transactions on Visualization and Computer Graphics, 29(1):875–885, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Ruan:2023:VVA

[RWJ^{+}23] Shaolun Ruan, Yong Wang, Weiwen Jiang, Ying Mao, and Qiang Guan. VACSEN: a visualization approach for noise awareness in quantum computing. IEEE Transactions on Visualization and Computer Graphics, 29(1):462–472, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Ren:2021:HSH

[RXX^{+}21] J. Ren, W. Xiang, Y. Xiao, R. Yang, D. Manocha, and X. Jin. Heter-Sim: Heterogeneous multi-agent systems simulation by interactive datadriven optimization. IEEE Transactions on Visualization and Computer Graphics, 27(3): 1953–1966, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Ren:2023:LAI

[RZH^{+}23] Yiming Ren, Chengfeng Zhao, Yannan He, Peishan Cong, Han Liang, Jingyi Yu, Lan Xu, and Yuexin Ma. LiDAR-aid inertial poser: Large-scale human motion capture by sparse inertial and LiDAR sensors. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2337–2347, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Ren:2024:EHT

[RZLX24] Yunlei Ren, Yan Zhang, Zhitao Liu, and Ning Xie. Eye-hand typing: Eye gaze assisted finger typing via Bayesian processes in AR. IEEE Transactions on Visualization and Computer Graphics, 30(5):2496– 2506, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Rautek:2024:VLI

[RZW^{+}24] Peter Rautek, Xingdi Zhang, Bernhard Woschizka, Thomas Theußl, and Markus Hadwiger. Vortex lens: Interactive vortex core line extraction using observed line integral convolution. IEEE Transactions on Visualization and Computer Graphics, 30(1):55–65, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Sakano:2022:CMV

[SA22] Yuichi Sakano and Hiroshi Ando. Conditions of a multiview 3D display for accurate reproduction of perceived glossiness. IEEE Transactions on Visualization and Computer Graphics, 28(10):3336– 3350, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Sun:2020:VVR

[SAB20] Q. Sun, M. Azmandian, and J. Baumeister. The 2019 VGTC Virtual Reality Best Dissertation Award. IEEE Transactions on Visualization and Computer Graphics, 26(5):xii–xiii, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Saint-Aubert:2023:IUP

[SACB^{+}23] Justine Saint-Aubert, Mélanie Cogné, Isabelle Bonan, Yoann Launey, and Anatole Lécuyer. Influence of user posture and virtual exercise on impression of locomotion during VR observation. IEEE Transactions on Visualization and Computer Graphics, 29(8):3507–3518, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Simonetto:2020:EBD

[SAK20] P. Simonetto, D. Archambault, and S. Kobourov. Event-based dynamic graph visualisation. IEEE Transactions on Visualization and Computer Graphics, 26(7):2373–2386, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Saint-Aubert:2023:EVI

[SAMB^{+}23] Justine Saint-Aubert, Julien Manson, Isabelle Bonan, Yoann Launey, Anatole Lécuyer, and Mélanie Cogné. Effect of vibrations on impression of walking and embodiment with first- and third-person avatar. IEEE Transactions on Visualization and Computer Graphics, 29(12):5579–5585, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Santini:2020:ARI

[San20] S. Santini. Assigning rated items to locations in nonlist display layouts. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1278–1291, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Swift:2023:VPT

[SAPW23] Melissa E. Swift, Wyatt Ayers, Sophie Pallanck, and Scott Wehrwein. Visualizing the passage of time with video temporal pyramids. IEEE Transactions on Visualization and Computer Graphics, 29(1):171– 181, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Splechtna:2023:IVA

[SaT^{+}23] Rainer Splechtna, Denis Gra anin, Goran Todorovi , Stanislav Goja, Boris Bedi , Helwig Hauser, and Kre imir Matkovi . Interactive visual analysis of structure-borne noise data. IEEE Transactions on Visualization and Computer Graphics, 29(1):778–787, January 2023. CODEN ITVGEA. ISSN 1077-2626.

South:2023:PAI

[SB23] Laura South and Michelle A.

Borkin. Photosensitive accessibility for interactive data visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(1):374–384, January 2023. CODEN ITVGEA. ISSN 1077-2626.

So:2021:HVA

[SBe^{+}21] W. So, E. P. Bogucka, S. epanovi , S. Joglekar, K. Zhou, and D. Quercia. Humane visual AI: Telling the stories behind a medical condition. IEEE Transactions on Visualization and Computer Graphics, 27(2):678– 688, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Spjut:2020:TSC

[SBK^{+}20] J. Spjut, B. Boudaoud, J. Kim, T. Greer, R. Albert, M. Stengel, K. Akşit, and D. Luebke. Toward standardized classification of foveated displays. IEEE Transactions on Visualization and Computer Graphics, 26(5):

2126–2134, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Schafer:2024:IIV

[SBk^{+}24] Marco Schäfer, Nicolas Brich, Jan By ka, Sérgio M. Marques, David Bedná , Philipp Thiel, Barbora Kozlíková, and Michael Krone. InVADo: Interactive visual analysis of molecular docking data. IEEE Transactions on Visualization and Computer Graphics, 30(4): 1984–1997, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Sakin:2023:TNT

[SBT^{+}23] Sayef Azad Sakin, Alex Bigelow, R. Tohid, Connor Scully-Allison, Carlos Scheidegger, Steven R. Brandt, Christopher Taylor, Kevin A. Huck, Hartmut Kaiser, and Katherine E. Isaacs. Traveler: Navigating task parallel traces for performance analysis. IEEE Transactions on Visualization and Computer Graphics, 29(1):788– 797, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Skarbez:2021:ICR

[SBW21] Richard Skarbez, Frederick P. Brooks, and Mary C. Whitton. Immersion and coherence: Research agenda and early results. IEEE Transactions on Visualization and Computer Graphics, 27(10):3839–3850, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Shen:2022:PCD

[SC22] I-Chao Shen and Bing-Yu Chen.

ClipGen: a deep generative model for clipart vectorization and synthesis. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4211–4224, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Sun:2023:ECD

[SCC^{+}23] Mengdi Sun, Ligan Cai, Weiwei Cui, Yanqiu Wu, Yang Shi, and Nan Cao. Erato: Cooperative data story editing via fact interpolation. IEEE Transactions on Visualization and Computer Graphics, 29(1):983–993, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2024:SGE

[SCC^{+}24] Yang Shi, Bingchang Chen, Ying Chen, Zhuochen Jin, Ke Xu, Xiaohan Jiao, Tian Gao, and Nan Cao. Supporting guided exploratory visual analysis on time series data with reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1172–1182, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Sperrle:2023:LPF

[SCEA23] Fabian Sperrle, Davide Ceneda, and Mennatallah El-Assady. Lotse: a practical framework for guidance in visual analytics. IEEE Transactions on Visualization and Computer Graphics, 29(1):1124–1134, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Shin:2023:SDP

[SCHE23] Sungbok Shin, Sunghyo Chung, Sanghyun Hong, and Niklas Elmqvist. A scanner deeply: Predicting gaze heatmaps on visualizations using crowdsourced eye movement data. IEEE Transactions on Visualization and Computer Graphics, 29(1): 396–406, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Schoenlein:2023:UED

[SCL^{+}23a] Melissa A. Schoenlein, Johnny Campos, Kevin J. Lande, Laurent Lessard, and Karen B. Schloss. Unifying effects of direct and relational associations for visual communication. IEEE Transactions on Visualization and Computer Graphics, 29(1):385–395, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Song:2023:NSD

[SCL^{+}23b] Liangchen Song, Anpei Chen, Zhong Li, Zhang Chen, Lele Chen, Junsong Yuan, Yi Xu, and Andreas Geiger. NeRF-Player: a streamable dynamic scene representation with decomposed neural radiance fields. IEEE Transactions on Visualization and Computer Graphics, 29(5):2732– 2742, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2024:NRC

[SCL^{+}24] Chuhan Shi, Weiwei Cui, Chengzhong Liu, Chengbo Zheng, Haidong Zhang, Qiong Luo, and Xiaojuan Ma. NL2Color: Refining color palettes for charts with natural language. IEEE Transactions on Visualization and Computer Graphics, 30(1):814–824, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Song:2021:DVD

[SCLK21] Chanyoung Song, Jehyun Cha, Mokwon Lee, and Deok-Soo Kim. Dynamic Voronoi diagram for moving disks. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2923–2940, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Sevastjanova:2023:VCL

[SCR^{+}23] Rita Sevastjanova, Eren Cakmak, Shauli Ravfogel, Ryan Cotterell, and Mennatallah El-Assady. Visual comparison of language model adaptation. IEEE Transactions on Visualization and Computer Graphics, 29(1):1178–1188, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Sarton:2020:IVD

[SCRL20] J. Sarton, N. Courilleau, Y. Remion, and L. Lucas. Interactive visualization and ondemand processing of large volume data: a fully GPU-based out-of-core approach. IEEE Transactions on Visualization and Computer Graphics, 26 (10):3008–3021, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Setlur:2024:HSC

[SCST24] Vidya Setlur, Michael Correll, Arvind Satyanarayan, and Melanie Tory. Heuristics for supporting cooperative dashboard design. IEEE Transactions on Visualization and Computer Graphics, 30(1):370– 380, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Saffo:2024:UDS

[SDC^{+}24] David Saffo, Sara Di Bartolomeo, Tarik Crnovrsanin, Laura South, Justin Raynor, Caglar Yildirim, and Cody Dunne. Unraveling the design space of immersive analytics: a systematic review. IEEE Transactions on Visualization and Computer Graphics, 30(1):495– 506, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Shen:2023:FRM

[SDK23a] Junxiao Shen, John Dudley, and Per Ola Kristensson. Fast and robust mid-air gesture typing for AR headsets using 3D trajectory decoding. IEEE Transactions on Visualization and Computer Graphics, 29(11):4622–4632, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Song:2023:HCC

[SDK23b] Zhaomou Song, John J. Dudley, and Per Ola Kristensson. HotGestures: Complementing command selection and use with delimiter-free gesture-based shortcuts in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4600– 4610, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2022:GSR

[SDMK22] Junxiao Shen, John Dudley, George Mo, and Per Ola Kristensson. Gesture spotter: a rapid prototyping tool for key gesture spotting in virtual and augmented reality applications. IEEE Transactions on Visualization and Computer Graphics, 28(11):3618–3628, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Shi:2022:LBI

[SDR22] Weiqi Shi, Julie Dorsey, and Holly Rushmeier. Learningbased inverse bi-scale material fitting from tabular BRDFs. IEEE Transactions on Visualization and Computer Graphics, 28(4):1810–1823, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Sisouk:2024:WDP

[SDT24] Keanu Sisouk, Julie Delon, and Julien Tierny. Wasserstein dictionaries of persistence diagrams. IEEE Transactions on Visualization and Computer Graphics, 30(2):1638– 1651, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Song:2022:IVP

[SDXR22] Huan Song, Zeng Dai, Panpan Xu, and Liu Ren. Interactive visual pattern search on graph data via graph representation learning. IEEE Transactions on Visualization and Computer Graphics, 28(1):335–345, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Sevastjanova:2022:PVC

[SEAB^{+}22] Rita Sevastjanova, Mennatallah El-Assady, Adam Bradley, Christopher Collins, Miriam Butt, and Daniel Keim. VisInReport: Complementing visual discourse analytics through personalized insight reports. IEEE Transactions on Visualization and Computer Graphics, 28(12):4757–4769, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Streeb:2021:WVU

[SEAKC21] D. Streeb, M. El-Assady, D. A. Keim, and M. Chen. Why visualize? Untangling a large network of arguments. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2220–2236, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Samuel:2024:VHV

[SEK^{+}24] Sara Samuel, Carmine Elvezio, Salaar Khan, Laureen Zubiaurre Bitzer, Letty Moss-Salentijn, and Steven Feiner. Visuo-haptic VR and AR guidance for dental nerve block education. IEEE Transactions on Visualization and Computer Graphics, 30(5):2839– 2848, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Singh:2020:EFD

[SES20] G. Singh, S. R. Ellis, and J. E. Swan. The effect of focal distance, age, and brightness on near-field augmented reality depth matching. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1385–1398, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Szafir:2023:WMV

[SES23] Danielle Szafir, David Ebert, and Hendrik Strobelt. Welcome: Message from the VIS 2022 general chairs. IEEE Transactions on Visualization and Computer Graphics, 29(1): xv, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Scimone:2024:MVT

[SESH24] Anna Scimone, Klaus Eckelt, Marc Streit, and Andreas Hinterreiter. Marjorie: Visualizing Type 1 diabetes data to support pattern exploration. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1216–1226, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Skaanberg:2022:TIF

[SFL^{+}22] Robin Skånberg, Martin Falk, Mathieu Linares, Anders Ynnerman, and Ingrid Hotz. Tracking internal frames of reference for consistent molecular distribution functions. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3126–3137, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Sarvghad:2023:SAI

[SFNRZ^{+}23] Ali Sarvghad, Rolando Franqui-Nadal, Rebecca Reznik-Zellen, Ria Chawla, and Narges Mahyar. Scientometric analysis of interdisciplinary collaboration and gender trends in 30 years of IEEE VIS publications. IEEE Transactions on Visualization and Computer Graphics, 29(7):3340–3353, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Spittle:2023:RIT

[SFPCW23] Becky Spittle, Maite Frutos-Pascual, Chris Creed, and Ian Williams. A review of interaction techniques for immersive environments. IEEE Transactions on Visualization and Computer Graphics, 29 (9):3900–3921, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Seinfeld:2022:IIP

[SFPM22] Sofia Seinfeld, Tiare Feuchtner, Johannes Pinzek, and Jörg Müller. Impact of information placement and user representations in VR on performance and embodiment. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1545–1556, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Shilpika:2022:VAA

[SFS^{+}22] Shilpika, Takanori Fujiwara, Naohisa Sakamoto, Jorji Nonaka, and Kwan-Liu Ma. A visual analytics approach for hardware system monitoring with streaming functional data analysis. IEEE Transactions on Visualization and Computer Graphics, 28(6):2338– 2349, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Stranner:2024:ISF

[SFSA24] Marco Stranner, Philipp Fleck, Dieter Schmalstieg, and Clemens Arth. Instant segmentation and fitting of excavations in subsurface utility engineering. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2319–2329, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Skarbez:2022:VRR

[SGB^{+}22] Richard Skarbez, Joseph L.

Gabbard, Doug A. Bowman, J. Todd Ogle, and Thomas Tucker. Virtual replicas of real places: Experimental investigations. IEEE Transactions on Visualization and Computer Graphics, 28(12):4594– 4608, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Sarma:2023:EUU

[SGH^{+}23] Abhraneel Sarma, Shunan Guo, Jane Hoffswell, Ryan Rossi, Fan Du, Eunyee Koh, and Matthew Kay. Evaluating the use of uncertainty visualisations for imputations of data missing at random in scatterplots. IEEE Transactions on Visualization and Computer Graphics, 29(1): 602–612, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Shi:2023:BFW

[SGJC23] Yang Shi, Tian Gao, Xiaohan Jiao, and Nan Cao. Breaking the fourth wall of data stories through interaction. IEEE Transactions on Visualization and Computer Graphics, 29(1): 972–982, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Snyder:2024:DDI

[SH24a] Luke S. Snyder and Jeffrey Heer. DIVI: Dynamically interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1): 403–413, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Subramonyam:2024:WCL

[SH24b] Hariharan Subramonyam and Jessica Hullman. Are we closing the loop yet? Gaps in the generalizability of VIS4ML research. IEEE Transactions on Visualization and Computer Graphics, 30(1):672–682, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Sun:2020:PVA

[SHC^{+}20] D. Sun, R. Huang, Y. Chen, Y. Wang, J. Zeng, M. Yuan, T. Pong, and H. Qu. PlanningVis: a visual analytics approach to production planning in smart factories. IEEE Transactions on Visualization and Computer Graphics, 26(1): 579–589, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schnorr:2020:FTT

[SHD^{+}20] A. Schnorr, D. N. Helmrich, D. Denker, T. W. Kuhlen, and B. Hentschel. Feature tracking by two-step optimization. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2219–2233, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Shen:2023:EMN

[She23a] Han-Wei Shen. Editorial: a message from the new Editorin-Chief. IEEE Transactions on Visualization and Computer Graphics, 29(4):1918– 1919, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2023:EN

[She23b] Han-Wei Shen. Editor’s note.

IEEE Transactions on Visualization and Computer Graphics, 29(4):1910–1917, April 2023. CODEN ITVGEA. ISSN 1077-

Shen:2024:MEC

[She24] Han-Wei Shen. Message from the Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 30(1): xiv, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Shi:2021:UVA

[SHL^{+}21] Lei Shi, Congcong Huang, Meijun Liu, Jia Yan, Tao Jiang, Zhihao Tan, Yifan Hu, Wei Chen, and Xiatian Zhang. UrbanMotion: Visual analysis of metropolitan-scale sparse trajectories. IEEE Transactions on Visualization and Computer Graphics, 27(10):3881– 3899, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Shao:2023:CLM

[SHM23] Han Shao, Libo Huang, and Dominik L. Michels. A current loop model for the fast simulation of ferrofluids. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5394–5405, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Savvides:2023:VDE

[SHOP23] Rafael Savvides, Andreas Henelius, Emilia Oikarinen, and Kai Puolamäki. Visual data exploration as a statistical testing procedure: Within-view and between-view multiple comparisons. IEEE Transactions on Visualization and Computer Graphics, 29(9):3937– 3948, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Saket:2020:IDM

[SHPE20] B. Saket, S. Huron, C. Perin, and A. Endert. Investigating direct manipulation of graphical encodings as a method for user interaction. IEEE Transactions on Visualization and Computer Graphics, 26(1):482–491, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Solah:2022:MDC

[SHS^{+}22] Michael Solah, Haikun Huang, Jiachuan Sheng, Tian Feng, Marc Pomplun, and Lap-Fai Yu. Mood-driven colorization of virtual indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2058–2068, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Shi:2022:PMM

[SHT^{+}22] Lei Shi, Junnan Hu, Zhihao Tan, Jun Tao, Jiayan Ding, Yan Jin, Yanjun Wu, and Paul M. Thompson. MVNet: Multivariate multi-view brain network comparison over uncertain data. IEEE Transactions on Visualization and Computer Graphics, 28(12):4640– 4657, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Suh:2020:PHG

[SHW^{+}20] A. Suh, M. Hajij, B. Wang, C. Scheidegger, and P. Rosen. Persistent homology guided force-directed graph layouts.

IEEE Transactions on Visualization and Computer Graphics, 26(1):697–707, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sicat:2023:RTV

[SIA^{+}23] Ronell Sicat, Mohamed Ibrahim, Amani Ageeli, Florian Mannuss, Peter Rautek, and Markus Hadwiger. Real-time visualization of large-scale geological models with nonlinear featurepreserving levels of detail. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1491–1505, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Sugimoto:2021:DDS

[SII^{+}21] Masatoki Sugimoto, Daisuke Iwai, Koki Ishida, Parinya Punpongsanon, and Kosuke Sato. Directionally decomposing structured light for projector calibration. IEEE Transactions on Visualization and Computer Graphics, 27(11): 4161–4170, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Somarakis:2021:VCC

[SIL^{+}21] A. Somarakis, M. E. Ijsselsteijn, S. J. Luk, B. Kenkhuis, N. F. C. C. de Miranda, B. P. F. Lelieveldt, and T. Höllt. Visual cohort comparison for spatial single-cell omics-data. IEEE Transactions on Visualization and Computer Graphics, 27(2):

733–743, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Shin:2023:RVA

[SJK^{+}23] DongHwa Shin, Jaemin Jo, Bohyoung Kim, Hyunjoo Song, Shin-Hyung Cho, and Jinwook Seo. RCMVis: a visual analytics system for route choice modeling. IEEE Transactions on Visualization and Computer Graphics, 29(3):1799– 1817, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Song:2023:FSA

[SJL^{+}23] Wenfeng Song, Xingliang Jin, Shuai Li, Chenglizhao Chen, Aimin Hao, and Xia Hou. FineStyle: Semantic-aware finegrained motion style transfer with dual interactive-flow fusion. IEEE Transactions on Visualization and Computer Graphics, 29(11):4361– 4371, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Sik:2020:SMC

[ŠK20] M. Šik and J. Křivánek. Survey of Markov Chain Monte Carlo methods in light transport simulation. IEEE Transactions on Visualization and Computer Graphics, 26(4):1821– 1840, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shen:2023:IVI

[SK23a] Han-Wei Shen and Kiyoshi Kiyokawa. IEEE VR 2023 introducing the special issue. IEEE Transactions on Visualization and Computer Graphics, 29(5):vii, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2023:MEC

[SK23b] Han-Wei Shen and Kiyoshi Kiyokawa. Message from the Editor-in-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 29(11):v, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2024:MECb

[SK24] Han-Wei Shen and Kiyoshi Kiyokawa. Message from the Editor-in-Chief and from the Associate Editor-in-Chief. IEEE Transactions on Visualization and Computer Graphics, 30(5):viii, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Saputra:2021:IDD

[SKA21] R. A. Saputra, C. S. Kaplan, and P. Asente. Improved deformation-driven element packing with RepulsionPak. IEEE Transactions on Visualization and Computer Graphics, 27(4):2396– 2408, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Soure:2022:CCV

[SKFZ22] Ehsan Jahangirzadeh Soure, Emily Kuang, Mingming Fan, and Jian Zhao. CoUX: Collaborative visual analysis of thinkaloud usability test videos for digital interfaces. IEEE Transactions on Visualization and Computer Graphics, 28(1):643– 653, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Shin:2023:RSA

[SKH^{+}23] Minjeong Shin, Joohee Kim, Yunha Han, Lexing Xie, Mitchell Whitelaw, Bum Chul Kwon, Sungahn Ko, and Niklas Elmqvist. Roslingifier: Semiautomated storytelling for animated scatterplots. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2980–2995, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Strobelt:2022:GHA

[SKK^{+}22] Hendrik Strobelt, Jambay Kinley, Robert Krueger, Johanna Beyer, Hanspeter Pfister, and Alexander M. Rush. GenNI: Human-AI collaboration for data-backed text generation. IEEE Transactions on Visualization and Computer Graphics, 28(1):1106–1116, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Strnad:2020:ERG

[SKNŽ20] Damjan Strnad, Štefan Kohek, Andrej Nerat, and Borut Žalik. Efficient representation of geometric tree models with levelof-detail using compressed 3D chain code. IEEE Transactions on Visualization and Computer Graphics, 26(11):3177– 3188, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Straub:2024:VAD

[SKR^{+}24] Alexander Straub, Nikolaos Karadimitriou, Guido Reina, Steffen Frey, Holger Steeb, and Thomas Ertl. Visual analysis of displacement processes in porous media using spatiotemporal flow graphs. IEEE Transactions on Visualization and Computer Graphics, 30(1): 759–769, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Skreinig:2023:GIA

[SKS^{+}23] Lucchas Ribeiro Skreinig, Denis Kalkofen, Ana Stanescu, Peter Mohr, Frank Heyen, Shohei Mori, Michael Sedlmair, Dieter Schmalstieg, and Alexander Plopski. guitARhero: Interactive augmented reality guitar tutorials. IEEE Transactions on Visualization and Computer Graphics, 29(11):4676– 4685, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Schott:2023:CIM

[SKW^{+}23] Danny Schott, Matthias Kunz, Tom Wunderling, Florian Heinrich, Rüdiger Braun-Dullaeus, and Christian Hansen. Cardio-Genesis4D: Interactive morphological transitions of embryonic heart development in a virtual learning environment. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2615–2625, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2021:ICC

[SLC21] L. Shi, R. S. Laramee, and G. Chen. Integral curve clustering and simplification for flow visualization: a comparative evaluation. IEEE Transactions on Visualization and Computer Graphics, 27(3): 1967–1985, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Shi:2023:SEF

[SLC^{+}23] Yang Shi, Pei Liu, Siji Chen, Mengdi Sun, and Nan Cao. Supporting expressive and faithful pictorial visualization design with visual style transfer. IEEE Transactions on Visualization and Computer Graphics, 29(1):236–246, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Sheng:2020:IID

[SLJ^{+}20] B. Sheng, P. Li, Y. Jin, P. Tan, and T. Lee. Intrinsic image decomposition with step and drift shading separation. IEEE Transactions on Visualization and Computer Graphics, 26(2): 1332–1346, February 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Snyder:2020:ILI

[SLK^{+}20] L. S. Snyder, Y. Lin, M. Karimzadeh, D. Goldwasser, and D. S. Ebert. Interactive learning for identifying relevant tweets to support real-time situational awareness.

IEEE Transactions on Visualization and Computer Graphics, 26(1):558–568, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Schloss:2021:SDV

[SLL21] K. B. Schloss, Z. Leggon, and L. Lessard. Semantic discriminability for visual communication. IEEE Transactions on Visualization and Computer Graphics, 27(2):1022– 1031, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Satyanarayan:2020:CRV

[SLR^{+}20a] A. Satyanarayan, B. Lee, D. Ren, J. Heer, J. Stasko, J. Thompson, M. Brehmer, and Z. Liu. Critical reflections on visualization authoring systems. IEEE Transactions on Visualization and Computer Graphics, 26(1):461–471, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Song:2020:DOS

[SLR20b] R. Song, Y. Liu, and P. L. Rosin. Distinction of 3D objects and scenes via classification network and Markov random field. IEEE Transactions on Visualization and Computer Graphics, 26(6):2204–2218, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Song:2021:MSW

[SLR21] R. Song, Y. Liu, and P. L. Rosin. Mesh saliency via weakly supervised classification-for-saliency CNN. IEEE Transactions on Visualization and Computer Graphics, 27(1):151– 164, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Srinivasan:2021:IMI

[SLS21] Arjun Srinivasan, Bongshin Lee, and John Stasko. Interweaving multimodal interaction with flexible unit visualizations for data exploration. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3519–3533, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Song:2023:VLE

[SLSW23] Sicheng Song, Chenhui Li, Yujing Sun, and Changbo Wang. VividGraph: Learning to extract and redesign network graphs from visualization images. IEEE Transactions on Visualization and Computer Graphics, 29(7):3169– 3181, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Sterzik:2024:PLA

[SLW^{+}24] Anna Sterzik, Nils Lichtenberg, Jana Wilms, Michael Krone, Douglas W. Cunningham, and Kai Lawonn. Perception of line attributes for visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):1041–1051, January 2024. CODEN ITVGEA. ISSN 1077-

Shen:2023:IID

[SLX^{+}23] Jingyi Shen, Haoyu Li, Jiayi Xu, Ayan Biswas, and Han-Wei Shen. IDLat: an importance-driven latent generation method for scientific data. IEEE Transactions on Visualization and Computer Graphics, 29(1):679–689, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Sizintsev:2021:LRA

[SMC^{+}21] Mikhail Sizintsev, Niluthpol Chowdhury Mithun, Han-Pang Chiu, Supun Samarasekera, and Rakesh Kumar. Longrange augmented reality with dynamic occlusion rendering. IEEE Transactions on Visualization and Computer Graphics, 27(11):4236–4244, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Sterzik:2024:PUC

[SMCL24] Anna Sterzik, Monique Meuschke, Douglas W. Cunningham, and Kai Lawonn. Perceptually uniform construction of illustrative textures. IEEE Transactions on Visualization and Computer Graphics, 30(1):1052– 1062, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Sridharamurthy:2020:EDB

[SMKN20] R. Sridharamurthy, T. B. Masood, A. Kamakshidasan, and V. Natarajan. Edit distance between merge trees. IEEE Transactions on Visualization and Computer Graphics, 26(3):

1518–1531, March 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Sermarini:2023:IIA

[SML^{+}23] John Sermarini, Robert A.

Michlowitz, Joseph J. LaViola, Lori C. Walters, Roger Azevedo, and Joseph T. Kider. Investigating the impact of augmented reality and BIM on retrofitting training for nonexperts. IEEE Transactions on Visualization and Computer Graphics, 29(11):4655– 4665, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Sakono:2021:RWU

[SMNK21] Hiroaki Sakono, Keigo Matsumoto, Takuji Narumi, and Hideaki Kuzuoka. Redirected walking using continuous curvature manipulation. IEEE Transactions on Visualization and Computer Graphics, 27(11):4278–4288, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Si-Mohammed:2020:TBB

[SMPJ^{+}20] H. Si-Mohammed, J. Petit, C. Jeunet, F. Argelaguet, F. Spindler, A. Évain, N. Roussel, G. Casiez, and A. Lecuyer. Towards BCI-based interfaces for augmented reality: Feasibility, design and evaluation. IEEE Transactions on Visualization and Computer Graphics, 26(3):1608–1621, March 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Streeb:2022:TBV

[SMS^{+}22] Dirk Streeb, Yannick Metz, Udo Schlegel, Bruno Schneider, Mennatallah El-Assady, Hansjörg Neth, Min Chen, and Daniel A. Keim. Task-based visual interactive modeling: Decision trees and rule-based classifiers. IEEE Transactions on Visualization and Computer Graphics, 28(9):3307– 3323, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Stanescu:2022:MFA

[SMSK22] Ana Stanescu, Peter Mohr, Dieter Schmalstieg, and Denis Kalkofen. Model-free authoring by demonstration of assembly instructions in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(11):3821– 3831, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Sui:2022:PQA

[SMYF22] Xiangjie Sui, Kede Ma, Yiru Yao, and Yuming Fang. Perceptual quality assessment of omnidirectional images as moving camera videos. IEEE Transactions on Visualization and Computer Graphics, 28(8): 3022–3034, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Summa:2022:W

[SN22] Brian Summa and Luis Gustavo Nonato. Welcome. IEEE Transactions on Visualization and Computer Graphics, 28(1): xiii, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Sridharamurthy:2023:CAM

[SN23] Raghavendra Sridharamurthy and Vijay Natarajan. Comparative analysis of merge trees using local tree edit distance. IEEE Transactions on Visualization and Computer Graphics, 29(2):1518–1530, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Sultanum:2023:CNL

[SNBC23] Nicole Sultanum, Farooq Naeem, Michael Brudno, and Fanny Chevalier. ChartWalk: Navigating large collections of text notes in electronic health records for clinical chart review. IEEE Transactions on Visualization and Computer Graphics, 29(1):1244–1254, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2023:MIV

[SNH^{+}23] Chuhan Shi, Fei Nie, Yicheng Hu, Yige Xu, Lei Chen, Xiaojuan Ma, and Qiong Luo. Med-ChemLens: an interactive visual tool to support direction selection in interdisciplinary experimental research of medicinal chemistry. IEEE Transactions on Visualization and Computer Graphics, 29(1):63– 73, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Sun:2022:TSD

[SNK^{+}22] Maoyuan Sun, Akhil Namburi, David Koop, Jian Zhao, Tianyi Li, and Haeyong Chung. Towards systematic design considerations for visualizing cross-view data relationships. IEEE Transactions on Visualization and Computer Graphics, 28(12):4741–4756, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Sun:2022:CDS

[SOL^{+}22] Yucheng Sun, Wenqing Ouyang, Zhongyuan Liu, Ning Ni, Yann Savoye, Peng Song, and Ligang Liu. Computational design of self-actuated deformable solids via shape memory material. IEEE Transactions on Visualization and Computer Graphics, 28(7):2577–2588, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Seebacher:2023:ISN

[SPJ^{+}23] Daniel Seebacher, Tom Polk, Halldor Janetzko, Daniel A. Keim, Tobias Schreck, and Manuel Stein. Investigating the sketchplan: a novel way of identifying tactical behavior in massive soccer datasets. IEEE Transactions on Visualization and Computer Graphics, 29(4): 1920–1936, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Schmidt:2024:VAU

[SPM24] Johanna Schmidt, Bernhard Pointner, and Silvia Miksch. Visual analytics for understanding draco’s knowledge base. IEEE Transactions on Visualization and Computer Graphics, 30(1): 392–402, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Subhash:2023:GPA

[SPN23] Varshini Subhash, Karran Pandey, and Vijay Natarajan. A GPU parallel algorithm for computing Morse–Smale complexes. IEEE Transactions on Visualization and Computer Graphics, 29(9):3873– 3887, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Sidenmark:2023:CGH

[SPNG23] Ludwig Sidenmark, Franziska Prummer, Joshua Newn, and Hans Gellersen. Comparing gaze, head and controller selection of dynamically revealed targets in head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 29(11):4740– 4750, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Sidenmark:2022:WPE

[SPW^{+}22] Ludwig Sidenmark, Mark Parent, Chi-Hao Wu, Joannes Chan, Michael Glueck, Daniel Wigdor, Tovi Grossman, and Marcello Giordano. Weighted pointer: Error-aware gazebased interaction through fallback modalities. IEEE Transactions on Visualization and Computer Graphics, 28(11): 3585–3595, November 2022.

CODEN ITVGEA. ISSN 1077-2626.

Saeedpour-Parizi:2024:PTR

[SPWW^{+}24] Mohammad R. Saeedpour-Parizi, Niall L. Williams, Tim Wong, Phillip Guan, Dinesh Manocha, and Ian M. Erkelens. Perceptual thresholds for radial optic flow distortion in neareye stereoscopic displays. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2570–2579, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Seraji:2024:AUB

[SPZS24] Mohammad Rajabi Seraji, Parastoo Piray, Vahid Zahednejad, and Wolfgang Stuerzlinger. Analyzing user behaviour patterns in a cross-virtuality immersive analytics system. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2613–2623, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Strauss:2020:SAR

[SRBP20] R. R. Strauss, R. Ramanujan, A. Becker, and T. C. Peck. A steering algorithm for redirected walking using reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1955–1963, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Schufrin:2021:VII

[SRKK21] M. Schufrin, S. L. Reynolds, A. Kuijper, and J. Kohlhammer. A visualization interface to improve the transparency of collected personal data on the Internet. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1840–1849, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Svirsky:2021:NLD

[SS21] Yonatan Svirsky and Andrei Sharf. A non-linear differentiable CNN-rendering module for 3D data enhancement. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3238–3249, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Shen:2024:PPS

[SS24] Jingyi Shen and Han-Wei Shen.

PSRFlow: Probabilistic super resolution with flow-based models for scientific data. IEEE Transactions on Visualization and Computer Graphics, 30(1): 986–996, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Sun:2022:SEC

[SSAZ22] Maoyuan Sun, Abdul Rahman Shaikh, Hamed Alhoori, and Jian Zhao. SightBi: Exploring cross-view data relationships with biclusters. IEEE Transactions on Visualization and Computer Graphics, 28(1): 54–64, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Schwab:2022:SSV

[SSB^{+}22] Michail Schwab, David Saffo, Nicholas Bond, Shash Sinha, Cody Dunne, Jeff Huang, James Tompkin, and Michelle A. Borkin. Scalable scalable vector graphics: Automatic translation of interactive SVGs to a multithread VDOM for fast rendering. IEEE Transactions on Visualization and Computer Graphics, 28(9):3219– 3234, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Stokes:2023:SBR

[SSC^{+}23] Chase Stokes, Vidya Setlur, Bridget Cogley, Arvind Satyanarayan, and Marti A. Hearst. Striking a balance: Reader takeaways and preferences when integrating text and charts. IEEE Transactions on Visualization and Computer Graphics, 29(1):1233– 1243, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Sohns:2022:ABE

[SSJ^{+}22] Jan-Tobias Sohns, Michaela Schmitt, Fabian Jirasek, Hans Hasse, and Heike Leitte. Attribute-based explanation of non-linear embeddings of highdimensional data. IEEE Transactions on Visualization and Computer Graphics, 28(1):540– 550, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Shen:2023:TNL

[SSL^{+}23] Leixian Shen, Enya Shen, Yuyu Luo, Xiaocong Yang, Xuming Hu, Xiongshuai Zhang, Zhiwei Tai, and Jianmin Wang. Towards natural language interfaces for data visualization: a survey. IEEE Transactions on Visualization and Computer Graphics, 29(6):3121– 3144, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Saktheeswaran:2020:TST

[SSS20] A. Saktheeswaran, A. Srinivasan, and J. Stasko. Touch? Speech? or touch and speech? Investigating multimodal interaction for visual network exploration and analysis. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2168–2179, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Spinner:2020:EVA

[SSSEA20] T. Spinner, U. Schlegel, H. Schäfer, and M. El-Assady. explAIner: a visual analytics framework for interactive and explainable machine learning. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1064–1074, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Shu:2020:SBS

[SSX^{+}20] Z. Shu, X. Shen, S. Xin, Q. Chang, J. Feng, L. Kavan, and L. Liu. Scribblebased 3D shape segmentation via weakly-supervised learning. IEEE Transactions on Visualization and Computer Graphics, 26(8):2671–2682, August 2020. CODEN ITVGEA. ISSN 1077-

Schwab:2021:VDE

[SSZ^{+}21] M. Schwab, D. Saffo, Y. Zhang, S. Sinha, C. Nita-Rotaru, J. Tompkin, C. Dunne, and M. A. Borkin. VisConnect: Distributed event synchronization for collaborative visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2):347–357, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Steed:2021:DUI

[STA^{+}21] Anthony Steed, Tuukka M. Takala, Daniel Archer, Wallace Lages, and Robert W. Lindeman. Directions for 3D user interface research from consumer VR games. IEEE Transactions on Visualization and Computer Graphics, 27(11):4171– 4182, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Song:2023:BAG

[STD^{+}23] Yu Song, Fan Tang, Weiming Dong, Feiyue Huang, Tong-Yee Lee, and Changsheng Xu. Balance-aware grid collage for small image collections. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1330–1344, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Sawada:2022:PTQ

[SUB^{+}22] Naoko Sawada, Makoto Uemura, Johanna Beyer, Hanspeter Pfister, and Issei Fujishiro. TimeTubesX: a query-driven visual exploration of observable, photometric, and polarimetric behaviors of blazars. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1917–1929, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Sabando:2021:CIV

[SUS^{+}21] M. V. Sabando, P. Ulbrich, M. Selzer, J. By ka, J. Mi an, I. Ponzoni, A. J. Soto, M. L. Ganuza, and B. Kozlíková. ChemVA: Interactive visual analysis of chemical compound similarity in virtual screening. IEEE Transactions on Visualization and Computer Graphics, 27(2):891–901, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Somarakis:2021:IVE

[SVK^{+}21] A. Somarakis, V. Van Unen, F. Koning, B. Lelieveldt, and T. Höllt. ImaCytE: Visual exploration of cellular microenvironments for imaging mass cytometry data. IEEE Transactions on Visualization and Computer Graphics, 27(1):98– 110, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Sra:2021:VVR

[SW21] Misha Sra and Folker Wientapper. The 2020 VGTC Virtual Reality Best Dissertation Award. IEEE Transactions on Visualization and Computer Graphics, 27(5):x, May 2021. CODEN ITVGEA. ISSN 1077-

Sereno:2022:CWA

[SWB^{+}22] Mickael Sereno, Xiyao Wang, Lonni Besançon, Michael J. McGuffin, and Tobias Isenberg. Collaborative work in augmented reality: a survey. IEEE Transactions on Visualization and Computer Graphics, 28(6):2530–2549, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Salagean:2024:UVS

[SWF^{+}24] Anca Salagean, Michelle Wu, George Fletcher, Darren Cosker, and Danaë Stanton Fraser. The utilitarian virtual self using embodied personalized avatars to investigate moral decisionmaking in semi-autonomous vehicle dilemmas. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2162–2172, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Smart:2020:CCA

[SWS20] S. Smart, K. Wu, and D. A. Szafir. Color crafting: Automating the construction of designer quality color ramps. IEEE Transactions on Visualization and Computer Graphics, 26(1):1215–1225, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Strobelt:2023:IVP

[SWS^{+}23] Hendrik Strobelt, Albert Webson, Victor Sanh, Benjamin Hoover, Johanna Beyer, Hanspeter Pfister, and Alexander M.

Rush. Interactive and visual prompt engineering for ad-hoc task adaptation with large language models. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1146–1156, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Son:2023:UBT

[SWSK23] Hyungki Son, Haokun Wang, Yatharth Singhal, and Jin Ryong Kim. Upper body thermal referral and tactile masking for localized feedback. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2211–2219, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Shu:2021:WMD

[SWT^{+}21] X. Shu, A. Wu, J. Tang, B. Bach, Y. Wu, and H. Qu. What makes a data-GIF understandable? IEEE Transactions on Visualization and Computer Graphics, 27(2):1492– 1502, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Shi:2022:FSL

[SWW^{+}22] Xuehuai Shi, Lili Wang, Jian Wu, Runze Fan, and Aimin Hao. Foveated stochastic lightcuts. IEEE Transactions on Visualization and Computer Graphics, 28(11):3684– 3693, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Shi:2021:FPM

[SWWY21] Xuehuai Shi, Lili Wang, Xiaoheng Wei, and Ling-Qi Yan.

Foveated photon mapping. IEEE Transactions on Visualization and Computer Graphics, 27(11):4183–4193, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Su:2022:PMM

[SWY^{+}22] Zhaoqi Su, Weilin Wan, Tao Yu, Lingjie Liu, Lu Fang, Wenping Wang, and Yebin Liu. MulayCap: Multi-layer human performance capture using a monocular video camera. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1862–1879, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Song:2023:LRG

[SWZ^{+}23] Zhengxi Song, Xue Wang, Hao Zhu, Guoqing Zhou, and Qing Wang. Learning reliable gradients from undersampled circular light field for 3D reconstruction. IEEE Transactions on Visualization and Computer Graphics, 29(12):5194– 5207, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2023:VSV

[SXL^{+}23] Neng Shi, Jiayi Xu, Haoyu Li, Hanqi Guo, Jonathan Woodring, and Han-Wei Shen. VDL-Surrogate: a view-dependent latent-based model for parameter space exploration of ensemble simulations. IEEE Transactions on Visualization and Computer Graphics, 29(1):820– 830, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Shi:2021:CAV

[SXS^{+}21] D. Shi, X. Xu, F. Sun, Y. Shi, and N. Cao. Calliope: Automatic visual data story generation from a spreadsheet. IEEE Transactions on Visualization and Computer Graphics, 27(2): 453–463, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Shi:2022:GSH

[SXW^{+}22] Neng Shi, Jiayi Xu, Skylar W. Wurster, Hanqi Guo, Jonathan Woodring, Luke P. Van Roekel, and Han-Wei Shen. GNN-Surrogate: a hierarchical and adaptive graph neural network for parameter space exploration of unstructuredmesh ocean simulations. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2301–2313, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Su:2023:DPI

[SYC^{+}23] Wanchao Su, Hui Ye, Shu-Yu Chen, Lin Gao, and Hongbo Fu. DrawingInStyles: Portrait image generation and editing with spatially conditioned StyleGAN. IEEE Transactions on Visualization and Computer Graphics, 29(10):4074– 4088, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2020:IBH

[SYHS20] Y. Shen, L. Yang, E. S. L. Ho, and H. P. H. Shum. Interactionbased human activity comparison. IEEE Transactions on Visualization and Computer Graphics, 26(8):2620–2633, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Shao:2020:HCS

[SYW^{+}20] T. Shao, Y. Yang, Y. Weng, Q. Hou, and K. Zhou. H-CNN: Spatial hashing based CNN for 3D shape analysis. IEEE Transactions on Visualization and Computer Graphics, 26(7):2403–2416, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Shu:2022:GBM

[SYX^{+}22] Yezhi Shu, Ran Yi, Mengfei Xia, Zipeng Ye, Wang Zhao, Yang Chen, Yu-Kun Lai, and Yong-Jin Liu. GAN-based multi-style photo cartoonization. IEEE Transactions on Visualization and Computer Graphics, 28(10):3376– 3390, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Shen:2024:QMG

[SYY^{+}24] Qiaomu Shen, Zhengxin You, Xiao Yan, Chaozu Zhang, Ke Xu, Dan Zeng, Jianbin Qin, and Bo Tang. QEVIS: Multigrained visualization of distributed query execution. IEEE Transactions on Visualization and Computer Graphics, 30(1): 153–163, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Shi:2023:RBD

[SZC^{+}23] Min Shi, Jia-Qi Zhang, Shu-Yu Chen, Lin Gao, Yu-Kun Lai, and Fang-Lue Zhang. Reference-based deep line art video colorization. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2965–2979, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Shen:2021:DDS

[SZF^{+}21] Yuefan Shen, Changgeng Zhang, Hongbo Fu, Kun Zhou, and Youyi Zheng. DeepSketchHair: Deep sketch-based 3D hair modeling. IEEE Transactions on Visualization and Computer Graphics, 27(7):3250– 3263, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Sun:2020:CPE

[SZH^{+}20] X. Sun, Y. Zhang, P.-C. Huang, N. Acharjee, M. Dagenais, M. Peckerar, and A. Varshney. Correcting the proximity effect in nanophotonic phased arrays. IEEE Transactions on Visualization and Computer Graphics, 26(12):3503–3513, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Shen:2024:DPA

[SZZW24] Leixian Shen, Yizhi Zhang, Haidong Zhang, and Yun Wang. Data player: Automatic generation of data videos with narration-animation interplay. IEEE Transactions on Visualization and Computer Graphics, 30(1):109–119, January 2024. CODEN ITVGEA. ISSN 1077-

Teng:2024:VVA

[TAL24] Xian Teng, Yongsu Ahn, and Yu-Ru Lin. Vispur: Visual aids for identifying and interpreting spurious associations in data-driven decisions. IEEE Transactions on Visualization and Computer Graphics, 30(1): 219–229, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Tandon:2023:MES

[TARB23] Sara Tandon, Alfie Abdul-Rahman, and Rita Borgo. Measuring effects of spatial visualization and domain on visualization task performance: a comparative study. IEEE Transactions on Visualization and Computer Graphics, 29(1): 668–678, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Tian:2024:WSY

[TB24] Nana Tian and Ronan Boulic. Who says you are so sick? An investigation on individual susceptibility to cybersickness triggers using EEG, EGG and ECG. IEEE Transactions on Visualization and Computer Graphics, 30(5):2379– 2389, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Tang:2020:SAA

[TBL^{+}20] Z. Tang, N. J. Bryan, D. Li, T. R. Langlois, and D. Manocha. Scene-aware audio rendering via deep acoustic analysis. IEEE Transactions on Visualization and Computer Graphics, 26(5):1991–2001, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Tran:2023:WAR

[TBW^{+}23] Tram Thi Minh Tran, Shane Brown, Oliver Weidlich, Mark Billinghurst, and Callum Parker. Wearable augmented reality: Research trends and future directions from three major venues. IEEE Transactions on Visualization and Computer Graphics, 29(11):4782– 4793, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Tong:2023:EIP

[TCX^{+}23] Wai Tong, Zhutian Chen, Meng Xia, Leo Yu-Ho Lo, Linping Yuan, Benjamin Bach, and Huamin Qu. Exploring interactions with printed data visualizations in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(1):418–428, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Tovanich:2022:GYI

[TDI22] Natkamon Tovanich, Pierre Dragicevic, and Petra Isenberg. Gender in 30 years of IEEE Visualization. IEEE Transactions on Visualization and Computer Graphics, 28(1):497–507, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Tyagi:2023:PEM

[TEK^{+}23] Anjul Tyagi, Tyler Estro, Geoff Kuenning, Erez Zadok, and Klaus Mueller. PC-Expo: a metrics-based interactive axes reordering method for parallel coordinate displays. IEEE Transactions on Visualization and Computer Graphics, 29(1): 712–722, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Tetzlaff:2024:HFS

[Tet24] Michael Tetzlaff. High-fidelity specular SVBRDF acquisition from flash photographs. IEEE Transactions on Visualization and Computer Graphics, 30(4): 1885–1896, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Tkachev:2021:LPM

[TFE21] Gleb Tkachev, Steffen Frey, and Thomas Ertl. Local prediction models for spatiotemporal volume visualization. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3091–3108, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Tkachev:2022:SSS

[TFE22] Gleb Tkachev, Steffen Frey, and Thomas Ertl. S4: Selfsupervised learning of spatiotemporal similarity. IEEE Transactions on Visualization and Computer Graphics, 28(12):4713–4727, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Tian:2024:SSV

[TG24] Xingze Tian and Tobias Günther. A survey of smooth vector graphics: Recent advances in representation, creation, rasterization, and image vectorization. IEEE Transactions on Visualization and Computer Graphics, 30(3):1652– 1671, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Tehrani:2021:AGR

[TGM21] M. A. Tehrani, M. Gopi, and A. Majumder. Automated geometric registration for multiprojector displays on arbitrary 3D shapes using uncalibrated devices. IEEE Transactions on Visualization and Computer Graphics, 27(4):2265– 2279, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Tovanich:2021:VBD

[THFI21] Natkamon Tovanich, Nicolas Heulot, Jean-Daniel Fekete, and Petra Isenberg. Visualization of blockchain data: a systematic review. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3135–3152, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Tao:2021:KAS

[THS^{+}21] W. Tao, X. Hou, A. Sah, L. Battle, R. Chang, and M. Stonebraker. Kyrix-S: Authoring scalable scatterplot visualizations of Big Data. IEEE Transactions on Visualization and Computer Graphics, 27(2):401– 411, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Tasnim:2024:IPT

[TIDQ24] Umama Tasnim, Rifatul Islam, Kevin Desai, and John Quarles. Investigating personalization techniques for improved cybersickness prediction in virtual reality environments. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2368–2378, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Tone:2020:FEO

[TIHS20] D. Tone, D. Iwai, S. Hiura, and K. Sato. FibAR: Embedding optical fibers in 3D printed objects for active markers in dynamic projection mapping. IEEE Transactions on Visualization and Computer Graphics, 26(5):2030–2040, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Takeuchi:2024:PMU

[TKIS24] Masaki Takeuchi, Hiroki Kusuyama, Daisuke Iwai, and Kosuke Sato. Projection mapping under environmental lighting by replacing room lights with heterogeneous projectors. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2151–2161, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Tian:2023:UVR

[TLBB23] Huayuan Tian, Gun A. Lee, Huidong Bai, and Mark Billinghurst. Using virtual replicas to improve mixed reality remote collaboration. IEEE Transactions on Visualization and Computer Graphics, 29(5):2785– 2795, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Tang:2021:PCE

[TLW^{+}21] T. Tang, R. Li, X. Wu, S. Liu, J. Knittel, S. Koch, L. Yu, P. Ren, T. Ertl, and Y. Wu. PlotThread: Creating expressive storyline visualizations using reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 27(2):294–303, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Tu:2023:SVQ

[TLW^{+}23] Yamei Tu, Olga Li, Junpeng Wang, Han-Wei Shen, Przemek Powa ko, Irina Tomescu-Dubrow, Kazimierz M. Slomczynski, Spyros Blanas, and J. Craig Jenkins. SDRQuerier: a visual querying framework for cross-national survey data recycling. IEEE Transactions on Visualization and Computer Graphics, 29(6):2862– 2874, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Trepkowski:2022:MPT

[TME^{+}22] Christina Trepkowski, Alexander Marquardt, Tom David Eibich, Yusuke Shikanai, Jens Maiero, Kiyoshi Kiyokawa, Ernst Kruijff, Johannes Schöning, and Peter König. Multisensory proximity and transition cues for improving target awareness in narrow field of view augmented reality displays. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1342–1362, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Thiel:2022:PCD

[TNJ^{+}22] Kevin Kennard Thiel, Florian Naumann, Eduard Jundt, Stephan Günnemann, and Gudrun Klinker. C.DOT — convolutional deep object tracker for augmented reality based purely on synthetic data. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4434–4451, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Tian:2024:ETC

[TNZ^{+}24] Feng Tian, Shuting Ni, Xiaoyue Zhang, Fei Chen, Qiaolian Zhu, Chunyi Xu, and Yuzhi Li. Enhancing tai chi training system: Towards group-based and hyper-realistic training experiences. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2713–2723, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Tsai:2022:FSV

[TPH22] Wan-Lun Tsai, Tse-Yu Pan, and Min-Chun Hu. Feasibility study on virtual reality based basketball tactic training. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2970–2982, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Tu:2024:PAB

[TQW^{+}24] Yamei Tu, Rui Qiu, Yu-Shuen Wang, Po-Yin Yen, and Han-Wei Shen. PhraseMap: Attention-based keyphrases recommendation for information seeking. IEEE Transactions on Visualization and Computer Graphics, 30(3):1787– 1802, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Tsandilas:2021:SFV

[Tsa21] T. Tsandilas. StructGraphics: Flexible visualization design through data-agnostic and reusable graphical structures. IEEE Transactions on Visualization and Computer Graphics, 27(2):315–325, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Tricoche:2021:EVP

[TSH21] X. Tricoche, W. Schlei, and K. C. Howell. Extraction and visualization of Poincaré map topology for spacecraft trajectory design. IEEE Transactions on Visualization and Computer Graphics, 27(2):765– 774, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Tovanich:2022:PMV

[TSHI22] Natkamon Tovanich, Nicolas Soulié, Nicolas Heulot, and Petra Isenberg. MiningVis: Visual analytics of the bitcoin mining economy. IEEE Transactions on Visualization and Computer Graphics, 28(1):868–878, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Tuffaha:2022:MMI

[TSS22] Mutaz Tuffaha, Øyvind Stavdahl, and Ann-Katrin Stensdotter. Modeling movement-induced errors in AC electromagnetic trackers. IEEE Transactions on Visualization and Computer Graphics, 28(3):1597– 1607, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Tong:2022:ILS

[TWA22] Jonathan Tong, Laurie M. Wilcox, and Robert S. Allison. The impacts of lens and stereo camera separation on perceived slant in virtual reality head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 28(11):3759–3766, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Troidl:2024:VVA

[TWC^{+}24] Jakob Troidl, Simon Warchol, Jinhan Choi, Jordan Matelsky, Nagaraju Dhanyasi, Xueying Wang, Brock Wester, Donglai Wei, Jeff W. Lichtman, Hanspeter Pfister, and Johanna Beyer. VIMO: Visual analysis of neuronal connectivity motifs. IEEE Transactions on Visualization and Computer Graphics, 30(1):748–758, January 2024. CODEN ITVGEA. ISSN 1077-

Tsao:2022:FPM

[TWT^{+}22] Chih-An Tsao, Tzu-Chun Wu, Hsin-Ruey Tsai, Tzu-Yun Wei, Fang-Ying Liao, Sean Chapman, and Bing-Yu Chen. Frict-Shoes: Providing multilevel nonuniform friction feedback on shoes in VR. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2026–2036, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Tang:2022:VRA

[TWW^{+}22] Tan Tang, Yanhong Wu, Yingcai Wu, Lingyun Yu, and Yuhong Li. VideoModerator: a risk-aware framework for multimodal video moderation in ecommerce. IEEE Transactions on Visualization and Computer Graphics, 28(1):846–856, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Tyagi:2023:NNV

[TXM23] Anjul Tyagi, Cong Xie, and Klaus Mueller. NAS-Navigator: Visual steering for explainable one-shot deep neural network synthesis. IEEE Transactions on Visualization and Computer Graphics, 29(1):299–309, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Tong:2020:SMS

[TYPC20] W. Tong, X. Yang, M. Pan, and F. Chen. Spectral mesh segmentation via ℓ_{0} gradient minimization. IEEE Transactions on Visualization and Computer Graphics, 26(4):1807– 1820, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Truong:2022:PMS

[TYW^{+}22] Nghia Truong, Cem Yuksel, Chakrit Watcharopas, Joshua A. Levine, and Robert M. Kirby. Particle merging-and-splitting. IEEE Transactions on Visualization and Computer Graphics, 28(12):4546– 4557, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Tang:2022:VAM

[TZT^{+}22] Junxiu Tang, Yuhua Zhou, Tan Tang, Di Weng, Boyang Xie, Lingyun Yu, Huaqiang Zhang, and Yingcai Wu. A visualization approach for monitoring order processing in e-commerce warehouse. IEEE Transactions on Visualization and Computer Graphics, 28(1):857–867, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Ullah:2023:EUP

[UDH23] A K M Amanat Ullah, William Delamare, and Khalad Hasan. Exploring users pointing performance on large displays with different curvatures in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(11):4535– 4545, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Ueda:2020:IFV

[UIHS20] T. Ueda, D. Iwai, T. Hiraki, and K. Sato. Illuminated focus: Vision augmentation using spatial defocusing via focal sweep eyeglasses and high-speed projector. IEEE Transactions on Visualization and Computer Graphics, 26(5): 2051–2061, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Unruh:2023:BTV

[UVL^{+}23] Fabian Unruh, David Vogel, Maximilian Landeck, Jean-Luc Lugrin, and Marc Erich Latoschik. Body and time: Virtual embodiment and its effect on time perception. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2626–2636, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Ulbrich:2023:SDM

[UWF^{+}23] Pavol Ulbrich, Manuela Waldner, Katarína Furmanová, Sérgio M. Marques, David Bednáł, Barbora Kozlíková, and Jan Byška. sMolBoxes: Dataflow model for molecular dynamics exploration. IEEE Transactions on Visualization and Computer Graphics, 29(1): 581–590, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Vallageas:2024:ESA

[VAWL24] Valentin Vallageas, Rachid Aissaoui, Iris Willaert, and David R. Labbé. Embodying a self-avatar with a larger leg: its impacts on motor control and dynamic stability. IEEE Transactions on Visualization and Computer Graphics, 30(5):

2066–2076, May 2024. CODEN ITVGEA. ISSN 1077-2626.

vanBeusekom:2022:SMO

[vBMS22] Nathan van Beusekom, Wouter Meulemans, and Bettina Speckmann. Simultaneous matrix orderings for graph collections. IEEE Transactions on Visualization and Computer Graphics, 28(1):1–10, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Valdivia:2021:ADH

[VBP^{+}21] P. Valdivia, P. Buono, C. Plaisant, N. Dufournaud, and J.-D. Fekete. Analyzing dynamic hypergraphs with parallel aggregated ordered hypergraph visualization. IEEE Transactions on Visualization and Computer Graphics, 27(1):1–13, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Vidal:2020:PWB

[VBT20] J. Vidal, J. Budin, and J. Tierny. Progressive Wasserstein barycenters of persistence diagrams. IEEE Transactions on Visualization and Computer Graphics, 26(1):151–161, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Volmer:2023:ERB

[VBV^{+}23] Benjamin Volmer, James Baumeister, Stewart Von Itzstein, Matthias Schlesewsky, Ina Bornkessel-Schlesewsky, and Bruce H. Thomas. Event related brain responses reveal the impact of spatial augmented reality predictive cues on mental effort. IEEE Transactions on Visualization and Computer Graphics, 29(12):4990– 5007, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Veras:2020:DTV

[VC20] R. Veras and C. Collins. Discriminability tests for visualization effectiveness and scalability. IEEE Transactions on Visualization and Computer Graphics, 26(1):749–758, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Vajiac:2023:TVO

[VCO^{+}23] Catalina Vajiac, Duen Horng Chau, Andreas Olligschlaeger, Rebecca Mackenzie, Pratheeksha Nair, Meng-Chieh Lee, Yifei Li, Namyong Park, Reihaneh Rabbany, and Christos Faloutsos. TrafficVis: Visualizing organized activity and spatiotemporal patterns for detecting and labeling human trafficking. IEEE Transactions on Visualization and Computer Graphics, 29(1):53–62, January 2023. CODEN ITVGEA. ISSN 1077-2626.

vandeRuit:2022:EDH

[vdRBE22] Mark van de Ruit, Markus Billeter, and Elmar Eisemann. An efficient dual-hierarchy t-SNE minimization. IEEE Transactions on Visualization and Computer Graphics, 28(1):614– 622, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Verbraeck:2021:IBH

[VE21] A. Verbraeck and E. Eisemann. Interactive black-hole visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2):796– 805, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Valdrighi:2024:MVS

[VFP24] Giovani Valdrighi, Nivan Ferreira, and Jorge Poco. MoRe-Vis: a visual summary for spatiotemporal moving regions. IEEE Transactions on Visualization and Computer Graphics, 30(4):1927–1941, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Venkat:2022:TRP

[VGK^{+}22] Aniketh Venkat, Attila Gyulassy, Graham Kosiba, Amitesh Maiti, Henry Reinstein, Richard Gee, Peer-Timo Bremer, and Valerio Pascucci. Towards replacing physical testing of granular materials with a topologybased model. IEEE Transactions on Visualization and Computer Graphics, 28(1):76– 85, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Vidal:2021:PAS

[VGT21] Jules Vidal, Pierre Guillou, and Julien Tierny. A progressive approach to scalar field topology. IEEE Transactions on Visualization and Computer Graphics, 27(6):2833–2850, June 2021.

CODEN ITVGEA. ISSN 1077-2626.

Vermast:2023:ITA

[VH23] Alissa Vermast and Wolfgang Hürst. Introducing 3D thumbnails to access 360-degree videos in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2547–2556, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Vieth:2024:MFE

[VKT^{+}24] Alexander Vieth, Thomas Kroes, Julian Thijssen, Baldur van Lew, Jeroen Eggermont, Soumyadeep Basu, Elmar Eisemann, Anna Vilanova, Thomas Höllt, and Boudewijn Lelieveldt. ManiVault: a flexible and extensible visual analytics framework for highdimensional data. IEEE Transactions on Visualization and Computer Graphics, 30(1):175– 185, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Volmer:2023:MLP

[VLM^{+}23] Benjamin Volmer, Jen-Shuo Liu, Brandon Matthews, Ina Bornkessel-Schlesewsky, Steven Feiner, and Bruce H. Thomas. Multi-level precues for guiding tasks within and between workspaces in spatial augmented reality. IEEE Transactions on Visualization and Computer Graphics, 29(11): 4449–4459, November 2023. CODEN ITVGEA. ISSN 1077-

Velinov:2023:CWV

[VM23] Zdravko Velinov and Kenny Mitchell. Collimated whole volume light scattering in homogeneous finite media. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3145–3157, July 2023. CODEN ITVGEA. ISSN 1077-2626.

vanOnzenoodt:2023:PFV

[vOVR23] Christian van Onzenoodt, Pere-Pau Vázquez, and Timo Ropinski. Out of the plane: Flower versus star glyphs to support high-dimensional exploration in two-dimensional embeddings. IEEE Transactions on Visualization and Computer Graphics, 29(12):5468–5482, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Vasiou:2022:MRM

[VSBY22] Elena Vasiou, Konstantin Shkurko, Erik Brunvand, and Cem Yuksel. Mach-RT: a many chip architecture for high performance ray tracing. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1585–1596, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Venkatakrishnan:2024:IEA

[VVC^{+}24] Roshan Venkatakrishnan, Rohith Venkatakrishnan, Ryan Canales, Balagopal Raveendranath, Christopher C. Pagano, Andrew C. Robb, Wen-Chieh Lin, and Sabarish V. Babu. Investigating the effects of avatarization and interaction techniques on near-field mixed reality interactions with physical components. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2756–2766, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Venkatakrishnan:2023:GMH

[VVR^{+}23a] Roshan Venkatakrishnan, Rohith Venkatakrishnan, Balagopal Raveendranath, Christopher C. Pagano, Andrew C. Robb, Wen-Chieh Lin, and Sabarish V. Babu. Give me a hand: Improving the effectiveness of near-field augmented reality interactions by avatarizing users’ end effectors. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2412–2422, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Venkatakrishnan:2023:HVH

[VVR^{+}23b] Roshan Venkatakrishnan, Rohith Venkatakrishnan, Balagopal Raveendranath, Christopher C. Pagano, Andrew C. Robb, Wen-Chieh Lin, and Sabarish V. Babu. How virtual hand representations affect the perceptions of dynamic affordances in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2258–2268, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Venkatakrishnan:2024:EST

[VVR^{+}24] Rohith Venkatakrishnan, Roshan Venkatakrishnan, Balagopal Raveendranath, Ryan Canales, Dawn M. Sarno, Andrew C. Robb, Wen Chieh Lin, and Sabarish V. Babu. The effects of secondary task demands on cybersickness in active exploration virtual reality experiences. IEEE Transactions on Visualization and Computer Graphics, 30(5):2745– 2755, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Wallinger:2022:EPB

[WAA^{+}22] Markus Wallinger, Daniel Archambault, David Auber, Martin Nöllenburg, and Jaakko Peltonen. Edge-path bundling: a less ambiguous edge bundling approach. IEEE Transactions on Visualization and Computer Graphics, 28(1):313–323, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Willett:2022:PIE

[WAC^{+}22] Wesley Willett, Bon Adriel Aseniero, Sheelagh Carpendale, Pierre Dragicevic, Yvonne Jansen, Lora Oehlberg, and Petra Isenberg. Perception! immersion! empowerment! superpowers as inspiration for visualization. IEEE Transactions on Visualization and Computer Graphics, 28(1):22–32, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:HVA

[WAP^{+}21] Q. Wang, W. Alexander, J. Pegg, H. Qu, and M. Chen. HypoML: Visual analysis for hypothesis-based evaluation of machine learning models. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1417–1426, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Weiss:2021:RCE

[WAV^{+}21] M. Weiß, K. Angerbauer, A. Voit, M. Schwarzl, M. Sedlmair, and S. Mayer. Revisited: Comparison of empirical methods to evaluate visualizations supporting crafting and assembly purposes. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1204–1213, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Weidner:2023:SRV

[WBA^{+}23] Florian Weidner, Gerd Boettcher, Stephanie Arevalo Arboleda, Chenyao Diao, Luljeta Sinani, Christian Kunert, Christoph Gerhardt, Wolfgang Broll, and Alexander Raake. A systematic review on the visualization of avatars and agents in AR & VR displayed using head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2596–2606, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Weissker:2020:GTT

[WBF20] T. Weissker, P. Bimberg, and B. Froehlich. Getting there together: Group navigation in distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 26(5):1860– 1870, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Weerasinghe:2022:PVL

[WBG^{+}22] Maheshya Weerasinghe, Verena Biener, Jens Grubert, Aaron Quigley, Alice Toniolo, Klen Čopič Pucihar, and Matja Kljun. VocabulARy: Learning vocabulary in AR supported by keyword visualisations. IEEE Transactions on Visualization and Computer Graphics, 28(11):3748– 3758, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Weissker:2023:GHG

[WBG^{+}23] Tim Weissker, Pauline Bimberg, Aalok Shashidhar Gokhale, Torsten Kuhlen, and Bernd Froehlich. Gaining the high ground: Teleportation to midair targets in immersive virtual environments. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2467–2477, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Williams:2020:VMT

[WBI20] K. Williams, A. Bigelow, and K. Isaacs. Visualizing a moving target: a design study on task parallel programs in the presence of evolving data and concerns. IEEE Transactions on Visualization and Computer Graphics, 26(1):1118– 1128, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2022:UVA

[WBL^{+}22] Xumeng Wang, Chris Bryan, Yiran Li, Rusheng Pan, Yanling Liu, Wei Chen, and Kwan-Liu Ma. Umbra: a visual analysis approach for defense construction against inference attacks on sensitive information. IEEE Transactions on Visualization and Computer Graphics, 28(7):2776–2790, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Westermeier:2023:EPP

[WBLW23] Franziska Westermeier, Larissa Brübach, Marc Erich Latoschik, and Carolin Wienrich. Exploring plausibility and presence in mixed reality experiences. IEEE Transactions on Visualization and Computer Graphics, 29(5):2680–2689, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Williams:2021:AAB

[WBM21a] Niall L. Williams, Aniket Bera, and Dinesh Manocha. ARC: Alignment-based redirection controller for redirected walking in complex environments. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2535–2544, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Williams:2021:RWS

[WBM21b] Niall L. Williams, Aniket Bera, and Dinesh Manocha. Redirected walking in static and dynamic scenes using visibility polygons. IEEE Transactions on Visualization and Computer Graphics, 27(11):4267– 4277, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Wijayanto:2023:CEV

[WBPC23] Ignatius Alex Wijayanto, Sabarish V. Babu, Christopher C. Pagano, and Jung Hong Chuang. Comparing the effects of visual realism on size perception in VR versus real world viewing through physical and verbal judgments. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2721–2731, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Wehrwein:2021:SSM

[WBS21] S. Wehrwein, K. Bala, and N. Snavely. Scene summarization via motion normalization. IEEE Transactions on Visualization and Computer Graphics, 27(4):2495–2501, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Westermeier:2024:ADP

[WBWL24] Franziska Westermeier, Larissa Brübach, Carolin Wienrich, and Marc Erich Latoschik. Assessing depth perception in VR and video see-through AR: a comparison on distance judgment, performance, and preference. IEEE Transactions on Visualization and Computer Graphics, 30(5):2140–2150, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Wu:2021:ARC

[WCC^{+}21] Nannan Wu, Qianwen Chao, Yanzhen Chen, Weiwei Xu, Chen Liu, Dinesh Manocha, Wenxin Sun, Yi Han, Xinran Yao, and Xiaogang Jin. AgentDress: Realtime clothing synthesis for virtual agents using plausible deformations. IEEE Transactions on Visualization and Computer Graphics, 27(11):4107–4118, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2024:TVP

[WCCS24] Miao Wang, Ze-Yin Chen, Wen-Chuan Cai, and Frank Steinicke. Transferable virtualphysical environmental alignment with redirected walking. IEEE Transactions on Visualization and Computer Graphics, 30(3):1696–1709, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:DIV

[WCH^{+}22] Yifan Wu, Remco Chang, Joseph M. Hellerstein, Arvind Satyanarayan, and Eugene Wu. DIEL: Interactive visualization beyond the here and now. IEEE Transactions on Visualization and Computer Graphics, 28(1): 737–746, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2023:CDD

[WCH^{+}23] Can Wang, Menglei Chai, Mingming He, Dongdong Chen, and Jing Liao. Cross-domain and disentangled face manipulation with 3D guidance. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2053–2066, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:JSC

[WCJW22] Yinqiao Wang, Lu Chen, Jaemin Jo, and Yunhai Wang. Joint t-SNE for comparable projections of multiple highdimensional datasets. IEEE Transactions on Visualization and Computer Graphics, 28(1): 623–632, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Weiss:2021:VIR

[WCTW21] Sebastian Weiss, Mengyu Chu, Nils Thuerey, and Rüdiger Westermann. Volumetric isosurface rendering with deep learning-based super-resolution. IEEE Transactions on Visualization and Computer Graphics, 27(6):3064–3078, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:FBF

[WCW^{+}22] Yunhai Wang, Da Cheng, Zhirui Wang, Jian Zhang, Liang Zhou, Gaoqi He, and Oliver Deussen. F2-Bubbles: Faithful bubble set construction and flexible editing. IEEE Transactions on Visualization and Computer Graphics, 28(1):422– 432, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:SPM

[WCWQ22] Qianwen Wang, Zhutian Chen, Yong Wang, and Huamin Qu. A survey on ML4VIS: Applying machine learning advances to data visualization. IEEE Transactions on Visualization and Computer Graphics, 28(12): 5134–5153, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:CRN

[WCX21a] Z. Wang, J. Chai, and S. Xia.

Combining recurrent neural networks and adversarial training for human motion synthesis and control. IEEE Transactions on Visualization and Computer Graphics, 27(1):14–28, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:RAE

[WCX21b] Z. Wang, J. Chai, and S. Xia.

Realtime and accurate 3D eye gaze capture with DCNN-based iris and pupil segmentation. IEEE Transactions on Visualization and Computer Graphics, 27(1):190–203, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:HVA

[WCX^{+}23] Xumeng Wang, Wei Chen, Jiazhi Xia, Zhen Wen, Rongchen Zhu, and Tobias Schreck. Het-Vis: a visual analysis approach for identifying data heterogeneity in horizontal federated learning. IEEE Transactions on Visualization and Computer Graphics, 29(1):310– 319, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:STW

[WCZ^{+}20] Y. Wang, X. Chu, K. Zhang, C. Bao, X. Li, J. Zhang, C. Fu, C. Hurter, O. Deussen, and B. Lee. ShapeWordle: Tailoring wordles using shape-aware Archimedean spirals. IEEE Transactions on Visualization and Computer Graphics, 26(1): 991–1000, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wei:2023:GGS

[WCZ^{+}23] Mingqiang Wei, Honghua Chen, Yingkui Zhang, Haoran Xie, Yanwen Guo, and Jun Wang. GeoDualCNN: Geometry-supporting dual convolutional neural network for noisy point clouds. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1357–1370, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2023:DVA

[WDC^{+}23] Aoyu Wu, Dazhen Deng, Furui Cheng, Yingcai Wu, Shixia Liu, and Huamin Qu. In defence of visual analytics systems: Replies to critics. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1026–1036, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Waldner:2020:CRL

[WDG^{+}20] M. Waldner, A. Diehl, D. Gračanin, R. Splechtna, C. Delrieux, and K. Matković. A comparison of radial and linear charts for visualizing daily patterns. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1033–1042, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wallinger:2023:LZC

[WDN23] Markus Wallinger, Alexander Dobler, and Martin Nöllenburg. LinSets.zip: Compressing linear set diagrams. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2875–2887, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Walchshofer:2024:TCD

[WDSM24] Conny Walchshofer, Vaishal Dhanoa, Marc Streit, and Miriah Meyer. Transitioning to a commercial dashboarding system: Socio-technical observations and opportunities. IEEE Transactions on Visualization and Computer Graphics, 30(1): 381–391, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2023:OVA

[WDX^{+}23] Yihong Wu, Dazhen Deng, Xiao Xie, Moqi He, Jie Xu, Hongzeng Zhang, Hui Zhang, and Yingcai Wu. OBTracker: Visual analytics of off-ball movements in basketball. IEEE Transactions on Visualization and Computer Graphics, 29(1):929–939, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:AIP

[WEM^{+}21] X. Wang, A. Eiselmayer, W. E. Mackay, K. Hornbaek, and C. Wacharamanotham. Argus: Interactive a priori power analysis. IEEE Transactions on Visualization and Computer Graphics, 27(2):432–442, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Weissker:2021:GNG

[WF21] Tim Weissker and Bernd Froehlich. Group navigation for guided tours in distributed virtual environments. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2524–2534, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Wei:2021:SGN

[WFC21] Mingqiang Wei, Yidan Feng, and Honghua Chen. Selective guidance normal filter for geometric texture removal. IEEE Transactions on Visualization and Computer Graphics, 27(12):4469–4482, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Weissker:2024:TSM

[WFK24] Tim Weissker, Matthis Franzgrote, and Torsten Kuhlen. Try this for size: Multi-scale teleportation in immersive virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2298– 2308, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Walny:2020:DCE

[WFW^{+}20] J. Walny, C. Frisson, M. West, D. Kosminsky, S. Knudsen, S. Carpendale, and W. Willett. Data changes everything: Challenges and opportunities in data visualization design handoff. IEEE Transactions on Visualization and Computer Graphics, 26(1):12– 22, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2021:PSV

[WFY^{+}21] Miao Wang, Xiao-Nan Fang, Guo-Wei Yang, Ariel Shamir, and Shi-Min Hu. Prominent structures for video analysis and editing. IEEE Transactions on Visualization and Computer Graphics, 27(7):3305– 3317, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Werner:2021:UTP

[WG21] Kilian Werner and Christoph Garth. Unordered task-parallel augmented merge tree construction. IEEE Transactions on Visualization and Computer Graphics, 27(8):3585–3596, August 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:PMS

[WGH20] B. Wang, L. Ge, and N. Holzschuch.

Precomputed multiple scattering for rapid light simulation in participating media. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2456–2470, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Wu:2024:SDS

[WGH^{+}24] Guande Wu, Shunan Guo, Jane Hoffswell, Gromit Yeuk-Yin Chan, Ryan A. Rossi, and Eunyee Koh. Socrates: Data story generation via adaptive machine-guided elicitation of user feedback. IEEE Transactions on Visualization and Computer Graphics, 30(1):131– 141, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wu:2024:RAB

[WGM^{+}24] Yifan Wu, Ziyang Guo, Michalis Mamakos, Jason Hartline, and Jessica Hullman. The rational agent benchmark for data visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):338–347, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Williams:2020:UMU

[WGO20] A. S. Williams, J. Garcia, and F. Ortega. Understanding multimodal user gesture and speech behavior for object manipulation in augmented reality using elicitation. IEEE Transactions on Visualization and Computer Graphics, 26(12):3479– 3489, December 2020. CODEN ITVGEA. ISSN 1077-2626. See errata [WGO22].

Williams:2022:EUM

[WGO22] Adam S. Williams, Jason Garcia, and Francisco Ortega. Errata to “Understanding Multimodal User Gesture and Speech Behavior for Object Manipulation in Augmented Reality Using Elicitation”. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2808, July 2022. CODEN ITVGEA. ISSN 1077-2626. See [WGO20].

Wurster:2023:DHS

[WGS^{+}23] Skylar W. Wurster, Hanqi Guo, Han-Wei Shen, Tom Peterka, and Jiayi Xu. Deep hierarchical super resolution for scientific data. IEEE Transactions on Visualization and Computer Graphics, 29(12):5483– 5495, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Wen:2024:VNR

[WGS^{+}24] Elliott Wen, Chitralekha Gupta, Prasanth Sasikumar, Mark Billinghurst, James Wilmott, Emily Skow, Arindam Dey, and Suranga Nanayakkara. VR.net: a real-world dataset for virtual reality motion sickness research. IEEE Transactions on Visualization and Computer Graphics, 30(5):2330–2336, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Wilson:2022:DPM

[WH22a] Austin Wilson and Hong Hua.

Design of a pupil-matched occlusion-capable optical seethrough wearable display. IEEE Transactions on Visualization and Computer Graphics, 28(12):4113–4126, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Worrallo:2022:ROB

[WH22b] Adam Grant Worrallo and Thomas Hartley. Robust optical based hand interaction for virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4186–4197, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:DSA

[WH23] Chaoli Wang and Jun Han. DL4SciVis: a state-of-the-art survey on deep learning for scientific visualization. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3714–3733, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2023:NDE

[WHBC23] Xue-Liang Wu, Huan-Chang Hung, Sabarish V. Babu, and Jung-Hong Chuang. Novel design and evaluation of redirection controllers using optimized alignment and artificial potential field. IEEE Transactions on Visualization and Computer Graphics, 29(11):4556– 4566, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:ENM

[WHC^{+}23] Qianwen Wang, Kexin Huang, Payal Chandak, Marinka Zitnik, and Nils Gehlenborg. Extending the nested model for user-centric XAI: a design study on GNN-based drug repurposing. IEEE Transactions on Visualization and Computer Graphics, 29(1):1266– 1276, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:MVE

[WHJ^{+}22] Xingbo Wang, Jianben He, Zhihua Jin, Muqiao Yang, Yong Wang, and Huamin Qu. M2Lens: Visualizing and explaining multimodal models for sentiment analysis. IEEE Transactions on Visualization and Computer Graphics, 28(1): 802–812, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2024:CVU

[WHJ^{+}24] Xingbo Wang, Renfei Huang, Zhihua Jin, Tianqing Fang, and Huamin Qu. Commonsense-VIS: Visualizing and understanding commonsense reasoning capabilities of natural language models. IEEE Transactions on Visualization and Computer Graphics, 30(1):273– 283, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wentzel:2020:CBS

[WHL^{+}20] A. Wentzel, P. Hanula, T. Luciani, B. Elgohari, H. Elhalawani, G. Canahuate, D. Vock, C. D. Fuller, and G. E. Marai. Cohort-based T-SSIM visual computing for radiation therapy prediction and exploration. IEEE Transactions on Visualization and Computer Graphics, 26(1):949–959, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2023:TNL

[WHS^{+}23] Yun Wang, Zhitao Hou, Leixian Shen, Tongshuang Wu, Jiaqi Wang, He Huang, Haidong Zhang, and Dongmei Zhang. Towards natural languagebased visualization authoring. IEEE Transactions on Visualization and Computer Graphics, 29(1):1222–1232, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:STM

[WHSZ21] H. Wang, E. S. L. Ho, H. P. H. Shum, and Z. Zhu. Spatiotemporal manifold learning for human motions via longhorizon modeling. IEEE Transactions on Visualization and Computer Graphics, 27(1):216– 227, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Walchshofer:2023:PEB

[WHX^{+}23] Conny Walchshofer, Andreas Hinterreiter, Kai Xu, Holger Stitz, and Marc Streit. Provectories: Embedding-based analysis of interaction provenance data. IEEE Transactions on Visualization and Computer Graphics, 29(12):4816– 4831, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:SMR

[WHY^{+}23] Liu Wang, Mengjie Huang, Rui Yang, Hai-Ning Liang, Ji Han, and Ying Sun. Survey of movement reproduction in immersive virtual rehabilitation. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2184–2202, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Wilson:2024:PPG

[WIP^{+}24] Ethan Wilson, Azim Ibragimov, Michael J. Proulx, Sai Deep Tetali, Kevin Butler, and Eakta Jain. Privacy-preserving gaze data streaming in immersive interactive virtual reality: Robustness and user experience. IEEE Transactions on Visualization and Computer Graphics, 30(5):2257–2268, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Weiss:2022:LAS

[WITW22] Sebastian Weiss, Mustafa IšIk, Justus Thies, and Rüdiger Westermann. Learning adaptive sampling and reconstruction for volume visualization. IEEE Transactions on Visualization and Computer Graphics, 28(7):2654–2667, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:VQI

[WJBB22] Yao Wang, Chuhan Jiao, Mihai Bâce, and Andreas Bulling.

VisRecall: Quantifying information visualisation recallability via question answering. IEEE Transactions on Visualization and Computer Graphics, 28(12):4995–5005, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wallinger:2021:RAS

[WJKN21] Markus Wallinger, Ben Jacobsen, Stephen Kobourov, and Martin Nöllenburg. On the readability of abstract set visualizations. IEEE Transactions on Visualization and Computer Graphics, 27(6):2821– 2832, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:DDL

[WJW^{+}20] Y. Wang, Z. Jin, Q. Wang, W. Cui, T. Ma, and H. Qu. DeepDrawing: a deep learning approach to graph drawing. IEEE Transactions on Visualization and Computer Graphics, 26(1):676–686, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wesslen:2022:EUV

[WKMD22] Ryan Wesslen, Alireza Karduni, Douglas Markant, and Wenwen Dou. Effect of uncertainty visualizations on myopic loss aversion and the equity premium puzzle in retirement investment decisions. IEEE Transactions on Visualization and Computer Graphics, 28(1):454–464, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Warchol:2023:VVS

[WKN^{+}23] Simon Warchol, Robert Krueger, Ajit Johnson Nirmal, Giorgio Gaglia, Jared Jessup, Cecily C. Ritch, John Hoffer, Jeremy Muhlich, Megan L. Burger, Tyler Jacks, Sandro Santagata, Peter K. Sorger, and Hanspeter Pfister. Visinity: Visual spatial neighborhood analysis for multiplexed tissue imaging data. IEEE Transactions on Visualization and Computer Graphics, 29(1):106–116, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:VVC

[WL20] Z. Wang and F. Lu. VoxSegNet:

Volumetric CNNs for semantic part segmentation of 3D shapes. IEEE Transactions on Visualization and Computer Graphics, 26(9):2919–2930, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:SRS

[WLC^{+}23] Ziyao Wang, Chiyi Liu, Jialiang Chen, Yao Yao, Dazheng Fang, Zhiyi Shi, Rui Yan, Yiye Wang, KanJian Zhang, Hai Wang, and Haikun Wei. Strolling in room-scale VR: Hex-Core-MK1 omnidirectional treadmill. IEEE Transactions on Visualization and Computer Graphics, 29(12):5538–5555, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:EPB

[WLF^{+}22] Zhibo Wang, Jingwang Ling, Chengzeng Feng, Ming Lu, and Feng Xu. Emotion-preserving blendshape update with realtime face tracking. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2364–2375, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:TVA

[WLG^{+}22] Jiang Wu, Dongyu Liu, Ziyang Guo, Qingyang Xu, and Yingcai Wu. TacticFlow: Visual analytics of ever-changing tactics in racket sports. IEEE Transactions on Visualization and Computer Graphics, 28(1): 835–845, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2023:RIP

[WLGW23] Jiang Wu, Dongyu Liu, Ziyang Guo, and Yingcai Wu. RASI-PAM: Interactive pattern mining of multivariate event sequences in racket sports. IEEE Transactions on Visualization and Computer Graphics, 29(1): 940–950, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2024:VVC

[WLH24] Yuxi Wang, Haibin Ling, and Bingyao Huang. Vi-Comp: Video compensation for projector-camera systems. IEEE Transactions on Visualization and Computer Graphics, 30(5):2347–2356, May 2024. CODEN ITVGEA. ISSN 1077-

Wang:2022:IVE

[WLS^{+}22] Yifang Wang, Hongye Liang, Xinhuan Shu, Jiachen Wang, Ke Xu, Zikun Deng, Cameron Campbell, Bijia Chen, Yingcai Wu, and Huamin Qu. Interactive visual exploration of longitudinal historical career mobility data. IEEE Transactions on Visualization and Computer Graphics, 28 (10):3441–3455, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Wen:2024:QVA

[WLT^{+}24] Zhen Wen, Yihan Liu, Siwei Tan, Jieyi Chen, Minfeng Zhu, Dongming Han, Jianwei Yin, Mingliang Xu, and Wei Chen. Quantivine: a visualization approach for large-scale quantum circuit representation and analysis. IEEE Transactions on Visualization and Computer Graphics, 30(1):573–583, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:WWH

[WLZ^{+}23] Junhong Wang, Yun Li, Zhaoyu Zhou, Chengshun Wang, Yijie Hou, Li Zhang, Xiangyang Xue, Michael Kamp, Xiaolong Luke Zhang, and Siming Chen. When, where and how does it fail? A spatialtemporal visual analytics approach for interpretable object detection in autonomous driving. IEEE Transactions on Visualization and Computer Graphics, 29(12):5033– 5049, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:DVC

[WM23] Jun Wang and Klaus Mueller. DOMINO: Visual causal reasoning with time-dependent phenomena. IEEE Transactions on Visualization and Computer Graphics, 29(12):5342– 5356, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Weidner:2023:ESS

[WMB23] Florian Weidner, Jana E. Maier, and Wolfgang Broll. Eating, smelling, and seeing: Investigating multisensory integration and (in)congruent stimuli while eating in VR. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2423–2433, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:TSB

[WMH^{+}22] Qianwen Wang, Tali Mazor, Theresa A. Harbig, Ethan Cerami, and Nils Gehlenborg. ThreadStates: State-based visual analysis of disease progression. IEEE Transactions on Visualization and Computer Graphics, 28(1):238–247, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:TTV

[WMH^{+}23] Jiachen Wang, Ji Ma, Kangping Hu, Zheng Zhou, Hui Zhang, Xiao Xie, and Yingcai Wu. Tac-Trainer: a visual analytics system for IoT-based racket sports training. IEEE Transactions on Visualization and Computer Graphics, 29(1):951–961, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wilson:2023:LRC

[WMMB23] Graham Wilson, Mark McGill, Daniel Medeiros, and Stephen Brewster. A lack of restraint: Comparing virtual reality interaction techniques for constrained transport seating. IEEE Transactions on Visualization and Computer Graphics, 29(5):2390–2400, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:PDV

[WMW^{+}22] Xingbo Wang, Yao Ming, Tongshuang Wu, Haipeng Zeng, Yong Wang, and Huamin Qu. DeHumor: Visual analytics for decomposing humor. IEEE Transactions on Visualization and Computer Graphics, 28(12):4609–4623, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wei:2020:EPB

[WMZ^{+}20] Y. Wei, H. Mei, Y. Zhao, S. Zhou, B. Lin, H. Jiang, and W. Chen. Evaluating perceptual bias during geometric scaling of scatterplots. IEEE Transactions on Visualization and Computer Graphics, 26(1): 321–331, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wald:2022:MEE

[WMZ22] Ingo Wald, Nate Morrical, and Stefan Zellmann. A memory efficient encoding for ray tracing large unstructured data. IEEE Transactions on Visualization and Computer Graphics, 28(1): 583–592, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wenskovitch:2021:EGS

[WN21] J. Wenskovitch and C. North. An examination of grouping and spatial organization tasks for high-dimensional data exploration. IEEE Transactions on Visualization and Computer Graphics, 27(2):1742– 1752, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Wall:2022:LRG

[WNC^{+}22] Emily Wall, Arpit Narechania, Adam Coscia, Jamal Paden, and Alex Endert. Left, right, and gender: Exploring interaction traces to mitigate human biases. IEEE Transactions on Visualization and Computer Graphics, 28(1):966–975, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:MLA

[WNV22] Hsiang-Yun Wu, Martin Nöllenburg, and Ivan Viola. Multi-level area balancing of clustered graphs. IEEE Transactions on Visualization and Computer Graphics, 28(7):2682–2696, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Williams:2022:IRD

[WO22] Adam S. Williams and Francisco R. Ortega. The impacts of referent display on gesture and speech elicitation. IEEE Transactions on Visualization and Computer Graphics, 28(11):3885–3895, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Wexler:2020:WIT

[WPB^{+}20] J. Wexler, M. Pushkarna, T. Bolukbasi, M. Wattenberg, F. Viégas, and J. Wilson. The what-if tool: Interactive probing of machine learning models. IEEE Transactions on Visualization and Computer Graphics, 26(1):56–65, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2022:SSV

[WPL^{+}22] Yifang Wang, Tai-Quan Peng, Huihua Lu, Haoren Wang, Xiao Xie, Huamin Qu, and Yingcai Wu. Seek for success: a visualization approach for understanding the dynamics of academic careers. IEEE Transactions on Visualization and Computer Graphics, 28(1):475– 485, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wu:2021:SVR

[WPNK21] Erwin Wu, Mitski Piekenbrock, Takuto Nakumura, and Hideki Koike. SPinPong — virtual reality table tennis skill acquisition using visual, haptic and temporal cues. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2566–2576, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Wetzels:2024:MTG

[WPTG24] Florian Wetzels, Mathieu Pont, Julien Tierny, and Christoph Garth. Merge tree geodesics and barycenters with path mappings. IEEE Transactions on Visualization and Computer Graphics, 30(1):1095– 1105, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:NHP

[WPZ^{+}23] Kangkan Wang, Sida Peng, Xiaowei Zhou, Jian Yang, and Guofeng Zhang. NerfCap: Human performance capture with dynamic neural radiance fields. IEEE Transactions on Visualization and Computer Graphics, 29(12):5097–5110, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Wu:2020:MAV

[WQ20] A. Wu and H. Qu. Multimodal analysis of video collections: Visual exploration of presentation techniques in TED talks. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2429–2442, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Weerasinghe:2022:PAE

[WQP^{+}22] Maheshya Weerasinghe, Aaron Quigley, Klen Čopič Pucihar, Alice Toniolo, Angela Miguel, and Matja Kljun. Arigat: Effects of adaptive guidance on engagement and performance in augmented reality learning environments. IEEE Transactions on Visualization and Computer Graphics, 28(11):3737– 3747, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2024:IVA

[WQQ^{+}24] Yifang Wang, Yifan Qian, Xiaoyu Qi, Nan Cao, and Dashun Wang. InnovationInsights: a visual analytics approach for understanding the dual frontiers of science and technology. IEEE Transactions on Visualization and Computer Graphics, 30(1): 518–528, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Woodward:2023:ARU

[WR23] Julia Woodward and Jaime Ruiz. Analytic review of using augmented reality for situational awareness. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2166–2183, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:IDC

[WRC^{+}22] Zezhong Wang, Hugo Romat, Fanny Chevalier, Nathalie Henry Riche, Dave Murray-Rust, and Benjamin Bach. Interactive data comics. IEEE Transactions on Visualization and Computer Graphics, 28(1):944– 954, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:DCR

[WRZ^{+}21] Z. Wang, J. Ritchie, J. Zhou, F. Chevalier, and B. Bach. Data comics for reporting controlled user studies in humancomputer interaction. IEEE Transactions on Visualization and Computer Graphics, 27(2): 967–977, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Walch:2020:LGI

[WSL^{+}20] A. Walch, M. Schwärzler, C. Luksch, E. Eisemann, and T. Gschwandtner. LightGuider: Guiding interactive lighting design using suggestions, provenance, and quality visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):569–578, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2024:OVV

[WSL^{+}24] Jialin Wang, Rongkai Shi, Xiaodong Li, Yushi Wei, and Hai-Ning Liang. Omnidirectional virtual visual acuity: a user-centric visual clarity metric for virtual reality headmounted displays and environments. IEEE Transactions on Visualization and Computer Graphics, 30(5):2033– 2043, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Wagner:2021:CCV

[WSN21] Jorge Wagner, Wolfgang Stuerzlinger, and Luciana Nedel.

Comparing and combining virtual hand and virtual ray pointer interactions for data manipulation in immersive analytics. IEEE Transactions on Visualization and Computer Graphics, 27(5):2513– 2523, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Wagner:2022:EEM

[WSN22] Jorge Wagner, Wolfgang Stuerzlinger, and Luciana Nedel. The effect of exploration mode and frame of reference in immersive analytics. IEEE Transactions on Visualization and Computer Graphics, 28(9):3252– 3264, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:DAG

[WSZ^{+}20] Y. Wang, Z. Sun, H. Zhang, W. Cui, K. Xu, X. Ma, and D. Zhang. DataShot: Automatic generation of fact sheets from tabular data. IEEE Transactions on Visualization and Computer Graphics, 26(1):895– 905, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2023:EFR

[WSZ^{+}23] Jialin Wang, Rongkai Shi, Wenxuan Zheng, Weijie Xie, Dominic Kao, and Hai-Ning Liang. Effect of frame rate on user experience, performance, and simulator sickness in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5):

2478–2488, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Wu:2021:MTW

[WTD^{+}21] A. Wu, W. Tong, T. Dwyer, B. Lee, P. Isenberg, and H. Qu. MobileVisFixer: Tailoring web visualizations for mobile phones leveraging an explainable reinforcement learning framework. IEEE Transactions on Visualization and Computer Graphics, 27(2):464–474, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2024:DFA

[WTL24] Chenglong Wang, John Thompson, and Bongshin Lee. Data Formulator: AI-Powered conceptdriven visualization authoring. IEEE Transactions on Visualization and Computer Graphics, 30(1):1128–1138, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:CEL

[WTS^{+}21] Z. J. Wang, R. Turko, O. Shaikh, H. Park, N. Das, F. Hohman, M. Kahng, and D. H. Polo Chau. CNN Explainer: Learning convolutional neural networks with interactive visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2):1396– 1406, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:WMK

[WTY^{+}22] Kui Wu, Marco Tarini, Cem Yuksel, James McCann, and Xifeng Gao. Wearable 3D machine knitting: Automatic generation of shaped knit sheets to cover real-world objects. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3180–3192, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:VCA

[Wu22] Eugene Wu. View composition algebra for ad hoc comparison. IEEE Transactions on Visualization and Computer Graphics, 28(6):2470–2485, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:EPP

[WW22a] Wencheng Wang and Shengchun Wang. Efficient point-in-polygon tests by grids without the trouble of tuning the grid resolutions. IEEE Transactions on Visualization and Computer Graphics, 28(12):4073– 4084, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Weiss:2022:DDV

[WW22b] Sebastian Weiss and Rüdiger Westermann. Differentiable direct volume rendering. IEEE Transactions on Visualization and Computer Graphics, 28(1): 562–572, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2021:TMV

[WWC^{+}21] Jiachen Wang, Jiang Wu, Anqi Cao, Zheng Zhou, Hui Zhang, and Yingcai Wu. Tac-Miner:

Visual tactic mining for multiple table tennis matches. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2770–2782, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Wu:2021:DOC

[WWG21] J. Wu, W. Wang, and X. Gao. Design and optimization of conforming lattice structures. IEEE Transactions on Visualization and Computer Graphics, 27(1):43–56, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:IRS

[WWL^{+}20] Y. Wang, Z. Wang, T. Liu, M. Correll, Z. Cheng, O. Deussen, and M. Sedlmair. Improving the robustness of scagnostics. IEEE Transactions on Visualization and Computer Graphics, 26(1): 759–769, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Woodin:2022:CMG

[WWP22] Greg Woodin, Bodo Winter, and Lace Padilla. Conceptual metaphor and graphical convention influence the interpretation of line graphs. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1209–1221, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Wen:2020:PAD

[WWR^{+}20] X. Wen, M. Wang, C. Richardt, Z.-Y. Chen, and S.-M. Hu. Photorealistic audio-driven video portraits. IEEE Transactions on Visualization and Computer Graphics, 26(12):3457– 3466, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Whitlock:2020:DMI

[WWS20] M. Whitlock, K. Wu, and D. A.

Szafir. Designing for mobile and immersive visual analytics in the field. IEEE Transactions on Visualization and Computer Graphics, 26(1):503–513, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wu:2022:PAS

[WWS^{+}22] Aoyu Wu, Yun Wang, Xinhuan Shu, Dominik Moritz, Weiwei Cui, Haidong Zhang, Dongmei Zhang, and Huamin Qu. AI4VIS: Survey on artificial intelligence approaches for data visualization. IEEE Transactions on Visualization and Computer Graphics, 28(12): 5049–5070, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:RBE

[WWSS20] K.-C. Wang, T.-H. Wei, N. Shareef, and H.-W. Shen. Raybased exploration of large timevarying volume data using per-ray proxy distributions. IEEE Transactions on Visualization and Computer Graphics, 26(11):3299–3313, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Wang:2024:TAN

[WWW^{+}24] Chen Wang, Guangshun Wei, Guodong Wei, Wenping Wang, and Yuanfeng Zhou. Tooth alignment network based on landmark constraints and hierarchical graph structure. IEEE Transactions on Visualization and Computer Graphics, 30(2): 1457–1469, February 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2022:MDA

[WWZ^{+}22] Aoyu Wu, Yun Wang, Mengyu Zhou, Xinyi He, Haidong Zhang, Huamin Qu, and Dongmei Zhang. MultiVision: Designing analytical dashboards with deep learning based recommendation. IEEE Transactions on Visualization and Computer Graphics, 28(1):162–172, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:LDM

[WWZ^{+}23a] Junpeng Wang, Liang Wang, Yan Zheng, Chin-Chia Michael Yeh, Shubham Jain, and Wei Zhang. Learning-from-disagreement: a model comparison and visual analytics framework. IEEE Transactions on Visualization and Computer Graphics, 29(9):3809– 3825, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:UFL

[WWZ23b] Shaodong Wang, Wencheng Wang, and Hui Zhao. Using foliation leaves to extract Reeb graphs on surfaces. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2117–2131, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Wu:2022:QFG

[WWZP22] Jian Wu, Lili Wang, Hui Zhang, and Voicu Popescu. Quantifiable fine-grain occlusion removal assistance for efficient VR exploration. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3154–3167, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:VAD

[WXC^{+}21] Q. Wang, Z. Xu, Z. Chen, Y. Wang, S. Liu, and H. Qu. Visual analysis of discrimination in machine learning. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1470–1480, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Wu:2024:LVA

[WXG^{+}24] Yuchen Wu, Yuansong Xu, Shenghan Gao, Xingbo Wang, Wenkai Song, Zhiheng Nie, Xiaomeng Fan, and Quan Li. LiveRetro: Visual analytics for strategic retrospect in livestream e-commerce. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1117–1127, January 2024. CO-DEN ITVGEA. ISSN 1077-

Wurster:2024:APM

[WXS^{+}24] Skylar W. Wurster, Tianyu Xiong, Han-Wei Shen, Hanqi Guo, and Tom Peterka. Adaptively placed multi-grid scene representation networks for large-scale data visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1):965–974, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Wang:2020:ISA

[WXW^{+}20] Y. Wang, M. Xue, Y. Wang, X. Yan, B. Chen, C. Fu, and C. Hurter. Interactive structure-aware blending of diverse edge bundling visualizations. IEEE Transactions on Visualization and Computer Graphics, 26(1):687–696, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2020:VGD

[WYC^{+}20] Q. Wang, J. Yuan, S. Chen, H. Su, H. Qu, and S. Liu. Visual genealogy of deep neural networks. IEEE Transactions on Visualization and Computer Graphics, 26(11):3340– 3352, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Wang:2024:VEA

[WYIS24] Shaoyu Wang, Hang Yan, Katherine E. Isaacs, and Yifan Sun. Visual exploratory analysis for designing largescale network-on-chip architectures: a domain expert-led design study. IEEE Transactions on Visualization and Computer Graphics, 30(4):1970– 1983, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:PDA

[WYS^{+}22] Xuanyu Wang, Hui Ye, Christian Sandor, Weizhan Zhang, and Hongbo Fu. Predict-and-drive: Avatar motion adaption in room-scale augmented reality telepresence with heterogeneous spaces. IEEE Transactions on Visualization and Computer Graphics, 28(11):3705– 3714, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2021:VND

[WYZ^{+}21] Y. Wang, G. Yan, H. Zhu, S. Buch, Y. Wang, E. M. Haacke, J. Hua, and Z. Zhong. VC-Net: Deep volume-composition networks for segmentation and visualization of highly sparse and noisy image data. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1301–1311, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2023:AGA

[WZC^{+}23] Shidong Wang, Wei Zeng, Xi Chen, Yu Ye, Yu Qiao, and Chi-Wing Fu. ActFloor-GAN: Activity-guided adversarial networks for humancentric floorplan design. IEEE Transactions on Visualization and Computer Graphics, 29(3): 1610–1624, March 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2020:TST

[WZD^{+}20] J. Wang, K. Zhao, D. Deng, A. Cao, X. Xie, Z. Zhou, H. Zhang, and Y. Wu. Tac-Simur: Tactic-based simulative visual analytics of table tennis. IEEE Transactions on Visualization and Computer Graphics, 26(1):407–417, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Weng:2021:TBB

[WZD^{+}21] D. Weng, C. Zheng, Z. Deng, M. Ma, J. Bao, Y. Zheng, M. Xu, and Y. Wu. Towards better bus networks: a visual analytics approach. IEEE Transactions on Visualization and Computer Graphics, 27(2): 817–827, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Wang:2020:DFR

[WZH20] Y. Wang, Z. Zhong, and J. Hua. DeepOrganNet: On-the-fly reconstruction and visualization of 3D / 4D lung models from single-view projections by deep deformation network. IEEE Transactions on Visualization and Computer Graphics, 26(1): 960–970, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Wang:2022:GVC

[WZL22] Zhimin Wang, Yuxin Zhao, and Feng Lu. Gaze-vergence-controlled see-through vision in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(11): 3843–3853, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Wald:2021:RTS

[WZU^{+}21] I. Wald, S. Zellmann, W. Usher, N. Morrical, U. Lang, and V. Pascucci. Ray tracing structured AMR data using ExaBricks. IEEE Transactions on Visualization and Computer Graphics, 27(2):625– 634, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Wen:2023:EVL

[WZW^{+}23] Zhen Wen, Wei Zeng, Luoxuan Weng, Yihan Liu, Mingliang Xu, and Wei Chen. Effects of view layout on situated analytics for multiple-view representations in immersive visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):440–450, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Wang:2022:VAR

[WZY^{+}22] Junpeng Wang, Wei Zhang, Hao Yang, Chin-Chia Michael Yeh, and Liang Wang. Visual analytics for RNN-based deep reinforcement learning. IEEE Transactions on Visualization and Computer Graphics, 28(12):4141–4155, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Wang:2023:RGW

[WZZ^{+}23] Chen Wang, Song-Hai Zhang, Yizhuo Zhang, Stefanie Zollmann, and Shi-Min Hu. On rotation gains within and beyond perceptual limitations for seated VR. IEEE Transactions on Visualization and Computer Graphics, 29(7):3380– 3391, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Xu:2021:FRD

[XCG^{+}21] L. Xu, W. Cheng, K. Guo, L. Han, Y. Liu, and L. Fang. FlyFusion: Realtime dynamic scene reconstruction using a flying depth camera. IEEE Transactions on Visualization and Computer Graphics, 27(1):68– 82, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Xu:2024:BUB

[XCG^{+}24] Sen-Zhe Xu, Fiona Xiao Yu Chen, Ran Gong, Fang-Lue Zhang, and Song-Hai Zhang. BiRD: Using bidirectional rotation gain differences to redirect users during back-and-forth head turns in walking. IEEE Transactions on Visualization and Computer Graphics, 30(5):2693–2702, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Xie:2020:VOD

[XCK20] W. Xie, O. Chkrebtii, and S. Kurtek. Visualization and outlier detection for multivariate elastic curve data. IEEE Transactions on Visualization and Computer Graphics, 26(11):3353–3364, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2020:BAP

[XCLF20] C. Xiong, C. R. Ceja, C. J. H.

Ludwig, and S. Franconeri. Biased average position estimates in line and bar graphs: Underestimation, overestimation, and perceptual pull. IEEE Transactions on Visualization and Computer Graphics, 26(1): 301–310, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiong:2022:RAS

[XCSJ22] Weidan Xiong, Chong Mo Cheung, Pedro V. Sander, and Ajay Joneja. Rationalizing architectural surfaces based on clustering of joints. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4274–4288, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Xing:2024:VHB

[XDBAR24] Yiwen Xing, Cristina Dondi, Rita Borgo, and Alfie Abdul-Rahman. Visualizing historical book trade data: an iterative design study with close collaboration with domain experts. IEEE Transactions on Visualization and Computer Graphics, 30(1):540–550, January 2024.

CODEN ITVGEA. ISSN 1077-2626.

Xu:2022:GDD

[XDH^{+}22] Jiayi Xu, Soumya Dutta, Wenbin He, Joachim Moortgat, and Han-Wei Shen. Geometrydriven detection, tracking and visual analysis of viscous and gravitational fingers. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1514–1528, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Xie:2021:VAA

[XDW21] X. Xie, F. Du, and Y. Wu. A visual analytics approach for exploratory causal analysis: Exploration, validation, and applications. IEEE Transactions on Visualization and Computer Graphics, 27(2):1448– 1458, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Xia:2024:LLE

[XEXW24] Menghan Xia, Jose Echevarria, Minshan Xie, and Tien-Tsin Wong. LF2MV: Learning an editable meta-view towards light field representation. IEEE Transactions on Visualization and Computer Graphics, 30(3):1672–1684, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Xu:2021:ADP

[XF21] Zhongwei Xu and Alessandro Foi. Anisotropic denoising of 3D point clouds by aggregation of multiple surfaceadaptive estimates. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2851–2868, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2023:VSD

[XFD^{+}23] Kai Xiong, Siwei Fu, Guoming Ding, Zhongsu Luo, Rong Yu, Wei Chen, Hujun Bao, and Yingcai Wu. Visualizing the scripts of data wrangling with somnus. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2950–2964, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2021:MPC

[XFF^{+}21] Guoming Xiong, Qiang Fu, Hongbo Fu, Bin Zhou, Guoliang Luo, and Zhigang Deng. Motion planning for convertible indoor scene layout design. IEEE Transactions on Visualization and Computer Graphics, 27(12):4413–4424, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Xu:2021:ALB

[XGS^{+}21] Jiayi Xu, Hanqi Guo, Han-Wei Shen, Mukund Raj, Xueyun Wang, Xueqiao Xu, Zhehui Wang, and Tom Peterka. Asynchronous and load-balanced union–find for distributed and parallel scientific data visualization and analysis. IEEE Transactions on Visualization and Computer Graphics, 27(6): 2808–2820, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Xu:2023:RLL

[XGS^{+}23] Jiayi Xu, Hanqi Guo, Han-Wei Shen, Mukund Raj, Skylar W. Wurster, and Tom Peterka. Reinforcement learning for load-balanced parallel particle tracing. IEEE Transactions on Visualization and Computer Graphics, 29(6): 3052–3066, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Xu:2024:SCB

[XHFZ24] Sen-Zhe Xu, Kui Huang, Cheng-Wei Fan, and Song-Hai Zhang. Spatial contraction based on velocity variation for natural walking in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5):2444– 2453, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Xia:2023:IVC

[XHL^{+}23] Jiazhi Xia, Linquan Huang, Weixing Lin, Xin Zhao, Jing Wu, Yang Chen, Ying Zhao, and Wei Chen. Interactive visual cluster analysis by contrastive dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 29(1):734–744, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Xiao:2024:LCS

[XHL^{+}24] Shishi Xiao, Suizi Huang, Yue Lin, Yilin Ye, and Wei Zeng. Let the chart spark: Embedding semantic context into chart with text-to-image generative model. IEEE Transactions on Visualization and Computer Graphics, 30(1):284– 294, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Xia:2024:PFS

[XHS^{+}24] Jiazhi Xia, Linquan Huang, Yiping Sun, Zhiwei Deng, Xiaolong Luke Zhang, and Minfeng Zhu. A parallel framework for streaming dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 30(1):142–152, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Xiang:2021:MOS

[XJZ^{+}21] Xiaojun Xiang, Hanqing Jiang, Guofeng Zhang, Yihao Yu, Chenchen Li, Xingbin Yang, Danpeng Chen, and Hujun Bao. Mobile3DScanner: an online 3D scanner for high-quality object reconstruction with a mobile device. IEEE Transactions on Visualization and Computer Graphics, 27(11):4245– 4255, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2023:RSD

[XLF^{+}23] Kai Xiong, Zhongsu Luo, Siwei Fu, Yongheng Wang, Mingliang Xu, and Yingcai Wu. Revealing the semantics of data wrangling scripts with comantics. IEEE Transactions on Visualization and Computer Graphics, 29(1): 117–127, January 2023. CO-DEN ITVGEA. ISSN 1077-

Xing:2020:ASV

[XLL^{+}20] G. Xing, Y. Liu, H. Ling, X. Granier, and Y. Zhang. Automatic spatially varying illumination recovery of indoor scenes based on a single RGB-D image. IEEE Transactions on Visualization and Computer Graphics, 26(4):1672– 1685, April 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xu:2022:MRT

[XLL^{+}22] Sen-Zhe Xu, Tian-Qi Liu, Jia-Hong Liu, Stefanie Zollmann, and Song-Hai Zhang. Making resets away from targets: POI aware redirected walking. IEEE Transactions on Visualization and Computer Graphics, 28(11):3778–3787, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Xu:2024:MUR

[XLW^{+}24] Sen-Zhe Xu, Jia-Hong Liu, Miao Wang, Fang-Lue Zhang, and Song-Hai Zhang. Multiuser redirected walking in separate physical spaces for online VR scenarios. IEEE Transactions on Visualization and Computer Graphics, 30(4): 1916–1926, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Xu:2022:SFI

[XLY^{+}22] Wenpeng Xu, Yi Liu, Menglin Yu, Dongxiao Wang, Shouming Hou, Bo Li, Weiming Wang, and Ligang Liu. A support-free infill structure based on layer construction for 3D printing. IEEE Transactions on Visualization and Computer Graphics, 28(12):4462–4476, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2024:PII

[XLZ24] Ningchang Xiong, Qingqin Liu, and Kening Zhu. PetPresence: Investigating the integration of real-world pet activities in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2559–2569, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Xie:2022:FVA

[XMK^{+}22] Tiankai Xie, Yuxin Ma, Jian Kang, Hanghang Tong, and Ross Maciejewski. Fair-RankVis: a visual analytics framework for exploring algorithmic fairness in graph mining models. IEEE Transactions on Visualization and Computer Graphics, 28(1):368–377, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Xie:2021:ASG

[XMT^{+}21] T. Xie, Y. Ma, H. Tong, M. T. Thai, and R. Maciejewski. Auditing the sensitivity of graph-based ranking with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1459–1469, February 2021. CO-DEN ITVGEA. ISSN 1077-

Xu:2023:PVA

[XNF^{+}23] Chaoqing Xu, Tyson Neuroth, Takanori Fujiwara, Ronghua Liang, and Kwan-Liu Ma. A predictive visual analytics system for studying neurodegenerative disease based on DTI fiber tracts. IEEE Transactions on Visualization and Computer Graphics, 29(4):2020– 2035, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Xue:2024:RAL

[XPK^{+}24] Yumeng Xue, Patrick Paetzold, Rebecca Kehlbeck, Bin Chen, Kin Chung Kwan, Yunhai Wang, and Oliver Deussen. Reducing ambiguities in linebased density plots by imagespace colorization. IEEE Transactions on Visualization and Computer Graphics, 30(1):825– 835, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Xu:2023:MSF

[XQXL23] Cheng Xu, Wei Qu, Xuemiao Xu, and Xueting Liu. Multiscale flow-based occluding effect and content separation for cartoon animations. IEEE Transactions on Visualization and Computer Graphics, 29 (9):4001–4014, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Xenopoulos:2023:CIA

[XRN^{+}23] Peter Xenopoulos, João Rulff, Luis Gustavo Nonato, Brian Barr, and Claudio Silva. Calibrate: Interactive analysis of probabilistic model output. IEEE Transactions on Visualization and Computer Graphics, 29(1):853–863, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2022:VAB

[XSB^{+}22] Cindy Xiong, Vidya Setlur, Benjamin Bach, Eunyee Koh, Kylie Lin, and Steven Franconeri. Visual arrangements of bar charts influence comparisons in viewer takeaways. IEEE Transactions on Visualization and Computer Graphics, 28(1):955–965, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2020:ICV

[XSHF20] C. Xiong, J. Shapiro, J. Hullman, and S. Franconeri. Illusion of causality in visualized data. IEEE Transactions on Visualization and Computer Graphics, 26(1):853–862, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiong:2023:SWY

[XSKF23] Cindy Xiong, Chase Stokes, Yea-Seul Kim, and Steven Franconeri. Seeing what you believe or believing what you see? Belief biases correlation estimation. IEEE Transactions on Visualization and Computer Graphics, 29(1):493–503, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Xu:2020:EED

[XTYL20] J. Xu, Y. Tao, Y. Yan, and H. Lin. Exploring evolution of dynamic networks via diachronic node embeddings. IEEE Transactions on Visualization and Computer Graphics, 26(7):2387–2402, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Xiong:2020:CKV

[XVF20] C. Xiong, L. Van Weelden, and S. Franconeri. The curse of knowledge in visual data communication. IEEE Transactions on Visualization and Computer Graphics, 26(10):3051– 3062, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Xia:2021:QVA

[XVW^{+}21] M. Xia, R. P. Velumani, Y. Wang, H. Qu, and X. Ma. QLens: Visual analytics of multi-step problem-solving behaviors for improving question design. IEEE Transactions on Visualization and Computer Graphics, 27(2):870– 880, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Xue:2023:TNA

[XWH^{+}23] Mingliang Xue, Yunhai Wang, Chang Han, Jian Zhang, Zheng Wang, Kaiyi Zhang, Christophe Hurter, Jian Zhao, and Oliver Deussen. Target netgrams: an annulus-constrained stress model for radial graph visualization. IEEE Transactions on Visualization and Computer Graphics, 29(10):4256– 4268, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Xie:2021:PTB

[XWL^{+}21] X. Xie, J. Wang, H. Liang, D. Deng, S. Cheng, H. Zhang, W. Chen, and Y. Wu. PassVizor: Toward better understanding of the dynamics of soccer passes. IEEE Transactions on Visualization and Computer Graphics, 27(2):1322– 1331, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Xu:2024:TCB

[XWPG^{+}24] Xiang Xu, Lu Wang, Arsène Pérard-Gayot, Richard Membarth, Cuiyu Li, Chenglei Yang, and Philipp Slusallek. Temporal coherence-based distributed ray tracing of massive scenes. IEEE Transactions on Visualization and Computer Graphics, 30(2):1489–1501, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Xu:2020:CIV

[XWY^{+}20] K. Xu, Y. Wang, L. Yang, Y. Wang, B. Qiao, S. Qin, Y. Xu, H. Zhang, and H. Qu. CloudDet: Interactive visual analysis of anomalous performances in cloud computing systems. IEEE Transactions on Visualization and Computer Graphics, 26(1):1107– 1117, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xiao:2022:IDC

[XWZ^{+}22] Yi Xiao, Jin Wu, Jie Zhang, Peiyao Zhou, Yan Zheng, Chi-Sing Leung, and Ladislav Kavan. Interactive deep colorization and its application for image compression. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1557–1572, March 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Xue:2023:TTU

[XWZ^{+}23] Mingliang Xue, Zhi Wang, Fahai Zhong, Yong Wang, Mingliang Xu, Oliver Deussen, and Yunhai Wang. Taurus: Towards a unified force representation and universal solver for graph layout. IEEE Transactions on Visualization and Computer Graphics, 29(1):886– 895, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Xu:2021:MSH

[XWZB21] C. Xu, R. Wang, S. Zhao, and H. Bao. Multi-scale hybrid micro-appearance modeling and realtime rendering of thin fabrics. IEEE Transactions on Visualization and Computer Graphics, 27(4):2409– 2420, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Xu:2021:PAS

[XXM^{+}21] X. Xu, M. Xie, P. Miao, W. Qu, W. Xiao, H. Zhang, X. Liu, and T.-T. Wong. Perceptual-aware sketch simplification based on integrated VGG layers. IEEE Transactions on Visualization and Computer Graphics, 27(1): 178–189, January 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Xu:2021:GBL

[XYF^{+}21] P. Xu, G. Yan, H. Fu, T. Igarashi, C.-L. Tai, and H. Huang. Global beautification of 2D and 3D layouts with interactive ambiguity resolution. IEEE Transactions on Visualization and Computer Graphics, 27(4):2355–2368, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Xia:2022:PGG

[XZF^{+}22] Qianwei Xia, Juyong Zhang, Zheng Fang, Jin Li, Mingyue Zhang, Bailin Deng, and Ying

He. GeodesicEmbedding (GE):

a high-dimensional embedding approach for fast geodesic distance queries. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4930–4939, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Xuan:2022:VCV

[XZKM22] Xiwei Xuan, Xiaoyu Zhang, Oh-Hyun Kwon, and Kwan-Liu Ma. VAC-CNN: a visual analytics system for comparative studies of deep convolutional neural networks. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2326–2337, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Xia:2022:RDR

[XZS^{+}22] Jiazhi Xia, Yuchen Zhang, Jie Song, Yang Chen, Yunhai Wang, and Shixia Liu. Revisiting dimensionality reduction techniques for visual cluster analysis: an empirical study. IEEE Transactions on Visualization and Computer Graphics, 28(1):529–539, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Xu:2023:GVS

[XZS^{+}23] Xinchi Xu, Yang Zhou, Bingchan Shao, Guihuan Feng, and Chun Yu. GestureSurface: VR sketching through assembling scaffold surface with non-dominant hand. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2499–2507, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Xiao:2020:NCB

[XZWY20] X. Xiao, Y. Zhou, H. Wang, and X. Yang. A novel CNNbased Poisson solver for fluid simulation. IEEE Transactions on Visualization and Computer Graphics, 26(3):1454– 1465, March 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Xin:2022:LBC

[XZXY22] Hanggao Xin, Shaokun Zheng, Kun Xu, and Ling-Qi Yan. Lightweight bilateral convolutional neural networks for interactive single-bounce diffuse indirect illumination. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1824–1834, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Yu:2020:PHD

[YB20] R. Yu and D. A. Bowman.

Pseudo-haptic display of mass and mass distribution during object rotation in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 26(5):2094– 2103, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Yue:2020:SSV

[YBL^{+}20] X. Yue, J. Bai, Q. Liu, Y. Tang, A. Puri, K. Li, and H. Qu. sPortfolio: Stratified visual analysis of stock portfolios. IEEE Transactions on Visualization and Computer Graphics, 26(1):601–610, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2024:VEM

[YBOB24] Jun Yuan, Brian Barr, Kyle Overton, and Enrico Bertini. Visual exploration of machine learning model behavior with hierarchical surrogate rule sets. IEEE Transactions on Visualization and Computer Graphics, 30(2):1470–1488, February 2024. CODEN ITVGEA. ISSN 1077-2626.

Yan:2023:GAM

[YBR^{+}23] Lin Yan, Talha Bin Masood, Farhan Rasheed, Ingrid Hotz, and Bei Wang. Geometryaware merge tree comparisons for time-varying data with interleaving distances. IEEE Transactions on Visualization and Computer Graphics, 29(8): 3489–3506, August 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Yao:2022:VMR

[YBVI22] Lijie Yao, Anastasia Bezerianos, Romain Vuillemot, and Petra Isenberg. Visualization in motion: a research agenda and two evaluations. IEEE Transactions on Visualization and Computer Graphics, 28 (10):3546–3562, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Ye:2023:VEP

[YC23] Zelin Ye and Min Chen. Visualizing ensemble predictions of music mood. IEEE Transactions on Visualization and Computer Graphics, 29(1):864– 874, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Yang:2021:ENI

[YCB^{+}21] Y. Yang, M. Cordeil, J. Beyer, T. Dwyer, K. Marriott, and H. Pfister. Embodied navigation in immersive abstract data visualization: Is overview + detail or zooming better for 3D scatterplots? IEEE Transactions on Visualization and Computer Graphics, 27(2): 1214–1224, February 2021. CO-DEN ITVGEA. ISSN 1077-

Ye:2021:SEA

[YCC^{+}21] S. Ye, Z. Chen, X. Chu, Y. Wang, S. Fu, L. Shen, K. Zhou, and Y. Wu. ShuttleSpace: Exploring and analyzing movement trajectory in immersive visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2): 860–869, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Ye:2023:PVR

[YCC^{+}23] Shuainan Ye, Zhutian Chen, Xiangtong Chu, Kang Li, Juntong Luo, Yi Li, Guohua Geng, and Yingcai Wu. PuzzleFixer: a visual reassembly system for immersive fragments restoration. IEEE Transactions on Visualization and Computer Graphics, 29(1):429–439, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Ye:2022:MPA

[YCH^{+}22] Shuquan Ye, Dongdong Chen, Songfang Han, Ziyu Wan, and Jing Liao. Meta-PU: an arbitrary-scale upsampling network for point cloud. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3206–3218, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Ye:2024:QA

[YCHL24] Shuquan Ye, Dongdong Chen, Songfang Han, and Jing Liao. 3D question answering. IEEE Transactions on Visualization and Computer Graphics, 30(3): 1772–1786, March 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Yang:2024:SPI

[YCM^{+}24] Fumeng Yang, Mandi Cai, Chloe Mortenson, Hoda Fakhari, Ayse D. Lokmanoglu, Jessica Hullman, Steven Franconeri, Nicholas Diakopoulos, Erik C. Nisbet, and Matthew Kay. Swaying the public? Impacts of election forecast visualizations on emotion, trust, and intention in the 2022 U.S. Midterms. IEEE Transactions on Visualization and Computer Graphics, 30(1):23–33, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Yeh:2024:AGV

[YCW^{+}24] Catherine Yeh, Yida Chen, Aoyu Wu, Cynthia Chen, Fernanda Viégas, and Martin Wattenberg. AttentionViz: a global view of transformer attention. IEEE Transactions on Visualization and Computer Graphics, 30(1):262–272, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Yang:2022:HBR

[YCZ^{+}22] Zhenjie Yang, Beijia Chen, Youyi Zheng, Xiang Chen, and Kun Zhou. Human bas-relief generation from a single photograph. IEEE Transactions on Visualization and Computer Graphics, 28(12):4558– 4569, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Yang:2021:TMI

[YDM^{+}21] Yalong Yang, Tim Dwyer, Kim Marriott, Bernhard Jenny, and Sarah Goodwin. Tilt map: Interactive transitions between choropleth map, prism map and bar chart in immersive environments. IEEE Transactions on Visualization and Computer Graphics, 27(12):4507– 4519, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Yeshchenko:2022:VDD

[YDMP22] Anton Yeshchenko, Claudio Di Ciccio, Jan Mendling, and Artem Polyvyanyy. Visual drift detection for event sequence data of business processes. IEEE Transactions on Visualization and Computer Graphics, 28(8):3050–3068, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Yu:2022:DRC

[YEP^{+}22] Kevin Yu, Ulrich Eck, Frieder Pankratz, Marc Lazarovici, Dirk Wilhelm, and Nassir Navab. Duplicated reality for co-located augmented reality collaboration. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2190–2200, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Yoon:2023:EAT

[YeSiK^{+}23] Boram Yoon, Jae eun Shin, Hyung il Kim, Seo Young Oh, Dooyoung Kim, and Woontack Woo. Effects of avatar transparency on social presence in task-centric mixed reality remote collaboration. IEEE Transactions on Visualization and Computer Graphics, 29(11):4578–4588, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2023:SSS

[YFM^{+}23] Ganzhangqin Yuan, Qiancheng Fu, Zhenxing Mi, Yiming Luo, and Wenbing Tao. SSRNet: Scalable 3D surface reconstruction network. IEEE Transactions on Visualization and Computer Graphics, 29(12): 4906–4919, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Yazdanpour:2022:MOD

[YFS22] Mahdi Yazdanpour, Guoliang Fan, and Weihua Sheng. ManhattanFusion: Online dense reconstruction of indoor scenes from depth sequences. IEEE Transactions on Visualization and Computer Graphics, 28(7): 2668–2681, July 2022. CODEN ITVGEA. ISSN 1077-2626.

Yu:2021:ATE

[YGE^{+}21] Kevin Yu, Gleb Gorbachev, Ulrich Eck, Frieder Pankratz, Nassir Navab, and Daniel Roth. Avatars for teleconsultation: Effects of avatar embodiment techniques on user perception in 3D asymmetric telepresence. IEEE Transactions on Visualization and Computer Graphics, 27(11):4129–4139, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Yu:2023:SAP

[YGH^{+}23] Piaopiao Yu, Jie Guo, Fan Huang, Zhenyu Chen, Chen Wang, Yan Zhang, and Yanwen Guo. ShadowMover: Automatically projecting real shadows onto virtual object. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2379–2389, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Yan:2024:TTR

[YGP^{+}24] Lin Yan, Hanqi Guo, Thomas Peterka, Bei Wang, and Jiali Wang. TROPHY: a topologically robust physics-informed tracking framework for tropical cyclones. IEEE Transactions on Visualization and Computer Graphics, 30(1):1249– 1259, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Yang:2023:MMD

[YGT^{+}23] Jie Yang, Lin Gao, Qingyang Tan, Yi-Hua Huang, Shihong Xia, and Yu-Kun Lai. Multiscale mesh deformation component analysis with attentionbased autoencoders. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1301–1317, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Yang:2024:IRM

[YGW^{+}24] Weikai Yang, Yukai Guo, Jing Wu, Zheng Wang, Lan-Zhe Guo, Yu-Feng Li, and Shixia Liu. Interactive reweighting for mitigating label quality issues.

IEEE Transactions on Visualization and Computer Graphics, 30(3):1837–1852, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Yin:2022:OMC

[YHC^{+}22] Tairan Yin, Ludovic Hoyet, Marc Christie, Marie-Paule Cani, and Julien Pettré. The one-man-crowd: Single user generation of crowd motions using virtual reality. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2245–2255, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Yin:2024:YEC

[YHC^{+}24] Tairan Yin, Ludovic Hoyet, Marc Christie, Marie-Paule Cani, and Julien Pettré. With or without you: Effect of contextual and responsive crowds on VR-based crowd motion capture. IEEE Transactions on Visualization and Computer Graphics, 30(5):2785– 2795, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Yu:2021:TRS

[YHD21] Guoxing Yu, Yongtao Hu, and Jingwen Dai. TopoTag: a robust and scalable topological fiducial marker system. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3769–3780, September 2021. CODEN ITVGEA. ISSN 1077-

Yasui:2024:PMB

[YIIW24] Masahiko Yasui, Ryota Iwataki, Masatoshi Ishikawa, and Yoshihiro Watanabe. Projection mapping with a brightly lit surrounding using a mixed light field approach. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2217–2227, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Yang:2021:TAS

[YISG21] Cong Yang, Bipin Indurkhya, John See, and Marcin Grzegorzek. Towards automatic skeleton extraction with skeleton grafting. IEEE Transactions on Visualization and Computer Graphics, 27(12): 4520–4532, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Yamamoto:2023:MPC

[YITS23] Kenta Yamamoto, Daisuke Iwai, Ikuho Tani, and Kosuke Sato. A monocular projectorcamera system using modular architecture. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5586–5592, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Yousef:2021:STA

[YJ21] T. Yousef and S. Janicke. A survey of text alignment visualization. IEEE Transactions on Visualization and Computer Graphics, 27(2):1149– 1159, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Ye:2022:CCA

[YKF22] Hui Ye, Kin Chung Kwan, and Hongbo Fu. 3D curve creation on and around physical objects with mobile AR. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2809–2821, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Yu:2023:PIP

[YKJ^{+}23] Yuncong Yu, Dylan Kruyff, Jiao Jiao, Tim Becker, and Michael Behrisch. PSEUDo: Interactive pattern search in multivariate time series with localitysensitive hashing and relevance feedback. IEEE Transactions on Visualization and Computer Graphics, 29(1):33–42, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Yu:2020:DEV

[YLF^{+}20] D. Yu, H. Liang, K. Fan, H. Zhang, C. Fleming, and K. Papangelis. Design and evaluation of visualization techniques of off-screen and occluded targets in virtual reality environments. IEEE Transactions on Visualization and Computer Graphics, 26 (9):2762–2774, September 2020. CODEN ITVGEA. ISSN 1077-2626.

Yan:2024:TPT

[YLGW24] Lin Yan, Xin Liang, Hanqi Guo, and Bei Wang. TopoSZ:

Preserving topology in errorbounded lossy compression. IEEE Transactions on Visualization and Computer Graphics, 30(1):1302–1312, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Yoo:2022:VHM

[YLJ^{+}22] Dongheon Yoo, Seungjae Lee, Youngjin Jo, Jaebum Cho, Suyeon Choi, and Byoungho Lee. Volumetric head-mounted display with locally adaptive focal blocks. IEEE Transactions on Visualization and Computer Graphics, 28(2):1415– 1427, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Yang:2020:HDS

[YLL^{+}20] S. Yang, B. Li, M. Liu, Y.-K.

Lai, L. Kobbelt, and S.-M. Hu. HeteroFusion: Dense scene reconstruction integrating multisensors. IEEE Transactions on Visualization and Computer Graphics, 26(11):3217– 3230, November 2020. CODEN ITVGEA. ISSN 1077-2626.

Ye:2022:PCB

[YLL^{+}22a] Zhichao Ye, Guanglin Li, Haomin Liu, Zhaopeng Cui, Hujun Bao, and Guofeng Zhang. CoLi-BA: Compact linearization based solver for bundle adjustment. IEEE Transactions on Visualization and Computer Graphics, 28(11): 3727–3736, November 2022. CODEN ITVGEA. ISSN 1077-

Yeh:2022:WDH

[YLL^{+}22b] Chih-Kuo Yeh, Zhanping Liu, I-Hsuan Lin, Eugene Zhang, and Tong-Yee Lee. WYSIWYG design of hypnotic line art. IEEE Transactions on Visualization and Computer Graphics, 28(6): 2517–2529, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Ye:2024:ILS

[YLLW24] Huayuan Ye, Chenhui Li, Yang Li, and Changbo Wang. InvVis: Large-scale data embedding for invertible visualization. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1139–1149, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Yan:2023:CED

[YLS^{+}23] Yukang Yan, Haohua Liu, Yingtian Shi, Jingying Wang, Ruici Guo, Zisu Li, Xuhai Xu, Chun Yu, Yuntao Wang, and Yuanchun Shi. ConeSpeech: Exploring directional speech interaction for multi-person remote communication in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5):2647– 2657, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2023:VAN

[YLTL23] Jun Yuan, Mengchen Liu, Fengyuan Tian, and Shixia Liu. Visual analysis of neural architecture spaces for summarizing design principles. IEEE Transactions on Visualization and Computer Graphics, 29(1):288– 298, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Young:2023:CSC

[YPW23] Jacob Young, Nadia Pantidi, and Matthew Wood. I can’t see that! considering the readability of small objects in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2567–2574, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Yang:2021:VRM

[YQN^{+}21] Fumeng Yang, Jing Qian, Johannes Novotny, David Badre, Cullen D. Jackson, and David H. Laidlaw. A virtual reality memory palace variant aids knowledge retrieval from scholarly articles. IEEE Transactions on Visualization and Computer Graphics, 27(12): 4359–4373, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Yan:2020:SPT

[YRL^{+}20] X. Yan, C. Rao, L. Lu, A. Sharf, H. Zhao, and B. Chen. Strong 3D printing by TPMS injection. IEEE Transactions on Visualization and Computer Graphics, 26(10):3037–3050, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Yu:2020:FNL

[YS20] B. Yu and C. T. Silva.

FlowSense: a natural language interface for visual data exploration within a dataflow system.

IEEE Transactions on Visualization and Computer Graphics, 26(1):1–11, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Ying:2023:MAG

[YSD^{+}23] Lu Ying, Xinhuan Shu, Dazhen Deng, Yuchen Yang, Tan Tang, Lingyun Yu, and Yingcai Wu. MetaGlyph: Automatic generation of metaphoric glyph-based visualization. IEEE Transactions on Visualization and Computer Graphics, 29(1):331– 341, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Ye:2020:UCD

[YSM^{+}20] Y. C. Ye, F. Sauer, K. Ma, K. Aditya, and J. Chen. A user-centered design study in scientific visualization targeting domain experts. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2192–2203, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Yan:2020:VAC

[YST^{+}20] J. Yan, L. Shi, J. Tao, X. Yu, Z. Zhuang, C. Huang, R. Yu, P. Su, C. Wang, and Y. Chen. Visual analysis of collective anomalies using faceted highorder correlation graphs. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2517–2534, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Yang:2023:VCE

[YTHL23] Fumeng Yang, James Tompkin, Lane Harrison, and David H. Laidlaw. Visual cue effects on a classification accuracy estimation task in immersive scatterplots. IEEE Transactions on Visualization and Computer Graphics, 29(12):4858– 4873, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Ying:2022:GTE

[YTL^{+}22] Lu Ying, Tan Tangl, Yuzhe Luo, Lvkeshen Shen, Xiao Xie, Lingyun Yu, and Yingcai Wu. GlyphCreator: Towards example-based automatic generation of circular glyphs. IEEE Transactions on Visualization and Computer Graphics, 28(1): 400–410, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Yuksel:2021:SLS

[Yuk21] Cem Yuksel. Stochastic lightcuts for sampling many lights. IEEE Transactions on Visualization and Computer Graphics, 27(10):4049–4059, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Yao:2024:DVM

[YVBI24] Lijie Yao, Romain Vuillemot, Anastasia Bezerianos, and Petra Isenberg. Designing for visualization in motion: Embedding visualizations in swimming videos. IEEE Transactions on Visualization and Computer Graphics, 30(3):1821– 1836, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2020:TPB

[YWB20] Y. Yuan, R. Wang, and H. Bao. Tile pair-based adaptive multirate stereo shading. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2303–2314, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Yang:2020:ONV

[YWL^{+}20] Y. Yang, M. Wybrow, Y. Li, T. Czauderna, and Y. He. OntoPlot: a novel visualisation for non-hierarchical associations in large ontologies. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1140–1150, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yang:2021:ISH

[YWL^{+}21] Weikai Yang, Xiting Wang, Jie Lu, Wenwen Dou, and Shixia Liu. Interactive steering of hierarchical clustering. IEEE Transactions on Visualization and Computer Graphics, 27 (10):3953–3967, October 2021. CODEN ITVGEA. ISSN 1077-2626.

Yan:2020:SAL

[YWM^{+}20] L. Yan, Y. Wang, E. Munch, E. Gasparovic, and B. Wang. A structural average of labeled merge trees for uncertainty visualization. IEEE Transactions on Visualization and Computer Graphics, 26(1):832–842, January 2020. CODEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Yuan:2023:VFC

[YWM^{+}23] Na Yuan, Peihui Wang, Wenlong Meng, Shuangmin Chen, Jian Xu, Shiqing Xin, Ying He, and Wenping Wang. A variational framework for curve shortening in various geometric domains. IEEE Transactions on Visualization and Computer Graphics, 29(4):1951– 1963, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Yang:2022:DSA

[YXL^{+}22] Leni Yang, Xian Xu, XingYu Lan, Ziyan Liu, Shunan Guo, Yang Shi, Huamin Qu, and Nan Cao. A design space for applying the Freytag’s pyramid structure to data stories. IEEE Transactions on Visualization and Computer Graphics, 28(1): 922–932, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Ye:2023:CEE

[YXS^{+}23] Zipeng Ye, Mengfei Xia, Yanan Sun, Ran Yi, Minjing Yu, Juyong Zhang, Yu-Kun Lai, and Yong-Jin Liu. 3D-CariGAN: an end-to-end solution to 3D caricature generation from normal face photos. IEEE Transactions on Visualization and Computer Graphics, 29(4):2203– 2210, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Yang:2023:EEL

[YXW^{+}23] Leni Yang, Cindy Xiong, Jason K. Wong, Aoyu Wu, and Huamin Qu. Explaining with examples: Lessons learned from crowdsourced introductory description of information visualizations. IEEE Transactions on Visualization and Computer Graphics, 29(3):1638– 1650, March 2023. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2021:ESM

[YXX^{+}21] J. Yuan, S. Xiang, J. Xia, L. Yu, and S. Liu. Evaluation of sampling methods for scatterplots. IEEE Transactions on Visualization and Computer Graphics, 27(2):1720–1730, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Yoghourdjian:2021:SNV

[YYD^{+}21] V. Yoghourdjian, Y. Yang, T. Dwyer, L. Lawrence, M. Wybrow, and K. Marriott. Scalability of network visualisation from a cognitive load perspective. IEEE Transactions on Visualization and Computer Graphics, 27(2):1677–1687, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Yoon:2022:PRV

[YYK^{+}22] Leonard Yoon, Dongseok Yang, Jaehyun Kim, ChoongHo Chung, and Sung-Hee Lee. Placement retargeting of virtual avatars to dissimilar indoor environments. IEEE Transactions on Visualization and Computer Graphics, 28(3):1619–1633, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Yang:2022:DEF

[YYZ^{+}22] Weikai Yang, Xi Ye, Xingxing Zhang, Lanxi Xiao, Jiazhi Xia, Zhongyuan Wang, Jun Zhu, Hanspeter Pfister, and Shixia Liu. Diagnosing ensemble few-shot classifiers. IEEE Transactions on Visualization and Computer Graphics, 28 (9):3292–3306, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Yu:2022:PBM

[YZEN22] Kevin Yu, Kostantinos Zacharis, Ulrich Eck, and Nassir Navab. Projective bisector mirror (PBM): Concept and rationale. IEEE Transactions on Visualization and Computer Graphics, 28(11):3694–3704, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2022:DCE

[YZF^{+}22] Lin-Ping Yuan, Wei Zeng, Siwei Fu, Zhiliang Zeng, Haotian Li, Chi-Wing Fu, and Huamin Qu. Deep colormap extraction from visualizations. IEEE Transactions on Visualization and Computer Graphics, 28(12):4048–4060, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Yang:2023:EIV

[YZF^{+}23] Chuang Yang, Zhiwen Zhang, Zipei Fan, Renhe Jiang, Quanjun Chen, Xuan Song, and Ryosuke Shibasaki. EpiMob: Interactive visual analytics of citywide human mobility restrictions for epidemic control. IEEE Transactions on Visualization and Computer Graphics, 29(8):3586–3601, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Yang:2020:MRT

[YZJ^{+}20] X. Yang, L. Zhou, H. Jiang, Z. Tang, Y. Wang, H. Bao, and G. Zhang. Mobile3DRecon: Real-time monocular 3D reconstruction on a mobile phone. IEEE Transactions on Visualization and Computer Graphics, 26(12):3446–3456, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Yu:2020:FOT

[YZN^{+}20] D. Yu, Q. Zhou, J. Newn, T. Dingler, E. Velloso, and J. Goncalves. Fully-occluded target selection in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 26(12):3402– 3413, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Yang:2023:RNE

[YZP^{+}23] Guo-Wei Yang, Wen-Yang Zhou, Hao-Yang Peng, Dun Liang, Tai-Jiang Mu, and Shi-Min Hu. Recursive-NeRF: an efficient and dynamically growing NeRF. IEEE Transactions on Visualization and Computer Graphics, 29(12):5124– 5136, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Yuan:2022:PII

[YZZ^{+}22] Lin-Ping Yuan, Ziqi Zhou, Jian Zhao, Yiqiu Guo, Fan Du, and

Huamin Qu. InfoColorizer:

Interactive recommendation of color palettes for infographics. IEEE Transactions on Visualization and Computer Graphics, 28(12):4252–4266, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2021:RBV

[ZA21] Jingbo Zhao and Robert S. Allison. The role of binocular vision in avoiding virtual obstacles while walking. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3277–3288, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:VUP

[ZAH22] Dongping Zhang, Eytan Adar, and Jessica Hullman. Visualizing uncertainty in probabilistic graphs with network hypothetical outcome plots (NetHOPs). IEEE Transactions on Visualization and Computer Graphics, 28(1):443–453, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2024:EGD

[ZBG^{+}24] Henan Zhao, Garnett W.

Bryant, Wesley Griffin, Judith E. Terrill, and Jian Chen. Evaluating glyph design for showing large-magnitude-range quantum spins. IEEE Transactions on Visualization and Computer Graphics, 30(4):

1868–1884, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Zong:2021:LDI

[ZBNS21] J. Zong, D. Barnwal, R. Neogy, and A. Satyanarayan. Lyra 2: Designing interactive visualizations by demonstration. IEEE Transactions on Visualization and Computer Graphics, 27(2): 304–314, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zytko:2023:DMW

[ZC23] Douglas Zytko and Jonathan Chan. The dating metaverse: Why we need to design for consent in social VR. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2489–2498, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2021:DEG

[ZCH^{+}21] M. Zhu, W. Chen, Y. Hu, Y. Hou, L. Liu, and K. Zhang. DRGraph: an efficient graph layout algorithm for large-scale graphs by dimensionality reduction. IEEE Transactions on Visualization and Computer Graphics, 27(2):1666– 1676, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2021:UOE

[ZCL^{+}21] M. Zhang, L. Chen, Q. Li, X. Yuan, and J. Yong. Uncertainty-oriented ensemble data visualization and exploration using variable spatial spreading. IEEE Transactions on Visualization and Computer Graphics, 27(2):1808– 1818, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2023:EVP

[ZCL23] Howe Yuan Zhu, Hsiang-Ting Chen, and Chin-Teng Lin. The effects of virtual and physical elevation on physiological stress during virtual reality height exposure. IEEE Transactions on Visualization and Computer Graphics, 29(4): 1937–1950, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhuang:2022:FED

[ZCM22] Mengdie Zhuang, David Concannon, and Ed Manley. A framework for evaluating dashboards in healthcare. IEEE Transactions on Visualization and Computer Graphics, 28(4): 1715–1731, April 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:GEE

[ZCM23a] Xinyu Zhang, Shenghui Cheng, and Klaus Mueller. Graphical enhancements for effective exemplar identification in contextual data visualizations. IEEE Transactions on Visualization and Computer Graphics, 29 (9):3775–3787, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:VBA

[ZCM^{+}23b] Kanglei Zhou, Ruizhi Cai, Yue Ma, Qingqing Tan, Xinning Wang, Jianguo Li, Hubert P. H. Shum, Frederick W. B. Li, Song Jin, and Xiaohui Liang. A video-based augmented reality system for human-in-the-loop muscle strength assessment of juvenile dermatomyositis. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2456–2466, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2021:ALE

[ZCR21] Junhong Zhao, Andrew Chalmers, and Taehyun Rhee. Adaptive light estimation using dynamic filtering for diverse lighting conditions. IEEE Transactions on Visualization and Computer Graphics, 27(11):4097– 4106, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Zheng:2023:TST

[ZCSS23] Youyi Zheng, Beijia Chen, Yuefan Shen, and Kaidi Shen. TeethGNN: Semantic 3D teeth segmentation with graph neural networks. IEEE Transactions on Visualization and Computer Graphics, 29(7):3158– 3168, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Zang:2024:DEE

[ZCX^{+}24] Zelin Zang, Shenghui Cheng, Hanchen Xia, Liangyu Li, Yaoting Sun, Yongjie Xu, Lei Shang, Baigui Sun, and Stan Z. Li. DMT-EV: an explainable deep network for dimension reduction. IEEE Transactions on Visualization and Computer Graphics, 30(3):1710– 1727, March 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2021:FPT

[ZCZ^{+}21] H. Zhou, K. Chen, W. Zhang, C. Qin, and N. Yu. Featurepreserving tensor voting model for mesh steganalysis. IEEE Transactions on Visualization and Computer Graphics, 27(1): 57–67, January 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:FER

[ZCZ22] Juyong Zhang, Keyu Chen, and Jianmin Zheng. Facial expression retargeting from human to avatar made easy. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1274–1287, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2023:AOA

[ZCZ^{+}23a] Song-Hai Zhang, Chia-Hao Chen, Fu Zheng, Yong-Liang Yang, and Shi-Min Hu. Adaptive optimization algorithm for resetting techniques in obstacleridden environments. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2080–2092, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:OSP

[ZCZ23b] Song-Hai Zhang, Chiahao Chen, and Stefanie Zollmann. One-step out-of-place resetting for redirected walking in VR. IEEE Transactions on Visualization and Computer Graphics, 29(7):3327–3339, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Zakopanova:2021:VVS

[ZeB^{+}21] K. Zákop anová, M. ehá ek, J. Bátrna, D. Plakinger, S. Stoppel, and B. Kozlíková. Visilant: Visual support for the exploration and analytical process tracking in criminal investigations. IEEE Transactions on Visualization and Computer Graphics, 27(2):881– 890, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2023:ACT

[ZFCG23] Yanhong Zeng, Jianlong Fu, Hongyang Chao, and Baining Guo. Aggregated contextual transformations for highresolution image inpainting. IEEE Transactions on Visualization and Computer Graphics, 29(7):3266–3280, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2022:CIS

[ZFF22] Jian Zhao, Mingming Fan, and Mi Feng. ChartSeer: Interactive steering exploratory visual analysis with machine intelligence. IEEE Transactions on Visualization and Computer Graphics, 28(3):1500– 1513, March 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2020:FGV

[ZG20] B. Zhou and S. Güven. Finegrained visual recognition in mobile augmented reality for technical support. IEEE Transactions on Visualization and Computer Graphics, 26(12):

3514–3523, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2021:RSA

[ZGL^{+}21] Hao Zhu, Mantang Guo, Hongdong Li, Qing Wang, and Antonio Robles-Kelly. Revisiting spatio-angular trade-off in light field cameras and extended applications in super-resolution. IEEE Transactions on Visualization and Computer Graphics, 27(6):3019–3033, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:BSS

[ZGX^{+}22] Long Zhang, Jianwei Guo, Jun Xiao, Xiaopeng Zhang, and Dong-Ming Yan. Blending surface segmentation and editing for 3D models. IEEE Transactions on Visualization and Computer Graphics, 28(8): 2879–2894, August 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Zhao:2023:AVA

[ZGX^{+}23] Ying Zhao, Luhao Ge, Huixuan Xie, Genghuai Bai, Zhao Zhang, Qiang Wei, Yun Lin, Yuchao Liu, and Fangfang Zhou. ASTF: Visual abstractions of time-varying patterns in radio signals. IEEE Transactions on Visualization and Computer Graphics, 29(1):214– 224, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zuenko:2020:WFC

[ZH20] E. Zuenko and M. Harders. Wrinkles, folds, creases, buckles: Small-scale surface deformations as periodic functions on 3D meshes. IEEE Transactions on Visualization and Computer Graphics, 26 (10):3077–3088, October 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2020:WSA

[ZHDX20] Z. Zhang, L. Hu, X. Deng, and S. Xia. Weakly supervised adversarial learning for 3D human pose estimation from point clouds. IEEE Transactions on Visualization and Computer Graphics, 26(5): 1851–1859, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2020:SAT

[ZHF^{+}20] L. Zhu, X. Hu, C. Fu, J. Qin, and P. Heng. Saliency-aware texture smoothing. IEEE Transactions on Visualization and Computer Graphics, 26(7): 2471–2484, July 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:SSE

[ZHH22] Wenxiao Zhang, Bo Han, and Pan Hui. SEAR: Scaling experiences in multi-user augmented reality. IEEE Transactions on Visualization and Computer Graphics, 28(5):1982– 1992, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:AOT

[ZHKY23] Yan Zhang, Xiaodan Hu, Kiyoshi Kiyokawa, and Xubo Yang. Add-on occlusion: Turning off-the-shelf optical seethrough head-mounted displays occlusion-capable. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2700–2709, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2020:LCO

[ZHL^{+}20] X. Zhao, R. Hu, H. Liu, T. Komura, and X. Yang. Localization and completion for 3D object interactions. IEEE Transactions on Visualization and Computer Graphics, 26(8): 2634–2644, August 2020. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2021:AAS

[ZHL^{+}21a] S. Zhang, Z. Han, Y.-K. Lai, M. Zwicker, and H. Zhang. Active arrangement of small objects in 3D indoor scenes. IEEE Transactions on Visualization and Computer Graphics, 27(4): 2250–2264, April 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2021:SSE

[ZHL^{+}21b] K. Zhu, X. He, S. Li, H. Wang, and G. Wang. Shallow sand equations: Real-time height field simulation of dry granular flows. IEEE Transactions on Visualization and Computer Graphics, 27(3):2073– 2084, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Zabel:2024:VVP

[ZHN24] Susanne Zabel, Philipp Hennig, and Kay Nieselt. VIPur-PCA: Visualizing and propagating uncertainty in principal component analysis. IEEE Transactions on Visualization and Computer Graphics, 30(4): 2011–2022, April 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhai:2020:FSA

[ZHQH20] X. Zhai, F. Hou, H. Qin, and A. Hao. Fluid simulation with adaptive staggered power particles on GPUs. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2234–2246, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:IEP

[ZHTR22] Xingdi Zhang, Markus Hadwiger, Thomas Theußl, and Peter Rautek. Interactive exploration of physically-observable objective vortices in unsteady 2D flow. IEEE Transactions on Visualization and Computer Graphics, 28(1):281–290, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2024:MTC

[ZIX^{+}24] Lixiang Zhao, Tobias Isenberg, Fuqi Xie, Hai-Ning Liang, and Lingyun Yu. MeTACAST: Target- and context-aware spatial selection in VR. IEEE Transactions on Visualization and Computer Graphics, 30(1): 480–494, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Zhao:2021:PMS

[ZJC^{+}21] Y. Zhao, H. Jiang, Q. Chen, Y. Qin, H. Xie, Y. Wu, S. Liu, Z. Zhou, J. Xia, and F. Zhou. Preserving minority structures in graph sampling. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1698–1708, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zhou:2021:SAD

[ZJJH21] Wen Zhou, Jinyuan Jia, Wenying Jiang, and Chenxi Huang. Sketch augmentation-driven shape retrieval learning framework based on convolutional neural networks. IEEE Transactions on Visualization and Computer Graphics, 27(8): 3558–3570, August 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zhu:2022:PTH

[ZJS^{+}22] Lifeng Zhu, Xudong Jiang, Jiangwei Shen, Heng Zhang, Yiting Mo, and Aiguo Song. TapeTouch: a handheld shapechanging device for haptic display of soft objects. IEEE Transactions on Visualization and Computer Graphics, 28(11):3928–3938, November 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2021:DDS

[ZJW21] L. Zhou, C. R. Johnson, and D. Weiskopf. Data-driven space-filling curves. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1591–1600, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2024:OCD

[ZJX^{+}24] Yu Zhang, Ruike Jiang, Liwenhan Xie, Yuheng Zhao, Can Liu, Tianhong Ding, Siming Chen, and Xiaoru Yuan. OldVisOnline: Curating a dataset of historical visualizations. IEEE Transactions on Visualization and Computer Graphics, 30(1): 551–561, January 2024. CO-DEN ITVGEA. ISSN 1077-2626.

Zhao:2020:MVA

[ZKS^{+}20] J. Zhao, M. Karimzadeh, L. S. Snyder, C. Surakitbanharn, Z. C. Qian, and D. S. Ebert. MetricsVis: a visual analytics system for evaluating employee performance in public safety agencies. IEEE Transactions on Visualization and Computer Graphics, 26(1): 1193–1203, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2022:PEE

[ZLC^{+}22] Danyong Zhao, Yijing Li, Siddhartha Chaudhuri, Timothy Langlois, and Jernej Barbi . ERGOBOSS: Ergonomic optimization of body-supporting surfaces. IEEE Transactions on Visualization and Computer Graphics, 28(12):4032– 4047, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:EUF

[ZLC^{+}23] Mingdong Zhang, Quan Li, Li Chen, Xiaoru Yuan, and Junhai Yong. EnConVis: a unified framework for ensemble contour visualization. IEEE Transactions on Visualization and Computer Graphics, 29(4): 2067–2079, April 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2021:PAA

[ZLG^{+}21a] J. Zhao, X. Liu, C. Guo, Z. C. Qian, and Y. V. Chen. Phoenixmap: an abstract approach to visualize 2D spatial distributions. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2000–2014, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zollmann:2021:VTA

[ZLG^{+}21b] Stefanie Zollmann, Tobias Langlotz, Raphael Grasset, Wei Hong Lo, Shohei Mori, and Holger Regenbrecht. Visualization techniques in augmented reality: a taxonomy, methods and patterns. IEEE Transactions on Visualization and Computer Graphics, 27 (9):3808–3825, September 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2020:VAE

[ZLL^{+}20] Y. Zhao, X. Luo, X. Lin, H. Wang, X. Kui, F. Zhou, J. Wang, Y. Chen, and W. Chen. Visual analytics for electromagnetic situation awareness in radio monitoring and management. IEEE Transactions on Visualization and Computer Graphics, 26(1): 590–600, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zeng:2021:RMA

[ZLL^{+}21] W. Zeng, C. Lin, J. Lin, J. Jiang, J. Xia, C. Turkay, and W. Chen. Revisiting the modifiable areal unit problem in deep traffic prediction with visual analytics. IEEE Transactions on Visualization and Computer Graphics, 27(2):839– 848, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2021:PPC

[ZLQH21] D. Zhang, X. Lu, H. Qin, and Y. He. Pointfilter: Point cloud filtering via encoder-decoder modeling. IEEE Transactions on Visualization and Computer Graphics, 27(3):2015– 2027, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2021:TPT

[ZLS21] Lu Zhao, Yue Liu, and Weitao Song. Tactile perceptual thresholds of electrovibration in VR. IEEE Transactions on Visualization and Computer Graphics, 27(5):2618–2626, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Zytek:2022:SUA

[ZLVV22] Alexandra Zytek, Dongyu Liu, Rhema Vaithianathan, and Kalyan Veeramachaneni. Sibyl: Understanding and addressing the usability challenges of machine learning in highstakes decision making. IEEE Transactions on Visualization and Computer Graphics, 28(1): 1161–1171, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2021:VEI

[ZLW21a] P. Zhang, C. Li, and C. Wang.

VisCode: Embedding information in visualization images using encoder–decoder network. IEEE Transactions on Visualization and Computer Graphics, 27(2):326–336, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2021:GBF

[ZLW^{+}21b] W. Zhao, X. Liu, S. Wang, X. Fan, and D. Zhao. Graphbased feature-preserving mesh normal filtering. IEEE Transactions on Visualization and Computer Graphics, 27(3): 1937–1952, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zheng:2023:SOA

[ZLX23] Chengwei Zheng, Wenbin Lin, and Feng Xu. A self-occlusion aware lighting model for realtime dynamic reconstruction. IEEE Transactions on Visualization and Computer Graphics, 29(10):4062–4073, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:DEB

[ZLY22] Mohan Zhang, Jing Liao, and Jinhui Yu. Deep exemplarbased color transfer for 3D model. IEEE Transactions on Visualization and Computer Graphics, 28(8):2926–2937, August 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2021:CPI

[ZLZ21] Yun Zhang, Yu-Kun Lai, and Fang-Lue Zhang. Contentpreserving image stitching with piecewise rectangular boundary constraints. IEEE Transactions on Visualization and Computer Graphics, 27(7):3198– 3212, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:CFV

[ZLZ^{+}23a] He Zhang, Fan Li, Jianhui Zhao, Chao Tan, Dongming Shen, Yebin Liu, and Tao Yu. Controllable free viewpoint video reconstruction based on neural radiance fields and motion graphs. IEEE Transactions on Visualization and Computer Graphics, 29(12):4891– 4905, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:NMP

[ZLZ^{+}23b] Yu-Wei Zhang, Ping Luo, Hao Zhou, Zhongping Ji, Hui Liu, Yanzhao Chen, and Caiming Zhang. Neural modeling of portrait bas-relief from a single photograph. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5008–5019, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2022:EFF

[ZMD^{+}22] Zehua Zeng, Phoebe Moh, Fan Du, Jane Hoffswell, Tak Yeon Lee, Sana Malik, Eunyee Koh, and Leilani Battle. An evaluation-focused framework for visualization recommendation algorithms. IEEE Transactions on Visualization and Computer Graphics, 28(1):346– 356, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Zenner:2020:IPM

[ZMK^{+}20] A. Zenner, A. Makhsadov, S. Klingner, D. Liebemann, and A. Krüger. Immersive process model exploration in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 26(5):2104– 2114, May 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2020:HPV

[ZNAN20] X. Zhu, M. A. Nacenta, Ö. Akgün, and P. Nightingale. How people visually represent discrete constraint problems. IEEE Transactions on Visualization and Computer Graphics, 26(8):2603–2619, August 2020. CODEN ITVGEA. ISSN 1077-

Zhang:2023:TFP

[ZNF^{+}23] Chenxu Zhang, Saifeng Ni, Zhipeng Fan, Hongbo Li, Ming Zeng, Madhukar Budagavi, and Xiaohu Guo. 3D talking face with personalized pose dynamics. IEEE Transactions on Visualization and Computer Graphics, 29(2):1438– 1449, February 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2023:MVS

[ZOF^{+}23] Guanghan Zhao, Jason Orlosky, Steven Feiner, Photchara Ratsamee, and Yuki Uranishi. Mitigation of VR sickness during locomotion with a motionbased dynamic vision modulator. IEEE Transactions on Visualization and Computer Graphics, 29(10):4089– 4103, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:SDS

[ZOS^{+}23] Xiaoyu Zhang, Jorge Piazentin Ono, Huan Song, Liang Gou, Kwan-Liu Ma, and Liu Ren. SliceTeller: a data slice-driven approach for machine learning model validation. IEEE Transactions on Visualization and Computer Graphics, 29(1):842– 852, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:DRV

[ZP23] Yuqi Zhou and Voicu Popescu. Dynamic redirection for VR haptics with a handheld stick. IEEE Transactions on Visualization and Computer Graphics, 29(5):2753–2762, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2024:CFS

[ZP24] Yuqi Zhou and Voicu Popescu.

CloVR: Fast-startup low-latency cloud VR. IEEE Transactions on Visualization and Computer Graphics, 30(5):2337– 2346, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhu:2022:LSA

[ZPF^{+}22] Peihao Zhu, Wamiq Reyaz Para, Anna Frühstück, John Femiani, and Peter Wonka. Large-scale architectural asset extraction from panoramic imagery. IEEE Transactions on Visualization and Computer Graphics, 28(2):1301– 1316, February 2022. CODEN ITVGEA. ISSN 1077-2626.

Zheng:2021:FTU

[ZPG21] J. X. Zheng, S. Pawar, and D. F. M. Goodman. Further towards unambiguous edge bundling: Investigating powerconfluent drawings for network visualization. IEEE Transactions on Visualization and Computer Graphics, 27(3): 2244–2249, March 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zong:2023:AVL

[ZPWS23] Jonathan Zong, Josh Pollock, Dylan Wootton, and Arvind Satyanarayan. Animated Vega-Lite: Unifying animation with a grammar of interactive graphics. IEEE Transactions on Visualization and Computer Graphics, 29(1):149–159, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2020:PHD

[ZRJW20] L. Zhou, M. Rivinius, C. R. Johnson, and D. Weiskopf. Photographic high-dynamic-range scalar visualization. IEEE Transactions on Visualization and Computer Graphics, 26(6): 2156–2167, June 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:TSH

[ZRPW23] Youjia Zhou, Archit Rathore, Emilie Purvine, and Bei Wang. Topological simplifications of hypergraphs. IEEE Transactions on Visualization and Computer Graphics, 29(7): 3209–3225, July 2023. CODEN ITVGEA. ISSN 1077-2626.

Zheng:2021:UCS

[ZS21] B. Zheng and F. Sadlo. Uncertainty in continuous scatterplots, continuous parallel coordinates, and fibers. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1819–1828, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zhao:2022:UML

[ZSCC22] Jian Zhao, Maoyuan Sun, Francine Chen, and Patrick Chiu. Understanding missing links in bipartite networks with MissBiN. IEEE Transactions on Visualization and Computer Graphics, 28(6):2457– 2469, June 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2023:EAR

[ZSCRB23] Yu Zhao, Jeanine Stefanucci, Sarah Creem-Regehr, and Bobby Bodenheimer. Evaluating augmented reality landmark cues and frame of reference displays with virtual reality. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2710–2720, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:VDP

[ZSG^{+}23] Yixuan Zhang, Yifan Sun, Joseph D. Gaggiano, Neha Kumar, Clio Andris, and Andrea G. Parker. Visualization design practices in a crisis: Behind the scenes with COVID-19 dashboard creators. IEEE Transactions on Visualization and Computer Graphics, 29(1): 1037–1047, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Zellmann:2021:BTC

[ZSL21] S. Zellmann, J. P. Schulze, and U. Lang. Binned k–d tree construction for sparse volume data on multi-core and GPU systems. IEEE Transactions on Visualization and Computer Graphics, 27(3):1904– 1915, March 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2022:EEB

[ZSL^{+}22] Ying Zhao, Jingcheng Shi, Jiawei Liu, Jian Zhao, Fangfang Zhou, Wenzhi Zhang, Kangyi Chen, Xin Zhao, Chunyao Zhu, and Wei Chen. Evaluating effects of background stories on graph perception. IEEE Transactions on Visualization and Computer Graphics, 28(12):4839–4854, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2020:DFR

[ZSLL20] B. Zhang, B. Sheng, P. Li, and T. Lee. Depth of field rendering using multilayer-neighborhood optimization. IEEE Transactions on Visualization and Computer Graphics, 26(8): 2546–2559, August 2020. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2021:IVA

[ZSM21] D. Zhang, A. Sarvghad, and G. Miklau. Investigating visual analysis of differentially private data. IEEE Transactions on Visualization and Computer Graphics, 27(2):1786– 1796, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2021:CAS

[ZSS^{+}21] Z. Zhou, C. Shi, X. Shen, L. Cai, H. Wang, Y. Liu, Y. Zhao, and W. Chen. Context-aware sampling of large networks via graph representation learning. IEEE Transactions on Visualization and Computer Graphics, 27(2): 1709–1719, February 2021. CO-DEN ITVGEA. ISSN 1077-2626.

Zeng:2021:EEO

[ZSW^{+}21] Haipeng Zeng, Xinhuan Shu, Yanbang Wang, Yong Wang, Liguo Zhang, Ting-Chuen Pong, and Huamin Qu. EmotionCues: Emotion-oriented visual summarization of classroom videos. IEEE Transactions on Visualization and Computer Graphics, 27(7): 3168–3181, July 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2024:SET

[ZSZ^{+}24] Tongyu Zhang, Yiran Shen, Guangrong Zhao, Lin Wang, Xiaoming Chen, Lu Bai, and Yuanfeng Zhou. Swift-Eye: Towards anti-blink pupil tracking for precise and robust highfrequency near-eye movement analysis with event cameras. IEEE Transactions on Visualization and Computer Graphics, 30(5):2077–2086, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:VRU

[ZTC^{+}23] Wei Zhang, Siwei Tan, Siming Chen, Linghao Meng, Tianye Zhang, Rongchen Zhu, and Wei Chen. Visual reasoning for uncertainty in spatio-temporal events of historical figures. IEEE Transactions on Visualization and Computer Graphics, 29(6):3009–3023, June 2023.

CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:SAR

[ZTL^{+}23] Shao-Kui Zhang, Hou Tam, Yi-Xiao Li, Tai-Jiang Mu, and Song-Hai Zhang. SceneViewer: Automating residential photography in virtual environments. IEEE Transactions on Visualization and Computer Graphics, 29(12):5523–5537, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2021:SFS

[ZTWW21] Jun Zhang, Jun Tao, Jian-Xun Wang, and Chaoli Wang. SurfRiver: Flattening stream surfaces for comparative visualization. IEEE Transactions on Visualization and Computer Graphics, 27(6):2783– 2795, June 2021. CODEN ITVGEA. ISSN 1077-2626.

Zenner:2021:CDP

[ZUK21] André Zenner, Kristin Ullmann, and Antonio Krüger. Combining dynamic passive haptics and haptic retargeting for enhanced haptic feedback in virtual reality. IEEE Transactions on Visualization and Computer Graphics, 27(5): 2627–2637, May 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:DSS

[ZWG^{+}23] Zhilan Zhou, Wenyuan Wang, Mengtian Guo, Yue Wang, and David Gotz. A design space for surfacing content recommendations in visual analytic platforms. IEEE Transactions on Visualization and Computer Graphics, 29(1):84–94, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:PLB

[ZWH^{+}22] Yunjin Zhang, Rui Wang, Yuchi Huo, Wei Hua, and Hujun Bao. PowerNet: Learning-based realtime power-budget rendering. IEEE Transactions on Visualization and Computer Graphics, 28(10):3486–3498, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:AJO

[ZWL23] Jingbo Zhang, Ziyu Wan, and Jing Liao. Adaptive joint optimization for 3D reconstruction with differentiable rendering. IEEE Transactions on Visualization and Computer Graphics, 29(6):3039–3051, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2021:PSC

[ZWP21] Zijing Zhou, Lili Wang, and Voicu Popescu. A partiallysorted concentric layout for efficient label localization in augmented reality. IEEE Transactions on Visualization and Computer Graphics, 27(11): 4087–4096, November 2021. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:CCE

[ZWP^{+}22] Yunjin Zhang, Rui Wang, Yifan Peng, Wei Hua, and Hujun Bao. Color contrast enhanced rendering for optical see-through head-mounted displays. IEEE Transactions on Visualization and Computer Graphics, 28(12):4490– 4502, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zahalka:2021:IIP

[ZWV21] J. Zahálka, M. Worring, and J. J. Van Wijk. II-20: Intelligent and pragmatic analytic categorization of image collections. IEEE Transactions on Visualization and Computer Graphics, 27(2):422– 431, February 2021. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2020:EVA

[ZWW^{+}20] H. Zeng, X. Wang, A. Wu, Y. Wang, Q. Li, A. Endert, and H. Qu. EmoCo: Visual analysis of emotion coherence in presentation videos. IEEE Transactions on Visualization and Computer Graphics, 26(1): 927–937, January 2020. CO-DEN ITVGEA. ISSN 1077-2626 (print), 1941-0506 (electronic), 2160-9306.

Zhao:2022:GCP

[ZWW22] Hui Zhao, Shaodong Wang, and Wencheng Wang. Global conformal parameterization via an implementation of holomorphic quadratic differentials. IEEE Transactions on Visualization and Computer Graphics, 28(3): 1529–1544, March 2022. CO-DEN ITVGEA. ISSN 1077-

Zeng:2023:GVA

[ZWW^{+}23a] Haipeng Zeng, Xingbo Wang, Yong Wang, Aoyu Wu, Ting-Chuen Pong, and Huamin Qu. GestureLens: Visual analysis of gestures in presentation videos. IEEE Transactions on Visualization and Computer Graphics, 29(8):3685–3697, August 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:CVA

[ZWW^{+}23b] Wei Zhang, Jason K. Wong, Xumeng Wang, Youcheng Gong, Rongchen Zhu, Kai Liu, Zihan Yan, Siwei Tan, Huamin Qu, Siming Chen, and Wei Chen. CohortVA: a visual analytic system for interactive exploration of cohorts based on historical data. IEEE Transactions on Visualization and Computer Graphics, 29(1):756– 766, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:FVA

[ZWW^{+}23c] Jiehui Zhou, Xumeng Wang, Jie Wang, Hui Ye, Huanliang Wang, Zihan Zhou, Dongming Han, Haochao Ying, Jian Wu, and Wei Chen. FraudAuditor: a visual analytics approach for collusive fraud in health insurance. IEEE Transactions on Visualization and Computer Graphics, 29(6): 2849–2861, June 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2023:DIP

[ZWW^{+}23d] Jiehui Zhou, Xumeng Wang, Jason K. Wong, Huanliang Wang, Zhongwei Wang, Xiaoyu Yang, Xiaoran Yan, Haozhe Feng, Huamin Qu, Haochao Ying, and Wei Chen. DPVis-Creator: Incorporating pattern constraints to privacypreserving visualizations via differential privacy. IEEE Transactions on Visualization and Computer Graphics, 29(1): 809–819, January 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Zeng:2023:NTD

[ZWXM23] Yan Zeng, Lu Wang, Yanning Xu, and Xiangxu Meng. Neural temporal denoising for indirect illumination. IEEE Transactions on Visualization and Computer Graphics, 29(12): 5569–5578, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2022:KBL

[ZWZ^{+}22] Yue Zhao, Yunhai Wang, Jian Zhang, Chi-Wing Fu, Mingliang Xu, and Dominik Moritz. KD-Box: Line-segment-based KDtree for interactive exploration of large-scale time-series data. IEEE Transactions on Visualization and Computer Graphics, 28(1):890–900, January 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:PIP

[ZWZ^{+}23] Chenyang Zhang, Xiyuan Wang, Chuyi Zhao, Yijing Ren, Tianyu Zhang, Zhenhui Peng, Xiaomeng Fan, Xiaojuan Ma, and Quan Li. PromotionLens: Inspecting promotion strategies of online e-commerce via visual analytics. IEEE Transactions on Visualization and Computer Graphics, 29(1):767–777, January 2023. CODEN ITVGEA. ISSN 1077-2626.

Zheng:2022:DDT

[ZX22] Chengwei Zheng and Feng Xu.

DTexFusion: Dynamic texture fusion using a consumer RGBD sensor. IEEE Transactions on Visualization and Computer Graphics, 28(10):3365– 3375, October 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2023:CAP

[ZXC^{+}23] Jian Zhao, Shenyu Xu, Senthil Chandrasegaran, Chris Bryan, Fan Du, Aditi Mishra, Xin Qian, Yiran Li, and Kwan-Liu Ma. ChartStory: Automated partitioning, layout, and captioning of charts into comic-style narratives. IEEE Transactions on Visualization and Computer Graphics, 29(2): 1384–1399, February 2023. CO-DEN ITVGEA. ISSN 1077-2626.

Zhao:2022:HLE

[ZXSR22] Zhenge Zhao, Panpan Xu, Carlos Scheidegger, and Liu Ren. Human-in-the-loop extraction of interpretable concepts in deep learning models. IEEE Transactions on Visualization and Computer Graphics, 28(1): 780–790, January 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2023:CIC

[ZYC^{+}23] Congyi Zhang, Lei Yang, Nenglun Chen, Nicholas Vining, Alla Sheffer, Francis C. M. Lau, Guoping Wang, and Wenping Wang. CreatureShop: Interactive 3D character modeling and texturing from a single color drawing. IEEE Transactions on Visualization and Computer Graphics, 29(12):4874– 4890, December 2023. CODEN ITVGEA. ISSN 1077-2626.

Zafar:2024:ECH

[ZYC24a] Adeel Zafar, Di Yang, and Guoning Chen. Extract and characterize hairpin vortices in turbulent flows. IEEE Transactions on Visualization and Computer Graphics, 30(1):716– 726, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2024:CAG

[ZYC^{+}24b] Yuxing Zhou, Weikai Yang, Jiashu Chen, Changjian Chen, Zhiyang Shen, Xiaonan Luo, Lingyun Yu, and Shixia Liu. Cluster-aware grid layout. IEEE Transactions on Visualization and Computer Graphics, 30(1):240–250, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2022:VIV

[ZYL^{+}22] Yizhong Zhang, Jiaolong Yang, Zhen Liu, Ruicheng Wang, Guojun Chen, Xin Tong, and Baining Guo. VirtualCube: an immersive 3D video communication system. IEEE Transactions on Visualization and Computer Graphics, 28(5): 2146–2156, May 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2024:PLV

[ZYL^{+}24] Guofeng Zhang, Jin Yuan, Haomin Liu, Zhen Peng, Chunlei Li, Zibin Wang, and Hujun Bao. 100-Phones: a large VI-SLAM dataset for augmented reality towards mass deployment on mobile phones. IEEE Transactions on Visualization and Computer Graphics, 30(5): 2098–2108, May 2024. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2024:TMC

[ZYM^{+}24] Zehua Zeng, Junran Yang, Dominik Moritz, Jeffrey Heer, and Leilani Battle. Too many cooks: Exploring how graphical perception studies influence visualization recommendations in Draco. IEEE Transactions on Visualization and Computer Graphics, 30(1):1063– 1073, January 2024. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2024:MWS

[ZYP^{+}24] Wenhua Zhang, Yating Yue, Hao Pan, Zhonggui Chen, Chuan Wang, Hanspeter Pfister, and Wenping Wang. Marching windows: Scalable mesh generation for volumetric data with multiple materials. IEEE Transactions on Visualization and Computer Graphics, 30(3):1728–1742, March 2024.

CODEN ITVGEA. ISSN 1077-2626.

Zhou:2020:EFT

[ZYR^{+}20] Q. Zhou, D. Yu, M. N. Reinoso, J. Newn, J. Goncalves, and E. Velloso. Eyes-free target acquisition during walking in immersive mixed reality. IEEE Transactions on Visualization and Computer Graphics, 26(12):3423–3433, December 2020. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2023:PGD

[ZZ23] Hui Zeng and Rong Zhao. Perceptually-guided dual-mode virtual reality system for motion-adaptive display. IEEE Transactions on Visualization and Computer Graphics, 29(5): 2249–2257, May 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2022:TDM

[ZZC^{+}22] Hang Zhou, Weiming Zhang, Kejiang Chen, Weixiang Li, and Nenghai Yu. Threedimensional mesh steganography and steganalysis: a review. IEEE Transactions on Visualization and Computer Graphics, 28(12):5006–5025, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:PCC

[ZZD^{+}23a] Wenxiao Zhang, Huajian Zhou, Zhen Dong, Jun Liu, Qingan Yan, and Chunxia Xiao. Point cloud completion via skeletondetail transformer. IEEE Transactions on Visualization and Computer Graphics, 29 (10):4229–4242, October 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:RPR

[ZZD^{+}23b] Wenxiao Zhang, Huajian Zhou, Zhen Dong, Qingan Yan, and Chunxia Xiao. Rank-PointRetrieval: Reranking point cloud retrieval via a visually consistent registration evaluation. IEEE Transactions on Visualization and Computer Graphics, 29(9):3840– 3854, September 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2023:SLC

[ZZS^{+}23] Hongkun Zhang, Kehong Zhou, Ke Shi, Yunhai Wang, Aiguo Song, and Lifeng Zhu. Smart-Spring: a low-cost wearable haptic VR display with controllable passive feedback. IEEE Transactions on Visualization and Computer Graphics, 29(11):4460–4471, November 2023. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2020:PRM

[ZZW^{+}20] Y. Zhang, C. Zhang, W. Wang, Y. Chen, Z. Ji, and H. Liu. Portrait relief modeling from a single image. IEEE Transactions on Visualization and Computer Graphics, 26(8):2659–2670, August 2020. CODEN ITVGEA. ISSN 1077-2626.

Zhang:2021:IAR

[ZZW21] Aijia Zhang, Yan Zhao, and Shigang Wang. An improved augmented-reality framework for differential rendering beyond the Lambertian-world assumption. IEEE Transactions on Visualization and Computer Graphics, 27(12):4374– 4386, December 2021. CODEN ITVGEA. ISSN 1077-2626.

Zeng:2022:DDC

[ZZW^{+}22a] Qiong Zeng, Yongwei Zhao, Yinqiao Wang, Jian Zhang, Yi Cao, Changhe Tu, Ivan Viola, and Yunhai Wang. Datadriven colormap adjustment for exploring spatial variations in scalar fields. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4902–4917, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhao:2022:RBP

[ZZW^{+}22b] Xi Zhao, Bowen Zhang, Jinji Wu, Ruizhen Hu, and Taku Komura. Relationship-based point cloud completion. IEEE Transactions on Visualization and Computer Graphics, 28(12): 4940–4950, December 2022. CODEN ITVGEA. ISSN 1077-2626.

Zhou:2022:DRE

[ZZW^{+}22c] Zixia Zhou, Xinrui Zu, Yuanyuan Wang, Boudewijn P. F. Lelieveldt, and Qian Tao. Deep recursive embedding for highdimensional data. IEEE Transactions on Visualization and Computer Graphics, 28(2): 1237–1248, February 2022. CO-DEN ITVGEA. ISSN 1077-2626.

Zhang:2022:FIS

[ZZX^{+}22] Song-Hai Zhang, Shao-Kui Zhang, Wei-Yu Xie, Cheng-Yang Luo, Yong-Liang Yang, and Hongbo Fu. Fast 3D indoor scene synthesis by learning spatial relation priors of objects. IEEE Transactions on Visualization and Computer Graphics, 28(9):3082–3092, September 2022. CODEN ITVGEA. ISSN 1077-2626.

Zheng:2024:IDO

[ZZZ24] Jinta Zheng, Eugene Zhang, and Yue Zhang. Interactive design and optics-based visualization of arbitrary non-Euclidean kaleidoscopic orbifolds. IEEE Transactions on Visualization and Computer Graphics, 30(1): 1292–1301, January 2024. CO-DEN ITVGEA. ISSN 1077-